A Complete Bibliography of Publications in Numerical Algorithms

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 May 2026 Version 2.159

Title word cross-reference

(0, 2) [MP08]. (1, 1) − q [MPC12]. (1 − x)^{α}(1 + x)^{β} + Mδ(x + 1) + Nδ(x − 1) [KW96]. (2, 2) [ST22]. (3, 3) [KCHD16]. (A, B) [KP96a]. (k, l) [Gut15]. (p, q) [JXNX24, JXX^{+}23]. (R_{1}, R_{2}, R_{3}) [IDAV09]. (w, p, u) [Oua99]. −2 < α [DJ18]. 0 < q < 1 [Mah10]. 1 [ATM19, Bea98, BCJ99, ICR06, JV98, Moh10, MK17, Pas92]. 1.5 [ST17]. 1/9 [MM00]. 11 [HvD93]. 13 [DL21]. 2 [AM98b, AEF^{+}14, CJ17, CG19, DSI11, DL21, EL01, JV98, LZlL20, Sal17, TC05, VMMD21, ZYX19, ZYLN18]. 28 [Mar04b]. 3 [CL00, FHC21, GH09c, LG08, MFK^{+}15, PMO05, ZJ08, ZS19]. 3(2) [RSKB17]. 3 × 3 [HN16, HM19b]. 4 [CVX16, NB16, PT18, Wri01, ZXRL11]. 4 × 4 [KH18]. 5 [LMUZ19]. 7 [MV02]. 8 [MV02]. [−1, 1] [The12, TV17]. [0, 1] [VC92]. ^{n} [Ple03]. ^{T} [Orb15]. \{_{1}Φ_{1}(\binom{1,1}{c,γ_{j}}andz)\}_{j=1}^{n} [dDL92]. _{3}F_{2}(−n, b, c,.e; z) [DJ02]. _{4}F_{3} [DL01]. _{σ} [LCZZ23]. A [But02a, FYYW19, MT13, MT14, MT19, ZD15, But96, Con93]. A(p)X = B(p) [DMD16, Pop18b]. A^{T}A [Tou98]. A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{ℓ}X_{ℓ}B_{ℓ} = C [hPwL09].

A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{l}X_{l}B_{l} = C [Pen13]. A_{i}XB_{i} = F_{i} [TPY14]. A_{n} [Sid20a]. A_{T,S}^{(2)} [Sol15, MGL20]. α [ASV23, Gar19, HÇK26, HY21, HS21, HSY23, IAH20, MA15, MP14, NUG25, PDRG19, PGP^{+}25, WZ26, YHL25]. a ⊗ 0 [KR07]. AX = B [ZMZ26]. AXB + C\overline{X}D + EX^{T} F + GX^{∗}H = J [SD26]. AXB = C [TDWL24, lLhYfD07]. AXB = CAXB = C [jHyPlZ06]. AXC = B [ZMZ26]. B [But19, Cha14, GEP16, HZX20, LL16, LDL17, Mil17, dS24, FJT94, Gre96, Lai92, MRS93, Rab92b, Rab92c, Str93]. β [AL15, CLQ26, MP14]. β < −1 [DJ18]. BiCGstab(l) [SvF94]. BN_{1} [SW26]. C^{0} [LY18, LWY26]. C^{1} [LRZ12, XZ25, Gal93]. C^{2} [ALY22, LM01, MS96]. C^{∞} [Zhu21]. C^{k} [MS01a, MST03, NS01]. C^{r} [ARTY20]. cos(z) [VC00]. d [MDH16, dR99]. D^{m} [GLRSG08]. D_{ω} [MM08a]. Δ^{2} [Sab92a]. Δ_{1} [MS01a]. Δ_{h} [CL13b]. E [Bag16, GS19a, MP92]. E_{ν}(x) [NP18]. ℓ_{0} [CS24]. ℓ_{1} [AR26]. ℓ_{p} [LGZ25]. ϵ [BRZ17, BRZ19, BRZS23, CGM93, GCFF95, Le 92, Mat92, MN11, Sal96]. η [CZL26]. F [AUA22, HLL22]. f(A)b [Dea15]. G [But15, Hil10, SCS18, YAT20]. G^{1} [Poc14, CLaL00]. G^{k,l} [GLW16]. γ [Kar00]. GC^{2} [HvD93]. H [Cve06, CK06, HS26, LZ14, WPL18, Bar13, BX19, LH23]. H(curl) [WPW25]. H^{−1} [Liu21]. H^{1} [FHC21, LR18, LZlL20, LCZZ23, ZJ14, ZZY24, GM06]. H^{2} [Ber93]. H^{∞} [DHS97, DI11]. H_{+} [RWTW19]. hp [CVX16]. i [Fue07]. I_{x}(a, b) [YA24]. ∞ [PW14]. k [BCL00, HVMT17, MKA26, OLB94, RALJ26, Wat93, Wat94]. L [BM14, CMRS00b, CRS04, HASI23]. L^{1} [ART14]. L^{2} [JZ25, LCH20, TZ24, MG11b, BM94]. L^{2}(I) ⊕ C^{2} [AT19]. L^{∞} [LKQ23]. L^{∞}[0, T] [DN24b]. l^{p} [BR21, LI10, Cro03]. l^{q} [BR21]. l_{0} [ZPX21, MH22]. L_{1} [GGNF17, GGN18, WYZ22, MG18, AGN07, Gaj05, PW04, ASW06, AS08, BL93, BL95, GH95]. L_{1}C^{1} [GGN14]. l_{2} [MG18, Meu05, KLZV95]. l_{2,p} [LZ22b]. L_{2}^{(m)}(0, 1) [BHS17]. L_{p} [BGS24a, Wat94]. LDM^{t} [TS92]. LL^{T} [BSB23, BRS26]. LU [WLW^{+}24, ZWLZ24]. M [GL04, Guo13, Laz99, Men26, PWCsL18, Rab92b, Rab92c, THS20]. R^{2} [Ave20]. R^{4} [Sal25]. R^{n} [AHL20, HYT26, MTTC22]. B [CHY19]. H [LLQ17, XZZ19a, XZZ19b, LW20]. H^{∈} [Bör26]. H_{+} [WCW20]. KS [WMCW21]. M [BHLZ21]. MB [CLWV15]. μ [CCZ23]. N [AGA24, BJT24, CDGC25, Don13, HKKN12, Lor19, Sha02, Sha19, WSZ21, LB93, Li95, Sid20a, WZ15a, dDL92]. N_{1} [SW26]. ν [KK00]. O(1/k^{2}) [ZXL25]. O(h^{6}) [RGJ10]. O(n) [Khe16, MR09, Pol10]. O(n^{2}) [Góm99, Góm01]. Ω [NN26, HÇK26, JM00]. Orthomin(k) [Tou98]. P [IUM^{+}19, KS20, LLY22, PFT98, ADN17, BHM05, BX19, CL11, CSFC04, Ern00, GL20, ML10, Sch25, Col92]. P_{∗}(κ) [Khe14, SD20]. P_{∗}(κ)(κ) [AM13]. P_{1} [OMW21]. P_{2m+1} [XZ25]. P_{n}^{(α,β)} [DJ18]. p ≥ 2 [LGZ25]. ϕ [Jaw22]. π [Kal00]. ψ [JM00]. q [AH17, CW08, GA08, GS19b, HLSLSL25, KG23, KJG23, LGA^{+}00, LWG18, LQ16, Mah10, MP14, MAK20, Ost07, SC18, dDL92, KL91]. q > 1 [Ost07]. qd [DMS09, DS12, EAB20]. QR [DM22, Hua18, SLZ25, ZZ22a, ZZW26, BH92b]. QZ [MVVV24]. R [Chk20, WK13, Gaj05, HW00]. r > 1 [ARTY20]. r[u] [ZAH26]. R^{3} [Caç10]. R^{d} [Gaj05]. R^{n} [Wat06]. r_{1}, r_{2}, r_{3} [ED22]. R_{II} [BPR20]. J [Hom98b]. S [SKK21, SCS18, MAL04, Tou98, ZWG18]. S^{k} [ZWL26]. SB [DLL13]. SDD_{k} [LRW25]. sin(z) [VC00]. t [ALL25, HHLS21a, HHLS21b, HZ24]. τ [GH25, lLHNS23, Reb97]. θ [LDL^{+}19, ZLM25, ZM26, Zag92]. θ_{2} [Sab92a]. d̃^{(m)} [Sid20a]. u [AL15, DLR24]. U(a, b, x) [GRAST23]. ε [Ber14, HLW26, Sal25, SGM02]. φ [AG25, JP14, LCL21, MWsC19]. |γ| = 1, ∑ d_{n} < ∞ [Thr92]. W [BBM08, Gau11a, Joh20]. w^{2} [AUA22]. W_{∞}^{r,d} [Kow00, Kow00]. X = Q + A^{H}(I ⊗ X − C)^{δ}A [YLD11]. x^{α} [VC92]. y′′ + My = f(y) [LDW18]. y′′ = f(x, y, y′) [JSF13]. y′′ = f(y(t)) [DEP12]. y′′(t) = f(t, y(t)) [LW16]. Y AY − Y B − CY + D = 0 [EH24b]. Z [Cou15a, Cou15b, GLS26, GLL19, San19]. Z × aZ^{n} [KR07].

-acceptability [Col92]. -accretive [THS20]. -adaptive [LH23]. -Algorithm [SGM02, Ber14, BRZ19, BRZS23, CGM93, EAB20, GS19a, GCFF95, HLW26, Le 92, MP92, Mat92, Sal96, Sal25, Zag92]. -algorithms [BRZ17]. -analogues [GA08]. -anti [CZL26]. -Appell [CL13b]. -Approximation [PW04, Gaj05]. -balanced [BCL00]. -band [Laz99]. -Baskakov [MP14]. -Bernstein [Mah10, MAK20, Ost07]. -Bernstein-like [Bar13]. -Blossoming [AH17, GS19b]. -body [AGA24, Sha02, Lor19, Sha19]. -bound-preserving [TZ24]. -classical [LGA^{+}00]. -coherent [MPC12]. -constrained [GLW16]. -convergence [MG11b, ZJ14]. -coupled [Don13]. -Curve [CRS04]. -cyclic [Ern00]. -D [MFK^{+}15, Sal17]. -dense [Gar19]. -dependent [ZWG18]. -dimensional [DSI11, Li95, MDH16]. -eigenpairs [GLL19]. -eigenvalue [San19]. -eigenvalues [CHY19]. -error [LCZZ23]. -exponential [AH17]. -feedbacks [HLL22]. -fractal [ASV23, PGP^{+}25]. -fractions [LB93]. -functions [AG25, Jaw22, Gau11a]. -Galerkin [GM06, LR18]. -gamma [GA08]. -generation [DLR24]. -gradient [Liu21]. -harmonic [Rab92b, Rab92c]. -Hermite [HLSLSL25]. -Hermitian [CZL26]. -horizontal [Khe14]. -hypergeometric [dDL92]. -instant [LMUZ19]. -inverses [BM94]. -iteration [SCS18]. -iterative [SKK21]. -Jacobi [KG23, KJG23]. -Laguerre [KJG23]. -Laplace [ML10, ZAH26]. -Laplacian [JP14]. -like [TS92]. -major [Wat94]. -matrices [Bör26, CLWV15, CK06, GLS26, Hil10, LW20, LZ14, LL16, LLY22, Men26, RWTW19, SW26, WPL18]. -Matrix [GL04, Cve06, DLL13, Guo13]. -means [RALJ26]. -Median [CSFC04]. -method [GH25, Reb97, HZX20]. -mode [CCZ23, BJT24]. -Nekrasov [GEP16, LDL17]. -nets [Lai92]. -nonexpansive [IAH20, MWsC19, NUG25, PDRG19, SCS18, YAT20]. -norm [LGZ25, BGS24a, FHC21, LZlL20, LCH20, MT13, MT14, MT19, ZZY24]. -order [WK13]. -orthogonal [Tou98]. -orthogonality [dR99]. -orthonormalization [ZD15]. -patches [Str93]. -pentadiagonal [JXNX24]. -periodic [WSZ21]. -Phillips [LQ16]. -point [DL21, WZ15a]. -preconditioner [lLHNS23]. -product [ALL25, HZ24]. -pseudocontractive [LCL21]. -rational [FJT94]. -regularization [MH22].

-regularized [ZPX21]. -representation [HS26]. -ribbon [CMRS00b]. -robust [CLQ26, HS21, YHL25]. -scheme [WYZ22]. -SCLCPs [SD20]. -SDD [ZWL26]. -semi-classical [MM08a]. -series [CW08, But19, Cha14, dS24]. -spaces [MAL04]. -sparse [PWCsL18]. -spline [Mil17]. -splines [Gre96, KK00, KL91, MRS93, Rab92b, Rab92c, BM14, GLRSG08, LRZ12, MS01a]. -stability [But96, But02a]. -stable [Con93, HY21, HSY23, WZ26, FYYW19, KS20, NN26, PFT98]. -stage [NB16]. -step [Gut15, HVMT17, Tou98]. -Stokes [Sch25]. -storage [Pol10]. -symplectic [But15]. -tensor [WMCW21]. -tensors [BHLZ21, LLQ17, WCW20, XZZ19a, XZZ19b]. -th [CVX16, Fue07, GL20, MKA26]. -theory [MA15]. -transform [BBM08, Cou15a, Cou15b]. -transformation [AL15]. -tridiagonal [JXX^{+}23]. -tuple [dDL92]. -Uniform [MN11]. -uniformly [SC18, IUM^{+}19]. -variate [PW14].

/Octave [MBR21].

1 [BRS92, DJM^{+}18, HM22a, LCZZ23]. 1-parameter [Uhl22a, Uhl22b]. 14 [Tov98]. 16-stage [Zha24]. 1994 [BV96]. 1D [BGS24b, CKO24, LZ23b]. 1st [DMRT03].

2 [HM22b]. 2-Toeplitz [SSC25]. 2022 [BC22]. 2023 [BJK^{+}24]. 2D [CS25, FMS25, HCBAEC23, KSJ26, Mil20, MP22, PM24, RNT26, SMM25, SzS21, WJC25, XSL22, ZL22b, ZZ22b, ZQzS22, SFS23]. 2D/ [SFS23]. 2D/3D [XSL22]. 2nd [Cor24]. 2nd-order [Cor24].

3.0 [Han99]. 3D [BGS24b, FLH25, FMS25, LAH22, QQX23, SFS23, XSL22]. 3DVAR [JS23].

4.0 [Han07].

5 [CP95b]. 5.2 [Han99]. 500 [AWL^{+}24].

7.3 [Han07].

8 [CZ95].

90th [BJK^{+}24]. 92j [BRS92]. 95i [CZ95]. ’97 [Ano98d]. 98c [Tov98].

=Helmholtz [SZ23].

A-posteriori [DBGB11]. A. [AD00]. A.D.I. [MS02, SMB02]. AAA [ZZX26]. AAA-type [ZZX26]. Abaqus [DN19]. Abel [CZZ^{+}25, CW08]. Aberth [Bin96]. ABS [KMA13]. abscissae [Not08]. absence [PPS25]. absolute [CHYH24, DW24, LLD23, LM25]. absorbing [ALZ21, WBY25, ZPEL25]. abstract [HPS20]. Accelerate [GR01]. Accelerated [And10, BM24b, DS20, Erb15, LY17, ML22, SM10, WMCW21, ZY13b, ZM16, And14b, BKR18, FC01, GH10, HT16, HS20, HHF22, HC24, KZ21, LG18, LLL26, PRK^{+}18, SS10, SBJC19, SLL22, TDWL24, WL24, YYL26, ZSWW25]. Accelerating [CL99, KBCG13, MCG^{+}04, Hom92, HL23a, LZOY22, RR00, Sab91, Vep08]. Acceleration [And19a, Bre00a, EGSV04, EN11, KP03, Mat91, MP92, Pas03, Pep23, Sid20a, TY21, UTO24, Wil12, AL15, And06, AS14, BDD20, CL00, CHHL18, CR20, Cro92, Esp05, GM96, HS16, Hom98a, Kza97, LDZ24, LCHH21, Mat92, MPS20, NG23, Now06, Now13, Now19, Osa12, PL99, PXH25, Pas92, Pas08, Pas11, SD26, SS98, WRM17, CDP16]. acceptability [Col92]. accessibility [EHVRV14]. accretive [THS20]. Accuracy [AEG02, BD04b, LZX23, Mat15, CC16b, CV92, FJ96, FS01, GST21, KKO17, LMV23, MPR24, Moh10, MDL15, Mok16, SMM25, SW05, TL24, YWYN22, Zag24b]. Accurate [AP21, DH04, FLH04, GLS26, LL05, MMV17, MPT21, MNS23, SS11a, YHZ20, YHZL21, AM12, BIMR19, CZ23, CFL19, CW21b, GH06, GKV23, GLW13, HJB18, HM26, KMRAR25, KGG25, LZM23, LZ23a, LZZ23, MH21, MPR22, NP18, PJ22, SS23b, SYLT14, VH92, WW14, YQM16, Zeg25a, Zeg25b, ZQS24, ZP23]. achievable [AHS22b, AHS22a]. acoustics [GX19]. across [Gor18, vSv94]. action [AH14, EHN17a, EHN17b, FAM25a, FAM25b, Fly22, ZWW25]. activated [MB09]. activation [XWC26]. Active [Góm01, ABM10, Góm99, GLV05, LG08, LL22a]. active-set [ABM10]. Adam [Iid24]. Adams [CVA01, DFF04, JXZ26, LRY18, MJ20, MG11a]. adapt [ART19]. adaptation [Pol10]. Adapted [LW13, CYM22, KK16, LZ25, Lin05, YW17]. Adapting [SH12]. Adaptive [AR24, ASGJ^{+}20, AKKT16, AKQ17, Bör26, CMR03, DZ01, FGP91, HSZ03, IJE15, JR10, Kac18, KLT03, KL22, KLS17, Kun05, LHZ^{+}21, LGC24, MRV23, Mal21, MSZ20, NR14, SGS25, SW00, SC03b, Söd02, SVZ05, Ste24, Sti18, WO00, YYZ22, ZS03, ZZL17, ABB15a, ABB15b, And18a, AV19, ASHF21, AEF^{+}14, BKR18, BKDA25, Bac20, BD17a, Bno21, BMV09, BGR25, CR96a, CHY19, CM99, CD25b, Das19, DDG05, DH26, FSY23, FS21, HHHN07, HOW95, Hei07, IL05, JL26, JA22, KE16, LY18, LLC20, LH23, MZG26, MAS17, MD21b, MKBY19, MC05a, MS23a, MP13, Moo07, NEY26, ODL21, PSZ23, PS22, PSW11, RFS23, RTTH22, RT24, RS25, Sch26, SSH^{+}19a, SSA25, TKSG23, TG20, WQ23, WBY25, XZZ22, YF22, YHS18, Ye96, YZLC24, YQM16, ZHSX23, ZH17, ZH19, ZW15, ZZWK12]. Adaptive [SGS26]. Adaptively [Lun23, BK08, WHS23]. Adaptivity [Coo03]. Addendum [BRS92, SS01a]. addition [GWBC20, MW16]. additional [GP05]. Additive [BCW13, CZ94, CZ95, CGV22, Axe99, BB14b, CZ96, CG19, KLL10, LZ19a, MKBY19, NG23, NT21, Prz16]. addressing [ST21]. ADI [BKS13, BPS23, DZ13, DW22, He16, KSJ26, QQX23, SzS21, YWWR12, ZWFY19, ZLM25, ZM26, ZZY24]. adjoint [BN18, Fly22, GNS22, JCH23, LG95]. Adjusting [CZH22]. adjustment [LHZ20b]. ADMM [AG23a, BBC21, BF20, BDD20, CWHL20, DCG25, JFSD25, PV23, SW24b, XYZ14, ZL24, ZL26b]. ADMM-based [BF20]. ADMM-TGV [ZL24, ZL26b]. Adomian [GNH10]. adopting [MZG26]. Advanced [Ano93, BvLP16, SAC18]. advection [AD22, ETY98, GM06, HZPW23, HLTA16, JBJB17, LLY26, MA13, PED15, RS25, RS26, SLA11, TL24, WZZQ24, dFO11]. advection-diffusion [AD22, ETY98, RS26, SLA11]. advection-diffusion-reaction [GM06, WZZQ24, dFO11]. advection-dispersion [HLTA16, JBJB17, MA13]. AE [DMD16]. AE-solution [DMD16]. aerial [DIM22]. aerodynamics [BZ18, Gau12a, Gau13a]. Affine [NKS04, dFS04, AK16, BCK06, CR23, LZ18a, MS26, Ska13, SH17]. affine-convex [BCK06]. AGE [Moh10]. agent [BWB24]. agent-based [BWB24]. agglomeration [CL00]. aggregate [JYLC21]. ahead [BRZ96, RSCH^{+}19]. AHSS [WZ13a]. aid [NR24b]. AIDS [DJM^{+}18]. AIDS-related [DJM^{+}18]. AINV [Meu02]. air [CV15, NZ19, HJ18b]. Airy [GST02, Tem97]. Airy-type [Tem97]. Ait [LGL23, ZWW21, JLWZ25, LCW25]. Aït-Sahalia [LCW25]. Aitken [BRZ19, CM92, GM92a, PC13]. al. [CN16]. algae [LLY26]. Algebra [ARJ03, De 02, AR18, BS25, BM96, BCGVS11, Cha14, CDW95, GHP^{+}00, Reb97]. Algebraic [HL03, Jbi03, KB02, Mai01, Tsu02, WLW^{+}24, AM98a, AM98b, Axe99, ANA14, BCL00, BBQO07, Bel94, BEQOR14, BDH^{+}13, CGPM00, Che16b, CNR15, Cou15a, Cou15b, CHS25, EH24b, EL01, GP99, GKRS22, Gau12a, Gau13a, GH09c, GW24, Guo13, GL23, Han22, HM22a, HM22b, HM26, HJ18b, IY15, ICR06, JH22, KPD25, LS14, LN95, LL26b, LT20, MS06, NK21, PT19, PV22b, Pié99, Pis16, Pog98, Sal96, Sid07, TFPG19, Yan95, ZY21a]. algebraic/logarithmic [Gau12a, Gau13a]. algebraically [Hil10, Kno23]. algebras [BM25, Kha13, Rob92]. Algorithm [BD02, BD04a, BS04, CM01, CCG01, CSFC04, DZ01, DR04, Dos03a, DL04, Góm01, HSZ03, HC03, KS14, KPFG04, LJW17, MP02, MN01, OKP21, PS01, PW04, Rob02, Sab03, SGM02, SBW98, SC03b, Ska13, Tir02, Van03, AGS08, AG15, AAU26, AA16b, ASW06, AS08, ALV20, AZ19a, AC94a, AAA25, Alt21, ART19, AKB15, AM98a, AT12, AHL20, AJMP11, And06, And08, And10, And14b, And15, And18b, Ant18, AABTB23, AM13, AE18, AB23, BPR22, BCCP25, BHLZ21, BBC21, BC94, BAK25, BG11, BD17a, BCK06, BGRS09, BE17, Ber14, BM09, BM24a, BH92b, BD00, BSF17, Bos21, BL93, BL95, BC16, BP93, BMV09, BS91, BZ94, BRZ19, BRZS23, BK16b, BK08, BF93, CL92, CDT10, CZL26, CMD19, CMM15, CFR06, CP95a, CM99, CF05, CE94, CGHH21]. algorithm [CS94, CC06, CHHL18, CHH^{+}20, CXC25, CPZ14, CC18, CCHH23, CS24, CGW25, CSC26, CM96, CRN19, CCJ99, Cho17, CW25, CDGC25, CGM93, Cor91, CHH93, CPS12, CYWH25, DDS93, DWZ14, DW15a, DW15b, DCG25, Dax09, DIM22, DB98, DY93, DTI09, Den14a, DN24a, DM21, Dey23, DEM94, DGL06, DLYH17, DMS09, DS12, DF01, EGT25, EG19, EAB20, EEM20, EM07, EO94, EG94, FZLL23, FLMR99, FZ26, FGL19, FM93, FC01, FDV13, FYI^{+}12, FS21, Gaj05, GGV96, GV99, GHM16, GS19a, GH06, GH09a, GJV17, GL12, GWW15, Gha18, Góm99, GCFF95, GM97, GI10, GLW13, GZP18, GLC22, Gug96, GH22, Guo16, GZD24, GEA20, Haj16a, HW00, HP18a, HHLS21a, HHLS21b, HDL23, HLW26, HA16, Hie18, Hie19, HRAH22, HN16, Hof21, Hom92, HL23a, HL23b, HM18b, HM19a, HFW^{+}21, IS22]. algorithm [IR13, JL12, Jbi93, JJ13, JL16, Jia20a, JXX^{+}23, JWZ23, JLZZ23, JXNX24, JWCZ21, JTW25a, JTW25b, JML25, JA22, JRRS08, KSV23, KST06, Kar10, Ke21, Khe12a, Khe12b, Khe14, Khe17, KMA13, Kim26, KLS17, KMS23, KLR07, La 17, LM11, LV15, Le 92, LDN16, LMV00, LL93, LLZ94, LWM10, LHZ20a, LCW23, LWLW24, LSL25, LWZ26, lLhYfD07, LHZ20b, LRL22, LYY12, LS07, LYL15, LLLD17, LQ20, LF25, LCW21, LS25b, MLM25, MS26, MZG26, MD21a, MR09, Md12, Mat91, MP92, Mat92, MS23a, Mel10, MZZ25, MS17, MSS18, MEJS19, MS23b, Meu97, Meu99, Meu05, Meu09a, MS14a, Meu20, Meu23, MMU20, MK94, MA22, Moh10, MN22, MN23c, MLM19, MvS09, MH22, MN92, MS24b, NIN12, NP18, NYL26, NS22, NEY26, gOM14, OOO11, PSS10, PSWE23, Pan20, PW22, PV23, PSZ23, Pei24, PYD23]. algorithm [PS06, Pet01, PH14, PDS^{+}23, PCDH20, PS09, Plo93, PW16, Pol10, PPV09, PP21, PSS22, PM24, QZG^{+}19, QXGZ20, QAS^{+}24, Rah11a, RT20, RTTH22, REE26, Rim25, Rob98, Sad99, Sal96, Sal25, SD20, SA14, SIE16, SDW24, SD26, SS23a, SDMMK18, SH17, SS23b, SJ14, SWJ24, Soo15, SZ99, SU14, mTLbJlL14, TOY26, Tom92, Ton25, Tou98, Tra93, VH10, VH12, VZ93, Vep08, VPA25, VMMD21, WZZ16, Wan17, WGZ18b, WZVJ22, WLW^{+}24, Wan24, WZW^{+}24, WGZ25, WBY25, WLG26, WLL26, WCH15, WZS14, WZ23b, XZL12, XYZ14, XTH07, XW17, XH21, Yak94, YIY22, Yan17, YCL17, YLL20, Yan22, YDW26, YJS25, Ye96, Ye22, YCW^{+}19, YA24, YP23, YY13, Zag92, ZH22, ZSGJ25, ZBDK23, Zeg25a, Zeg25b, ZYGQ17, ZLQT19, ZLWZ21, ZZ22a, ZFH23, ZJJW24, ZGLH24, ZHW25, ZZX26, ZH19, ZZX^{+}23, ZDSY20]. algorithm [ZS22, lZfLqL26, ZLLC11, ZZWK12, ZFH26, dCOS21, dAR06, dC20, dC22, van93]. Algorithmic [Kub23, LM11, SAE19]. Algorithms [Ano93, Ano95c, AHKW04, BFGM03, BHM05, Bou17, Bre04, CGR12, CZ95, CLaL00, CP95b, DHJJ10, DF94, DHMS16b, EGSV04, GPP01a, Gau11b, HR03a, HSS04, JMS16, MCG^{+}04, NRS12, Sas93, Tov98, dSCS04, AH23, AN17, ART14, AL09, AV19, AMR15, Ari98, AHKW05, BKFMA11, BS21, BBQO07, BFK^{+}09, BFK11, BG24, BvLP16, BWC22, BESC22, BEL23, Bra96, BRS91, BRS92, BRZ98, BRZ17, BN18, BM24b, BKL10, Buo17, CSI16, CSI17, Cal20, CPP14, CB13, CK20, CJK22, CD25b, CH11, CvPS15, CCW21, CMWP20, DM98b, DG17, DHF21, DZS21, DLL^{+}24, DB06, DZH23, Dur93, El 18, FT24, FHH05, FH05, GA08, GGN14, GMZ19, GPGC98, GST17, GL21, GLS^{+}18, GS26, HLL22, HS20, HH05, HIS26, HWD25, HZ95, HR00, HC24, Ihs07, Iva17, IUM^{+}19, JU22, JW25, JLX22]. algorithms [JY23, JCH23, JK19, JH22, JM18b, KP96a, Kar09, Kar13, KJC18, KR23, KPC20, KD18, KGG25, KSW09, Kuh13, KLW^{+}23, LY18, LL22a, LLWC24, LJX25, LB25, LL20b, LCHH21, Luc06, LHW13, MP99, MVV05a, Mer92, MDR23, Mil18, Oar94, PKC18, PT24, PP16, PT17, PLH20, PW14, Pop18a, RTD^{+}21, RT24, RR00, RW11, SKA23, Sab91, SCD^{+}21, SKK21, Sal94, Sal05, SW10b, SI18, SSN^{+}12, Śmi09, SW05, SJW21, Spr01, SBJC19, SLZ25, TKSG23, TQY21, TQW24, TCO26, TG20, Tas93, Tem97, TH19a, TH19b, Thr92, TLD22, TMS25, Vul97, WZ11, WYP23, WQ23, Wim99, WWM21, XZZ19b, Yan18, YAY26, YHR26, YZX25, YWS20, YYZ22, ZTW19, Zag24b, ZLH22, ZFZ19, ZMZ26, le 91, Pop19, PP17]. ALI [GW24]. aligned [CT21]. alignment [XQZ24]. Alikhanov [TTP26]. All-at-once [PQS22, GZL24, HFDSC24]. Allen [CS22, CLY25, HZX21, LZM23, LGC24, LHlLG24, SZQS23, SZ24, TZ24, XF25, ZYQ^{+}21, ZY23, ZQS24]. Almost [DM21, GMP92, GL19, LM17b, AAAA^{+}18, BDH^{+}13, CEX14, DFJP10, FHC21, HvD93, Van07]. almost-uniform [FHC21]. Along [CG03, DF94]. alternate [OdZdRV13]. Alternated [TQW24]. Alternating [BCN^{+}16, BF17, HR05, JHLL15, LZZ19, Lui02, ZFZ19, AC19, Bno21, Bru93, CTS09, Che19, Cui13, GZD24, HS16, JN99, KL25, KPFF24, LWZ18, LRL19, Ma20, NSM20, Now19, RWTM21, TYSY20, WLMA21, Zha15, ZZY18]. alternating-direction [WLMA21]. alternative [LCW21]. Alternatives [HHLS21a, SU14, HHLS21b]. Always [GA15, Gal18, SH26]. American [CL10a, SPB26, SMR26]. AMG [DDT25]. Ampère [BS19, SG10]. amplitude [Lin98, Lyn08]. amplitude-phase [Lin98]. amplitudes [Pan96]. analogue [HT16]. analogues [GA08]. analyses [LXZZ21, ZJ08]. Analysis [ADP26, ARS26, ABI20, Bac21, Bel03, BH01, BV96, CA22, DL08, DAM16, DZ13, DFF04, Flo03, FEK^{+}23, GL04, HK06, Ila20, KS18a, KK16, LV01, Li96, LWY26, LRL22, Lin09, MD21b, MM12, Mor25, RS02, Sad05, San25, SS24a, SJ14, SJW21, SWJ24, TPLB22, Wei18, ZLW^{+}13, ZKD04, AS10, AJ13, AG15, AAN25, ALZ21, AR09, AH10, AMKV96, BD09, BD10, Bac24, Bac25, BVV14, BC05a, BM22, Bel99, BF00, BBd95, CWZ13, CR00, CM24a, Cao12, CM99, CWL16, CCJC18, CWHL20, CDS20, Cou15a, Cou15b, Cui13, Das19, DJM^{+}18, EG18, EHN17a, EHN17b, FY13, Fan26, FHH05, FYI^{+}12, GM20, GKL21, GS16b, Van12, Van17, GS21, GLS^{+}18, GSF26, HÇK26, HZPW23, Han94, HSTW14, Hie18, HD18, HZ20, HM19b, HS21, Ibr25, Ihs07, Iid24, Iva17, JSZ22]. analysis [KXXW21, KCBT21, KR23, Ke21, KL17, KPC20, KGH14, KPS22, KXD25, KKA17, LLY26, LK20, LSX10, LW14, LWLW24, LCZ26, LM14, LZ15, LZ18b, Lin16, LLL22, Lis25, LGP11, LLLD17, LRY18, LWS18, LS20, LHlLG24, LZCW24, LGZ25, LGM25, LRM16, LZ22d, MA16, MP99, MV13, MV14, Man10, MR12, MM09b, MZ19, Moo20, MSMS12, Mot14, MRS25, MGG25, NIN12, NS24, NUG25, NV21, NEMS14, NAA19, NR24b, OL23, PKC18, PH14, PK22a, PK22b, Pis16, PZ20, PSS22, QAS^{+}24, RR08, RA12, RZZ25, RK11, RR22, Rou25, SSR25, SP21, SLW13, SKD26, SMK14, SL21b, SW22, SK25, SDMMK18, Smi97, SPS26, ST98, SLL22, TT21, Van19, VH92, WSK14, Wan15a, WZQ17, WXQ20, WYZ22, WZ22a, WLZ22, Wim00, WWD^{+}12, WC13, XXW17, YLYZ23, YHL25, YCZ26, YY24, YZL25, YWS20, ZCT19, ZJZ20, ZD21, ZYJY22]. analysis [ZW22, ZLLH22, ZWLZ24, ZGK^{+}25, ZZW26, ZFX14, ZWX19, ZL26a, ZL17, ZLWQ09, ZZY24, ZAC26, ZLTA16, dC20, dC22]. Analytic [CCLi16, KK23, SVZ01, Ash19, AP21, AAH24, GS19b, Van17, HW18, HSL19, Pas99, Pet95, XCLA15, Yak94]. Analytical [AG03, MN23a, WXQ20, Wan25, YX11, AS11, ALB^{+}18, AHC05, BC17, CS12, GPAA14, HV15, KM09, LJ11, MFK^{+}15, MSM12, QLZX11]. analytical-numerical [GPAA14, HV15]. Analyticity [Cro03]. analyzing [ZSGJ25]. Anchoring [Hag13]. and/or [FHH96, GLRSG08]. Anderson [And19a, HS16, LZOY22]. André [Bre06a]. Andrei [DW15b, DW15a, ZDSY20]. angle [HDL23]. anharmonic [EAGS20]. Anisotropic [vLV02, BM24b, CTS09, CS22, LHW13, MP25, VMMD21, Yua21, ZY13a]. annihilator [Cou15a, Cou15b]. annihilators [MJF09]. Announcement [Ano00c]. annular [BQ19]. ANODE [BJK^{+}24]. anomalous [CLTA10, CC16b, MA12, SLLA15]. ansatz [HM22a]. ANSI [DCM^{+}13, DCMM13]. Anti [Bou03, MSM25, AR18, CZL26, FFSC25, Spa24, YLY12]. anti-bisymmetric [YLY12]. anti-Gauss [AR18]. Anti-Gaussian [Bou03, MSM25, Spa24]. anti-reflective [FFSC25]. antilinear [SD26]. antitriangular [BK16b]. any [MW24, MW26]. Anymatrix [HM22c]. AOR [BH22]. appearing [Gha24]. Appell [CL13b, CGN22]. applicability [AH11b, AGS20, HL06]. Application [CDF99, CP93, CP95b, DH04, GM03, HR14, KH11, PG05, RBN14, Rec01, SC03b, TC05, AAM24, AIA^{+}24, AWH^{+}24, AA15, AB98, ÁCL11, ALRT16, ABK22, ABKD23, AG00, BGRS09, BQ19, BHS14, BH17, Ber14, BGS24a, Bou17, BZ18, Bru93, CJKL23, CL24b, CHH93, Cro92, CBGVPP09, CZX26, DPR23, DPP19, DNR15, Fly22, FDFM23, Gás99, Gau12a, Gau13a, Gha24, GHP^{+}00, HS26, HM26, HVMT17, IAAH24, Ila20, JBJB17, Jia20b, Kri26, LSL25, LLL26, LPXX19, MMV19, MK17, MP13, MSS11, NH24, NRV23, NAE22, Orb15, Pen98, PLH20, PS25, PPR15, PR93, RVF07, Sat24a, Sat24b, SD25, SK19, USAF14, VH92, VL19, WLL12, WLS25, WLG26, XyJl16, XZW13, XW17, Yak95, ZWL26, ZHT15, ZCGS24, BE98]. Applications [HLW26, IMT02, MAL04, PL04, dFS04, vdHS02, AHJ17, ALW98, ARS26, AR18, ABT07, AMKV96, Arn97, BGR23b, BBL22a, BJNKR20, BEJS21, BF18, Ber10, BZ91, BS92, BM96, BRS08, BRZ17, CM98, Ceg24, CS24, CRV91, CK06, Dah93, DG17, DJS20, EDAH12, FLG08, GGJL24, GCFF95, GL23, GSF26, HKP25, HLV26, HL17, HSY23, JS15, JLJ22, KBA26, KADE18, LS14, LG08, Leo08, LWwCL13, LWLW24, LRW25, LZZ19, LZX22, LSY^{+}23, MJJ^{+}23, MBJ17, MA15, MRS93, MM08b, MM09a, MZ19, MS11, MvS09, MA12, Oar94, OOOR12, OB16, RFS23, RB21, ROB18, SKK21, SKK25, SRK^{+}25, SIO20, SD26, SS11b, SKJ^{+}18, TQY21, TG20, TOY26, Tas93, TA96, TRSI23, Ton25, TTA26, VS19, Wal94, WZVJ22, WYP23, WZ22b, YAY26, YJJ^{+}21, ZAGD22, ZJJW24, ZLZ22, ZL24, ZL26b, dAR06]. applications [dBGKR08, Dra96]. Applied [CR03a, Jay02, LM04, AM01, BM22, DW21, GS16b, HHST19, MPS20, MS24a, MRS25, PV23, PRVI20, RSZ20, SSH^{+}19a, SSH20, ZH25]. Approach [CGN03, Kun01, MKO04, NPP04, NW04, Rob02, SR04, TDKB24, TRRD02, WHL24, AAN25, AAN14, And97, ABV23, AD00, BDN17, BBO21, BQ19, BL92, BE17, Ber14, BR07a, BMS24, BN18, BWB24, CCZ23, CKS24, Cia94, CL96b, CGV26, DOR26, DZW17, DLC14, FPP05, FJ96, FGBP21, GPAA14, Gor18, GM96, GSF26, Hal14, HZX21, HFZ19, HZX20, JKC26, KS25, KPD25, Kar07, KLT95, KM24b, Kha25, KADE18, KE16, KXD25, KV07, LAG05, LZOY22, LCZ26, LJ11, Liu21, LZZ23, MWDD24, MS92, MdR08, MFBB23, MMLM20, OL21, PV22a, PSB26, Pat25, PZ20, QW08, Rab23, Ria16, RB17, Sal96, Sat24a, Sat24b, SL18, ST18, ST22, SN24, SKD26, SW22, SR24a, SKM25, SS16, SWG20, TCOA19, TLR^{+}25, TCW14, VDVJB12, WZ11, WJW14, WZL26, WZ22b, XCLA15, YXL18, YP09, ZHFW21, ZW20]. approach [ZE10, dFG93]. Approaches [MHZ05, AAB13, AE18, CB16, Kuh13, LP18, LL20c, LL22c, WMCW21]. Approaching [Rec01]. Approximant [DV01, Wen03, Van92]. Approximants [ACO03, Bou03, Dra02, MC05b, MC05c, AH17, BL92, BBPV12, BC00, Bre99a, BGVHN96a, BGVHN96b, CGV92, CAB22, CJTW96, DM97, GCGVH92, GMT92, IT93, Mat96, PP92, Sab14, TBA94, VB91, dDL92]. Approximate [AHC05, CA07, ÇD25a, PM22, Sau07, vdHS02, AM24, ABG97, BD17b, CM99, CDD21, DBH21, EV22, FGM19, FP18, Gar24, GK20, HN94, HK14, KST21a, KST21b, LL14, LL18, MKG24, Mil20, SW00, Śmi06]. Approximated [GCPG99, BRY14]. Approximately [DLYH17, GLW13]. approximates [SH21b]. Approximating [AH03, AF94, AT17, Bag16, DV01, FGS24, Gar19, HC03, JB22, MT19, PDRG19, SR06, VBG96, Yak94, ASS11, Ash19, AP21, AK26c, CM05, FRR07, GK24, LX17, Rim25, SCTP00, SRK^{+}25]. Approximation [All03, Ano95c, BG03a, CN01, CP01b, Cro03, CMP22, Dab04, GAM24, GCGF03, GLRSG08, GM03, Gro93, HL02, KLT03, Kva01, LM00, LRZ12, MN01, MAK20, PP05, PW04, PV03, RKM25, RS93, SBR26, SVZ05, Str02, VGM96, Wit96, ZKD02, AHM21, AMH10, AMA21, AAPR21, ALRÁY25, ALZ20, And97, ABV23, AG00, AGRT05, AAD14, BDS25, BBL22a, BC06, BF20, BEJ20, BEJR23, Ber93, BK97, BG13, BM96, BM23, CCV07, CT93, CB13, CR20, CLT^{+}13, CGW25, Cuy00, CY10, Dar99, DFR26, Des17, Die08, Dri93, Dun94, ED22, FGJ00, FZ07, Faz23, FM19, FM93, FT05b, Gaj05, GGN14, GGNF17, GH23a, GNS22, GA20, GH09b, GH95, Gla01, Gug96, GTA19, GI97, HH11, Haj16b, HW00, HIS26, HGVPA92, Hof05, Hof21, HL23b, HG25, IAH20]. approximation [IDAV09, IL05, JK18, Jaw22, JSZ22, JJK97, Kal00, KJ18, KR23, KM24a, KADE18, KGG25, KLS17, KGN^{+}24, KP25, KLR07, LGL23, LWK12, LKW17, LXX23, LGHS25, LCW20, LSM16, LWZ21, Lya97, MSCB93, MRV23, Maj13, ME95, Mel24, MS23c, MS15, NCC11, Ost07, PV22a, Pan26, hPwL09, PPS25, PGGC97, PS22, PK25b, PK25a, PGC24, PS00, Prz09, Prz16, RIAA19, RS20, RS06, Reb97, Rei97, SI13, SS11b, SR22, Sle25, SR24b, Ste24, SWG20, The12, TV19, Thi93, THS20, VH10, VC92, WC24, Wat92, Wat93, Wat94, WKY25, XyJl16, XHZ07, XLC93, XC25, YD09, YH21, ZJR25, ZWWW20, ZZ18, ZLCW23, ZZY26, ZLZ22, dBGKR08, van93, Ano92, Ano93, DH18]. Approximations [EKM03, Sab03, Zil01, AB99, AR26, ART19, AGG17, Ber10, But96, CS12, Col92, CT21, Ell93, ED13, FW13, GV99, GH25, GO21, GGS22, HV15, JKNR13, JLP20, JL15, JLWZ25, LLWC24, Lig93, LM15, Mac96, MK17, RB21, Riz18, Rob97, RRZ21, Sch14, SLA11, SHGL22, TV17, UTO07, WQL20, Xu19, YZ21, YX11, ZAGD22, Zha23, ZWX19]. aquifer [LCW20, SMNZ20]. aquifers [SFZ22]. Arbitrarily [ABI22, MKO04]. Arbitrary [BWC22, ANI^{+}17, ABV23, VV07a, VV07b, FT14, Joh20, KKV22, SZ99, TB25]. arbitrary-precision [Joh20]. arc [BD25, FRR07, KOK21, SD20, YIY22, Yan22]. arc-search [KOK21, SD20, YIY22, Yan22]. arcs [AS08]. area [CKS16, FRR07, MW24, MW26]. Argument [Mat04, Bic24]. arguments [BS17, GRAST23, GWY26, SSL24, ZA24, Zag24b, ZY21a, ZSLZ24]. arise [FS20]. arises [Ila20]. arising [ABG97, BBO21, BC06, DI11, Fab16, FM99, GM23, GNT24, Haj16b, JH22, LM14, LXZZ21, lLHNS23, Lin98, MM99, RCW22, SGO22, SG10, SMA99, SPS26, Wim00, YDW26, ZZY18, Zha26, ZWY22]. Arithmetic [ASS04, ACM04, DHL^{+}04, KL04, MAL04, NKS04, Vig04, dFS04, vGK04, BBZ95, BMR19a, CM96, GLM15, Ihs07, Joh20, NAA19, OO22, RR23, ST26, Sv95]. Arithmetical [Nov26]. ARK [But98]. Armijo [DW12, SW11]. Arnoldi [AHJ17, BOR23, Che04, CGW25, DBGB11, FEL15, NR14, TM14]. Arnoldi-based [BOR23]. array [ZLL^{+}21a]. Arrow [XY25]. artery [LC21]. artery-like [LC21]. artifacts [HOW95]. Artificial [Ano95b, JY23, LZ24, CK22, LHR20, PJ22, ZP23]. Asian [AAA25, BN25, Che22, Che24]. ASOR [FGC19]. ASOR-like [FGC19]. aspects [BD95, BBB^{+}06, DL97, KBA23, KBA26, LMD25, Mil13, Mon01, Sal05, SGO22, ZAGD22]. asset [ZZ22b]. assignment [PV23, YZX25]. assimilation [BH11, BGS24a, NZ19]. Associated [BBR03, ALV20, BEGG91, BHJTM92, BPR20, BGVHN92c, BGVHN96a, CRV91, GAM24, LM12, Mar92, PK25b, PK25a, PGC24, QQX23, Røn92a, SS16, VP23, YC22, ZBDK23]. assumption [PSS25]. assumptions [CMP22]. ASTRA [BvLP16]. astronomical [LP08]. asymmetric [JS21, JCH23, lZfLqL26]. Asymptotic [AA16a, BG03b, CG03, DD99, DHM12, KBP17, LMMD05, PZL15, Sab92a, Wal94, CF05, Che13, CS12, FLMR00, GEP19, LZZ24, QLZX11, QW08, Sid20a, Tem97, VH92, WQL20]. Asymptotically [AHR21, Bal11, CSI16, CSI17, Den14b, EM10, GMZ19, Hom98a, PKC18, PS22, SKK21, VA20]. Asymptotics [GG01, Han96, JNW92, Nor00, Che16a, GG08, Wil12]. Asynchronous [AAIT94, BMR97, CFS21, SME03, AAII96, BBP17, HD18, JRB17, SX00]. attachment [BM00]. attenuation [Ste95]. attraction [GHPMGRR14, Mor25, Par16]. attractor [MMU20]. attractors [JMS16, KS97, PK25b, PK25a, PGC24, dCOS21]. augmentation [Cao12, LJbL21]. Augmented [BE98, HSS04, AHS22b, AHS22a, BRR13, BW13, FGC19, HZ15, JCL16, LZZ24, LGZ25, TYSY20, Wit96, XWY19, ZXL25]. Author [Ano01a, Ano01b, Ano02a, Ano02b, Ano04a]. auto [LLY25]. auto-convolution [LLY25]. Automatic [Coo03, FRJT09, HR03b, Hof16, Maj18, Mar93, PS22, PSW11, Söd02, TRRD02, CPZ14, EO94, EG94, HS96, JL12, Moo07, RC14, SL15b, SL15a, TY21]. automatically [OOO11]. autonomous [Jay21]. autoregressive [LDN16]. auxiliary [BVV14, GZL^{+}25, LL22c, Sat24a, Sat24b, YK22]. average [ART14, ALL25, AMR15, AH11b, JSZ22]. Averaged [DFD23, AC19, DDRS23, Fan15, MN17, Spa07, Spa20, SH26, ZAC26]. averaged-rule [SH26]. averaging [BRZ18, BGR23b, DXY18, JS23, RZ16, TL23, XTH07]. avoidance [DIM22]. Avoiding [BS91, BRS91, BRS92, GMS99]. axisymmetric [CJK22].

B [AZ19a, AG19, ALRÁY25, BLS06, DLR12, FW25, Gau17b, MST03, Maz05b, Maz09b, MT06, NS01, PZL15, Rad08, Rei97, WZ19, XL14]. B-spline [AZ19a, ALRÁY25, DLR12, FW25, Gau17b, Maz05b, PZL15, Rei97, WZ19, XL14]. B-Splines [NS01, BLS06, MST03, Maz09b, Rad08]. backtracking [And06]. Backward [IM02, Pot19, TTV21, WK20, dC22, AH21, AHP25, ALV20, BC16, CCN25, CS24, Cho16, HYT26, JCH23, LZ22d, Mil19, NBJA17, Pan18, SBG24, SB21, SCF23, SSP15, SZ20, Ton25, YW17, YJ21, YLYZ23, YAY26, ZZW26, ZWW21, ZL17]. backward-adjoint [JCH23]. backward/forward [ZL17]. backwards [Sza92]. Bacteria [Car01, SL21b]. Bagley [Mok16]. Bakhvalov [LZ23b, LZ23c, NV21, TKS24, ZL22b, ZL23]. Bakhvalov-type [LZ23b, LZ23c, TKS24, ZL22b, ZL23]. Balanced [CYM22, MS24a, AH15, BCL00, CMM17, HH12, Kao20, LS15b, ZXL25]. Balanced-norm [CYM22]. ball [ACH14, LPP21, CZH22]. balls [CZH22]. Banach [Jai17, CSI16, CSI17, CRN19, Cho16, DN24b, HCD26, HM14, IAH20, Jai16, JA22, JM18b, LLL22, LCL21, MWsC19, OAMA22, PG12, PDRG19, SSS14, SIE16, SC18, SCS18, TAM21, Tak17, TB19, THT19, VA20, WGK11, WKG11, WK12, WK17, XCD23, YAT20, ZG12a, ZG12b]. Band [BO03, SC03a, ANI^{+}17, AAAGAD23, DS12, FC01, Laz99, RSCH^{+}19, Soo15]. band-limited [AAAGAD23]. Banded [BM12, IMT02, JBJB17, LLM25, AAAA^{+}18, CGSS24, DMT13, DPP19, EG19, HN94, JRB17, KD14, WSLL24]. Bandlet [MP07]. bandlimited [Gor18, IP16]. Bandwidth [DR01, FLT09]. banks [Tur94]. barrier [CWZ13, Khe12a, Rou25]. Barycentric [Bal11, JKK^{+}08, Tia21, AH21, AAP24, AHP25, Ber93, Ber00, DHWL25, Hof21, LC14, LX23, ZH25]. Barzilai [AX26, HLZ14, HL15b, LLS11, MR96, MPB16]. base [KC23]. Based [CR02, CPV04, Dos03a, GPP01b, KPFG04, MT04, WDY04, dADdRC04, vGK04, AD22, AHP20, AH21, AAU26, AF13, AFG26, AZ19a, ALRÁY25, AKB15, AK15, ABKD23, AT12, ASGJ^{+}20, AE18, AAD14, AGG17, AKT15, BKFMA11, Bac25, BZ13, BAK25, BBQO07, BF20, BuR24, Bra96, BN18, BOR23, BWB24, BK08, CWZ13, CR12, CW19a, CJSZ14, CK05, CF05, CM16, Che16a, CZM21, Che19, CJKL23, CDS20, DDS93, DWZ14, DW24, DG17, Das19, DP21, DR07, DMW23, EH97a, Erb15, ES19, FJ25, Fan22, Fan26, FGM19, FLV14, FM93, FC01, FYI^{+}12, GHM16, GLLJ12, GGJL24, Gar20, GWW15, GS14, Gha24, GS16a, HP18a, HHHN07, HMdÁES08, HA16, Hof21, HFDSC24, HST15, HS15, HLV26, HC24, ITA24, IDS16, IS17, IMT23, JL12, Jaw22, JXX^{+}23, JHLL15, JFSD25, KPD25, KST21b, KLT95, KMZ18]. based [Kha25, Khe12a, KH20, KBA23, KBA26, Lev95, LLS11, LY17, LHZ^{+}21, LD21, LLY22, LZ22c, LPRÁ24, LJX25, LZ25, LCZ26, LQ26, LX23, LDZ24, LZL22, LZL23, Lin01, LYY12, Liu11, LYL15, Liu21, LL22c, LTP18, Loh22, MKG24, MBG19, MdR08, MRS93, MR12, MT13, MT14, MG20, MG21, Mez22, MNS23, MS23c, MSZ20, MN22, MN23c, NKM25, NAA19, gOM14, OL23, OKB23, OR24, PKR20, PSWE23, Pan26, PM25, PH14, QXGZ20, ROB17, REM21, RWTW19, RS25, SBA26, Sad99, SKK21, SR25, SAE19, SL16, SL18, SB21, SW10b, SWS22, SD25, SD26, SR24a, SJ14, SJW21, SR24b, Soo15, SA23, Str05, SLL22, TY96, TA13, TD09, VH12, VZ93, VDVJB12, WT08, WL22a, WD23, WC24, WPW25, WL25, WCZ25, WZ22b, WL00, WPL18, WCD21, WL22b, WZ23b, XY25, XYZ14, XCLA15, XW17, XZP^{+}20]. based [XSL22, YYK24, YD09, YP09, ZRZ11, Zha11, Zha15, ZD15, ZSF18, ZLWZ21, ZY21b, ZZ22a, ZSWW25, ZHW25, ZWZ25, ZLL21b, ZY13b, ZLV17, ZV19, ZV21, ZZLV23, ZJSS25, ZLLC11, ZLZ22, ZH26, DF01, KST21a]. Bases [Osw01, CT21, DLR12, GS16a, LM00, Liu11, MPR22, MPR24, Maz99, Maz05b]. Bashforth [MG11a, MJ20]. basic [AAH18]. basics [HM22a]. basins [GHPMGRR14, Mor25, Par16]. basis [Bar13, BZS22, Bro05, BX19, CKP19, CLMM05, DS09b, Dze15, GCJ26, HM06, KMA13, KP22, Low05, MMV17, Mel10, Nar05b, NJ13, Pea13, TS18a, UTO07, WZ15b, YH97, ZJ08, ZZL17]. Baskakov [MP14]. BB [KP09]. BB-form [KP09]. BBVMs [ZY21a]. BCP [LZM23]. BDF [EG10, Eba18, Hea10, HWZ24, MWZL23, WLZ22]. BDF2 [JZ25, MS24a, SZ23, YSLL23]. be [Cat24a, Cat24b, ICR06, JB22, TM20]. beam [QLZX11, ZP23]. Beavers [DMW23, WDL23]. bed [CDF99]. behavior [Bel99, HLSLSL25, KBP17, Meu20, MT23, TM14, VVL25]. Behaviour [KZ03, AA16a, KGN^{+}24, Sab92a]. Bell [XZ25]. BEM [ADL05, ADG10, AEF^{+}14, VMMD21]. BEM-Fading [VMMD21]. Benjamin [ZJZ20, ZPEL25, ZLS24]. Bernoulli [BBQO07, CD01, Rab23, ROB17, ZP23]. Bernstein [Bar13, BZS22, CAV23, CW21b, DPS18, Dra00, DMS09, GS16a, Her96, IR13, JBJB17, KADE18, LWAG08, Mah10, MPR22, MMV17, Maz99, Maz09a, MAK20, Not08, Ost07, Pej14, TS18a]. Bernstein-type [Maz09a].

BertiniLab [BNN16]. Bessel [BBZ95, HGVPA92, JH26, KXXW21, LM08, Mül00, OM26, Saf10, SS23b, SH23, ZAGD22, ZuI19]. Bessel-trigonometric [JH26]. Best [GGNF17, GH95, Van92, Zha95, AHS22b, AHS22a, BDIR18, Car91, ED22, GH23a, Hof21, IDAV09, IT93, Kow00, VC92, Wat92, Wat94, Zha23]. beta [EGT25, GST17, KSV23, YA24]. better [GTA19, SH26, ZLLH22]. between [Cha14, DHS09, FLV14, Her96, MdR08, MSM12, MKA26, TAM21, Wen92, YZH21, ZS19]. beyond [BM09, CRZ26, OIM21]. Bézier [CFR06, GS16a, GLW16, LWK12, LKW17]. BFGS [AIA^{+}24, AB99, And18a, And22, BKFMA11, BK16a, BKA19, Deh20, LZ22c, LL07, Liu14, LTP18, VL19, WZ23b, You16, YLL22, YZLC24]. BFGS-based [WZ23b]. Bi [YWWR12, BC94, CBGVN07, PLVB11, SvF94]. Bi-CG [SvF94]. bi-conjugate [BC94]. bi-objective [PLVB11]. bi-orthogonal [CBGVN07]. Bi-parameter [YWWR12]. bias [MG22, YCW^{+}19]. bias-compensated [YCW^{+}19]. bibliography [Gau07]. BiCGStab [GM02, Sv95]. biconfluent [FS16]. biconjugate [ZD15]. bicubic [Bia94, CLaL00]. bidiagonal [CDW95, NIN12, YKY15]. bidiagonalization [BR06, BPP23, BJNKR20, HLJ25, JL21, JXX^{+}23]. bidiagonalization-based [JXX^{+}23]. BIEs [MSS11]. bifurcation [Bea98, CCJ99, Smi97]. Bifurcations [ARJ03, JJK97]. Biharmonic [AEG02, LGM25, AB06, ALZ20, BAEF25, Bia12, BFKM20, BFK22, DG17, Gás99, HS21, Kha25, LY18, LWY26, LKK21, MK17]. Bilevel [ARSS19, dSCS04, Anh19, AT21, TQY21, TTLD20]. bilinear [CXC25, WSZ21]. binary [GI10, KZ21, Str05, XCY21]. binomial [ASVC21a]. Binormalization [LG04]. Biographic [All08a]. biomechanical [AB98]. Biorthogonal [Ise96, Laz99, Bre99a, IN95]. biorthogonality [Da 92]. biquaternion [BAS25, HS26]. Birkhoff [BDS00, BCGVS11, GM92b, MT18, NB16, REE26, dS00, dDS00, dBD05]. birthday [BJK^{+}24]. Bisection [PL99, El 18, Mer92]. bisymmetric [hPwL09, YLY12]. Bivariate [CH95, CGN22, GM92b, Par16, PS01, SVZ05, AB24, ASV23, ÁFP07, AP21, CG05, CAB22, DL08, Len93, MMV19, Sch08, Thi93, VC10]. BKM [DR04]. Black [AAN25, AWL^{+}24, Cia94, KN23, ON26, Rou25, Val14, Val15, ZZ22b]. BLAS [CCZ23]. blender [CKO24]. blending [BDIR18, Zhu21]. Blind [DMT13, CPN14, MAF25, ZLS25]. Bloch [CCJC18, Pas99, SKD26, SYLT14]. Block [Bre02, BZ02, DM21, DP01, EJR02, GL04, GGV02, IS17, Jbi03, KKR25, LS03b, LM21, NJ13, RW24, TT21, TMS25, ZS08, AEH20, AHJ17, ALJLYJ24, ALL25, Bag00, BR06, BCW13, BL23, BH22, BGL26, BC06, Bel08, BH92b, BP19, BHS23, Bör26, BK16b, CR99, Cao12, CW19b, CH22, CK06, DW24, DCG25, DS12, DDRT97, EHTSM21, FP18, FDV13, GLG26, Hem94, HL17, HM19b, Hua20, JSF13, JXZ26, gK25, LRC19, LZM23, LLWC24, LB25, LL26a, LZ16a, LZ19a, LA22a, LZ19b, Lin01, LSW16, LXQ15, Lun23, LZ22d, Meh11, NK16, OM26, RKMS16, RR20, RN21, Saa23, SS99, ST22, Sol23, Soo15, SW24b, TL23, Uhl22a, Uhl22b, VL19, WGZ18a, XC25, Ye96, ZD17, ZFG18, ZSWW25, Zha26, ZWY22, JCF15, CO94]. block-centered [LRC19]. Block-circulant [RW24]. block-coordinate [LB25]. block-decomposability [Uhl22a, Uhl22b]. block-diagonal [BL23]. block-diagonalize [Bel08]. block-Lanczos [CR99, XC25]. block-relative [Bör26]. block-tridiagonal [Hem94]. blocked [CK20]. Blocking [BDS25]. blocks [Hem94, HL17]. Blossoming [AH17, AHM21, GS19b, Maz02, Maz05a, FJ96, GS16a]. blow [Cho17, CW25, HZX20, QL12, WYZ22, WWL24]. blow-up [Cho17, CW25, HZX20, QL12, WYZ22, WWL24]. blowup [CD15]. blue [FS21]. blurring [ED13]. BNS [VL19]. Bodies [Mar04c]. Body [Sha02, Lor19, PPR15, Sha19, AGA24]. Boltzmann [FEK^{+}23, CEZ26]. Bona [ZJZ20, ZPEL25, ZLS24]. bond [CS12]. Book [Ano95a, Ano97, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano02c, Ano03a, Ano04b, Ano05a, Ano06, Bre96, Bre97b, Bre97a, Bre97c, Bre97d, Bre00b, Bre03b, Bre06b]. Boolean [ALW98, BD95]. Bordered [DM21]. bordering [CD96, CM99]. Borges [TS18b]. Borsuk [CL24b]. Borwein [AX26, HLZ14, HL15b, LLS11, MR96, MPB16]. both [ASZ23]. Bound [AAAS03, DR01, Dos03a, GA20, JML25, LPGL16, MS26, MZG26, MT23, Odl00, PYD23, SZ24, TZ24]. Boundaries [KGD03, GK24, KZ21, ZLLH22]. Boundary [AKW02, AKKW03, Boy05, ED05, FHH05, GPHAPR24, GWY26, MP25, MT04, Nac03, NMM18, RS06, SC03a, SA23, TA26, ALQ17, BK18a, ATM19, Ahu09, AR13, AHC05, ABL… 12, ABI20, ACH19, AKPW05, AKQ17, Bac18, Bac20, Bac21, Bac23, Bac24, BD17a, BQ19, BAEF25, BBB22, BKF20, BFK11, Bic24, BEG25, BSF17, BCI14, CS99, CW14, Che24, CH11, CK22, CD07, Cro92, DD20, DD21, DNV25, DMC20, DL21, EH97b, EM07, FY13, FHAL15, FFSC25, GD15b, GX19, Ghe13, Ghe16, Ghe18, Van12, HJB18, HCL21, JL12, JP14, JWY21, JTW25a, JTW25b, Kar15, KCHD16, KS12, KKS22, KKD25, Kun05, LLY26, LZ24, LWY26, LJ11, LZ09, LS15b, Lis25, LSW16, LLC20, LM21, LRM16, LF25, LKKM15, LHW13, MS20a, ML20, MKA14, MMW20, MFK^{+}15, MO19, MN08, ML10, Mot14]. boundary [NZF11, NM14, PJ22, Rad08, RS20, RR20, RGJ10, RT19, RB17, SCF23, SSH19b, TS18a, TBY13, TPLB22, VT10, WZ19, WCM94, XWX24, YSLH19, YH21, Yse99, Zah09, ZE12, ZA20, ZJ08, ZW12a, ZZY18, ZBX21, ZZ23, ZFX14, ZXLF15, ZS22, ZP23, ZPEL25, ZZB20, ZS19, ZS13]. boundary-domain [RT19]. boundary-layer [Mot14]. boundary-type [CD07]. boundary-value [Bac18, Bac20, Bac23, Bac24, KS12]. Bounded [LER03, PW04, Ber10, CGM25, Coo09, Gaj05, Mül00, OKB23, ZLH22]. boundedness [Lóc18]. Bounding [MKA26, CR23]. Bounds [ADGP15, CRS04, Kol06, MW16, AH13, CMRS00a, CLWV15, CCL16, DLL12b, DLL12a, DLL13, Des17, DH18, DPS18, DBGB11, GWL18, GGJL24, GEP14, GEP16, GEP19, GS14, IDS16, JLMP16, JLP20, JJC24, KG23, KK22b, LZ14, LS15a, LL16, LDL17, LZ18a, LCVL18, LWG18, LL20a, LYH^{+}20, LRW25, LZX22, MS24a, Meu97, Meu99, MT13, MT14, MT19, PV22b, Pej14, PCDH20, RG10, Rum14, SS98, SW26, WW19, XC25, Yal01, YKY15]. Boussinesq [Bra07, CM24a, HEOS04, LZW20, Nat07, SW14, SLT20, YSLL23]. Box [Góm01, KM04, Fra26, HMS96, JJ24, KP09, Kim21, Kob97, KLB10, Lai92, LWM10, SEG14, Yan22, de 93, CFR06]. box-spline [Kim21, CFR06]. box-splines [KP09, Kob97]. bracketing [FLMR99]. brain [Ila20]. Branch [VR04, JML25, LCW21, MS26, MZG26, PYD23]. branch-and-bound [JML25, MZG26]. branch-and-cut [LCW21]. Branch-and-Prune [VR04]. branch-reduction-bound [MS26]. branched [WQ10]. branching [MZG26]. Bratu [KPA20, RMT13]. Brauer [San19]. Brauer-type [San19]. Breakdown [CM96, BS91, BRS91, BRS92, BZ94, DL97, GMS99, Jia20a, JXNX24, TY96]. breakdown-free [Jia20a, JXNX24]. Breakdowns [GM02]. Bregman [Alt21, BPR21, CE94, God26, GZD24, KS23, Ngu16, XCD23]. Brent [Hua94]. Brezinski [CHHL18]. bridge [CFM15]. Briggs [AM12]. Brinkman [Gha24, KL25, YJ18]. Brown [Hua94]. Brownian [CFM15, LD20, NM26, NT21, QQX23]. Broyden [Man21, Som05, TZ25]. Broyden-like [Man21, TZ25]. Bruijn [Odl00]. BSDEs [PKR20]. BTTB [LW12]. bubble [CJKL23]. bulk [FMS25]. bulk-surface [FMS25]. bundle [LPXX19, LXP20, MN22, MN23c, PW22, PL25]. bundle-based [MN22]. Burgers’ [KK22b, MP22, ZQzS22, AZ19a, AH18, Hei07, MDL15, PV03, RK11, SSA25, ZLS24]. Burmeister [ZLLC11]. Busemann [Bag16]. Butcher [BJK^{+}24, Kha13]. BVODEs [EM10]. BVP [Fab16, PSW11]. BVPs [AK12, Ghe15, MS13].

C [Das26, Kub23, OBAHK^{+}19, Ple03, YJS25]. C-FISTA [YJS25]. C.B.S. [AAAS03]. C1 [CDF99]. C2 [FKMS01]. C90 [DCM^{+}13, DCMM13]. cable [KXD25, LDL^{+}19, ZLTA16]. CADNA [Tou98]. CAGD [PMO05]. Cahn [ASGGRG23, BVV14, CS22, CLPY23, CLY25, HZX21, LZM23, LGC24, LS20, LHlLG24, SZQS23, SZ24, TZ24, WPY25, XF25, ZWWW20, ZYQ^{+}21, ZY23, ZQS24]. Cahn-type [LZM23]. Calculating [KKK22, Coo09]. Calculation [Lau07, MBM25, ZA24, BDL^{+}12, EAB20, MVV05b, Pet01, Zag24b]. calculations [CL00, HK14]. Calculus [vGK04, Gha16, Loh22, OB16, SM25a]. Calderon [BBB22]. call [ZLT^{+}17]. Camassa [YYW21]. Can [GR01, MKO04, Cat24a, Cat24b, ICR06, TM20]. cancer [DJM^{+}18]. candidate [KKV22]. canonical [AAAGAD23, AK26c, EDAH12, KW00, MdR08, VVL25]. cantilever [QLZX11]. capacitance [ÁCL11]. Capacitated [CSFC04]. capacity [LNS23]. capillarity [Wit96]. capturing [AC17]. Caputo [BZV16, Die08, DN24b, DNH^{+}25, GK21, GHZ25, LGHS25, OCWV24a, OCWV24b, PGP^{+}25, Rou25, SLA11, SKM25, VLCL16]. Caputo-type [Die08]. cardiac [GS19c]. Cardinal [Sun94a, TM05, AA15, Gau17b, Mil17, Rab92b, Rab92c]. cardinality [PL25, XP23]. cardinality-constrained [XP23]. Carlo [Mil20]. cascade [Dur93, MB09]. Cascadic [MClXzJe16, MRS10, WL22a, WPW25, XH20, YZ11]. Case [CP01a, Rev03, BDL^{+}12, CM98, DL97, Gau11a, GS21, HL06, Len93, Leo07b, Leo08, Mah10, MP13, Ost07, Smo99, Thr92]. cases [GGN18, Ste20]. CAT [LP18, PKC18]. Categorification [Gla04]. Cattaneo [RhG15, Wan19, Wei18]. Cauchy [BV93, BK18b, BCJ22, BM24b, CMD19, CS08, Fas23, HLN25, JN99, KL25, LKK21, Mel24, VMMD21, XZW13, XH25, XLG22, YHZ20, ZBDK23, ZW12a, ZW14, ZZ19]. Cauchy-like [Fas23]. Cauchy-polynomial-Vandermonde [YHZ20]. caused [Mot14]. cautious [LL07]. Cayley [Coo09]. CBC-based [GLLJ12]. CCD [He16]. Cebysev [Ber93]. Cell [AMR15, BDH^{+}13, ART14]. cell-average [ART14]. cells [DJM^{+}18]. center [Saf10, ZCS14]. center-Hölder [ZCS14]. centered [LRC19, WSLL24]. centering [KH20]. Central [GC04, CKP22, GST17, KSV23, LC21, iV12]. central-line [LC21]. centralities [BKM25]. Centrality [DMR21]. Centro [HR03a, LRT19, lLhYfD07]. centro-invertible [LRT19]. Centro-Skewsymmetric [HR03a]. Centro-Symmetric [HR03a, lLhYfD07]. centroid [Ron08]. centrosymmetric [GNT24, LHZ10]. century [Ste20]. Certain [ED05, SK04, BM25, CBGVPP09, CS08, KST06, Lu15, NW17, RA12]. Certified [Kie23]. CEV [YW17]. CFD [MHZ05]. CG [BH22, LP08, LLL26, MT19, MPT21, MT23, PT24, Rah11a, SvF94, TT21, TMS25]. CG-like [PT24]. CGME [Jia20b]. CGMY [CGSS24]. CGS [BS91, BZ94, CM96]. chain [GM23]. chained [LL16]. challenging [Gau08a]. Chandrasekhar [SPS26]. change [Ari98, CMR93, FJT94, LEK21, VPA25]. channel [MSMS12]. chaos [AE18, BK97, FM19]. chaotic [DJM^{+}18, GNH10, PK25b, PK25a, PGC24]. characterisation [BCM19]. characteristic [ALW98, HZPW23, LDX23, LZ18c, RZZ25, XH25, Zha19, ZYJY22]. characteristics [BEM99, HHST19, Hu22, Si12]. Characterization [Bar91, KAL22, KGD03, CG09]. characterizations [SSC25]. charge [LNS23, XY25]. charge-conservation [XY25]. charged [ZLQT19]. Charlier [HSL19]. Chebfun [AMR23]. Chebyshev [ABI23, Ano05b, AM16, AS14, Boy05, CCTV16, CKKT16, DG05, DDRS23, Dri93, Dze15, Fun01, GGNF17, Gem97, GS14, GI15, Gla01, Gu20, Gug96, IT93, JVH15, KD14, KSCS08, KL06, KRS19, KGMH21, KP03, LMM97, LL05, LSW16, ME95, MC05b, MC05c, Mas05, Maz09a, Meu09a, Mok16, Not95, OPSM22, Pet95, PG05, REM21, RS97, Rob98, Sar06, SW14, SS12a, SH17, SSP15, SVZ05, Som05, Spa24, Tas93, Thi93, WK13, WK14, WK15, WLJ24, WRM17, WK93, ZJZ20, Zha23, dR99, dAR06, dC23]. Chebyshevian [BCM19, Maz05b, Maz09b, Maz11a, Maz11b, Maz12, Maz18]. Chelyshkov [NRV23]. Chemical [KV04, BWB24, Hol98, MB09, San14]. Chen [GNH10]. Chipot [KKPT16]. Choice [Alt21, AL15, BKA19, BVV14, BL95, BGR25, CHK14, HRY16, HRY19, LKK21, RP26, RR13, SR25, Van92, YHS18, ZL24, ZL26b]. choices [HM22a, HJ21]. Cholesky [BKS13, Bre06a, DDRT97, ST26, TT06]. choose [FEL15]. Choosing [CMRS01, CKL16, FZ07]. Christoffel [Mas95]. Ciarlet [LS20]. CIR [YW17]. circle [BCM07, BPR20, BGVHN92b, BGVHN92d, BGVHN96a, BC09, BSL18, BD25, CGM12, CMR16, DHS09, GHM08, JK18, MP08, MA95, VM17]. Circles [Mar04b, DHS09]. circuit [CDS20, Pen98]. Circulant [CO94, DNR17, FNS19, HW18, Lin01, Lu15, RW24, vdMRS06]. circulant-block [RW24]. circular [AS08, ALZ20, HM06, Kar13, Kar15]. circular/spherical/elliptical [ALZ20]. circumcentered [BOW21, BBCS21]. circumcentered-reflection [BBCS21]. Circumcentering [BCS18]. Civil [CCK04]. Clarke [SR06]. Class [All03, ABQ04, DL03, GPP01b, NS01, Rec01, AD22, ALY22, AE09, AH15, AAH18, ATT22, Arg10, Bai97a, BH22, BHT16, CCJ10, CP95a, CZ20, CRN19, CKKT16, CRHTV24, Cro92, DS20, DZH23, Eba18, EY10, ES19, Fan22, FT24, FAM25a, FAM25b, Gar20, GS95, HH12, HT19, HSE16, Hie19, HL23b, HZ15, HM18a, HM19a, Hua20, IY15, JWCZ21, JSZ22, JML25, KMH24, KS18a, KNBGV18, LYW14, LY17, LM19, LLY22, LZ16a, LA22a, LT20, LHL11, LPP21, LSSS15, LZ22d, MN23a, Maj14, MM08a, MVV25, MM12, MS01a, Maz11b, Men26, MG21, MN08, MGG25, NLT21, NEMS14, RS20, RP26, RA12, RB17, SR25, SCDM20, SS24b, SH21a, Spa20, TTXZ23, USAF14, VSA12, WG13, WK15, Wan15b, WYP23, WL24, WC25, Wat92, Wat94, WZS14, WL17, XLW20, XZP^{+}20, YWX14, YBK^{+}21, YDW26, Zah09, ZYW21, ZWXX24, ZSWW25]. class [ZH23, ZZZ22, Zhu21]. classes [Gau10, IAAH24, KCBT21]. classic [Eft15a, Eft15b]. Classical [FPP05, LW04, dB07b, BD17b, FLV14, GAM24, LGA^{+}00, LM12, MM08a, Wal94, Yak95, dC23]. Classification [Bre03a, CG09]. Clenshaw [MC05c, SVZ08, SH21b, SH22]. Clifford [Rob92]. close [Iid24]. Closed [CS08, AAAA^{+}18, Ber93, YHR26, Zas22, ZLLH22]. Closed-form [CS08]. closest [BE17, MHR23, YHR26]. cloud [XW17]. clustered [NP22]. Clustering [KPFG04, ADL05, And18b, RALJ26, ZCC26]. CMRH [AEH20, Dum13, DHMS16b, Hey01, Sad99]. co [Erb15, Mar92, ZJSS25]. co-coercive [ZJSS25]. co-dilated [Erb15]. co-recursive [Mar92]. coagulation [Str05]. coalescence [OKP21]. Coalescing [DPP19, DPP22]. Coarse [Yse99, Cal20, LC21, Spi25]. coarsening [FS21, Not22]. Cocoercivity [MG22]. coconvex [NCC11]. cocycle [KS97]. Code [AKKW03, DV01, BGI25, BCJ99, DMC20, Moo07]. Codes [CMR03, CMR93]. codimension [GH23b]. Coefficient [ABQ04, BCW13, BQ19, BK13, Bou24, DP16, GD15a, HFDSC24, KS06, Lis25, MD15, WZZQ24]. Coefficients [CR03b, LW04, SVZ01, WDY04, AK12, ASVC21a, ASVC21b, BK18b, BHS17, BEH24, CLQ26, DHMS16a, DL09, FYYW19, FHV15, Gau09d, Hag25, HP18a, Ibr25, Ixa19, Ixa21, Jat15, KL22, KP25, LGA^{+}00, Lie00, LZQ26, Lóc18, MdR08, MdR13, Moo20, MAFN16, Pet01, PP92, Riz18, SL18, SS14, Ter22, TPLB22, TTP26, VC10, Wan15b, Wim99, Wri95, YYK24, ZW14]. coercive [GM20, ZJSS25]. cognitive [XW17]. coherent [HWD25, MPC12]. cohesiveness [But15]. Coiflet [AK12]. Coiflets [YXL18]. coiled [BQ19]. collection [HM22c]. colliding [KN21a, KN21b]. collinear [Ari98]. Collocation [AKW02, AKKW03, BK04, DM03, Fun01, HSZ03, ZXF14, AB06, ATM19, ASS13, ABI23, BBC25, BB14a, Bhr16, Bia94, BK13, BF17, BFKM20, BFK22, BCI14, CKP19, CS99, CL24a, CCD10, DFJP10, Fab16, FW25, FHS12, FH15, Ghe13, Ghe16, Gu20, GK21, Han22, HM22a, HM22b, Hei06, KPD25, KP96b, Kar13, KR11, LM21, ML20, MH21, MH23, MD21b, Mic23, ON26, OM26, Sal17, SCTP00, SCDM20, SPS26, WZ19, WLJ24, ZH25]. color [CKM^{+}25, DMZ20, HS26, LSL25, SD26]. coloured [ZYW17]. Column [AK26b, NK21, EHN17a, EHN17b]. column-action [EHN17a, EHN17b]. Column-oriented [NK21]. columnar [YSLH19]. columns [MP92]. Combination [You16, BCJ24, HHHN07, WHD22]. combinations [FHH05]. combinatorial [Wim00]. combined [FT05b, HPS13, LDL^{+}19, MClXzJe16, SKK25, YZ23, ZYGQ17, Zha19]. combining [BF93, LZ18a, NB16]. Comment [AAFL23, Don16, ZDSY20]. Comments [And19a, CB16, DW15a, DW15b, PK25b, PK25a]. common [Den14b, FGM19, LCL21, PKC19, PKC18, PPPN23, RZ16, RT20, RT22, RTTH22, SDMMK18, TAM21, Tak17, THT19, ZH19]. communicability [EHNR23]. Communication [NR25a]. commutative [HLW26, Sal94, WLS25, YBK^{+}21]. commutator [CO19]. commuting [MZ99]. Comonotone [NCC11]. Compact [Che24, MA13, SS26, BKF20, CZZ^{+}25, CMWP20, Cui13, DMA09, DZ13, DW22, KM24a, KMV17, LlLZ21, LKQ23, LCW23, LZW20, LNS23, LZQ26, LJbL21, MClXzJe16, MDL15, QQX23, RSCH^{+}19, SMM25, SzS21, TV25, Wan15b, Wan19, WLG26, YJX15, ZP17, ZQzS22, ZLM25, ZM26, ZZB20]. companion [FDV13]. Comparative [CN17, SS12b]. Comparison [AFN16, AFN17, HR03b, Sab91, SU14, ANA14, BBBC20, CS12, FLV14, HMS96, MP00, MFPG07, PZ20, WZC23]. Comparisons [Zha23]. compensated [YCW^{+}19]. compensation [LLM25]. competitive [KA26, MAS17]. Compiler [ASS04]. complement [BBL23, Cao12, ST18, TA13, Zaf22]. complementarity [AJMP11, AM18, AM13, BD06, BZ13, CW19a, CLWV15, CHK14, DLL12b, DLL12a, DLL13, Don10, Fan22, GWL18, GEP14, GEP16, GEP19, GLG26, GZL^{+}25, HT16, HLV26, HLZ14, IJSS16, JSZ22, JRRS08, KMZ18, Ke21, Khe14, Khe16, LLS11, LZ14, LL16, LY17, LDL17, LL20a, LYH^{+}20, LZOY22, LLY22, LZL22, LZL23, LHL11, LZX22, LWZ21, MG20, MG21, Mez22, Nie93, RWTW19, SWS22, SW24a, SG17, SG18, SW26, VPA24, VPA25, WCW20, WL25, WGZ25, WCZ25, WPL18, WK20, WL22b, XZP^{+}20, YSXY19, ZWL26, Zha11, Zha15, ZZY18, ZY13b, ZLV17, ZV19, ZV21, ZZLV23, ZLLC11]. Complementary [CDD13, EL08, Zag24a, Zag26]. Complete [Caç10, HYT26, KM13, LP18, NV21, Sut09]. Completely [MP22, Xu19]. completion [Dax17, DC17, GS19a, GWW15, LHZ^{+}21, LQ26, NH24, WC23].

Complex [FKMS01, FH04, GST02, GJVW24, IMT23, JM93, Mat04, MKO04, Neh04, SS01a, ZZ17, AMH10, AMA21, ANA14, BBC11, BOP98, Bel99, BKS13, BPV13, CL11, Car95, CC16a, CW19b, DG17, DS20, DM92, FHH05, FLV14, FM93, For21, For22, FDFM23, GH25, Har20, HSE16, HC24, IS17, IT93, JLMP16, Joh20, LYW14, LM19, LA22a, LZ19b, MP08, MBM25, Mor25, NMM25, NR25b, PR93, Reb97, Ree92, Riz18, Sas93, SS01b, SD25, SLZ25, SSH20, WGZ18a, WZVJ22, WL17, XW18, XLW20, YWX14, ZA24, Zag24b, ZD17, ZFG18, ZYW21, ZJJW24, ZWXX24, ZM16, ZZZ22, ZCG15, ZMZ26]. Complex-extrapolated [ZZ17]. complex-symmetric [WL17]. Complex-Valued [Mat04]. complexes [FS20]. Complexity [AG15, CWZ13, Cor02, PH14, AG19, Ber00, GS21, KD14]. complicated [JYLC21, Yse99]. component [BDL^{+}12, CHS19, HNY^{+}18, JKC26, PGGCGF11, Ter23]. component-by-component [BDL^{+}12]. component-wise [CHS19, JKC26]. components [HJ25, LNS23]. Componentwise [AMKV96, Rum14]. Composite [BE03, GS14, AABTB23, BC94, CS94, HL15b, LB25, LX24, RW06, Wan24, ZZZ20]. Composition [McL02, BC05b, BC01a, CHMT10b]. Compound [KM09, HKKN12, MK94, MK97]. Comprehensive [ZCC26]. compressed [CZX26, HLL22, LWLW24, LLL26, MJJ^{+}23, SIO20]. compressibility [JY23]. Compressible [Jan03, CL00, GB21, HMS96, Hu22, KLB10, LRC19, LCW20, Son93, YZ23, ZYGQ17, Zha19]. Compression [BKW25, DV01, Som05, ART14, ALRT16, DF93, DBV23, FT05b, LHR20]. compressive [HIS26]. Computable [CMRS00a]. Computation [CGL01, CJTW96, CCK04, DVJBN03, DMRT03, DI11, DGST15, EDAH12, Esp05, EAGS20, FH04, FLH04, GH25, GRAST23, HR03b, INR01, JC04, LS03a, Lor95, MZ99, Sab03, SER02, YA24, AM24, ALW98, AM98a, ASVC21a, ASVC21b, Bec96, BBL22a, Bel94, BBP17, Bin96, BMV09, BZ91, BGVHN92a, BC09, BGZ20, CR00, CC07, Car95, CBGVPP09, DN24a, DHS97, DG94, Eba18, EGT25, EGG08, EL01, FT24, FRJT09, FLV14, FC01, Gau11a, GH09b, GST21, HM26, HPS97, HMS96, Ixa21, Joh15, Kar00, KMRAR25, KV00, KHM20, LZ12, Luc06, MdR08, Meu97, Meu12, Mil17, MK94, MK97, NZ19, OR17, POP17, Sas93, SS23b, Sto93, TBY13, TM10, Uhl09, VB91, WLS25, YZ17, Zag24a, Zag26, dC22, vdMRS06]. Computational [BD95, Cor02, Cuy00, DH04, LMD25, Ste20, WLL12, ZAGD22, dSCS04, ABG97, ALZ21, BN25, BBC25, BHS23, BZ18, DSI11, DP21, DEM94, ES19, Jai16, Jai17, JCF15, KPD25, Kha25, MWDD24, Mil13, MMLM20, Mon01, MWWY13, NM26, SGO22, WCB15, Zas22]. computationally [Liu11, Yan18]. Computations [CR03a, Gla04, ARY17, BESC22, CHYZ98, CM98, CFL19, FT05a, GS14, KFK^{+}24, KM09, MT98, MPR22, MdR13, Tsi07, YHZ20]. Compute [WH04, AT19, GJV17, HYW20, HN16, Jia06, VC10, Van07]. computed [ICR06]. Computer [ARJ03, Reb97]. computerized [NAE22]. Computers [IM02, PH20, Con93]. Computing [AA12b, BM19, BBPV12, BM06, BIM^{+}23, BS04, CKO24, DY93, FS20, FHV15, GST02, GST03, GKS04, Gra03, HS03, Joh20, KP96a, LMV23, LNS23, LER03, MT06, Sch08, Sut09, TA96, Wal06, AG25, AT12, AHL20, BCCP25, BH92b, BK16b, BSB23, CDT10, CE17, CHHL18, Che94, Eft15a, Eft15b, EG19, ED22, EEM20, FGM19, FDV13, GLM15, GI10, GLC22, HHST19, HKP25, HS16, HIK17, HJ21, HJ25, IT93, Iva17, Jia20a, JXX^{+}23, KGG25, LMV24, LMV25, LLZ18, LZ22a, MGL20, Maj18, MVV05a, MVV25, MSS18, MEJS19, MS23b, Meu99, MT13, MT14, Mül00, NIN12, REE26, Saf10, SL18, SS12a, SS12b, Sol15, TLD22, Vep08, WHS23, ZJJW24]. Concave [KPFG04, JWCZ21]. Concepts [dFS04, Cha14]. Concerning [CJ04, dB07a]. concerns [SPV20]. concurrent [BBP17]. condensed [Bos21]. Condition [MZW20, TDKB24, AK26a, And10, ACH19, Bag16, CWL16, CL10b, DL21, Dea15, DS09b, DBV23, DWX17, DP16, DMW23, FGP91, Gon16, GO21, Ish26, KP96a, KKS22, KN21a, KN21b, LLC20, LJ22, LF25, MKA14, MV17, Ovi22, PJ22, Pie96, SA23, WK12, WSK14, WK14, WK16, WK17, WDL23, XWX24, XP23, YYZ22, Zha20, ZBX21, ZW22, ZP23, ZPEL25, ZLWQ09, ZCS14]. conditional [GO20]. Conditionally [Str02]. Conditioned [RST03, Ant22, BJNKR20, DHS97, Ria16]. Conditioning [CS18, DP01, MT04, Ant18]. Conditions [CG03, ED05, KPT03, Mat01, Ahu09, AABM17, AH09, AH11b, AKQ17, BKF20, BSF17, CCN25, CCC25, CK22, DDP14, FFSC25, GZ20, Ghe18, GLV05, HM26, HVM15, Jay21, JP14, JTW25a, JTW25b, Kar15, KCHD16, Kol06, LLY26, LAG05, LG08, LWY26, LZ09, LLL22, Lin05, MS22, MFK^{+}15, MSZ20, MKS18, TBY13, WK15, WZ19, YZY^{+}14, ZYW17, ZS22, ZS13]. conduction [BM24b, GD15b, HYT26, MP25, MS14b, VMMD21, WXQ20, ZLH21]. conductive [dOS07]. conductivity [Bou24]. cone [GNT24, KMZ18, LGZ25, SMZMA18, YYL15]. cones [LZOY22, YZLP16]. conference [BV96]. confined [SFZ22]. confirmation [ZQL^{+}19]. confluent [AGS08, GRAST23, POP17]. Conformable [WX24, AVCO26, CCTV23]. conformal [PTW22, Tru24]. conforming [KP96b]. Congress [Ano92]. Conic [FM04, TO21]. conical [DGST15]. conjecture [KM13, Sal25]. Conjectured [GL07, Gau08b, Gau09c, Gau09b, Gau11b, Kou07]. Conjectures [CD00, LR14, MS22, dB07a]. conjugacy [CL10b, LK20]. Conjugate [Ben99b, DLL04, WZ16, WSY04, AF23, AK19, ABK22, ABKD23, And08, And10, And14b, And15, And18b, AK00, BKFMA11, BKG15, BC94, BGS24a, BF14, BE20, BE98, BF99b, BK08, CZL26, CMRS00a, CS94, CL10b, Cui26, DW12, DW15a, DW15b, DMA19, Don16, FHH96, GS14, Haj16a, HM18b, HC24, JLJ22, JM93, Jón93, KBA26, LM97, LLL18, LZ22c, Liu11, LWQR15, LSY^{+}23, LJY24, LTP18, MJJ^{+}23, Meu97, Meu99, Meu05, MT13, MT14, Meu20, Meu23, NH24, NKA24, NA24, OAR22, PPPN23, Pla99, SKK25, SLL22, SLL23, TTXZ23, UPS25, WYP23, WHS20, YCL17, YHS18, YJJ^{+}21, YWS20, Zha09, ZW12b, Zha20, ZWG18, ZLL21b, ZDSY20]. conjugate-gradient [FHH96]. conjugate-gradient-type [OAR22]. connected [DM98a, ZLH22]. Connecting [SER02, DFK97, LZ12].

Connection [MdR08, MdR13, Wim99, CLaL00]. connections [MKG24, Tov97, Tov98]. connectivity [CL96b, CHS25]. conquer [KKM20, MVV05a]. consensus [Wan24]. conservation [AC17, JB24, MCDS24, Sar06, XY25]. Conservative [CG20, CJKL23, HZPW23, HVMT17, LRC19, LHZ20a, LKQ23, LDX23, LHlLG24, WH15]. conserved [LZM23, YL22]. conserving [ABI22, BS14, LZW20, Sat24a, Sat24b]. consistency [Kub23]. Consistent [HL03, Sch09, AM98a, EL01, HK14, Hom94, Hom98a, SL18]. Constancy [TD09]. Constant [AAAS03, AH11c, BEH24, Che16a, Dra00, HFDSC24, Kar00, Odl00, Sîn07, TRCL24, Yak95]. constants [CYIB12]. Constrained [Ano95c, CT93, CPN14, Dos03a, Góm01, LKW17, ME95, SC03b, AWH^{+}24, ABM10, AFN16, BGRS09, BGS24b, BMR21, BP93, BN18, CB16, CW21a, Car91, CB00, DEM94, Fab16, GO20, GLW16, GWL20, Gug96, HHF22, KLT95, Kha25, LWZ18, LZL20, LWLW24, LSY^{+}23, LJY24, LPXX19, MBG19, MJJ^{+}23, MAS17, MS20b, MN22, NK21, OMW21, Orb15, OL23, OKB23, ODL21, PV23, Pea13, PSZ23, Rei98, Rog95, SL18, SEG14, Sei98, Sla06, SG10, mTLbJlL14, Wan18, WCH15, WZS14, WZZ15, XP23, YIY22, YJJ^{+}21, ZLWZ21, ZZ22a, ZLZ23, ZXL25, ZFH26, AFN17]. constraining [CPP14]. constraint [BPR21, CL96b, DHMS16a, HBP13, LHZ10, LCH20, LJ22, LHlLG24, MAF25, PL25, Pen13, PN21, WXT22, ZAH26, ZRZ11, ZLCW23]. Constraints [Str02, AAFL23, CLGS17, FGBP21, Fra26, Gal93, HNY^{+}18, JR20, JYLC21, Jos22, LWK12, LWM10, LZ25, LF19, LZT25, NPS09, PPPN23, WZ23b, Yan22, Zha95, ZW20, ZZX^{+}23]. construct [Hal14]. constructed [BD20]. Constructing [FVV25, SBW98, YZ17, ZYBJ23]. Construction [BJ98, Bel94, BHS17, BJ04, CHH^{+}20, DNV25, KPR03, KPT03, LM97, NS01, TS18b, ZWL26, ANI15, ANI^{+}17, AUD18, BCL00, BDL^{+}12, Bou17, DR12, God15, KP22, Laz99, LSG15, MS01a, Mic91, MRU91, Pas95, Tur94, WLG26, Wri01]. constructions [KSB08]. Constructive [MS96, SCTP00, XLC93]. contact [MZ19, ZZZ20]. Contents [Ano04c]. contiguous [Seg08]. Continuation [CCJC18, lLXhL22, lLXZ24, BZ18, CR99, CCJ99, DC17, DFK97, HV98, Lin16, PR25, YYD14]. Continued [HLM04, Pas03, Van03, Arn97, Lor95, LM15, LSM16, Mor11, Now06, Now13, Pas92, WQ10, dB07b]. continuity [WK12, WK17, WXT22, YYZ22]. Continuous [CZ20, DJ10, JZ16, ZXRL11, ZLL^{+}21a, AM24, ABB15a, ABB15b, AV19, BRW11, BX19, CG07b, EH97a, GGN18, GWY26, Hag25, HY21, JT96, LG19, LT20, Maz09b, SI18, SSL24, Sun94b, YC22, ZY21a, ZSLZ24]. continuously [LGW14, TPLB22]. contour [IDS16, LG08]. contours [GLV05]. contraction [CL96a, Dey23, DLYH17, TVC20, VA20, WZC26]. contractions [PK25b, PK25a, PGC24]. contractive [SRK^{+}25]. contractive-type [SRK^{+}25]. Contractivity [Cro03, KNBGV18]. contrastive [ZCC26]. contribution [XXW17]. Control [BD03, BD04a, Cha04, Van12, HÇK26, MV13, Söd02, AJ13, ATM19, ART14, AMR15, ASHF21, BKPS93, Bla15, Bör26, CB16, CW21a, CZH22, CM96, CGV26, DOR26, DTI09, DI11, EDAH12, EH97a, EM10, Fab16, FWC16, FJT94, Hof05, Kha25, KS06, KA26, Kun05, KLR07, LWZ18, LZL20, LAN18, LCH20, LW17, LP20, MV14, MB06, MKA26, NPS09, NEMS14, NAA19, OMW21, Pot19, PSW11, Ria16, SS24a, SKM25, Sla06, SR24b, Ste24, TQY21, VH10, VH12, Van19, VS19, WZ22a, WZS14, WZ23b, YF22, ZM94, ZZ24, ZZZ20, ZZ23, ZZ18, ZLCW23, Sla06]. controlled [AM24]. controller [DHS97]. Controlling [NK16]. controls [PLH20]. Convection [Bog02, BPR22, BS17, BEM99, COSE22, Che19, CEK21, CGL99, CJ17, FMD23, GLLJ12, GH09c, GO06, GO21, GS26, JTW25a, JTW25b, Kno23, LL22b, LZ23c, LZQ26, MS20a, Ngo23, NV21, PM24, RBN14, RT19, RZZ25, SKA23, SK25, Wan15b, Wan19, YM24, YJ18, ZZH15, ZLH21]. convection-diffusion [BPR22, Che19, CJ17, FMD23, GH09c, GO06, LL22b, LZ23c, NV21, RBN14, RT19, RZZ25, SK25, Wan19, YM24]. convection-diffusion-reaction [BS17, COSE22, PM24]. convection-diffusion-wave [Wan15b]. convection-dominated [MS20a]. convective [ASGGRG23]. Convergence [AUA22, AHVR17, BD10, BRZ18, BCN^{+}16, BRW11, BE03, CCN25, CC12, CWHL20, CJ04, Cui13, Dos03a, DL09, Dur93, EHN17a, Fan19, Fan22, Gal22, GR01, Hag25, HKE97, Hie18, HC03, JSZ22, Ke21, KR07, KGMH21, KP03, LGZ25, LWZ21, Ma20, Man21, Mat92, MM09b, NUG25, Now19, PG12, Pas92, Pas03, Pas08, PLH20, Pop18a, RS02, SP21, SHLY18, Sid94, SDMMK18, Śmi09, SBJC19, SX00, THS20, Val15, VS19, WSK14, WK14, WK15, WC13, XZZ22, YLL20, ZY21a, ZSLZ24, Zil01, dDL92, AJ13, AL15, AC17, AABM17, AU08, AH09, AR09, AH10, AM16, AG17, AM18, Ari98, AC19, AK00, BBBC23, BOW21, Bic24, Bog13, BGVHN96a, CSI18, CK24, Ceg25, CZ96, CHHL18, CC15, CDG23, CL13a, CWL16, CW25, CJ17, CHMT10a, CKKT16, Cro92, Cui26, CKS16, DW12, DEC24]. convergence [DD21, Das19, Deh20, FLMR99, FT14, FS23, GS14, GO20, GM96, Guo13, HLL22, HVM15, HVMT17, HKKN12, Hom92, Hom98a, Hua94, HZ15, HM14, Iva17, Jai16, Jai17, JLP20, JZ25, JWY21, KPC20, Kno23, KL06, KLL10, Kza97, Le 92, LKQ23, LWLW24, LLY25, LZ15, LZ18b, LL26b, Lin16, LLL22, LT20, Lin05, LQ20, LCHH21, LZ23c, LZCW24, LSSS15, MA16, MM17, Mat96, Meu20, MG11b, MS13, MGG25, NEMS14, NAA19, Not22, NT21, Now06, Now13, Osa12, OL23, Ovi22, PT19, PKC18, PK21, Pas11, PHI98, PP06, PR14, PSS22, RTD^{+}21, RWB09, RA12, RK11, Sab91, Sad05, Sch25, SS10, SA14, SI17, SSL24, SS98, Sid20a, Sid20b, SK19, Sv95, SHGL22, SPS26, Stu97, SLL22, SYZ22, SSH19b, SSH20, TT21, TA26, TH18b, TVC20, TTLD20, TSI20, TDC21, TLD^{+}23]. convergence [USM24, Vep08, WGK11, WKG11, WK12, WZ13b, WK13, WZZ16, WK16, Wan17, WZQ17, WK17, WCHK21, WZ26, Wei17, XCD23, XWL25, XSL22, Yak95, YWX14, YL19, YH24, YCZ26, YWS20, YLL22, ZW12b, ZJ14, ZD21, ZXL25, ZWG18, ZG12a, ZG12b, ZL26a, ZG09, ZCS14, ZW15, ZAC26, dB07b, EHN17b, Pop19]. convergence-control [AJ13]. convergence-guaranteed [USM24]. Convergent [BKM03, CCG01, FMD18, Rec01, Rob97, AH08, CEX14, CL93, CG19, CJ20, GA15, Gal18, GLW13, Hai08, Har18, KSB08, KK17, MFBB23, MA22, NKA24, PLZ^{+}24, Pie96, Sab92a, Sid07, WB17, wYN18]. converges [KH18]. convert [KCHD16]. Convex [DM92, EGSV04, KPFG04, KL04, Mar04c, Zil01, AF23, AAFL23, AG23a, AN17, AHL20, AW23, AGG17, BBBC23, BCK06, BBCS21, CWZ13, CF96, CKS24, CD25b, El 18, GPGC98, GWW15, GH95, GWL20, HHF22, Hua96, IUM^{+}19, JWCZ21, KJO23, Kri26, LP18, LF19, LZT25, LCW21, MBG19, MWZL23, MS96, OAR22, PKC18, SPV20, SW24b, Tan20, TOY26, WZ11, WS24a, WZ22b, Yan22, YZL25, ZAH26, ZJR25, Zas22, ZXL25, Zha95, ZZX^{+}23, ZFH26]. convex-concave [JWCZ21]. convexification [PW22]. Convexity [LN95, AT12, BBBC20, BV21, GQ09]. Convolution [Lev95, AA15, Ave20, BHH24, DL18, LLY25, Lig93, Mon96, MSS11, WWL24]. convolutions [JLP20]. Cooperative [BBP17, KW04]. coordinate [BGRS09, BGRS12, LB25, RALJ26]. coordinates [ÁCL11, BW15, JKK^{+}08, dOS07]. copolymer [LZM23]. cordial [KMH24]. corner [MP99]. corners [MD15]. correct [EM10, WLL12]. corrected [AAIT94]. corrected-asynchronous [AAIT94]. Correction [AHS22a, AHKW04, Cat24a, CUY24a, DH18, EDAM13, FAM25a, Gau19, HHLS21b, IZ24, Jan03, JTW25a, KST21a, KN21a, MW26, Meu26, OCWV24a, PK22a, Pop19, PK25b, Sat24a, SGS26, Uhl22a, Zag26, ZS03, Zeg25a, ZM26, ZL26b, AHKW05, CH26, FLR01, GS16b, GO06, Gug96, HCL21, HK06, MBG19, Meu23, SGS25, Si12, VdR13, ZS22]. corrections [GPHAPR24, Iva24]. Corrector [IW04, Bra06, DZW17, Khe14, MJH17, NA26, SMZMA18, SD20, WO00, YZLP16, ZLS24]. correlation [HS16, Zhu15]. corresponding [LWG18, LCHH21, SLZ25, WLW^{+}24]. corrupted [LP12, ZWZ25]. Cosine [HS03, Col92, DPR23, DKL15, HH05, LZlL20, Mac96, Saf10]. cosparse [LLLD17]. cost [EY10, FLMR00, Rab23, ZFH23]. Coulombic [EAGS20]. countable [CAV23]. countably [Ste24]. Coupled [KV00, BDV18, BGS24b, CZL26, COSE22, CGSS24, CDS20, DW21, DW22, DCW23, DN19, Don13, DZ21, FZLL23, Haj16a, HR05, Hol98, HKCW24, HM18b, JKC26, LHZ20a, Lin09, LCW20, SZ23, SSA25, TPLB22, TLD22, WSZ21, WDL23, XH20, YSLL23, ZD15, ZDZ26, ZD18]. Coupling [GX19, BBB22, BDF^{+}25, CK22, CGM25, KK22a, Zaf22]. coupon [CS12]. CQ [KPC20, SCD^{+}21, Wan17]. CQSCO [Khe17]. Crank [LS25a, LBF24, LWJ21, LW22, MS24a, QQX23, TL24, Wan19]. criss [DLR12]. criss-cross [DLR12]. Criteria [Cse04, KGD03, CL19, KS25, LLQ17, RSZ20, XZZ19a]. Criterion [CMRS01, PT19, SDW24]. Cross [CDG23, BC92, BL93, BL95, BR21, DLR12, Jia06, MRV23, SWB08]. Cross-points [CDG23]. cross-product [Jia06]. crosses [FKP06]. crosslet [GSZ22]. Crude [WH04]. crystal [LL20c, LZZ23, RCW22, YK22, YZL25, Zha26, ZWY22]. CS [KLSS14, Sut09]. CT [MN23b]. Cubature [PS01, AB24, BD95, BC00, CDSV11, DFD23, SVZ08]. Cube [Coo03]. Cubic [CPV04, CMRS01, AB06, AGN07, BK13, BV21, CG07a, CW14, DFR26, DHS09, Gha16, GGS22, Kva14, Liu21, PSZ23, PZL15, PTW22, RA12, WZ19]. cubically [AH08, KSB08, wYN18]. Cubics [LM01, BFK^{+}09]. cuboid [WZZ26]. Cummings [Kar07]. cumulative [GST17, KSV23]. CUR [WKY25]. curl [Cal20]. Current [Tsu02]. Curtis [SVZ08, SH21b, SH22]. Curtis-type [SH22]. Curvature [CRS04, CEK21, MAF25]. Curve [CRS04, Flo03, BM06, FJT94]. curved [ZZ23]. CurveLP [Yan17]. Curves [CP01b, KPR03, Neh04, BG91a, Bar13, BV99, BMV09, DF94, FHH05, Gar19, GS16a, GLW16, HSSB13, LN95, LWK12, YYD14]. curvilinear [BBBC23, KZ21, dOS07]. Cut [BKM25, fLxX12, LCW21, SS24d]. Cut-edge [BKM25]. cuts [AL23]. cutters [CRZ26]. cutting [MP99, PR93, Ree92]. CVaR [PL25]. Cycle [CEZ26]. Cycle-free [CEZ26]. cycles [CZLS18]. Cyclic [BGRS12, Bia94, TPY14, BGRS09, BM97, BM09, Ern00, JJ13, JL16, KH18, LKKM15, RT22, RZ23b, Zas22]. cycling [RVF07]. cylinder [Seg98, SCF23]. cylinders [Wat06].

d [CG19, ATM19, AEF^{+}14, CL00, CJ17, DL21, FHC21, GH09c, LG08, LZlL20, MFK^{+}15, Moh10, MK17, PMO05, PT18, Sal17, TC05, VMMD21, ZJ08, ZYX19, ZYLN18, ZS19]. D-BEM [AEF^{+}14]. DAE [Ano98d, CB16, MW98]. DAEs [AAP24, AB98, Bea98, BCT15, CHYZ98, CM98, Fab16, HCH18, JV98, LM11, Pen98, Pry98, SL15b, SL15a, SL16, SL18, SL21a, Jay02, MT98, Sch09, Sim98]. Dahlquist [CKM19]. Dai [AK19, ABK22, And18b, BKDA25, KBA23, Zha09]. damped [BD06, DS21]. Danckwerts [LLY26]. Darboux [BuR24]. Darcian [CW17]. Darcy [COSE22, CLPY23, CJKL23, DZ21, DZ22, DMW23, GG22, LRC19, LHR20, LHW13, RRZ21, WDL23, WQ23, XH21, YJ18, ZD18]. dark [ZTZZ19]. Data [CMRS01, KPFG04, BK18a, AUA22, AS08, ACM93, AG00, AF94, BW93, BQ19, BH11, BGS24a, Ber10, BCJ22, Bro05, BF93, CF96, CK05, CG07b, CM24b, CAV23, CP93, CP95b, CDD13, CM05, CV92, DF93, DdAF^{+}20, DC17, Dri93, Ell93, FT24, FM16, GGN14, GZ18, GLRSG08, GI97, Hof05, IL05, KBCG13, KPT23, LW20, LWC^{+}21, MFBB23, Mic93, MN22, Nar05b, NZ19, OKB23, PS25, Rip93, Ros97, Sch14, Sch08, Wat06, WD96, ZXL23]. data-bounded [Ber10, OKB23]. Daubechies [Dur93]. Davidson [HPS97, HJ25, SS99, VdR13]. Dawson [Zag24a, Zag24b, Zag26]. DB2 [ADN17]. DC [GWW15]. DCA [GWW15]. DCACO [YP23]. DCT [BP19, Ihs07]. DDEs [CE17, CMP22]. Dealing [Bag00]. Deblurring [NPP04, CPZ14, DHMS16a, FAMA20, FFSC25, LP08, LP12, LZ25, MAF25, WCH15]. decay [GL19]. decentralized [Wan24]. decision [GK24, ON26]. decision-making [ON26]. decomposability [Uhl22a, Uhl22b]. decomposable [BRSY21]. Decompose [CGV26]. Decompose-then-optimize [CGV26]. decomposing [LV15]. Decomposition [Bog02, Dah93, HPS20, KZ03, APST21, AA15, BCCP25, BDH^{+}13, BFK^{+}09, BFK11, CKM^{+}25, CZ94, CZ95, CCJ99, CV22, Cic20, CGV26, CK06, DHV22, ETY98, FLV14, GHM16, HG93, HN16, HJ21, IL05, JNS19, KLSS14, KST06, KH11, KP26, LCW23, Lor19, MS92, MHR23, RR23, Ria16, SB21, Spi25, Sut09, UTO24, VZ93, VH92, VVL25, WZVJ22, WLW^{+}24, WLG26, YHR26, ZWL26, ZZ22a, ZJJW24, ZWLZ24, ZLS25, ZZW26, ZDZ26]. decomposition-compact [LCW23]. decomposition-type [CK06]. Decompositions [Amo02, DPP22, AD00, BMA16, BKS23, DBV23, FH97, FHH99, LKKM15, PQ95]. decompositon [SLZ25]. Deconvolution [Han02, CPN14, DMT13]. Decoupled [ZZH15, CLPY23, HKCW24, LZM23, XCY21, XH21]. decoupling [SL16, SSYL20]. DECUHR [EG94]. Deep [BS25, XLL26, ZCC26]. Defect [AHKW04, AKT15, FLR01, AHKW05, EH97a, EM10, GO06, HK06, Jaw22, Si12, ZS22]. Defect-based [AKT15, Jaw22]. defect-correction [GO06, Si12, ZS22]. deferred [SGS25, SGS26]. Deficient [CW14, WSY04, HG93, ZL26a]. defined [AC94b, Mar96, Van19]. defining [BDV18, LZ16b]. Definite [ABQ04, Str02, AA16b, AL97, BMA16, CR96b, CLMM05, CS08, DPP22, tFZyZ16, Hag13, Har18, HM18a, HV12, KPS14, LG17, LDC10, Orb15, PG15, Str97, Sun94a, Sza92, Tan17, VGV06, XM16, YLD11, Zha15]. deflated [BF14, EHTSM21, JGMC25, LM99, ZTW19]. deflating [Oar94]. deflation [lLXZ24, SS99, SA26]. deflection [YXL18]. deformation [LM14]. Degenerate [HvD93, LZZ24, PN93, RKM25, SKJ^{+}18]. Degree [CG03, FS20, FJ96, GLW16, HHLS21a, HHLS21b, HvD93, KLZV95, LL14, Poc14, Wal06]. degrees [BR17, RRZ21]. Delaunay [AT12]. Delay [Car01, HCH18, SER02, WC25, ZFX14, ZXLF15, ABB15a, ABB15b, AE09, AA09, BPR22, BB14a, CCN25, Das19, EH97a, EA12, ES19, GMY18, GHZ25, HKE97, KGH14, KK17, KKS24, LT20, MD21b, MMLM20, NS24, PLZ^{+}24, PM26, ROB17, RS20, RF23, SAC18, SCDM20, SHGL22, SPS26, WZL26, XyJl16, YY24, YZX25, ZY21a, ZE10]. Delay-dependent [HCH18, WC25, ZFX14, ZXLF15]. delay-differential [ZY21a]. delay-integro-differential [WZL26]. delayed [HDP18, PM24, RK11, YX11]. Delays [LER03, BCT15, HCH18, ZJ23]. deletion [BG91a]. Demand [Cha04]. demicontractions [Ceg24]. demicontractive [HS20, PKC19]. demonstrations [MdR13]. dengue [LWLT19]. Dennis [MS22]. Denoising [BP03, HR14, ZY13a]. denominator [KC23, KLZV95]. denominators [CJTW96, Gug96]. Dense [BG03b, ABG97, CDW95, Dum13, DHMS16b, Gar19, HL17, SAE19, ST18]. densifiable [Gar24]. density [LD20, Low05, She00, SS11b]. dependence [AUA22, Bag00, Pop21]. dependency [CCLi16]. dependent [ATM19, BS17, Bou24, COSE22, DZS21, EG18, FYYW19, Ghe18, HV22, HCH18, HY21, HS12, JWZ23, KGH14, Lam09, LSZW19, Lor19, LC25, MZ19, MDL15, NM14, NPS09, Rip93, SSYL20, SFS23, Si12, SG23, TBY13, TV25, TTP26, WS24a, WC25, WZ26, YJ18, YK22, ZSF18, ZFX14, ZXLF15, ZWG18]. depending [DPP19, DPP22, HDL23, SR16]. depletion [NAR05a]. derandomization [Gor18]. Derivation [ALZ21, TÖ17, ST21]. Derivative [CHS19, CGHH21, CMRS01, FT14, AH21, AHJ22, AAFL23, AAM24, AAB13, AMH10, AH08, AM16, AG23b, BRW11, BH05, CG07b, CT10, CQLY15, CC18, CKKT16, DNH^{+}25, FYM14, GRAST23, IAAH24, JL12, Jat15, KKB16, KMH25, KS14, LZ21, LGP11, LF19, MAH22, MKS18, MHA16, NBK17, Ngo23, OI14, OL23, PS16, PGP^{+}25, RAH11b, SA14, SHD25, SGJ15, SKM25, SWG20, TCW14, TÖ17, WZQT15, WZ16, WK16, WGZ25, WB17, YZBJ21, YYL26, Zha20, ZBX21, ZLLC11]. Derivative-free [CHS19, CGHH21, AAFL23, AAM24, AAB13, AH08, IAAH24, KKB16, KS14, LZ21, LF19, OL23, PS16, WZQT15, WZ16, WGZ25, WB17, YZBJ21, YYL26, Zha20, ZLLC11]. Derivatives [PW04, AR20, AMA21, AH11c, Ash16, Ash19, AP21, AK26c, BDF^{+}25, Die08, DN24b, Gaj05, GKL21, GPHHAPR18, Joh15, Kim21, MBM25, Mor25, SKTGR19, VDVJB12, VLCL16, ZLZ22]. derived [USM24]. Descartes [BHNS16]. Descent [Bre02, WHS20, AK19, AIA^{+}24, ABK22, And06, AB23, BGRS09, BGRS12, BG24, Bno21, BE20, BEL23, BN18, CTY25, CL10b, DMZ20, HSK20, KBA23, KBA26, LWQR15, LTP18, Maj13, PRK^{+}18, RALJ26, SW11, SS12b, SM10, WYP23, You16, ZW12b]. descent-based [BN18]. described [MSS11]. descriptor [BHS14, BH17]. Design [BCM19, BF00, BM14, Cou15a, Enr02, GLS^{+}18, CDW95, CGV26, GNH10, Lo97, Maz18, PR93, RW06, Son93, ZLL^{+}21a, Cou15b]. designing [YP23]. DESIRE [BD98]. destabilization [ST23]. destruction [MO10]. Detailed [DFF04, LRY18]. Detecting [CG07b, CM05, GK24, Pan20, Sza03, Mit11]. Detection [GI97, Meu23, Riz18, CKO24, CH26, MS23c, Ros97]. DeTemple’s [Che16a]. deteriorated [HWXC17, LZ18b, SCW17]. determinant [AC94b, FJ25, JL16, Sol23]. determinants [GLS26, GM92b, Jia20a, JXNX24, KV00, Wim00]. Determination [AJ13, Car01, Cox93, LM11]. determined [Kub15, McL25]. Determining [KL04, Ber00, van93]. Deterministic [PWCsL18, CH22, LGM25, PW16, Str05, Stu97]. Deuflhard [YA24]. developable [HSSB13]. developed [CN16]. Developing [ST26]. developments [BRZ19, Ise96, Nar05b]. deviating [Bic24]. Deviation [DM22]. deviations [SBG24]. DFO [Kim26]. DG [ASGGRG23, YL22]. Diagnosis [SL15b, SL15a]. Diagonal [DM21, FG03a, LEK21, And19b, BCW13, BL23, CT21, JB22, JJ13, LZ19b, MVV05a, MA22, MA95, MVG21, TA24]. Diagonal-Plus-Semiseparable [FG03a, MVV05a]. Diagonalization [Fue07, ZHY^{+}20]. diagonalize [Bel08]. diagonally [BD98, Con93, GLS26, LL16, LZX22, PFT98]. diamond [MMW20]. Dichotomy [LS03a]. Difference [BS04, AC94a, AE09, AA09, AAIT94, Bar91, BC14, Bia12, BW15, BRMG18, CFR06, CL10a, CXL16, CHH^{+}20, CLA11, CZZ^{+}25, CS12, CV15, CL93, DL21, DW21, DCW23, DK00, lDzS21, EA12, FHC21, FLH25, For22, GHC15, hGzS17, GGS22, HZPW23, HCL21, HKE97, HP18b, Hoa25, HKCW24, HTVY13, HL20, JL26, KCHD16, LWD23, LRC19, LMUZ19, LKQ23, LCW23, LZW20, LCW20, LWZ23, LZQ26, MY22, Mil13, MA13, ML10, MA12, NV21, PKR20, Pan18, QQX23, QXGZ20, RF23, RZZ25, SS26, SHD25, SS16, SMA99, Spr01, ST98, SzS21, VLCL16, Wan15b, WH15, Wan19, WZC23, WSLL24, WQZH24, Wei18, Xu19, YHQ19, YQM16, ZP17, ZBX21, iV12, LWJ21]. difference-summation [CFR06]. difference/compact [QQX23]. difference/local [Wei18]. Differences [Sab03, AR16, AR20, Arg09, EGSHVN15, HVMMV24, Müh99, SR06, Tur26, Wal94, dC20]. differencing [BKF20, SFMK23]. different [BF93, DJM08, HMS96, KJG23, KCHD16, LAH22, XH21]. differentiability [AH09, Pie96, Rim25]. differentiable [AHVR17, CD99, HVYMS23, HVMMV24, KGMH21, LGW14]. Differential [AAAS03, AM24, AKKW03, But02c, BH02, CL05, Che02, DL03, DOS03b, DM03, ED05, Fun01, HL03, IDAV09, INR01, KB02, LJW17, LER03, NW04, Pog98, Rec01, SKA23, SER02, TRRD02, Tsu02, ZYLN18, vLV02, ABB15a, ABB15b, ARS26, ÁFP07, AE09, AA09, AM98a, ABI23, AR99, AM98b, Asc97, AHC13, BJ98, BDV18, BT14, BD17b, BK13, DOT21, BEH24, Bre99b, BX19, BWB24, But10, CCN25, CA07, ÇD25a, CP00b, Che99, Che14, CQLY15, Che16b, CDLW21, Che22, CZZ^{+}25, CCC25, CCD10, CDP16, CPS12, DJ10, DFJP10, VV07a, VV07b, DBTG24, DSI11, Das19, DS26, DP21, DZW17, DW97, DB06, DGST15, Dze15, EG10, EG18, EH97a, EA12, ES19, EL01, EED19, FY13, FYM14, FHL21, FGJ00, FH15, FS01, GKRS22, GCPG99, Van17, Gug96, GMY18, GL19]. differential [GHZ25, GWY26, HJ18a, HH12, Hal14, Han22, HM22a, HM22b, HM26, HP18a, HZ20, HL20, HCXL20, IY15, ICR06, IJE15, IMT23, JT96, JVH15, JM18a, JWL20, JXZ26, Kam15, KJ18, KL17, KSCS07, KS06, KK17, KKS24, KLR07, LLAL21, Le 98, LW13, LW16, LRL19, LLY25, LGHS25, LZ16b, LTFL10, LT20, LM17b, LLC20, LA22b, LG95, LWZ21, MKG24, MH21, MH23, MN23a, MAH22, MKA14, Mal21, MWDD24, MS06, MPS20, Mil19, MG11b, MR96, NM26, ND21, NRV23, NA26, OCWV24a, OCWV24b, PT19, PM25, PM26, PS09, PS17, Prz16, ROB17, ROB18, RS20, RSKB17, RT12, REM21, RS97, RM26, SAC18, San25, SFMK23, ST17, ST23, SCDM20, SSL24, SR24a, SS16, SMN24, SHGL22, SPS26, Sti18, SZ20, SZ21, SG23, SS24d, TX19, Tem08, TS15, TY21, TN10, Tuo98, VH10]. differential [Van19, VRM23, VLCL16, WG13, WCHK21, WYZ22, Wan22, WC25, WZL26, WLL26, WZ26, WZ22b, WK93, Wri01, WCD21, WZ23b, XyJl16, XT16, Yan95, YBK^{+}21, YH21, ZY21a, ZSLZ24, ZXF14, ZFX14, ZXLF15, ZL17, ZJ23, ZZY26, ZE10, aZ19b]. Differential-algebraic [Pog98, AM98a, AM98b, Che16b, EL01, Han22, HM22a, HM22b, HM26, IY15, LT20, MS06]. Differential-geometric [IDAV09]. Differentiation [ACG20, Hof16, IM02, TRRD02, AH21, AAP24, AHP25, Cou15a, Cou15b, GZ18, Kub23, LM11, MJF09, NBJA17, SAE19, SL15b, SL15a, SS14, Tia21, TY21]. diffuse [Wan15a]. Diffusion [AG03, Bog02, KV04, Sch02, SME03, ZL25, AD17, AD22, BPR22, BS17, BKF20, Bhr16, BZV16, BV99, BBL22b, COSE22, CTS09, CLTA10, CX20, CL21, Che22, Che24, CL24a, CGSS24, CLQ26, Che19, CYM22, CGL99, CJ17, CG19, CJ20, Cui13, DWZ14, DAM16, DGM26, DMZ20, lDzS21, ETY98, FNS19, FZLL23, FZL^{+}16, FMD23, GZL24, GD15a, GM06, GK20, GH09c, GO06, GO21, GCZH24, Hag25, HT23, HZPW23, HCL21, HDP18, HV22, HA16, HTVY13, HL17, HYJ20, HG25, JKC26, JKNR13, JZF^{+}20, JWZ23, KS18a, KZS21, KCHD16, Kaz24, KR11, KK22a, Kno23, KMS05, KB20, KP25, LLY26, LR18, LSZW19, LL22b, LLM25, LCZ26, LS15b, LWJ21, lLHNS23, Lin05, Lin09, LRZ12, Lis25, LJWW21, LZ23c, LF25, LV21, LC25, Mal21, MS23a, MM09b, MS24a, Moo07, Moo20, MM11, MA12, MAFN16, Nac99, Ngo23]. diffusion [NMM18, NV21, PNW17, PQS22, PED15, PSB26, PM24, RF23, RW24, RS06, RBN14, RT19, RZZ25, RS25, RS26, SKA23, Sal17, SN24, SLA11, SCF23, SK25, Smi97, ST98, TH23, VT10, Wan15b, Wan19, WLMA21, WZC23, WSLL24, WZZQ24, WJC25, Wei17, WLY^{+}21, YM24, YYK24, YHQ19, YZL20, YPL21, YHL25, YCZ26, YWZ19, YD26, YQM16, ZWFY19, ZP17, ZYX19, ZBX21, ZL22a, ZYW23, ZZ18, ZZB20, ZYLN18, ZXL23, dFO11]. diffusion-reaction [Kno23, Ngo23]. diffusion-wave [AD17, BZV16, DWZ14, DAM16, HDP18, HTVY13, HYJ20, KP25, Sal17, SCF23, YZL20]. diffusive [CEK21, YJ18]. diffusivity [LL22b, RT19]. Digamma [Wen03]. Digital [Sar06, AK26b, FGL19, God15, Li95, PR93]. digits [ZA24]. dilated [Erb15]. Dimension [Maz11a, ASV23, CA22, GO06, HKPW19, KMS05, KA26, Sau07, SB21]. Dimensional [AEG02, BD02, Boy05, HR03b, AD17, AD22, AAU26, ALJLYJ24, ASS13, AVCO26, AAD14, BX17, Bhr16, BK13, BZS22, Bra06, Bra07, CMD19, CAB22, CLTA10, CLA11, CY19, CvPS15, CCW21, CK22, CGM25, CCD10, CJ20, Cui13, DSI11, DS09b, DW21, DHWL25, DDRT97, FMD23, hGzS17, GK20, GPHAPR24, GK24, HCL21, He16, Hoa25, HST15, HS15, HL20, HG25, JJC24, JJK97, KKPT16, KXD25, KPA20, LWD23, Li95, LRL19, LKBF17, LZW20, Lie00, lLHNS23, Lin05, LSW16, LD20, LW22, LWZ23, LWC^{+}21, MP00, MP25, Mar92, Mic23, MDH16, MS24b, Nat07, NV21, PTW22, PTSB01, PV98, Pis16, QXGZ20, RW24, RP26, RhG15, RR22, SH12, SKD26, SK25, SW07, SSA25, Ste24, Ter22, TD09, WWBM21, WHD22, YD09, YJX15, ZR17, ZBDK23, ZLW^{+}13, ZHFW21, ZZY24]. Dimensions [GGV02, VP23, AGA24, AH23, AK12, BBL22b, Cal20, GH06, GH09a, Gha18, Hem96, HKCW24, KKV22, KSW07, Moo07, Rab23, Ros97, WZ19, WLMA21, WQZH24, XH25, ZLM25, ZM26]. Diminution [CG03]. DIMSIM [BCJ99]. DIMSIMs [IJ19, Wri01, BCJ97]. Dirac [AT17, LZ23a, MY22, SS11b]. Direct [AW23, De 02, FG03a, GRT97, TBC^{+}23, Ari98, AVI97, BCT15, BW15, CB16, CFL19, CDLW21, Cic20, DG17, DB06, GK20, JRB17, KP96b, Liu11, MS20a, Meh11, MV17, SDL^{+}23, lZqL25, CIP10, MS14b]. Directed [vGK04, GWL20]. Direction [MST03, BF17, Bno21, CFR06, Cui13, Haj16a, JHLL15, KPFF24, Kim21, LWZ18, LRL19, LCW21, Ma20, RWTM21, UPS25, WLMA21, Zha15, ZZY18, ZFZ19]. direction-type [UPS25]. Directional [NS01, CFR06, Lai92, MS01a]. directions [Ben99b]. directly [BDV18, LZ16b]. director [RCW22]. directors [Zha26, ZWY22]. Dirichlet [Bia12, BFKM20, BFK22, CDG23, JP14, Kun05, LWY26, NPR08, OM18, Rab23, SJW21, ZZ23]. DIRK [FG03b, HR00, NB16]. disc [BW15, KAL22]. discontinuities [ART19, AF94, AT17, GI97]. discontinuity [CG07b]. Discontinuous [CMR03, CG20, EKM03, Moo20, AMM16, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BGR23a, BGS24b, Che19, CYM22, CJKL23, DWZ14, DDG05, DL09, FMD23, GD15a, GO21, KS26, LGC24, MS20a, Mal21, MWDD24, MM09b, MM23, MM11, MAFN16, SN22, TBY13, WZZQ24, WBY25, Wei17, Wei18, Wri95, YZ21, YZ23, ZZ18, ZWX22]. discovered [Zha24]. discrepancy [Dam08]. Discrete [ACM04, BZS22, CR02, DV01, GKQ24, Kva01, Kva14, LSX10, MN01, PL25, SDL^{+}23, Vig04, WZC23, WLS25, ZLG^{+}13, Ave20, BN18, BDF^{+}25, CMRS00b, CMM15, CW14, CZ20, CCHH23, Cou15a, Cou15b, DWZ14, DC17, DO25, DNR17, Fan19, Fue07, FYI^{+}12, GGN18, GM06, GLS^{+}18, HM22b, Han94, HWZ24, HLSLSL25, JRS09, Jia20b, JZ16, KR20, KA26, Li96, LMUZ19, LLM25, lLLVZ17, fLxX12, Loh22, LWZ21, MM22, MP22, MRS06, MRS10, NRS12, OR17, Pan18, PZ20, RR13, RSZ20, Roh07, SD26, SK25, Tas93, TYSY20, The12, USM24, WPY25, Wei17, YY24, YH97, ZXRL11, ZLL^{+}21a, ZH26]. Discrete-time [ZLG^{+}13, GLS^{+}18, SD26]. discretisation [MS24a]. Discretization [ABMV03, BBHM03, And14a, CB16, Che99, CX20, Don13, GLLJ12, GS26, HJB18, KNBGV18, KS97, LW17, LN22, Moh10, NT21, PNW17, PR25, SDL^{+}23, TY21, TCW14, WZ22a, XSL22, ZHT15]. discretizations [BM22, Che16b, DMW23, FSY23, KLL25, LD21, LYL15, MC05a, NMM18, PV22b, SW10b, Spr01, WD23, ZLT25]. discretized [Asc97, BRY14, DSS14, HM18c, TH23, ZW20]. Discussions [Zaf22]. Disguised [JKM18]. disjoint [GHM24]. disk [ACH14, CG14, CMP21, CYIB12, Gla01, SX96]. disks [PP06]. Dispersion [MR12, FGR01, HLTA16, HFZ19, JBJB17, MA13]. displacement [AR10, HFZ19, Hu22, LDX23, YZ23, ZYGQ17, Zha19]. dissipation [MV13, Met19, YK22]. dissipation-preserving [YK22]. Dissipative [SFT03, AMCM06, KLB10]. Distance [DL04, SC03b, BF20, HZ93, HZ95, MKA26, UPY26]. Distributed [BD03, BD04a, LER03, TFPG19, AD17, AH23, CLQ26, hGzS17, GWL20, HDP18, HLTA16, Kha25, KXD25, LR18, LRL19, LCZ26, lLLVZ17, LC25, MFBB23, PLZ^{+}24, PPPN23, REM21, TA24, WLY^{+}21, dS00, dBD05]. distributed-order [AH23, hGzS17, HLTA16, LR18, LRL19, lLLVZ17, TA24, WLY^{+}21]. Distribution [HY21, DD99, EGT25, GST17, HHLS21a, HDL23, JWL20, KSV23, KK22a, Roh07, VDVJB12, WZ15b, ZZY26, HHLS21b]. Distribution-dependent [HY21]. Distributions [FLH04, TB25]. div [Ahu09]. Divergence [Mil19, Osw01, PM26, KP22, LN22, YH97]. Divergence-Free [Osw01, KP22, YH97]. diverse [SGO22]. Divide [MVV05a, KKM20]. divide-and-conquer [KKM20]. Divided [Sab03, Arg09, EGSHVN15, HVMMV24, Müh99, SR06, Tur26, Wal94, dC20]. division [ZGLH24]. Divisor [PR03, FGM19]. divisors [Gau13c]. DLN [Pei24]. DMLPG [MS14b]. Does [CL96a]. dogleg [ZYBJ23]. Domain [BD04a, Bog02, Kun01, Lor19, ZKD02, ZKD04, AR99, AK26c, BQ19, BF18, CZ94, CZ95, CCJ99, CDGC25, CV22, CGV26, Cro92, CYWH25, DHV22, ETY98, FHC21, GHM16, HLS10, IL05, KH11, MHR23, NPR08, NNCN23, REM21, RT19, Ria16, SB21, Spi25, YSLH19, YHR26, YL16, ZQL^{+}19, ZDZ26]. domain/multiresolution [YL16]. Domains [BD03, CP01a, HR03b, AN17, ALZ20, ABV23, AG17, AAD14, CKS24, CJK22, CK22, CGM25, DM98a, GH25, GL15, Kar09, Kar13, Kar15, KJC18, KZ21, Lee94, LSX10, LJ11, LC21, MD15, QL12, YL16, Yse99, YWZ19, ZLH22, ZZ23, ZYLN18]. dominance [LXP20]. Dominant [ABQ04, GLS26, LL16, LZX22, MVV05b, VBG96]. dominated [ETY98, MS20a]. Double [DMZ20, AGS08, AK26a, AKB15, BAK25, CEK21, CFRV23, DS20, GNT24, LL26a, MBJ17, OMU^{+}24, PRVI20, RR23, RCW22, Rou25, WL25, YJ18, YAY26, Zha26, ZWY22]. double-cone [GNT24]. double-diffusive [CEK21, YJ18]. double-projection-based [AKB15]. double-relaxation [WL25]. double-step [DS20, MBJ17]. doubling [CZLS18, Guo16]. doubly [DM98a, GLS26, HHHN07, LZX22]. doubly-adaptive [HHHN07]. Douglas [AG19, BCN^{+}16, BCS18]. downdating [LLZ18]. DPMHSS [WGZ18a]. DQ [AD22, CJK22, Gha18, Tom11, ZLZ22]. DQAINF [EO94]. Drift [DMYT23, Hag25, HY21, Mal21, Nac99, ST23, SYZ22]. Drift-implicit [DMYT23]. driven [DMYT23, HSY23, MKBY19, NT21, Ste24, WZ26]. DSS [XLW20]. DTZD [QZG^{+}19]. Dual [BEL23, NØ96, SC03b, SLZ25, Alt21, AABTB23, BGS24a, BC16, CWZ13, CXC25, CL19, CW21b, DCG25, DGL06, Gás99, GWL20, HLN25, HHF22, JWCZ21, JFSD25, Khe12a, MBG19, MLM25, MFBB23, MBM25, NAHZ21, RWTM21, SMZMA18, WZ11, WXQ20, WLW^{+}24, Wan25, WLG26, ZXL25, ZLS25, ZDZ26, ZFH26]. dual-complex [MBM25]. Dual-orthogonal [NØ96]. dual-permeability-Stokes [NAHZ21, ZDZ26]. dual-phase-lag [WXQ20]. dual-power [Wan25]. Duality [BF99b, BR17, BV95, MSP25]. Duffing [CZLS18, YX11]. duplexer [ZZX26]. Durbin [CHHL18]. during [Bag00, WXQ20]. Durrmeyer [MP14]. dwarf [SPS26]. DWR [LH23]. DWTPer [FC01]. DWTPer-based [FC01]. Dyadic [Mer94]. Dym [BAV18]. Dynamic [Cha04, LPV03, PP18, BRZ18, Fly22, HFW^{+}21, PPPN23]. Dynamical [Man10, GH22, GH23b, Hoa25, LK20, LWLT19, RM13, THF21, VTV22, WMCW21]. Dynamics [CCV23, AABM17, AM16, BCST14, BCM16, BCMT18, CZ20, CB00, GKQ24, Ila20, JJC24, JZ16, KGMH21, LZ23a, LSSS15, MCC26, OKP21, PPR15, Rei98, SSH^{+}19a, SSH20, ZXRL11, ZLQT19, ZHY^{+}20, ZLL^{+}21a, ZZY26].

E-algorithm [Mat91]. E-univexity [MSP25]. early [BRZ19, Osa12, TM14]. Easy [BvLP16, TZH24]. easy-to-implement [TZH24]. economic [BL23]. Economics [GCGF03]. ECT [Müh99, MT06, TM05]. ECT-B-splines [MT06]. ECT-splines [TM05]. ECT-systems [Müh99]. Eddy [Tsu02].

Edge [DMR20, NR25b, WPW25, BKM25, LXZZ21]. Edge-based [WPW25]. edgewise [KS26]. effect [JCF15, NAR05a, WS24b, ZJ08]. Effective [BFGM03, BBL22b, JH22, MT98, BKFMA11, BD98, BC99, BC01a, BI14, Cam19, GJV17, Haj16b, IP16, KPD25, MZZ25, RW06, YZ17]. effectiveness [BMR19b]. Effects [NZF11, SC03a, CR14]. efficiency [BCST14, Che14, EH24a, EGGSH13, HR07, Jai16, NR24a, Jai17]. Efficient [AMM18, AH18, AKW02, BDL^{+}12, BBQO07, BHS11, BKS13, BKF20, BCI14, CX20, CLPY23, CHMT10a, CRHTTN24, DB06, EGG08, FKMS01, GST17, HH05, HV15, HYJ20, JKC26, KXXW21, KKB16, Kar09, Kar10, Kar13, LM19, LWZ26, LL20c, MCG^{+}04, MP13, NW04, OM26, PW14, RhG15, SFMK23, SRK^{+}25, TRRD02, WZ15a, WKY25, XCY21, XLG22, YM24, Zag24a, Zag24b, ZBX21, ALQ17, AAU26, AA12a, AAA25, AHL20, BN25, BG11, BMV09, CJ20, CLBT15, CRHTV24, DM98a, DGM26, Dzu13, FZ26, Gha18, GH22, HZX21, JXNX24, JML25, Kac18, KKV22, KP96b, KNBGV18, KMA13, KXD25, KP25, KHM20, LDN16, LWZ18, lLhYfD07, LL18, LZZ23, MGL20, MD21a, MBR21, Mel24, MDH16, NBK17, PKR20, PXH25, PV22b, PSB26, PS16, Plo93, PM24, RNT26, RS26, SKD26, SGS13, SLD20, SSK23, SS24b, SS23b]. efficient [Sol15, VZ93, Vep08, Yan18, Yan22, ZH22, lZfLqL26, ZXL23, ZE10, Zag26]. Efficiently [WH04, WWM21]. Eftekhari [Eft15a]. Ehrenfests [BGL07]. eigen [LJbL21]. eigen-problems [LJbL21]. eigendecomposition [MVV05a]. Eigenfunctions [INR01]. Eigenmodes [LL05]. eigenpairs [BM19, GLL19, Pat25]. eigenparameter [Ghe18]. eigenproblem [BOP98, LL93]. eigenproblems [CVLX19, Ema96, MM99, SS99, VRM23]. eigensolver [HK14, IDS16]. eigensolvers [IS17]. eigenspaces [KP96a]. Eigenvalue [FG03a, TC05, Van07, AMM11, ALZ20, BLW09, BESC22, Bos21, Cve06, DBGB11, EGG08, EP97, FG07, FM19, FYI^{+}12, GLG26, HKP25, IDS16, IS17, IMT23, IJSS16, JKM18, JRRS08, KW00, Kol06, KSW09, KLW^{+}23, LLZ94, LY18, LDC10, LG95, LS25b, Mia19, OdZdRV13, PL99, Pan20, RSCH^{+}19, Saa23, San19, She15, WBY25, XHZ07, wYN18, YSXY19, ZLH22, ZS08]. eigenvalue-free [ZLH22]. Eigenvalues [IMT02, AAAA^{+}18, ANI15, ANI^{+}17, AM21, And18b, AA12b, AT19, AT17, BIM^{+}23, CHY19, CD96, DPP19, DPP22, EG19, EV22, EDAM13, EAGS20, FDV13, HKP25, HM18c, KP96a, LWG18, MT15, NP22, OO22, TBY13, YHZ20, YZX25]. eigenvector [JU22]. eigenvectors [CD96, Gar20]. eight [LSSS15, Sho18]. eight-order [Sho18]. eighteenth [Ste20]. eighth [AR13, BCMT18, CN16, CN17, CKKT16, FR18, Jai16, Jai17, LK20, LRM16, SS10, SZQS23]. eighth-order [BCMT18, CN17, FR18, Jai16, Jai17, LK20, SZQS23]. Einstein [BJT26, EJR26, Hua21]. Einstein-multidimensional [BJT26]. Ekman [TPLB22]. Elastic [BE03, MD21a, MR12, RW06, Sim98]. Elasticity [TC05, BK18b, DN19, KST06, WG25]. elastodynamics [AMM16]. electrical [CHH93]. electro [ÁCL11]. electrolyte [Met19]. Electromagnetic [ZKD02, ZKD04, AF13, GH06, GH09a, MDR23]. electromagnetics [ABG97]. electromagnetism [BR17]. electron [SSH^{+}19a]. Element [BBHM03, CP01a, DG05, PG05, Tsu02, ACF99, BBB22, BFK^{+}09, BX19, CLWH20, CLQ26, CVX16, CDS20, DMC20, Das26, DN19, DZS21, DZ19, DZ21, DMW23, DZ24, FZLL23, Far20, FZL^{+}16, FHC21, FLH25, GS25, GG22, GX19, GA20, Gha24, GLRSG08, GRT97, GCZH24, HCBAEC23, HFZ19, Hu22, HS21, Ila20, JKNR13, JCL16, JZYY23, KPFF24, KLL10, KBP17, KP22, LLAL21, LHW17, LR18, LLX20, LD21, LDX23, LGC24, LBF24, LWY26, LXZZ21, LTFL10, LS15b, LCH20, LYL15, LDL^{+}19, LN22, LH23, LZ23c, LCGH23, LZCW24, LAH22, LKKM15, MS20a, MC05a, MR12, MDR23, NAHZ21, ND21, NMM18, NT21, OMW21, OBAHK^{+}19, PV22b, PR10, PLZ^{+}24, Rah11a, SLW13, Sch08, SGO22, SS24a, SW10b, She15, Si12, SLT20, SR24b, SS94, Sut17, TKS24, Val15, WSY12, WYZ21, WZ22a, WD23, WDL23, WG25, WLY^{+}21, XSL22, YJ18, YZ23, YHL25, YH97]. element [YYW21, ZZH15, ZYGQ17, ZSF18, Zha19, ZZZ20, ZHFW21, ZD21, ZL22b, ZF22, ZYJY22, ZL23, ZGK^{+}25, ZLT25, ZWX22, ZYW22, ZLCW23, ZLTA16]. element-free [Ila20]. Elementary [Mai01, Rab92b, BP23, FLMR00, Fue07]. Elementary-Algebraic [Mai01]. Elements [FMS25, Kol04b, Mat01, Osw01, AMM16, AW23, AL23, BR17, CM99, Dri93, Rip93, WZZ26, XZ25]. elevation [FJ96, Maz11a]. eliminate [ddS24]. elimination [LZX23, MGL20]. ellipsoid [MW24, MW26]. Elliptic [ABMV03, BD02, NW04, SK04, AAU26, ACE99, ACH17, ACH19, AVI97, AK09, BFK11, BK08, Car95, CK24, CZ94, CZ95, CFL19, CLWH20, CJK22, DBAE09, Gha18, Van12, HP18a, Huc92, JK19, Kar10, Kar13, Kar15, KJC18, Kun05, LYL15, LW17, MM00, Mil20, MS20b, Moo20, MWWY13, OMW21, Pat25, PPS25, She15, SHD25, ST99, WL22a, WZ22a, XZZ22, XH20, YWWR12, YL16, YZ11, ZZ24, ZW14, ZZ19, ZZ23, ZWX19, ZLT25, MRU91]. elliptical [ALZ20]. EM-like [Wan15a]. embedded [AL23, BDJ11, CD07, RSKB17]. Embedding [BD03, BD04a]. EMC [LL22b]. EMC-HDG [LL22b]. Emden [SSR25, ITA24, SS26, SPS26]. emission [BG11, RVF07]. employing [CS08, SAE19]. Enabling [PTW22]. Enclosing [Pop18b, Wat06]. Enclosure [BL04, Eft15a, Eft15b]. encryption [LSL25]. end [HV98]. endpoint [Pop21, Zha23]. energetic [ADG10]. Energy [CY19, DCW23, LM04, LG19, YK22, Yue25, ZLQT19, ABI22, BWC22, CZH22, CLPY23, CLY25, Dam08, DW21, EAGS20, HKCW24, JTW25a, JTW25b, LZM23, LZW20, LL20c, LS20, Liu21, LL22c, LT24, MS24a, Miy24, Rab23, Rei98, Rip93, XF25, YZZL17, YZLZ22, YL22, ZYQ^{+}21, ZY23]. energy-conserved [YL22]. energy-conserving [ABI22, LZW20]. Energy-preserving [DCW23, LG19, ZLQT19, BWC22, DW21, HKCW24, Miy24, YZZL17, YZLZ22]. energy-stable [LS20]. enforcing [Kub23]. Engineering [CCK04]. Enhanced [SMN24, BC01a, GS19c, JRS09, SWJ24]. Enhancement [LHZ20b]. enhancements [Nov26]. Enhancing [NR24a, ZH26]. Enlarged [AR26]. ENO [ABT07, ADN17, Ber10, HMS96, OKB23]. ENO-based [OKB23]. Ensemble [JTW25b, Fan26, JY23, YYLX23, JTW25a]. entire [CD00, MCC26]. entries [CT06]. entropic [Hua96]. enumeration [CL24b]. envelope [Luc06]. Environment [LS03a, MJF09, NPS09]. enzymatic [SMK14]. epidemic [LWLT19, YH24]. Epsilon [Rob02, CHH^{+}20, CCHH23, JS15]. epsilon-type [JS15]. Equal [Mar04b]. equalities [GZP18]. equality [CCL18, Che94, CRN19, zDYG18, HHF22, JR20, LJ22, THF21, Wan17, ZLWZ21, ZZ22a, ZLZ23, ZFH26]. Equation [BK04, Boy05, Cro03, DOS03b, Gra03, HL03, KV04, Lui02, MN01, MS01b, ZL25, AGS08, AD17, AD22, AM24, AAH20, AM01, AZ19a, ASS13, AWL^{+}24, And14a, AMM16, Ant18, Ant22, ACH19, BQ19, BC14, BEHS20, BX17, BH92a, BZV16, BW15, BM09, BS19, Bou06, Bra06, Bra07, BX19, CZ23, CMD19, CM24a, CXL16, COSE22, CLTA10, CLA11, CLT^{+}13, CC16b, Che16b, CCJC18, CY19, CCHH23, CGSS24, CZZ^{+}25, CLY25, CSC26, CLQ26, CJKL23, CMWP20, Cui13, CYWH25, DWZ14, DL21, DMA09, DS09b, DAM16, DMD16, DZ13, DCW23, DH26, DI11, DLR24, DGST15, EH97a, EH24b, FYYW19, FZL^{+}16, FHC21, FSY23, FRS21, FS16, FYI^{+}12, GS25, GX19, GK20, Guo16, HT23, HS26, He16, HP18b, Hei07, HV22, HA16, jHyPlZ06, HZX21, HYT26, HLTA16, HTVY13, HM18a, HCXL20, HS21, HWZ24, Ila20, JBJB17]. equation [JZYY23, JWZ23, JJC24, JWY21, JH26, KKPT16, KS18a, KM24b, KCHD16, KR11, Kha25, KPFF24, KS26, KV07, KB20, KK22b, KP25, Lam09, Lee94, LWD23, LSZW19, LRL19, LHZ20a, LL22b, LKQ23, LZ23a, LGC24, LSL25, lLhYfD07, LYL25, LTFL10, lLHNS23, LDL^{+}19, LL20c, LD20, LP20, LLC20, LL20b, LS20, LW22, LWZ23, LZ23c, LCGH23, LHlLG24, LZQ26, LA22b, LF25, LHW13, LM25, LC25, MY22, MB09, MD21a, MBR21, MKA14, MV14, MWDD24, MS23a, MM09b, MM22, MWZL23, Mic91, MRU91, MClXzJe16, MDL15, MK17, MA13, Mok16, MMLM20, MM11, MA12, NS24, NPR08, Ngo23, NA26, OKP21, Oua99, PED15, hPwL09, Pen13, PV00, Pop18b, QQX23, RM11, RT19, RhG15, RK11, RR22, RS25, RS26, Sal17, San14, SN24, SB21, SKP20, SLA11, SDL^{+}23, SCF23, SD26, SKTGR19, SW22, SYLT14, SJW21, SM25b, SSA25]. equation [Str05, SzS21, SZQS23, SZ24, TA24, TZ24, Ter22, TDWL24, Val14, Val15, VdR13, WH15, WLMA21, WWBM21, WLZ22, WZC23, WS24a, WZZQ24, WC24, Wan25, WPY25, WZZ26, Wei18, WZ22b, XyJl16, XM16, XWC26, XZW13, XF25, YZZL17, YZLZ22, YSLH19, YZL20, YPL21, YC22, YXS22, YHL25, YLD11, YYLX23, YZL25, YL22, YJX15, YYW21, YD26, ZWFY19, ZWWW20, ZW12a, ZJ14, ZW14, ZHT15, ZP17, ZJWF18, ZZ19, ZYX19, ZJZ20, ZBX21, ZYQ^{+}21, ZZ22b, ZL22a, ZQzS22, ZF22, ZLLH22, ZLH22, ZY23, ZQS24, ZGK^{+}25, ZPEL25, ZZ18, ZLS24, ZZWK12, ZTZZ19, ZXL23, ZLTA16, ZS19]. equation-by-equation [DLR24]. equation-dependent [FYYW19]. equation-error [KV07]. equation-free [MBR21]. Equations [AAAS03, AG03, AEG02, AKKW03, Boy05, But02c, BH02, CM01, CC03, Che02, CGL01, DL03, DM03, ED05, EJR02, GF02, GSA03, HEOS04, Jbi03, KM04, KB02, LJW17, LER03, MO04, NW04, PL04, PG05, Rec01, RS02, SS03, SGM02, SER02, TDKB24, TRRD02, WHL24, vLV02, AS11, AR24, AAFL23, AAM24, AK19, AIA^{+}24, AB06, ATC16, AWH^{+}24, AAB13, ALJLYJ24, ABB15a, ARS26, AABM08, ASS11, AKB15, AE09, AA09, AM98a, ABI23, AR99, ALL25, ASGJ^{+}20, AVCO26, AC11, Ano17, ALZ21, Arg10, AH11c, AGS20, AG23b, Arn97, AAA24, AH18, AAD14, AHC13, ACH17, AKKT16, AKQ17, Awa10, AB23, BDN17, Bai97a, Bai97b, BRY14, BGR23a, BGL26, BAK25, Bar91, BBQO07, BJ98, BNN16, BDV18, BJNKR20, Bel94, BB14a, BAEF25, BQO99, BEQOR14, BHS14, BH17]. equations [BPS23, BHW23, BW13, BT14, BD17b, Bhr16, BK13, BK18b, Bic11, BV99, DOT21, BSF17, BEH24, BZ24, BS92, Bre99b, Bru93, BDR25, BB14b, BWB24, But10, CQ16, Cai22, CCN25, CZL26, CGPM00, CCJ10, CA07, CIP10, CS99, ÇD25a, CS25, CHH^{+}20, CP00b, CL05, CL06, Che14, CQLY15, CX20, CL21, CSZ22, Che22, CS22, CLPY23, CHYH24, CCC25, CH11, CC13, Che19, CCD10, CDGC25, CN17, CNR15, CHMT10a, CLBT15, CPS12, DJ10, DFJP10, VV07a, VV07b, Dah93, DWC18, DW24, DEC24, DBTG24, DSI11, Das19, DL18, DS26, DP21, DZW17, DW21, DW22, DCW23, DHWL25, DGM26, DW97, DZS21, DEM93, DB06, Don13, lDzS21, DM92, Dze15, Dzu13, EG10, Eft15a, Eft15b, EG18, EA12, ES19, EL01, EHV19, EED19, FY13, FWC16, FW25, FYM14, FNS19, FHL21, Fas25]. equations [FGJ00, FH15, FLH25, FS01, FMD23, Fra26, FS23, GA15, Gal18, GZL24, GM06, hGzS17, GKRS22, GG22, GM23, GH09b, Gon16, GCPG99, GLdO09, Van17, GKV23, GMT92, GM97, Gu20, GCZH24, GW24, Guo13, GMY18, GL19, GL23, GJVW24, GHZ25, GWY26, HJ18a, HH12, Hai09, Haj16a, Haj16b, Hal14, HZPW23, Han22, HM22a, HM22b, HM26, HCL21, HP18a, HV15, Hei06, HVYMS23, HVMMV24, HSTW14, Hey99, HSE16, HASI23, HST15, HZ20, HKCW24, HWC25, HL17, HM18c, HM18b, HM19a, HYJ20, HL20, HC24, HZ24, HZX20, Ibr25, IY15, IAAH24, ITA24, IZ24, ICR06, IJE15, JT96, JVH15, Jbi93, JZ25, JLFL19, JZF^{+}20, JM18a, JWL20, JWY21, JY23, JXZ26, KMH24, KPD25, Kam15, KJ18, KKB16, KCBT21, KT07, KSB08, KL25, Kaz24, KM17, KL17, KLF17, KGH14]. equations [KSCS07, KSCS08, KK22a, KE16, KLL10, KKM20, KS06, KKS24, KLR07, La 17, LS25a, LLAL21, Le 92, Le 98, LMMH11, LW13, LWwCL13, LW16, LHW17, LG17, LG18, LWZ18, LR18, LZL20, LHM20, LZlL20, LD21, LZ21, LDH23, LLD23, LZ24, LWLW24, LLM25, LLY25, LGHS25, LCZ26, LM14, LKBF17, LZZ19, LXZZ21, LDZ24, LL26b, LZ16b, lLLVZ17, LZW20, Lie00, LZ09, LWJ21, LT20, Lin09, LYL15, LSW16, LM17b, LW17, LF19, LM21, LJWW21, LSY^{+}23, LCZZ23, LZZ24, LJY24, LZCW24, LZT25, LWN13, LGM25, Lor19, LKK21, LHW13, lLXhL22, LV18, LV21, MKG24, ML20, MH21, MH23, MJJ^{+}23, MBJ17, MN23a, MAH22, Maj14, Mal21, Man21, MCW22, MPS20, MV17, Mic23, Mil19, MNS23, MP22, MDH16, MA22, Moh10, MG11b, MR96, Mon96, Moo07, MLM19, MWWY13, MAFN16, Nac99, NBK17, NM26, Nat07, ND21, NHP06, NLT21, NRV23, NNCN23, NAA19]. equations [OL23, OCWV24a, OCWV24b, OM26, PT19, PNW17, PQS22, PP24, PSB26, Pei24, PLZ^{+}24, PM25, PS16, PM26, Pie96, PV98, Pis16, PS09, PS17, Pog98, PR25, Prz16, ROB17, ROB18, RS20, RF23, RW24, RT12, REM21, RZZ25, RS97, RM26, RG10, SAC18, SS26, SSS14, SP21, San25, SS11a, Sch25, ST17, SW10b, SW14, SGS13, SA14, SMM25, SSYL20, SY20, Shi96, SCDM20, SL21b, SFS23, SZ23, SSL24, SHD25, SD25, SGJ15, SLT20, SS24c, SR24a, SK25, Sla06, Śmi06, Śmi09, Śmi13, Smi97, SS16, Sol11, SMN24, SHGL22, SMA99, SPS26, Sti18, SZX11, SZ20, SZ21, SG23, SS24d, SSH20, TZH24, TPY14, TX19, TH23, TZ25, Tem08, TS15, TY21, TL24, TN10, UA09, VH10, Van19, VLCL16, WLL12, WG13, Wan15b, WCB15, WZ15a, WCLW16, WGZ18a, Wan19, WW19, WZ19]. equations [WQL20, WMCW21, WCHK21, WSZ21, WHD22, WYZ22, Wan22, WD23, WQ23, WSLL24, WWL24, WQZH24, WLJ24, WC25, WZL26, WLL26, WZ26, WHS20, Wei17, WZ22b, WK93, WB17, Wri01, WC10, WC13, WCD21, XT16, XLW20, XL24, Xu19, XH20, XSL22, Yan95, YJ18, YHQ19, YBK^{+}21, YH21, YZ21, YJ21, YSLL23, YY24, YL16, YJJ^{+}21, YYL26, YXL18, YWZ19, Yua21, YP09, YQM16, ZTW19, ZR17, ZA20, ZBDK23, Zeg25a, Zeg25b, ZLW^{+}13, ZLG^{+}13, ZSF18, Zha20, ZY21a, ZHFW21, ZYW23, ZSLZ24, ZLM25, ZHW25, ZM26, ZXF14, ZFX14, ZXLF15, ZH23, ZCG15, ZL17, ZZB20, ZYW22, ZJ23, ZZY26, ZMZ26, ZZ10, ZYLN18, ZE10, aZ19b, dAFPR23, dFO11, FHS12, ABB15b]. equidistant [AAAGAD23, BSL18, Car10, SS94]. equidistribution [KMS05]. equilibrium [AAH18, AT21, Den14a, DHF21, HS20, HMA16, Hie18, Hie19, JA22, JM18b, KAF18b, KD18, MN17, NEY26, PT17, PK22a, PK22b, SSSS22, TLR^{+}25, VTV22, YHQ19]. equispaced [BT23]. Equivalence [ZAC26, HT21, Mit11]. Equivalent [AK09, Ern00, ZLZ23]. Erdélyi [PS17]. Ericksen [MWZL23]. ERKN [WW14, YZ17, ZYW17]. eRPIM [HS15]. Erratum [ABB15b, CZ95, CP95b, Cou15b, DCM^{+}13, DLL12a, DW15b, Eft15a, EHN17b, Gau13a, Gau17a, Jai17, MT14, PP17, Tov98, WZZ07a, dC16a]. Error [ATM19, Amo02, BEM99, BBHM03, BRS08, BRS09, Bro05, BJ02, CM24a, CLWV15, DLL12b, DFF04, EG18, FYI^{+}12, GEP14, GL04, Haj16b, HL02, IDS16, KPS22, KK22b, LW14, LL20a, LGP11, MY22, MV02, MP99, MM23, NIN12, OOOR12, Pej14, RRS09, SS03, SK25, SW26, Tir02, VPL97, WDY04, ZWX19, AD17, AHIJ22, AHJ22, AM24, AAN25, ART14, ALZ20, ASGJ^{+}20, AMM17, AMR15, ASHF21, AMKV96, AAA17, AKPW05, AKT15, Bac24, Bac25, BMA16, BGR23a, BDL^{+}12, Ber07, Ber11, Bör26, Bün24, CMRS00a, COSE22, CSZ22, Che19, CYM22, CY10, DLL13, DW22, DPS18, EL08, FLH25, GWL18, GEP16, GEP19, GA20, Van12, GLM15, HV15, HWC25, HCXL20, HS21, Ihs07, JL12, Jaw22, JV98, JLMP16, JL15, JJC24, KXXW21, KL17, KXD25, KV07, KKS22, KKS24, KKD25, KBP17]. error [LZ14, LS15a, LL16, LDL17, LWG18, LYH^{+}20, LZlL20, LCH20, LRY18, LN22, LW22, LCZZ23, LGM25, LRM16, LZ22d, MV13, MV14, MCW22, Mas95, MS24a, Meu97, Meu99, Meu05, MT13, MT14, MT19, MPT21, MT23, MN11, MAK20, Nar05b, NK16, Not12, PV22b, PT24, PZL15, QZG^{+}19, Roh07, RG10, Rum14, SSR25, SS26, SPB26, SS24a, SKP20, She15, SFS23, SZ23, SLT20, SN22, SH17, TTV21, Van19, WW19, WZ22a, WPY25, WJC25, WBY25, XC25, YJ21, Zag24a, Zag24b, Zag26, ZA20, ZZL17, ZLH21, ZWLZ24, ZZW26, ZWX22, ZZY24, ZLTA16, dC20, DLL12a]. error-controlled [AM24]. Error-free [OOOR12]. Errors [Dab04, Mat01, BD09, CH26, KL17, LDN16, MW16, Meu23, Tam10, WK20, Zas22]. ESIRK [BC01b, Che02]. ESPIRA [DPR23]. ESPRIT [DPR23]. essentially [LRY26]. Estimate [HSY23, SS03, AD17, ALZ20, BEM99, CSZ22, Che19, CYM22, JL12, LZlL20, LCZZ23, MN11, SFS23, SLT20, WPY25, WJC25]. estimated [EDAM13]. Estimates [AR20, BDF^{+}25, GS94, HL02, Meu05, Meu09a, WDY04, AR16, AAA17, AKPW05, BGR23a, BM94, BRS08, BRS09, Bro05, CMRS00a, COSE22, CY19, DL09, FLH25, Haj16b, HWC25, Jaw22, JL15, KBP17, LCH20, LZ18c, LN22, LW22, MY22, Nar05b, RZ23b, RRS09, SPB26, SS24a, SKP20, She15, SZ23, SN22, Tac12, WBY25, YJ21, ZA20, ZLH21, ZWX22]. Estimating [AH13, BG91b, Dea15, FDFM23, PT24, SSH19b, BQ19, VDVJB12, YC22]. Estimation [BJ02, DVJBN03, GD15b, MK98, Tir02, AHIJ22, AHJ22, Ali23, AWL^{+}24, ASHF21, BQ19, DWX17, Dra00, FZ26, FGP91, HHLS21a, HHLS21b, HCXL20, JV98, JFSD25, KP96a, KKS22, KKS24, LWLT19, LGM25, LRM16, Mel10, MPT21, Mic93, MAK20, Roh07, SS26, WWL24]. estimations [DW22, DGM26, KKD25, MCW22, NC94]. Estimator [KW04]. Estimators [ABMV03, MV02, ASGJ^{+}20, AKT15, LGP11]. Eta [SM25a]. etc [dC20]. Euclidean [Dam08]. Euler [ABB15b, ABB15a, BP22, CCN25, Che16a, CMP22, DMYT23, GMY18, GHZ25, Hag25, HSY23, Kar00, KL22, LDH23, LGHS25, LM17b, LCW25, Maj14, Mil19, NT21, PM26, Sab92b, Sab14, Sîn07, SMA99, SYZ22, TA26, WZ26, YW17, YJ21, YLYZ23, ZWW21, ZP23, ZG09, ZJ23]. Euler-type [LCW25, NT21]. European [AAN25, BN25, KN23, KMV17, Rou25, ZLT^{+}17, ZZ22b]. Evaluate [CCG01, BBZ95, CP95a, CD96]. Evaluating [ACH14, CC12, FLMR00, JXNX24, Ter23]. Evaluation [CR03b, GA08, LZRJ92, Pas06, RS97, BV93, BF99a, CW21b, FJ25, FHH05, FKP06, Gau08a, GHM24, GS95, HNSH09, HPS13, Hun95, IP16, JL16, Kie23, KP09, Kim21, KK23, Kob97, Lyn08, PSS10, Pas95, Pow93, SHF15, XX16, ZuI19, de 93]. Evaluations [SMB02, CS08, GP05]. even [Ber11, GN12, KSW09]. even-grade [GN12]. every [KH18]. evolution [AKQ17, WZ23b]. Evolutionary [CC03, LPV03, ST21, ASZ23, HFDSC24, LHR20]. EW [FW25]. Exact [AL23, HEOS04, KS18b, LA22a, Lóc18, AMR15, GL19, MN23c, SD26]. examples [Maz12]. exchange [BS21]. exclusion [DY93, Yak94]. exhibiting [SK25]. Existence [DD20, IAH20, LA22b, MS01a, NS01, BH09, KM19, Maz05b, Zaf22]. expanded [Hu22, JZYY23, LYL15, SKP20]. Expansion [FH05, LL05, LMMD05, ME95, Ter22, VH92, ZFZ19]. Expansions [CD01, LW04, MC05b, MC05c, GL20, HIK17, KV12, LGA^{+}00, LI10, Mül00, Pow93, QW08, Sid20a, Tem97, VPL97, Wal94, WGZ18b, WWL24, WL00, Zha23]. expectation [JR20]. expected [DGM26]. expensive [Liu11].

Experimental [MP00, CB13, Kno23]. Experiments [BCJ99, BM03, BHS23, CK05, Meu99]. Explicit [CJ12, CSC26, DK25, FHL21, GCJ26, IN95, JLWZ25, RW06, SW19, SFT03, Ver14, Wri02, ZHZ25, AHIJ22, AHJ22, Bla15, BHT16, CJSZ14, CT10, CDLW21, FLY25, FR18, GM20, IJ19, JA22, KCHD16, KLF17, KS20, Kha14, LW14, LDW18, LT20, LCW25, MCDS24, Sal25, SW10a, SL21a, SAHM25, SQG13, SSL24, ST92, SHGL22, SZ20, TX19, Ver24, WW14, WYZ22, YYK24, YH24, ZYQ^{+}21, Zha24]. explicitly [YJ18]. exploiting [Vos00]. exploration [DMZ20]. explored [RKM25]. Exploring [CL24a]. exponent [BEL23, LDH23, LZQ26]. Exponential [BL04, CR02, CIP10, DP01, Ixa19, JLP20, MVVA08, Sab03, Zah09, AH17, BKF20, BR07b, CCN25, CM24a, CL06, DBTG24, DMR20, DMR21, DZH23, FHL21, Gau15, Gau22, HWC25, Ixa21, JJC24, JWY21, KC23, LZ23a, LGHS25, LKBF17, LDZ24, NP18, OB16, PW14, Rad08, SFMK23, Sat24a, Sat24b, SMM25]. exponentially [VV07a, FR18, NZF11, VV11]. exponentially-fitted [VV07a, VV11]. exponentiation [CS18]. expressions [Ixa19, Pas99]. Extended [AB24, BEJ20, FLY25, HEOS04, LDW18, LWS18, AHJ17, ALL25, AT21, BD98, CCL18, CM25, GA20, GQ09, HHHN07, HZX21, HZ24, Ila20, McL25, Men26, MS24b, NBJA17, PP16, PP17, Pre93, RR23, Sch17, SH21a, WW19, Wu22, YW17, ZY21a]. Extending [AH11b, AGS20, BBC21, SH23, Sza92]. extensible [HM22c, OBAHK^{+}19]. Extension [TO21, AH11a, Eft15a, Eft15b, GL21, Han93, MA15, MFK^{+}15]. Extensions [Amo02, Osa92a, SH22, ABK22, Bou17, LSY^{+}23]. exterior [BF18, Loh22, Rab23]. extinction [YH24]. extra [Ter23]. extra-component [Ter23]. Extracting [KW00]. extraction [Fas25]. Extragradient [TCO26, DHF21, DJG18, HMA16, JA22, KD18, LQ20, NEY26, OAMA22, OR24, PT17, TQC22, TH18a, TH19a, TH19b, TTLD20, VS19, YLL20]. extragradient-like [LQ20, PT17]. extragradient-type [VS19]. extragradient-viscosity [TH19b]. Extrapolated [PN21, LKBF17, ZZ17]. Extrapolation [And19a, Ano92, BCK06, BBJ25, BT23, Bre04, CJSZ14, CH11, Coo03, FRR07, Osa92b, Rab92a, ABL… 12, BEJS21, BJT24, BJT26, BR07a, BC05b, BESC22, BRZ96, BRZ20, CXC25, CLTA10, DW97, DSS14, EG19, GPGVS92, JRS09, JS15, Kul10, KR20, LHW13, MRS25, MN92, NP22, PDS^{+}23, RVF07, Saf10, Sal94, Sid94, Sid17, Sid20b, SS08, SS12b, Tai92, Ton25, Wan19, dC20]. Extrapolation-based [CJSZ14]. extrapolations [YAY26]. Extremal [Thi93]. extreme [CHY19, MT19].

F [AG19]. F-B [AG19]. Fabrizio [SKM25]. facility [Xue95]. FACR [Bia94]. factor [BKS13, CWHL20, EK94, HZ15, SZ24]. factored [BEQOR14]. Factorization [BFGM03, vdHS02, ACSD16, BKDA25, BSB23, BRS26, CMM17, DLC14, DLDW21, DDRT97, EJR26, GHM16, Hua18, KBA26, KC24, KRS19, MVG21, Orb15, PLH20, SST92, TS92, TT06, XQZ24, ZCC26]. Factorizations [Tur94, BH92b, BK16b, DM22, GNT24, HHLM23, MP00, ST26, TK94]. factorized [FP18]. factors [Ste95]. Faddeev [PS06]. Fading [ZBDK23, CMD19, DC17, VMMD21]. Faedo [BT14]. fail [KSW07]. Fair [DHS09]. Falk [Har20, Mat15]. Falkner [LW13]. Falkner-type [LW13]. fall [CGM93]. falsi [CL06]. Families [Iva24, McL02, AG00, BD09, BE20, Den14b, FW13, KT07, NHP06, PP18, Rob97, XZ25]. Family [CPP14, NSS03, YWS20, ARTY20, ACM93, BD10, BCM16, BCMT18, CCTV16, CN16, CLBT15, CRHTTN24, EGSHVN15, FYYW19, GMZ19, HM14, Iva17, JLJ22, KKB16, KAF18a, KBA26, LK20, LSY^{+}23, MCC26, Mer92, Miy24, NBK17, NBJA17, OAR22, PM05, SS10, SKTGR19, Sto93, TV25, TA96, WZ13b, WSK14, WK14, WK17, ZCT19, ZJR25, ZCGS24, ZLM25, ZM26, ZG09]. Fan [Wat93]. Farey [DM92]. Fast [ASVC21a, ASVC21b, AM18, AAA24, Ave20, BD02, BD04a, BFGM03, BESC22, CMM15, CLMM05, CF05, CL21, CW21b, DH04, DZW17, DZH23, DSS14, EED19, FKP06, FM99, GST21, God15, GKV23, HG93, HIS26, HR03a, HL20, KP09, Kim21, KGG25, KRS19, Lee94, LMV00, LHZ20a, LZL20, LDH23, LX17, LJW17, Luc06, MS14a, NP18, NC94, PNW17, Pet01, RT19, Rum12, SWB08, Sle25, SVZ01, Som05, Tas93, WLY^{+}21, YQM16, Zeg25b, ZZ22b, ZFH26, vdMRS06, AA15, AR10, AMKV96, AVI97, BDL^{+}12, BC06, Bel08, BW15, BP19, BD00, CDSV11, CH22, CDGC25, DMC20, DM98b, DG17, DN24a, Ehr97, FZLL23, FLG08, GCZH24, GHZ25, HLL22, Hem94, HL17, HYJ20, Ihs07, JZF^{+}20, JWZ23, JW25, JWY21, Kaz24, KPT23, LlLZ21, LGHS25, LX23, LCZZ23, LWN13, LXQ15, LF25, Luc97, LKKM15, LV21, MFBB23, MDH16]. fast [OOOR12, PJ22, PPV09, Pow93, SzS21, Ter23, VV25, WLG26, WCH15, Yal01, YYW21, ZYW23, ZP23, ZPEL25, ZLS24, Ave20, Zeg25a]. Faster [Luc97, TL23, Cat24a, Cat24b, PDRG19, SI13]. fastly [ML22]. fault [CM05, GLRSG08, MS23c]. Favard [BGVHN92b, dR99]. FB [HBP13]. FCC [KK23]. FD [KLL25, LPRÁ24, YYK24]. FDE [BGI25]. FDEM [DdAF^{+}20]. FDEMtools [DdAF^{+}20]. FDTD [JLFL19]. FE [YSLL23]. Feasibility [EGSV04, BCK06, BBCS21, Buo17, DLLD17, GPGC98, KPC20, KS23, MWsC19, MG18, RT24, SPV20, SCD^{+}21, SIE16, TKSG23, TQW24, Wan18, ZW15]. Feasible [KPFG04, LLS11, AFG26, MLM19, PW22, mTLbJlL14, YZ17, ZFH23, Zhu15]. feature [YDW26]. feedback [YX11]. feedbacks [HLL22]. Fejér [Chk20]. Fekete [KST21a, KST21b]. Feller [ZL25]. FEM [CK24, GM20, KSJ26, LHM20, Lin09, MS24a, SKP20, TTP26, WLZ22, YWYN22]. Femlab [Sla06]. Femlab-Control [Sla06]. FEMs [CLY25, SZ23, YJ21]. Fermi [SS11b, ZZWK12]. fever [LWLT19]. few [Maz09b, Wri95]. fewnomial [BHNS16]. fewnomials [BHNS16]. FFT [KKV22, PW16, PWCsL18]. FFTRR [DG17]. FFTRR-based [DG17]. FFTs [AMKV96]. fhbvm [BGI25]. Fickian [RF23]. Fictitious [Kun01, YL16]. Fictitious-Domain-Lagrange-Multiplier-Approach [Kun01]. Fiedler [CFRV23]. field [CEK21, CDS20, Han93, HK14, LL20c, LZZ23, NAR05a, SZ20, SZ21, WG99, XCY21, YK22, YZBJ21, YZL25]. field/circuit [CDS20].

Fields [Che01, ZZX^{+}23, le 91]. Fifth [BM03, LMMH11, GB21, RGJ10, SSK23, ZG12a]. Fifth-order [LMMH11, RGJ10, SSK23, ZG12a]. filled [LGW14, fLxX12, PSWE23]. filling [CLMM05]. Film [TE03]. Filon [ZH17]. Filon-type [ZH17]. filter [Cou15a, Cou15b, GZ11, JS23, KK22a, KE16, Laz99, LZ18a, LKK21, PSZ23, PR93, SR25, Tur94, YY13, Zhu15]. filter-based [SR25]. filter-SQP [GZ11]. Filtered [DOT21, Mia19, GV99]. Filtering [Bre04, CR02, EKM03, Cic20, Sar06, ZZX26]. Filters [WDY04, LM97]. finance [KN23]. financial [ON26]. find [LMUZ19]. finders [ARY17, BCMT18, LK20]. Finding [FM04, MS01b, NAA19, NN26, PSWE23, AK15, BDN17, BEGG91, BC17, CCV23, CDGC25, CN17, GLL19, HKP25, LMMH11, LRL22, MCC26, OIM21, Pan96, PR14, PH20, QAS^{+}24, SDMMK18, TLD^{+}23, THS20, WK17, WMCW21, ZXRL11]. Finite [ABMV03, BBHM03, BRMG18, CP01a, FZL^{+}16, For22, HM19a, JKNR13, KLT03, LTFL10, Mat01, ML10, Osw01, OB16, SW22, SMA99, SS94, Tsu02, ZL22b, ZLCW23, ANI15, ACF99, AE09, AAIT94, AHL20, AW23, AL23, Ber07, Ber11, BBB22, BFK^{+}09, Bia12, BW15, BR17, BX19, CWZ13, CM99, CL10a, CXL16, CHH^{+}20, CLA11, CLWH20, CS12, CV15, CDGC25, CL93, CVX16, DL21, DAN26, DS15, DW21, DZS21, DK00, DZ19, DZ21, DMW23, DZ24, EA12, FZLL23, Far20, FHC21, FLH25, GS25, GHC15, GMZ19, GG22, GX19, GS14, GA20, Gha24, GLRSG08, GGS22, GRT97, HLL22, HCL21, HKE97, Hoa25, HFZ19, Hu22, HTVY13, HL20, HS21, JL26, JCL16, JZYY23, KCHD16, KAF18a, KPFF24, KLL10, KBP17, LLAL21, LWD23, LHW17, LR18, LRC19, LMUZ19, LLX20, LD21, LDX23, LGC24, LWY26, LJ11, LZW20, LS15b]. finite [LDL^{+}19, LCW20, LH23, LWZ23, LZ23c, LZZ24, LZCW24, MS20a, MY22, Maj13, MC05a, MDR23, Mil13, MA13, MA12, NAHZ21, NW19, ND21, NMM18, NV21, NT21, OMW21, PNW17, PV22b, PLZ^{+}24, QQX23, QXGZ20, Rah11a, RP26, RS93, SS26, Sch08, SS24a, SW10b, SHD25, Si12, SLT20, Sv95, Śmi06, SS16, SR24b, Spr01, SSA25, ST98, TKS24, VW08, Val14, Val15, VLCL16, WSY12, Wan15b, WYZ21, WZC23, WD23, WDL23, WQZH24, WG25, WZZ26, Wei18, WLY^{+}21, XX16, XSL22, XZ25, YJ18, YHQ19, YZ23, YHL25, YH97, YQM16, ZZH15, ZSF18, Zha19, ZZZ20, ZHFW21, ZD21, ZYJY22, ZL23, ZGK^{+}25, ZLT25, ZYW22, ZLTA16, iV12, ZLT^{+}17]. Finite-difference [BRMG18, CS12, DK00, NV21]. finite-element [GRT97, NMM18, Sch08]. finite-part [DS15]. finite-step [ANI15]. FIR [Cou15a, Cou15b]. First [ACE99, LGL23, Mar04b, AGG17, CC18, DHJJ10, EGSHVN15, FHL21, FGL19, HVMMV24, JWCZ21, LM21, MS06, NS24, PP24, PP18, RSKB17, WHZ^{+}18, YCZ26, dAFPR23]. first- [WHZ^{+}18]. first-kind [LM21, NS24]. first-order [AGG17, DHJJ10, FHL21, FGL19, HVMMV24, JWCZ21]. Fischer [ZLLC11]. Fisher [Ila20, TA26]. FISTA [YJS25]. fit [BWB24]. fits [GGN18]. Fitted [CV15, EA12, OCWV24b, Val14, VV07a, FYM14, FR18, JSF13, Li17, LG19, MVVA08, NBJA17, SS15, VV11, OCWV24a]. Fitting [ASS03, Wat06, AS08, AG00, CIP10, CK05, CV92, Ell93, Ixa21, MKS18, Sch08, WD95, WK93, Wri95, lZfLqL26]. fittings [Ixa19]. Fitzpatrick [XZL12]. five [But98, MA95, QZG^{+}19]. five-diagonal [MA95]. five-step [QZG^{+}19]. Fix [GZ20]. Fixed [GF02, XWC26, AVCO26, Bag16, BMR97, BRZ19, CSI16, CSI17, CSI18, CCL18, CD99, CRN19, Den14b, Fan15, GM23, HS20, HCD26, Ihs07, IS22, IAH20, IKMS24, JM18b, KAF18a, KAF18b, KLZV95, Lei15, LP18, LLD23, LRL22, LCL21, LM25, Lyn08, MWsC19, MN17, NEY26, OMU^{+}24, Osa92b, PKC19, PM22, PKC18, PG12, PPPN23, PDRG19, PN21, RZ16, RT22, RTTH22, Sab91, SRK^{+}25, SIE16, SI18, SC18, TAM21, TH19a, TH19b, ZFC18, ZH19, ZAC26]. fixed-amplitude [Lyn08]. Fixed-Point [GF02, AVCO26, JM18b, PM22, PPPN23, PN21]. Fixed-time [XWC26]. Fixing [BBC21]. flat [YXL18]. Flattened [JYLC21]. Fletcher [BKG15]. Flexible [BDR25, JGMC25, PV00]. flights [JS21]. Flipped [FFSC25]. Floater [dC16a, AHP20, MNS23, TB25, dC16b]. Floating [DHL^{+}04, DH04, GWBC20, GLM15, JLMP16, KGG25, OO22]. Floating-Point [DHL^{+}04, GWBC20, GLM15, JLMP16, OO22]. Floquet [PR25]. Flow [AA03, CGN03, Jan03, BBL23, BM24a, BBB^{+}06, BGL07, CL00, CLPY23, CJKL23, CW17, HHST19, HMS96, Hol98, KLB10, KBP17, Liu21, Met19, MSMS12, MSM12, NZF11, NAHZ21, Pan18, Son93, ZD21]. flows [ASZ23, CEK21, DG17, FT24, FHAL15, GB21, GS26, LC21, LL22c, MDR23, Mot14, RRZ21, TCOA19, Uhl22a, Uhl22b, WSY12, WCB15, Yue25, ZS22]. FLR [HHHN07]. fluid [ASZ23, ABV23, BBL23, BGL07, CC18, Hol98, KBP17, NZF11, NAHZ21, Pan18, WSY12, XL24, ZD21]. fluids [FHAL15]. Flux [Jan03, LDX23]. fluxes [Che19]. FMM [KZ21]. Foerster [HT23]. Fokker [LD20, MM22]. Following [MP02, CVLX19]. FOM [EHTSM21, Ess98, RS02]. Föppl [YXL18]. forced [CZLS18]. Forchheimer [LRC19, LCW20]. forecast [NZ19]. Form [GL04, Neh04, AAAA^{+}18, Bos21, BGR25, CDW95, CS08, FLT09, GHC15, GCPG99, HRY19, JR20, KCHD16, KMRAR25, KP09, KKD25, MRV23, Mar96, MK94, MK97, REM21, Sal25, Sch09, SMM25, TPY14, WZZ07a, WZZ07b, WCY25, XZL12, ZW12b]. Formal [Arn97, BV95, Dra96, Dra02, FT05a, Bec96, CGV92, Dra00, KP22, MT98, TBA94]. formally [WZQ17]. forming [KS14, MT15]. Forms [Che01, Sab03, CD99, Her96, Loh22, MZ99, Rob92, RSCH^{+}19, Ska13, SH17]. formula [AB24, Ash19, AP21, AAH24, BK16a, BCM07, Ber93, Ber07, Ber11, BG13, Cam95, CKP22, CM92, CY10, HMS11, LMUZ19, LM15, LSM16, OPSM22, SS14, SH21b, SSH19b, YKY15, ZQL^{+}19, ZLL^{+}21a]. Formulae [KP03, AH21, AHP25, CS12, DPS18, Gau00, Gau09a, LJX25, Not95, Not08, Not12, Spa24, SH23, SH26, SSH20]. Formulas [IM02, AAP24, BHS17, BGVHN96a, BC09, BD25, Che13, CD07, CBGVN07, For22, GM92a, GS94, Hai08, Kal00, KS06, Lau07, lLLVZ17, MM08b, MM09a, MG91, MG94, NBJA17, Spa07, Spa20, YZH21]. formulated [BSF17]. Formulation [FM04, BH92a, DAN26, GM20, JM00, KPS22, Mez22, MG11b, Oua99, SFZ22, dOS07]. formulations [Cdv98, HJ21, LS25a, SGO22]. Fort [DCW23]. Fortran [Lai92]. Forward [YAY26, BC16, CS24, Cho16, JCH23, SBG24, SB21, SSP15, SZ20, Ton25, ZWLZ24, ZL17, dC22]. forward-backward [CS24, Cho16, SBG24, SB21, SSP15]. forward-backward-forward [BC16]. Forward-reflected-backward [YAY26]. Four [FGR01, MC05b, MC05c, NS01, AG19, AABM17, Bos21, CFR06, HMS96, Lai92, MDL15, PTW22, Sab91, VV11]. four-dimensional [PTW22]. four-direction [CFR06]. Four-Directional [NS01, Lai92]. four-operator [AG19]. Four-stage [FGR01, VV11]. Fourier [Ave20, AVI97, BD03, BBM08, BHS17, CG07b, CCW21, DHMS16a, Hom92, Hom98a, HIK17, JWY21, KPT23, LGA^{+}00, LSX10, LZ23a, Mel10, PPV09, SI13, SW14, TZ24, WDY04, Wri95, ZHT15, ZKD04, dFG93]. fourteenth [SS11a]. fourteenth-order [SS11a]. Fourth [DMA09, FG03b, KMV17, NA26, Vul97, AHC05, Bac21, Bac23, Bac24, BCM16, Bia12, BW15, BV09, BGZ20, CRHTTN24, CRHTV24, DD20, DNV25, DZ13, DDRS23, DZH23, FW25, HZX20, KCBT21, KM24a, KLF17, LKQ23, jLyLqW17, LA22b, ME95, Miy24, ON26, SSR25, SP21, SGS13, SMM25, She15, WGK11, ZCT19, ZP17, ZYX19, ZF22]. fourth- [KCBT21]. Fourth-Order [FG03b, DMA09, KMV17, NA26, Bac21, Bac23, Bac24, BCM16, BGZ20, CRHTTN24, CRHTV24, DD20, DZ13, FW25, HZX20, KM24a, KLF17, LKQ23, LA22b, Miy24, ON26, SSR25, WGK11, ZCT19, ZP17, ZYX19, ZF22]. Fowler [SSR25, SS26]. Fractal [HAN24, KA26, Man07, PK25b, PK25a, PGC24, ASV23, CA22, CAV23, GHM24, PGP^{+}25, VP23, dCOS21]. fractals [GHM23]. Fraction [Van03, LM15, LSM16, Mor11, SKJ^{+}18, WL00]. Fractional [DFF04, DM03, KK22a, LGM25, ML20, NRV23, SKM25, TDKB24, WSLL24, WHL24, ZL25, AS10, AD17, AD22, ALB^{+}18, ADP26, AH23, AAN25, AWL^{+}24, AVCO26, ALZ21, AH18, BD17b, BKF20, Bhr16, BZV16, BuR24, BGI25, BX19, CCTV23, CMM15, CG20, CX25, CLA11, CC18, CY19, CX20, CL21, CDLW21, CS22, CGSS24, CLQ26, CJ17, Cui13, DWZ14, DOR26, DAM16, DS26, DP21, DZW17, DHWL25, DGM26, DW97, Die08, lDzS21, DN24b, DNH^{+}25, FAMA20, FNS19, FZLL23, FZL^{+}16, FS01, GZL24, hGzS17, GKL21, GK21, GCZH24, GHZ25, HZPW23, HCL21, HP18b, HV22, HA16, HZX21, HLTA16, HZ20, HTVY13, HL17, HYJ20, HL20, HS21, HWZ24, ITA24, JBJB17, JZF^{+}20, JWZ23, JJC24, JB24, JXZ26, Kam15, KS18a, KZS21, KN23, Kaz24, KR11, KLF17, KXD25, KKA17, KB20, KP25, LLAL21, LWD23, LHW17, LR18, LSZW19, LWLT19]. fractional [LHZ20a, LHM20, LLX20, LLM25, LGHS25, LJX25, LCZ26, LKBF17, LYL25, lLLVZ17, LWJ21, lLHNS23, LRY18, LDL^{+}19, LD20, LJWW21, LWZ23, LF25, LV18, LV21, LC25, MKG24, MH23, MZG26, MD21a, MN23a, MCW22, MS23a, MSP25, MA13, MG11b, MM11, MAFN16, NM26, ND21, Ngo23, NT21, OCWV24a, OCWV24b, PNW17, PQS22, PED15, PSB26, PLZ^{+}24, PS17, PM24, PGP^{+}25, QXGZ20, ROB17, ROB18, RF23, RW24, REM21, RhG15, RR22, Rou25, RS25, RS26, RM26, Sal17, San25, SFMK23, SR16, SN24, SM25a, SKD26, SLA11, SCDM20, SCF23, SHD25, Sid20a, SS24c, SR24a, SK25, SMN24, SYLT14, SzS21, TA24, TZH24, TH23, TTP26, VLCL16, Wan15b, WH15, Wan19, WXQ20, WLMA21, WWBM21, WYZ22, WZC23, WZ23a, WQZH24, WC24, WZL26, Wei17, Wei18, WCD21, YZL20, YPL21, YH21, YXS22, YWYN22, YHL25, YY24, YCW^{+}19, YJX15, YWZ19, YQM16, ZE12]. fractional [ZA20, ZSGJ25, ZWFY19, ZWWW20, ZLW^{+}13, ZP17, ZJWF18, ZJZ20, ZBX21, ZHFW21, ZYQ^{+}21, ZZ22b, ZL22a, ZF22, ZYW23, ZY23, ZGLH24, ZXF14, ZZ18, ZZB20, ZYW22, ZLCW23, ZLS24, ZYLN18, ZLZ22, ZXL23, ZLTA16, aZ19b]. fractional-multiplicative [MZG26]. Fractional-order [NRV23, FAMA20, ROB18]. Fractions [HLM04, Pas03, Arn97, LB93, Lor95, Now06, Now13, Pas92, Røn92a, WQ10, dB07b]. fractures [LCW20]. fragmentation [SW22]. frames [BF93, VW08]. framework [BDD20, GWW15, HPS20, VB92, WCB15, lZqL25]. France [BV96]. Frankel [DCW23]. Fréchet [AMH10, AMA21, AH08]. Fredholm [AS11, AR24, ASS11, Bic11, Bru93, DS26, EHV19, HST15, KM24b, MWDD24, OdZdRV13, PP24, PS09, RM26, WZL26, ZA20]. Free [Osw01, AAFL23, AAM24, AAB13, Ahu09, ALY22, AH08, AM16, AG23b, BRZ98, BZ18, CEZ26, CHS19, CGHH21, Cdv98, CKKT16, CRHTTN24, CO19, FT14, GLLJ12, IAAH24, Ila20, Jia20a, JXNX24, KKB16, KS14, KP22, LZ21, LF19, LN22, MBR21, MCDS24, NBK17, OL23, OOOR12, PS16, RAH11b, SBN25, SA14, SGJ15, TV25, WZZ07a, WZZ07b, WZQT15, WZ16, WGZ25, WB17, YKY15, YZBJ21, YH97, YYL26, ZZY18, Zha20, ZLH22, ZLLC11]. free-form [WZZ07a, WZZ07b]. freedom [BR17, HHLS21a, HHLS21b, RRZ21]. freezing [VVL25]. frequencies [Ave20, Pan96]. frequency [BMR19b, Cou15a, Cou15b, KKV22, KPT23, SW19, WW14, YZ17, ZYW17]. Fresnel [ZA24]. friction [PPS25, ZS22]. Friedrichs [ZZ10]. friendly [Ixa19]. Frobenius [CVLX19, CUY24a, CUY24b, HM18b, LEK21]. Front [Lie00, CL96b]. front-oriented [CL96b]. frontier [PLVB11]. Frontini [BD10]. FSAI [BM12, JCF15]. Full [CS99, Lui02, vGK04, AG15, AF13, APST21, Khe12b, Khe16, KH20, LYY12, LS07, MR09, WZVJ22, Khe17]. full-Newton [AG15, Khe16, KH20, LYY12, LS07, MR09]. full-NT [Khe17]. full-rank [APST21]. full-wave [AF13]. Fully [HKCW24, Pan18, WF23, CGPM00, DWZ14, DD20, DD21, GM06, HWZ24, KP96b, LRC19, lLLVZ17, SK25, Van19, WPY25, Wei17, YY24]. Function [DHL^{+}04, LM04, Mat04, Sab03, WDY04, Wen03, vGK04, AH17, AMH10, AMA21, AZ19a, Alz08, AGN07, BBQO07, BF20, BQO99, BEQOR14, BEJ20, BEJR23, Bic24, Bro05, Cau22, CAV23, Che13, CDD21, Col92, CFK^{+}20, CZX26, EGT25, EV22, EL08, FW13, GA08, GRAST23, GCFF95, GMT92, GCJ26, HM22a, HA16, HKPW19, JMS16, JYLC21, Joh15, Joh20, KL94, Khe12a, KH20, KW96, KKK22, LZRJ92, LGW14, fLxX12, LYY12, LX17, LD20, LM15, LSM16, MCC26, Mar96, Mel10, MMU20, MS15, OPSM22, Ost07, PSWE23, Pea13, PH14, Riz18, SS11b, SMNZ20, Sun94b, Van92, VGM96, WZ11, XWC26, XL14, XC25, YA24, YP09, Zag24a, Zag26, ZSGJ25, ZZL17, ZLL^{+}21a, ZLLC11]. Function-Based [vGK04]. function-valued [GMT92]. Functional [LM01, AR99, Bic11, CQ16, GWY26, LTFL10, MKG24, MAF25, NS24, Rab23, SS11b]. functionals [AAPR21, BM96, GAM24, Hag13, LM12, PP21, Sch14]. Functions [ASS03, Amo02, Ano95b, Bre04, But02a, CD01, Dab04, GG01, GST02, GST03, GKS04, GPP01b, GF02, KM04, Kol04b, PW04, Str02, WH04, AF23, AG25, ALW98, ASV23, AAN25, AAAGAD23, Ash19, AP21, AAH24, AK26c, BBZ95, BBL22a, Bel94, BT23, BHJTM92, BP93, BC17, BP23, BX19, VVV22, BGVHN92a, BGVHN92b, BGVHN92c, BGVHN92d, BGVHN96b, CM15, CL11, CLMM05, CG07b, CM24b, ÇD25a, CGW25, CR23, CJTW96, CBGVN07, CD00, Dar99, DS09b, DBH21, DP21, DPP19, DPS18, DL01, DGST15, EAB20, EG94, EHV19, FVV25, FT05a, FLMR00, FM19, Flo16, FDFM23, Gaj05, GGNF17, GA15, GV99, GD15b, GG08, Gau09a, Gau10, Gau11a, Gau14, Gau17b, Gau22, GM22, GK20, GH95, Gla01, GS19b, GSF26, HGVPA92, HW18, HSK20, HL15b, HM06, Hun95]. functions [IP16, Jaw22, JL15, Jos22, KSB08, KCHD16, KD14, Kel07, Kie23, LV15, LWZ26, Liu11, jLyLqW17, LM08, Low05, MBR21, MM00, Mas95, MS14a, MCMX20, MS11, Mül00, Nar05b, POP17, Pep23, PGGC97, Pet95, PK25b, PK25a, PGC24, Rab05, Rei98, Rim25, Røn92a, Saf10, Sas93, Sch26, SHF15, SM25a, Seg98, Seg08, She00, SS23b, Sle25, Śmi06, Spa20, Str97, SR06, SH23, Sun94a, TA24, Ter23, UTO07, Vep08, VP23, WCY25, WZ15b, Wil12, Wri95, XLC93, XX16, Yak94, YWS20, YYZ22, YLL22, Zag24b, ZuI19, ZJ08, Zha23, Zhu21, ZS19, dC22]. Fundamental [BS04, LL05, Ant22, BCJ22, BCJ24, Kar15, KL25, RT19, ZLLH22, ZLH22]. fundamentality [Sun94b]. Further [Che01, LRT19, ZDSY20, KM13]. fused [JFSD25]. future [CZ20, QZG^{+}19, ZQL^{+}19]. fuzzy [DHWL25, dCOS21].

GADMM [ZSWW25]. GADMM-based [ZSWW25]. Gale [BHNS16]. Galerkin [LR18, MS14b, SZ23, WLZ22, ADL05, ADG10, AK12, ALZ20, AMM16, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BGR23a, BQ19, BFK^{+}09, BK18b, BX19, BGZ20, CQ16, Cai22, CG20, Che19, CYM22, DWZ14, DMC20, DDG05, DB06, DBAE09, DOS03b, EG18, FMD23, GM06, Gha24, GPP01b, Ila20, JWY21, KMH24, KM24b, Kha25, KPFF24, KS26, LHW17, LHM20, LGC24, LLY25, LSW16, LW22, LZ23b, LZCW24, LI10, Lya97, MS20a, MCW22, MFPG07, MM09b, MM23, Moo20, MWWY13, MM11, MAFN16, ND21, RNT26, RMT13, RBN14, RM26, SHF15, SLT20, SN22, SA23, TZH24, TKS24, WLMA21, WYZ21, WZZQ24, WG25, WBY25, WZZ26, Wei17, Wei18, XZW13, YZ21, YJ21, YZ23, ZJWF18, ZHFW21, ZZ18, ZLTA16, dFO11]. Galilei [AD22]. game [DI11]. games [BK18a, HV98]. Gamma [Alz08, Mat04, Cau22, Che13, FW13, GA08, LM15, LSM16, MS15]. gap [Rab23]. gaps [AM21]. Gating [GS19c]. Gating-enhanced [GS19c]. Gatteschi [All08a, All08b, GG08]. Gauge [ZSF18]. Gauss [EDAM13, AAPR21, AR18, AH11b, AT17, AAH24, BEJR23, BK04, BCI14, CR03a, CFK^{+}20, DLL^{+}24, DPS18, DDRS23, FG07, Gau00, Gau09a, Gau10, Gau14, GST21, Gon16, GS16b, JFSD25, KP03, MGL20, MVVA08, MM08b, MM09a, MT23, MS24b, Not95, OPSM22, POP17, Pej14, PP21, RWTM21, SH23, SH26, VV11, YWYN22, ZH22]. Gauss-type [Pej14]. Gaussian [AB24, ATM19, Ash16, Ash19, AK26c, BC00, Bou03, DK00, Hag13, KN18, Kza97, LMV23, LMD25, LZ25, LZX23, Man07, MSM25, Mil95, Mil17, Pet95, RN21, SS01a, SS01b, Spa07, Spa20, Spa24, UL18, Van25, XW17, XLL26, ZAH26, dABR01]. Gautschi [Kou07, LR14, SW19]. Gautschi-type [SW19]. GB [Kva01]. GB-Splines [Kva01]. GC [dB07a]. GC-sets [dB07a]. GCD [LL14]. GD [ZLG^{+}13]. Gegenbauer [AH23]. Gene [Meu09b]. GenEO [Spi25]. General [ASS03, AHKW04, BJ04, DEP12, DL03, SR16, WL25, Wri02, AH11a, AHIJ22, AHJ22, AP26, AF23, ABS19, BS17, BX17, BBd95, BDD20, BGR25, BH09, CQ16, CHYZ98, CM98, CJSZ14, CGN22, CYWH25, CV92, DEC24, DZH23, FLY25, GS16a, GL15, GZ11, GL19, GEA20, Haj16a, HWCR19, Hil10, HRY19, Jbi93, Kno23, Leo07b, Leo08, LWM10, LZL23, LWS18, MS26, MZG26, MRV23, Maz11b, Meh11, MCDS24, MHA16, OI14, PP21, RKMS16, RGJ10, REM21, RWTW19, SD26, SK19, SDMMK18, Son93, USAF14, VB92, WK12, WSK14, WCB15, WPY25, WPL18, YL16, YZY^{+}14, ZYW17, ZW12b, Zha20, DDP14]. generalised [CDS20]. Generalization [Che99, IM02, CM16, Hua94, LQ16, Sîn07, SZ99, WT08]. Generalizations [CKY99, LCHH21, BL92, CHH^{+}20, Yue25]. Generalized [AH14, Ash16, BCN^{+}16, BR21, But02a, CM01, CCTV23, CG07a, CCHH23, Cse04, DMS09, DS12, For93, FLG08, GG22, Haj16a, Hey99, Ibr25, JH26, MAH22, Mas95, SZX11, TV17, UA09, VA20, WG25, ZZ19, APST21, Arn97, AH18, AAA17, BH22, BLW09, BBQO07, BGR23b, BQO99, BV21, Bos21, Bru93, CCN25, CZL26, CA07, CG20, CZ14, CC15, CC16a, CC18, CVLX19, Che19, Cho16, Coo09, Cor91, DW24, DAN26, Den14a, DHWL25, DDRS23, DLDW21, FWC16, tFZyZ16, FJ25, Fas25, FSY23, FRS21, Flo16, GZ20, Gar20, Gau09a, GM21, GLG26, GS21, GW24, HL23a, HNY^{+}18, HWXC17, HM18b, HWXC18, HJ21, IDS16, IS17, IAH20, JMS16, JXZ26, KM17, LRT19, LDN16, LWAG08, LL93, LLZ94, LCVL18, LMUZ19, LLY22, LKQ23, LLD23, lLhYfD07, LZL22, LLL22, jLyLqW17, LCW25, Lu15, LS25b, MMV19, MM17, Mat96]. generalized [MMU20, Müh99, NP18, NUG25, NW17, ON26, hPwL09, PDRG19, Pre93, RF23, RT22, Sab92b, Sab14, SSS14, SSS21, SN24, SCW17, SDL^{+}23, SCF23, Sid17, SWB08, Sle25, Śmi09, Śmi13, Sol15, Spa07, Spa20, Tak17, Tam10, TB25, TM10, Val14, Val15, VPA24, VPA25, Wal94, WK16, WCY25, WZ22b, Wil12, XM16, XLL26, YDWL15, YHZL21, YSXY19, ZRZ11, ZWW21, ZH23, ZCTD24, ZLLC11, dCOS21, dB07b]. generalized-Sylvester [SDL^{+}23]. Generalizing [ADN17, CUY24b, For21, CUY24a]. generate [dCOS21]. generated [CGSS24, GP05, Kim21, KA26, Man10, Mor25, TM20, VW08]. generates [MMU20]. Generating [Ter22, CFR06, JMS16, KCHD16, Lig93]. Generation [VVV22, CG05, OO22, Tan20, DLR24, KGN^{+}24]. generators [BR17, Hua18, SBR26]. genesis [BRZ19]. Genetic [CSFC04]. genus [Kie23]. Geodesic [LG08, GLV05, PTSB01]. geodesics [PTSB01]. Geometric [Ber14, JK19, RDdRC04, Uhl09, AAA25, BC05b, CG09, Cas17, Dri93, HK06, IDAV09, KSB08, LWG18, MRS25]. Geometrical [CR20, GLV05, LG08, LAH22, MP07]. geometrically [Maz09b]. Geometry [DH04, ST02, BHH24, DS23, KL94, MN23b]. Geophysical [De 02]. Georges [Le 19]. Geronimus [CDT10, DM14]. Gerschgorin [PP06]. Gerschgorin-like [PP06]. GFR [ZWG18]. Gibbs [ALY22, Bre04, DF01]. gift [AHL20]. Ginzburg [HP18b, HWZ24, WS24a]. GIPSCAL [lZfLqL26]. given [BCCP25, MLM19, VGM96, Wri95]. Givens [ZZW26]. GJK [DL04]. GKB [AR26]. GKO [Pol10]. Global [AHIJ22, AHJ22, BBBC23, CN01, Cse04, Cui26, DW12, FS16, HASI23, HL23b, JV98, LPV03, LM04, MS26, MV02, MCG^{+}04, Ovi22, QW08, Sch25, Tai92, Tir02, VR04, WW19, WCW20, Zil01, dSCS04, AGS08, AHJ17, AE18, BEJ20, BEJR23, CR23, Deh20, GL12, GO20, Van17, Hey01, IKMS24, JJK97, KS06, LL26a, LGW14, lLXZ24, NR24a, PSWE23, Prz16, Rog95, Seg98, Ste24, WB17, YLL22, YY13, ZTW19, ZW12b]. globalization [Shi96]. Globally [BKM03, CX25, Har18, MA22, NKA24, WCY25, CCC25, HL23a, KL22, Pie96, PR93]. Glowinski [CWHL20]. GMRES [BHS23, CKY99, EHTSM21, Ess98, GP14, JGMC25, JR10, LM99, MT15, NG23, RS02, Smo99, SA26, TM14, TM20, ZTW19]. GMRHSS [BBS20]. Goeken [Kri26]. Goertzel [SW05]. Golay [– DK15]. golden [MLM25, OR24]. Golub [BJNKR20, BEJR23, Meu09b, MS14a, Van25]. good [KM13, SW14]. Gordon [CZ23, CG20, LHZ20a, BF18, Bra06, DMA09, DCW23, Don13, FSY23, HKCW24, JJC24, LV18, RM11, ZHFW21, ZLM25, ZM26]. Goursat [HNSH09]. governed [ZZ18]. governing [KN23, LD20, YXL18, ZZ22b]. GPBi [TT21]. GPBiCGstab [HASI23]. GPIU [LZ15]. GPS [CX25]. GPU [CDP16, DEC24, MPS20]. GPU-acceleration [CDP16]. GPUs [BHS23, DM21]. grad [Ahu09]. grad- [Ahu09]. grade [CC18, GN12]. graded [GCZH24, LRY18, Maj14, RS25, ZZY24]. Gradient [CT21, DZ01, DLL04, LRGH02, MCG^{+}04, WSY04, AF23, AK19, ALV20, ABK22, ABKD23, AJMP11, And06, And08, And10, And14b, And15, And18b, ABM10, AV19, AK00, BKFMA11, BKG15, BC94, BGS24a, BF14, BE20, BE98, BF99b, BK08, CZL26, CMRS00a, CFL19, CL10b, Cui26, DW12, DW15a, DW15b, DMA19, DAN26, Don16, El 18, EE18, FH25, FHH96, GS14, God26, GO20, GWL20, HC24, JLJ22, JM93, Jón93, JRRS08, KPC20, KBA26, LLL18, LZ22c, LWLW24, LB25, Liu11, LWQR15, LL18, Liu21, LL22c, LSY^{+}23, LJY24, LT24, LTP18, MBG19, MJJ^{+}23, Meu97, Meu99, Meu05, Meu20, Meu23, MN23c, MG22, NH24, NA24, ODL21, OAR22, PPPN23, PRK^{+}18, Pla99, SKK25, SV26, SS06, SD25, SD26, SM10, SBJC19, SLL22, SLL23, TTXZ23, WCB15, WZ16, Wan18, WYP23, WL24, WHS20, YCL17]. gradient [YHS18, YJJ^{+}21, YSXY19, YWS20, YYZ22, Yue25, Zha09, ZW12b, ZCGS24, ZWG18, ZLL21b, ZDSY20, ZCTD24, ZJSS25]. gradient-based [HC24, SD25, SD26]. gradient-CQ [KPC20]. gradient-type [ALV20]. Gradients [GF02, CS94, GKQ24, JVH15, MT13, MT14, NKA24, USM24]. gradual [Dax17]. Graf [MW16]. Gram [MPR22, Sal05]. graph [BS25, BKM25, JCF15, LKKM15, YAT20]. graphs [CL96b, DMR20, DMR21, GWL20, SCS18]. Gray [LWD23]. Grcar [Meu25, Meu26]. greatest [FGM19]. Greedy [HSZ03, Sch14, Sch26, CM25, DH26, FS23, LL26a, LLLD17, SW00]. green [FS21, Bic24, BG13, PJ22]. Grid [CLGS17, KMS05, Bou17, CLWH20, CZZ^{+}25, CV22, DZ19, DMW23, GS26, HL06, HL15a, HFZ19, Hu22, LD21, LYL15, LW17, LLC20, LS20, MDH16, NAHZ21, PSW11, QXGZ20, SW10b, SSYL20, SL21b, Str05, UL18, WD23, YZ21, YZ23, Yse99, ZDZ26, ZYW22, ZXL23, ZD18]. Grids [GGV02, MM04, AHP25, BBC25, CJKL23, CYWH25, DDG05, GSZ22, GG98, GL21, HP18a, JK19, KR07, LPRÁ24, Spr98, Spr01, WQ10]. Gross [CCJC18]. groundwater [ASZ23]. groundwater-surface [ASZ23]. Group [CO03, DL03, CO19, Dam08, LZ22b, PPV09]. groups [VW08]. groupwise [Aih17]. GRPIA [MEJS19]. Grüneisen [Pas99]. Grünwald [CSZ22, LWJ21]. Grünwald-difference [LWJ21]. GS [mTLbJlL14]. GSAV [HKCW24]. GSOR [HZ15]. GSOR-like [HZ15]. GSVD [DNR15, HLJ25, HJ25]. Guaranteed [KW04, FLMR99, PV22b, PR14, USM24, ZW12b]. guarantees [SI13]. Guide [Hof16]. guided [Gau07]. Gummel [BR07a]. gyrocenter [ZLQT19].

H [Meu09b, Cal20]. Haar [AAU26, All18, PSS10]. Hadamard [CMM15, ER19, KM13, LGHS25, Mit11, OCWV24a, OCWV24b, RFS23, TCO26, ZZX^{+}23]. Hager [AWH^{+}24, Ovi22]. Hager-type [Ovi22]. half [KM24a, LMV24, LT20, ST26]. half-explicit [LT20]. half-range [LMV24]. half-step [KM24a]. Halley [AM16, BRW11, CCTV16, CKKT16, SS12a, Ste20, WGK11, WK13, WSK14, WK15, WK16, ZG09]. Halley-type [AM16]. Halpern [ATT21, HWCR19, TAM21]. Halpern-type [TAM21]. HAM [QLZX11, XCLA15]. HAM-based [XCLA15]. Hamilton [Haj16b, SS03, ZZ10]. Hamiltonian [ABI20, BCW13, Ben99a, BWC22, BCI14, BS14, BMR19b, BIMR19, BB14b, FT24, GWY26, HM18b, Jay21, LG19, LC19, MM99, McL25, MZ99, Sei98, WW14, XHZ07]. Hamiltonians [Haj16b]. Hammerstein [AAA24, DS26, GH09b, HVYMS23, HVMMV24, Mic23, NNCN23, SMN24, WWL24]. Hammerstein-type [HVYMS23]. Hand [MS01b, BEHS20, BF14, BDR25, DMD16, Hey01, HE05, KBCG13, TT21]. handle [Das26]. Handling [Md12, BKS13, CZ20]. Hankel [AH03, Bel08, BV95, Cam95, HR03a, LMV00, LV01, LXQ15, PSS10, Sas93, Wim00, XX16, ZuI19]. Hankel/Toeplitz [BV95, LMV00, LV01]. Hansen [KKB16]. hard [PSS22]. hard-sphere [PSS22]. Hardware [ACM04]. Harmonic [Che04, JLZZ23, BPR22, Dar99, DC17, HJ25, LAN18, LXZZ21, Rab92b, Rab92c, RS06, TM20]. harmonics [BBC25]. Harris [LL05]. Harrison [Pas06]. Harten [AD00]. having [ARTY20, BF93, LNS23, MZW20]. HBVMs [GWY26, BGI25]. HDG [LL22b, WPY25, ZWX19]. HDG-IEQ [WPY25]. Heat [Cro03, And14a, BQ19, BCJ22, BM24b, CK22, CGM25, FHC21, GD15b, HYT26, MP25, MS14b, RS06, SB21, SJW21, VMMD21, WXQ20, YSLH19, YXS22, YYLX23, ZJ14]. heat-diffusion [RS06]. heated [CC18]. heating [WXQ20]. Heaviside [AGN07, GGNF17]. Heaviside-type [GGNF17]. Hele [CLPY23, CJKL23]. Helmholtz [Ant18, AMKV96, BK04, Boy05, CMD19, DL21, GM20, JWY21, LWZ18, LL20b, MCW22, RT19, XZW13, Yal01, YJ21, ZLH22]. hemivariational [MZ19]. Henrici [Bel03, Rha22]. hereditary [TN10]. Herglotz [BM23]. Hermite [AH17, AHM21, AA12b, AAAGAD23, AT17, AAA17, AAH24, BC92, BL92, Bia94, BFK^{+}09, CF96, Cor24, FHS12, GM92b, Gau14, Her96, HLSLSL25, Iva15, KK22b, LMV24, Man17, MTTC22, Mer92, Mer94, MSS18, NB16, Plo93, Poc14, RKM25, REE26, Sab03, SST92, Sun94a, VB91, Wal07, YWZ19, dDL92]. Hermitian [CZL26, BM19, CM16, EGG08, Ema96, tFZyZ16, HM18b, LN10, PG15, Tan17, WL17, XM16]. Hessenberg [CDW95, FLT09, GL04, GLC22, Hey99, Hey01, Sad99, Sch09, Sol23]. Hessenberg-type [GLC22]. Hessian [AB99, CT21, FGS24, GGS22, JB22, LL07, Pan26]. Hessians [GF02, KD14]. Hestenes [DW12, TOY26]. Heston [iV12]. Heterogeneous [AA03, GHM16, GM20, GL15]. Heun [AGS08, FS16]. Heuristic [DR01, Hua96, DLC14, LL14]. Hexahedra [Mat01, Kee94]. hexahedral [SKJ^{+}18]. hierarchical [AGA24, Fan15, KAF18a, KAF18b]. hierarchically [Hom94]. hierarchy [Hom98a]. hierarchy-consistent [Hom98a]. High [AG03, AABM08, AEG02, AMM16, BOP98, Bla15, BR17, BC26, BGZ20, CCW21, CGL99, DR04, Gau00, Gau09a, Gau14, Hoa25, KM24a, LZZ24, LZQ26, MPR24, McL02, Rab92c, Sat24b, SMM25, SQG13, SS15, SG23, TZ24, VLCL16, XT16, AAII96, ABI22, AKT15, AKQ17, BCST14, BJ98, BBPV12, BKF20, BWC22, Bou17, BGL07, CC16b, CvPS15, Cui13, DNV25, EH97b, EH24a, FS20, FT14, FGR01, GH09a, GM20, GLLJ12, GZ20, Gau11a, GST21, GH23b, HCBAEC23, Iva24, JJC24, JHLL15, Joh15, KKV22, KNBGV18, Kaz24, KD14, KGH14, KSW07, LMV23, LBF24, LPRÁ24, LWC^{+}21, MS23a, MS06, Moh10, Mok16, Ngo23, OKB23, Pry98, RRZ21, RR22, Rou25, SFMK23, SK25, Sti18, Tsi07, TD09, WLL12, YD09, YZZL17, ZWFY19, ZYX19, ZHFW21, ZYQ^{+}21, ZLM25, ZM26, Sat24a]. high-accuracy [GST21]. high-degree [FS20]. High-dimensional [CCW21, JJC24, LWC^{+}21]. high-index [Pry98]. High-Order [AG03, McL02, BR17, Gau00, Gau09a, Hoa25, KM24a, LZQ26, Sat24b, TZ24, AAII96, ABI22, AKT15, AKQ17, BKF20, BWC22, Bou17, Cui13, EH97b, GH09a, GM20, GZ20, HCBAEC23, Iva24, JHLL15, KNBGV18, Kaz24, MS23a, Mok16, Ngo23, OKB23, RRZ21, RR22, Rou25, SFMK23, SK25, Sti18, YZZL17, ZWFY19, ZYX19, ZHFW21, ZYQ^{+}21, Sat24a]. high-performance [WLL12]. High-precision [BGZ20, Gau14, Gau11a, Joh15]. high-Reynolds [BGL07]. Higham [KZ03]. Higher [HHLM23, AMA21, BDL^{+}12, BN25, BM22, CM24a, DGP15, Gha18, God15, Han22, HM22a, HM22b, HM26, HKKN12, HLSLSL25, LM14, Loh22, MS13, PM05, PMM11, PP92, Rim25, San25, TY21, Tur94, ZLT25, ZAC26]. higher-index [Han22, HM26]. Higher-order [HHLM23, BN25, CM24a, HLSLSL25, MS13, PMM11, San25]. Highly [DHV22, HZX21, IP16, BMR19b, Cai22, CRHTV24, FT24, HJB18, HS15, JH26, Kub15, LW14, LZZ23, Maj13, OM26, Pen98, PM25, WW14, Wan22, WWM21, XX16, XLG22, ZH23, ZJ23]. HILBERT [AEF^{+}14, ATT21, AT21, Buo17, CCG01, CC12, CFM15, DHF21, GS95, KD18, NEY26, Pot19, QW08, RZ16, RT20, RTCL21, RTTH22, RZ23b, RT24, SCD^{+}21, SLD20, Tan20, TLD^{+}23, WZC26, YL19, Zas22, ZCTD24]. Hilbertian [Amo02]. Hilliard [ASGGRG23, BVV14, CLPY23, CLY25, LS20, WPY25, ZWWW20]. Hilliard-Darcy [CLPY23]. historical [KK22a]. history [MZ19, Osa12, SBG24]. history-dependent [MZ19]. histosplines [SBW98]. Hitchhiker [Hof16]. HIV [DJM^{+}18]. HIV-1 [DJM^{+}18]. HODLR3D [AGA24]. Hohenberg [Liu21, YK22]. Hölder [BRW11, CA07, CWL16, CR23, DMYT23, Hag25, HY21, ZCS14]. hole [CLMM05]. Holm [YYW21]. holomorphic [AK26c]. Homoclinic [Tov97, Tov98, BK97]. Homogeneous [LW25, AB24, BC00, CAB22]. Homogenization [BE03, CF05, Haj16b]. homographic [FJT94]. Homotopy [AS10, LG95, AJ13, BVV14, BZ18, DJM^{+}18, FY13, FY19, Van12, Van17, HSTW14, LL93, LM14, MV13, MV14, MSMS12, Mot14, NEMS14, NAA19, RW11, SMK14, Van19, WZQ17, WC10, WWD^{+}12, Yak95, YYL15, YXL18, ZL17]. Honour [AL04, BJK^{+}24]. Hopf [Cha14]. horizontal [AM18, AM13, Khe14, Khe16, LZL22, LZL23, MG20, MG21, SWS22, SW24a, WL25, ZV21]. Hormann [dC16a, AHP20, MNS23, TB25, dC16b]. Horner [BD04b, SW05]. Householder [ZCGS24]. hp [LCH20, ZWX22, MH23]. hp-error [ZWX22]. hp}-version [MH23]. HPM [RK11]. HSS [AGS20, CZ14, CC16a, CL13a, CM16, CWL16, CL21, HYW20, LG17, LG18, SM17, SB24, WC13, ZRZ11, ZYW21]. HSS-based [CM16, ZRZ11]. huge [DHV22]. hull [AHL20, DM92, Kub23]. hull-consistency [Kub23]. Hulls [KL04]. hungry [FYI^{+}12, WLS25]. Hurwicz [XY25]. Hurwitz [Joh15, Vep08]. Huxley [AZ19a, MDL15]. Hybrid [AH03, BPP23, BBB22, Bre02, DMRT03, EG10, HS20, HSS04, IY15, KAF18a, LL22a, MT04, ABK22, And08, And10, BKFMA11, BCCP25, BBO21, CR96b, CXL16, DW15a, DW15b, Dax09, Dey23, DHV22, Eba18, ETY98, EHVRV14, Fan15, FR18, GMZ19, GK20, GH09c, HMT17, HMA16, HSK20, JJ24, JLJ22, KXD25, LW16, LWLW24, LLL26, LYL25, LZZ24, LTP18, MBG19, MN17, MN11, NJ13, PSB26, PTW22, RKMS16, SBA26, SV26, SKD26, SvF94, THS20, WC23, WJW14, YJJ^{+}21, YYL26, ZDSY20]. hybrid-inertial [LLL26]. hybridizable [LW25, ZWX22]. Hybridization [PRK^{+}18, BKG15, PRVI20]. Hydrodynamic [TE03]. hydrothermal [BGRS12]. hyperbolic [AUA22, AC17, DZ13, FKP06, FGL19, Gar20, KL94, Lie00, MCDS24, NS22, Pan20]. Hypercomplex [RK11]. hypercubes [CL24b]. Hypergeometric [GKS04, AHM21, CFK^{+}20, DD99, DL01, GRAST23, NW17, Pas95, POP17, Pep23, Seg08, Sle25, Wil12, ZSGJ25, dDL92]. hypergraph [ZCGS24]. hyperparameters [Iid24]. hyperplane [AK15]. hyperrectangular [EG94]. hypersingular [ADG10, CCJ10]. hyperspectral [EJR26, RWTM21]. hypersphere [Gen12]. hypothesis [CRV91].

I.F.S. [Man07, Man10]. ICTM [dADdRC04]. ideals [Sau07]. Identifiability [DVJBN03]. Identification [LSZW19, Nac03, SMNZ20, ZAH26, GCFF95, NAR05a, PGGCGF11, SA23]. identifying [LYL25, XZZ19a, XZZ19b, YSLH19, YZL20, YPL21, YXS22]. identities [CW08, LWAG08]. identity [FGJ00, KL94, Szy06]. IEQ [CLY25, WPY25]. IEQ-FEMs [CLY25]. II [AHKW05, CM98, GM92a, GMT92, HJ18b, Leo07b, MC05c, Mil95, Now13, Rob92]. IIA [CX20]. III [Leo08, MRU91, MG94]. IIIA [BCT15]. IIM [ÁCL11]. Ill [CR02, RST03, AMR23, AC11, Ant18, BJNKR20, BH11, CMRS00b, DNR15, DNR17, FH00, Han96, Han94, HR14, JRS09, Jia20b, MRS06, MRS10, NRS12, OR17, Pla99, RRS09, RR13, RSZ20, Ria16, SGJ15]. Ill-Conditioned [RST03, BJNKR20, Ria16]. ill-conditioning [Ant18]. Ill-Posed [CR02, AMR23, AC11, BH11, CMRS00b, DNR15, DNR17, FH00, Han96, Han94, HR14, JRS09, Jia20b, MRS06, MRS10, NRS12, OR17, Pla99, RRS09, RR13, RSZ20, SGJ15]. illustrated [AR26]. ILU [BRY14]. Image [BP03, DHMS16a, NPP04, XTH07, AIA^{+}24, AR26, ALRT16, AMR15, BKPS93, BC05a, BH11, CKM^{+}25, CC06, CPZ14, CGYZ19, DHJJ10, DMT13, EHN17a, EHN17b, FAMA20, FFSC25, FM99, FLG08, HLS10, HS26, JLJ22, KBA26, LP08, LP13, LAG05, Li96, LZ22b, LSL25, LZ25, MP07, MMU20, MAF25, NR14, SKK25, SD25, SD26, SMA99, WYP23, WZW^{+}24, XQZ24, XYZ14, XW17, YAY26, ZY13a, ZL24, ZL26b, ZH26]. images [CTS09, EJR26, JMS16, LP12, Li95, SST08, WHZ^{+}18, dCOS21]. imaging [AH14, BEL23, CZX26, HOW95]. IMEX [Che22, Che24, GS19c, JZ25, Kaz24, SMR26, TTP26]. IMEX-Alikhanov-FEM [TTP26]. Immersed [KV04, Li97, JCL16, RW06]. immobile [QXGZ20]. Impact [Dum03, Moo07]. impedance [GHC15]. implement [TZH24]. Implementation [ACM04, CO03, Da 92, DHL^{+}04, DN19, DEM94, El 18, Enr02, Mel14, Nac03, PL04, SS99, Tur26, AG15, ACF99, AMM18, AEF^{+}14, BL23, BW13, BF00, BvLP16, BRZ96, BCI14, BC01b, CWZ13, Cam19, DEC24, Dum13, HM22a, HM22b, HJ18b, HCBAEC23, KK22b, Loh22, MH08, Nie93, Pol10, SLW13, SGO22, SGJ15, Yan17]. implementations [Han22, HMS96, HYJ20, OM18, SI13]. Implications [ME92]. Implicit [JU22, Jat15, KJ18, KSW09, Mar04a, SAHM25, AMM17, AMM18, BF17, BWC22, BD98, BC99, CSI18, CGPM00, CW19a, CJSZ14, CLA11, Che14, Che16b, CDLW21, Con93, Cui13, DWZ14, DMYT23, FHH05, HMT17, HZ93, HLTA16, IJ19, KCHD16, KLF17, KPFF24, KLB10, Le 98, LRL19, ME92, MC08, PXH25, PFT98, Pei24, Rim25, Sat24a, Sat24b, SL21a, SMM25, SYLT14, SZ99, TA26, TB19, VA20, WLMA21, WCZ25, WL22b, YYK24, YZLZ22, ZYQ^{+}21, ZV19, ZHZ25]. implicit-explicit [CJSZ14, IJ19, KLF17, YYK24]. Implicitly [Che04, DB98, LM99, FEL15, GSV96, Van19]. importance [DMR20, NR25b]. improve [KADE18, Ria16, YWYN22]. Improved [AH23, AB99, BM97, CGYZ19, DZ24, JJC24, Kol04a, LZ22a, MA16, Nac03, RA12, RCW22, Rum14, SI13, WDL16, AD17, AR09, BCST14, CKS24, CL21, CKKT16, DW15c, HJ18b, Khe16, Kim26, KGMH21, LP12, MSMS12, NH24, Odl00, TL24, WSK14, WK16, YHS18, ZYW17, ZZ24, ZW12a, ZFG18]. Improvement [Dra00, Gra03, Yal01, AHL20, MBJ17]. Improvements [CTY25, LPGL16, SCW17, VV25]. improves [EHVRV14]. Improving [ART14, EH97b, SW05]. improvised [FW25]. impulse [LZ22b, LZ25]. Impulses [DOS03b]. impulsively [Mot14]. IMR [CR14]. inaccurate [Mic93]. inclined [MSM12]. including [SPS26, Wal94]. inclusion [BC16, Cho16, HRAH22, HCD26, KAF18a, Kol06, Mil18, PM05, PMM11, RFS23, San19, SSS21, SDMMK18, SC18, TAM21, TG20, TDC21, Ton25]. inclusions [AG19, AABTB23, Dey23, OMU^{+}24, RIAA19, TRCL24]. incoherent [YP23]. Incomplete [EGSV04, GPGC98, GNT24, Mat04, WCLW16, ACSD16, CMM17, FJ25, MP00, SEG14, ST26, SS23b, YA24]. Incompressible [PG05, BRY14, BW13, DG17, Hei06, JZ25, JY23, KM17, LDX23, PV98, ZYJY22, ZS22]. inconsistent [ALL25, BF17, CM25, FHH96, HZ24, ZHSX23]. Incorporating [SBG24, DJM^{+}18]. increasing [Dax17]. Incremental [DL04, SW07, BG24, BE03, LXP20, WXT22, YWWR12]. indefinite [BHW23, BEGG91, CR96a, CWHL20, FHH96, FH00, Kel07, KPS14, MZZ25, MXL25, SR16, SGK^{+}99, WL17, ZD17, ZFG18, ZZZ22, ZLT25]. independent [LX23, LN22, Lor19, SBJC19]. Index [Ano01b, Ano02a, Ano02b, Ano04a, HL03, AWL^{+}24, Ano01a, AM98b, BCST14, Bea98, EL01, Fab16, GKRS22, Han22, HM22a, HM22b, HM26, HR07, ICR06, JV98, LM11, NSM20, Pry98]. index- [Bea98, EL01, ICR06, JV98]. index-1 [Fab16]. Index-2 [HL03]. indicator [ZZL17]. Indicators [BBHM03]. indices [EH24a, Le 98]. indirectly [YF22]. INDSCAL [lZqL25]. induced [Not95]. induction [AH12]. inductionless [XY25]. Industrial [The97]. Inequalities [Che16a, KM04, LR14, Alz08, AV19, ATT21, AH10, Bno21, BHT16, CSI16, CSI18, CRV91, Dax09, DLYH17, DHV22, Dra00, DMS09, Fan15, FY19, GL07, Gau09c, Gau09b, Gau11b, GZP18, GEA20, HT19, JM18b, Lei15, LL22a, MZ19, RTD^{+}21, RT22, SI17, SI18, SLD20, SK19, TQY21, TQC22, TCO26, TSI20, TLD^{+}23, UA09, WZZ16, WXT22, WZZ15, YL19, YLL20, YJS25, YAY26, Ye22, ZFC18, ZFH23, ZJSS25]. Inequality [AAAS03, CGR12, Anh19, ATT22, CSI17, CZ20, DMT22, DJG18, Gau08b, Kou07, LRL22, NPS09, OIM21, OAMA22, OR24, PPPN23, PN21, RTCL21, TH18a, TH18b, TH19a, TH19b, TVC20, TTLD20, TRSI23, TTA26, VS19, WL22a, WZC26, XCD23, XWL25, ZLL^{+}21a, ZWX19]. inertia [LZT25].

Inertial [CD25b, God26, HHF22, HWD25, OIM21, RTCL21, SCD^{+}21, Sha97, TH19a, WXT22, ALV20, BAK25, BC16, DCG25, Dey23, GZD24, HS20, Hie19, LLL26, LRL22, LSY^{+}23, LJY24, MJJ^{+}23, OMU^{+}24, Rob97, SIO20, TQY21, TQW24, TOY26, TVC20, TLD^{+}23, WZC26, YAY26, ZFC18]. inertial-double-projection-based [BAK25]. Inertial-type [WXT22]. Inexact [AG23b, BKM03, Gon16, HSS04, JCH23, PSS25, AG23a, AG19, AAN14, AGS20, AM18, BP22, BK08, CC15, DYW16, FGBP21, GO20, GZP18, HNY^{+}18, JWCZ21, KM17, Lei15, Lin16, LM13, LM25, MsC20, MC08, MN22, MN23c, Śmi09, Śmi13, UPY26, WZW^{+}24, WZZ15, YDWL15, ZZ24, ZZY18, ZYBJ23]. Inexactly [BN18]. Inexpensive [PV22b]. infeasibility [Md12]. Infeasible [IW04, Khe12b, Khe16, Khe17, KH20, KOK21, LYY12, LS07, LPXX19, MR09, SD20, YIY22, YZLP16, Yan17, Yan18, Ye22]. infeasible-interior-point [SD20, YZLP16]. infected [ZHZ25]. infiltration [SLLA15]. Infinite [BO03, HR03b, JC04, BIM^{+}23, Cat24a, Cat24b, CP95a, CL99, Cro92, Den14b, EO94, Esp05, FLMR99, FM93, HS96, HNSH09, KH11, LPXX19, Men26, PW22, Sid20a, SS12b, SH21b, SH22, SH23, TO21, WWD^{+}12, XCLA15, Zah09]. infinitely [LNS23]. Infinity [LRW25, SVZ01, GGJL24, KOK21, LPGL16]. inflection [PSWE23]. Influence [GHPMGRR14]. Information [MCG^{+}04, BP22, KP16, MN23c, Prz09, Rim25, WT08]. informed [BM24a]. Inherent [Wri02, MHA16]. Inheritance [BG03a]. inhomogeneous [DG17, ST21]. Initial [CN01, Mar04a, ADA07, AE09, AM98a, AHKW05, BKA19, Bac25, BAV18, BVV14, CEX14, Cor24, DK25, EL01, GO21, HM26, JL12, KS20, KK12, Li17, LYL25, LWS18, NBJA17, NJ13, NA26, OQ12, PGGCGF11, RKMS16, RSKB17, SL18, SW19, SS15, SS98, SK25, SA23, SZ99, Stu97, WZQ17, YSLH19, YZL20, YD26, vSv94, vdHM98]. initial-value [Bac25, Cor24, RKMS16, RSKB17]. Initialization [HL03, MMW20, Sch09]. initialize [MS11]. injective [Gon16]. injective-overdetermined [Gon16]. inner [DGP15, HHLM23, Ron92b, SBN25]. inner-product [SBN25]. Innovative [NM26, SSR25]. input [YZX25]. inputs [YYLX23, YCW^{+}19]. Instability [Car01]. instant [LMUZ19, YZH21]. integer [DGST15, GLM15, SSC25, XyJl16]. Integrability [DS23, MT98]. Integrable [FSY23, ZHT15, CHHL18]. Integral [DV01, FHS12, JC04, Med10, MO04, SM25a, WHL24, AS11, AR24, AAH20, ASS11, ASS13, AAD14, BB14a, Bic11, BD00, BG13, BZ24, Bru93, CQ16, Cai22, CCJ10, CIP10, CS99, CS25, CA22, CH11, CS08, DD20, DM98b, DS09b, DHWL25, Esp05, EHV19, FRS21, Gau08a, Gau22, GX19, GH09b, GKV23, GMT92, Gu20, HNSH09, HVYMS23, HVMMV24, HSTW14, HST15, IZ24, IDS16, JWY21, JH26, KMH24, KPD25, KM24b, KGH14, KSCS07, KSCS08, KKS22, KKD25, LDH23, LLC20, LM21, LWN13, LJbL21, LHW13, Lyn08, Maj14, MKA14, MV17, Mic23, MNS23, MMLM20, NS24, NP18, NLT21, NNCN23, NAA19, Nor00, OdZdRV13, OM26, Pas06, PP24, Pis16, RS06, RT19, SP21, TZH24, VdR13, WQL20, WHD22, WWL24, WLJ24, XL14, YY13, Zag24a, Zag24b, Zag26, ZA20]. integral [Zeg25a, Zeg25b, ZHW25, ZH23, ZLCW23, ZS19, dAFPR23]. integral-algebraic [Pis16]. integral-based [IDS16]. Integrals [HR03b, ADG10, CC07, Car95, CP95a, CBGVPP09, CS08, DS15, Gau11a, Gau15, GHM23, GHM24, HS96, HS15, KR23, KHM20, Mac96, MM00, Maj13, Maj18, Mel10, MZ99, Pas99, Prz09, Saf10, SR16, SS12b, SH21b, SH22, SH23, XLG22, ZA24, ZuI19, ZH17]. integrand [KHM20]. integrands [CvPS15, CBGVN07, XLG22]. integrated [BPS23, NPS09]. integrating [Li95, SZ24]. Integration [Coo03, Dab04, DL03, PW04, Sch02, ADG10, AL23, BuR24, BLS92, BL98, Bou17, CM15, DDS93, Dam08, DKL15, Dze15, EO94, EG94, FHL21, Gau95, Gau12a, Gau13a, GG98, God15, HW00, HS96, HKKN12, HS15, HFW^{+}21, Kel07, LS11, Mal21, MKBY19, Man07, MFPG07, MO19, MCMX20, Pet95, PW14, Rab92a, ROB18, SP21, SGK^{+}99, Tov97, Tov98]. integrator [CM24a, Don13, JJC24, LZ23a, PM25, SAE19, Sha19]. integrators [AM98b, BC05b, CZ23, CO19, HK06, LWS18, MO10, Miy24, Sat24a, Sat24b, SW19, SSH^{+}19a, WW19, Wan22]. integro [DOT21, CA07, Che22, CZZ^{+}25, CPS12, DSI11, DS26, EED19, FH15, GHZ25, HCXL20, KSCS07, KKS24, LLY25, LA22b, MH21, MH23, MN23a, MKA14, MWDD24, MG11b, PS09, PS17, RT12, RM26, SAC18, San25, SMN24, WZL26, ZXF14, ZFX14]. integro-differential [DOT21, CA07, Che22, CZZ^{+}25, CPS12, DSI11, DS26, EED19, FH15, GHZ25, HCXL20, KSCS07, KKS24, LLY25, LA22b, MH21, MH23, MKA14, MWDD24, PS09, PS17, RT12, RM26, SAC18, SMN24, ZXF14, ZFX14]. integro-partial [MN23a, San25]. integrodifference [PR25]. integrodifferential [QQX23, WQZH24]. Intelligence [Ano95b]. Interaction [Car01, CG20, GA20]. intercomparison [BBC25]. interconnecting [TFPG19]. interest [LGL23]. Interface [KV04, AL23, BNN16, BV99, CFL19, CLWH20, DMW23, HHST19, JK19, Li97, LPRÁ24, RW06, WDL23, WG25]. Interior [IW04, AG15, AJMP11, AM18, AM13, BBO21, CWZ13, DGL06, Fra26, JKC26, Khe12a, Khe12b, Khe14, Khe16, Khe17, KH20, KOK21, LZ18a, Lin16, LYY12, LS07, LW25, MR09, MN11, PH14, RS20, SMZMA18, SD20, SS94, WO00, WZ11, WZZQ24, WGZ25, YIY22, YZLP16, Yan17, Yan18, Yan22, ZW20]. Interior-Point [IW04, AG15, AJMP11, AM13, CWZ13, DGL06, Khe12a, Khe12b, Khe14, Khe16, KH20, KOK21, LYY12, LS07, MR09, PH14, SMZMA18, WZ11, WGZ25, YIY22, Yan17, Yan18, Yan22]. Interlacing [DJM08, KG23, Seg08, KJG23]. intermediate [CSI17, QLZX11, Sid94]. intermittent [DMZ20]. internal [LPRÁ24, MN92]. Internality [DDRS23]. international [BV96, Ano92, Ano95b]. interpolant [ASS13, AGN07, KMRAR25, Mon96, RB21, dC16a, dC16b]. Interpolants [NSS03, AHP20, AH21, ASS11, AC94b, Ber00, BCGVS11, CF96, CM92, EM10, Gen12, LC14, MTTC22, Mer92, Mer94, TF00, Tsi07, Ver10]. Interpolated [WZZ26]. interpolating [ALY22, BM00, BGVHN92a, CDF99, Cox93, DDS93, Ila20]. Interpolation [All03, Amo02, CG14, CG17, Gen12, HM06, LM01, LWC^{+}21, Mat01, OS25, PS06, WDY04, Wen92, AD22, AL18, ACM93, AAAGAD23, Bal11, BDIR18, BC92, BF18, BM00, Ber07, Ber11, BV21, BESC22, BDS00, DOT21, BZ24, BLS06, Bra96, Bro05, CDSV11, Car91, Car92, CG05, Car10, CKP19, CKP22, CM24b, Cau22, CAV23, CXC25, CMMP19, CT21, Cor91, CL13b, CDD13, CGN22, CY10, DF93, DHWL25, DZH23, EG19, EEM20, FM16, FP20, Gal93, GZ18, GM92b, GM92a, Gás99, Gem97, GI15, GLV05, GP05, Hof21, HvD93, HAN24, HM07, IZ24, Iva15, IN21, JKK^{+}08, KK00, KST21a, KST21b, KK22b, LW20, LAG05, LN95, LX23, lLLVZ17, MP08, Man17, MMV17, ME95, Maz18, MT18, MS17, MSS18, MEJS19, MS23b, MNS23, MG91, MG94, MS96]. interpolation [Nar05b, NKM25, NCC11, NL97, OKB23, PR10, PZL15, PS22, Plo93, Plo94, Pre93, PQ95, REE26, Rip93, SST92, SX96, Sid07, Spr98, Str97, Sun94a, SWG20, SG23, Tan20, TB25, Tia21, Tra93, UL18, VB92, Van07, WQ10, XZL12, ZZL17, Zha95, dB07a, dS00, dDS00, dBD05, dC22, ddS24]. interpolation-based [AD22]. interpolation-extrapolation [BESC22, EG19]. interpolation-type [lLLVZ17]. interpolations [AA12b, Bic11, JCL16]. Interpolatory [BLS92, BSL18, CH95, DL08, FMD18, Not08]. Interpretation [NG23, BCM10]. Intersection [WZZ07a, WZZ07b, BM06, CKO24, PN21, SKJ^{+}18]. Interval [ASS04, BL04, Cha04, Cse04, DL04, Gla04, Kol04a, KPFG04, KL04, Mar04a, MJ18, MJ20, MCG^{+}04, MKO04, NKS04, Pop04, RDdRC04, Rev03, VR04, ZGLH24, dADdRC04, vGK04, BBZ95, Ben99b, Ber07, CC07, CKP19, CP95a, Eft15a, Eft15b, Gla01, HM07, Ibr25, Joh20, Kub15, MJH17, Pop18b, Pop21, Rum12, SM21, SH21a, SH21b, SH22, TV25, UPS25, Val14, Val15]. interval-valued [UPS25]. Intervals [Kol04b, vGK04, Maj13, Mül00]. Intrinsic [ASS04, GL23]. introduced [Dur93]. introduction [CFM15, JJ24, MSCB93]. intrusion [SMNZ20, SFZ22]. invariant [AD22, BEH24, CKO24, Dam08, DBV23, ED13, JJK97, LWG18, LL20c, LL22c, SBR26, Yue25]. invariants [Asc97, BS14, Sat24a, Sat24b]. Inventory [Cha04]. Inverse [CKP22, FG03a, LN10, MS11, NR12, ZL25, ABG97, ADN17, BDV18, BI00, BAS25, BZS22, CM24b, CR20, CDD21, DS09a, EP97, FP18, GWBC20, GHC15, GGJL24, GD15b, GK20, GP05, HLN25, HN94, Hua21, JLX22, JS23, JN99, KZS21, Kuh13, KLW^{+}23, Li95, LPGL16, LCVL18, LRW25, LA22a, LZ16b, LDC10, LL20b, LZX22, LGM25, LS25b, MGL20, MsC20, Meu09a, MT93, MH22, PGSU25, PLH20, SBN25, SHLY18, Sol15, XHZ07, YLY12, ZJZ20]. inverses [AG25, APST21, BM94, BD17b, GLS26, JXX^{+}23, LMUZ19]. inversion [CAB22, CUY24a, CUY24b, DdAF^{+}20, DW15c, FM99, GST17, JLMP16, KSV23, WGZ18b, YD26]. invertible [LRT19]. investigating [Pat25]. investigation [Die08]. Investigations [ZWFY19, HHST19]. invex [Jos22]. inviscid [GB21]. involution [LYL25]. involve [Sid20a]. involving [AV19, Ash16, AKT15, HS96, Jos22, KXD25, LQ20, MP08, MMV19, Pep23, SKD26, SKM25, ZuI19]. IPG [LY18]. IR [GHN19]. IRKS [MH08]. IRM [DS26]. irreducible [wYN18]. Irregular [MM04, Cal20, EAGS20, Lee94, SH12, ZYLN18]. Irregularity [Ros97]. Isogeometric [HZ20, Moo20, MRS25]. isometry [BOW21]. Isotropic [Hem96, HMdÁES08, Meu12, MP13, WWD^{+}12]. issue [Ano17]. Issues [But02c, CDG23, KM09]. Itô [AAH20, ES19, HH11, MNS23, Prz09, TX19, YBK^{+}21, Zeg25a, Zeg25b, ZHW25]. Italy [BT14]. Iterate [JS23]. iterated [Alt21, BOR23, BGR25, CFM15, Don12, GLM15, HRY16, JMS16, KR23, KR20, MMU20, RP26, Sab92a, SWG20, TL24, YXS22, vSv94]. iterates [Bog13]. iterating [Mor25]. Iteration [AG19, ABQ04, BCV03, CL11, CPV04, Rev03, ZL22a, ZL25, Axe99, AFN16, AFN17, BBS20, BBC11, BZ13, BH22, Bel94, BEQOR14, BPR21, BHT16, BK08, CR96a, CW19a, CW21a, CL13a, CZM21, CHYH24, Chk20, CKS16, DW24, DS20, DLL^{+}24, Don10, DG94, EH24b, Fan22, FGJ00, Le 19, GA15, GKL21, GNH10, GW24, GLL19, HWCR19, KD14, KMZ18, Ke21, Lei15, LYW14, LY17, LXX23, LLD23, LZ16a, LZ18b, LZL22, LZL23, LM25, MM17, Mor17, NUG25, OC24, PDRG19, RWTW19, SPV20, SAE19, SS21, SB24, SCS18, SSH19b, SSH20, TLD22, TDWL24, WT08, WZ13a, WGZ18a, WL25, WCZ25, WHS23, WL17, WPL18, XW18, XWY19, XZP^{+}20, YAT20, Yan18, ZZ17, Zha11, ZYW21, ZY13b, ZM16, ZLV17, ZV19, ZV21, ZZZ22, ZZLV23, lZqL25]. iteration-based [SAE19]. iteration-inexact [BK08]. Iterations [GF02, AMM18, Bag00, BM06, CGV22, CVLX19, CL19, CFS21, Elf25, GM23, Hig97, KL94, NSM20, PM22, Pié99, SS98, TM14, ZG09]. Iterative [AAA^{+}18, Bog02, BKS23, Buo17, CSI16, Cic20, CGL01, DF93, De 02, GP05, Jay02, LP18, LDZ24, Nac99, Nac03, NPP04, RR00, SI18, SIO20, SS16, SC18, TRCL24, YP09, AAFL23, AUA22, ARS26, AABM08, AVCO26, Ano17, AG13, AS14, Awa10, ANA14, BPP23, Bai97a, Bai97b, BGR23b, BHS14, BH17, BE17, Bre00a, BE98, BM24b, BDR25, CSI17, CZ23, CR96b, CGPM00, Cas17, CPP14, CCL18, CKY99, CSC26, CRN19, CHMT10a, CRHTTN24, DD20, Den14a, Ern00, EGSHVN15, Fan15, FT14, FAM25a, FAM25b, GA08, GHN19, GRT97, GEA20, HT16, HT21, Hai08, HJ18b, HVM15, HVMT17, HR07, HSE16, Hom94, Hom98a, jHyPlZ06, HM18b, HM18a, HM19a, HWXC19, HC24, HM14, Iva17, Jia20b, KSV23, KKR25, KS25, KCBT21, KT07, KSB08, KAF18a, KAF18b, KS26, KGN^{+}24, LMMH11, LZZ19]. iterative [LF19, LSG15, MN17, MM19, MGG25, NHP06, NRS12, NC94, NK16, NK21, PNW17, PS16, QAS^{+}24, RFS23, RIAA19, RT22, RAH11b, RA12, RR98, SKK21, SKSS21, SRK^{+}25, SIE16, SCF23, SD26, SGJ15, SDMMK18, Sol15, SZX11, TAM21, TPY14, UTO24, USAF14, WO00, WZ15a, WZQT15, WDL16, Wan22, XY25, XL24, XZZ19b, XCY21, XSL22, YWWR12, YZL20, ZW14, ZY21b, ZH19, ZAC26]. ITR [CR14]. IVP [SGK^{+}99, Cor02, MN92]. IVPs [BGI25, BC01b, Eba18, KP16, Kac18, Sho18, SKTGR19, CMR03, SFT03].

Jacobi [Gau11b, Gau17a, Gau19, Bhr16, DB06, DBAE09, DMS09, DJM08, DJ18, EK94, EKM03, GL07, Gau08b, Gau09c, Gau09a, Gau09b, Gau12b, Gau18, GST21, Haj16b, Har18, HD18, HJ25, KMH24, KG23, KJG23, KM24b, Kou07, KH18, Leo07a, LGP11, LI10, LR14, NLT21, PP24, Pre93, RM26, Saa23, SS03, SSN^{+}12, SH23, VdR13, YH21, YWYN22, ZA20, ZZ10]. Jacobi-type [SH23]. Jacobian [CW19b, CRHTTN24, LG17, MCDS24, WGZ18a, XLW20, ZCG15]. Jacobian-free [CRHTTN24, MCDS24]. Jagerman [GPGVS92]. Jain [AA16a]. Jaiswal [AAN25]. Jarratt [AABM17, CHMT10b, RWB09, Sol11, WKG11, WK12]. Jarratt-like [AABM17]. Jaynes [Kar07]. Jean [AL04]. Jeffrey [NZF11]. Jitter [Dab04, CHH93]. JMP [ZLL^{+}21a]. John [BJK^{+}24]. Johnson [EN11, Kri26]. joining [BESC22]. Joint [Ano95b, CB13, HKP25, HLJ25, JL21]. Jordan [MGL20]. Joseph [DMW23, WDL23]. Josephy [AH10]. journey [CR00]. Julia [KR23, KA26, KGN^{+}24, SSH19b]. jump [Che22, Che24, RW06, YYK24, YW17, YCZ26]. jump-adapted [YW17]. jump-diffusion [Che22, Che24, YYK24, YCZ26]. jump-extended [YW17]. jumps [BT23, LGL23, OKP21, SPB26, SZ21, ZWW21]. just [Wri95].

Kaczmarz [PP17, Pop19, ALL25, CH22, CM25, DH26, EN11, FAM25a, FAM25b, HDL23, HWD25, HZ24, gK25, LL26a, LHZ20b, NYL26, PP16, Pop18a, RN21, TL23, Wu22, ZHSX23, ZWZ25]. Kaczmarz-type [Pop19, Pop18a]. Kahan [BJNKR20, BEJR23]. Kalman [JS23]. Kansa [Kar10, KJC18]. Kansa-type [Kar10]. Kantorovich [AH10, GTA19, LWN13, LMMD05, UA09]. Kantorovich-type [AH10]. Kármán [YXL18]. karst [LCW20]. Kawahara [CMWP20, Wan25]. Kawasaki [LZL20]. KdV [LKQ23, SW14]. Keeler [VA20]. Keller [KKA17]. Kelvin [KBP17, Pan18, ZD21]. Kernel [BLS06, FH00, KST21b, LRM16, MO04, AR13, AF13, BAV18, BRZS23, CZZ^{+}25, DR07, HSY23, Khe12a, KH20, LX23, LZ09, LYY12, Maj14, MS23c, NS24, Sal25, SN24, Tan20, Tru24, WZ11, WZ15b, WCD21, KST21a]. Kernel-based [KST21b, DR07, KST21a]. kernel-independent [LX23]. kernels [ASS11, Ave20, Cai22, CFM15, GZ20, HS96, JH26, KSCS08, Lev95, Lig93, ML20, MH23, OM26, SP21, San25, TA96, WHD22, Xu19, ZXF14]. Kerr [JLFL19]. Kerr-type [JLFL19]. key [FEL15]. kinase [MB09]. kind [AR24, AZ19a, AAD14, Cai22, CS25, DS26, Den14a, FNS19, HVMT17, HSTW14, Hun95, IR13, JXZ26, LM21, ML20, MH21, NS24, NLT21, PP24, PP18, REM21, SP21, WLJ24, YZ11, dAFPR23]. Kinds [MC05b, MC05c, DDRS23, GMZ19, ME95]. kinetic [OKP21]. kinetics [Fly22]. Kirchhoff [JZYY23, KKPT16, SKP20]. KKT [ZL07]. Klein [CZ23, CG20, DMA09, Don13, LHZ20a, LV18, ZHFW21]. knock [Rou25]. knock-out [Rou25]. Knots [CMRS01]. knowledge [Bea96, TRCL24]. known [AAAA^{+}18]. Kober [PS17]. Kogbetliantz [NS22, Nov26]. Kogbetliantz-type [NS22]. Kohonen [Hof05]. Koiter [PPS25]. Kolmogorov [Ila20]. Koornwinder [KW96]. Korobov [DKL15]. Korpelevich [TLR^{+}25]. Korteweg [TA24, YZZL17, YZLZ22]. KP [AR26]. KPP [Bou06]. Krasnoselski [KAF18b]. Kravchuk [ADGP15]. Kriging [AD22]. Kronecker [CK20, Cor91, FT05b, MMV19, YHZL21]. Kronrod [Bou17, DPS18, Lau07, Mon01, Not95, OPSM22]. Krylov [Aih17, AFN16, AFN17, BBS20, BM19, BLW09, BPS23, BEHS20, BMS24, Bre99a, BPR21, CH26, Cas17, CK24, EJR02, ED22, HE05, IS17, Jaw22, Jbi03, JL15, Jia20b, Lam09, LC19, LLWC24, Lun23, MB09, Mia19, MR02, PK21, PV98, RR00, SBN25, Sch17, SWB08, SWJ24, TY96, ZS08]. Krylov-like [Mia19]. Krylov-type [IS17]. Kulisch [AL04]. Kuramoto [MK17]. Kurchatov [EGGSH13, HVYMS23]. Kurchatov-type [HVYMS23]. Kutta [AMCM06, Con93, FLY25, Jat15, LG19, PFT98, SQG13, VH12, AH15, AMM17, AMM18, BJ98, BS14, BB14b, BD98, BC99, BJ04, CGPM00, CMR03, CT10, Che14, CQLY15, Che16b, CDI14, DJ10, EH97a, FYM14, FYYW19, FHL21, GCJ26, JT96, Kha14, KM19, KFK^{+}24, Kri26, Li17, LT20, ME92, MV02, MC08, McL25, MKS18, MHA16, ST21, SZ24, TX19, Tir02, TCW14, Tsi07, TÖ17, VV11, Ver10, Ver14, Ver24, WG13, WC25, Wri02, XT16, ZYQ^{+}21, Zha24, dS24, vSv94]. Kutta-Nyström [CQLY15]. Ky [Wat93].

L2 [LCZZ23]. L2-1 [LCZZ23]. Lacunary [Mas05]. lag [SS15, SKTGR19, WXQ20]. lagoon [CDF99]. Lagrange [BZ24, CF96, DP21, DHWL25, EEM20, GL15, IZ24, JKK^{+}08, KMRAR25, KS26, Kun01, LC14, MS23b, MG91, Pre93, SX96, SLT20, ddS24]. Lagrangian [AHS22a, AHS22b, BW13, HSS04, LGZ25, SS03, TYSY20, Wit96, ZXL25]. Laguerre [Cam19, CRV91, DHM12, Gau09a, Gau14, Ghe13, KJG23, LMV25, jLyLqW17, MM08b, MM09a, QW08, Ter22, XyJl16, YWZ19, YYW21]. Lambert [Gau11a, Joh20]. laminar [NM14]. Lanczos [BRS92, Bag00, BR06, BPP23, BEJ20, BEGG91, BRS91, BRZ98, CR99, CR02, CR12, Cdv98, CPRZV23, DB98, FGP91, GGV96, GV99, GLG26, GM97, GSV96, HLJ25, JNS19, JLZZ23, PP21, Soo15, Van02, Vos00, XC25, Ye96]. Lanczos-Based [CR02]. Lanczos-Type [Van02, BRZ98, Cdv98, CPRZV23]. Landau [HP18b, HWZ24, WS24a]. landscape [DMZ20]. Landweber [Ceg25, Erb15, YZL20, YPL21]. Landweber-type [Ceg25]. Lane [ITA24, SSR25, SPS26]. Langemeyer [Har20, Mat15]. Langevin [WC24]. language [MdR13]. Laplace [Ant22, CAB22, DW15c, HV15, KH11, Kuh13, LM14, ML10, NPR08, WGZ18b, WQL20, WF23, ZAH26, ZW12a, ZLLH22, ZL25, DCM^{+}13, DCMM13]. Laplace-space [Kuh13]. Laplacian [ALZ21, JP14, LWD23, ZLCW23]. Large [EJR02, GSA03, HL03, HSZ03, Jbi03, LWwCL13, AAB13, AN17, Ali23, AKB15, And22, Ano17, AGS20, AG23b, Bai97b, BLW09, BAK25, BKW25, BPS23, BG91b, BEH24, CKO24, CDT10, CR99, CMRS00b, CH22, CM25, CC13, CL96b, CDP16, ED22, Ema96, FZ26, FGS24, GGN14, GHN19, GH09c, Guo16, GZL^{+}25, HJ18a, HPS97, HRY19, HJ21, HZ24, HJ25, JVH15, JNS19, JML25, KW00, Kim26, La 17, Lei15, LZ21, fLxX12, LL18, Md12, OR17, Pan96, RRS09, RR98, RR00, Rog95, RALJ26, SBN25, SS99, SW00, ST22, Sti18, VBG96, WCLW16, WYP23, Wu22, YCL17, YXL18, ZTW19, Zha20, ZQS24, ZW20, ZZLV23, ZS08]. Large-Scale [GSA03, LWwCL13, AAB13, AN17, Ali23, AKB15, And22, AG23b, BLW09, BAK25, BPS23, BEH24, CH22, CC13, FZ26, FGS24, GHN19, Guo16, HJ18a, HRY19, JNS19, JML25, Kim26, La 17, LZ21, LL18, Md12, RRS09, RR00, Rog95, SBN25, ST22, Sti18, WYP23, YCL17, Zha20, ZW20]. larger [Ma20]. largest [BG91b, GL07, Gau08b, HM18a, wYN18]. laser [WXQ20]. lasso [JFSD25]. lateral [BCJ22]. lattice [BDL^{+}12, WLS25, FEK^{+}23]. lattices [CG05, CG07a, JKK^{+}07, JKK^{+}08, LSX10]. Laurent [AAPR21, BHJTM92, CGV92, GCGVH92, MC05b, MC05c, Pow93, Riz18, VGM96]. Laurent-type [AAPR21]. Lavrentiev [SGJ15]. law [Wan25]. laws [AC17, JB24, MCDS24, Sar06]. Lawson [JJC24]. Lawson-type [JJC24]. Lax [ZZ10]. Layer [Boy05, SC03a, ADA07, CYM22, DL21, FY13, FHAL15, Ghe13, Ghe16, KR11, KK16, LPRÁ24, Lin05, Mot14, NZF11, NM14, NMM18, RS20, TPLB22]. layer-adapted [CYM22, Lin05]. layers [ALZ21, JKC26, LZ18c, LZ23c, MN11]. LBMAT [FEK^{+}23]. LCCO [AG15]. LCP [CKS16, LYY12, PPR15]. LDG [ADP26, WJC25, ZLW^{+}13]. LDL [Orb15]. Lead [MKO04]. leading [MD15]. leap [HMT17]. Learning [CMMP19, BS25, BWB24, TOY26]. Least [BCV03, DG05, Ell93, Hei06, LMV00, LV01, MMV19, PG05, SK04, AAH20, Ari98, AAD14, BGS24a, BRS09, CM24b, CX25, CR20, CL24a, Dax09, DMA19, DWX17, FGS24, HS26, Han22, HM22a, HM22b, HYW20, Hei07, HZ95, jHyPlZ06, HM18b, HM19a, KC23, LDN16, Lei15, LEK21, LL22a, LJ22, MMGH17, MRS93, MZW20, MZZ25, MXL25, MS24b, NC94, NK21, NA24, OC24, OLB94, PT24, Pen13, PS25, RAH11b, RALJ26, Rum14, SEG14, ST18, SM21, SMN24, SXHZ20, TV19, VZ93, VPA25, Wu22, XXW17, YCW^{+}19, Yua21, ZH22, ZLWZ21, ZW22, ZZ22a, ZLZ23, ZHW25, ZLWQ09, ZMZ26, ZWW25]. least-change [Ari98, VPA25]. least-norm [PT24]. Least-Squares [DG05, PG05, SK04, Hei06, CR20, CL24a, Han22, HM22a, HM22b, Hei07, PT24, RALJ26, ST18, Wu22, ZWW25]. Lebesgue [BEL23, CYIB12, dC22]. leech [AB98]. Leffler [WZ23a]. Left [MS01b]. Left-Hand [MS01b]. leg [KS06]. Legendre [Ahu09, BZV16, BK04, BK18b, CQ16, Cai22, EDAM13, ES19, GK20, HIK17, Hun95, ITA24, Kha25, KSCS07, LW22, Luc97, MH21, MG11b, Pan26, ROB18, SHF15, SKM25, SPS26, SH26, YYW21, ZJWF18, ZJZ20]. Legendre-based [ES19]. Leja [Chk20]. Lemaître [Le 19]. lemma [BS21, CW08, EN11]. Lemniscates [PMO05, Ple03]. length [BKA19, CKP19, FRR07, NR24b]. lens [RKM25]. Leopardi [Kou07]. Lerch [DN24a]. Leslie [MWZL23]. less [And22, BESC22, EG19, EV22]. Letnikov [CSZ22]. Level [FAMA20, ARSS19, BMV09, DW22, GLV05, Kim26, LAG05, MBR21, MT12, NG23, PL25, Rab92c, SMM25, ST99, Wan18, XL24, ZZH15, ZLT25, vdMRS06]. level-bundle [PL25]. levelling [AE18]. Levenberg [SG17, SG18]. Leverrier [PS06]. Levin [Hom92, Osa92a]. Levin-type [Hom92]. Levinson [AA16b]. Lévy [JS21]. Li [MKA26]. Liao [AK19, ABK22, And18b, BKDA25, KBA23]. library [BWB24, CDW95, Das26, MHR23, OBAHK^{+}19, RZ99b, Tou98]. Libration [Lo97]. Lidstone [CDD13]. Lie [CO03, CO19]. life [Bre06a]. lifting [AABTB23, BCL00, Laz99]. Light [Lev05]. Like [CPV04, AIA^{+}24, AABM17, Arg10, AG17, BD09, Bar13, BKW25, CCV23, CHYH24, CW17, ED22, Fas23, FGC19, Gon16, GLdO09, Guo13, HRAH22, HZ15, LQ20, LC21, LS25b, MA16, Man21, Mia19, MA22, PT24, PP06, PT17, Pol10, SSS14, Sal05, SHLY18, SSH20, Tan17, TS92, TZ25, Wan15a, XZL12]. Limit [Thr92, AHS22b, AHS22a, AH13, CZLS18, Lor95, MKA26]. Limited [FH25, Orb15, AB99, AAAGAD23, Liu14, SLL23, VL19, WZ23b]. Limited-memory [Orb15, VL19, WZ23b]. Limits [ST92]. Lindenstrauss [EN11]. line [AHR21, AAN14, AK15, BKR18, CBGVPP09, DW12, Deh20, DMR20, DMR21, FZ26, GS21, GZ11, HWC15, KJ15, LZ18a, LLL18, LM13, LC21, MP00, NKA24, Ovi22, SW11, SM10, TH19a, WZ16, YLL22]. line-search [TH19a]. Linear [AZMJ04, AAAS03, ABQ04, BBD03, BCN^{+}16, Bre02, BZ02, BJ04, CGV22, DDP14, De 02, DL03, GL04, GC04, GM03, IW04, INR01, Kol04b, LJW17, Pop04, Prz09, RST03, SGM02, SK04, Wat92, Wat94, Wri02, dSCS04, AEH20, AH11a, AHIJ22, AHJ22, AHP25, AP26, AR16, AR20, ABS19, ALJLYJ24, ABG97, AL97, AR18, ALRÁY25, ASGJ^{+}20, AAAGAD23, ASHF21, AR10, AC19, AM13, AK26c, AKT15, ANA14, BS25, BMA16, Bag00, BD06, BBC11, BCW13, BZ13, BH22, BC94, Bar91, BBO21, BOW21, BVV14, BC06, Ben99b, BGS24a, BBP17, Bla15, BMS24, BM96, BRZ98, BRS08, BDR25, BV09, BH09, CR96a, CR96b, CMRS00b, CJSZ14, CF96, CL00, CMR16, Ceg25, CS94, Cdv98, CC16a, CL05, CLWV15, CZ20, CK20, CLWH20, CH22, CLPY23, CM25, CDW95, CCD10, Cor24]. linear [CPS12, CHK14, DEP12, DL08, DLL12b, DLL12a, DLL13, DEC24, DCG25, Dax09, DS20, DMD16, DZ13, DLL^{+}24, Don10, DP16, Dum13, DHMS16b, Dze15, EHTSM21, Elf25, Ess98, FGC19, FP18, FGL19, FM93, FHH96, FT05b, FMD23, FGS24, FS23, GWL18, GEP14, GEP16, GEP19, GP99, Gau13c, Van12, GHP^{+}00, GM97, GZL^{+}25, HT16, Hag25, Han22, HM26, HYW20, He16, HHF22, HWD25, HDP18, HR05, HSTW14, Hey01, HE05, HSE16, Hil10, HKPW19, HASI23, HL23a, HL23b, HCH18, HKCW24, HWC25, HLV26, HLZ14, HZ24, IKMS24, JRB17, Jbi93, JRS09, JJ13, Jia20b, JW25, JML25, JS23, KBCG13, KKR25, gK25, KST06, Kaz24, KMZ18, Ke21, KPFF24, Khe14, Khe16, KKO17, KLL25, KS18b, KGN^{+}24, LLAL21, Lei15, LLS11, LYW14, LZ14, LL16, LDL17, LM19, LC19, LL20a, LYH^{+}20, LZOY22]. linear [LLY22, LL22a, LZ16a, LZ19a, LA22a, LZ19b, LZL22, LZL23, LTFL10, lLHNS23, LS07, LHL11, LJ22, LZX22, LX24, LAH22, Lóc18, LM12, Luc97, LWZ21, MZG26, MAH22, MKBY19, Man21, MW98, MS06, MG20, Mez22, Mic93, Moh10, MK17, MMLM20, MCDS24, MRS06, MRS10, Mor17, MHA16, MM11, NS24, NMM25, NRS12, Nie93, NNCN23, NYL26, OI14, OR17, OOO11, PXH25, PYD23, PH14, PLH20, Pla99, Pop21, PG15, PP21, RWTW19, RALJ26, Rum14, Sad99, Sch14, SL16, ST18, ST23, SCTP00, SWS22, Sha97, SMM25, SM21, SDW24, ST21, SH21a, SS24d, SW26, TT21, TLD^{+}23, TY96, Tur94, WO00, WCB15, WQL20, WWBM21, WL25, WG25, WB17, WL17, WPL18, WK20, Wu22, XW18, XWY19, XZP^{+}20, YWX14, YZLP16, Yan17, Yan18, YHZL21, YBK^{+}21, YJX15, ZR17, ZM94, ZZ17, ZHSX23, ZWL26, Zha11]. linear [Zha15, ZD17, ZZY18, ZFG18, ZZ19, ZYW21, ZL23, ZGLH24, ZWZ25, ZFX14, ZY13b, ZM16, ZLV17, ZV21, ZZZ22, ZZLV23, ZL26a, ZFH26, ZWW25]. linear-nonlinear [Man21]. linear-quadratic [Bla15]. linear-time [Luc97]. linearisation [Mot14]. Linearization [Kol04a, PM22, XL24, ZGLH24]. Linearized [BPR21, BRY14, DW22, GCZH24, GZD24, HP18b, KLW^{+}23, LV18, LV21, PJ22, SZ23, WH15, YJ21, ZHFW21, ZY23, ZPEL25, ZD18]. Linearly [YZLZ22, ZQzS22, APPR14, ABM10, BWC22, DW21, DEM94, MBG19, NKA24, Sat24a, Sat24b, TA26, YLYZ23, ZXL25, ZLM25, ZM26]. Lines [CG03, ACF99, ACM93, CL24a, CM05, DS09a, HG25, Sut17]. Linesearch [Anh19, MLM25]. Liouville [AS10, AA12b, BX19, CA22, GHC15, JLX22, LJX25, lLHNS23, Mar93, TBY13, VLCL16]. Lipschitz [AV19, AH11b, CCC25, Dab04, GA15, HLC15, Iid24, KL22, LLL22, RTD^{+}21, TSI20, TRCL24, WXT22, Wan24, XWL25, YYZ22]. Lipschitz-type [AH11b]. Lipschitzian [DHF21]. liquid [RCW22, Zha26, ZWY22]. liquidity [KMV17, SMR26]. Liquids [PP05]. List [All08b, Ano99e, Ano03b, Ano04d]. little [KG23, KJG23]. Lobatto [BCT15, Gau00, Gau09a, Lau07, MVVA08]. LOBPCG [KPS14, KMS23]. Local [AABM17, ASHF21, Ari98, CG20, DZS21, DZ21, DZ22, DMW23, GS16b, GS26, Han93, KV04, Kul10, LD21, MM17, MDR23, MS14b, SW10b, WD23, XWX24, AC17, AR09, AM16, AKT15, Bac18, Bac20, Bac21, Bac23, Bac24, BI00, Che19, CYM22, DWZ14, FMD23, GG22, KAL22, KK22a, KS06, KB20, Len93, LZ25, MV14, MT12, MFPG07, Mot14, QAS^{+}24, RC14, RA12, Sch26, WDL23, Wei17, Wei18, ZAH26, ZW12a, ZG09, ZLZ22]. local-global [KS06]. localization [CX25, Cve06, MT93]. localized [HG25]. Locally [CJKL23, Rab05, BL95, LKBF17, LYY12, LZX22]. locally-kernel [LYY12]. Locating [KPFG04, DFK97]. location [CKL16, HHLS21a, HHLS21b, Xue95]. locations [CG07b]. Log [ZLT^{+}17, DKL15]. log-cosine [DKL15]. log-Korobov [DKL15]. logarithm [AM12]. Logarithmic [CKM^{+}25, CZX26, Gau10, Gau12a, Gau13a, GQ09, KSCS08, LNS23, Ost07, Sab91, Zha23]. logistic [ÇD25a]. Long [CKP97, MKBY19, FT24, Hol98, HWZ24, JJC24, LZ23a, LS11]. Long-term [MKBY19]. long-time [FT24, HWZ24, JJC24, LZ23a]. Look [RSCH^{+}19, BRZ96, Gau22, GM97, Mic93, San14]. Look-ahead [RSCH^{+}19, BRZ96]. look-around [GM97]. loop [BESC22]. Lopsided [LYW14]. Lord [MC05c]. Lorenz [GPAA14, WLL12]. Lotfi [CN16]. Lotka [CCHH23]. Love [FRS21]. Low [BHS14, BG13, CMM17, FH97, GSA03, KKM20, BBL23, BKS13, BHW23, BEHS20, EGG08, EKPU23, Faz23, GK24, HIS26, HV15, Jay21, Kie23, LHZ^{+}21, LLWC24, LQ26, Lun23, MVVA08, MDR23, PK21, RS21, SYZ22, WC23, WKY25, ZJR25, ZFH23, Zhu15]. low-cost [ZFH23]. low-dimensional [GK24]. low-error [HV15]. low-genus [Kie23]. low-memory [HIS26]. Low-Rank [GSA03, BHS14, BG13, CMM17, FH97, KKM20, BBL23, BKS13, BHW23, EKPU23, Faz23, LHZ^{+}21, LLWC24, LQ26, PK21, WC23, WKY25, ZJR25, Zhu15]. low-synchronization [Lun23]. lower [ARSS19, Ber00, HL17, LM17a, LZCW24, Ver14, YKY15]. LOWLAD [BL95]. LP [TO21]. LQ [DTI09]. LQP [Bno21]. LS [AIA^{+}24]. LSMR [Jia20b]. LSQR [BRR13, PS25, RSZ20]. LU [SZ99, Deh20]. Lubich [MSS11]. Lubkin [BBM08]. Luigi [All08a, All08b, GG08]. Luminy [BV96]. lumped [QLZX11]. Lyapunov [DL97, BQO99, BHS14, BH17, BPS23, CSC26, FWC16, GSA03, HJ18a, LWwCL13]. Lyapunov-type [DL97]. Lyapunov/Stein [FWC16]. lying [PS20].

m [SA26, VB92]. M-Padé [VB92]. Machine [RDdRC04, BWB24, MBM25, TOY26]. macro [Zaf22]. Maehly [MC05b]. magnetic [CEK21, CZX26, MT93]. magnetohydrodynamic [MDR23, XY25]. Magnus [YBK^{+}21]. Magnus-type [YBK^{+}21]. Mahony [ZLS24, ZJZ20, ZPEL25]. Maintaining [Sv95]. major [Wat94]. majorant [EHV19, Gon16]. majorization [WZW^{+}24]. majorization-minimization [WZW^{+}24]. majorizing [AH13, AG13]. making [ON26]. Mandelbrot [GJV17, KGN^{+}24]. Manifold [CO03, LZ12, ODL21, WCY25, ZCC26]. manifolds [CKO24, JJK97, RFS23, Rob97, Sha97, SR22, TCO26, UPY26, YHR26, ZZX^{+}23]. Mann [HCD26, KAF18b, MM17]. Mann-type [HCD26]. many [DBV23, JYLC21, LNS23, Szy06]. MAOR [CHK14]. map [BR07a, JVH15, LK20]. Mapped [Mat01]. mapping [BDV18, LZ16b, MWsC19, Tru24, VA20]. mappings [Bag16, BI00, CSI16, CSI17, CRN19, DM98a, Den14a, Fan15, GMZ19, GEA20, HS20, HMA16, IAH20, KAF18a, LQ20, LCL21, MN17, NUG25, PKC18, PDRG19, RZ16, RTD^{+}21, SKK21, SRK^{+}25, SIO20, SCS18, TB19, TSI20, TTA26, YAT20, ZAC26]. maps [AH08, CD99]. Marchenko [GM21]. marching [AF13, BZS22, FLG08, GD15b, GNH10]. marching-on [AF13]. Markov [dC23, BBZ95, BBL22a, Dra00, DMS09, GM23, vGK04]. Markovian [PM26, WZ26]. Marquardt [SG17, SG18]. Martensen [DS15]. Maruyama [ABB15b, ABB15a, GMY18, GHZ25, Hag25, HSY23, LDH23, LGHS25, LM17b, PM26, VH12, WZ26, ZJ23]. Mascheroni [Che16a]. Mass [Sza03, DFP^{+}10, Fly22, LDX23, MvS09, QLZX11, YL22, ZYJY22]. mass- [YL22]. mass-conservative [LDX23]. mass-preserving [ZYJY22]. massic [FJT94]. Massively [AL97, DM21]. massless [MY22]. master [MB09, San14]. matched [DL21]. matching [ASW06, BKPS93, LPRÁ24, Sch17]. material [RW06]. Materials [BE03, RW06]. Mathematica [MdR13]. Mathematical [Ano92, CR00, DHL^{+}04, Jos22, DJS20, JR20, KN23, PH20, SW22, WZZ15]. mathematics [RK11]. MATLAB [AEF^{+}14, BNN16, BMR19a, DdAF^{+}20, FMS25, GHN19, HL03, HCBAEC23, HM22c, Sut17, Yan17, ACF99, CDSV11, CFR19, FHH99, Han94, Han99, Han07, KR23, LP13, LLZ18, MBR21, MS14a, Mil20, NPP04, RZR12]. Matlab/Octave [CDSV11]. Matrices [BO03, BG03b, CPV04, DL03, DR01, FG03a, Han02, HR03a, IMT02, LV01, The97, AGA24, AAAA^{+}18, ADL05, ANI15, ANI^{+}17, ACG20, BCW13, BM19, BM94, BBL22a, BKW25, BM12, BG91b, BAS25, BIM^{+}23, BM25, Bör26, Bos21, VVV22, BK16b, Cas17, CD96, CT93, CO94, CP00b, CLWV15, CW19b, CUY24a, CUY24b, CT06, CK06, DLL12b, DLL12a, DPP19, DP16, DS09c, DS12, DR12, EGG08, EG19, EV22, EK94, Ern00, ED13, Fas23, FRJT09, Faz23, FLV14, FM99, FGS24, FDV13, GWL18, GGJL24, GLS26, GEP14, GEP16, GMP92, GV00, GNT24, Haj16a, HN94, HG93, HPS97, Hil10, HW18, HM18c, Hua18, JBJB17, Jia06, JL16, Jia20a, JXX^{+}23, JXNX24, KH18, KV00, KM09, KM13, KRS19, KN21a, KN21b, LW20, LRT19, LZ14, LPGL16, LL16, LG17, LDL17, LL20a, LYH^{+}20, LLY22, LRW25]. matrices [LZX22, LXQ15, Lu15, MPR22, MVV05a, MW98, Men26, MT15, MK97, Mit11, MA95, NSM20, NR12, Orb15, OO22, PS06, Pol10, RZR12, RWTW19, Sch17, SSC25, Sto93, SW26, Tan17, TT06, TA13, TK94, Uhl22a, Uhl22b, VGV06, VBG96, Vos00, WGZ18a, WZVJ22, WZ15b, WPL18, XHZ07, XLW20, YHZ20, YP23, YLY12, ZWL26, ZCG15]. matricial [BC09]. Matrix [AH03, ABQ04, BP03, BFK^{+}09, BFK11, BCM10, BZ18, CP01a, DP01, GS19a, GL04, GCGVH92, HE05, HS03, JS15, PGGC97, ST02, AG25, AMH10, AMA21, AAPR21, AM21, AL09, BKDA25, BMA16, BBQO07, BL92, Bel08, BQO99, BEQOR14, BEJ20, Ber00, BHM05, BMR19a, BM25, BESC22, BSF17, Bos21, BEH24, Bre00a, CKM^{+}25, CZL26, CR00, CW19a, CS18, CR20, CZ20, CZM21, CGW25, CSC26, CDD21, CsL24, Cve06, DLL13, DMT13, Dax17, DMD16, DMR20, DMR21, DPP19, DH26, DM97, DLLD17, EKPU23, EG19, EV22, EH24b, FJ25, Fan22, Fas25, FPP05, FFSC25, FDFM23, FYI^{+}12, GM23, GWW15, GI10, Guo13, GL20, GSF26, HJ18a, HT16, Haj16a, HS26, HH05, Har18, HHLS21a, HHLS21b, HHLM23, Hig97, HS16, HN16, HM22c, HASI23, HST15]. matrix [jHyPlZ06, HM18b, HM18a, HM19a, HJ21, HC24, HJ25, JL15, KP96a, KST06, KMZ18, Ke21, KBA26, KPS14, KKM20, KKK22, Lei15, LY17, LMUZ19, LLWC24, LWZ26, LQ26, lLhYfD07, LZL22, LZL23, Low05, MS92, MPS20, MS23b, Meu09a, Meu25, Meu26, MG20, MG21, MK94, MG11b, Mor17, MVG21, NH24, NZ19, NR24b, OOOR12, hPwL09, Pen13, PLH20, PZ20, Pop18b, ROB17, ROB18, RWTW19, Rum12, SY20, SDL^{+}23, SD25, SD26, SS21, Sol23, SZ99, SZ24, TPY14, TH23, Ter23, TDWL24, TM10, Uhl09, Uhl22a, Uhl22b, VB92, VC10, VV25, WC23, WCB15, WLW^{+}24, WLS25, WL25, WCY25, WCZ25, WLG26, Wat93, WHS23, WPL18, WL22b, WKY25, XHZ07, XC25, YKY15, YDW26, YLD11, ZTW19, Zha11, ZHY^{+}20, ZLL^{+}21a, ZJJW24, ZLS25, ZYBJ23, ZCC26, ZY13b, ZLV17, ZV19, ZV21, ZMZ26, Zhu15, BBPV12]. Matrix-free [BZ18]. matrix-function [CDD21]. matrix-less [BESC22, EG19, EV22]. matrix-sequences [FFSC25]. matrix-type [BL92]. matrix-valued [AM21, Low05, SZ24]. matrix-vector [Ter23]. Mauro [BT14]. Max [BHH24, fLxX12, Śmi06]. Max-convolution [BHH24]. max-cut [fLxX12].

Maximal [RALJ26, Bic24, CV92]. maximization [DM22, TD09]. maximizing [WCY25]. Maximum [SZ24, ASGGRG23, CS22, FM19, GLM15, SZQS23, XF25, ZYQ^{+}21, ZY23, ZQS24]. maximum-principle-preserving [SZQS23, ZQS24]. Maxwell [JLFL19, LXZZ21, Yua21]. may [KSW07]. McCallig [Dun94]. McKendrick [HT23]. mdLVs [NIN12]. MDPMHSS [CW19b]. Mead [Gal22]. Mean [Neh04, WCHK21, ALRT16, Cas17, Flo16, FP20, GQ09, MAF25, PL25, ST23, SZ20, SZ21, YCW^{+}19]. mean-CVaR [PL25]. mean-field [SZ20, SZ21]. Mean-square [WCHK21, ST23]. Means [BCV03, CD01, Bin96, BM06, GS95, Mil20, Pié99, RALJ26, The12]. measles [ALB^{+}18]. measurable [Dam08]. measure [GM21, NT21]. Measurement [Dab04, CHH93, LXX23]. Measurements [MKO04, HIS26, Zak17]. Measures [BBR03, Sza03, Dam08, DMR21, DFP^{+}10, DDRS23, EK94, FMD18, GHM08, Man07, Man10, Not95, TTV21, VP23]. Measuring [LAH22]. mechanical [CH11, LHW13]. mechanics [GHP^{+}00, GJVW24, MZ19]. mechanism [LLL26]. Media [AA03, CGN03, ASZ23, GM20, Gha24, JLFL19, SLLA15, WBY25, Yua21, ZZZ20, dOS07]. Median [CSFC04]. medical [CTS09, HOW95]. Meir [VA20]. Mellin [DL18, Mon96]. membrane [PPS25]. memory [AB99, And22, BD17b, CLBT15, FT14, FH25, HIS26, JKNR13, KKB16, Liu14, LSG15, LXQ15, Orb15, PS16, SN24, SS06, SLL23, VL19, WZ15a, WZ23b]. memory-less [And22]. memoryless [BK16a, BKA19, LZ22c, LTP18]. Menten [SMK14]. merit [ZLLC11]. meromorphic [EAB20, PS00]. Mesh [MST03, MT04, NS01, SC03b, ST02, WCM94, BPR22, CEZ26, CD15, Das19, FZLL23, FHC21, FSY23, Ish26, Kac18, LZ23b, LZ23c, Maj14, MS01a, Moh10, MK17, MKA26, NV21, QL12, RS26, ST98, WLY^{+}21, YHQ19, YF22, ZL22b, ZZY24, ZH26]. mesh-based [ZH26]. meshes [CL00, CZ94, CZ95, CZ96, CYM22, CL96a, CGL99, FLR01, FS21, Gha24, GCZH24, Kno23, KK16, Lai92, LGC24, LGHS25, Lin05, LRY18, LZ18c, LCGH23, MT12, MS23a, PR10, PTSB01, Ple12, RS25, SS94, SKJ^{+}18, TKS24, WZZ26, XZ25, ZL23, ZWX22]. meshfree [BS19, SB21]. meshing [PTW22]. Meshless [Flo03, MS14b, AD17, AF13, Ant18, AAD14, CMD19, DS09b, DAM16, HG25, KB20, LCZ26, MFPG07, Mil13, Mil20, Sal17, SH12, SW14]. meshless-type [SH12]. message [GSF26]. message-passing [GSF26]. meta [Kub23]. meta-programming [Kub23]. metal [NAR05a]. Method [AZMJ04, AG03, ASS03, AX26, BD03, BD04a, BBD03, Bog02, BE03, BM03, Che04, CN01, CGL01, DLL04, DG05, DS09a, DFF04, ED05, FEK^{+}23, Fun01, GPP01b, GSA03, JC04, KV04, KZ03, LRGH02, Nac03, PG05, Rec01, SC03b, TC05, VR04, WSY04, ZL25, Zil01, ZKD04, ALQ17, AS10, AJ13, AD17, AEH20, AR24, AF23, AAFL23, AAM24, AK19, AIA^{+}24, AG25, ADP26, ATC16, AAH20, AWH^{+}24, AA12a, ADL05, AK12, AM01, ABB15a, ABB15b, AB99, AM12, ARS26, Ali23, ASS13, AG19, AC17, AABM17, ALRÁY25, ABK22, ADA07, AA09, ALL25, And18a, And19b, ABM10, AVCO26, AC11, ATT22, Ant22, AU08, AH08, AR09, Arg09, AH10, AH11b, AH13, AHVR17, AGS20, AM18, AG23b, AC19, AT17, AAD14, AGG17, AHC13, ACH17, ACH19, AVI97]. method [Awa10, BKDA25, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BRR13, BD06, BL23, BGR23a, BN25, BEM99, BDV18, BCS18, BBCS21, Bel08, BB14a, BKS13, BHW23, Ben99b, BEJ20, BEHS20, BEJR23, BGS24b, BX17, BBB22, BKF20, Bhr16, BZV16, BRW11, BFK^{+}09, Bia12, BK18b, BFKM20, BFK22, Bic11, Bic24, Bin96, BF14, Bno21, BS19, BCJ22, BCJ24, BEH24, BX19, BRMG18, BHT16, BDF^{+}25, CQ16, CCN25, CR96a, CR96b, CR99, CMRS00a, Cam19, CM24a, CCTV23, CM15, CHY19, CW21a, CG20, CP95a, CD96, CL10a, CZ14, CC15, CC16a, CDG23, CCL18, CKY99, CL06, CLA11, CD15, CM16, CC16b, CWL16, CFL19, CW19b, CDLW21, CH22, CL24a, CHYH24, CM25, CTY25, CZZ^{+}25, CLQ26, CC13, CYM22, CR23, Cho16, CK22, CW17, Cic20, CJ17, CVX16, Cor24, CFK^{+}20, Cui26, DW12]. method [DWZ14, DWC18, DW24, DMT13, zDYG18, DD20, DMC20, DSI11, DM98a, Das26, DL18, DS09b, DAM16, DS20, DBH21, Deh20, DAN26, DP21, DC17, DZ13, DW22, DHWL25, DN19, DDG05, DO25, DH26, DW15c, DMT22, DEM93, DBAE09, Don16, DJG18, DS09c, DHV22, DYW16, DLDW21, DL09, DN24b, DNH^{+}25, Dum13, DHMS16b, EMS25, Eft15a, Eft15b, EV22, EDAM13, EG18, EE18, EN11, ED22, Ema96, EH97a, EA12, EH24a, ES19, EGGSH13, EHVRV14, EED19, FAMA20, FY13, Fan15, FY19, FW25, Fan22, FEL15, FGM19, FZL^{+}16, FGC19, FHC21, FSY23, FLH25, Le 19, FAM25a, FAM25b, Fly22, FLG08, FMD23, Fra26, FR18, FR12, Gal22, GS25, GM06, GHC15, Gar24, Gás99, GKL21, GN12, GH23a, GH25, Gha24, GK20, Gla01, GNH10, GO20, GLG26, GLdO09, Van17, GLRSG08, GO06]. method [GGS22, GP05, GMS99, GL15, GSV96, GZ11, Gu20, GWL20, GK21, GCZH24, GW24, GMY18, GL20, GJVW24, GHZ25, GZL^{+}25, GHPMGRR14, HLS10, HHST19, HT16, HLN25, Hag25, Hai09, HS26, HL06, HZPW23, Han22, HYW20, Har20, HHHN07, He16, Hei06, HZ93, HCBAEC23, HSTW14, Hey99, HSE16, HS16, HK06, HFDSC24, HASI23, HSK20, HST15, jHyPlZ06, HYT26, HLTA16, Hu22, HWC25, HNY^{+}18, HLZ14, HL15b, HLC15, HL17, Hua20, HS21, HWZ24, HCD26, HZX20, HG25, IAAH24, ITA24, Ila20, Ish26, IDAV09, JL26, Jai16, Jai17, JL12, JSF13, Jat15, JR20, JCL16, JZYY23, JNS19, JZF^{+}20, JJC24, JHLL15, JWL20, JYLC21, JWY21, JH26, JXZ26, Jón93, JN99, KMH24, KKPT16, KPD25, KJ18, Kao20, KP96b, Kar15, KL25, KCHD16, KAF18a, KAF18b, Kaz24, Kel07, KR11, KS20, KH11, KGH14, Kha25, KPFF24]. method [Khe16, KH20, KOK21, KK22a, KBA23, KP26, KS26, Kno23, KKO17, KS23, KL06, KPA20, KH18, KLL10, KJO23, KK12, KB20, KGN^{+}24, KHM20, KP22, LS25a, LLAL21, LP08, LP12, LLY26, LRY26, Li95, LMMH11, LLS11, LWwCL13, LYW14, LHW17, LG18, LZ18a, LWZ18, LLL18, LRC19, LHR20, LLX20, LLY22, LZ22a, LZ22c, LDH23, LDX23, LLD23, LZ23a, LZ24, LBF24, LGHS25, LWY26, LL26a, LLL26, LM14, LZ15, LZ16a, LZ18b, LX23, LYL25, LL26b, LZ16b, LZL22, LZL23, LNS23, LGW14, LS15b, fLxX12, Lin16, LLL22, LHL11, LM13, Liu14, LWQR15, LSW16, LM17b, LRY18, LL18, LF19, LDL^{+}19, LLC20, LS20, LW22, LH23, LZ23b, LZ23c, LCGH23, LJY24, LHlLG24, LX24, LGZ25, LZT25, LZQ26, LTP18, LWN13, LA22b, Lor19, LW25, LKK21, LG95, LKKM15, LCW21, LPXX19, LXP20, LM25, LV21, MKG24, MBG19, Ma20, MS20a, MGL20, MsC20, MH21, MH23, MJJ^{+}23, MBJ17]. method [MAS17, MC08, MN17, Maj13, Maj14, Man21, Mat15, MHR23, MFPG07, MM09b, MB06, Meh11, MXL25, Mia19, MZ19, Mil13, Mil20, Mil19, MNS23, MS14b, MS23c, MDH16, MDL15, Mok16, MR96, MN23c, Mor25, MRS06, MPB16, MSMS12, MSM12, MS13, Mot14, MWWY13, MM11, MA12, MAFN16, NH24, NAHZ21, NMM25, ND21, NRV23, NJ13, Nie93, NP22, NEMS14, NAE22, NAA19, NA24, Nov26, Now06, Now13, NW17, NA26, OMU^{+}24, OdZdRV13, OBAHK^{+}19, OL23, OAMA22, OOO11, Pan96, PQS22, PL25, Pan18, PFT98, PG12, Par16, PC13, Pea13, PLZ^{+}24, PPS25, PPPN23, PRK^{+}18, PM26, Pie96, PV00, PK22a, PK22b, Pla99, PR93, PN21, QL12, QXGZ20, RFS23, RNT26, RKMS16, RR20, RMT13, RBN14, RT19, Reb97, RN21, Ree92, RT22, RZ23b, RSZ20, RWB09, RAH11b, RWTW19]. method [RWTM21, RZZ25, RM13, RS26, RM26, RALJ26, RW06, Saa23, SBA26, SS99, Sad99, SAC18, SSS14, SP21, Sal17, SLW13, SSS21, Sch25, SHF15, Sch08, Sch17, SEG14, Seg98, SV26, SMZMA18, SB21, SW14, SWS22, SW24a, SS12a, SGS13, SMM25, SLD20, SIO20, SHLY18, SSYL20, SS06, SW11, SCDM20, SDL^{+}23, SCF23, SFS23, SHD25, SD25, Sho18, SGJ15, Si12, SLT20, SS21, SR22, SSK23, SS24c, SK25, Śmi06, Śmi13, SB24, Sol11, Sol15, SMN24, SJ14, SYLT14, SG17, SG18, SXHZ20, SJW21, Soo15, SA23, SSSS22, SGK^{+}99, ST98, SZX11, SG23, SLL23, SZ24, SW24b, Sut17, TT21, Tai92, TZH24, THF21, TA26, TO21, TH23, TZ25, TYSY20, Ter23, TBY13, TH18a, TVC20, TLD^{+}23, TLR^{+}25, THS20, TDWL24, TY96, TKS24, TA13, TL24, THT19, UPS25, UPY26, Val14, Val15, VC10]. method [VdR13, VRM23, VDVJB12, VPA24, VL19, VLCL16, Vos00, VT10, VN18, VS19, WC23, WT08, WGK11, WKG11, WK12, WSY12, WZ13a, WW14, Wan15b, Wan15a, WZQT15, WZ16, WCLW16, WZQ17, WGZ18a, Wan18, Wan19, WLMA21, WHD22, WXT22, WL22a, WZC23, WDL23, WQ23, WLW^{+}24, WS24a, WX24, WLJ24, WPW25, WG25, WJC25, WBY25, WCZ25, WZC26, Wei17, Wei18, WRM17, WLY^{+}21, Wit96, WB17, WC10, WWD^{+}12, WC13, WZZ15, WL17, WPL18, WCD21, Wu22, WKY25, XY25, XW18, XLW20, XZZ22, XCD23, XWL25, XL24, XZW13, XH25, XL14, XH20, XWX24, XC25, Yak95, YAT20, YWX14, YYL15, YZLP16, YW17, wYN18, YJ18, YSLH19, YZL20, YPL21, YZ21, YF22, YK22, YXS22, YLYZ23, YSLL23, YZ23, YHL25, YHS18, YZBJ21, YHR26, YL22, YL16, YJJ^{+}21, YYL26, YZ11, YJX15, YYW21, Yua21, YLL22, YZLC24, Yue25, ZR17, Zaf22, ZA20, Zas22, ZZ17, ZHSX23]. method [ZZ24, ZWWW20, Zha11, ZW12a, ZY13a, ZJ14, ZHT15, Zha15, ZD15, ZZY18, ZFC18, ZSF18, ZZ19, ZYX19, Zha19, ZJZ20, Zha20, ZBX21, ZY21b, ZYW21, ZD21, ZL22b, ZZ22b, ZL22a, ZF22, ZLLH22, ZLH22, ZZ23, ZL23, Zha24, ZWXX24, ZSWW25, ZHW25, ZDZ26, ZH17, ZWG18, ZWW21, ZH23, ZYBJ23, ZG12a, ZG12b, ZLV17, ZV19, ZV21, ZZZ22, ZCG15, ZL17, ZS08, ZCS14, ZW15, ZPX21, ZWX22, ZJ23, ZLS24, ZZY24, ZCTD24, ZJSS25, Zhu15, ZYLN18, ZXL23, ZLTA16, ZS13, ZD18]. Methods [AH03, Ano05b, BCN^{+}16, BKM03, Bre02, BZ02, Bre03a, BJ02, BO02, But02a, BH02, BJ04, CVA01, CO03, Che02, De 02, DOS03b, EJR02, FG03b, GPP01a, GWY26, GR01, IW04, Jay02, Jbi03, KB02, LS03a, LM04, Lui02, Mar04a, McL02, MR02, NPP04, NKS04, PR03, PL04, RS02, SFT03, SME03, Tir02, Van02, Wri02, AAA^{+}18, AD22, AH11a, AHP20, AHIJ22, AHJ22, AP26, AHJ17, AMM11, ABS19, Ahu09, Aih17, ASZ23, AMCM06, AK15, ABKD23, AH15, ABI20, ABI22, ABI23, AH14, And22, AAH18, Anh19, ATT21, AT21, Ano17, ALZ21, AMM16, Ant18, ASHF21, AH09, Arg10, AH12, AM16, AG17, Ari98, AKT15, AKKT16, AKQ17, Axe99, AK00, ANA14, AFN16, AFN17, BKG15, BKR18, BKA19, BD09, BD10, BBS20, Bag00, BR06, BPP23]. methods [BMR97, Bai97a, Bai97b, BBC11, BZ13, BRY14, BH22, BLW09, BGL26, BRZ18, BGR23b, BBBC20, BBO21, BJ98, BHS11, BOW21, BCM16, BEJS21, BHS14, BH17, BJT24, BBJ25, BJT26, BR07a, Ber10, BBB22, BCT15, Bia94, Bla15, BE20, BMS24, BRZ96, Bre99a, Bre99b, Bre00a, BRZ19, BRZ20, BF99b, BCI14, BS14, BMR19b, BDR25, BV09, BGZ20, BB14b, BBL22b, But98, BD98, BC99, BC01b, BC01a, BH05, BH09, But10, BI14, But15, CWZ13, Cai22, CE17, CGPM00, CR12, CCV23, CW19a, CCJ10, CIP10, CJ12, CJSZ14, CH26, Cas17, CK24, Ceg25, CS99, CZ94, CZ95, CZ96, Cdv98, CT10, CTS09, CL13a, Che14, CQLY15, CGYZ19, CLWH20, CZM21, CCC25, CL10b, CH11, Che19, CGM25, CN16, CN17, CV22, CEK21, CL93, CGL99, CGV26, Con93, CDI14, CDP16, CHMT10a]. methods [CKKT16, CRHTV24, CPN14, Cuy00, CK06, CHK14, CKS16, DJ10, DFJP10, DEP12, DDP14, VV07a, VV07b, DHJJ10, DEC24, DBTG24, DD21, DNV25, DK25, DMA19, DW21, Die08, DLL^{+}24, Don10, DBGB11, DR07, DZ21, lDzS21, DMW23, DZ24, DLLD17, Dzu13, EG10, Eba18, EKPU23, EHTSM21, EHN17a, EHN17b, Erb15, EH24b, Ern00, EGSHVN15, Fan19, FYM14, FYYW19, FHL21, FLY25, Fan26, Far20, FHS12, FH15, FT14, FGP91, FH25, FAM25a, FAM25b, FHH96, FRR07, FGR01, GH10, GA15, Gal18, GD15a, GP99, Gar20, GG22, GHN19, GX19, Gon16, GPHHAPR18, GLV05, GS19c, GNT24, GCJ26, Guo13, Gut15, HT21, HH11, HH12, Hal14, HJ18b, Har18, HT19, HHF22, HKP25, Hea10, HR14, HVYMS23, HPS97, Hey01, HE05, HMA16, HR00, Hil10, Hoa25, HCH18, Hu22, HLV26, Hua94]. methods [HZ15, HM18a, HWXC19, HYJ20, HZ24, HJ25, HM14, Ibr25, IMT23, Iva24, JT96, JKM18, Jay21, JS15, JLJ22, JB24, JM93, KKR25, gK25, KCBT21, KT07, KNBGV18, KM24b, KD14, KMZ18, Kee94, KL22, KMH25, Kha14, KS14, KMA13, KBA26, KC24, KJ15, KKM20, Kri26, Kul10, KK16, KGMH21, KV12, LS14, Lam09, LP13, LEK21, Li97, LW13, LW14, LW16, LY17, LG17, LM17a, Li17, LDW18, LG19, LC19, LRL19, LD21, LZ21, LGC24, LPRÁ24, LLY25, LZZ19, LDZ24, LCH20, Lin98, LT20, Liu11, LSW16, LX17, jLyLqW17, LD20, LSY^{+}23, LH23, LZZ24, LZCW24, Lóc18, Lóc20, Loh22, LJbL21, LSG15, LSSS15, Lun23, LHW13, lLXhL22, lLXZ24, ML20, MY22, ME92, Mac96, MA16, ML22, MAH22, MP07, MCC26, MCW22, MJH17, MJ18, MJ20, MVVA08, MM19, MR12, MT18, McL25, MN23b, MV17]. methods [MG20, MG21, MClXzJe16, MG11a, MSZ20, MH08, MP13, MKS18, MCDS24, MRS10, MG22, MRS25, MHA16, MS96, MGG25, NBK17, NW19, NHP06, NRS12, NC94, NB16, NK16, NK21, OQ12, OIM21, OMW21, OI14, OLB94, OM18, Osa92b, Osa12, OM26, ODL21, OAR22, Ovi22, OR24, ÖRBB14, PK21, PNW17, PR10, Pas11, PP24, POP17, PHI98, PMM11, PR14, PS16, PH20, PS20, PSS25, PV98, PG15, Rab92a, RSKB17, RTCL21, RA12, RhG15, RS97, RS93, Rog95, RVF07, RS25, SBN25, Sad05, Saf10, SPB26, SKK25, SS11a, Sar06, SFMK23, SW10a, SL21a, SH12, Sha97, SAHM25, SS10, SM21, SL21b, SS15, SKTGR19, SS98, SWB08, SMR26, ST99, ST92, SvF94, Sv95, SS08, SS12b, SS16, SM10, Ste20, SS94, SZ21, SLL22, SZQS23, SC18]. methods [SX00, TQC22, TPY14, TX19, TTXZ23, TCOA19, Tem08, TTLD20, TRSI23, TRCL24, TCW14, TÖ17, USAF14, USM24, VH12, VV11, WO00, WZ13b, WK13, WSK14, WK14, WK15, WZ15a, WDL16, WK16, WK17, WZ19, WCHK21, WYZ21, WD23, WL24, WC25, WL25, WHS20, WF23, WL22b, XT16, XLC93, XZP^{+}20, XLG22, XSL22, YM24, YYK24, Yan95, YWWR12, YDWL15, YZ17, YBK^{+}21, YH97, YZL25, YZY^{+}14, You16, YYD14, YWZ19, YSXY19, YD26, Yue25, YP09, YQM16, ZCT19, ZYW17, ZWFY19, Zha09, ZW12b, ZLW^{+}13, ZW14, ZZH15, ZCGS24, ZXL25, ZXF14, ZFX14, ZXLF15, ZLL21b, ZH25, ZY13b, ZM16, ZZLV23, ZL26a, ZLT25, ZZB20, ZYW22, ZHZ25, ZWW25, dR99, vSv94, vdHM98]. metric [CS24, IUM^{+}19, PSS25, lZqL25]. metrological [CP93, CP95b]. metrology [For93]. MEW [FW25]. MFS [CKL16, KST06, Kar09, Kar10, SK04, ZBDK23]. MHD [MSMS12, SFS23, XSL22, ZSF18]. MHSS [BBC11, YWX14, ZZ17, ZCG15]. Michaelis [SMK14]. micropolar [XL24]. microscopic [SR22]. microscopy [MvS09]. microwave [ZZX26]. midpoint [AU08, HWC25, TB19]. Migration [De 02]. mild [AH09, HVM15]. Milne [MJ18]. Milstein [JLWZ25, KJ18, TS15, ZSLZ24]. Milstein-type [JLWZ25]. MIMD [AAIT94]. MIMO [KV07]. min [Ska13]. min-range [Ska13]. minima [PSWE23]. Minimal [AM01, FM04, MS01b, SC03b, AABTB23, BJT24, Liu11, LXQ15, PP21, Sad05, Sid94, WCD21, YKY15, ZD15, dAFPR23]. minimal-norm [dAFPR23]. Minimality [Mas05]. minimax [CT93, Jón93, KLT95, ME95, PR93, Ree92, ZGLH24]. minimisation [CC06]. Minimization [GR01, LM04, AAA^{+}18, ALV20, And14b, Ari98, AGG17, BR21, CKM^{+}25, CZX26, DR12, FGBP21, GLS^{+}18, GZD24, HSK20, IUM^{+}19, JL26, LP18, LLL18, Liu14, LX24, MBG19, ML22, MH22, OAR22, ÖRBB14, PKC18, QZG^{+}19, SG10, SLL22, SLL23, VL19, WL24, WZW^{+}24, ZFZ19, ZLL21b]. minimizing [HL15b]. Minimum [ZYW21, DIM22, HYW20, HZ15, LRL22, OIM21, Rip93, SH17, TLD^{+}23]. minimum-error [SH17]. minimum-norm [LRL22, OIM21, TLD^{+}23]. minimum-time [DIM22]. mining [PS25]. Minkowski [FH04, FHH05]. MINRES [HFDSC24, Soo15]. Mirakyan [AUD18]. mirror [BG24]. miscible [HFZ19, Hu22, LDX23, YZ23, ZYGQ17, Zha19]. Missing [BK18a, lZfLqL26]. mitogen [MB09]. mitogen-activated [MB09]. Mittag [WZ23a]. Mittag-Leffler [WZ23a]. Mixed [CHYZ98, CM98, DM03, Far20, Kee94, OC24, PW04, SFZ22, XLL26, ASZ23, AAN25, AR26, AW23, BD17b, CFR06, CA22, Den14a, lDzS21, DZ22, DMW23, EH97b, Gaj05, Gha24, GPHHAPR18, GZP18, HFZ19, Hu22, HS21, JZYY23, Kar15, KAF18b, KMS23, KN18, LR18, LDX23, LTFL10, LYL15, Man21, Oua99, SGO22, SS24a, SKP20, Si12, SSP15, SA26, WG99, WZL26, WLY^{+}21, WZ23b, YZ23, ZSGJ25, ZYGQ17, Zha19, ZF22, ZYJY22, ZW22, ZGK^{+}25]. mixed-order [EH97b]. mixed-type [LTFL10]. Mixing [And19a, ABV23]. mixture [XLL26]. Mizuno [YZLP16]. ML [HHLS21b, HHLS21a]. MLS [AAD14]. MN [WGZ18a]. MN-DPMHSS [WGZ18a]. mobile [QXGZ20]. mobile/immobile [QXGZ20]. Möbius [LK20, MP08, dDS00]. mode [BJT24, CCZ23]. Model [BW93, Car01, Cha04, GSV96, RDdRC04, Str09, dADdRC04, ALB^{+}18, AHJ17, ASGGRG23, AAN25, AAA25, ACSD16, BKDA25, BMA16, BHS14, BH17, BHS23, Bün24, CX25, CPZ14, CGSS24, CV15, DZ21, DMW23, DJM^{+}18, FAMA20, FZLL23, GGV96, GM20, GA20, GS16b, GCFF95, JKNR13, JLWZ25, Kao20, KMV17, KXD25, KLZV95, KV07, KN18, KKA17, KA26, KBP17, LLY26, LGL23, LWLT19, LHZ^{+}21, LZM23, LLLD17, LCW20, Liu21, LZZ23, LCW25, NM26, NAHZ21, ON26, OKP21, Pan18, PPS25, PRVI20, QXGZ20, RC14, Rou25, SPB26, SKD26, SMK14, SLL22, TA26, TD09, VSA12, WJW14, WDL23, WZZQ24, WLY^{+}21, WCH15, XWC26, XW17, XH21, YYK24, YD09, YLYZ23, YH24, ZLT^{+}17, ZD21, ZWW21, ZLL21b, ZHZ25, ZH26]. model-hybrid [WJW14]. modeled [CMWP20]. modeling [DJS20, DR07, GM23, Ila20, NPS09, Pog98, RCW22, XLL26, Zha26, ZWY22]. Modelling [Hol98, CCLi16]. models [AH23, ARS26, Bün18, BWB24, CDLW21, Che22, CT21, CCLi16, HÇK26, LCW23, PGGCGF11, SKA23, SKK25, SFMK23, SSK23, TCOA19, XCLA15, YW17, YK22, ZY13a, ZHY^{+}20, ZLL^{+}21a]. modes [BC26, GH23b]. modification [BC92, MP14, OLB94, TS92, Ver99]. Modifications [GN12, CKY99, Gau13c, MS13, SSN^{+}12]. Modified [AAM24, AB06, AAH18, AHKW04, AHKW05, BK13, CW19b, DWC18, DYW16, GB21, GSA03, GTA19, HWXC19, HFW^{+}21, HC24, KT07, Maj14, Rei98, SS12a, SW11, Spa24, TH18a, TL24, WZC26, WRM17, WL17, XLW20, XWL25, ZWXX24, ZL25, dAR06, AF23, AAN14, And10, AC19, AVI97, BKFMA11, BK16a, BKDA25, BD06, BF18, BW13, BG91b, BRW11, Bra06, BGVHN96b, CSI18, CZL26, Cam19, CD96, CC16a, CL13a, CWL16, CHMT10b, DW12, DL18, Deh20, DW21, DMT22, DDRS23, DJG18, Eft15a, Eft15b, EK94, FGC19, Guo16, GLL19, Hem94, HYT26, HWXC18, HCD26, KM17, KM19, KR20, KK22b, KZ21, LMV24, LG17, LG18, LZ22c, LL26b, LLLD17, LL20c, LZZ23, MJJ^{+}23, MP00, MMGH17, Meu09a, OPSM22, gOM14, RT19, RZZ25, Rha22, Sab92a, SS10, Si12, SCS18, TQY21, TY96, VBG96, WK12, WK13]. modified [WK14, WK15, WDL23, WL24, Wan25, WBY25, WC13, XM16, XZZ22, XZP^{+}20, XCY21, YAT20, YZZL17, YZLZ22, YLL22, Zha09, ZYW21, ZV19, ZCG15, ZCS14, ZCTD24, ZXL23]. Modify [AA16b, Hea10, MH08]. modifying [CZH22]. modular [BGR23b, RZ16]. module [VB92]. module-theoretic [VB92]. moduli [RW06]. Modulus [BZ13, DW24, MG20, MG21, ZZLV23, CW19a, CZM21, Fan22, GZL^{+}25, HLV26, KAL22, KMZ18, LY17, LZOY22, LLY22, LZL22, LZL23, Mez22, RWTW19, WL22a, WL25, WCZ25, WPL18, WL22b, XZP^{+}20, Zha11, ZY21b, ZY13b, ZLV17, ZV19, ZV21]. Modulus-based [BZ13, DW24, MG20, MG21, ZZLV23, CW19a, CZM21, Fan22, HLV26, KMZ18, LY17, LLY22, LZL22, LZL23, Mez22, RWTW19, WL22a, WL25, WCZ25, WPL18, WL22b, XZP^{+}20, Zha11, ZY21b, ZY13b, ZLV17, ZV19, ZV21]. MODUS [MSCB93]. Moisil [Caç10]. molecular [CB00, GKQ24, Rei98, Saf10]. Molecule [MN01]. Molecules [LM04]. mollification [BZS22]. mollified [GD15b]. Moment [BH92a, ZLT^{+}17, BGVHN92c, ÇD25a, GAM24, Hag13, Hua96, IS17, IMT23, JLP20, RS93, Sch17, Str09]. moment-based [IS17, IMT23]. moments [BG91b, CP95a, LMV24, Mah10, Roh07, VBG96]. momentum [LT24, Sei98, SD26]. Monge [BS19, SG10]. monitored [Bal11]. monitoring [AMM17, SL16]. monodomain [GS19c]. monolithic [BZ18]. Monotone [Guo13, KLT03, AAFL23, AAM24, AIA^{+}24, AWH^{+}24, AAB13, AG19, AKB15, AT21, ATT22, AABTB23, BAK25, Bog13, BC16, Dey23, ER19, GS21, GZ11, HCD26, KAF18a, La 17, LZ21, LYY12, LF19, LQ20, LSY^{+}23, LJY24, LZT25, Ngu16, OMU^{+}24, OL23, RIAA19, RW11, SBA26, SI17, SI18, SLD20, TAM21, TB19, TTLD20, TSI20, TRCL24, Ton25, WHS20, WD96, YL19, YLL20, YJJ^{+}21]. Monotonic [LM08, VN18, Xu19]. monotonicities [Leo07a]. monotonicity [CCN25, OIM21]. Monte [Mil20]. Moore [Hua21]. Moreau [Luc06]. Morley [She15]. Mortar [BBHM03, Rah11a]. most [PR93]. motion [AB98, HOW95, LD20, MSCB93, NM26, NT21, Poc14, QQX23, YX11, ZTZZ19]. Motions [DL04]. Motzkin [NYL26]. Moulton [MG11a]. Moving [QL12, RS26, YHQ19, AD22, AAD14, CD15, Che24, FSY23, Li97, SMN24]. MP [LH23]. MPRK [IKMS24]. MR [BRS92, CZ95, Tov98, XYZ14]. MSSOR [Zha15]. MSSOR-based [Zha15]. MSTMAOR [CKS16]. Much [GR01]. Müller [WT08]. Multi [ALJLYJ24, BC05a, DCG25, LW16, LG17, AD22, ATC16, AAU26, AH23, BD09, BD17a, Bhr16, BMR19b, CTY25, CV22, DP21, Den14a, DLYH17, lDzS21, EH24a, FMD23, GZL24, GPHAPR24, GLW16, HS15, JLZZ23, KKR25, KPD25, KT07, Li17, LSZW19, LWLT19, LKBF17, lLHNS23, LWZ23, LHlLG24, LF25, MBJ17, NHP06, PV23, PSB26, RS26, Saf10, Sal17, San25, Sha97, SW19, SKM25, SPS26, SzS21, TZH24, WK12, WW14, WK15, WK16, WK17, WLMA21, WWBM21, WHD22, WZC23, Wei17, WHS23, WCD21, YZ17, YZX25, Zaf22, Zag24b, ZYW17, ZSWW25, ZCC26]. multi-accuracy [Zag24b]. Multi-block [DCG25, ZSWW25]. multi-center [Saf10]. multi-constrained [PV23]. multi-degree [GLW16]. multi-dimensional [AD22, AAU26, Bhr16, FMD23, GPHAPR24, HS15, LKBF17, lLHNS23, LWZ23, WWBM21, WHD22]. multi-dimensions [AH23]. multi-frequency [BMR19b, SW19, WW14, YZ17, ZYW17]. multi-grid [CV22]. multi-input [YZX25]. multi-linear [KKR25]. multi-objective [CTY25]. multi-order [WCD21]. multi-pantograph [SPS26]. Multi-parameter [ALJLYJ24]. multi-physical [LHlLG24]. multi-point [BD17a, KT07, NHP06, WK12, WK15, WK16, WK17]. multi-precision [Zag24b]. Multi-resolution [BC05a]. multi-scale [SKM25]. multi-singularities [PSB26]. Multi-step [LW16, LG17, ATC16, BD09, EH24a, KPD25, Li17, MBJ17, Sha97, WHS23]. multi-symplectic [JLZZ23]. multi-term [DP21, lDzS21, GZL24, LSZW19, LWLT19, LF25, RS26, Sal17, San25, SzS21, TZH24, WLMA21, WZC23, Wei17]. multi-time [SKD26, Zaf22]. multi-valued [Den14a, DLYH17]. multi-view [ZCC26]. multibody [Pog98, Sim98]. Multicriteria [SR04]. multiderivative [MCDS24, Sho18, SKTGR19]. multidiagonalization [KLSS14]. Multidimensional [VR04, AAH24, BGRS09, BJT26, CL92, CCZ23, GZ11, MZW20, Ngo23, ST23, VPL97, WW14, YZ17, lZqL25, lZfLqL26]. Multidomain [CFK^{+}20]. Multiextremal [MS01b]. Multigrid [Gás99, Jan03, JB24, PV98, vLV02, BH01, CL00, CC06, JK19, JH22, LW25, MClXzJe16, MRS25, Not22, PV00, WL22a, WPW25, XH20, XWX24, YZ11, ZY13a]. Multigroup [AG03]. multilayer [EHNR23]. Multilevel [CK05, HL02, HMT17, IL05, LW20, LJbL21, Meu02, Osw01, Spr01, UL18, Axe99, CZ96, LXQ15, MRS10]. multilinear [BHLZ21, BBJ25]. multinomial [ASVC21b]. multiobjective [ARSS19, EE18, MN22, MPB16, UPS25, UPY26]. Multiparameter [CC03, HKP25]. Multiple [Boy05, BS14, EE18, HPS13, Rev03, AS10, BCM16, BCMT18, BRW11, BF14, BCM10, BDR25, Buo17, CLaL00, zDYG18, DMD16, FHV15, GH10, GH09a, Gás99, Hey01, HE05, HCH18, HKCW24, KBCG13, KD18, LK20, LMMH11, LZ22a, Pat25, PM05, PV00, RTTH22, RG10, SS12a, SHLY18, SS24b, SSH20, TKSG23, TT21, VSA12, WLL12, XH25, ZCT19, Zag24a, Zag26, ZCS14, ZJ23, ZWW25]. multiple-precision [Zag24a, Zag26]. multiple-root [LK20]. multiple-sets [Buo17, zDYG18]. multiples [Bea96]. multiplex [NR24a, NR25a]. multiplication [Bör26, Bün24, OOOR12, Rum12, Ska13, SH17, Ter23, VV25]. multiplications [GLM15]. Multiplicative [MG11a, BuR24, HL23a, HL23b, JML25, MZG26, ÖRBB14, OB16, SDW24, WG13]. multiplicity [GHPMGRR14, LWG18, Tur94]. Multiplier [Kun01, GL15, KS26]. multipliers [LWZ18, PR25, RWTM21, ZZY18]. multiply [ZLH22]. multipoint [AG13, BD10, BGVHN96a, CCTV16, LDX23, WGK11]. multipole [DMC20, LX23, LX17]. multiprecision [BF00]. multiprocessors [AAIT94]. multiprojection [CE94]. Multiquadric [GZ18, CM24b, SG23]. Multirate [NW19]. Multiresolution [ASS03, Rad08, ABT07, ALRT16, RR08, YL16, dCOS21]. Multiscale [CsL24, GZ20, AD00, CF05, LWN13, MM12]. Multisearch [TBC^{+}23]. multishift [DG94]. Multisplitting [SME03, BMR97, Bai97b, BZ13, SX00, WRM17, XZP^{+}20, ZZLV23]. multistage [JSZ22]. Multistep [CVA01, FHS12, Mor17, Yan95, ASHF21, CMR93, FH15, HLW26, HCH18, KMH25, Lóc18, Lóc20, MAH22, MJ18, Meh11, NB16]. multiterm [GHZ25, HV22]. multithreaded [Kub15]. Multivalue [KB02]. Multivalued [TTA26, ATT21, ATT22, IAH20]. Multivariate [ACO03, All03, AGRT05, BC00, BL04, CV92, FM16, GSZ22, GM92a, LWAG08, MG91, MG94, NKM25, PW04, dB07a, dBGKR08, All18, AC94a, AC94b, BM94, BC06, BP23, CCS05, CLTA10, CY10, EAB20, EEM20, Gaj05, HKPW19, LMMD05, PV22a, PW14, Ste95, van93]. Multiwavelet [BP03]. multiwavelets [BCL00]. Müntz [Maz99, ROB18]. MUSIC [FZ26]. mussel [LLY26]. my [Gau07].

Naghdi [PJ22]. nano [FHAL15]. nano-fluids [FHAL15]. Nash [BK18a]. NATO [Ano93]. Natural [CDI14, CEK21, GS26, Tra93, ZZH15]. Navier [LN22, BRY14, BGR23a, BW13, BK18b, DZ22, DZ24, DSS14, FLH25, Hei06, HM18c, JZ25, JY23, KM17, LS25a, LD21, LBF24, Pei24, PV98, PV00, PG05, Si12, Sla06, XM16, ZD18]. Near [BV21, ACM93, BDIR18, BRS91, BRS92, BZ94, CD99, CM96, DHS97, ME95]. near-best [BDIR18]. near-breakdown [BRS91, BRS92, BZ94, CM96]. near-minimax [ME95]. Near-optimal [BV21]. nearby [Gau13b]. nearest [HS16, NN26, Zhu15]. Nearly [Boy05, Ple12, KN21a, KN21b, SKK21, ddS24]. nearly-colliding [KN21a, KN21b]. Nearly-Singular [Boy05]. nearness [LHZ10]. Necessary [Mat01]. negative [AFG26, DGST15, EJR26, IKMS24, KOK21, LV15, PLH20, XyJl16, ZCC26]. negativity [Zaf22]. neighborhood [KOK21, SMZMA18]. neighbourhoods [Len93]. Nekrasov [LDL17, GGJL24, GEP14, GEP16, LPGL16, LDL17, LYH^{+}20]. Nelder [Gal22]. Nemes [LM15]. Nernst [SSYL20]. Nested [ABQ04, DB98, BK08, CZ23, CL19, Wan22]. Nesterov [BG24, Khe12b]. net [CFR06]. nets [AK26b, God15, Lai92]. Network [Cha04, NR24b, DMR20, GSF26, Hof05]. networks [BS25, DMR21, EHNR23, FDFM23, LM00, NR25a, NR25b]. Neumann [SJW21, ACH19, CDG23, DBAE09, HGVPA92, Med10, MW16, OM18, PM25, SJW21, WZ19, XWX24]. Neumann/Neumann [SJW21]. neural [BS25, CZ20, Hof05, LM00, ZHY^{+}20]. neurodynamic [OL21]. neutral [CCN25, MD21b, PM26, WC25]. Neutralizing [Gau13b]. Neville [CM92, GM92a, XZL12, dC20, dC22]. Neville-like [XZL12]. Newman [Odl00]. Newton [AG15, ATC16, ABKD23, And19b, AC11, AMM18, AH09, AH10, AH11b, AH13, AG17, AGS20, AG23b, BD09, BD06, BEQOR14, BCV03, BKM03, Bre03a, BKL10, CCV23, CCTV23, CCJ10, CKP19, CKP22, CWL16, CW19b, CPV04, CHMT10b, DWC18, DBH21, DO25, Dra02, FGJ00, Fra26, GH10, GKL21, GN12, GH23a, Gon16, GS16b, GO20, GLdO09, GGS22, Gug96, Guo13, GLL19, GHPMGRR14, HVM15, Hey99, HFW^{+}21, IDAV09, INR01, JKM18, Khe16, KH20, KMRAR25, KBA23, KBA26, Kie23, LP08, LP12, LP13, LEK21, LG17, LG18, LLY22, LZ22a, LL26b, LLL22, LYY12, LS07, LHL11, lLXhL22, lLXZ24, MsC20, MBJ17, MA16, MPR24, MR09, Mor25, MG91, Pan26, Par16, PC13, PL04, PSS25, Rev03, SSS14, SSS21, Sch25, Seg98, SW24a, SGS13, SSK23, Śmi06, Śmi09, Śmi13]. Newton [Sol15, SSSS22, Ste20, SGK^{+}99, SZX11, TF00, TO21, VPA24, WCLW16, WC13, WZZ15, XLW20, XL24, Yak95, ZH22, ZWXX24, ZYBJ23, ZCG15, ZCS14, ZPX21, dC20]. Newton-iterative [SZX11]. Newton-Like [CPV04, AG17, BD09, CCV23, GLdO09, Guo13, MA16, SSS14]. Newton-type [AH09, CCTV23, HVM15, MsC20, SSS21, Sol15, XL24]. Next [NKS04]. NI [ZLG^{+}13]. Nicolson [LWJ21, LS25a, LBF24, LW22, MS24a, QQX23, TL24, Wan19]. Nicolson-type [Wan19]. Nicolson-weighted-shifted [LWJ21]. Ninomiya [HHHN07]. Nitsche [BBB22, Ish26, ZZ23]. no [BRS92, CZ95, CP95b, Tov98]. Noda [CVLX19]. Nodal [AG03, AB06, BK13, Sch14]. Node [dABR01, DMR21]. Node-Weight [dABR01]. node-weighted [DMR21]. nodes [Ber11, BSL18, Car10, KRS19, KN21b, Plo93, Pre93, SH12, SH23, TB25, KN21a]. Noise [BPV13, Elf25, MRS10, AK16, BB14b, CHH93, HY21, HFW^{+}21, HSY23, JHLL15, Kim26, KLL10, LZ22b, MKBY19, NK16, Prz16, SBJC19, VH12, WG13, WHZ^{+}18, WZW^{+}24, WC24, XT16, XLL26, ZAH26]. Noise-reducing [MRS10]. noises [DMYT23]. noisy [GZ18, KP16, Kim26, LGP11, MJF09, Ros97, YCW^{+}19, ZWZ25]. Non [AFG26, ALRÁY25, AM98b, BR07b, BE03, DEC24, EJR26, KS12, MAF25, Sal94, ZR17, ARTY20, AV19, AHVR17, AAD14, Ave20, BBBC20, BBBC23, BM22, BGS24a, BF17, BC17, CM16, CD25b, CCC25, CMWP20, CW17, DL08, DLR12, DHF21, DGST15, Dze15, El 18, EH24b, tFZyZ16, FHAL15, FP18, GWW15, GHM24, Van17, GS21, GZ11, HLN25, HLW26, HV22, HVYMS23, HVMMV24, HLC15, IKMS24, Jay21, KM24b, Kaz24, KL22, KGMH21, KGN^{+}24, Lam09, LV15, LRY26, LLY22, LPRÁ24, LCZ26, LN10, lLHNS23, Lin16, LCW21, MH21, ML22, MV14, MM12, MS23a, MT15, Moh10, MMLM20, NSM20, NNCN23, PLH20, PG15, RF23, RTD^{+}21, RS26, ST22, ST23, SB21, Śmi06, TZH24, Tan17, TSI20, TTP26, Tru24, VB91, Wan24, WLS25, WB17, XWL25, XH25, XZZ19b, XCY21, XH21, YHQ19, YBK^{+}21, YJX15, ZAH26]. non [Zaf22, ZJR25, ZW12a, ZCC26, dS00]. non-analytic [Van17]. non-autonomous [Jay21]. Non-blind [MAF25]. non-characteristic [XH25]. Non-commutative [Sal94, HLW26, WLS25, YBK^{+}21]. non-compact [CMWP20]. non-convex [BBBC23, El 18, GWW15, LCW21, ZAH26, ZJR25]. non-convexity [BBBC20]. non-Darcian [CW17]. non-differentiable [AHVR17, HVYMS23, HVMMV24]. non-disjoint [GHM24]. non-equilibrium [YHQ19]. non-Fickian [RF23]. non-global [IKMS24]. non-globally [CCC25, KL22]. non-Hermitian [CM16, tFZyZ16, LN10, PG15, Tan17]. Non-incremental [BE03]. non-interior [Lin16]. non-iterative [XCY21]. Non-linear [ALRÁY25, BGS24a, DL08, Dze15, Kaz24, KGN^{+}24, LLY22, Moh10, MMLM20, NNCN23, PLH20, WB17, YJX15]. non-Lipschitz [AV19, HLC15, RTD^{+}21, TSI20, Wan24, XWL25]. non-Lipschitzian [DHF21]. non-local [MV14, ZAH26, ZW12a]. non-matching [LPRÁ24]. non-monotone [GS21, GZ11]. Non-negative [AFG26, EJR26, DGST15, LV15, PLH20, ZCC26]. non-negativity [Zaf22]. non-normal [MT15, ST23]. non-oscillatory [ARTY20, LRY26]. non-overlapping [SB21]. non-parameter [XZZ19b]. non-perfect [VB91]. Non-polynomial [KS12, ZR17, BC17]. non-rectangular [AAD14, BF17]. non-separable [MM12]. non-similarity [FHAL15]. non-singular [NSM20]. non-smooth [CD25b, HLN25, HV22, KM24b, Lam09, LCZ26, MH21, ML22, MMLM20, Śmi06, TZH24, Tru24]. Non-stationary [DEC24, XH21]. Non-stiff [AM98b, BM22]. non-symmetric [EH24b, FP18, lLHNS23]. Non-uniform [BR07b, Ave20, DLR12, MS23a, RS26, TTP26]. non-uniformly [dS00]. nonabelian [VW08]. Nonautonomous [KS97]. noncentral [EGT25]. nonclassical [BEG25, Die08]. nonconforming [GG22, KP22, LN22, XSL22]. nonconvex [ALV20, BMR21, BE20, DCG25, GZD24, LB25, Liu14, LX24, LPXX19, MN23c, PW22, PDS^{+}23, SLL22, mTLbJlL14, WL24, You16, YWS20, YYZ22, YLL22, YZLC24, ZSWW25]. nondecreasing [Don12]. nondegeneracy [HBP13]. nondense [AR99]. Nondifferentiable [MS01b, Śmi09]. nondiscrete [AH12]. nonequidistant [PKR20]. nonequispaced [Ave20, KPT23, PPV09]. nonexpansive [CSI16, CSI17, CRN19, Den14b, GEA20, HMA16, IAH20, KAF18a, MWsC19, NUG25, PKC18, PDRG19, RZ16, SKK21, SIO20, SCS18, VA20, YAT20]. nonexpansive-type [CRN19]. nonhomogeneous [BKF20]. nonintegrable [Ben99a]. Nonlinear [AD00, Ber10, BKM03, BE03, CGL01, DOS03b, EHV19, Gra03, HL03, KV04, KW04, Kol04a, Kol04b, PL04, SME03, TRRD02, WHL24, WK93, ZY13a, Zil01, ALQ17, AAFL23, AAM24, AK19, AIA^{+}24, ATC16, AAU26, AAB13, ALRT16, AKB15, ABK22, AVCO26, AC11, Arg10, AGS20, AG23b, AAD14, ACH17, ACH19, AKKT16, AKQ17, Awa10, AB23, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BDN17, BMR97, Bai97a, Bai97b, BGL26, BAK25, BGR23b, BDV18, BC14, Bel99, BF18, BI00, Bhr16, BV99, Bog13, BZ24, BC26, BKL10, BRMG18, CCTV23, CCJ10, CC06, CGV22, CL06, CLT^{+}13, CW19b, CZZ^{+}25, CL10b, CH11, CC13, Che19, CDGC25, CN17, CNR15, CHMT10a, CLBT15, CRHTV24, DS21, DW12, DWC18, DD20, DD21, DNV25, DSI11, Das19, DK25, DMA09, DMA19, DBH21, Den14a]. nonlinear [DW21, DW22, DEM93, Don13, DXY18, Dzu13, Eft15a, Eft15b, EH24a, EL01, EGGSH13, FY13, Fan22, FZLL23, Faz23, Fra26, GA15, Gal18, GM06, Gar19, Gar24, Ghe15, GO20, GLdO09, GPAA14, Van17, GZ11, GZP18, Gu20, GCZH24, GLS^{+}18, Hai09, HJB18, HT23, HV15, HKE97, HZ95, HSTW14, Hey99, HST15, HM18a, HYJ20, HFW^{+}21, Ibr25, IAAH24, IZ24, JKNR13, JKM18, JLFL19, JWL20, JYLC21, JSZ22, JZ16, KS25, KKB16, KCBT21, KT07, KSB08, KLF17, KE16, Kub15, KB20, KKD25, KSJ26, La 17, LLY26, Le 92, LW95, LMMH11, LS15a, LY17, LHW17, LG17, LG18, LHM20, LZ21, LWLW24, LCZ26, LJ11, LKBF17, LZ16b, LZ09, LYL15, LF19, LDL^{+}19, LLC20, LSY^{+}23, LWZ23, LCZZ23, LCGH23, LZZ24, LJY24, LZT25, LA22b, lLXhL22, LV18, LV21, MJJ^{+}23, MD21a, MBJ17, Man21, MK98, Md12, MM12]. nonlinear [MG21, MA22, MG11b, MP13, MLM19, MS13, MH22, NBK17, ND21, NHP06, NKA24, NEMS14, NA24, OQ12, OL23, PV23, PED15, PSZ23, PLZ^{+}24, PS16, Pie96, QLZX11, QZG^{+}19, QXGZ20, RSKB17, RGJ10, RM11, RT12, RR98, RR00, RM13, RM26, RG10, SBA26, SPV20, SS11a, SFMK23, SN24, SW24a, SGS13, SA14, SY20, Shi96, SW19, SCDM20, SZ23, Sid20b, SSK23, SS24c, Śmi09, SS16, Sol11, SG18, SHGL22, SX00, TZ25, USAF14, Van19, VSA12, WG13, WH15, WZ15a, WZQT15, WCLW16, WGZ18a, WZ19, WWBM21, WLZ22, WX24, WSLL24, WQZH24, WGZ25, WCZ25, WHS20, WZ22b, WLY^{+}21, WC10, WWD^{+}12, WC13, WZ23b, XLW20, XWC26, XWX24, YIY22, YYL15, YH21, YZ21, YJ21, YCZ26, YJJ^{+}21, YYL26, YP09, ZR17, ZA20, Zha09, ZLG^{+}13, ZHT15, ZJWF18, Zha20, ZY21a, ZHFW21, ZYW23, ZWXX24, ZW20, ZCG15]. nonlinear [ZYW22, ZJ23, ZLS24, ZLLC11, ZTZZ19]. nonlinearities [JU22, Wan25]. nonlinearity [Liu21, MP22, ZQzS22]. nonlinearly [KLT95]. Nonlocal [TDKB24, BSF17, HKCW24, KK22a, LZ09, LHlLG24, ZS13]. Nonmonotone [ABKD23, CC13, FGBP21, GPP01a, KJ15, AA12a, AAN14, BKR18, BKA19, DHF21, DMT22, HWC15, LLL18, LM13, PZ20, SW11, WZC26, Ye22, ZLLC11]. Nonnegative [DLC14, WD95, ANI15, ANI^{+}17, BKDA25, DLDW21, GLL19, HNY^{+}18, KBA26, KLW^{+}23, LP08, LP12, LP13, MVG21, NK21, SXHZ20, wYN18, ZYBJ23]. nonnegativity [BHLZ21, BPR21]. nonnormal [AC94b]. nonoverlapping [ETY98, WRM17]. nonpositivity [Hua18]. nonrelativistic [CZ23, MY22]. nonresonance [ACE99]. nonseparable [DCG25]. nonsingular [BHLZ21, FS23, HWXC19]. Nonsingularity [HBP13, Kol06]. nonsmooth [BGRS12, DCG25, DEM94, GZD24, KJO23, LB25, LX24, MAS17, MC08, MN22, MN23c, PW22, PDS^{+}23, PSS22, SW24a, Śmi13, SG17, mTLbJlL14, TOY26, VPA24, WL24, You16, ZXL23]. nonstandard [CMP22, Gau09d, Hoa25, HSL19, PV00]. Nonstationary [Cha04, SKSS21, BCJ24, Cic20, Don12, HRY16, LD21, WQ23]. Nonsymmetric [BEGG91, AAPR21, AL09, BBS20, BC94, BEJ20, BRZ98, CS94, Cdv98, Don10, DBGB11, Ess98, GW24, HWXC17, HWXC18, LZ15, LL26b, Sad99, Sch17, TT21, YWX14, ZLT25]. Nontensorial [SVZ08]. Nonuniform [KPT23, PS01, AHP25, LCZZ23]. nonuniformly [dBD05]. Nordsieck [BCJ97, BJ02, BO02, CJ12]. Norm [BM94, ASW06, AS08, BGS24a, CKM^{+}25, CYM22, Coo09, FHC21, GGJL24, HYW20, HM18b, JZ25, LEK21, LPGL16, LZlL20, LRW25, LQ26, LCH20, LRL22, LZX22, LGZ25, LI10, MS24a, Meu97, Meu99, Meu05, MT13, MT14, MT19, MT23, Not12, OIM21, PT24, Szy06, TLD^{+}23, WC23, ZFZ19, ZZY24, dAFPR23]. Normal [CD99, Che01, CD96, FLV14, MT15, MZ99, MK97, ST23]. normalized [YCW^{+}19]. norms [Bea96, CMMP19, MT15, PT24, RS21, TM20, Wat92, Wat93, Wat94, ZJR25]. Note [CVA01, LS25a, LYH^{+}20, ASV23, And22, Cao12, CC15, DM14, FHH96, HLC15, Huc92, Iva15, KW96, LB93, LZ19b, MVV05b, Meu25, Meu26, Nie93, OdZdRV13, PGGCGF11, PRVI20, Pol10, Sab92b, Spa07, Spa20, TB25, VGV06, ZJ14]. notes [All08a]. notion [GS16a]. Novel [ALB^{+}18, GS16a, LWLT19, Liu21, RS25, SSN^{+}12, TDKB24, TRSI23, ALQ17, BN25, BX19, CLA11, CGSS24, DM98a, GPAA14, Hal14, KN23, KE16, LKQ23, LCZ26, MKG24, REM21, RR22, SA14, SDL^{+}23, TCOA19, TCW14, WLMA21, XW17, YJX15, YXL18, Yue25]. NPTool [LP13]. NS [HÇK26]. NS- [HÇK26]. NSS [DWC18]. NT [Khe17]. nuclear [LQ26, ZFZ19]. null [KMA13, RT20, ST22, Tak17, TM10, THT19]. null-space [ST22]. Nullspaces [Ver24]. number [Ber11, BGL07, Dea15, DBV23, DM92, KN21a, KN21b, LBF24, MDR23, XCLA15, YZH21]. numbers [AK26a, CL11, DL97, DP16, Kal00, LJ22, MZW20, MBM25, ZW22, ZLWQ09]. Numer [BRS92, CZ95, CP95b, Gau11b, Tov98]. numeralgo [RZ99b]. numerator [PP92]. numeric [EAB20, Mil17, MCMX20]. Numerical [AHP20, AA03, ALW98, AAH20, ADG10, AR13, AK16, BAEF25, BK97, Bin96, BV99, BEG25, BBB^{+}06, BV96, Bre99b, BGI25, BRS26, BH02, CE17, Car01, Car95, CGN03, CLTA10, CLT^{+}13, CC16b, CC18, Coo03, CCK04, DWZ14, DV01, DL03, DVJBN03, DKL15, DW97, DLLD17, DLDW21, FWC16, FW25, Fan26, FH15, FRS21, Gau12a, GNS22, GG98, GH09b, GHM23, GHM24, GST02, GO21, HJ18a, HT23, HCL21, HNSH09, Hof05, HS15, HFZ19, Hua21, ICR06, Ixa21, JP14, JS21, JL16, JW25, JLX22, Kam15, KGH14, KSCS07, KZ03, LJX25, LZ12, LD20, Lyn08, Maj13, Mat04, MJF09, MN23b, Meu99, Mic23, Mok16, MO04, NW04, NLT21, PP05, PED15, POP17, PPS25, PM25, PL04, PS20, Pis16, PS01, PS17, PV03, Rab23, RS20, RR20, SST92, SER02, Sch02, SLA11, SLLA15]. Numerical [SY20, SMR26, SS08, Smi97, SM25b, SA26, SZ21, TA24, TE03, TDKB24, Tem97, TBY13, Tom11, USAF14, Vig04, WYZ22, WHL24, WZS14, YLYZ23, YY24, ZM94, ZWWW20, ZJ08, ZLT^{+}17, ZYJY22, ZZY26, aZ19b, ZKD02, ZKD04, AAP24, ALB^{+}18, AB24, AG15, AAU26, AZ19a, ASS13, AA15, ADA07, ABI20, Ant18, AH18, AAD14, BPR22, BS17, BV93, BCM19, Bel99, BT14, BCM07, BuR24, Bic11, BC05b, BZS22, BR17, Bra07, BMR19b, Bru93, BBL22b, But10, CWZ13, CHYZ98, CM98, CM15, CC07, CP95a, CKS24, Che19, Cho17, CW25, CDGC25, CPN14, DS21, DDS93, Dam08, DD21, DNV25, DM98a, DM98b, DK25, DOR26, DS09b, DTI09, Des17, DH18, DGM26, DL97, Die08, DW15c, Dra00, EG10, Eba18, EGT25, FHL21, FT05a, FS01, GZ18, GMZ19, Gar24, Gau95]. numerical [Gau08a, Gau13b, GL12, Ghe18, GK20, God15, GPAA14, GO06, GL19, HÇK26, HH11, Hal14, Han96, HV15, HCBAEC23, HLTA16, HL17, ITA24, IJE15, IJSS16, JLP20, JXX^{+}23, KS18a, KN23, KM24a, KL17, KS20, KKO17, KXD25, KM09, Kuh13, KK12, KK16, KK17, LLZ18, LHZ20a, Lin16, Lin98, Lis25, LA22b, Luc06, LG95, LV21, MN23a, Mal21, MKBY19, Man10, MS23a, MFPG07, MB06, Mel10, MWZL23, MZ19, MNS23, MR96, MN11, NS24, Ngo23, OQ12, OKP21, PSS10, PSB26, Pat25, Pen98, PV98, PV00, PR25, PM24, Rab92a, RC14, RGJ10, RMT13, RhG15, Riz18, RB17, RR22, Rou25, RS25, RS26, SAC18, SSR25, Sal05, SHF15, SN24, SW14, SKD26, SDL^{+}23, SSL24, SS15, Sho18, SKTGR19, SR22, SR24a, SS12b, SYLT14, SZ20, TS18a, Tov97, Tov98, Uhl09]. numerical [VC92, VH92, VT10, VN18, WXQ20, WCHK21, WSZ21, WWBM21, WZ23a, Wan25, WZL26, WF23, XCY21, XLG22, YM24, YCZ26, YYLX23, YJX15, Yue25, Zaf22, Zag24b, ZuI19, ZSGJ25, Zeg25a, Zeg25b, ZYX19, Zha24, ZHW25, ZXL23, ZE10, dC16a, dC16b, Gau13a]. Numerically [Ver10, DGST15, JXNX24, MD21a]. numerics [BHH24]. Numerov [JSF13, Moh10, MK17, SFT03]. Numerov-Type [SFT03]. NURBS [CLaL00, KGD03]. Nyström [AMCM06, CQLY15, Con93, DDP14, DL18, FLY25, Jat15, LG19, Mon96, PFT98, PR25, SQG13, Som05].

Object [NPP04, CV92, OBAHK^{+}19]. Object-Oriented [NPP04, OBAHK^{+}19]. objective [CTY25, PLVB11]. objectives [Hua96]. oblique [HR05, SEG14, Szy06, Tru24]. Obrechkoff [VV07a, VV07b, KS20, SS15]. Obreshkov [Cor24]. observations [GP14]. obstacle [DIM22, PPS25, PS20, ZWX19, ZW20]. obstacle-avoidance [DIM22]. obtain [AGS08, GPAA14, GSF26]. obtaining [WLL12]. Octave [CDSV11, MBR21]. ODE [ABI20, Cor02, Cor24, Enr02, EY10, FGM19, KLT95, KPR03, MM99, MT04, MN92, gOM14, PLVB11, ST92, SGK^{+}99]. ODE-based [FGM19, KLT95, gOM14]. ODE-IVP-PACK [SGK^{+}99]. ODEs [AH21, AAP24, AHP25, AP26, ATM19, AK16, BM22, BP22, CE17, DK25, EH97b, Meh11, NB16, WZ23a, WWM21, AHKW04, AHKW05, BV09, MC08]. Off [BO02, MVG21, AMM17, EY10, KL17]. Off-diagonal [MVG21]. Off-Step [BO02]. Ohta [LZL20]. Old [Nie00, RR13, BRZ96]. Oldroyd [WSY12]. Ollongren [Ben99a]. once [GZL24, HFDSC24, PQS22]. One [Boy05, KS06, AVCO26, Arg09, BX17, BP19, BZS22, Bra07, CLA11, CK22, CGM25, CCD10, CJ20, DS09b, DHS09, hGzS17, GO06, HR07, Hoa25, Iid24, KKPT16, KNBGV18, KMS05, LXX23, Lie00, Lin05, MM08a, Mat15, MZW20, Nat07, OL21, Ros97, TS92, Ter22, WW19, WZ19, ZLW^{+}13]. one-block [BP19]. One-Dimensional [Boy05, AVCO26, BX17, BZS22, Bra07, CLA11, CK22, CCD10, CJ20, DS09b, hGzS17, Hoa25, KKPT16, Lie00, Lin05, Nat07, Ter22, ZLW^{+}13]. One-leg [KS06]. one-measurement [LXX23]. one-point [HR07]. one-stage [WW19]. one-step [KNBGV18]. online [HFW^{+}21, XWC26]. OPED [XTH07]. open [BPV13, JTW25a, JTW25b, Xue95]. Operational [KMH24, JBJB17, ROB17, ROB18, REM21]. Operator [FMD23, GS25, MDL15, TRRD02, Ahu09, AG19, AB23, AK09, BGR23b, BVV14, CT06, Dah93, Fue07, GNS22, HS12, KK22a, KXD25, KSJ26, LJX25, LX24, ML10, Pla99, PS17, RNT26, SSS14, SPB26, SKD26, TCOA19, TRCL24, ZWWW20]. Operator-splitting [FMD23, GS25, SPB26]. Operators [All03, BG03a, BS04, DV01, Fun01, GPS01, INR01, LPV03, ZL25, AA16a, AR16, AR20, AM21, AR99, AV19, AT19, AUD18, AHVR17, AKT15, BDIR18, BvLP16, BG13, Den14b, ER19, GNS22, GTA19, IR13, KADE18, KGMH21, LW95, LQ16, LJbL21, LMMD05, LG95, Mah10, Maz09a, Mic93, MP14, MP13, MRS07, MAK20, PKC19, RKM25, SPV20, SI18, Tam10, THS20, VdR13, XWL25]. optical [Wan15a]. optics [Kar07]. optima [KAL22]. Optimal [AN17, Amo02, AC19, BD03, BD04a, BCN^{+}16, BHS23, CKP19, CHYH24, CGW25, Che19, CKM19, DDT25, DJM^{+}18, FY13, FLH25, GWBC20, HÇK26, HWCR19, LZCW24, Lóc20, MV14, Mar04b, Mat01, Mic93, NT21, Plo94, Prz16, SFS23, Sid07, Van19, WPY25, XSL22, AJ13, AHR21, ALZ20, Bac21, BL23, BVV14, BCM16, BCMT18, Ber93, BM00, BFK22, Bla15, BV21, BHS17, CB16, CT93, CSC26, CN16, CGV26, CMMP19, CKKT16, CRHTV24, DL21, DOR26, DTI09, DO25, Ehr97, EDAH12, Fab16, FZ07, GA20, HW00, HMS11, KP16, Kha25, KS14, Kno23, LK20, LWZ18, LZL20, LAN18, LCH20, LW17, LL18, LP20, LW22, LZ23c, LCZZ23, Lóc18, LSSS15, MB06, NEMS14, NAA19, OMW21, OAR22, hPwL09, PS22, Ple12, QAS^{+}24, Ria16, SS24a, SLT20, SS24b, SKM25, SR24b, TQY21, TS18b, Van07, Ver10]. optimal [VS19, WZ22a, WZS14, WZ23b, YJ21, YF22, ZCT19, ZZ24, ZZZ20, ZZ18, ZLCW23]. Optimality [MSP25, GEP19, XP23]. optimised [GL15]. Optimization [CG13, Cse04, LPV03, MCG^{+}04, MS02, NM14, SR04, SMB02, VR04, YZX25, ZS03, AG23a, AA12a, AN17, AAN14, ARSS19, And06, And08, And10, And14b, And15, And18a, And19b, And22, ABM10, AE18, AFN16, AFN17, BBBC20, BBBC23, BGRS09, BGS24b, BMR21, BM24a, BE20, BN18, CWZ13, CW21a, CKS24, CGYZ19, CS24, CD25b, CTY25, CR23, CPN14, Cui26, DW15a, DW15b, DBH21, DEM94, El 18, EE18, FH25, GL12, GWL20, HHF22, HWC15, HLC15, HFW^{+}21, IS22, JR20, JZ16, Khe12a, KOK21, KC24, KJO23, Kri26, LZ18a, LLL18, LB25, LGW14, LS07, LM13, LL18, LGZ25, lLXZ24, LPXX19, MAS17, MR09, Md12, MS20b, MN22, MN23c, MPB16, NN26, Orb15, gOM14, ODL21, PL25, Pan26, Pea13, PSZ23, PYD23, PH14, PLVB11, PSS22, Rha22, Rog95, SL18, SV26]. optimization [SMZMA18, SS06, SDMMK18, SLL23, Tan20, mTLbJlL14, TTXZ23, TOY26, TYSY20, TD09, UPS25, UPY26, USM24, WJW14, WZ16, WYP23, Wan24, XP23, YIY22, YCL17, YDW26, You16, YZLC24, YY13, ZAH26, ZQL^{+}19, ZSWW25, ZWG18, ZFZ19, ZLL21b, ZDSY20, ZPX21, ZFH26, Ano95c]. Optimizations [DMC20]. optimize [CGV26]. Optimized [CGM25, GD15a, YHR26, CK22, DL09, LS14, RKMS16, RR20]. optimum [DHS97, ZQL^{+}19]. option [AAN25, AAA25, AWL^{+}24, Che22, Che24, CCLi16, KMV17, Rou25, SPB26, YYK24, ZLT^{+}17, ZZ22b]. options [BN25, CL10a, CDLW21, KN23, SMR26]. orbital [FR18]. Orbits [CCTV16, SER02, DFK97, LZ12, MO10]. Order [AG03, BM03, Che02, CJ04, DDP14, FG03b, Fun01, MV02, McL02, PW04, SFT03, YZY^{+}14, AD17, ADP26, ABS19, AH23, AK12, AR13, AHC05, AMA21, AABM08, AC17, AAII96, ABI22, ACE99, AMM16, ACSD16, Arg09, Arn97, AGG17, AKT15, AKQ17, BD10, Bac21, Bac23, Bac24, Bac25, BDL^{+}12, BN25, BJ98, BM22, BCM16, BCMT18, BKF20, Bia12, BK13, BW15, Bic24, BC05b, Bla15, BWC22, BR17, Bou17, Bra07, BV09, BGZ20, But98, BD98, BC99, BCJ99, BC01a, BI14, CM24a, Cat24a, Cat24b, CEX14, CHH^{+}20, CLT^{+}13, CW14, CQLY15, CC16b, Che22, CS22, CLPY23, CLQ26, CN16, CN17, CGL99, CJ17, Col92, CVX16, CHMT10a, CKKT16, CRHTTN24, CRHTV24, Cor24, Cui13, VV07a, VV07b, DHJJ10, DBTG24, DD20, DD21, DNV25, DK25, DMA09, DZ13, DW97, DB06]. order [DK00, lDzS21, DZH23, DGP15, EH97b, EM07, FAMA20, FW25, FHL21, FLY25, FT14, FGL19, FGR01, FR18, GB21, GH09a, GM20, GLLJ12, hGzS17, GZ20, Gau00, Gau09a, God15, HH11, Hai08, Hai09, HP18a, He16, Hea10, HDP18, HV22, HHLM23, HVMMV24, HCBAEC23, HKKN12, HMS96, Hoa25, HLTA16, HLV26, HWZ24, HLSLSL25, HZX20, Iva24, Jai16, Jai17, JL12, Jay21, JZF^{+}20, JWZ23, JHLL15, JWCZ21, KSV23, KCBT21, KNBGV18, KM24a, Kaz24, KD14, KMZ18, KLF17, KS20, KS12, KGH14, KADE18, KMV17, KM19, KXD25, KL06, KLZV95, KKA17, KK17, KKD25, LK20, LGL23, LMMH11, LW13, LW16, LR18, LRL19, LHR20, LHZ^{+}21, LlLZ21, LZOY22, LKQ23, LZM23, LPRÁ24, LCZ26, LM14, lLLVZ17, LWJ21, jLyLqW17, LWS18, LDL^{+}19, LW22, LWZ23, LZ23c, LZZ24, LZQ26, Loh22, LA22b, LSSS15, LRM16, LXP20, LV18, LC25, ME92, MWDD24]. order [MVVA08, MS23a, Meh11, MS06, MWZL23, MBM25, MClXzJe16, Miy24, Mok16, Mor25, MS13, MWWY13, Mül00, NBK17, NM26, NP18, NRV23, Ngo23, Nov26, NA26, ON26, OKB23, PKR20, PQS22, Pan18, PG12, PM05, PMM11, PP92, ROB18, RKMS16, RSKB17, RF23, RGJ10, REM21, RWB09, RZZ25, RS97, Rim25, RRZ21, RR22, Rou25, SSR25, SS26, San25, SS11a, SFMK23, Sas93, Sat24a, Sat24b, SAE19, ST17, SS10, SGS13, SA14, SMM25, SQG13, She15, SW19, SHD25, SS15, Sho18, Si12, SSK23, SS24b, SK25, SYLT14, SR24b, Sti18, SZ20, SzS21, SZQS23, SG23, TA24, TX19, Tem08, TZ24, TY21, TTP26, Ver14, VTV22, VLCL16, Vul97, WGK11, WKG11, WZ13b, WK13, WZQT15, WHZ^{+}18, WWBM21, Wan22, WQ23, WSLL24, WZZQ24, WHL24, WQZH24, WC24, WZL26, WLY^{+}21, Wri01, WWM21, WCD21, XT16]. order [XYZ14, XZZ22, Xu19, XH21, YZZL17, YZLZ22, YYL15, YCZ26, YZL25, YCW^{+}19, ZCT19, ZE12, ZYW17, ZWFY19, ZP17, ZYX19, ZHFW21, ZYQ^{+}21, ZF22, ZYW23, Zha24, ZQS24, ZXLF15, ZG12a, ZG12b, ZPEL25, ZLT25, ZLS24, ZAC26, vdHM98]. ordering [CKP22, Cia94, DF94]. Orderings [GGV02]. orders [DGST15, EGSHVN15, KFK^{+}24, van93]. Ordinary [AKKW03, But02c, BH02, DL03, DOS03b, ED05, BJ98, BWB24, But10, CQLY15, DJ10, DFJP10, Dze15, EG10, JT96, KL17, KMV17, KS06, MAH22, Mil19, NA26, RSKB17, Tem08, Tuo98, WCHK21, WYZ22, WK93, Wri01]. orientation [CV92]. Oriented [NPP04, CL96b, NK21, OBAHK^{+}19]. Orthogonal [Amo02, BBR03, BD20, BGVHN92d, BGVHN96b, But96, DFP^{+}10, Dra02, Fas23, FT02, GM21, GM22, GR01, LW04, MRS93, MRS07, Rec01, Ron08, AB24, APPR14, ÁFP07, BMA16, BDJ11, Bec96, BC00, Bia94, BF17, BLS92, BEGG91, BHJTM92, BP23, BZ91, VVV22, BGVHN92a, BGVHN92b, BV95, CGM12, CMR16, CCL16, CCHH23, CJTW96, CBGVN07, DM14, Dra96, DM97, DHM12, DGP15, Erb15, FLT09, FPP05, FMD18, FHV15, Gau09d, Gau13c, Gau15, Gau17b, Gau22, GCGVH92, Han96, HG93, HZ93, HSL19, HLSLSL25, KG23, KLA24, LGA^{+}00, LWAG08, LCVL18, LCW23, LI10, MS92, Mar92, MdR08, MM08a, MdR13, MVV25, Mil17, MD15, MA95, NØ96, Pas08, Ron92b, Sad05, SS23a, Tou98, VC10, VZ93, Wal06, Wim99, BV96]. Orthogonality [Van02, BCM10, CYIB12, FGBP21, MSM25, dR99]. ORTHOMIN [AZMJ04]. orthonormal [Caç10, FVV25]. orthonormalization [ZD15]. Oscillating [SFT03, EO94, NJ13, Sho18, SKTGR19]. Oscillation [Cai22, WWM21, GLLJ12]. oscillation-free [GLLJ12]. Oscillation-preserving [Cai22, WWM21]. oscillations [ddS24]. oscillator [CZLS18]. oscillators [DS21, MKBY19, YZ17]. oscillatory [ARTY20, BMR19b, Cai22, FYM14, FT24, FLY25, FR18, HS96, HS15, JH26, KXXW21, Kel07, KS20, KHM20, LRY26, LW13, LW14, Li17, LDW18, LG19, LWZ26, LWS18, Maj13, Maj18, NBJA17, OM26, Pen98, PM25, SW19, Vep08, WW14, Wan22, WWM21, XX16, XLG22, YZY^{+}14, ZuI19, ZYW17, ZH17, ZH23]. Oseen [DZS21, Far20]. Ostrowski [Eft15a, Eft15b, SS10]. other [DHS09, GN12, Kal00, SvF94]. outcome [JML25]. outcome-space [JML25]. Outer [BAS25, HL23a, MGL20]. outliers [AL18]. Output [MZG26, RTTH22, SAE19, TKSG23]. Output-space [MZG26]. over-relaxation [DLL^{+}24]. overcoming [DF01]. overdetermined [DEM93, Gon16, dAFPR23]. overlapping [Cal20, Hei06, LZ22b, SB21, ZJ14]. Overloading [TRRD02]. overrelaxation [HT16, Nie93]. oversampling [CL24a]. overspecification [DS09a]. overview [AGRT05, Mon01, RR08, dBGKR08].

P [SKTGR19, VV07a, VV07b, FGR01]. P-stable [SKTGR19, VV07a, VV07b, FGR01]. Pack [FH05, SGK^{+}99]. Package [Sha02, CDF99, CFR19, DCM^{+}13, DCMM13, DdAF^{+}20, FMS25, GHN19, Han94, LLZ18]. Packing [Mar04b]. Padé [VB92, ACO03, AH17, AHM21, ABV23, BL92, BBPV12, BC00, Bou03, BM96, Bre99a, BGVHN96a, BM23, But96, But02a, CL92, CAB22, CJTW96, DM97, Dra02, DF01, Ema96, EKM03, GCGVH92, GCGF03, GMT92, GM03, KL94, KC23, MC05b, MC05c, Mat96, MN01, PGGC97, PP92, PV99, Sab03, Sab14, TBA94, VB91, Van92, dDL92]. Padé-based [DF01]. Padé-Rayleigh [Ema96]. Padé-Type [ACO03, Bre99a, BGVHN96a, Mat96, PP92, Sab14, Van92]. Padua [CDSV11]. Padua2DM [CDSV11]. PageRank [BBJ25, GLC22, TLD22, WHS23]. pair [DGST15, HJ21, HJ25, KP96a, RSKB17]. Pairs [MV02, Har18, KFK^{+}24, KN21a, KN21b, MPC12, SQG13, ST21, Ver10, Ver14, Ver24, – DK15, dS24]. Pál [MP08, dBD05]. Pál-type [MP08, dBD05]. palindromic [GN12, KSW09]. Panel [ADL05]. pantograph [EED19, GL19, ROB18, SPS26, WZL26, YH21]. paper [ZW15]. Para [BHJTM92, CMR16]. Para-orthogonal [BHJTM92, CMR16]. Parabolic [BBHM03, LJW17, AAII96, AHC13, BPR22, Bog13, BSF17, BC26, BDF^{+}25, CCD10, CJ17, CG19, CJ20, DS09a, GPHHAPR18, HKE97, HYT26, HWC25, HZX20, JKC26, JZYY23, KM24a, KMV17, KLL10, LAN18, LYL25, LSW16, LZ23c, LZCW24, MM23, Moh10, MN11, PXH25, PLZ^{+}24, RNT26, RBN14, Ria16, SS24a, SKP20, SMM25, SN22, SR24b, TY21, YM24, ZYW22, ZS13]. parabolic-ordinary [KMV17]. parabolic-parabolic [BDF^{+}25]. Parallel [AC11, Bai97b, Bos21, CL00, CTS09, CB00, Fab16, GH09c, GHP^{+}00, HHST19, HPS97, HMA16, IM02, KD18, LS03a, NAR05a, SW10a, ST99, SME03, TBC^{+}23, Tir02, ZYGQ17, AL97, AAII96, AT12, BW13, BM12, Ber14, BF99a, BD00, CDW95, Con93, CK06, DDT25, DEC24, DM21, DZS21, DEM93, DZ19, DZ21, DZ22, DMW23, DZ24, Dum13, DDRT97, EG19, GL12, GZL^{+}25, GS26, HD18, JCF15, Lee94, LD21, LKKM15, MC05a, MHR23, MDR23, SKA23, SW10b, SSN^{+}12, Tai92, Wan15a, WD23, WDL23, WRM17, XWX24, ZD18]. Parallelism [KPT03, vSv94, ST92]. Parallelization [MMW20]. Parameter [AWL^{+}24, BS17, DVJBN03, KK12, LP20, MS02, OQ12, AJ13, ALJLYJ24, AABM17, ADA07, BK16a, BKA19, BG11, BG91a, BGR25, CCL18, CPZ14, CHYH24, CSC26, Cou15a, Cou15b, CCLi16, Dzu13, FJT94, GJV17, GO06, HT21, HWCR19, Hua20, LWLT19, LKK21, MXL25, MP14, Moo07, NAA19, RP26, RR13, SR25, SS21, TBY13, TV25, Uhl22a, Uhl22b, WC10, XZZ19b, YM24, YWWR12, YHS18, ZG09, ZWY22, ZL24, ZL26b, dFO11]. Parameter-robust [KK12, LP20, YM24]. Parameter-uniform [OQ12, ADA07, GO06]. Parameterization [Flo03]. Parameterized [YLY12, ATC16, AA09, CC15, DP16, DYW16, JTW25a, JTW25b, KK17, KKD25, LM17a, LM19, TDWL24, WZZ15, YDWL15, YYD14]. Parameters [SMB02, Alt21, BKS13, BV21, BL95, CRZ26, CHK14, DPP19, DPP22, DH26, Don12, DJM08, FZ07, IKMS24, gK25, LMM97, LZ25, MK98, MN23b, Plo94, PM24, TKS24, ZL22b, ZXLF15]. Parametric [DMD16, Góm01, LM01, Pop04, CLaL00, CLBT15, CRHTTN24, Góm99, HMdÁES08, JK18, KS18b, LSG15, Pop15, Pop18b, Pop21, SH21a, ZS19]. parametrised [CKO24]. parametrization [MB06]. parametrized [DF94]. Parareal [GKRS22, FT24, NG23, SJ14, SJW21, SWJ24, WLL26]. paraunitary [Tur94]. Pareto [PLVB11]. ParNes [GLW13]. Part [ABQ04, DMRT03, MS01b, DS15, IJ19, LPP21, SAHM25, AHKW04, AHKW05, HM22a, HM22b, MC05b, MC05c]. Partial [AAAS03, AK26a, ABM10, LHL11, LJW17, PV99, TRRD02, vLV02, ARS26, AP21, BCCP25, BT14, BD17b, BK13, Bre99b, Che16b, CDLW21, Che22, CCD10, EG18, FGJ00, God26, HZ20, HLJ25, JL21, JM18a, KW00, LLAL21, LG95, MN23a, MR96, Nor00, PP21, San25, SG23, TY21, TN10, VC00, WG13, WL00, ZWLZ24, aZ19b]. partial-differential [TN10]. partially [CM25, Hag25, TTA26]. Participants [Ano03b, Ano04d, Ano98d, Ano99e]. particles [ZLQT19]. Particular [GPP01b, Zag92, GS95]. partition [DZ19, DZ22, MS23c]. partitioned [DBTG24, Jay21, WQ23, ZSF18]. partitioning [CL96b, JCF15, ZCGS24]. partitionings [DDRT97]. Partitions [PS01, BF17]. parts [CW08]. passing [GSF26]. past [Ton25]. Pastur [GM21]. PAT [MP02]. Patankar [KM19]. Patch [MW24, MW26]. patches [HvD93, PN93, Str93]. Path [MP02, BHS11, DIM22, NR24b, WZ16]. Path-Following [MP02]. Patrick [KKB16]. Pattern [GPP01a, QZG^{+}19]. patterns [Li95]. PBS [ZWXX24]. PC [Røn92a]. PC-fractions [Røn92a]. PDE [AAA25, AFN16, AFN17, BGS24b, BM24a, CW21a, GPHHAPR18, KN23, MS20b, NPS09, Pea13, PZ20, Sla06, SG10, ZAH26, ZS13, iV12]. PDE-constrained [AFN17, AFN16, BGS24b, CW21a, MS20b, Pea13, Sla06]. PDE-informed [BM24a]. PDE-W-methods [GPHHAPR18]. PDEs [ATM19, AAU26, AAIT94, BWC22, BC26, BIMR19, BK08, FMS25, GPHAPR24, HFDSC24, KM24a, MHR23, MM23, RNT26, TTP26, TCW14, USAF14, YM24, AAII96, DL09, Sim98]. peer [SW10a, SAHM25]. Pell [SR24a]. Pellet [Mel14]. penalized [PV23, ZLLC11]. Penalty [JR20, Fan26, FLH25, LW25, MAS17, PPS25, PSS22, SFS23, Ton25, WZZQ24, ZW20]. penalty-based [Fan26]. pencil [JKK^{+}07, NN26]. pencils [AL09, CG07a, CsL24, DPP22, KPS14]. Penrose [Hua21]. penta [JJ13]. penta-diagonal [JJ13]. pentadiagonal [JL16, JXNX24]. per-iteration [KD14]. percolation [CLA11]. perfect [BM19, MVVV24, VB91]. perfectly [DL21]. Performance [CMWP20, HD18, BOP98, CEK21, EH97b, LS14, Moo07, PSS22, Sha19, WLL12, Wan15a]. peridynamic [JH22]. period [CCTV16, CZLS18]. period-doubling [CZLS18]. Periodic [FG03b, MO04, Spr98, AMM11, AKQ17, Ben99a, BHS14, BH17, CCC25, CKP97, CBGVN07, CZLS18, FJ25, GL21, HM19a, Jia20a, JW25, LDN16, Lor95, MO10, Plo93, Plo94, PR25, SS15, SJ14, SWJ24, Thr92, WSZ21, YC22, YX11, ZZY26, – DK15, van93]. Periodization [KSW07]. periods [Hol98]. permeability [NAHZ21, ZDZ26]. Perron [CVLX19, EHNR23, Kol06]. Perry [WHS20, YHS18]. persistent [Fly22]. perspective [BS25, GZ18]. Perturbation [Boy05, ZW22, ZLWQ09, CCL16, CGL99, El 18, LCVL18, LXX23, Vul97, XXW17]. perturbations [BRZ18, CHS19, EGG08, LGM25, MW98, Pla99]. Perturbed [Bog02, ED05, Leo03, Leo07b, Leo08, ABL… 12, AE09, AA09, BPR22, BS17, Bog13, BDS00, CEX14, CXL16, CL93, CJ17, CG19, CJ20, CVX16, Das19, EA12, FM16, GO06, GO21, HSK20, HCXL20, JKC26, Kow00, KK12, KK16, KK17, KKS22, KKS24, KKD25, KSJ26, LS15b, Lin09, LRZ12, LLC20, LZ23b, LZ23c, MWDD24, MS24a, MN11, NMM18, NV21, OQ12, PM24, Rad08, RS20, RNT26, RBN14, SN22, ST99, Śmi13, Tem08, TKS24, WJC25, YM24, ZL22b]. Peter [Bre19]. Petrov [MS14b, BX19, MCW22, MFPG07, dFO11]. pfaffians [CHHL18, LCHH21]. PH [FKMS01]. Phage [Car01]. PHAM [VSA12]. Phan [PP05]. Phan-Thien [PP05]. Phase [AA03, MKO04, BZS22, HHST19, Lin98, LWS18, LL20c, LZZ23, Lyn08, McL25, Met19, SS15, SKTGR19, Tur94, WXQ20, XCY21, YK22, YZL25]. phase-field [LL20c, XCY21, YK22]. phase-lag [SS15, SKTGR19].

Phenomenon [Bre04, ALY22, DF01]. Philip [Gau95]. Phillips [HS12, LQ16]. PHSS [HWXC19, WDL16]. physical [DR07, LHlLG24, XCLA15]. physics [SPS26]. Picard [CZ23, GEA20, NUG25, SAE19, Wan22]. Picard-type [GEA20]. Picone [BT14]. picosecond [CHH93]. PIDE [Che24]. PIDEs [TTP26]. Piecewise [CF96, DOR26, Maz18, MM11, WZQ17, All18, And97, BCM19, Bia94, BFK^{+}09, CG07b, CR23, Ell93, GH22, GH23b, GWY26, Hag25, HL23b, Maz05b, Maz09b, SMK14, SSL24, SS94, ZY21a, ZSLZ24]. Piecewise-linear [MM11]. piecewise-smooth [GH22, GH23b]. PIES [KZ21]. Pillow [WH04]. Pipeline [OM18]. pipelined [CH26]. Pitaevskii [CCJC18]. pivot [KH18]. Pivoting [TS92, KM13, ZWLZ24]. pivots [KM13]. planar [AL23, LSX10]. Planck [LD20, MM22, SSYL20]. Plane [MST03, NS01, SS01a, BPV13, CM15, For21, For22, HAN24, MP08, PR93, Reb97, Ree92, SS01b, Wal07, Yua21]. planes [GJV17]. planning [DIM22]. plate [BKPS93, CM24b, Han93, HM07, Lev95, LCW23, Mot14, Pow93, WWD^{+}12, YXL18]. play [Das26, MT15, XLL26]. plug [XLL26]. plug-and-play [XLL26]. Plus [FG03a, HR03a, LZ25, MVV05a]. PMHSS [CW21a, LYW14, ZM16, ZZZ22]. POD [SR24b]. Poincaré [Pas92]. Point [ACO03, BO02, DZ01, DHL^{+}04, DH04, GF02, IW04, LS03b, WSY04, AG15, AK26a, ASW06, AABM17, AJMP11, AVCO26, AM18, AM13, Bac18, Bac20, BBS20, BMR97, BKPS93, BBO21, BD17a, BE17, Bic24, BF99a, BRZ19, BM23, CWZ13, CSI16, CSI17, CSI18, CAB22, CXC25, CZ14, CC15, CCL18, CD99, CM16, CRN19, Cro92, DL21, DS09a, Den14b, Dey23, DGL06, DYW16, EMS25, EAB20, ER19, Fan15, tFZyZ16, Fra26, GWBC20, GM23, GCGVH92, GLM15, HT21, HS20, HR07, HPS13, HWXC17, HM18c, HWXC18, HWXC19, HM19b, Hua20, HCD26, Ibr25, Ihs07, IS22, JLMP16, JWCZ21, JCH23, JM18b, Kac18, KT07, KAF18a, KAF18b, KM17, Khe12a, Khe12b, Khe14, Khe16, Khe17, KH20, KOK21, KKO17, KGG25, LK20, LP18, LM17a, LLD23, LZ15, LZ18b, LXZZ21, LS15b, LRL22]. point [LYY12, Lis25, LS07, LCL21, Lo97, LSSS15, LZ22d, LM25, MWsC19, MS20a, MP00, MN17, MR09, MHR23, NHP06, NEY26, OMU^{+}24, Osa92b, OO22, PKC19, PM22, PKC18, PSWE23, PR14, PH20, PPPN23, PH14, PN21, RGJ10, RZ16, RT20, RT22, RTTH22, RCW22, RB17, Sab91, SKK21, Sal17, SM17, ST22, SMZMA18, SD20, SIE16, SI18, SCW17, SSA25, SS94, SC18, TAM21, Tak17, Tan20, TH19a, TH19b, TBA94, THT19, UPY26, VT10, WO00, WZ11, WK12, WZ13a, WK15, WZ15a, WDL16, WK16, WK17, WGZ25, WCM94, XM16, YIY22, YDWL15, YZLP16, Yan17, Yan18, Yan22, YHR26, Zah09, ZRZ11, ZFC18, Zha26, ZH19, ZAC26, ZWY22]. point-based [PSWE23]. Points [GC04, Sza03, AHL20, AHS22b, AHS22a, AH13, AT17, Bag16, CDSV11, CG05, CDG23, CL93, zDYG18, DF94, DFP^{+}10, DPP19, GI15, IAH20, IKMS24, JB22, KST21b, MM19, PKC18, PG12, PR10, PDRG19, SX96, SL15b, SL15a, SRK^{+}25, Van07, WZZ15, ZH17, dBD05, KST21a]. Pointwise [LZ18c, WJC25]. Poisson [ABI22, BQ19, BW15, BM09, CYWH25, JHLL15, KS26, KN18, LP12, Lee94, LGL23, MFBB23, MZ99, MClXzJe16, MDH16, NT21, Prz16, SSYL20, WZZ26, YL22, ZAH26, ZWW21]. Poisson-corrupted [LP12]. Pol [CZLS18]. Polak [BKG15, MA22, YZBJ21]. Polar [KZ03, ÁCL11, BW15, HN16]. pole [Car92, Van92]. Poles [BBD03, Bal11, BM00, CM92, EAB20, Gre96, Van07]. pollution [CV15]. Pólya [BG91a]. polyadic [VVL25]. Polyak [BKG15, MA22, Wan18, YZBJ21]. polygon [FJT94]. polygonal [Gha24, Kar09, KJC18, LGC24, LCGH23]. polygons [AW23, FP20, LHW13]. Polyharmonic [HL02, RR08, Flo16, KR07]. Polyhedral [HV98, Wat92]. polylogarithm [Vep08]. Polynomial [AM13, BC14, BFGM03, BCGVS11, CR03b, CPV04, DV01, FM19, FT02, INR01, LWK12, LM12, MC05b, MC05c, Maz12, Ple03, Rec01, RZ23b, VM17, WDY04, ABL… 12, AL23, BDL^{+}12, BLW09, BNN16, BM00, Ber10, Bin96, BF99a, BF00, BC17, Caç10, Cam19, CGW25, CJTW96, DDT25, DF93, DY93, EEM20, FM16, GGN14, GH10, GM92a, GH25, GI15, GNH10, GS16a, Her96, HvD93, Hua96, HV98, ICR06, IN21, Iva17, JK18, JS15, JNW92, KAL22, KS12, LV15, LKW17, LZ22a, Man07, MMV17, MMV19, MS17, MSS18, MK94, MS11, MG91, MG94, NKM25, Pan26, PHI98, PS06, PMM11, PR14, RVF07, SST92, Sid94, SS23a, SR24a, Sol23, SSH20, TS18a, TV25, Yan18, YHZ20, ZR17, dB07a]. polynomial-iteration [Yan18]. Polynomials [Ano05b, BBR03, BL04, BD04b, BV96, CD01, DG05, Des17, Dra02, DJ02, FH04, Gau15, Gau17b, HLM04, LW04, Mas05, Maz02, PG05, Rec01, AB24, ALW98, APPR14, All18, ÁFP07, ACG20, ADGP15, ACH14, BDJ11, Bec96, BC00, BD20, BLS92, BEGG91, BHJTM92, BZ24, BP23, BZ91, BM96, Bre99a, Bru93, BV95, But96, CG17, CMP21, CGM12, CMR16, CCHH23, CW21b, Cox93, CBGVPP09, DFP^{+}10, DM14, DB06, Dra96, DM97, DD99, DL01, DJM08, DJ18, DHM12, DGP15, DLR24, EDAH12, Erb15, FS20, FGM19, FKP06, FPP05, FMD18, FHV15, FH00, GAM24, GL07, Gau08b, Gau09c, Gau09b, Gau09d, Gau11b, Gau12b, Gau13c, Gau18, GM21, Gau22, GM22, GN12, GS14, GCGVH92, Han96, HSL19, HLSLSL25, IZ24, IN95, Ise96, JBJB17]. polynomials [KG23, KJG23, KLA24, KSCS07, KSCS08, KW96, Kou07, LMV25, Leo07a, LGA^{+}00, LWAG08, LL14, LI10, LPP21, LR14, Man17, MS92, Mar92, MdR08, MM08a, MdR13, ME95, MVV25, MEJS19, MS23b, Mil17, MD15, MA95, MS11, NØ96, Not95, Ost07, PM05, Pet95, PCDH20, QW08, RKM25, REE26, Rim25, Røn92a, Ron92b, Ron08, SSP15, TTV21, TV19, VC10, Wal06, Wal07, Wim99, Wim00, ZAGD22, dFG93, DH18, Gau17a, Gau19]. polynomiography [QAS^{+}24]. polytopal [CEZ26]. polytopes [PT18]. population [OKP21]. Porous [AA03, CGN03, ASZ23, Gha24, MSMS12, SLLA15]. portfolio [LXP20, PL25]. portrait [HPS97]. Posed [CR02, AMR23, AC11, BH11, CMRS00b, DNR15, DNR17, FH00, Han96, Han94, HR14, JRS09, Jia20b, Mok16, MRS06, MRS10, NRS12, OR17, Pla99, RRS09, RR13, RSZ20, SGJ15, WZQ17]. posedness [CDG23]. Position [Sei98]. Positive [All03, ABQ04, BD25, GPS01, Leo07a, Str02, YLD11, AR16, AR20, AA16b, AL97, BMA16, CR96b, CLMM05, CWHL20, CV15, DFD23, tFZyZ16, GMP92, GRAST23, Hag13, Har18, HM18a, HV12, LG17, Liu11, Mas95, NP18, PG15, Str97, Sun94a, Sza92, Tan17, VGV06, WMCW21, XM16, ZM94, Zha15]. positive-definite [PG15, XM16, Zha15]. positive-indefinite [CWHL20]. Positivity [Not08, JLWZ25, LLY26, LCW25, MPR22, OKB23, YH24]. positivity-preserving [JLWZ25, LLY26, LCW25, OKB23]. positron [BG11, RVF07]. Posteriori [ABMV03, SS03, ASGJ^{+}20, AKPW05, Bac24, Bac25, COSE22, Das19, DBGB11, HWC25, HCXL20, JL15, KL17, KKS22, KKS24, KKD25, LCH20, LRM16, LKK21, MCW22, SR25, SS24a, She15, WZ22a, WBY25, XC25, ZLH21, LGM25]. Postface [Meu19a]. postprocessing [Sar06, ZL23, ZH25]. Potential [LM04, Lui02, Rog95, BX17, EAGS20, KZ21, Rei98]. potentials [Lam09, LZ23a]. Poussin [The12, TV17]. Powell [Rip93, WD95, WD96]. Power [Bou06, Maz02, SVZ01, AAA25, GL20, KP26, Sab92b, Wan25]. powering [GI10]. powers [DM92, GLM15, Sid20a, SSC25]. PPM [RZZ25]. PPS [SM21]. PPS-methods [SM21]. PQser [CFR19]. practicability [USM24]. Practical [BKL10, Len93, SC03a, CY10, Mot14, ZL07]. practice [CFS21]. Prandtl [DOT21, CM01]. pre [GV99, Mic91, MRU91]. pre-filtered [GV99]. pre-wavelets [Mic91, MRU91]. Preassigned [BBD03]. precise [EED19, Wil12]. Precision [Rev03, AR26, BBPV12, BGZ20, FLMR99, Gau09d, Gau11a, Gau14, GS14, Joh15, Joh20, KFK^{+}24, KMS23, MBM25, OC24, RR23, ST26, Sv95, SA26, Tsi07, WLL12, YZH21, Zag24a, Zag24b, Zag26]. Preconditioned [BBS20, LRGH02, LM25, MLM25, MR96, WZ13a, XWY19, ZZB20, AAAA^{+}18, AFN16, AFN17, BBC11, CK24, CXC25, CZ14, CDLW21, EG19, GM20, HYW20, HFDSC24, HM19b, LZOY22, LZ18b, LZL23, Meu99, PNW17, PV98, Rah11a, RWTW19, SA26, WHS23, WPL18, ZZ24, ZCG15]. Preconditioner [Meu02, ALJLYJ24, BW13, BOR23, Che16b, CL21, CGSS24, tFZyZ16, HN94, HWXC17, HWXC18, KM17, LM19, LAN18, lLHNS23, LC21, MS20b, NG23, RCW22, RR98, SM17, SCW17, Tan17, XM16, ZD17, ZFG18, Zha26, ZWY22]. Preconditioners [LXZZ21, Lin01, LS03b, AL97, ACSD16, BBO21, BBL23, Cao12, CW21a, CMM17, CT93, CO94, CX20, DNR17, FNS19, FC01, HW18, Huc92, KLL25, LZ19a, LZ19b, LN10, LW12, MMGH17, MHR23, NMM18, RW24, WSLL24, ZRZ11, vdMRS06]. Preconditioning [GZL24, GP99, MT18, MS02, SC03a, SMB02, ABG97, AK09, BCW13, BRY14, BL23, BDS25, BM12, CGV22, GNS22, GV00, HPS20, JCF15, JM00, LLM25, LP20, Mor17, ST18, SH21a, TH23, TA13]. predator [GA20]. Prediction [PV03, LL22a, MBG19, PV99]. prediction-correction-based [MBG19]. predictions [BM24a]. Predictor [IW04, Bra06, DZW17, Khe14, MJH17, NA26, SMZMA18, SD20, WO00, YZLP16, ZLS24]. predictor-corrector [NA26, SMZMA18, ZLS24]. Preface [Ano98e, Ano99f, Ano00b, Ano01h, Ano02d, Ano03c, Ano05c, BCN06, Bre25, BJP^{+}19, BMV25, BSSJ14, But02b, BJPW10, CP00a, DEV97, DRZV07, MY04, Meu19b, The97, RZ99a, RZ23a]. prefiltering [PZL15]. prefilters [Ehr97]. preprocessing [ddS24]. Prescribed [CG03, ANI^{+}17, CM92, Gre96, PP92, Van07, ZYBJ23]. Prescribing [TM14, Meu20]. presence [ART19, BRZ18, CEK21, CFRV23, Gau12a, Gau13a, MZ99, Zas22]. Presentation [Kub15, BZ91]. Preservation [MO10, DS21, LZ23a, TA26, Tur26, WS24b]. preserve [CL96a]. Preserving [CP01b, LM01, ASGGRG23, BHLZ21, BH17, Bla15, BWC22, BV21, BB14b, Cai22, CS22, CMWP20, DW21, DCW23, DLL^{+}24, Ell93, Guo16, GJVW24, Hoa25, HKCW24, HLJ25, HWZ24, IJ19, JLWZ25, KK00, KNBGV18, KMH25, LLY26, LN95, LG19, LHlLG24, LCW25, MZZ25, Miy24, OKB23, SZQS23, SLZ25, TZ24, WZVJ22, WLW^{+}24, WWM21, XF25, YZZL17, YZLZ22, YK22, YSLL23, YH24, ZLQT19, ZYQ^{+}21, ZYJY22, ZZ22a, ZY23, ZJJW24, ZQS24]. Pressure [LN22, HLS10, PV00]. Pressure-independent [LN22]. pressure-release [HLS10]. prewavelet [HP18a]. prewavelet-based [HP18a]. prey [GA20]. price [AAA25, CCLi16]. pricing [AAN25, BN25, CL10a, CDLW21, Che22, Che24, CS12, KMV17, SPB26, YYK24, ZZ22b]. Primal [Khe12a, Alt21, AABTB23, BC16, CWZ13, CC06, CXC25, CL19, DGL06, HLN25, HHF22, JWCZ21, MBG19, MLM25, SMZMA18, WZ11, ZXL25, ZFH26]. Primal-dual [Khe12a, AABTB23, BC16, CWZ13, CXC25, CL19, DGL06, HLN25, HHF22, MBG19, SMZMA18, WZ11, ZFH26]. primary [GI10]. primitive [PMO05]. principal [CG07a, HNY^{+}18, XLG22]. principle [ASGGRG23, BD17b, CGHH21, CS22, LEK21, She00, SZQS23, SZ24, WZC23, XF25, ZYQ^{+}21, ZY23, ZQS24]. principles [PZ20]. prior [LZ22b, TRCL24, XLL26]. priori [BGR23a, KKS24, KKD25, LKK21, SKP20, WZ22a, ZA20, LGM25]. Probabilistic [Stu97]. probability [GM21, HDL23, LD20, WZ23b]. Problem [ABMV03, BBHM03, Bog02, CGN03, CSFC04, CN01, LV01, Mar04a, SR04, SME03, ZL25, AMM11, ADP26, ALV20, ADA07, ACE99, ARSS19, BKDA25, BAV18, BEM99, BVV14, BBCS21, BH92a, Bia12, BFKM20, BFK22, BuR24, BZS22, BCJ24, BM24b, BGVHN92c, Buo17, BGZ20, CMD19, CEX14, CXC25, CLGS17, CCL18, CC18, CZM21, CS24, CFRV23, Cor91, CKM19, CL13b, CGN22, CRV91, CHK14, DLL13, zDYG18, DD20, DD21, DS09a, Den14b, DWX17, Don16, DJG18, DZ22, DZ24, DLDW21, El 18, ER19, FAMA20, Far20, GD15a, GHC15, GWL18, GPGC98, GD15b, Gha24, GLG26, GO06, GO21, HT16, HT21, HLN25, HJB18, HM22b, HZ95, HYT26, HFZ19, Hu22, HNY^{+}18, HM19b, HCD26, IJSS16, IUM^{+}19, JHLL15, JM00, JN99, JRRS08, KL25, KZS21, KAF18a, KAF18b, KPC20, Kha25, Khe14, Khe16]. problem [KMA13, KPS22, KS18b, KPA20, KLZV95, KLW^{+}23, KSJ26, Kun05, KP22, LMV00, LLZ94, LLS11, LY18, LDX23, LWY26, LL20b, LCL21, LJ22, LZ23b, LG95, LXP20, MWsC19, MM00, MM19, Med10, MS24a, MZZ25, MXL25, Mez22, MAF25, MT93, MK17, ML10, NPR08, OIM21, OMW21, OR17, PKC18, PL99, Pan20, PV23, PYD23, PPPN23, PM24, Rab23, Rad08, RS06, RZ16, RT20, RT22, RTTH22, RT24, Ria16, RS93, RVF07, Saa23, SCD^{+}21, SR16, SW24a, She15, SCF23, SMNZ20, SFZ22, SW07, SR24b, SG10, SPS26, Str09, SC18, TKSG23, TAM21, Tak17, TPLB22, THT19, VVL25, VPA24, VPA25, Vul97, Wan17, Wan18, WZ22a, WL25, WJC25, WBY25, WZC26, Wit96, WK20, WL22b, XHZ07, XZZ22, XZW13, XH25, XWX24, Xue95, YZL20, YPL21, YZ23, YDW26, YZ11, YLY12, ZE12, ZM94, ZBDK23, ZW12a]. problem [ZW14, ZZH15, ZYGQ17, ZZ19, Zha19, ZJZ20, ZZZ20, ZLWZ21, ZLH21, ZL22a, ZW22, ZZ22a, ZLZ23, ZWX19, ZW20, ZY13b, ZLT25, ZW15, ZZ18, ZLCW23, lZqL25, lZfLqL26, Zhu15, ZWY22, ZFH26, ZD18, CGR12]. Problems [AKW02, AKKW03, BD02, Bel03, BCV03, CR02, EGSV04, FG03a, FG03b, GC04, GLC22, HSZ03, Kol04a, LS03b, MT04, Nac03, The97, PV03, SK04, TC05, Tsu02, WSY04, ALQ17, AGA24, AAA^{+}18, AS10, AK26a, AR13, AHC05, AMCM06, AMR23, ÁCL11, ABL… 12, AE09, ABI20, ABI22, ALZ20, AJMP11, AAH18, Anh19, AT21, ATT22, AA12b, AM18, AM13, AHKW05, AKPW05, AK09, AFN16, AFN17, BPR22, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BBS20, BMR97, BD06, BZ13, BS17, BLW09, BGR23b, BBBC20, BD17a, BCK06, BGRS09, BGRS12, BBL23, BF18, BAEF25, BH11, BGS24b, BPV13, BFK11, Bic24, BM97, Bla15, Bog13, BEG25, BC16, BRS09, BE98, BS14, BMR19b, CSI16, CSI17, CSI18, CCZ23, CR99, CMRS00b, Cal20, CB16, CW19a, CW21a, CK24, CKS24, CZ94, CZ95]. problems [CZ14, CC15, CR20, CW14, CLWV15, CD15, CM16, CFL19, CGYZ19, CLWH20, CJK22, CD25b, CYM22, CRN19, CCJ99, Cho17, Cho16, CW17, CDD21, CL93, CGL99, CJ17, CGV26, CVX16, Cor24, CDS20, Cro92, CsL24, DLL12b, DLL12a, DNV25, DCG25, DG17, DK25, DOR26, DMA19, DBH21, DC17, Den14a, DHF21, DO25, DMT22, DEM94, DBAE09, Don10, DBGB11, DYW16, DI11, DLLD17, DSS14, DR12, DNR15, EMS25, EDAH12, EE18, EH97b, EP97, EM07, ETY98, Fab16, Fan15, tFZyZ16, FYYW19, Fan22, FG07, For93, FR18, FH00, GEP14, GEP16, GEP19, GHN19, GX19, Gha16, Gha18, Ghe13, GH09c, GS16b, GPHHAPR18, Van12, GL15, GRT97, GK21, GZD24, GZL^{+}25, HÇK26, HS20, Han96, Han94, HHF22, HKP25, HR14, HVMT17, HMA16, Hie18, Hie19, HRAH22, HKPW19, HL23a, HLV26, Hua96]. problems [HLZ14, HWXC17, HWXC18, HRY19, HWXC19, Hua20, Huc92, IDS16, IS17, IMT23, JKM18, JL12, JRS09, JR20, Jia20b, JWCZ21, JSZ22, JLX22, JCH23, JTW25a, JTW25b, JK19, JA22, Jón93, JS23, JM18b, KW00, KS25, KST06, Kar10, Kar13, Kar15, KJC18, KLT95, KM17, KMZ18, Ke21, KS20, KS12, KD18, Kim26, Kno23, KMS05, KS23, KJO23, KSW09, KK12, KK17, KKS22, KKD25, KZ21, LP18, Li97, LWM10, LHZ10, LZ14, LL16, LY17, LDL17, LM17a, Li17, LWZ18, LL20a, LZL20, LYH^{+}20, LLX20, LlLZ21, LZOY22, LLY22, LBF24, LPRÁ24, LJ11, LZ15, LAN18, LZ18b, LZL22, LZL23, LDC10, LS15b, LCH20, fLxX12, LRL22, Lin05, LRZ12, Lis25, LHL11, jLyLqW17, LW17, LWS18, LZ18c, LZX22, LGZ25, LJbL21, LGM25, LRM16, LKK21, lLXZ24, LS25b, LPXX19, LZ22d, MS20a, MsC20, MSCB93, MN17, MK98, MMV19]. problems [MMGH17, Mar93, MRS93, MO19, MB06, MZW20, MG20, MG21, Mia19, Mil20, Mil18, MS20b, MS14b, MSP25, MN08, MSS11, MN22, Moo20, MRS06, MRS10, MPB16, MG18, MN11, MH22, MS24b, NM14, NBJA17, NRS12, NJ13, NMM18, NV21, Nie93, NEMS14, NK21, NR12, NA24, NA26, NEY26, OQ12, OMU^{+}24, OC24, OdZdRV13, OLB94, gOM14, OAMA22, OR24, PKC19, PXH25, PT24, PGSU25, Pea13, PT17, PLH20, PDS^{+}23, PS25, PK22a, PK22b, Pla99, PPR15, PR93, PLVB11, Pot19, PSS22, QL12, RFS23, RKMS16, RR20, RGJ10, RMT13, RBN14, Ree92, RTCL21, RRS09, RR13, RSZ20, RAH11b, RWTW19, RCW22, RR98, RR00, RSCH^{+}19, RB17, RALJ26, Rum14, SBN25, SPV20, SM17, SEG14, ST18, SV26, SS24a, SWS22, SM21, SIE16, SI18, SCW17, SHLY18, SDW24, SW19, SS15]. problems [Si12, ST21, SN22, SKM25, ST99, SDMMK18, SJ14, SG17, SG18, SXHZ20, SWJ24, SA23, SSSS22, Stu97, SS94, SLL22, SW26, TS18a, THF21, TQY21, TQW24, TO21, TG20, TYSY20, TBY13, Thi93, TH18a, TH18b, TH19a, TH19b, TVC20, TTLD20, TDC21, TRSI23, TLR^{+}25, Tom11, Ton25, TKS24, TTA26, UPS25, UPY26, Van07, VTV22, VMMD21, VT10, VN18, VS19, WZ13a, WJW14, WDL16, WZQ17, WCW20, WL22a, WYP23, WL24, WPW25, WG25, WGZ25, WCZ25, Wat93, WCM94, WZS14, WPL18, Wu22, XM16, XXW17, XCD23, XWL25, XZP^{+}20, XP23, YWWR12, YDWL15, YH21, YF22, YWYN22, YCZ26, YSXY19, YZLC24, Zah09, ZA20, ZH22, ZSGJ25, ZZ24, ZWL26, ZRZ11, Zha11, Zha15, ZZY18, ZFC18, ZL22b, ZZ23, ZSWW25, ZH19, ZLV17, ZV19, ZV21, ZZLV23, ZS08, ZLWQ09, ZZY24, ZLLC11, ZWW25, vSv94, vdHM98]. procedure [ABT07, BGR23b, CPRZV23, FVV25, HS96, LSY^{+}23, RZ16, Shi96, ZD15, ZYGQ17, ZYJY22]. Procedures [Bre02, DLL04, AG13, Bre00a, VVV22, CP93, CP95b, SRK^{+}25]. Proceedings [BV96]. process [BRZ19, BE98, Dax17, FJ96, GGN18, GLRSG08, HVM15, HVMT17, JL21, NUG25, PRVI20, PDRG19, Sid17, Ste24, SCS18, TAM21, TA26, TH19a, WZ26, YLYZ23]. processes [BM09, DMYT23, GKL21, HR07, SKSS21]. processing [ABK22, ABKD23, BKPS93, BH11, FM99, OB16, YAY26]. producing [JJ24]. Product [BG03a, GGV02, ALL25, ACM93, BH01, CRZ26, CE94, CK20, DGP15, FT05b, HPS13, Hua21, HZ24, Jia06, LN10, MS15, OB16, Rab05, SBN25, SP21, XQZ24, YHZL21]. Products [DL01, Str93, CL99, HL15a, HHLM23, MMV19, Ron92b, Str97, TK94, YHZ20]. profile [MM23]. program [Góm99, KLZV95, ZGLH24]. Programmable [LLQ17]. Programming [Dos03a, GC04, IW04, Zil01, dSCS04, AFG26, BBO21, DGL06, FM93, GWW15, GZ11, HBP13, JYLC21, KS18b, Kub23, LWM10, LDC10, Lin16, LCW21, LPXX19, MS26, MSP25, OL21, PW22, SPV20, SDW24, SW24b, TO21, WO00, WZ22b, YYL15, YZLP16, Yan17, Yan18, Yan22, ZM94, ZXL25]. Programs [Góm01, HL23b, JML25, Jos22, MZG26, WZZ15]. progress [Ano17, KM13]. Progressive [NL97]. Projected [BGR25, SL21a, AJMP11, ABM10, BHS14, BH17, Jón93, JRRS08, KJO23, LP08, LHL11, MRS06, YSXY19, ZCTD24]. projected-gradient [AJMP11]. Projection [BZ02, DHF21, EGSV04, AAM24, AAB13, AKB15, AK15, AAH18, AV19, ATT21, AT21, BAK25, BRZ18, BEHS20, BBB22, zDYG18, DEM93, DLYH17, GWBC20, GPGC98, GP99, GG22, HHST19, HT19, Jbi93, KR20, LP12, LP13, LCVL18, LC19, LWLW24, LLL26, LHZ20b, LRL22, LSY^{+}23, LJY24, LZT25, MJJ^{+}23, MT12, MRS07, MvS09, OL23, ODL21, PP24, PP16, PP17, PCDH20, RTCL21, SBA26, SLD20, SLT20, SXHZ20, THF21, TQY21, TVC20, TRSI23, TLD^{+}23, THT19, WZZ16, WXT22, WZC26, XCD23, XWL25, YJS25, Ye22, YJJ^{+}21, YYL26, YZLC24, ZFH23, ZZX^{+}23, ZJSS25]. projection-based [AK15, OL23, SBA26]. projection-type [RTCL21, SLD20, TLD^{+}23]. projection/Lagrange [SLT20]. Projections [BCN^{+}16, CE94, CCL16, HR05, HS16, KS23, NKM25, RZ23b, SEG14, Sei98, Szy06, Tru24, Zas22]. projective [HAN24]. projectivities [GM92a, MG91, MG94]. projector [SL16, SL18]. projector-based [SL18]. prolate [SHF15, Tia21]. proof [Eft15a, Eft15b, Fue07, GQ09, Not22]. proofs [Stu97, Szy06, YZH21]. propagation [AMM17, BCJ22, Elf25, LRC19, MR12, ZYJY22]. Proper [Oar94, LCW23]. properly [SL15a]. Properties [Bar13, BO03, BG03a, BSL18, BJ04, GL12, GC04, IMT02, Mas05, RB21, USM24, VL19, AHJ17, ÁFP07, BRZ18, Bea98, BKW25, BGVHN92a, CHYZ98, CMP21, Ceg24, DHM12, GTA19, Hoa25, HSL19, Jia20b, KM09, LWS18, Man21, Mat92, MAK20, NW17, Oar94, Sv95, SS14, YWX14, ZW12b]. property [BE20, Flo16, KM13, LWQR15, MP92, WS24b]. proportional [MMLM20]. Proportioning [Dos03a]. Proposal [DHL^{+}04]. protein [MB09]. Proving [Pop21]. Proximal [HSS04, IUM^{+}19, LX24, MH22, PKC18, SKK21, AAA^{+}18, AG23a, ATT22, CWHL20, Dey23, DEM94, God26, GZD24, HRAH22, JFSD25, KLW^{+}23, LB25, Ma20, MN22, MN23c, PW22, PLH20, PDS^{+}23, SBJC19, SW24b, TYSY20, UPY26, WL24, WZW^{+}24]. proximal-like [HRAH22]. Proximal-type [IUM^{+}19]. proximity [PH14]. PRP [AIA^{+}24, YYZ22]. PRP-BFGS-like [AIA^{+}24]. Prudnikov [GM22]. Prune [VR04]. Pseudo [PTSB01, TC05, CM24a, FHL21, FLY25, LDW18, LZlL20, LZ23a, LYL25, LQ20, LSY^{+}23, LJY24, LZT25, Mar96, Mor25, TZ24, TTLD20, TSI20, ZHT15, ZLLH22]. pseudo-boundaries [ZLLH22]. pseudo-function [Mar96]. pseudo-geodesic [PTSB01]. Pseudo-geodesics [PTSB01]. pseudo-monotone [LQ20, LSY^{+}23, LJY24, LZT25, TTLD20, TSI20]. pseudo-parabolic [LYL25]. Pseudo-Spectral [TC05, CM24a, LZlL20, LZ23a, TZ24, ZHT15]. pseudocontractions [Ceg24]. pseudocontractive [GMZ19, LCL21]. pseudomonotone [Anh19, ATT21, AB23, HMA16, JA22, JM18b, PT17, RTCL21, TQY21, TLD^{+}23, TLR^{+}25, TTA26, ZFH23, ZZX^{+}23]. pseudomontone [Hie18]. pseudoparabolic [LZ09]. Pseudospectra [LS03a]. Pseudospectral [BBD03, Ghe15, HLS10, Don13, JVH15, RS97, Sar06, Tia21]. pseudostress [Gha24, KPS22]. pseudostress-based [Gha24]. PSS [HWXC17, LZ18b, SCW17]. PT [AMM11]. PT-symmetric [AMM11]. publications [All08b]. Publisher [Ano00c]. Puiseux [WGZ18b]. pulse [Roh07, WXQ20]. purpose [FH05]. pursuit [HNY^{+}18, LLLD17]. put [CL10a]. Py [PGSU25]. pyramid [WQ10, dC22]. pyramid-typed [WQ10]. pyramids [PTW22]. Python [PGSU25].

QD [Van03]. QDR [APST21]. QLP [HC03]. QMGI [LSL25]. QN [GWL18, LL20a]. QN-matrices [GWL18]. QR [DG94, KSW09, ZLS25]. QSVD [ZLWZ21]. Quadratic [CCD10, Dos03a, FKMS01, Góm01, MS92, ZE12, AFG26, AABM08, ASS13, BHW23, BFKM20, BFK22, Bla15, Bog13, CWZ13, CJ12, CT21, DLR12, DGL06, GS94, Góm99, GS16b, HLC15, IJSS16, KKM20, Le 92, LWM10, LDC10, LRZ12, Liu21, LCW21, Mic23, OL21, OAR22, Pan20, Sat24a, Sat24b, UPS25, WZ11, Wan24, WCY25, Yan22, Zak17]. quadratic-cubic [Liu21]. Quadratically [Rec01, KSB08, SG10]. quadratization [LL20c, LL22c, Yue25]. Quadrature [BPR20, CR03a, CvPS15, GST02, GST03, KP03, SS01a, WL00, dABR01, AHR21, AAPR21, AR18, BD10, BEJR23, BCM07, BT23, BK18b, BHS17, BG13, BGVHN92d, BGVHN96a, BC09, BSL18, BD25, Cam95, CC12, Che99, CH11, CD07, CBGVN07, DPS18, DDRS23, FG07, FMD18, GM06, GSZ22, Gau10, Gau13b, Gau14, GHM23, GST21, GS95, Hag13, HHHN07, HMS11, HL15a, JH26, KXXW21, Kza97, LMV23, LMD25, Lau07, LM97, Leo07a, LHW13, Mas95, MM08b, MM09a, MV17, MT13, MT14, MSM25, Mon01, Not95, Not08, Not12, OdZdRV13, OPSM22, PP21, SKA23, SS01b, Spa24, SH26, TS18b, Van25, YWYN22, ZH23, ZYLN18, dC20, dC23, CIP10]. quadrature-based [MT13, MT14]. quadratures [Mil95, Mil17, MSS11, Pej14]. quadrilaterals [Kee94]. quadruple [KFK^{+}24]. Qualifying [BD04b]. qualitative [WS24b]. quality [AHS22b, AHS22a, Bra96, LAH22, NZ19]. Quantification [MHZ05, KBCG13]. Quantifying [ABV23]. Quantile [ZWZ25]. Quantile-based [ZWZ25]. Quantitative [BG03a, PK22b, PK22a]. quantum [FGL19, GJVW24, Kar07]. quartic [Kim21, MS96, QZG^{+}19]. Quasi [ABKD23, AL18, BMR19a, BKM03, BC17, Bre03a, Fra26, IMT02, LRW25, LI10, NSS03, PL04, ASS11, ASS13, ALY22, And19b, AG23b, BDIR18, BD20, BIM^{+}23, BM25, Cho16, CZLS18, DDS93, DBH21, Den14b, GZ18, GMZ19, GO20, HSSB13, JKM18, KG23, KBA23, KBA26, LEK21, LCL21, MWsC19, Maz11b, Men26, MK17, Ngu16, Orb15, PKC18, PXH25, Pan26, PSS25, RB21, RZ16, SKK21, SMM25, SCF23, SS23a, SSSS22, SWG20, SG23, SC18, UPY26, UL18, WRM17, WZZ15, YSLH19, YJS25, ZD15, ZZB20, BKL10]. Quasi- [LRW25, LI10, LCL21, MWsC19]. Quasi-analytical [BC17]. quasi-asymptotically [Den14b, GMZ19]. quasi-boundary [SCF23, YSLH19]. quasi-Bregman [Ngu16]. quasi-Chebyshev [WRM17]. quasi-compact [ZZB20]. quasi-definite [Orb15]. quasi-distance [UPY26]. Quasi-Eigenvalues [IMT02]. quasi-interpolant [ASS13, RB21]. Quasi-Interpolants [NSS03, ASS11]. quasi-interpolating [ALY22, DDS93]. Quasi-interpolation [AL18, BDIR18, GZ18, SWG20, SG23, UL18]. quasi-linear [MK17, PXH25, SMM25]. quasi-minimal [ZD15]. Quasi-Newton [BKM03, Bre03a, Fra26, PL04, And19b, AG23b, DBH21, GO20, JKM18, KBA23, KBA26, LEK21, Pan26, PSS25, SSSS22]. Quasi-Newton-based [ABKD23]. quasi-nonexpansive [PKC18, RZ16, SKK21]. quasi-orthogonal [BD20, KG23, SS23a]. quasi-periodic [CZLS18]. Quasi-Toeplitz [BMR19a, BIM^{+}23, BM25, Men26]. quasi-variational [SC18, WZZ15, YJS25]. quasiconformal [DM98a]. quasiconvex [UPY26]. quasideterminants [HLW26]. Quasilinear [Ben97, KK16, KKS22, Lie00, Pat25]. quasilinearization [ITA24, MS13]. Quasipower [Pas95]. quasiseparable [BAEF25, DP16, EGG08, Hua18]. Quasivector [Mar04c]. Quaternion [CZL26, ZJR25, BC05a, CKM^{+}25, CUY24a, CUY24b, DLL^{+}24, DH26, HLJ25, JNS19, JLZZ23, LLWC24, LSL25, LZX23, MZZ25, SLZ25, WZVJ22, WLW^{+}24, WLG26, WKY25, ZLWZ21, ZZ22a, ZLZ23, ZJJW24, ZWLZ24, ZLS25, ZZW26, ZMZ26]. quaternionic [GJVW24]. questions [Lin98]. queueing [BM97]. quicker [LM15, LSM16]. Quintic [FKMS01, XL14, KS12]. quotient [AC94a, AL09, OC24]. quotient-difference [AC94a].

Rabinowitz [Gau95]. Rachford [AG19, AG19, BCN^{+}16, BCS18]. Radau [CX20, Gau09a, Lau07, MVVA08, MT23, SH26]. Radial [CYIB12, Ave20, Bro05, CLMM05, DS09b, GZ20, GCJ26, HV15, LM00, LPP21, Low05, Nar05b, Pea13, SKTGR19, WZ15b, XLC93, YXS22, ZJ08, ZZL17]. radiation [NZF11, YHQ19]. radiative [GS25]. radius [Ali23, CB13, KE16, Mel24, Uhl09, ZCS14]. Radix [DR04]. Radon [NKM25]. Raising [BC05b, Wal06]. random [AK16, CCC25, DPP19, FS20, Fas25, Gor18, LWZ18, LL22b, LM17b, LGM25, Mil20, NT21, WCHK21, YYLX23, ZWZ25, ZL26a]. Randomized [ALL25, HKP25, JGMC25, LLWC24, SBN25, ZHSX23, BP22, CM25, DH26, EN11, FS23, HDL23, HWD25, HZ24, Jia20b, Kim26, KS23, LB25, MS24b, TL23, Wu22, ZWW25]. randomly [CM05]. randomness [DGM26]. Range [WH04, Gau12b, Gor18, GKQ24, HS96, LMV24, NRS12, SBW98, Ska13, Zah09, Gau17a, Gau19]. Rank [GSA03, OLB94, WSY04, APST21, AH11c, BBL23, BKS13, BHS14, BHW23, BEHS20, BH92b, BG13, CMM17, Dax17, DM22, DLC14, EGG08, EKPU23, ED22, Faz23, FH97, FHH99, FH05, HG93, HIS26, IDAV09, KKM20, LLZ18, LHZ^{+}21, LLWC24, LQ26, PK21, RS21, Sid94, Sid20b, TS92, TM10, VZ93, VVL25, WC23, WZVJ22, WKY25, ZJR25, ZL26a, Zhu15]. Rank- [OLB94, ED22, IDAV09]. rank-1 [VVL25]. rank-one [TS92]. rank-revealing [BH92b, DM22, FHH99, FH05, VZ93]. RankRev [LLZ18].

Raphson [Mor25, Par16, Ste20]. Rapidly [HIK17, AG00, GLRSG08, GLW13]. Rate [Dos03a, ART14, AK00, CCN25, GL19, LGL23, SLL22, ZWW21]. Rates [BCN^{+}16, CK24, AC19, Gor18, JLP20, Kno23, NT21, Pop18a, Pop19, SS98, SBJC19, WZ26, ZXL25]. ratio [ART14, MLM25, OR24, WCY25]. Rational [Ano92, BBD03, BM00, CGV92, Car92, CCV07, Col92, Dar99, GV99, Gre96, HGVPA92, KLT03, Mac96, MVVV24, Reb97, SS11b, Wen03, AH21, AAP24, AHP25, AC94b, Bal11, BC92, BBL22a, Ber00, BHJTM92, VVV22, BGVHN92a, BGVHN92b, BGVHN92c, BGVHN92d, BGVHN96b, BC09, CM92, CGW25, CH95, Cor91, CV92, Cuy00, CY10, EK94, FVV25, FWC16, FG07, Le 19, FJT94, GGV96, Gem97, Gen12, GH23a, GH25, Gla01, GCPG99, Gug96, Hof21, IT93, LZRJ92, LV15, LWK12, LKW17, LX23, MCMX20, NCC11, PGGC97, Poc14, Rob92, SST92, Sch17, She00, Sle25, TF00, VB92, Van07, VC92, Ver99, WQ10, XZL12, YD09, ZH25, dC16a, dC16b]. rationals [Dun94]. ratios [MS26, PYD23, YA24]. Raviart [LS20]. Rayleigh [ADP26, AL09, Ema96, IDS16, OC24, RR98]. RBF [CJK22, CL24a, FZ07, HSZ03, KJC18, KLL25, SW00, YYK24]. RBF-DQ [CJK22]. RBF-FD [YYK24]. RBFs [ZLZ22]. RBFs-based [ZLZ22]. Reaction [Sch02, BS17, BKF20, COSE22, CLT^{+}13, CSZ22, CLQ26, CYM22, CG19, CJ20, GM06, GK20, HG25, JKC26, Kaz24, Kno23, KMS05, LLY26, LZlL20, LCZ26, LS15b, Lin05, Lin09, LRZ12, LZQ26, MS24a, Moo07, Ngo23, NMM18, PM24, RF23, RR22, SKA23, Smi97, VT10, WZZQ24, WJC25, YJX15, ZYX19, dFO11]. reaction-diffusion [BKF20, CYM22, CJ20, JKC26, LCZ26, LS15b, Lin09, LRZ12, MS24a, Moo07, RF23, VT10, WJC25]. reaction-subdiffusion [CLT^{+}13, LZlL20, YJX15]. Reactions [KV04, BWB24, Hol98, SMK14]. reactive [VSA12]. Real [CD01, GKS04, Mai01, ZS03, ZMZ26, AHR21, Ali23, BDH^{+}13, Car95, CDGC25, CBGVPP09, CHS25, DY93, DFD23, EV22, HAN24, JZ16, KKK22, MZZ25, MLM19, NP18, ZA24, ZZ22a, ZLZ23, DCM^{+}13, DCMM13]. Real-Time [ZS03, JZ16]. realization [Lin98, PP21]. realizations [GMZ19]. Reciprocal [ABL… 12]. reciprocals [LMUZ19]. Reciprocity [SC03b, Gás99]. recoarsening [Hem96]. Reconstructing [Bea96]. Reconstruction [Flo03, HSSB13, LL20b, AB23, Bou24, CLMM05, CsL24, DHJJ10, DAN26, DPR23, EHN17a, HJ18b, HOW95, HV22, MP25, MFBB23, MN23b, Wan15a, XYZ14, XTH07, XL14, Zak17, ZJ08, EHN17b]. recovered [ZHZ25]. Recovering [LC25, ST02, WDY04, Gor18, KK22a, LJX25]. Recovery [ZLH21, BK18a, AFG26, EKPU23, GLW13, Han93, Kow00, RS21, SBA26, Sch26, TY96, ZY21b]. Rectangle [BK04]. Rectangular [CP01a, DM97, AAD14, BF17, KRS19, WZZ26]. Recurrence [Leo03, LGA^{+}00, LW04, ZG12a, Aih17, BPR20, Da 92, DS23, FHV15, Gau09d, HHLM23, Leo07b, Leo08, Maz09b, Müh99, Sch17, VC10]. recurrences [ZD15]. recursion [SSP15]. Recursive [CK20, MS17, TA13, BM12, CP95a, FGP91, IR13, Mar92, MVV05b, OLB94, SS23a, SS23b, TZH24]. recursively [MT06]. red [FS21, Cia94]. Red-Black [Cia94]. Redivo [CHHL18]. reduce [Ant18]. Reduced [GF02, HOW95, SR24b, AK26b, BAS25, El 18, EE18, HS26, KLZV95, LL07, LCW20, MG22, Sid94, Sid20b, SH17, WZZQ24]. Reduced-order [SR24b, KLZV95]. Reducing [AMM17, Bün24, VVL25, Met19, MRS10]. Reduction [Che01, CMRS01, DR01, AHJ17, ACSD16, BW93, BHS14, BH17, Bia94, BM97, BM09, Bos21, CDW95, CP93, CP95b, CJ17, DLC14, DR12, FLT09, GGV96, GLW16, GSV96, Kao20, KV07, MS26, ME92, RSCH^{+}19, Sad99, Sau07, Str09]. redundant [YP23]. Reeves [BKG15]. reference [BF93, PV22a, WLL12]. referential [PV22a]. referred [TN10]. Refinable [GPS01, Dah93]. Refined [Che04, HJ25, ARS26, BPP23, BE17, FS21, LCVL18, MT12]. refinement [AS14, BKS23, Hem96, Mic91, MRU91, Str05, UTO24, YF22]. reflected [YAY26, ZCTD24]. reflection [BBCS21]. reflections [AC19]. reflective [FFSC25]. reflexive [HCD26, Pen13, XCD23]. regime [AKKT16, CZ23, CDLW21, Che22, Che24, CGSS24, MY22, RS06, SPB26, YYK24]. regime-switching [CDLW21, Che22, Che24, CGSS24, SPB26, YYK24]. Region [CR03a, Car01, AA12a, ARY17, BKR18, BE17, CHY19, EG94, HZ93, HZ95, KE16, LZ18a, LM13, lLXhL22, MLM19, NAR05a, SS06, ZW14, lZqL25]. regionally [LQ26]. Regions [SSH20, ASZ23, BBBC20, BF17, Lóc20, SH12, SSH19b, Tru24]. registration [BE17, CGYZ19]. regression [For93, HZ93, KN18, LDN16, Len93, MMV19]. regula [CL06]. Regular [SX96, Bar91, CT21, HMdÁES08, HJ25, Ish26, KP96a, Kar09, KJC18, KR07, Mar93, Mar96, SYZ22]. regularisation [MRS93]. Regularity [CRZ26, CCS05, ARTY20, BM22, LZCW24, PSS25]. Regularization [BG11, CRS04, CDD21, Han02, Jia20b, Alt21, AK26c, BMR21, BEL23, BRS08, BRS09, BOR23, BGR25, CMD19, CK05, CPZ14, DC17, Don12, DN24b, DNH^{+}25, DR12, FR12, GHN19, GGS22, HR14, HYT26, HLC15, HRY16, HRY19, HS12, KZS21, LP12, LQ26, LYL25, LGM25, LKK21, MRV23, MRS07, MG18, MH22, NR12, NR14, PGSU25, PSZ23, RP26, SR25, SCF23, SGJ15, SLL22, VMMD21, WC23, Wan24, WZ22b, XH25, YXS22, ZBDK23, ZZ19, ZL22a, ZLL21b, ZFH26, dFG93, Han94, Han99, Han07]. Regularized [AX26, GWL20, PPR15, dAFPR23, AC11, AT19, BG11, BGS24a, LZ22b, Liu14, Sch25, SLL23, ZH22, ZCC26, ZPX21]. Reich [PK25b, PK25a, PGC24]. Reid [LL05]. Related [BBR03, Boy05, Góm01, APPR14, Ari98, BM25, BRZ19, BRZS23, CL13b, DJM^{+}18, KR11, LM08, LM15, LSM16, MSM25, SH23, WZ15b, ZA20]. relation [BPR20, Müh99, Sch17, VC10]. Relations [Cha14, Leo03, LW04, dABR01, Che94, Da 92, HHLM23, Leo07b, Leo08, LGA^{+}00, Maz09b, ZG12a]. Relationship [YZH21]. relative [Bör26, GM21, Gau22, GM22, GLM15, LMV23, MPR24, DCMM13]. ReLaTIve. [DCM^{+}13]. Relativistic [MN01]. Relaxation [MS02, ZZ10, vLV02, BX17, CRZ26, CC06, CWHL20, CGM25, CDP16, CG13, CK06, DEC24, DLL^{+}24, Fan19, FS23, GZL^{+}25, JM18a, gK25, LHM20, LJWW21, LCW21, NRV23, OM18, PT19, SJ14, SJW21, TPLB22, WL25, ZJ14, ZLV17]. relaxation-type [LHM20]. Relaxed [BCN^{+}16, zDYG18, GM23, OMU^{+}24, BBL23, CRZ26, DCG25, DH26, tFZyZ16, HC24, KM17, LSY^{+}23, PSS25, SCD^{+}21, SM17, SW24b, WK15, WK17]. relaxed-dual [DCG25]. relaxed-inertial [LSY^{+}23]. release [HLS10]. reliability [EY10]. reliability/cost [EY10]. Reliable [Cox93, CCK04, Mar04b, MP02, dADdRC04, BC92, EM07, GST21, HM22a, HM22b, HA16, JW25, ZSGJ25]. remainder [Wil12, MG94]. Remark [Gau11b]. Remarkable [All18]. Remarks [ACF99, Dum03, Osw01, CC07, Le 98, Maz09b, ZW15]. Reminiscences [Ano00d, Bre19]. remote [WZW^{+}24]. removal [AL18, CHH93, JHLL15, WHZ^{+}18, WZW^{+}24, XLL26]. reorthogonalization [JL21]. Repeated [Cia94, Gau13c]. representation [BCJ97, CKM^{+}25, HS26, IN21, MP07, Med10, OB16, TO21, TD09, YD09]. Representations [APST21, Mat04, AC94b, BAEF25, Ber00, IN95, MM12, SM25a]. Representing [Sza03, CDF99]. reproducing [AR13, BAV18, LZ09, LRM16, Tan20, TA96, WCD21]. repulsive [OKP21]. Requirements [Sha02]. Rescale [MH08]. Rescaling [DNR15, CW25]. research [Ben99a, Lev05, PH20]. Residual [Bac25, GR01, Rob02, AM01, DBGB11, GS16b, lLXhL22, MT15, RALJ26, Sad05, TM20, ZD15, Zha20, ZYW21]. Residual-based [Bac25]. residuals [CKM19]. resistance [ZZZ20]. resolution [BC05a, FZ26, HLN25, Tuo98]. resolved [MvS09]. resolvents [Tak17]. resolving [SKM25]. resonance [CZX26]. Respect [LW04, Gau15, Gau17b, LGA^{+}00, Mil17, Müh99]. response [Cou15a, Cou15b]. Restart [DLL04, HLJ25, LLL26, LSY^{+}23]. Restarted [BR06, BLW09, Che04, Bag00, BPP23, CR12, EHTSM21, FEL15, GSV96, LM99, Lun23, TM20]. restarting [DB98, JGMC25]. Restoration [SST08, AIA^{+}24, AR26, FGBP21, HS26, JLJ22, KBA26, LP13, LZ22b, NR14, SKK25, SD25, SD26, WYP23, ZL24, ZL26b, ZH26]. restrained [QLZX11]. restricted [AG17, LY17, NRS12, SBW98, CGV22]. restriction [ADL05]. result [AG00, Che16a, SI17, YL19]. resulting [FJT94]. Results [Vig04, AUA22, BBC21, BRZ20, CSI18, DD20, DEM94, LRT19, Lin16, LA22b, Lu15, Mat91, Mat96, MSP25, NNCN23, RR08, SI13, Smo99, VC92]. resurgent [FT05a]. Retarded [AR99, MD21b]. retinex [CZM21]. Retraction [JM18b]. Reusing [GI15]. revealing [BH92b, DM22, FH97, FHH99, FH05, VZ93]. reversible [BL98]. Review [De 02, Kuh13, MP07]. Reviews [Ano01f, Ano01g, Ano02c, Ano03a, Ano04b, Ano06, Bre96, Bre03b, Ano95a, Ano97, Ano98a, Ano98b, Ano98c, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano01c, Ano01d, Ano01e, Ano05a, Bre97b, Bre97a, Bre97c, Bre97d, Bre00b, Bre06b]. revised [BKDA25]. Revisit [Sol11]. Revisited [vGK04, BDS00, Bün18, Cho17, Dra96, TMS25]. Revisiting [Gut15, Saa23, TQC22]. Reynolds [BGL07, LBF24, MDR23]. RHBPIA [REE26]. rho [CHH^{+}20]. ribbon [CMRS00b, HSSB13]. Ribiére [YZBJ21, BKG15, MA22]. Riccati [Arn97, BBZ95, BEQOR14, BHW23, CXL16, DI11, EH24b, FWC16, GW24, Guo13, Guo16, GL23, Jbi03, LL26b, Sti18]. Richard [Ano00d, BC22]. Richardson [ABL… 12, CR96a, Chk20, CFS21, KR20, Sid17, XWY19, dC20]. ridge [XLC93]. Riemann [BX19, CA22, CRV91, LK20, lLHNS23, NPR08, QW08, VLCL16]. Riemannian [EKPU23, IS22, KC24, KLW^{+}23, NH24, NN26, Ovi22, TTXZ23, UPY26, YZBJ21, ZYBJ23]. Riesz [BM23, Caç10, CY19, LLAL21, LLM25, LCZ26, LWJ21, SKD26, SLA11, WSLL24, ZYW23, ZL25, ZLZ22]. right [BEHS20, BF14, BDR25, DMD16, Fue07, Hey01, HE05, KBCG13, TT21]. right-hand [BEHS20, BDR25, DMD16, Hey01, HE05, KBCG13, TT21]. rigid [PPR15, TCOA19]. Rigorous [Dze15, Joh15, Not22]. rise [CGM93]. rising [TA24]. Ritz [Ema96, IDS16, MT19, RR98, TM20]. river [BM24a]. RK [BH05, IKMS24, SAHM25]. RKN [LDW18, LWS18, WW19, YZY^{+}14]. RKN-type [LWS18]. RLW [LKQ23]. RMPIA [MS23b]. RMVPIA [EEM20]. RNN [SDL^{+}23]. Robin [BDF^{+}25, Rab23, ZBX21]. robotic [SKK25]. Robust [BE17, KN18, LQ26, LZ19a, SKJ^{+}18, YYD14, BPR22, BDN17, CK24, CL10a, CLQ26, Cou15a, Cou15b, DFR26, HS21, JFSD25, KP16, KK12, LAN18, LP20, LLC20, NH24, ST26, SPS26, YM24, YHL25, YQM16, ZZY24, dFO11]. robustness [KC23]. Rogers [HLM04]. Role [Fas25, MT15]. Romanovski [ZAGD22]. Romanovski-Bessel [ZAGD22]. Romberg [JC04, dC20, dC23]. Root [MS01b, BCMT18, BHM05, BC17, CCV23, GL20, HKP25, Hig97, KGG25, Kol06, LK20, MCC26, PR14, PH20, QAS^{+}24, SS24b, WK17, ZXRL11]. root-finding [BC17, MCC26, PR14, PH20, QAS^{+}24, WK17]. rooted [ZYW17]. rooted-tree [ZYW17]. rootfinder [BF00]. rootfinding [LC14]. Roots [Rec01, BCM16, BDS00, Cam19, CL11, CN17, DY93, FS20, GH10, Gar19, GI10, GHPMGRR14, LMMH11, MLM19, RG10, SS12a, SSH20, TTV21, ZCT19, ZCS14, dS00, dDS00]. Rosenau [LKQ23]. Rosenbrock [DBTG24]. Rosenbrock-exponential [DBTG24]. ROSEXP [DBTG24]. rotated [BL23, LM19, NMM25, SST08]. rotation [PPV09, Saa23]. Rothe [BCJ24, MZ19]. rough [Bro05, HLS10]. round [AMM17, KL17]. round-off [AMM17, KL17]. Rounding [GL04, dC20, ZWLZ24, ZZW26]. Roundoff [Ihs07]. routines [Gau10, OOOR12]. row [AH14, EKPU23, FAM25a, FAM25b, ZWW25]. row-action [AH14, FAM25a, FAM25b, ZWW25]. row-sparse [EKPU23]. rows [ST18, Sid94]. RPIA [MS17]. RSVD [AR26]. Rubenchik [LZ24]. Rule [BD04b, Boy05, BD10, BHNS16, BC92, Bou17, CC12, For21, MZG26, NAE22, SH26, TB19, YWYN22]. Rule-of-Thumb [Boy05]. ruled [WZZ07a, WZZ07b]. Rules [GST03, INR01, dABR01, AHR21, AAPR21, AC94a, AR18, BDL^{+}12, BLS92, BPR20, CSI18, DDRS23, DK00, FMD18, GSZ22, Gau14, GPGVS92, GS95, HKKN12, KXXW21, LMV23, MFPG07, MSM25, Mon01, RR13, SR25, SH22, SH23, TS18b, VPL97, WL00, Zag92]. runaway [SSH^{+}19a]. Runge [KM19, AMCM06, AH15, AMM17, AMM18, BJ98, BS14, BB14b, BD98, BC99, BJ04, CGPM00, CMR03, CT10, Che14, CQLY15, Che16b, Con93, CDI14, DJ10, EH97a, FYM14, FYYW19, FHL21, FLY25, GCJ26, JT96, Jat15, Kha14, KFK^{+}24, Kri26, Li17, LG19, LT20, ME92, MV02, MC08, McL25, MKS18, MHA16, PFT98, SQG13, ST21, SZ24, TX19, Tir02, TCW14, Tsi07, TÖ17, VH12, VV11, Ver10, Ver14, Ver24, WG13, WC25, Wri02, XT16, ZYQ^{+}21, Zha24, dS24, vSv94].

S [Ano00d, GLRSG08, RS20, LL20a]. S&P [AWL^{+}24]. S-QN [LL20a]. S.S.O.R. [MS02, SMB02]. Sabin [Rip93, WD95, WD96]. Saddle [LS03b, WSY04, AK26a, BBS20, CXC25, CZ14, CC15, CM16, DYW16, EMS25, tFZyZ16, HT21, HWXC17, HM18c, HWXC18, HWXC19, HM19b, Hua20, JWCZ21, JCH23, KM17, KKO17, LM17a, LZ15, LZ18b, LXZZ21, LZ22d, RCW22, SM17, ST22, SCW17, WZ13a, WDL16, XM16, YDWL15, ZRZ11, Zha26, ZWY22]. saddle-point [DYW16, EMS25, HT21, LZ15, LXZZ21, YDWL15, ZRZ11]. Safe [PHI98]. Safouhi [SS14]. Sahalia [JLWZ25, LGL23, LCW25, ZWW21]. Sahalia-type [LGL23, ZWW21]. sample [DWX17, JB22, JR20, JSZ22, KKV22, KJ15]. sampled [BW93]. sampled-data [BW93]. samples [AAAGAD23, BT23]. sampling [Ash16, AAA17, Ash19, AP21, AAH24, AK26c, Gor18, Kow00, LX17, Man07, MW24, MW26, MN23c, NYL26, Tam10, UTO07]. Sandro [BT14]. sandwiched [DMYT23]. Sard [Ehr97, HMS11]. Sard-optimal [Ehr97]. satisfactory [DGST15]. satisfying [Bag16, BCCP25, CG09]. SAV [HZX21, JY23, LZZ23]. SB [DLL12b, DLL12a]. SB-matrices [DLL12b, DLL12a]. scalable [DHV22]. scalar [CZ20, DHV22, Kac18, LMUZ19, LL22c, PGGCGF11, Sat24a, Sat24b, XT16]. Scale [Boy05, GSA03, Hea10, AAB13, AN17, Ali23, AKB15, And22, AG00, AG23b, BLW09, BAK25, BPS23, BEH24, CH22, CC13, DS20, FZ26, FGS24, GHN19, Guo16, HJ18a, HRY19, JNS19, JML25, KW00, Kim26, La 17, LWwCL13, LZ21, fLxX12, LL18, Md12, RRS09, RR98, RR00, Rog95, SBN25, ST22, SKM25, Sti18, WCLW16, WYP23, XLL26, YCL17, Zha20, ZW20]. Scaled [DMA19, KZ03, LZ21, And18a, BKA19, BE20, CL21, CS24, ODL21, Zag24a, Zag26, ZW22, ZLWQ09]. scales [Zaf22]. Scaling [GPP01b, LG04, AG25, Ari98, BK16a, Fly22, LZ18a, LZ22c, VGV06, ZWL26, lZqL25, lZfLqL26]. scaling-and-squaring [AG25]. scan [BKPS93, FZ26]. SCAN-MUSIC [FZ26]. scatter [HHLS21a, HHLS21b]. Scattered [Rip93, AF94, Bro05, CF96, CK05, CM24b, CDD13, CM05, GI97, Hof05, IL05, LW20, Nar05b, Sch08, WD96]. scattered-data [Nar05b]. Scattering [ZKD02, ZKD04, BE98, GH06, GH09a, HLS10]. Scheme [Dum03, EGSV04, SS03, ASGGRG23, AUA22, AE09, BCL00, BS17, BF20, BW15, BZS22, Bra06, Bra07, CL10a, CEX14, CXL16, CSZ22, Che22, Che24, CV15, CJKL23, CJ20, CMP22, DL21, DMYT23, DZ19, DZ22, ETY98, GLLJ12, GD15b, HHST19, HKE97, HP18b, Hei07, HMS96, HL20, HSY23, JJ24, JZ25, JLFL19, KN23, KLF17, KS26, KLB10, KK17, Laz99, LWD23, LZlL20, LlLZ21, LL22b, LKQ23, LKBF17, LW17, LS20, LWZ23, LCZZ23, lLXhL22, LV18, MMW20, MWZL23, MClXzJe16, MA13, MN11, Ngo23, ON26, PKR20, Pen98, RF23, RP26, RR22, Rou25, SSR25, San25, Son93, SSA25, SPS26, Str05, SZ20, SzS21, SYZ22, Ton25, VA20, WLL12, WG13, WH15, WYZ22, WQZH24, WC24, WPY25, WZ26, XCY21, XF25, YZZL17, YH24, YCZ26, YYLX23, YY24, ZP17, ZSF18, ZQzS22, ZYW23, ZY23, ZSLZ24, ZLS25, ZZ10, dFO11]. Schemes [AKW02, ABQ04, CC03, MM04, ADG10, AABM08, ABT07, ALRT16, ARTY20, AHS22b, AMM17, AMM18, AK16, AH18, BCST14, BF18, BP22, BC26, BC01a, CF05, CCS05, CLTA10, CS22, CLPY23, CG19, CRHTTN24, Cui13, DL08, DAN26, DCW23, DFD23, DK00, DXY18, FS23, GB21, hGzS17, Haj16b, HHHN07, HZX21, HKCW24, HTVY13, IKMS24, JLWZ25, KP96b, Kie23, KMV17, KM19, Kun05, LR18, LZM23, lLLVZ17, LZW20, LWJ21, LDL^{+}19, Liu21, LL22c, LCW25, MPS20, Met19, MP22, NT21, OCWV24a, OCWV24b, PXH25, PS17, QQX23, SS26, SR25, ST17, SSL24, SM25b, Sti18, TZ24, TS15, WWBM21, WSLL24, YZLZ22, ZJWF18, ZYQ^{+}21, ZQS24, ZLM25, ZM26, ZAC26, dC20, iV12, AHS22a]. Schmidt [CFM15, Sal05]. Schmidt-like [Sal05]. Scholes [AAN25, AWL^{+}24, KN23, ON26, Rou25, Val14, Val15, ZZ22b]. Schröder [GL20, SSH19b, SSH20]. Schröder-like [SSH20]. Schröder’s [PP18]. Schrödinger [AM21, AT19, AKKT16, BX17, BH92a, CZ23, CG20, GJVW24, KLF17, KPFF24, Lam09, LHW17, LHZ20a, LHM20, LKBF17, LZW20, LWZ23, Lor19, SZ23, SKTGR19, WH15, WZ19, WWBM21, WLZ22, YJ21, YSLL23, YY24, YL22, ZHT15, ZJWF18, ZTZZ19]. Schrödinger-=Helmholtz [SZ23]. Schur [BBL23, BKS23, Cao12, ED22, GI10, ST18, Thr92, TA13, Zaf22, ZS08]. Schur-like [ED22]. Schwarz [CZ95, BX17, Cal20, CZ94, CZ96, CGV22, CTS09, CK22, CGM25, CG13, DL09, GD15a, GL15, Hei06, LS14, Lui02, NG23, Rah11a, ST99, SME03, TPLB22, YHR26, ZJ14]. science [Ste20]. scientific [CR00, Che94, YZ17, ZL24, ZL26b]. SCLCPs [SD20]. Scott [LWD23]. SDD [ZWL26]. SDE [AH15]. SDEs [HH11, Hag25, HY21, HSY23, Ste24, SYZ22, VH12]. SDFEM [LZ18c]. SDIRKN [FGR01]. SDO [Khe12b, WZ11]. Search [GPP01a, LPV03, AAN14, AK15, BKR18, BBBC23, DW12, Deh20, GZ11, HWC15, KOK21, KJ15, LZ18a, LLL18, Liu11, LM13, MS11, NKA24, Ovi22, SD20, SW11, SM10, TH19a, WZ16, WCD21, YIY22, Yan22, YLL22, Zha24]. searches [GS21]. seawater [SMNZ20, SFZ22]. Secant [AH12, KGMH21, SXHZ20, And10, Arg10, Ari98, VPA25, Zha20, AR09]. secant-like [Arg10]. Secant-type [AH12, KGMH21]. Second [AH21, BH05, Che22, CS22, DBTG24, Fun01, KKD25, MHA16, The97, OI14, PW04, Si12, SFT03, WQ23, Xu19, AR24, AHJ22, AK12, AM16, Arn97, AAD14, Bac25, BK13, Bra07, BV09, Cai22, CEX14, CS25, CW14, CQLY15, CC18, CLPY23, CKKT16, VV07a, VV07b, DD21, DK25, DS26, DB06, DK00, lDzS21, EGSHVN15, FLY25, HH11, HP18a, Hea10, HSTW14, HWZ24, Hun95, IR13, JL12, KMZ18, KMH25, KS20, KK17, LW13, LW16, LR18, LHR20, LHZ^{+}21, LlLZ21, LZOY22, LZM23, jLyLqW17, LWS18, LDL^{+}19, LW22, LWZ23, LXP20, LV18, ML20, MAH22, MWDD24, ME95, MWZL23, Mic93, MBM25, MWWY13, Pan18, PP18, RKMS16, RF23, REM21, RZZ25, SW19, Sho18, SYLT14, SZ20, SzS21, TX19, TTP26, VTV22, WK16, WHZ^{+}18, WWBM21, Wan22, WQZH24, WLJ24, WLY^{+}21, WWM21, XYZ14, XZZ22, XH21, YZLZ22, YYL15, YZY^{+}14, ZP17]. second [ZYW23, ZXLF15, ZPEL25, ZLS24, vdHM98]. second- [DD21]. second-derivative [CKKT16]. Second-Order [Fun01, Che22, CS22, DBTG24, KKD25, WQ23, Xu19, Bac25, CQLY15, CLPY23, VV07a, VV07b, DK25, DB06, lDzS21, FLY25, HP18a, HWZ24, KMZ18, KS20, LW13, LW16, LR18, LHR20, LHZ^{+}21, LlLZ21, LZOY22, LZM23, LWS18, LDL^{+}19, LW22, LWZ23, LXP20, LV18, MWDD24, MWZL23, MWWY13, Pan18, RF23, RZZ25, SW19, Sho18, SYLT14, SZ20, SzS21, VTV22, WHZ^{+}18, WWBM21, Wan22, WQZH24, WLY^{+}21, WWM21, XYZ14, XZZ22, XH21, YZLZ22, YYL15, YZY^{+}14, ZP17, ZYW23, ZPEL25, ZLS24, vdHM98]. seepage [CW17]. Segel [KKA17]. Segmentation [GLV05, dADdRC04, FLG08, GLRSG08, LAG05, RC14, SMA99, XW17]. segments [CCV07]. Seidel [DLL^{+}24, JFSD25, MS24b, RWTM21]. SEIR [ALB^{+}18]. Selected [Mil13]. selecting [OOO11]. Selection [CVA01, Cse04, MT04, ABT07, BG11, CPZ14, Kac18, Len93, LZ25, LXP20, Moo07, SGS25, SGS26, SS21, WCM94, YDW26]. selections [HWCR19]. Selective [HT19, RC14]. Self [AV19, NEY26, TKSG23, Bno21, CHY19, CD25b, FSY23, GNS22, GHM24, HK14, JA22, LZ22c, LG95, PV22a, RFS23, RTTH22, RT24, TG20, TA13, TN10, ZH19, ZW15]. Self-adaptive [AV19, NEY26, TKSG23, Bno21, CHY19, CD25b, FSY23, JA22, RFS23, RTTH22, RT24, TG20, ZH19, ZW15]. self-adjoint [GNS22, LG95]. self-consistent [HK14]. self-referential [PV22a]. self-referred [TN10]. self-scaling [LZ22c]. self-similar [GHM24]. Semi [LW04, LZ15, LZ18b, LDC10, SS03, BIM^{+}23, CC15, CL13a, CLWH20, Cro92, DO25, tFZyZ16, FSY23, FM93, Van12, HDP18, Ish26, KH11, KR07, LPXX19, Maj13, MM08a, MFK^{+}15, MSMS12, PW22, Pei24, SS12b, SH21b, SH22, SH23, SS24d, TO21, TB19, YWX14, YBK^{+}21, Zah09, ZZ19]. semi-analytical [MFK^{+}15]. Semi-Classical [LW04]. Semi-convergence [LZ15, LZ18b, CC15, CL13a, YWX14]. Semi-definite [LDC10, tFZyZ16]. semi-discrete [DO25]. semi-discretizations [FSY23]. semi-finite [Maj13]. semi-implicit [Pei24, TB19]. semi-infinite [BIM^{+}23, Cro92, FM93, KH11, LPXX19, PW22, SS12b, SH21b, SH22, SH23, TO21, Zah09]. Semi-Lagrangian [SS03]. semi-linear [CLWH20, Van12, HDP18, SS24d, YBK^{+}21, ZZ19]. semi-porous [MSMS12]. semi-regular [Ish26, KR07]. semiaxis [DFD23]. semicircle [MSM25]. semiclassical [AKKT16]. semicoarsening [PV00]. semiconductor [NAR05a]. semidefinite [HBP13, Ke21, Khe12a, KOK21, Lin16, LCW21, MR09]. semilinear [CJ20, HWC25, HZX20, KLL10, KP25, LSW16, Vul97, VT10, WW19, XH20, XWX24, YM24, YZ11]. Semilocal [CWL16, HVMT17, HM14, Jai16, WGK11, WKG11, WK12, WK13, WK16, WK17, ZG12b, HVM15, Hua94, Iva17, LLL22, QAS^{+}24, ZG12a, Jai17]. seminorm [Zha95]. seminorms [HS12]. Semiseparable [FG03a, MVV05a, VGV06]. semismooth [SR06, SZX11]. sense [CSI17, Dax09, HMS11, KC23, Lei15, LL22a, TS18b]. sensing [CZX26, HLL22, HIS26, LWLW24, LLL26, MJJ^{+}23, SIO20, WZW^{+}24]. sensitivities [CE17]. Sensitivity [Kow00, MW98, EHNR23, Han22, Hua18]. Separable [FLMR00, AG23a, CsL24, MM12, SY20, SXHZ20, SW24b]. separated [BDN17]. separating [AK15, CL10a]. separation [WLJ24, YC22, ZS19, dC23]. September [BV96]. septic [KK22b]. sequence [BDJ11, Ben97, BS92, CHHL18, DM92, GM96, Hom94, Hom98a, MdR08, Mia19, SS12b, Wen92]. Sequences [CP01a, AH13, AG13, BD20, BHJTM92, BCJ24, Chk20, CL13b, Don12, EHTSM21, FFSC25, KJG23, LM08, LM12, Ngu16, Osa92a, Osa92b, Pep23, PP06, Sab91, Sab92a, Sid20a, Sza92]. sequential [AFG26, CM16, PSZ23, PN21, XP23]. serendipity [AW23]. seriation [CFR19, CFRV23]. Series [Boy05, Bre04, FHAL15, LW04, Mas05, Maz02, Mül00, NKS04, SVZ01, AHM21, BBM08, Bou06, But19, CGV92, Cha14, CW08, Esp05, GL20, HGVPA92, Hom92, Hom98a, LGA^{+}00, Mil95, Mor11, NW17, Now19, Pas95, Pas08, PM25, Pry98, Sab92b, SL21a, Sid07, TBA94, VGM96, Vep08, VSA12, WWL24, dDL92, dS24]. Set [BL04, Góm01, AHL20, ABM10, AH08, Góm99, GLV05, JB22, KD18, KA26, LAG05, LL22a, Nor00, PCDH20, Pop15, Pop18b, RT22, San19, SSH19b, TS18b]. set-valued [AH08]. Sets [FH04, KLT03, MKO04, ASW06, Buo17, BF93, CKO24, Caç10, CG09, CDD13, zDYG18, DMD16, FHH05, GGN14, Gar24, GJV17, GHM24, KKV22, Kol06, KGN^{+}24, LNS23, Mar92, Ple12, PN21, RTTH22, TKSG23, Tan20, Wan18, Zas22, dB07a]. setting [ARS26, HL06]. seven [Kim21]. seven-direction [Kim21]. Seventh [WZQT15, NBK17, SA14, WZ13b]. Seventh-order [WZQT15, NBK17, WZ13b]. Several [SLL22, APPR14, AT17, AAH24, BBC25, Che13, DFP^{+}10, DGP15, Gal18, HJ25, Iva24, ZWFY19]. SGEM [LT24]. shadowing [CKP97]. Shallow [vdHS02, PS20]. Shanks [BRZ19, LCHH21]. Shannon [Tam10]. Shape [BG91a, CP01b, LM01, CKS24, Ell93, FZ07, KK00, LMM97, Tur26, ZS19]. Shape-Preserving [CP01b, LM01, KK00]. shapedeness [DO25]. shapes [SST08]. Sharp [CSZ22, JLMP16, LZlL20]. Shaw [CLPY23, CJKL23]. SHBVMs [ABI20]. shear [WWD^{+}12]. shell [PPS25]. shells [PS20]. Shepard [LRY26]. Shift [ED13, LLD23, TK94, BS17, BKS13, Fue07, HWXC18, LM19, LLM25, NMM25, Plo94, SBR26]. Shift-invariant [ED13, SBR26]. Shift-splitting [LLD23, HWXC18, LM19, NMM25]. shift-variant [ED13]. Shifted [Don10, Tan17, EHTSM21, KM24b, LWJ21, MAK20, Plo93]. Shifts [Che04, BF14, Bro05, CDT10, DG94, MVVV24]. Shishkin [CGL99, FLR01, LZ18c, ST98]. Shishkin-type [FLR01]. shock [AC17, VN18]. shocks [KMV17, SMR26]. Shohat [dR99]. Shooting [AMM11, CW17]. short [Aih17, ACF99, BD17b, CG13, DPR23, GKQ24, LS11, Sab92b, Sch17, WXQ20]. short-memory [BD17b]. short-pulse [WXQ20]. short-range [GKQ24]. short-recurrence [Aih17, Sch17]. should [JB22]. Shrink [Bün18]. shrinking [THT19]. shrinks [SSH19b]. Side [BKPS93, BEHS20]. Side-scan [BKPS93]. sided [BZV16, HZPW23, HLTA16, RSCH^{+}19, Sch17, VZ93]. sides [BF14, BDR25, DMD16, Hey01, HE05, KBCG13, TT21]. sideways [ZL22a]. Sidi [VPL97]. sign [BBQO07, BQO99, BEQOR14, KL94, Mit11, SWS22]. sign-based [SWS22]. signal [AFG26, ABK22, ABKD23, AB23, Bea96, OB16, SBA26, ZY21b]. signals [Cic20, CHH93, Gor18, GLW13, LGP11]. signatures [Thi93]. significant [ZA24]. signum [YP09]. silent [CH26, Meu23]. Similar [Mel10, GHM24]. similarity [FHAL15]. Simple [BD04b, DD21, FGL19, KSB08, NS01, RSZ20, dABR01, AN17, BESC22, CD15, CN17, DN24a, GQ09, LZ22a, NP22, SSK23]. simple-loop [BESC22]. simpler [AEH20, JR10]. Simplex [CD01, CD07, JKK^{+}08]. simplices [CvPS15]. simplicial [AMM16, FS20]. Simplified [AR18, Khe12b, AMM18, BRZ17]. simulate [AD22]. Simulating [AB98, BBL22b, LRC19]. Simulation [AA03, TE03, AF13, BGL07, GS19c, Gro93, KS18a, KR23, LNS23, MBR21, NPS09, Pen98, SLLA15, ZLT^{+}17, ZZ22b, ZTZZ19]. Simulations [MHZ05, Sha02, GKQ24]. simulators [SR22]. Simultaneous [All03, BG03a, KLSS14, LXX23, BL92, Bel94, BM06, CCL18, Iva24, LMD25, PHI98, PM05, PMM11, Pié99, TPY14, Van25, dDL92]. Simultaneously [Rec01, Iva17, LYL25]. Sinc [AT19, SGK^{+}99, Ash16, Ash19, Ber07, Ber11, EG18, HA16, KPT23, TBY13, RMT13, RBN14, ZS13]. sinc-Galerkin [EG18]. sinc-Gaussian [Ash16, Ash19]. sinc-interpolation [Ber07]. Sinc-regularized [AT19]. sine [AA15, Bra06, CM24a, DCW23, FSY23, HH05, Hem94, HFDSC24, HKCW24, JJC24, Mac96, RM11, Saf10, ZLM25, ZM26]. sine-Gordon [Bra06, DCW23, FSY23, HKCW24, JJC24, RM11, ZLM25, ZM26]. Single [PP16, Car92, FZ26, KR11, TRCL24, XW18, ZFC18, PP17]. single-step [XW18]. single-valued [TRCL24, ZFC18]. singly [BD98, BC99, Che14, PFT98]. singly-implicit [BC99, Che14]. Singular [AZMJ04, AC94a, AKW02, AKKW03, Bel03, Boy05, DM98b, MO04, RR23, AWH^{+}24, ABL… 12, AAA24, AKPW05, BCCP25, Bar91, BG91b, BD00, BKL10, BRMG18, CM15, CA07, CS25, CZ14, CC15, CC16a, CvPS15, CGL99, DK25, DTI09, DYW16, EG94, FLV14, Ghe15, GHM23, GHM24, GKV23, HT21, HYW20, HWXC19, HJ21, JP14, JNS19, JLZZ23, KM24b, Kel07, KP25, KHM20, LS15a, LDH23, LZ15, LZ18b, LWN13, ML20, MsC20, MH23, Maj14, Mar93, MVV05b, MN08, MMLM20, ML10, Mor17, NIN12, NS24, NSM20, Osa92b, RB17, RS25, SP21, San25, SL15b, SL15a, SHLY18, SY20, Smo99, TZH24, UTO24, VBG96, Vul97, WZ13a, WQL20, WHD22, WLJ24, WLG26, YKY15, YWX14, YDWL15, YLY12, ZZ17, ZJJW24, ZXF14, ZH25]. singular-value [AWH^{+}24]. singularities [EAGS20, Gau12a, Gau13a, Gau13b, PSB26, Riz18, SL16, Tuo98, Wan24, XCLA15, Zha23]. Singularity [WLJ24, HCL21, Kol06, SK25, CL10a]. singularity/nonsingularity [Kol06]. Singularly [Bog02, AE09, AA09, BPR22, BS17, Bog13, CEX14, CXL16, CL93, CJ17, CG19, CJ20, CVX16, Das19, EA12, GO06, GO21, HCXL20, JKC26, KXXW21, KS20, KK12, KK16, KK17, KKS22, KKS24, KKD25, KSJ26, LS15b, Lin09, LRZ12, LLC20, LZ23b, LZ23c, MWDD24, MS24a, MN11, NMM18, NV21, OQ12, PM24, Rad08, RS20, RNT26, RBN14, SN22, ST99, Tem08, TKS24, WJC25, YM24, ZL22b]. SIS [YLYZ23, YH24]. Sites [KV04]. Sivashinsky [MK17]. six [KS20, Poc14]. six-step [KS20]. sixth [CHMT10a, He16, KCBT21, KL06, MClXzJe16, RWB09, WKG11, ZG12b]. sixth-order [He16, KCBT21, KL06, RWB09, WKG11, ZG12b]. Size [CVA01, ASHF21, BHS23, CMR93, CO19, FEL15, JR20, KJ15, KS06, KLR07, Lóc18, Ma20, Mal21, NAE22, PRVI20, Ste24, TG20, VH10]. sizes [MJ20]. SJSGM [KM24b]. skeletons [Lun23]. sketch [ZL26a]. sketching [RN21]. skew [BCW13, Don10, tFZyZ16, GV00, HM18b, PG15, WL17, XHZ07, XM16]. skew-Hamiltonian [BCW13, XHZ07]. skew-Hermitian [tFZyZ16, PG15, WL17, XM16]. skew-symmetric [Don10, GV00]. skewcirculant [Huc92]. Skewsymmetric [HR03a]. Slevinsky [SS14]. slicing [CR12]. Sliding [GH23b, GGN18]. slightly [LCW20]. slowly [Sab92a]. Smale [MA15]. Small [DWX17, FLT09, LZ23a, PW16, PM24, ST22, TKS24, VH12, VGV06]. Smaller [KLL25]. Smith [GSA03, LWwCL13, MK94, MK97]. Smoluchowski [Str05]. Smolyak [Gaj05, Pet01, PW04]. Smooth [CHS25, BDH^{+}13, CD25b, DLR12, GH22, GH23b, HLN25, HV22, KM24b, Lam09, LLS11, LCZ26, MH21, ML22, MMLM20, PV22a, SBR26, Śmi06, SC18, TZH24, Tru24, Wan19, YL16, ZA20, ZLLH22]. smoothers [DDT25]. Smoothing [CR14, DXY18, LB25, WD96, BG24, BL93, BL95, BP93, BC26, DFR26, DZH23, FY19, HLC15, LX24, MRS93, SG18, Wan24]. smoothness [God26, Iid24, MG22, Zhu21]. smt [RZR12]. snapshot [FZ26]. Sobolev [ADA07, BBR03, BHS17, CGM12, DM14, DHM12, DGP15, FVV25, GAM24, God15, HHLM23, HSL19, HLSLSL25, KLA24, Kow00, LMV25, LCGH23, LPP21, MVV25, MP22, Rei97, Ron92b, Sch26, YZ21, ZQzS22]. Sobolev-type [DGP15, HHLM23, HSL19, HLSLSL25]. Software [BHNS16, But02c, Coo03, Enr02, Vig04, ATM19, AM24, BRZ17, CDW95, DCM^{+}13, DCMM13, DHJJ10, KP96a, LP13, Mac96, Mil20, RZ99b, Stu97]. Solar [Sha02]. solely [HV22]. solenoidal [Han93]. Solitary [Nat07]. solitons [ZTZZ19]. Solodov [DMT22]. Solution [AMR23, AEG02, Bog02, Boy05, CM01, CMR03, FKMS01, GMT92, Jay02, KL25, Mar04b, MN08, PV03, AR24, AAH20, AR13, AHC05, ASS13, ABI20, ASGJ^{+}20, AH11c, AH18, AAD14, BOP98, BOL98, BC06, BAEF25, BEQOR14, BHW23, BT14, BBP17, BZS22, BEG25, BEH24, Bra07, BE98, BMR19b, BIMR19, BGI25, Bru93, CGPM00, CB16, CA07, CLGS17, CL05, CW14, CFL19, CCC25, CNR15, DMA09, DS09b, DMD16, DW97, DB06, DHV22, DI11, Fab16, FWC16, FH15, FS01, FH00, Gha16, Gha18, GK20, Góm99, HT16, HT23, Han96, Han94, HYW20, HCL21, HP18a, jHyPlZ06, HFZ19, HRY16, HM18a, IJE15, IJSS16, JVH15, JW25, JRRS08, KP16, Kac18, Kam15, KM24b, KL17, KS20, KS12, KS18b, KR20, Lee94, lLhYfD07, LM14, LA22a, LTFL10, LG95, MB09, MN23a]. solution [MD21b, MKA14, Mar93, Med10, MZW20, Mic23, Mil20, MDL15, MA13, MMLM20, MR96, ML10, Nac99, NLT21, PED15, PPS25, PV00, Pis16, PPR15, Pop15, Pop18b, Rab23, RC14, RR20, RGJ10, RM11, RMT13, RT19, RT22, RR98, RR00, RS25, Sal25, SW00, SLA11, SY20, SD26, SS15, Sho18, SKTGR19, SDMMK18, SMA99, SM25b, TS18a, TA24, TTA26, USAF14, WLL12, WMCW21, WK93, XWC26, YHQ19, YLD11, YYLX23, YYD14, ZR17, Zaf22, ZE12, ZM94, ZZY18, ZZ22b, ZL07, lZqL25, ZZY26, ZE10, dAFPR23]. Solutions [BS04, CGL01, EKM03, Gra03, HEOS04, MO04, NW04, Rec01, SFT03, WWD^{+}12, AGS08, AS10, AM24, AAIT94, AAII96, Ant22, Arn97, BDN17, BDV18, BC14, Bel94, Ben99a, BGS24b, BCJ22, Bou06, Caç10, CZL26, CLWH20, CMWP20, CDGC25, CZLS18, CMP22, DS09a, DGST15, EG10, Eba18, Eft15a, Eft15b, FW25, FHAL15, FS16, Gar24, Ghe13, Ghe15, Ghe16, GCPG99, GPAA14, Van17, GL19, HJ18a, HS26, HV15, HV22, Hua96, HM18b, HM19a, HZX20, ICR06, JP14, JS21, Kar15, KL25, KSCS07, KKO17, LS15a, LCZ26, LRL22, LM21, MH21, MV13, MV14, MW98, Met19, MS11, Nat07, NJ13, OIM21, OKP21, hPwL09, Pen13, PV98, PR25, QLZX11, SS24a, Sho18, SKTGR19, TZH24, TLD^{+}23, TY21, Tom11, TN10, VSA12, VN18, Wan19, WSZ21, WZ23a, WWL24, YHZL21, YXL18]. solutions [Zah09, ZA20, ZLLH22, ZLH22, ZMZ26, aZ19b]. Solvability [SD26, Le 98, WCW20]. Solve [PG05, AS11, AAN25, AAA25, ANA14, DP21, DW21, EH24b, JKC26, KSB08, MD21a, NAA19]. solved [SDL^{+}23]. Solver [Pop04, Som05, AF13, AR10, AVI97, BPS23, BW15, BBd95, EH97a, JRB17, Kub15, LZL20, MC05a, PV00, TY96, YYW21]. solvers [AMKV96, AAA24, DSS14, EY10, GH09c, GKV23, GRT97, HM26, HL20, Jia20b, MHR23, MRS25, NAR05a, NB16, PSW11, SS24b, Yal01]. Solving [ASS11, BAV18, BQO99, BH11, BD17b, BCV03, BKM03, BRSY21, CSFC04, DO25, EGSV04, EJR02, FY19, FT05b, GL04, Góm01, HVMMV24, HSZ03, HLV26, KM04, KSCS08, Kol04a, Kol04b, LZ09, Mai01, MM99, Mar04a, MS14b, PL04, Pié99, Pry98, SGM02, SM21, SD25, TRRD02, VRM23, ZS13, AAA^{+}18, AD17, AAM24, AIA^{+}24, AG23a, ATC16, AK12, AZ19a, AG19, ABI23, AVCO26, ATT21, ATT22, Arg10, AH10, AGS20, Awa10, BKDA25, BBS20, BMA16, BLW09, BGL26, BGR23b, BD17a, BNN16, BB14a, Ben99b, Bhr16, Bic24, BM97, Bla15, BS19, DOT21, BSF17, BC16, BEH24, BZ24, BMS24, BK08, CSI16, CQ16, CMD19, CCTV23, CX25, CKS24, CS25, CS94, CHH^{+}20, CCL18, CL06, CM16, CC16b, CC18, CW19b, CH22, CHYH24, CM25, CD25b, CH11, CRN19, Cho16, CRHTV24]. solving [CPS12, DWC18, DW24, zDYG18, DD21, DNV25, DSI11, DK25, Dax09, DOR26, DS20, DS26, Den14a, DHF21, DW22, DCW23, DHWL25, DH26, DMT22, DLYH17, DJG18, lDzS21, DLLD17, DLR24, Dzu13, EMS25, Ema96, EH97b, EH24a, ES19, EM07, Ess98, EGGSH13, EED19, FYM14, FHS12, Fra26, FR18, FS23, GS25, hGzS17, GPGC98, GP99, GNH10, GM97, GH22, GHZ25, GZL^{+}25, Hai09, Haj16a, HZPW23, HT19, HWD25, Hem94, HVMT17, HSTW14, Hey99, HSE16, HRAH22, HASI23, HL23a, HWXC19, Hua20, HCD26, Ibr25, IY15, IAAH24, IZ24, IDS16, IS17, Jbi93, JR20, JJ13, JWY21, JXZ26, JML25, JM18b, KBCG13, KKR25, KPD25, gK25, KKB16, KCBT21, KT07, Kha25, KXD25, KPA20, KLZV95, KZ21, LLAL21, Lei15, LMV00, LP18, LW13, Li17, LZ21, LLY22, LL22a, LLD23, LLM25, LSL25, LCZ26]. solving [LJ11, LZZ19, LDZ24, LZL22, LZL23, fLxX12, LWS18, LL22c, LA22b, LHW13, LM25, MZG26, MBJ17, MWDD24, MB06, Meh11, MXL25, Mil18, MDH16, Moo07, MLM19, MS13, MS24b, NBK17, NM14, NM26, NHP06, NRV23, NEMS14, NEY26, OMU^{+}24, OAMA22, OR24, PKC18, Pea13, PYD23, PS16, Pie96, PS09, Pop21, PSS22, ROB17, ROB18, RKMS16, RSKB17, REM21, Ree92, RZ16, RT20, RTCL21, RTD^{+}21, RT22, RTTH22, RT24, RAH11b, RZZ25, RB17, RALJ26, SKA23, SBA26, SS99, Sad99, SCD^{+}21, SP21, SSS21, SS11a, SEG14, SB21, SKD26, SA14, SIE16, SI18, SDW24, Shi96, SCDM20, SS24c, SR24a, SH21a, Śmi09, Śmi13, Sol11, SMN24, SXHZ20, SG23, SC18, TKSG23, TT21, TAM21, THF21, TQC22, TCO26, TG20, TH23, Ter22, TH19a, TH19b, TTLD20, TSI20, TDC21, TRSI23, TRCL24, TLR^{+}25, TDWL24]. solving [Ton25, THT19, VPA24, VPA25, WZ15a, WZQT15, WDL16, WHD22, WXT22, WX24, WZC26, WZL26, WB17, WC13, WWM21, WCD21, Wu22, XW18, XLW20, XCD23, XWL25, XH25, wYN18, YL19, YLL20, YF22, YDW26, YYL26, YZY^{+}14, YJX15, YP09, ZTW19, ZH22, Zeg25a, Zeg25b, ZLG^{+}13, ZY13a, ZFC18, ZYX19, ZLL^{+}21a, ZLH22, ZWXX24, ZHW25, ZLV17, ZV19, ZLT25, ZWW25]. Some [BO03, BS92, BRZ20, CC07, Cro03, Dum03, GP14, HM18a, HSS04, IMT02, Le 98, LZ14, LCVL18, LDL^{+}19, LM15, Lu15, MAL04, MM08b, MM09a, Mat96, MS13, Pas99, Pas03, PV03, SKK25, SGO22, Smo99, TH19b, VC92, XZZ19a, ZW15, – DK15, AHJ17, BD09, BBC21, BF17, BV99, BDS00, BRZ96, CK05, CS99, DL08, DSI11, Die08, Dra00, DLLD17, EDAH12, EP97, GEP19, Gau11a, Ghe15, GCFF95, HH11, HR00, Hua96, Jay21, JJK97, KC23, KK16, Leo08, Lin98, LSG15, MSCB93, Mon96, NHP06, NT21, Now06, Now13, OQ12, Osa92b, PS16, PP92, RR08, Sab92a, ST18, SKTGR19, Sid20a, SX00, Tam10, Tem08, WG99, Wim00, YYK24, Yal01, dAR06]. sonar [BKPS93]. SOR [CHYH24, MSZ20, Nie00]. SOR-like [CHYH24]. Sormani [BD10]. sound [DR07]. source [HV22, KLB10, LSZW19, LYL25, LL20b, LGM25, LC25, MWDD24, MT93, PED15, SMNZ20, YPL21, YXS22, YD26, ZR17]. sources [CKL16]. SP [NUG25]. Space [And14a, CP01b, DG05, ZL25, AD22, ADG10, AR13, Bag16, BAV18, BGS24a, BKF20, BZV16, BZS22, BHS17, Bou24, BIMR19, Cal20, CE94, CY19, CX20, CL21, Dam08, DN24b, FZL^{+}16, GSV96, HZPW23, HCL21, HL23a, HLTA16, HL17, HL20, JZF^{+}20, JWZ23, JJC24, JML25, Kaz24, KLF17, KMA13, Kuh13, LLAL21, LLX20, LLM25, LCZ26, LKBF17, LZ09, LWJ21, LRM16, LF25, LC25, MZG26, McL25, MDL15, MSS11, Not22, PNW17, Pot19, RZ16, RTCL21, RT24, ST22, SS24a, SKD26, SLA11, SDW24, SYLT14, TM10, VLCL16, WLMA21, WSLL24, WZC26, WCD21, YL19, YZL20, YPL21, YHL25, Zas22, ZWFY19, ZJWF18, ZJZ20, ZBX21, ZYQ^{+}21, ZZ22b, ZYW23, ZY23, ZZB20, ZYLN18, ZXL23]. space-dependent [Bou24, LC25]. space-fractional [BKF20, CL21, HZPW23, HCL21, HLTA16, JZF^{+}20, Kaz24, LLM25, LKBF17, PNW17, ZYQ^{+}21, ZYW23, ZZB20, ZYLN18]. space-Riesz [CY19]. space-time [BIMR19, FZL^{+}16, LLAL21, Not22, SS24a, ZXL23]. space-time-dependent [JWZ23]. space/multi [SKD26]. space/multi-time [SKD26]. Spaces [Mar04c, AUA22, All18, ATT21, AT21, BEL23, Buo17, CSI16, CSI17, CA07, CRN19, Cho16, DLR12, Dah93, DL18, DHF21, DKL15, ER19, GGNF17, GAM24, God15, GK24, HKPW19, HCD26, HM14, IAH20, IUM^{+}19, Jai16, Jai17, JA22, JM18b, KD18, Kow00, LP18, LCZ26, LLL22, LCL21, MWsC19, MAL04, Maz99, Maz09a, NEY26, OAMA22, PKC18, PG12, PDRG19, RT20, RTTH22, RZ23b, Rei97, SSS14, SCD^{+}21, Sch26, SBR26, SIE16, SLD20, Spi25, SC18, SCS18, TAM21, Tak17, Tan20, TB19, TLD^{+}23, THT19, VA20, WGK11, WKG11, WK12, WK17, XCD23, YAT20, Yse99, ZG12a, ZG12b, ZCTD24, van93]. SPARK [Jay02]. Sparse [ABG97, BP19, CP00b, De 02, DR01, HL03, HL15a, NZ19, The97, AFG26, AA15, BDN17, Bai97b, BG91b, BF99a, Bou17, BRSY21, CKM^{+}25, CM25, CCW21, CL96b, DDRT97, EKPU23, ED22, FS20, FP18, FGS24, GSZ22, GG98, GHP^{+}00, GLW13, GZL^{+}25, HL06, HP18a, HIK17, Jón93, KKV22, KD14, KMA13, LWC^{+}21, PLH20, PW16, PWCsL18, RWTM21, SS99, Sch14, ST18, SI13, Spr98, Spr01, TYSY20, TL23, UL18, VBG96, ZY21b, ZWZ25, ZZLV23, ZXL23]. sparse-grid [Bou17]. sparsity [DHMS16a, LZ22b]. Spatial [CZZ^{+}25, KK00, DZ13, FNS19, FHC21, He16, KPT23, MS24a, OKP21, PYD23, RW24, YWYN22, YY24]. spatial-temporal [YY24]. Spatially [LZ25, GA20, MM22, SYLT14]. spatio [HOW95]. spatio-temporal [HOW95]. SPD [Spi25]. SPDEs [NT21]. Special [Ano95b, Ano17, DJS20, FH05, GST03, CP93, CP95b, CJTW96, GG08, HVMT17, IY15, LHZ10, LW16, Ter23, Tom92, XX16, ZMZ26]. Special-purpose [FH05]. specified [ANI15, KP96a, LM97, OO22]. Speckle [WHZ^{+}18]. speckled [CTS09]. Spectra [BG03b, PP05]. Spectral [ABI23, ALZ20, Ano05b, BD09, BO03, BK04, BFGM03, DG05, DR01, EKM03, GHM08, Ghe16, GK21, HM19b, KJO23, LS03a, jLyLqW17, LRGH02, MR02, NEMS14, PR10, PG05, SSC25, TC05, Tem08, XZW13, YSXY19, AR24, AH23, Ahu09, AAN25, AM21, ABI20, ABM10, AHC13, ACH17, ACH19, BD10, Bhr16, BZV16, BK18b, BMR19b, CM24a, CMR16, CG07b, CB13, CC13, CFR19, CFK^{+}20, Cui26, DB06, DBAE09, DS09c, DK00, EV22, EED19, FZ26, GS16b, GNT24, Gu20, Hei06, Hei07, HPS97, IAAH24, JRRS08, KP96b, KM24b, KBA23, KBA26, La 17, Lam09, LRL19, LZlL20, LZ23a, LLL26, LCH20, LSW16, LJY24, Lya97, MR12, MT18, MG11b, MSMS12, Mot14, NA24, PP24, RM26, SBA26, SKK25, SLW13, SGS25, SGS26, SMK14, Spi25, TZH24, TTXZ23, TZ24, VRM23, WLMA21]. spectral [WC24, WZ15b, YH21, YWZ19, YYW21, ZA20, ZAGD22, ZHT15, ZJWF18, ZJZ20, ZWX22]. Spectral-Galerkin [ALZ20, DB06, TZH24]. Spectrally [BIMR19, GH06]. spectrometry [MvS09]. spectroscopy [MvS09]. spectrum [ACE99, CR12, Cao12, EV22, GNS22, KKK22, ZYBJ23]. Speed [FS01, BGRS09]. SPH [MFK^{+}15]. Sphere [ASS03, NSS03, IP16, LK20, Leo07a, PSS22, TV19]. spheres [HL15a, Sun94b]. Spherical [BBC25, CM24b, ALZ20, BBL22b, Xue95]. spheroid [TCOA19]. spheroidal [SHF15]. Spline [CMRS01, NSS03, Wri95, AB06, AZ19a, ASS11, ASS13, ART19, ALRÁY25, ACM93, AGN07, BDIR18, Bia94, BK13, BF17, BFKM20, BFK22, BL93, BL95, Bra96, CFR06, CM24b, CW14, CCD10, DLR12, DFR26, FW25, GGN14, GGN18, GZ18, Gau17b, Gha16, Han93, HM07, JJ24, KS18a, KR11, KS12, Kim21, Maz05b, Mic23, Mil17, PZL15, Plo93, Plo94, Poc14, PGP^{+}25, RB21, RM11, Rei97, Tra93, UTO07, WZ19, WD96, XL14, ZR17, Zah09, ZE12, van93]. Splines [FKMS01, HL02, Kva01, MSCB93, NS01, PS01, AHR21, ALY22, BKPS93, BCM19, BV21, BR07b, BLS06, BM14, CH95, DDS93, DS15, DZH23, GLRSG08, Gre96, KK00, KP09, Kim21, KK23, Kob97, KR07, KL91, Kva14, Lai92, LMM97, Lev95, LRZ12, MRS93, MS01a, MST03, Maz09b, Maz11a, Maz11b, Maz12, Maz18, MRU91, MT06, MS96, Pow93, Rab92b, Rab92c, Rad08, Sch08, TM05, WD95, de 93]. Split [PKC19, WZ19, AAA^{+}18, ALJLYJ24, Anh19, Buo17, CCL18, CRN19, zDYG18, Den14b, HH12, HMT17, HS20, HRAH22, IUM^{+}19, KAF18a, KAF18b, KPC20, KD18, LCL21, MWsC19, MN17, MG18, OIM21, RIAA19, RT20, RT22, RTTH22, RT24, SCD^{+}21, SIE16, SDMMK18, TKSG23, TAM21, Tak17, THF21, TQW24, TG20, TDC21, THT19, Wan17, Wan18, ZH19, ZW15, ZMZ26, CGR12]. Split-step [WZ19, HH12, HMT17]. Splitting [EH24b, Li95, LHW13, PT17, vdHM98, AG19, AABTB23, AKT15, AKKT16, AKQ17, BC16, CW19a, CXC25, Che16b, CZM21, CS22, Cho16, CHK14, DW24, DS20, tFZyZ16, Fan22, FMD23, GS25, GS19c, HWXC18, JKC26, JCH23, KM17, KMZ18, Ke21, KSJ26, LY17, LM17a, LM19, LLD23, LKBF17, LZL22, LZL23, MPS20, Met19, MG20, MG21, MWZL23, Mor17, NMM25, OMU^{+}24, PG15, RNT26, RWTW19, RCW22, SBG24, SPB26, Sti18, SS24d, Tai92, TH23, TCOA19, WS24a, WL25, WCZ25, WS24b, WHS23, WL17, WPL18, WL22b, XM16, YZL25, ZWWW20, Zha11, ZFG18, ZGK^{+}25, ZY13b, ZLV17, ZV19, ZV21]. Splitting-integrating [Li95]. splittings [GV00]. Spot [BvLP16]. Spurious [ZLH22, IKMS24]. SQP [GZ11, GZP18, LL07, mTLbJlL14]. Square [LL05, Mar04b, MST03, DM97, Gau12a, Gau13a, Hig97, KGG25, ST23, SHLY18, WCHK21, YCW^{+}19]. Squared [RVF07, CX25, CS94]. Squares [BCV03, DG05, LV01, PG05, SK04, AAH20, AAD14, BGS24a, BRS09, CM24b, CX25, CR20, CL24a, Dax09, DMA19, DWX17, Ell93, FGS24, HS26, Han22, HM22a, HM22b, HYW20, Hei06, Hei07, jHyPlZ06, HM19a, KC23, LV15, LDN16, Lei15, LL22a, LJ22, MMV19, MMGH17, MRS93, MZW20, MZZ25, MXL25, MS24b, NC94, NK21, NA24, OC24, OLB94, PT24, Pen13, PS25, RAH11b, RALJ26, Rum14, SEG14, ST18, SM21, SMN24, SXHZ20, TV19, VZ93, Wu22, XXW17, Yua21, ZH22, ZLWZ21, ZW22, ZZ22a, ZLZ23, ZHW25, ZLWQ09, ZMZ26, ZWW25, LMV00]. squares-scaled [ZW22]. squaring [AG25]. SR1 [And22]. SSOR [LZ16a, Tan17]. SSOR-like [Tan17]. SSS [Sha19]. stabilisation [MT12]. Stabilised [GM03]. Stability [Bea98, BBd95, But02a, BH02, BJ04, CMP21, LC14, lLLVZ17, LER03, RW11, SPB26, SKD26, VV25, Wei17, Wri02, ZCT19, ZD21, ZLLH22, ZKD04, iV12, ASV23, AUA22, Ali23, BRS26, But96, BH05, CCN25, CC12, CJ12, CCJC18, DL18, GL19, HH11, HKE97, Her96, Hom98b, HCH18, HWZ24, IKMS24, IJ19, KCBT21, KMH25, LLY26, LCZ26, LM17b, LWS18, LW22, LZX23, Lóc20, Lya97, MM22, MV17, MHA16, SAHM25, Smi97, SHGL22, Tom11, WZ23a, WC25, XF25, Yal01, YY24, ZY21a, ZYQ^{+}21, ZY23, ZSLZ24, ZFX14, ZXLF15, dC16a, dC16b, dC22]. stability-preserving [HWZ24]. Stabilization [Asc97, Axe99, FT24, CEK21, YJ18, YZX25]. Stabilized [WSY12, ASZ23, GG22, Kno23, LBF24, LL20c, TX19, WD23, ZD21]. stabilizers [WYZ21]. stabilizing [DI11]. Stable [AMCM06, BOL98, BGL07, CCG01, Hig97, Kob97, LPRÁ24, LJW17, PT18, Ver99, Bec96, BQO99, CLPY23, CLY25, Con93, CKKT16, VV07a, VV07b, FYYW19, FGR01, He16, HP18b, Hil10, HNY^{+}18, HY21, HSY23, IP16, JTW25a, JTW25b, KS20, KP09, Kim21, LZM23, LS20, Liu21, Met19, MNS23, Mon96, NL97, NN26, PSS10, PJ22, PFT98, RF23, RhG15, SCDM20, SKTGR19, SR22, Sle25, SSP15, Sol15, Tou98, WZ26, WCD21, ZLT^{+}17]. stage [Bai97a, Bai97b, BZ13, CM16, FGR01, JSZ22, KM19, LG19, NB16, SKTGR19, VV11, Ver14, WW19, Zha24, ZJSS25]. stage-order [Ver14]. Stancu [LQ16]. standard [WT08]. Standardization [DHL^{+}04]. star [DO25]. starlike [NPR08, Røn92a]. start [SW10a]. Starting [HR00, CL19, MM19]. State [GF02, KW04, BM24b, CL21, GSV96, KGH14, LCH20, MB09, MP25, MB06, OMW21, PNW17, QZG^{+}19, WZ26, ZLCW23]. state-constrained [OMW21]. state-dependent [KGH14, WZ26]. stated [SL15a]. states [Smi97]. static [JH22, KKPT16, PS20, Uhl22a, Uhl22b, ZZ10]. Stationary [JWL20, YXS22, CV15, DEC24, DZ24, Elf25, MP13, SKSS21, SFZ22, SW07, VN18, WZZ15, XY25, XL24, XH21, ZH17]. Statistical [YD26, DWX17, TA96]. Statistics [KPFG04]. Statistics-Based [KPFG04]. steady [BRY14, BM24b, CL21, FLH25, MB09, MP25, PNW17, QZG^{+}19, Smi97, ZZH15]. steady-state [BM24b, CL21, MB09, MP25, PNW17, QZG^{+}19]. steady-states [Smi97]. steepest [BN18, CTY25, HSK20, Maj13, SS12b, You16]. steering [Hag13]. Stefan [BZS22]. Steffensen [AHVR17, EHVRV14, MCC26, SS24b, SS24c, WZ13b]. Steffensen-type [SS24b, SS24c]. Stein [FWC16, LWwCL13]. Steiner [FM04]. stellar [SPS26]. stencil [ABT07, KLL25]. Step [BO02, CVA01, LL22c, AG15, ATC16, ANI15, AMH10, AMA21, ASHF21, BKA19, BD09, BCST14, BC94, BJ98, BFK22, BV09, CMR93, CW19a, CS94, CZM21, CJ17, CHMT10a, CO19, DJ10, DFJP10, DS20, EH24a, FHL21, FLY25, FR18, GZD24, GZL^{+}25, Gut15, HH12, HMT17, HVMT17, JT96, JZ25, KPD25, KNBGV18, KM24a, KS20, Khe12b, Khe16, Khe17, KH20, KS97, KS06, KLR07, LW16, LG17, Li17, LDW18, LZ22a, LL26b, LLL22, LYY12, LS07, Lóc18, Ma20, MsC20, MBJ17, Mal21, MR09, MJ20, Mat15, Meh11, NSM20, NW19, NB16, NAE22, PRVI20, Pot19, PM24, QZG^{+}19, RKMS16, RR20, RWTW19, SGS25, SGS26, SW10a, ST17, Sha97, SAHM25, SSH^{+}19a, SS15, Sho18, SKTGR19, SSK23, Ste24, TG20, TZ25, TS15, Tou98, VH10, WZ19, WQ23, WCZ25, WHS23, XW18]. step [YAT20, Zha11, ZV21, ZS22]. Step-by-step [LL22c]. step-size [Lóc18, Ste24, VH10]. Stepping [Söd02, CEK21, IJE15, lLXhL22, MM23, SN22, ZQS24, ZZ18]. steps [ALV20, BGL07, HLL22, HVM15, KH20, LlLZ21, MVVV24, OMU^{+}24, WZZ16, Wan17, XH21, vSv94]. Stepsize [DZ01, ZQL^{+}19, AP26, ABS19, BL98, BC01a, JT96, JA22, LM17b, LL18, MN92, PLH20, VH12]. stepsizes [ZHSX23]. Stetter [GPGVS92]. Steven [BC22]. Stewart [HC03]. Stiefel [CO03, DW12, ODL21, TOY26, WCY25]. Stieltjes [KJG23, Not95, TBA94, Van92, Van03, dC23]. Stiff [AHKW05, Che02, FG03b, AHP20, AAP24, AHP25, AMCM06, AH15, AM98b, BM22, BC01b, Che14, FYYW19, HH12, NBJA17, NB16, SS24d, TX19, WCM94, vSv94]. stiff-cut [SS24d]. Stiffness [ST02]. Stirling [Mor11]. Stochastic [ACM04, BG24, CCC25, MAL04, MHZ05, Vig04, WLL26, ZS03, ABB15a, ABB15b, AAA25, AH15, BGL26, BD17a, BM09, BZ24, BB14b, CCN25, ÇD25a, CSC26, CM96, DS21, DI11, El 18, ES19, FWC16, GZ18, GMY18, GL19, GZD24, GHZ25, HH12, Haj16b, Hal14, HLZ14, IS22, IZ24, IJE15, JLP20, JWL20, JSZ22, Kam15, KJ18, KR23, KL22, KS23, KLL10, KLS17, KLR07, LLS11, LDH23, LXX23, LGHS25, LHL11, LM17b, LT24, LWZ21, LXP20, Mal21, MKBY19, Mil19, MNS23, MG22, NM26, NAA19, PM26, Prz16, ST17, ST23, SSL24, SR24a, SS16, SMN24, SHGL22, SZ20, SZ21, TX19, TS15, VH10, VH12, WG13, WZ26, XT16, YBK^{+}21, YLYZ23, YH24, Zeg25a, Zeg25b, ZSLZ24, ZHW25, ZL17, ZJ23, ZHZ25, ZJSS25, ZZY26, aZ19b]. Stokes [DZ22, ZD18, BK18a, ADP26, BRY14, BGR23a, BW13, BCJ24, CW21a, CC18, DZ21, DMW23, DZ24, DSS14, FRJT09, FLH25, Hei06, HM18c, JZ25, JY23, JM00, KM17, KPS22, KP22, LS25a, LL05, LHR20, LD21, LBF24, LC21, LN22, Med10, NAHZ21, Oua99, Pei24, PV98, PV00, PG05, Sch25, SW10b, Si12, Sla06, WD23, WDL23, WQ23, XM16, XH21, ZBDK23, ZDZ26, ZS19]. Stokes/ [DZ21]. Stokes/Darcy [XH21]. Stopping [KS25, CL19, RSZ20]. storage [LXQ15, Pol10]. stories [Maz05a]. Strang [CS22, GZ20]. Strategies [CR12, HR03b, NPS09, CR20, RP26, TBC^{+}23]. Strategy [MT04, Aih17, AAN14, AK15, ABM10, BuR24, BRZ96, KZS21, KH18, KSW07, LL26a, TS92, Wan15a, WS24b]. Stratonovich [XT16]. stream [ZHY^{+}20]. streamed [HIS26]. streamline [ST98]. Strengthened [AAAS03]. stretching [FHAL15, Mot14, NZF11]. Strict [Ceg24, Cat24a, Cat24b, LCL21, Zaf22]. strictly [GLS26, GMP92, LZX22, OAR22]. string [BRZ18, BGR23b, MD21a, RZ16]. string-averaging [BRZ18]. strip [ZW14]. stripe [WZW^{+}24]. Strong [CSI18, KMH25, KLL10, LQ20, SI17, SSL24, TTLD20, TLD^{+}23, WC24, WZ26, XCD23, YL19, YH24, ZW15, Hag13, IJE15, IJ19, JLP20, LGL23, LLQ17, NT21, ST17, SHGL22, TH18b, TVC20, TSI20, TDC21, XT16, XWL25, XZZ19a, XZZ19b, XP23, YCZ26, ZW12b]. stronger [SI13]. strongly [Hie18, Jos22, LYL15, PT17, QLZX11]. structural [CDS20, GHP^{+}00, Pis16]. Structure [BH17, BB14b, HLJ25, KL04, YSLL23, ANI^{+}17, AR10, Bla15, CK20, CMWP20, CV92, DS21, DLL^{+}24, Guo16, GL23, GJVW24, JL16, KW00, LZ23a, LHlLG24, MZZ25, Men26, SLZ25, WZVJ22, WLW^{+}24, ZZ22a, ZJJW24]. structure-preservation [LZ23a]. Structure-preserving [BB14b, HLJ25, YSLL23, DLL^{+}24, Guo16, GJVW24, LHlLG24, MZZ25, SLZ25, WZVJ22, WLW^{+}24, ZZ22a, ZJJW24]. Structured [CPV04, DP16, FG07, Faz23, LMV00, LV01, LZ22d, Ali23, Bno21, VVV22, CCZ23, CJKL23, EP97, ED13, FFSC25, JK19, LAN18, LDC10, ML22, NA24, PDS^{+}23, RZR12, SL15b, SKJ^{+}18, YHZL21]. structures [BDS25, DCG25]. Student [HHLS21b, HHLS21a]. studies [DW21, JJK97, Wan25]. Study [CGN03, QLZX11, dSCS04, AWH^{+}24, AM16, CM98, CB13, CN17, DL97, FS20, GH10, Gau11a, GL20, Hua21, JCF15, Jaw22, KLA24, Kar07, KK22b, Ma20, Man10, MWWY13, PP18, PH20, SHF15, SW14, Sid20b, SS12b, SA26, WSZ21]. studying [Bel99, SD26]. Sturm [AS10, AA12b, GHC15, JLX22, LJX25, LR14, Mar93, TBY13]. Sub [Gau12b, Bhr16, CLTA10, LR18, LJWW21, MM09b, MA12, PSB26, WLY^{+}21, YWZ19, ZP17, Gau17a, Gau19]. sub-diffusion [Bhr16, CLTA10, LR18, LJWW21, MM09b, MA12, PSB26, WLY^{+}21, YWZ19, ZP17]. Sub-range [Gau12b, Gau17a, Gau19]. subanalytic [Ple12]. subcells [CL24b]. subclasses [CK06]. subdifferentiable [HSK20]. subdiffusion [CLT^{+}13, CC16b, CSZ22, LZlL20, LlLZ21, lLLVZ17, LCZZ23, LZQ26, RR22, YJX15, ZF22, ZZY24]. subdiffusive [MM22]. Subdivision [CJ04, MM04, ARTY20, CCS05, DL08, DXY18, JJ24, MKA26, PT18, le 91]. subdomain [AL23, XH21]. subdomains [Cal20, GL15]. Subgradient [AT21, AN17, ATT22, DJG18, JA22, KJO23, NEY26, OR24, PDS^{+}23, PK22a, PK22b, TQC22, TH18a, TH19a, WZC26, YLL20, ZFC18]. subgradient-type [PK22a, PK22b, ZFC18]. subgradients [ML22, SR06]. Subinterval [Cse04]. subintervals [XL14]. subject [LZ18a]. subjected [PPS25, RS25]. sublinear [AK00]. submatrix [LHZ10, Pen13]. Subquadrature [KV12]. subrange [Gau18]. subroutines [Lai92]. subset [Sab91]. subsets [Dam08, Lóc20]. Subsonic [Lui02, KLB10]. Subspace [Ali23, EJR02, Jbi03, KP26, MR02, Aih17, And14b, AFN16, AFN17, BBS20, BLW09, BEHS20, BMS24, Bre99a, BPR21, CH26, FEL15, HE05, HRY16, HRY19, JL15, Lam09, LLL18, LLWC24, Lun23, Pan20, PV98, Saa23, Sch17, SWB08, SWJ24, SLL22, SLL23, TY96, Wu22, YCL17, ZLL21b]. Subspaces [Amo02, BCN^{+}16, AC19, BM19, HPS20, LWG18, MVV05b, Oar94, RR00]. Substructured [CV22, CGV22]. subsurface [Hol98]. subtraction [YKY15]. subtraction-free [YKY15]. Successive [DFK97, Bic11, DLL^{+}24, Nie93, XL14]. Sufficient [CL10b, Lin05, Mat01, AIA^{+}24, BE20, LWQR15, MV17]. sufficiently [AB23]. suitable [OOO11]. Sum [INR01, Ber14, Kar07, LV15, MS26, PYD23]. Summation [DH04, Mil95, SVZ01, CFR06, CW08, NW17, Sab92b]. sums [DPR23, LI10, Nor00, VC00]. super [CAV23, FZ26, Hag25, WGK11, WSK14, WK16]. super-Halley [WGK11, WSK14]. super-Halley-type [WK16]. super-linear [Hag25]. super-resolution [FZ26]. Supercloseness [LZ23b, ZL23]. Superconsistent [Fun01]. Superconvergence [FHC21, LS20, NNCN23, SL21b, SN22, WLZ22, Bac21]. Superconvergent [RM26, AAA24, Bac18, Bac23, PLZ^{+}24]. superfast [CL92, KV00]. superimplicit [FH15]. superiorization [CHS19, CGHH21, HSK20, NAE22]. superlinear [AK00, CK24, Cat24a, Cat24b]. supernodal [FP18]. Support [ASS04, MST03, BP19, PW16]. supported [GHM08]. supports [HPS13]. suppression [HFW^{+}21]. sure [GL19, LM17b]. Surface [KGD03, ASZ23, AG00, BBL22b, CLMM05, CL24a, FRR07, FMS25, Gal93, HLS10, JJ24, MW24, MW26, MKA26, NZF11, NPR08, PN93, WZZ07a, WZZ07b, WD95, WD96, ZJ08]. Surfaces [KPT03, BE17, BDH^{+}13, BMV09, CZH22, CDF99, CLaL00, FHAL15, GLRSG08, HMdÁES08, HSSB13, LN95, LKW17, MHR23, Moo20, PTSB01]. surfactant [XCY21]. Survey [FAM25a, FAM25b, BFK11, Lor95, ST23]. susceptible [ZHZ25]. susceptible-infected-recovered [ZHZ25]. Svaiter [DMT22]. SVD [CFM15, HC03, JCL16, Jia06, LXQ15, MRS06, NS22, Nov26, OR17, RSCH^{+}19]. SVDs [Jia20b]. swamp [VVL25]. sweep [LZL22, WL25]. sweeping [CEZ26, ZZ10]. Swift [Liu21, YK22]. switched [WZS14]. Switching [BBO21, CDLW21, Che22, Che24, CGSS24, PM26, SPB26, WZ26, YYK24]. Sylvester [BEHS20, BEH24, CZL26, EJR02, HM18b, HM19a, Mit11, RS02, SDL^{+}23, TM10, XWC26, YC22, ZTW19, ZLL^{+}21a]. Sylvester-conjugate [HM18b]. Sylvester-transpose [ZLL^{+}21a]. Sylvester-type [CZL26]. Symbolic [Bel99, Ben99a, DVJBN03, INR01, JJ13, Mil17, MCMX20, AB24, CHYZ98, CM98, EAB20, MdR08, MdR13]. symbolic-numerical [CHYZ98, CM98]. Symmetric [ABQ04, FG03b, HR03a, AMM11, AG23a, AAAA^{+}18, ADL05, AA16b, AL97, ADGP15, BBC11, BOP98, BBP17, BP23, BK16b, CR96b, CGV92, CHY19, CG17, CC16a, CW19b, CC13, DMC20, DS20, DPP22, Don10, EG19, EH24b, FLV14, FP18, FGL19, GV00, GLG26, HT21, Haj16a, HS26, HSE16, jHyPlZ06, HM19a, JFSD25, JRRS08, KP26, KKO17, KH18, LL93, LLZ94, LHZ10, LYW14, LM19, lLhYfD07, LZ19b, lLHNS23, Man17, MM08a, MVV05a, Meu09a, MS14a, Mia19, MVG21, NMM25, Orb15, PL99, RWTM21, RSCH^{+}19, Saa23, ST22, SMZMA18, Sho18, TT06, VGV06, Vos00, WL17, XW18, XHZ07, XLW20, YZLP16, YSLH19, ZZ17, Zha15, ZD17, ZFG18, ZHY^{+}20, ZYW21, ZWXX24, ZFX14, ZXLF15, ZYBJ23, ZM16, ZZZ22, ZCG15, ZS08]. Symmetries [Kha14]. symmetrization [AE18]. symmetrized [Fue07]. symmetry [Vos00]. Symplectic [BI14, MO19, SSH^{+}19a, VV11, ZTZZ19, AMM17, AMM18, BSB23, BRS26, BH09, But15, DS09c, FGR01, JLZZ23, MKBY19, MO10, Sal05, WW14]. Symplecticness [Jay21]. Symposium [Ano95b, The97]. synchronization [Lun23]. synchronous [BZ13, XZP^{+}20, ZZLV23]. synthesis [DR07]. System [DVJBN03, FKMS01, Sha02, WHL24, BK18a, ALJLYJ24, AA16b, ACE99, AG23b, AT17, BMA16, Ben99a, CG20, COSE22, CW14, Che22, Che24, CDGC25, Das19, DS26, Den14a, FY19, Fra26, GNH10, GLdO09, GPAA14, GM97, Gu20, HBP13, HYW20, HDP18, HSE16, HL17, HM19b, HG25, JH22, KS12, KE16, KMV17, KK12, LHR20, LS15b, lLHNS23, Lin09, LCW20, MBR21, MBJ17, MV13, MW98, Med10, MMU20, MLM19, NRV23, PJ22, Pat25, Pog98, SD25, SK19, TY96, TN10, VTV22, WB17, XY25, XCY21, YX11, YZX25, ZDZ26, ZS19, dAFPR23, CHYZ98]. Systematic [Loh22]. Systems [AZMJ04, AAAS03, Amo02, ABQ04, BKM03, Bre02, BZ02, CNR15, DMRT03, GL04, GM03, HL03, KM04, KW04, Kol04b, Mai01, PL04, RT12, RST03, SGM02, Sch02, Tsu02, AEH20, ATC16, AAB13, ABG97, AL97, AKB15, AH15, ASGJ^{+}20, Ano17, AH11c, AGS20, AR10, AM98b, Asc97, AKQ17, Awa10, ANA14, BDN17, Bai97a, Bai97b, BBC11, BCW13, BHLZ21, BH22, BC94, BOL98, BAK25, BW93, Bar91, BBO21, BNN16, BHNS16, BDV18, BC06, Bel94, BHS14, BH17, Ben99b, BKF20, BBP17, BBd95, BMS24, BS92, BRZ98, BRS08, BRSY21, BDR25, BKL10, BRMG18, BWB24, CL92, Caç10, CR96a, CR96b, CCTV23, CS94, Cdv98, CAV23, CC16a, CL05, CW19b, CK20, CH22, CM25, CG19, CJ20, CDP16, CRHTV24, CBGVN07, DWC18, DEC24, DK25, DS20, DMD16, DTI09]. systems [DLL^{+}24, DEM93, DP16, DLR24, Dum13, DHMS16b, DNR17, EHTSM21, EH24a, Ess98, EGGSH13, FLY25, FGC19, FP18, FGL19, FHH96, FT05b, FS23, GZL24, GP99, Gar19, Gar24, Gar20, Ghe16, Gon16, GO20, GCPG99, GSV96, GZP18, GH22, GH23b, Hai09, HKE97, HWD25, Hem94, HR05, Hey99, Hey01, HE05, Hoa25, HFDSC24, HZ15, Ibr25, JKC26, JRB17, JMS16, Jay21, Jbi93, JJ13, JW25, JH22, KBCG13, KKR25, gK25, KCBT21, KL22, Kha13, KS97, KKO17, KLL25, Kub15, KK16, La 17, Le 92, LW14, LYW14, LS15a, LG17, LG18, LDW18, LM19, LG19, LC19, LZ22a, LZ16a, LZ19a, LXZZ21, LA22a, LZ19b, Lin01, LN10, LW12, LWZ21, MD21b, Man21, Man10, MM99, McL25, MS06, MZ99, Mil18, MS11, Mor17, Müh99, NBK17, NMM25, NYL26, OOO11, PS16, Pié99]. systems [Pop21, PG15, RSKB17, RM13, RW11, RG10, Rum14, SKA23, SBA26, Sad99, SSR25, SW00, ST22, Sei98, ST23, SCTP00, SGS13, SA14, Shi96, SD26, Sid20b, Sim98, SH21a, SS16, SS24d, TT21, THF21, Tom92, Tuo98, USAF14, WG99, WW14, WZQT15, WGZ18a, WHD22, WX24, WHS20, WC13, WZS14, WL17, WZ23b, XW18, XWY19, XLW20, YWX14, YHZL21, YZY^{+}14, YYD14, ZA20, ZZ17, ZYW17, ZHSX23, ZD17, ZFG18, ZYW21, ZWXX24, ZWZ25, Zha26, ZL07, ZM16, ZZZ22, ZL26a, ZCG15, DM21]. Szász [AUD18]. Szego [BGVHN92c, CBGVPP09, DPS18, GHM08, HLM04, JNW92, Not08, Pej14, Røn92a, Tru24].

tables [Sid94]. tackle [MS26]. tail [LS11, SS08]. tailored [VdR13]. tails [EO94]. Takagi [KC24]. Talbot [DW15c]. tale [LS11]. tamed [JLP20]. tamed-truncated [JLP20]. Tanabe [gK25]. tangent [DHS09, KL94, YZBJ21]. tangential [Kao20]. Tanner [PP05]. target [BM12]. Tau [ED05, AH23, BZV16, HMT17, RM13, SW14, EDAM13, GH09b, MG11b, Mok16, VRM23]. tau-leap [HMT17]. Taylor [BB14a, Bün18, Bün24, Her96, MTTC22, NKS04, Pry98, SL18, SL21a, ZFZ19]. Taylorian [le 91]. tearing [TFPG19]. Technique [RST03, AS11, BuR24, CDD21, Deh20, Gás99, Han93, HL23b, HWC15, ITA24, JFSD25, LZ18a, NM14, OR24, PXH25, RIAA19, SPV20, SKK21, SS24d, WZ16, YD09, YZ17, YZX25]. technique-based [JFSD25]. Techniques [DR01, Kun01, Lig93, The97, PGSU25, ALB^{+}18, AT19, BW93, Cro92, GZL24, JCF15, Li96, LWLT19, LA22a, LDC10, lLXZ24, MB09, MN23a, PLVB11, RS26, SKK25, SW00, SCTP00, SLA11, TFPG19, ZH25]. telegraph [AH23, He16]. telegrapher [SM25b]. temperature [NM14, ZZZ20]. tempered [CDLW21, CGSS24, DZW17, LD20, QQX23, TH23, WQZH24]. template [Kub23]. templates [FHH99]. Temporal [lDzS21, ALB^{+}18, HOW95, KP25, LR18, Rou25, SzS21, YY24]. temporal-fractional [Rou25]. Tension [RM11, AZ19a, BV21, BR07b, KS18a]. Tensor [ACM93, BEJS21, BG03a, GGV02, HZ24, OS25, RS21, XQZ24, ALL25, BJNKR20, BEHS20, BEJR23, BJT24, BH01, CCZ23, DBV23, DLDW21, EJR26, GLL19, HIS26, HPS13, HLV26, KP26, LHZ^{+}21, LZZ19, LDZ24, LM25, NYL26, PS25, Rab05, WCW20, WMCW21, XWC26, wYN18, YSXY19, ZJR25]. Tensor-Product [GGV02, ACM93]. tensor-structured [CCZ23]. tensors [BHLZ21, CHY19, CPRZV23, ED22, Hua21, IDAV09, LLQ17, San19, WCW20, XZZ19a, XZZ19b]. tenth [EM07]. tenth-order [EM07]. Teodorescu [Caç10]. term [AAM24, And14b, And15, BE20, DP21, Don16, lDzS21, GZL24, HFZ19, IAAH24, JLJ22, Khe12a, KBA26, KLB10, LSZW19, LWLT19, LZ21, LYL25, LSY^{+}23, LF25, LC25, MJJ^{+}23, MKBY19, MWDD24, PED15, RS26, Sal17, San25, SSP15, SzS21, TZH24, TLD^{+}23, WLMA21, WYP23, WZC23, Wei17, YPL21, YXS22, YJJ^{+}21, YYZ22, YD26, ZR17, ZD15]. terminal [GK21]. terminating [CW08, TM14]. terms [BHW23, KC23, Oua99, SGO22, SIO20, Śmi09]. Tessellation [dADdRC04, AT12]. Tessellation-Based [dADdRC04]. Test [BD04b, LM04, BX19, CKM19, GHN19, Mac96, Mit11, OO22, Ste20]. Testing [LW95, GL12]. Tests [WH04, HÇK26]. TGNN [XWC26]. TGV [ZL24, ZL26b]. th [BHM05, CL11, CVX16, Fue07, Gaj05, GL20, MKA26, ZXRL11]. Their [Gla04, Mas05, AR20, BCST14, BL92, BF18, BHS14, BH17, Ber10, BZ91, BGVHN92a, Ceg24, GMZ19, GL07, GG08, Gau11a, HYJ20, KXXW21, KCBT21, Kim21, LXZZ21, LSY^{+}23, Mar92, MSS11, Ron08, SH26, YWS20, ZAGD22, ZWL26]. Theorem [MKO04, BGVHN92b, CL24b, CS08, Low05, LR14, MA15, Mel14, SK19, Tac12, TVC20, UA09, WZZ16, XWL25, dC23]. theorems [AAAGAD23, CD00, MW16, RTD^{+}21, TH18b, TSI20, dB07b]. theoretic [DI11, VB92]. Theoretical [Iid24, Lis25, NS24, Sal05, ZAGD22]. Theory [AHKW04, MT98, BGR23b, BC06, BESC22, CR00, CF05, CZ96, CVLX19, CFS21, CD00, Cuy00, Cve06, Guo16, KP96a, KADE18, Kul10, LG08, LL26b, MKG24, MA15, MA12, SS11b, ZYW17, dBGKR08]. there [Lev05]. thermal [Bou24, NZF11]. thermo [DN19]. thermo-elasticity [DN19]. theta [CCC25, JWL20, Kie23, RS21]. thick [HLJ25]. thick-restart [HLJ25]. Thickness [TE03, WWD^{+}12]. thickness-shear [WWD^{+}12]. Thiele [Gen12, LB93, TF00]. Thien [PP05]. Thin [HM07, ZKD02, ZKD04, BKPS93, CM24b, Han93, Lev95, Pow93, YXL18]. thin-plate [CM24b, Han93, Lev95]. Third [CMRS01, ZQS24, AC17, BD10, DD21, DDRS23, GB21, Hai08, Hai09, Hea10, HMS96, HLV26, JL12, Jat15, KSV23, KS12, KM19, MH21, ME95, Meh11, NLT21, PKR20, PG12, RS97, SS26, Tem08, YZL25]. third-kind [MH21, NLT21]. Third-order [ZQS24, DD21, Hai08, Hai09, HMS96, HLV26, KSV23, KM19, Meh11, PKR20, SS26, YZL25]. Thomas [ZZWK12]. Three [AEG02, GMZ19, GGV02, JKK^{+}07, LZM23, MST03, NSM20, XZ25, AGA24, AAM24, ALJLYJ24, AK12, And14b, And15, AKT15, BCST14, BK13, BE20, BBL22b, CMD19, Cal20, CHMT10a, DZ13, Don16, GH06, GH09a, HT21, HVM15, Hua20, IAAH24, JLJ22, KBA26, KM19, Lai92, LK20, LZ21, LSY^{+}23, LX24, LSSS15, LZ22d, MJJ^{+}23, MP00, MS01a, Moo07, PH20, PTSB01, PV98, Rab23, SH12, SKTGR19, SSP15, SW07, TS18b, TÖ17, WG99, WZ19, WLMA21, WYP23, WQZH24, XSL22, YAT20, YJJ^{+}21, YYZ22, ZXLF15, ZS22, ZZY24, lZfLqL26]. three- [Lai92]. three-by-three [ALJLYJ24, Hua20, LZ22d]. three-composite [LX24]. three-derivative [TÖ17]. Three-Dimensional [AEG02, BK13, CMD19, MP00, PTSB01, PV98, SW07, ZZY24]. Three-Direction [MST03]. three-directional [MS01a]. three-field [WG99]. three-parameter [HT21]. Three-pencil [JKK^{+}07]. three-point [LK20, LSSS15, PH20]. three-stage [KM19, SKTGR19]. Three-step [NSM20, BCST14, CHMT10a, YAT20, ZS22]. three-term [AAM24, And14b, And15, BE20, Don16, IAAH24, JLJ22, KBA26, LZ21, LSY^{+}23, MJJ^{+}23, SSP15, WYP23, YJJ^{+}21, YYZ22]. three-way [lZfLqL26]. Threshold [LLY26, BCCP25]. Thresholding [BP03, EKPU23, HLL22]. Thresholds [PR03]. Thumb [Boy05]. Tight [VW08]. Tikhonov [BOR23, BGR25, CRS04, Don12, DNH^{+}25, DR12, FR12, HYT26, HRY16, HRY19, HS12, LP12, LGM25, MRV23, NR14, RP26, YXS22, ZZ19, ZFH26]. Tikhonov-type [DNH^{+}25, ZZ19]. Time [BL98, DG05, DGM26, Söd02, TDKB24, ZS03, ZZ18, ZKD02, ZKD04, AD17, ATM19, AH23, ADG10, AF13, AAN25, AWL^{+}24, And14a, AH18, BPR22, BS17, BM22, BKF20, BZV16, BEH24, BC26, BIMR19, BDF^{+}25, COSE22, CX20, CLQ26, CW25, CEK21, CG13, Cui13, DIM22, DAM16, DZS21, lDzS21, EG18, Fan19, FZLL23, FT24, FZL^{+}16, Fly22, GZL24, hGzS17, GNH10, GCZH24, GLS^{+}18, HLS10, HKE97, HDP18, HV22, HA16, Hol98, HZX21, HLTA16, HZ20, HTVY13, HL17, HYJ20, HL20, HS21, HWZ24, ICR06, IJE15, JWZ23, JJC24, JB24, JZ16, KS18a, KNBGV18, KN23, Kaz24, KR11, KS97, KXD25, KB20, KP25, LLAL21, Lam09, LMUZ19, LSZW19, LLX20, LlLZ21, LZ23a, LAN18, LXZZ21, LWZ23, Lor19, LF25, Luc97, lLXhL22, LV18, LV21, LC25, MN23a, MM23, MS24a, MDL15, MA13, MSS11, MvS09, MAFN16, NPS09]. time [Ngo23, NG23, Not22, PQS22, PSB26, PLZ^{+}24, PM24, QXGZ20, RF23, RS06, REM21, Ria16, RK11, RR22, RS25, RS26, Sal17, San25, SFMK23, SGS25, SGS26, SN24, SS24a, SKD26, SLA11, SSYL20, SSH^{+}19a, SCF23, SFS23, SD26, Si12, SN22, SK25, SJ14, SYLT14, SWJ24, SG23, TTP26, VLCL16, Wan15b, Wan19, WXQ20, WLMA21, WWBM21, WZC23, WZ23a, WQ23, WS24a, WWL24, Wei17, WLY^{+}21, XWC26, XH21, YJ18, YZL20, YPL21, YZH21, YK22, YXS22, YHL25, YY24, YWZ19, YZX25, Zaf22, ZWFY19, ZXRL11, ZLW^{+}13, ZLG^{+}13, ZSF18, ZLL^{+}21a, ZBX21, ZHFW21, ZZ22b, ZL22a, ZF22, ZQS24, ZYW22, ZJ23, ZLS24, ZXL23, aZ19b]. time-delayed [PM24, RK11]. time-dependent [ATM19, COSE22, DZS21, EG18, HV22, Lam09, LSZW19, Lor19, NPS09, SSYL20, SFS23, SG23, TTP26, WS24a, YJ18, YK22, ZSF18]. time-domain [HLS10]. Time-fractional [DGM26, AD17, AH23, AAN25, AWL^{+}24, CLQ26, FZLL23, GZL24, GCZH24, HV22, HZX21, HZ20, HS21, JWZ23, KN23, KR11, KXD25, KP25, LSZW19, LWZ23, LV21, MN23a, Ngo23, PSB26, PLZ^{+}24, QXGZ20, RF23, RR22, RS25, RS26, San25, SN24, SK25, TTP26, WWBM21, YXS22, YY24, YWZ19, ZLW^{+}13, ZL22a, ZF22, ZYW22, ZLS24, ZXL23, aZ19b]. time-harmonic [LAN18, LXZZ21, RS06]. time-independent [Lor19]. time-instant [YZH21]. time-periodic [SJ14, SWJ24]. time-resolved [MvS09]. time-scaling [Fly22]. time-space [HL17, HL20, LF25, LC25, VLCL16, WLMA21, YZL20, YPL21, ZWFY19, ZBX21, ZZ22b]. time-step [KS97]. Time-Stepping [Söd02, ZZ18, IJE15, lLXhL22, ZQS24]. time-steps [LlLZ21]. Time-transformations [BL98]. time-varying [GLS^{+}18, LMUZ19, ZXRL11, ZLG^{+}13, ZLL^{+}21a]. times [Cas17]. timestepping [NT21]. Toda [FYI^{+}12, MN01, WLS25]. Todd [YZLP16, Khe12b]. Toeplitz [AAAA^{+}18, AA16b, BM94, BBL22a, BKW25, BC06, BCV03, BMR19a, BIM^{+}23, BM25, BESC22, BBd95, BG03b, BV95, CT93, CO94, DNR17, EG19, EV22, GL23, HN94, HG93, Han02, HR03a, Hem94, HW18, HL17, IMT02, JL16, JW25, KV00, LMV00, LV01, Lin01, LN10, Men26, MA95, SC03a, Sol23, SSC25, TT06, TA13, Vos00]. Toeplitz-block [CO94, BC06]. Toeplitz-like [BKW25]. Toeplitz-Plus-Hankel [HR03a]. tolerable [Pop15]. Tomographic [MFBB23, AH14]. tomography [ÁCL11, BG11, BvLP16, NAE22, RVF07, Wan15a]. Tool [FEK^{+}23, CKO24]. toolbox [BMR19a, BvLP16, KR23, MBR21, RZR12]. Tools [GHN19, Han99, Han07, HJ18b, Han94, FHH99]. topics [BRZS23]. Topographic [dADdRC04]. Topological [KGD03, BRZ17, FHH05, Le 92]. toroidal [Par16]. Torrey [SKD26, SYLT14]. torus [DL97]. Torvik [Mok16]. Total [LMV00, LV01, MPR22, Zak17, BG11, CC06, CPZ14, DHJJ10, Den14b, DWX17, FAMA20, Hua18, JL26, JHLL15, LHZ^{+}21, LZ22b, LJ22, MZW20, OC24, Sar06, VZ93, WHZ^{+}18, WCH15, XXW17, ZH22, ZW22, ZLWQ09]. totally [ANI15, ANI^{+}17, GMP92]. tour [Gau07]. TPS [SB21]. trace [FDFM23, GSF26, LJX25, Meu09a, WG13]. Tracing [PLVB11, YYD14]. tracking [BHS11, Lie00]. trade [EY10]. trade-off [EY10]. Traffic [FLH04, PV23]. Trajectories [LPV03]. trajectory [Lo97]. transcendent [DN24a]. transcendental [Kal00]. transcription [BCT15, CB16]. transfer [BQ19, GV99, GS25, GCFF95]. transfer-function [GCFF95]. Transfinite [FP20]. Transform [CCG01, AMR15, ADN17, AK26c, BHNS16, BC05a, BBM08, Cam95, CC12, CMM15, CAB22, Cou15a, Cou15b, DW15c, Ehr97, Hem94, HFDSC24, Kuh13, LM14, Luc97, Roh07, Sab92b, WQL20, WF23, XX16, dDS00, Ave20, DCM^{+}13, DCMM13]. Transformation [Bel03, MS06, AL15, AG00, BRZ19, CDT10, CHHL18, CHH^{+}20, DM14, GHM08, HLS10, Hom94, Hom98a, Hom98b, HV12, Li95, Li96, LCHH21, Rha22, Rog95, Sid20a, SS12b, XW17, ZCGS24]. transformational [YD09]. transformations [Ben97, BL98, BS92, CMR16, GHM08, Her96, Osa92a, OOOR12, Pep23, VC10, VPL97, Wen92]. Transformed [IJ19, PRVI20, SS21, Val14, Val15, XQZ24, YW17]. transformers [BS25]. transforms [AAAGAD23, BV93, BD00, BM23, Coo09, DM98b, GS95, HIK17, KXXW21, KPT23, MP08, Nor00, PSS10, PPV09, Sab92a, Sab14, Tas93, WGZ18b]. Transient [CGN03, SW10b, YYLX23]. transistors [NAR05a]. transition [DHS09, HSY23]. translated [AGG17]. translates [Sun94b]. transmission [GX19, WBY25]. transport [AM01, CEZ26, DO25, Guo16, HHST19, LL26b, QXGZ20, Ter22, VSA12]. transportation [NR25a]. Transpose [BRZ98, Cdv98, CZ20, HC24, ZLL^{+}21a]. Transpose-free [BRZ98, Cdv98]. trapezoidal [For21, OdZdRV13, VPL97]. Traub [AG13, SS24b, SS24c, WX24]. Traub-type [AG13]. traveling [CMWP20]. traversing [BBBC20]. Treating [Bre04]. Treatment [BZ94, The97, Che13, FRS21, Ghe18, Mok16, SS08]. treatments [MMW20]. tree [KZ21, ZYW17]. treecode [CYWH25]. Trees [But10, But19, FM04]. Trefftz [Kar13]. trended [LDN16]. tri [LSG15, ZYW17]. tri-coloured [ZYW17]. tri-parametric [LSG15]. triangle [KLA24, MR12, SLW13, SKJ^{+}18, Wal06]. triangle-based [MR12]. triangles [HvD93]. Triangular [DP01, CL96a, DW24, HL17, LKW17, LM17a, LZ19a, LZ19b, LZ18c, MT12, Pié99, SLW13, TT06, XZ25, Zha26, ZWX22, ZWY22, ZH26]. triangulation [GGJL24, HMdÁES08, Mer94]. triangulation-based [GGJL24]. triangulations [DLR12, JKK^{+}07, Rip93, Son93]. Tribute [Bre03c]. trichotomy [DS09c]. Tricomi [ZLW^{+}13]. Tricomi-type [ZLW^{+}13]. Tridiagonal [FKMS01, BOL98, CDW95, FJ25, Hem94, Jia20a, JXX^{+}23, JW25, LLZ94, PL99, SSC25]. tridiagonalization [FJ25]. tridiagonalization-based [FJ25]. Trigonometric [PQ95, CBGVN07, Don13, FKP06, GZ18, Jat15, JJ24, JH26, Khe12a, KH20, NJ13, SG23]. Trigonometrical [MKS18, Ixa19]. Trigonometrically [FYM14, JSF13, Li17, LG19, NBJA17, SS15]. Trimmed [KGD03]. trinomial [DM92]. triplets [JLZZ23, VBG96]. TRIPs [PGSU25]. TRIPs-Py [PGSU25]. Trivariate [BDIR18, CFR06]. tropical [BHH24]. Trummer [Pol10]. Trummer-like [Pol10]. Truncated [Pow93, DDRS23, GMY18, Hag25, JLP20, Jia20b, KC24, KLL10, LQ26, MM08b, MM09a, MRS06, OR17, SH21b, TA26, WC23, YLYZ23, ZJ23]. truncating [dOS07]. Truncation [AAA17, HKPW19, Bün24, Kao20, MW16, Tam10]. truncations [PK21]. Trust [CR03a, HZ95, AA12a, ARY17, BKR18, BE17, CHY19, HZ93, KE16, LZ18a, LM13, lLXhL22, SS06, lZqL25]. trust-region [AA12a, ARY17, KE16, LZ18a, lLXhL22, lZqL25]. Tseng [AG19, OMU^{+}24, OAMA22]. TSERKN [LW14]. TSRK [KV12]. TSVD [JRS09]. TT-LSQR [PS25]. tubes [BQ19]. Tucker [DBV23, HIS26]. Tucker-rank [HIS26]. tumor [Ila20]. tuned [Kub15]. tuning [CSC26, Cou15a, Cou15b, Kim26]. tuple [dDL92]. Turán [Gau14, GS95]. turbulence [HÇK26]. turning [CL93, Lis25, SS94, VT10]. TV [BBB^{+}06, MFBB23, XYZ14, ZH26]. TV-based [XYZ14]. TVD [GLLJ12]. TVD/CBC [GLLJ12]. TVD/CBC-based [GLLJ12]. twined [GN12]. twisted [SS08]. Two [ACO03, AA03, AAB13, BKFMA11, BKR18, BKA19, BD02, BS21, BCN^{+}16, BE20, BV09, BM23, CW19a, CAB22, CQLY15, CLWH20, CZM21, DFJP10, DZ01, DDRT97, FAMA20, FW13, hGzS17, GGV02, Góm01, GL21, HR03b, Hu22, HTVY13, IAAH24, JJK97, LR18, LRL19, LZW20, Liu11, MsC20, Mar92, Mil18, NAHZ21, ODL21, OR24, RT24, RS26, ST17, SCW17, SH23, TE03, TQY21, TS15, TBA94, TCW14, WWBM21, WCZ25, WDY04, WC10, XZZ19b, Yan18, Yue25, Zha09, Zha11, ZW14, ZDZ26, ZH25, ZLT25, ZYW22, AD17, ASS13, AABM17, AC19, AAD14, Bac18, Bac20, Bai97a, Bai97b, BCW13, BZ13, BH22, BJ98, BZV16, BFK22, Bic24, BZS22, Bra06, CB16, CCTV16, Cas17, CRZ26, CT10, CLTA10, CM16, CY19, CW19b, CS24, CZZ^{+}25, CV22, CLBT15, Cro92, Cui13, DJ10, zDYG18, DW21, DW22]. two [DHWL25, DHS09, DZ19, DMW23, Dzu13, EAB20, EGGSH13, FYM14, FHL21, FZLL23, FLY25, FPP05, FR18, GKRS22, Gau10, GK20, GCGVH92, GL15, GLS^{+}18, GZD24, GZL^{+}25, GS26, HHST19, HZPW23, HCL21, He16, Hem96, HST15, HLTA16, HFZ19, HL20, HG25, JT96, JSZ22, KXD25, KPA20, LS25a, LWD23, LWAG08, LL14, LDW18, LD21, LZ22a, LZ16a, LZ19a, LA22a, LL26b, LZL22, LS15b, LLL22, LRL22, LYL15, LSW16, LW17, LD20, LS20, LW22, MS20a, MP25, MS22, MT12, Met19, Mic23, MP14, MDH16, MDL15, MKS18, MSM12, MS24b, NG23, NV21, Nov26, PR14, PCDH20, Pis16, PM24, QXGZ20, Rab23, RKMS16, RR20, RW24, RGJ10, RhG15, RWTW19, RSCH^{+}19, Ros97, RB17, RR22, SW10a, Sch17, SB21, SW10b, SKD26, SAHM25, SMM25, SSYL20, SL21b, SS15, Sho18, SKTGR19, SSK23, SK25, ST99]. two [SJW21, SSA25, SH26, Tak17, TZ25, TS18b, TKS24, VZ93, WGZ18a, WZ19, WSZ21, WD23, WL25, WCY25, WLY^{+}21, WCM94, Wu22, XL24, YZ21, YZ23, YJX15, ZR17, Zah09, ZBDK23, ZZH15, ZD15, ZHFW21, ZL22b, ZZ22b, ZV21, ZJSS25, ZWY22, ZD18, vdMRS06]. two-asset [ZZ22b]. two-by-two [BCW13, BH22, CW19b, LZ16a, LZ19a, LA22a, WGZ18a]. Two-derivative [CQLY15, TCW14, CT10, FYM14]. two-dimension [SB21]. Two-Dimensional [BD02, HR03b, CAB22, DDRT97, JJK97, Mar92, AD17, AAD14, Bra06, CLTA10, CY19, Cui13, DW21, DHWL25, GK20, HCL21, He16, HL20, HG25, KXD25, KPA20, LWD23, LRL19, LZW20, LSW16, LD20, LW22, MP25, Mic23, MS24b, NV21, Pis16, QXGZ20, RW24, RhG15, RR22, SKD26, SK25, SSA25, YJX15, ZR17, ZBDK23, ZHFW21]. Two-grid [CLWH20, Hu22, NAHZ21, ZDZ26, ZYW22, CZZ^{+}25, CV22, DZ19, DMW23, GS26, HFZ19, LD21, LYL15, LW17, LS20, MDH16, QXGZ20, SW10b, SSYL20, SL21b, WD23, YZ21, YZ23, ZD18]. Two-Hydrodynamic [TE03]. two-Lagrange [GL15]. Two-Level [FAMA20, ZLT25, DW22, MT12, NG23, SMM25, ST99, XL24, vdMRS06]. two-mesh [FZLL23, WLY^{+}21]. two-metric [CS24]. Two-parameter [BKA19, WC10, Dzu13, ZWY22]. two-parametric [CLBT15]. Two-Phase [AA03, BZS22, HHST19, Met19]. Two-Point [ACO03, DZ01, BM23, TBA94, AABM17, Bac18, Bac20, Bic24, CAB22, Cro92, EAB20, GCGVH92, LS15b, MS20a, PR14, RB17, WCM94]. two-sided [BZV16, HZPW23, HLTA16, RSCH^{+}19, Sch17, VZ93]. two-stage [Bai97a, Bai97b, BZ13, CM16, JSZ22, ZJSS25]. Two-step [BV09, CW19a, CZM21, DFJP10, MsC20, ST17, TS15, WCZ25, Zha11, BJ98, BFK22, DJ10, FHL21, FLY25, FR18, GZD24, GZL^{+}25, JT96, LDW18, LZ22a, LL26b, LLL22, RKMS16, RR20, RWTW19, SW10a, SAHM25, SS15, Sho18, SKTGR19, SSK23, TZ25, ZV21]. two-subspace [Wu22]. two-sweep [LZL22, WL25]. two-term [ZD15]. Type [ACO03, MC05b, MC05c, SFT03, Van02, AWH^{+}24, AAPR21, ALV20, AH09, AH10, AH11b, AH12, AG13, AM16, AH18, BL92, BPR20, BRS91, BRS92, BRZ98, Bre99a, BS14, BGVHN96a, BCJ99, CZL26, CCTV23, Ceg25, Cdv98, CRN19, CPRZV23, CGL99, CG19, CJ20, CKKT16, CD07, CK06, CHK14, DL97, Die08, DPS18, DNH^{+}25, DHM12, DGP15, EED19, FHH96, FR18, FLR01, GGNF17, GM22, Gen12, Ghe13, Ghe16, GLC22, GEA20, HHLM23, HVM15, HVYMS23, HPS97, HMS96, Hom92, HLZ14, HL15b, HCD26, HSL19, HLSLSL25, ITA24, IDS16, IS17, IUM^{+}19, JSF13, JS15, JZYY23, JLFL19, JJC24, JLWZ25, Kar10, KAF18b, KK22a, KGMH21, KSJ26, KGN^{+}24, LGL23, LW13, LHM20, LZM23, lLLVZ17, LTFL10, LQ16, LWS18, LS20, LZ23b, LZ23c, LCW25, LMMD05, LV18, MsC20, MP08, MTTC22, MJ20]. type [Mat96, Maz09a, MP14, MP22, Moh10, MK17, NT21, NS22, OPSM22, OAR22, Ovi22, Pas92, PC13, Pej14, PS06, PP92, PK22a, PK22b, Pop18a, Pop19, RTCL21, Ron92b, Sab14, SSR25, San19, SSS21, SH12, SS12a, SKP20, SRK^{+}25, SLD20, SW19, SS24b, SS24c, Sol15, SMN24, SG10, SSSS22, Spa24, SH22, SH23, TAM21, Tem97, TLD^{+}23, TKS24, UPS25, Van92, VV11, VdR13, VS19, WZ13b, WG13, WK16, Wan19, WXT22, WPY25, XL24, XH20, YM24, YBK^{+}21, YJS25, Ye22, ZAH26, ZLW^{+}13, ZFC18, ZZ19, ZLH21, ZL22b, ZQzS22, ZL23, ZZX26, ZH17, ZWW21, ZG09, dBD05, dC23]. typed [WQ10]. types [LAH22].

UGS [BRY14]. Ulam [CL24b]. Ulm [Arg09, LS25b, SHLY18, WCLW16]. Ulm-like [LS25b, SHLY18]. Ulrich [AL04]. ultra [Bac23, Bac24, Bac25]. ultra-weak [Bac23, Bac24, Bac25]. ultrasound [WHZ^{+}18]. ultraspherical [BD20]. umbrella [HMdÁES08]. unbalanced [GWL20]. unbiased [ZL24, ZL26b]. unbounded [CC07, CK22, NNCN23, PM26, Pop15, QL12, YWZ19]. uncertain [YYLX23]. Uncertainty [Cha04, MHZ05, KBCG13, WLL12]. Unconditional [HWZ24, LW22, SZ23, SLT20, XF25, YJ21]. Unconditionally [CLY25, JLWZ25, LCZZ23, LCW25, PLZ^{+}24, CLPY23, He16, HP18b, LLY26]. unconfined [SMNZ20]. unconstrained [AA12a, AAN14, And06, And08, And10, And14b, And15, And18a, And19b, And22, Cui26, DW15a, DW15b, DBH21, Don16, FH25, HWC15, Kim26, Kri26, LLL18, LM13, Liu14, LL18, gOM14, SS06, SLL23, VL19, WJW14, WZ16, WYP23, YCL17, YZLC24, YY13, ZLZ23, ZWG18, ZFZ19, ZLL21b, ZDSY20]. uncountable [Man10]. uncoupled [YJ18]. Uncoupling [ASZ23]. underdetermined [CMRS00b, Kub15, Rum14]. Undirected [vGK04, BKM25]. unequal [LlLZ21]. unequally [DLR12]. Unfitted [ZZZ20]. unfixed [CZH22]. Unified [AG13, BM22, Che13, MFK^{+}15, MGG25, BMS24, Iva17, OL23, WZ11, ZL26a, ZE10]. Uniform [AA09, Bog13, Dun94, LZ23c, MM22, MN11, Rei97, The12, TV19, AHR21, ADA07, Ave20, BS17, BL92, BR07b, CJ17, DLR12, FHC21, GH23a, GO06, GL21, JJC24, LX23, Lin05, MW24, MW26, MS23a, OQ12, RS25, RS26, Tem97, TTP26, VC92, VT10]. Uniformly [CZ23, CL93, CG19, CEX14, CJ20, IUM^{+}19, KK17, LZ23a, MT12, SI18, SC18, dS00]. unifying [dFG93]. unilateral [CLGS17]. uniparametric [LK20]. Unique [DC17, VSA12, CNR15, WMCW21]. uniqueness [WCW20, Zaf22]. Unit [Mar04b, ACH14, BCM07, BW15, BPR20, BGVHN92b, BGVHN92d, BGVHN96a, BC09, BSL18, BD25, CGM12, CMR16, GHM08, Gla01, HM07, LPP21, MA95, SX96, VM17]. unitary [Uhl22a, Uhl22b, VC10]. unity [BDS00, DZ19, DZ22, MS23c, dS00, dDS00]. univalent [Røn92a]. Univariate [Rev03, ACG20, LL14]. universal [Yak95]. univexity [MSP25]. unknown [CZ20, GD15b, MN23b, Roh07, YXS22]. unknown-transpose [CZ20]. unknowns [SW07, YWWR12]. unmanned [DIM22]. unmixing [EJR26, RWTM21]. unorganized [DF94]. unresolved [MSCB93]. Unstable [BG03b, SR22]. unsteady [DMW23, KPS22, KSJ26, LBF24, MK17, Mot14, RT19]. unstructured [CL00, CZ94, CZ95, CZ96, CYWH25, DAN26, PR10, ZWX22]. unsupervised [YDW26]. unsymmetric [DB98]. Untypical [Pas11]. unusual [BRZ20]. update [Aih17, Ari98, BBL23, LZ22c, LL07, LTP18, VL19]. updates [BKFMA11, CMM17, KKM20]. Updating [ACSD16, MMGH17, And19b, BK16a, FVV25, LLZ18, LEK21]. updating/downdating [LLZ18]. UPML [dOS07]. Upper [AAAS03, GGJL24, LZX22, SS98, AH13, LM17a, MT19, MT23, PV22b, YKY15, Zha26]. upwind [ASGGRG23, GLLJ12, Gar20, RZZ25, Son93]. Usable [Enr02]. Usage [HSTW14]. Use [Che04, DG05, KPFG04, MCG^{+}04, Ash19, Cor91, DR07, KFK^{+}24, Moh10, Mot14, WO00]. used [JB22]. Useful [SS14, Vep08]. user [Ixa19]. user-friendly [Ixa19]. uses [EGSHVN15]. Using [BD03, BD04a, BBZ95, BWB24, DBH21, EGSV04, FM04, GST03, HL02, JC04, Jia06, KCHD16, KGD03, KL04, LAG05, Mic91, MRU91, PG05, WH04, YWYN22, Zil01, BK18a, ASW06, AS08, AAN25, ASS13, ABT07, ALRT16, AAN14, ADN17, Arg09, AG17, AT17, Ave20, BMA16, BKPS93, BBO21, BC05a, BH11, BE17, BD17b, BBB22, BvLP16, BG13, Bro05, BGL07, BGI25, CSI18, CLMM05, CM24b, CFM15, CE94, CAB22, CGW25, CR23, CM96, CT21, Cou15a, Cou15b, DMT13, DS09b, DC17, DB06, DLYH17, DSS14, EAB20, Ell93, Esp05, FW25, FT24, FHH05, FT05b, FGS24, GD15b, GG98, Gha16, GNH10, GLV05, GLRSG08, GMT92, GL20, HLS10, HYW20, HZX21, HL23b, HFZ19, Iid24, JCL16, JK19, KL25, KK00, KSCS07, Khe12b, KMA13, Kub23, LM11, Laz99]. using [LW20, LL26a, jLyLqW17, LKKM15, MMV17, ME95, MB06, Meu09a, MBM25, MS96, NPS09, NLT21, NJ13, OOOR12, OO22, PSS10, Pan20, PZL15, PS22, QLZX11, Rab23, RR20, RT19, Reb97, Rim25, Rip93, RW11, SL15b, SL15a, SL18, SW14, SD25, SWB08, SMR26, Sla06, SS08, Str09, Sun94a, TCOA19, TBY13, TCW14, WGZ18b, Wat92, Wat94, XX16, YZX25, ZJR25, ZZL17, Zha20, ZLL^{+}21a, ZCGS24, ZAC26, ZH26, ZS13]. utility [Van17]. UTV [FH97, FHH99, FH05]. Uvarov [HV12]. Uzawa [CC15, DYW16, EMS25, HNY^{+}18, Hua20, MXL25, YDWL15, ZZ24, ZSF18].

Validated [IW04, dS24, BL95]. Validating [Vig04]. validation [BL93, BR21, SWB08]. validity [CL96b]. Vallée [The12, TV17]. valuation [SMR26]. Value [AKW02, AKKW03, CN01, GWY26, Mar04a, MT04, Neh04, ALQ17, AJ13, ATM19, AWH^{+}24, AR13, AHC05, AE09, ABI20, AHKW05, AKPW05, Bac18, Bac20, Bac21, Bac23, Bac24, Bac25, BCCP25, BAV18, BD17a, BVV14, BAEF25, BFK11, Bic24, BEG25, BCI14, CEX14, CW14, CHYH24, CFRV23, Cor24, Cro92, DW24, DD20, DD21, DNV25, DK25, EH97b, EM07, FLV14, Flo16, FP20, Van12, GK21, HJB18, HM26, HJ21, JL12, JNS19, KS20, KS12, KK12, KKD25, Li17, LLD23, LJ11, LYL25, LS15b, LSW16, LWS18, LM21, LRM16, LM25, MS20a, ML20, MsC20, MO19, MN08, ML10, NBJA17, NJ13, NA26, OQ12, RKMS16, RSKB17, RR20, RGJ10, RR23, RB17, SHLY18, SW19, SCF23, SS15, SA23, Stu97, TS18a, UTO24, WZQ17, WLG26, WCM94, XLG22, YKY15, YSLH19, YZL20, YH21, YLY12]. value [Zah09, ZE12, ZA20, ZW12a, ZJJW24, ZFX14, ZXLF15, ZZB20, vSv94, vdHM98]. Valued [KM04, Mat04, AM21, AH08, ANA14, Den14a, DLYH17, GMT92, LA22a, Low05, Rob92, SZ24, TRCL24, UPS25, ZFC18]. Values [FH04, AM98a, BG91b, DGM26, EL01, GQ09, LX17, Lóc18, MT19, NIN12, SL18, SHLY18, TM20, XL14, YD26, lZfLqL26]. Vandermonde [KN21a, GM92b, KV00, KRS19, KN21b, Tas93, Tom92, YHZ20]. vanished [SKTGR19]. vanishing [Jos22, ML22, NS24]. Varga [BC22, Ano00d]. Variable [AP26, ABS19, BC01a, Che02, CO19, Gau09d, Hua20, JT96, MXL25, Ngu16, WHL24, AK12, BK13, BK18b, BL98, BEL23, BCJ99, Car92, CLT^{+}13, CS24, CLQ26, GZL^{+}25, HP18a, HFDSC24, JR20, JZ25, JZF^{+}20, JWZ23, KS97, KJ15, KS06, KP25, LDH23, Lie00, LWJ21, Lis25, LM17b, LL22c, LX24, LZQ26, Lyn08, MJ20, Meh11, Moh10, MK17, MvS09, Mül00, MAFN16, NM26, NRV23, NB16, PQS22, RT19, SS99, Sat24a, Sat24b, SAE19, SHD25, SXHZ20, TPLB22, VH10, Wan15b, WZZ16, Wan17, WSLL24, WZZQ24, YYK24, YK22, ZW14]. variable-coefficient [HFDSC24, KS06]. Variable-Order [Che02, WHL24, BCJ99, CLT^{+}13, JZF^{+}20, LWJ21, NRV23, PQS22, SAE19, SHD25, WSLL24]. variable-phase [Lyn08]. Variable-precision [Gau09d]. variable-step [JZ25]. variables [APPR14, AAH24, BMA16, DFP^{+}10, DZ13, DGP15, FPP05, Gal18, LWAG08, LZ18a, WZ23b]. Variance [ZJSS25, MG22]. Variance-based [ZJSS25]. variance-reduced [MG22]. Variant [AZMJ04, BEQOR14, CCTV16, DFR26, ED13, HWXC17, KL06, KPS14, LZL22, ZZZ22]. Variants [Aih17, MR02, DDRS23, EGGSH13, KKB16, LRL22, MS14a, RWB09, SS12a]. variate [PW14]. variation [BG11, CC06, CPZ14, DHJJ10, FAMA20, JL26, JHLL15, LHZ^{+}21, LZ22b, Sar06, WHZ^{+}18, Zak17]. variation-based [JHLL15, LHZ^{+}21]. variation-regularized [BG11]. Variational [CGR12, Alt21, Anh19, AV19, ATT21, ATT22, AH10, BM22, Bno21, BHT16, CSI16, CSI17, CSI18, DOR26, Dey23, DMT22, DLYH17, DJG18, DHV22, Fan15, GNH10, GEA20, HT19, HRAH22, JM18b, KS25, KAF18a, KPA20, LRL22, MD21a, OIM21, OAMA22, OR24, PPPN23, PN21, RFS23, RC14, RIAA19, RTCL21, RTD^{+}21, RT22, SR16, SI17, SI18, SLD20, SK19, SDMMK18, SMA99, SC18, TAM21, TQY21, TQC22, TCO26, TG20, TH18a, TH18b, TH19a, TH19b, TVC20, TTLD20, TSI20, TDC21, TRSI23, TLD^{+}23, TRCL24, TTA26, UA09, VS19, WZZ16, WXT22, WL22a, WZC26, WCH15, WZZ15, XCD23, XWL25, XH25, YL19, YLL20, YJS25, YAY26, Ye22, ZSGJ25, ZFC18, ZFH23, ZWX19, ZJSS25, ZH26]. variations [Gha16]. varieties [CHS25, Sau07]. various [ARS26, BGR23b, GKL21, Le 98]. varying [AG00, BLS92, FMD18, GLRSG08, GLS^{+}18, JZ16, LMUZ19, ZXRL11, ZLG^{+}13, ZLL^{+}21a]. Veamy [OBAHK^{+}19]. Vector [Che01, GM03, JRS09, MRS25, Rob02, SGM02, Van02, Van03, Ber14, BR07a, BS92, BRZS23, Cas17, CCS05, CGM93, DSS14, Han93, Mar93, Mat91, Mat92, Osa92a, REM21, Rob92, Rob98, Sal96, Sal25, SV26, Ter23, VB92, WX24, YZBJ21, ZZX^{+}23]. Vector-Orthogonality [Van02]. Vector-Padé [GM03]. vector-valued [Rob92]. vectorial [CRHTTN24]. vectorized [Sid17]. vectors [Ber14, BP19, GM96, Meu12, PW16, PWCsL18]. vehicles [DIM22]. velocity [LN22]. Vem [Das26]. VEMcomp [FMS25]. Venice’s [CDF99]. Verification [MO04, NW04, LL14, OOO11, RW11]. verifications [MdR13]. Verified [LS15a, LWG18, RG10, MC08, Rum14]. verifying [KKO17]. versal [Sto93]. Version [DL04, Han99, AG23a, BD00, BX19, Han07, JR10, MH23, MBJ17, Rha22, Sid17, Śmi13, Van25, ZCTD24, DCM^{+}13, DCMM13]. versions [BKDA25, MJ18]. versus [ART14, DPR23, FS01, Maz18, Sei98]. vertical [CCV07, GZL^{+}25, Mez22, SG17, ZZLV23]. via [AWH^{+}24, APST21, BCL00, BAEF25, BEQOR14, BG91b, Cau22, CHHL18, CMP22, DHJJ10, DMR20, DMR21, DLC14, El 18, EN11, FVV25, FY19, GGN18, GV99, GKL21, GK20, GSV96, HVMMV24, HIK17, HLV26, Hua21, KJ18, KS18a, KP09, LM01, LEK21, Lig93, LCHH21, LR14, LCW21, MT18, Mic23, Nar05b, OB16, PPS25, Ria16, Sid20a, SGK^{+}99, TA24, THF21, TO21, The12, VBG96, VRM23, WQL20, XQZ24, XLL26, ZM94]. vibration [LCW23, YZX25]. vibrations [QLZX11, WWD^{+}12]. Vibro [Dum03]. Vibro-Impact [Dum03]. VIEs [AHP20, ZH25]. View [Cor02, ZCC26]. Vignes [AL04]. violated [PR93]. Virtual [FMS25, LCGH23, CLQ26, Das26, DN19, GCZH24, HCBAEC23, KP22, LBF24, LN22, OBAHK^{+}19, RW06, SGO22, Sut17, WZ22a, ZF22]. viscoelastic [KBP17, WSY12, ZD21, ZGK^{+}25]. Viscosity [CSI17, CSI18, KGN^{+}24, MN17, NM14, TH19b, THS20, VA20]. viscous [TCOA19]. visual [GKL21]. visualization [JJK97]. VLSI [Gro93]. VMS [YJ18]. Voigt [KBP17, Pan18, ZD21]. volatility [AAA25]. Volterra [AAH20, CQ16, CIP10, CS25, CCHH23, CZZ^{+}25, CDI14, CPS12, DEC24, DSI11, DHWL25, EED19, FHS12, FH15, GH09b, Gu20, HST15, HCXL20, JH26, KMH24, LDH23, LLY25, LM21, ML20, MH21, MH23, Maj14, MV17, MNS23, MMLM20, NS24, NLT21, OM26, ÖRBB14, RT12, RM26, SMN24, TZH24, WQL20, WHD22, WHL24, WWL24, WLJ24, WZL26, Xu19, ZA20, Zeg25a, Zeg25b, ZHW25, ZH23]. Volume [ABMV03, Ano04a, Ano04c, Ple03, CV15, DAN26, DJS20, FZLL23, LZZ24, PNW17, SW22, SKJ^{+}18, Val14, Val15]. Vorobyev [Str09]. Voronoi [KK23, SSH19b]. Voronoi-splines [KK23]. Voronovskaja [Tac12]. Vorst [GMS99]. VPAStab [GM03]. Vries [TA24, YZZL17, YZLZ22]. vs [ABL… 12, Cve06, SKSS21].

W [GPHHAPR18]. walk [Dam08, Mil20]. walks [Fas25]. wall [MFK^{+}15]. walls [MSM12]. Walter [Gau11b]. Wang [MA15]. Water [vdHS02, ASZ23]. watermarking [ZLS25]. Wave [CK22, AD17, AF13, BZV16, BCJ22, CM24a, CY19, CGM25, CMWP20, DWZ14, DEC24, DS09b, DAM16, DW21, DW22, lDzS21, hGzS17, HDP18, HTVY13, HYJ20, KB20, KP25, LLX20, LZ23a, LP20, LW22, LZZ24, MS23a, MR12, Nat07, Sal17, SHF15, SCF23, SzS21, Tom11, Wan15b, WW19, WSZ21, YZL20, Yua21, ZR17, ZGK^{+}25]. Wave-heat [CK22, CGM25]. Waveform [JM18a, LJWW21, PT19, vLV02, BX17, CGM25, CDP16, CG13, Fan19, OM18, SJ14, SJW21, TPLB22, ZJ14]. Wavefront [ZS03]. Wavelet [JM00, KZS21, Kun01, ADN17, BC05a, BK08, CK05, Ehr97, FT05b, ITA24, Kun05, MP13, NRV23, PQ95, ROB18, SKM25, SA23, TK94, VH92, YXL18]. wavelet-based [BK08, SA23]. wavelet-homotopy [YXL18]. Wavelets [DDG05, FT02, AAU26, Mic91, MRU91, NLT21, PSS10, ROB17, REM21, Spr98, TV25, Tur94, VM17]. wavenumber [GM20]. wavenumber-explicit [GM20]. waves [CM99, CCJC18]. waves-bordering [CM99]. way [Lev05, lZfLqL26]. Ways [MKO04]. Weak [SYZ22, TH18b, TSI20, WYZ21, AUA22, ABB15a, ABB15b, AABM17, Bac23, Bac24, Bac25, BGZ20, GGNF17, Gha24, HH11, KLR07, LZ23b, LZCW24, MWWY13, RNT26, SK25, TX19, TKS24, UA09, USM24, VH10, WK14, WK16, WG25, WZ26, YWS20]. weak-Chebyshev [GGNF17]. Weakly [LL16, AAA24, Bai97a, Bai97b, CS25, Gar20, GKV23, KM24b, KP25, LDH23, LWN13, ML20, MH23, Maj14, MMLM20, NS24, RS25, SP21, San25, TZH24, WQL20, WHD22, WLJ24, ZXF14, ZH25]. Wei [Deh20]. Weight [dABR01, DPS18, DMS09, Gau09a, Gau10, Gau14, Gau17b, Gau22, GM22, KW96, Leo07a, MS14a, OPSM22, PW14, Spa20, TD09, Ver99]. Weighted [Car10, Ess98, HKKN12, LRY26, RB21, WC23, ZTW19, BL95, CBGVPP09, DL18, DMR21, Gaj05, God15, GP14, HW00, KV07, Kva14, LEK21, LQ26, LWJ21, ME95, OMW21, OAR22, RP26, SGS13, SSK23, SH26, YXS22, ZLZ23, ZLM25, ZM26]. weighted- [ZLM25, ZM26]. weighted-Newton [SGS13, SSK23]. weights [Ber93, BLS92, BEGG91, LMV24, Not08, Pej14]. Well [AC94b, DHS97, Ant22, BDN17, CDG23, CZ20, Kub15, Mok16, WZQ17]. Well-conditioned [DHS97, Ant22]. Well-defined [AC94b]. well-determined [Kub15]. well-posed [Mok16, WZQ17]. well-posedness [CDG23]. well-separated [BDN17]. Welsch [MS14a, Van25]. WENO [ALRÁY25, GB21]. WG [KSJ26, XZZ22]. WG-FEM [KSJ26]. Where [Lev05]. Which [LL05, Ila20, JB22, PR10]. white [MKBY19, SPS26, WC24]. Whitham [MV14]. Whitney [Loh22, RRZ21]. Whittaker [GG01]. Whose [Kol04b, Sid20a]. wide [Gor18, SMZMA18]. wider [CKS16]. Wiener [Ste24, TA26, YLYZ23]. Will [Lev05]. window [GGN18]. windows [CG13]. Windschitl [LSM16]. Wire [ZKD02, ZKD04]. wise [CHS19, JKC26]. within [CL24b]. without [DG94, GZL^{+}25, Iid24, JCL16, KH20, NKA24, Pie96, Rim25, TRCL24, WZ16, WK16, WYZ21, WXT22, YYZ22, ZH17]. Wixom [BF18]. Wolfe [MSZ20, NKA24]. Wong [KJ18]. Work [DR01, Bre06a, Gau95, GG08]. Work-Bound [DR01]. Workload [FLH04]. Workshop [Ano93, Ano95c]. wormhole [LRC19, ZYJY22]. worst [BDL^{+}12, GS21, HL06]. worst-case [BDL^{+}12, GS21]. wrapping [AHL20, Bün18]. Wright [TA26]. WWP [YLL22]. Wynn [Bre19, CHH^{+}20, DS23]. WZ [Wim99]. WZ-algorithms [Wim99].

Ye [YZLP16]. yield [Ver24]. Yosida [RIAA19, WZ22b]. Yosida-regularization [WZ22b]. Yuan [Deh20, Zha09]. Yukawa [CG20].

Zaglia [CHHL18]. Zakai [KJ18]. Zakharov [LZ24, MV13]. ZD [GLS^{+}18, ZLG^{+}13]. ZeaD [YZH21, ZQL^{+}19]. Zero [BL04, SFT03, ARY17, CS12, DD99, ER19, Gau08b, HLSLSL25, OL21, ST22, SHLY18, TB19]. zero-finders [ARY17]. zero-one [OL21]. zeroes [Ron08]. zeroing [CZ20, ZHY^{+}20, ZLL^{+}21a]. Zeros [BDJ11, CGM12, CMR16, DJ02, DJ18, GG01, GKS04, INR01, VC00, ADGP15, BRW11, Bin96, DL01, DJM08, EL08, GL07, GG08, Gau09c, Gau09b, Gau11b, Gau18, Iva17, JNW92, KG23, KJG23, LMV25, LM97, LZ22a, LM08, LR14, MVV25, PHI98, PM05, PMM11, PR14, Seg98, Seg08, SH23, THS20, Yak94]. zeta [Joh15, Vep08]. Zhang [AWH^{+}24, JZ16, Ovi22, ZXRL11]. Zhangtype [AWH^{+}24]. Zigzag [TA24]. Zones [KPFG04]. Zonotopes [Mar04c].

References

Afif:2003:NST

[AA03] M. Afif and B. Amaziane. Numerical simulation of two-phase flow through heterogeneous porous media. Numerical Algorithms, 34(2–4):117–125, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

              ipsapp007.kluweronline.com/content/getfile/5058/48/
              12/abstract.htm;   http://ipsapp007.kluweronline.com/
              content/getfile/5058/48/12/fulltext.pdf.

Amiraliyeva:2009:UDM

[AA09] I. G. Amiraliyeva and G. M. Amiraliyev. Uniform difference method for parameterized singularly perturbed delay differential equations. Numerical Algorithms, 52(4):509–521, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

              9265 (electronic). URL http://www.springerlink.com/
              openurl.asp?genre=article&issn=1017-1398&volume=52&
              issue=4&spage=509.

Ahookhosh:2012:ENT

[AA12a] Masoud Ahookhosh and Keyvan Amini. An efficient nonmonotone trust-region method for unconstrained optimization. Numerical Algorithms, 59(4):523–540, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

              URL http://www.springerlink.com/openurl.asp?genre=
              article&issn=1017-1398&volume=59&issue=4&spage=523.

Annaby:2012:CES

[AA12b] Mahmoud H. Annaby and Rashad M. Asharabi. Computing eigenvalues of Sturm–Liouville problems by Hermite interpolations. Numerical Algorithms, 60(3):355–367, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=3&
spage=355.

Alouges:2015:SCS

[AA15] François Alouges and Matthieu Aussal. The sparse cardinal sine decomposition and its application for fast numerical convolution. Numerical Algorithms, 70(2):427–448, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9953-6.

Abel:2016:ABJ

[AA16a] Ulrich Abel and Octavian Agratini. Asymptotic behaviour of Jain operators. Numerical Algorithms, 71(3):553–565, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0009-3.

Akhoundi:2016:MLA

[AA16b] Nasser Akhoundi and Iman Alimirzaei. Modify Levinson algorithm for symmetric positive definite Toeplitz system. Numerical Algorithms, 71(4):907–913, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0029-z.

Asharabi:2017:TEE

[AAA17] R. M. Asharabi and H. S. Al-Abbas. Truncation error estimates for generalized Hermite sampling. Numerical Algorithms, 74(2):481–497, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0159-y.

Abbas:2018:IMS

[AAA^{+}18] M. Abbas, M. AlShahrani, Q. H. Ansari, O. S. Iyiola, and Y. Shehu. Iterative methods for solving proximal split minimization problems. Numerical Algorithms, 78(1):193–215, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Arrai:2024:FSS

[AAA24] Mohamed Arrai, Chafik Allouch, and Abderrahim Aslimani. Fast superconvergent solvers for weakly singular Hammerstein equations. Numerical Algorithms, 97(1):345–372, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01706-7.

Alsenafi:2025:EAS

[AAA25] Abdulaziz Alsenafi, Fares Alazemi, and Javad Alavi. An efficient algorithm to solve the geometric Asian power option price PDE under the stochastic volatility model. Numerical Algorithms, 98(1):287–306, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01794-z.

Ahmad:2018:EPB

[AAAA^{+}18] Fayyaz Ahmad, Eman Salem Al-Aidarous, Dina Abdullah Alrehaili, Sven-Erik Ekström, Isabella Furci, and Stefano Serra-Capizzano. Are the eigenvalues of preconditioned banded symmetric Toeplitz matrices known in almost closed form? Numerical Algorithms, 78(3):867–893, July 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0404-z.pdf.

Annaby:2023:HIT

[AAAGAD23] M. H. Annaby, I. A. Al-Abdi, A. F. Ghaleb, and M. S. Abou-Dina. Hermite interpolation theorems for band-limited functions of the linear canonical transforms with equidistant samples. Numerical Algorithms, 94(3):1281–1308, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01535-8.

Achchab:2003:UBC

[AAAS03] B. Achchab, S. Achchab, O. Axelsson, and A. Souissi. Upper bound of the constant in strengthened C.B.S. inequality for systems of linear partial differential equations. Numerical Algorithms, 32(2–4):185–191, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
45/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/45/4/fulltext.pdf.

Ahookhosh:2013:TDF

[AAB13] Masoud Ahookhosh, Keyvan Amini, and Somayeh Bahrami. Two derivative-free projection approaches for systems of largescale nonlinear monotone equations. Numerical Algorithms, 64(1):21–42, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9653-z.

Alarcon:2008:HOI

[AABM08] Virginia Alarcón, Sergio Amat, Sonia Busquier, and Fernando Manzano. High order iterative schemes for quadratic equations. Numerical Algorithms, 48(4):373–381, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=4&
spage=373.

Amat:2017:LCD

[AABM17] S. Amat, Ioannis K. Argyros, S. Busquier, and Á. Alberto Magreñán. Local convergence and the dynamics of a two-point four parameter Jarratt-like method under weak conditions. Numerical Algorithms, 74(2):371–391, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0152-5.

Aragon-Artacho:2023:PDS

[AABTB23] Francisco J. Aragón-Artacho, Radu I. Bot, and David Torregrosa-Belén. A primal-dual splitting algorithm for composite monotone inclusions with minimal lifting. Numerical Algorithms, 93(1):103–130, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01405-

Assari:2014:MMB

[AAD14] Pouria Assari, Hojatollah Adibi, and Mehdi Dehghan. A meshless method based on the moving least squares (MLS) approximation for the numerical solution of two-dimensional nonlinear integral equations of the second kind on non-rectangular domains. Numerical Algorithms, 67(2):423–455, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9800-1.

Abdullahi:2023:CSD

[AAFL23] Muhammad Abdullahi, Auwal Bala Abubakar, Yuming Feng, and Jinkui Liu. Comment on: “A derivative-free iterative method for nonlinear monotone equations with convex constraints”. Numerical Algorithms, 94(4):1551–1560, December 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01546-5. See [LF19].

Anh:2018:MBP

[AAH18] Pham Ngoc Anh, Tran T. H. Anh, and Nguyen D. Hien. Modified basic projection methods for a class of equilibrium problems. Numerical Algorithms, 79(1):139–152, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ahmadinia:2020:NSI

[AAH20] M. Ahmadinia, H. Afshari A., and M. Heydari. Numerical solution of Itô–Volterra integral equation by least squares method. Numerical Algorithms, 84(2):591–602, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00770-2.

Asharabi:2024:MHG

[AAH24] Rashad M. Asharabi and Felwah H. Al-Haddad. A multidimensional Hermite–Gauss sampling formula for analytic functions of several variables. Numerical Algorithms, 96(1):105– 134, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01641-7.

Amitai:1996:PAH

[AAII96] Dganit Amitai, Amir Averbuch, Moshe Israeli, and Samuel Itzikowitz. On parallel asynchronous high-order solutions of parabolic PDEs. Numerical Algorithms, 12(1–2):159–192, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amitai:1994:ACA

[AAIT94] Dganit Amitai, Amir Averbuch, Samuel Itzikowitz, and Eli Turkel. Asynchronous and corrected-asynchronous finite difference solutions of PDEs on MIMD multiprocessors. Numerical Algorithms, 6(3–4):275–296, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abdullahi:2024:MTT

[AAM24] Muhammad Abdullahi, Auwal Bala Abubakar, and Kanikar Muangchoo. Modified three-term derivative-free projection method for solving nonlinear monotone equations with application. Numerical Algorithms, 95(3):1459–1474, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01616-8.

Amini:2014:ILS

[AAN14] Keyvan Amini, Masoud Ahookhosh, and Hadi Nosratipour. An inexact line search approach using modified nonmonotone strategy for unconstrained optimization. Numerical Algorithms, 66(1):49–78, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9723-x.

Alazemi:2025:SAU

[AAN25] Fares Alazemi, Abdulaziz Alsenafi, and Alireza Najafi. A spectral approach using fractional Jaiswal functions to solve the mixed time-fractional Black–Scholes European option pricing model with error analysis. Numerical Algorithms, 98(1): 347–371, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01797-w.

Abdi:2024:BRN

[AAP24] Ali Abdi, Martin Arnold, and Helmut Podhaisky. The barycentric rational numerical differentiation formulas for stiff ODEs and DAEs. Numerical Algorithms, 97(1):431–451, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01709-4.

Alahmadi:2021:GLT

[AAPR21] J. Alahmadi, H. Alqahtani, M. S. Pranić, and L. Reichel. Gauss–Laurent-type quadrature rules for the approximation of functionals of a nonsymmetric matrix. Numerical Algorithms, 88(4):1937–1964, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01101-0.

Ahmad:2026:ENA

[AAU26] Masood Ahmad, Muhammad Ahsan, and Zaheer Uddin. An efficient numerical algorithm based on Haar wavelets for multidimensional nonlinear elliptic PDEs. Numerical Algorithms, 101(3):1505–1535, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02053-
5.

Alscher:1998:SML

[AB98]  Christian Alscher and Wolf-Jürgen Beyn.  Simulating

the motion of the leech: a biomechanical application of DAEs. Numerical Algorithms, 19(1–4):1–12, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/16/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

16/4/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Al-Baali:1999:IHA

[AB99] Mehiddin Al-Baali. Improved Hessian approximations for the limited memory BFGS method. Numerical Algorithms, 22(1):99–112, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
21/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/21/2/fulltext.pdf.

Abushama:2006:MNC

[AB06] Abeer Ali Abushama and Bernard Bialecki. Modified nodal cubic spline collocation for biharmonic equations. Numerical Algorithms, 43(4):331–353, December 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=4&spage=331.

Awwal:2023:NSD

[AB23] Aliyu Muhammed Awwal and Thongchai Botmart. A new sufficiently descent algorithm for pseudomonotone nonlinear operator equations and signal reconstruction. Numerical Algorithms, 94(3):1125–1158, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01530-z.

Abouir:2024:EHB

[AB24] Jilali Abouir and Brahim Benouahmane. Extended homogeneous bivariate orthogonal polynomials: symbolic and numerical Gaussian cubature formula. Numerical Algorithms, 97(4): 1535–1561, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01761-8.

Akhtari:2015:AWC

[ABB15a] B. Akhtari, E. Babolian, and A. Foroush Bastani. An adaptive weak continuous Euler–Maruyama method for stochastic delay differential equations. Numerical Algorithms, 69(1):29–57, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9880-6. See erratum [ABB15b].

Akhtari:2015:EAW

[ABB15b] B. Akhtari, E. Babolian, and A. Foroush Bastani. Erratum to: An adaptive weak continuous Euler–Maruyama method for stochastic delay differential equations. Numerical Algorithms, 69(2):471–472, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-015-9966-9.pdf. See

[ABB15a].

Alleon:1997:SAI

[ABG97] Guillaume Alléon, Michele Benzi, and Luc Giraud. Sparse approximate inverse preconditioning for dense linear systems arising in computational electromagnetics. Numerical Algorithms, 16(1):1–15, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
9/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/9/2/fulltext.pdf. Sparse ma-

trices in industry (Lille, 1997).

Amodio:2020:ASH

[ABI20] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. Analysis of spectral Hamiltonian boundary value methods (SHB-VMs) for the numerical solution of ODE problems. Numerical Algorithms, 83(4):1489–1508, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amodio:2022:AHO

[ABI22] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. Arbitrarily high-order energy-conserving methods for Poisson problems. Numerical Algorithms, 91(2):861–894, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01285-z.

Amodio:2023:SCC

[ABI23] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. (Spectral) Chebyshev collocation methods for solving differential equations. Numerical Algorithms, 93(4):1613–1638, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01482-w.

Aminifard:2022:DLE

[ABK22] Zohre Aminifard and Saman Babaie-Kafaki. Dai–Liao extensions of a descent hybrid nonlinear conjugate gradient method with application in signal processing. Numerical Algorithms, 89(3):1369–1387, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01157-y.

Aminifard:2023:NQN

[ABKD23] Zohre Aminifard, Saman Babaie-Kafaki, and Fatemeh Dargahi. Nonmonotone quasi-Newton-based conjugate gradient methods with application to signal processing. Numerical Algorithms, 93(4):1527–1541, August 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01477-7.

Amat:2012:RPE

[ABL… 12] Sergio Amat, Sonia Busquier, María José Legaz, and Fernando Manzano …. Reciprocal polynomial extrapolation vs Richardson extrapolation for singular perturbed boundary problems. Numerical Algorithms, 61(4):631–647, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9555-0/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=4&spage=631-647.

Andretta:2010:PSP

[ABM10] Marina Andretta, Ernesto G. Birgin, and J. M. Martínez. Partial spectral projected gradient method with activeset strategy for linearly constrained optimization. Numerical Algorithms, 53(1):23–52, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=1&spage=23.

Afif:2003:PEF

[ABMV03] M. Afif, A. Bergam, Z. Mghazli, and R. Verfürth. A posteriori estimators for the finite volume discretization of an elliptic problem. Numerical Algorithms, 34(2–4):127–136, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/23/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
23/fulltext.pdf.

Axelsson:2004:CNI

[ABQ04] Owe Axelsson, Zhong-Zhi Bai, and Shou-Xia Qiu. A class of nested iteration schemes for linear systems with a coefficient matrix with a dominant positive definite symmetric part. Numerical Algorithms, 35(2–4):351–372, April 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/9/abstract.htm.

Ahmad:2019:VOS

[ABS19] Saghir Ahmad, J. C. Butcher, and Winston L. Sweatman. Variable order and stepsize in general linear methods. Numerical Algorithms, 81(4):1403–1421, August 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amat:2007:MSU

[ABT07] Sergio Amat, Sonia Busquier, and J. Carlos Trillo. On multiresolution schemes using a stencil selection procedure: applications to ENO schemes. Numerical Algorithms, 44(1):45–68, January 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=1&spage=45.

Anderson:2023:QMA

[ABV23] Thomas G. Anderson, Marc Bonnet, and Shravan Veerapaneni. Quantifying mixing in arbitrary fluid domains: a Padé approximation approach. Numerical Algorithms, 93(1):441– 458, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01423-7.

Allouche:1994:SRM

[AC94a] H. Allouche and A. Cuyt. Singular rules for a multivariate quotient-difference algorithm. Numerical Algorithms, 6(1–2): 137–168, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Allouche:1994:WDD

[AC94b] H. Allouche and A. Cuyt. Well-defined determinant representations for nonnormal multivariate rational interpolants. Numerical Algorithms, 6(1–2):119–135, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anh:2011:PRN

[AC11] Pham Ky Anh and Cao Van Chung. Parallel regularized Newton method for nonlinear ill-posed equations. Numerical Algorithms, 58(3):379–398, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=3&spage=379.

Amat:2017:CLT

[AC17] Sergio Amat and Vicente F. Candela. On the convergence of a local third order shock capturing method for hyperbolic conservation laws. Numerical Algorithms, 74(4):1011–1028, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0182-z.

Artacho:2019:ORL

[AC19] Francisco J. Aragón Artacho and Rubén Campoy. Optimal rates of linear convergence of the averaged alternating modified reflections method for two subspaces. Numerical Algorithms, 82(2):397–421, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anane:1999:FOS

[ACE99] A. Anane, O. Chakrone, and Z. El Allali. First order spectrum for elliptic system and nonresonance problem. Numerical Algorithms, 21(1–4):9–21, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/20/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
20/fulltext.pdf.

Alberty:1999:RAL

[ACF99] Jochen Alberty, Carsten Carstensen, and Stefan A. Funken. Remarks around 50 lines of Matlab: short finite element implementation. Numerical Algorithms, 20(2–3):117–137, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/6/fulltext.pdf.

Amiraslani:2020:DMU

[ACG20] Amirhossein Amiraslani, Robert M. Corless, and Madhusoodan Gunasingam. Differentiation matrices for univariate polynomials. Numerical Algorithms, 83(1):1–31, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Atkinson:2014:EPU

[ACH14] Kendall Atkinson, David Chien, and Olaf Hansen. Evaluating polynomials over the unit disk and the unit ball. Numerical Algorithms, 67(4):691–711, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9817-5.

Atkinson:2017:SMN

[ACH17] Kendall Atkinson, David Chien, and Olaf Hansen. A spectral method for nonlinear elliptic equations. Numerical Algorithms, 74(3):797–819, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0172-1.

Atkinson:2019:SME

[ACH19] Kendall Atkinson, David Chien, and Olaf Hansen. A spectral method for an elliptic equation with a nonlinear Neumann boundary condition. Numerical Algorithms, 81(1):313– 344, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Alvarez:2011:IPC

[ÁCL11] Juan Ruiz Álvarez, Jinru Chen, and Zhilin Li. The IIM in polar coordinates and its application to electro capacitance tomography problems. Numerical Algorithms, 57(3):405–423, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=3&
spage=405.

Anderson:1993:TPS

[ACM93] I. J. Anderson, M. G. Cox, and J. C. Mason. Tensor-product spline interpolation to data on or near a family of lines. Numerical Algorithms, 5(1–4):193–204, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Avot-Chotin:2004:HID

[ACM04] Roselyne Avot-Chotin and Habib Mehrez. Hardware implementation of discrete stochastic arithmetic. Numerical Algorithms, 37(1–4):21–33, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/3/abstract.htm.

Abouir:2003:MTP

[ACO03] J. Abouir, A. Cuyt, and R. Orive. Multivariate two-point Padé-type and two-point Padé approximants. Numerical Algorithms, 33(1–4):11–26, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/2/fulltext.pdf.

Anzt:2016:UIF

[ACSD16] Hartwig Anzt, Edmond Chow, Jens Saak, and Jack Dongarra. Updating incomplete factorization preconditioners for model order reduction. Numerical Algorithms, 73(3):611–630, November 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0110-2.

Arandiga:2000:NMD

[AD00] Francesc Aràndiga and Rosa Donat. Nonlinear multiscale decompositions: the approach of A. Harten. Numerical Algorithms, 23(2–3):175–216, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
25/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/25/3/fulltext.pdf.

Abbaszadeh:2017:IMM

[AD17] Mostafa Abbaszadeh and Mehdi Dehghan. An improved meshless method for solving two-dimensional distributed order timefractional diffusion-wave equation with error estimate. Numerical Algorithms, 75(1):173–211, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abbaszadeh:2022:CMK

[AD22] Mostafa Abbaszadeh and Mehdi Dehghan. A class of moving kriging interpolation-based DQ methods to simulate multi-dimensional space Galilei invariant fractional advectiondiffusion equation. Numerical Algorithms, 90(1):271–299, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01188-5.

Amiraliyev:2007:PUN

[ADA07] G. M. Amiraliyev, Hakki Duru, and I. G. Amiraliyeva. A parameter-uniform numerical method for a Sobolev problem with initial layer. Numerical Algorithms, 44(2):185–203, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=44&
issue=2&spage=185.

Aimi:2010:NIS

[ADG10] Alessandra Aimi, Mauro Diligenti, and Chiara Guardasoni. Numerical integration schemes for space–time hypersingular integrals in energetic Galerkin BEM. Numerical Algorithms, 55(2–3):145–170, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=145.

Area:2015:BZS

[ADGP15] Iván Area, Dimitar K. Dimitrov, Eduardo Godoy, and Vanessa Paschoa. Bounds for the zeros of symmetric Kravchuk polynomials. Numerical Algorithms, 69(3):611–624, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9916-y.

Aimi:2005:PCM

[ADL05] A. Aimi, M. Diligenti, and F. Lunardini. Panel clustering method and restriction matrices for symmetric Galerkin BEM. Numerical Algorithms, 40(4):355–382, December 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=4&
spage=355.

Arandiga:2017:GED

[ADN17] Francesc Aràndiga, Rosa Donat, and José J. Noguera. Generalizing the ENO–DB2 p transform using the inverse wavelet transform. Numerical Algorithms, 74(1):175–198, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0144-5.

Agrawal:2026:ALM

[ADP26] Ayush Agrawal, Raksha Devi, and Dwijendra Narain Pandey. Analysis of LDG method for fractional order Rayleigh–Stokes problem. Numerical Algorithms, 101(1):301–341, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-02006-4.

Amiraliyev:2009:FDS

[AE09] Gabil M. Amiraliyev and Fevzi Erdogan. A finite difference scheme for a class of singularly perturbed initial value problems for delay differential equations. Numerical Algorithms, 52(4):663–675, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=4&spage=663.

Aslimani:2018:NCO

[AE18] N. Aslimani and R. Ellaia. A new chaos optimization algorithm based on symmetrization and levelling approaches for global optimization. Numerical Algorithms, 79(4):1021–1047, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Aurada:2014:HMI

[AEF^{+}14] Markus Aurada, Michael Ebner, Michael Feischl, Samuel Ferraz-Leite, Thomas Führer, Petra Goldenits, Michael Karkulik, Markus Mayr, and Dirk Praetorius. HILBERT — a MAT-LAB implementation of adaptive 2D-BEM. Numerical Algorithms, 67(1):1–32, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-013-9771-2.
pdf.

Altas:2002:HAS

[AEG02] Irfan Altas, Jocelyne Erhel, and Murli M. Gupta. High accuracy solution of three-dimensional biharmonic equations. Numerical Algorithms, 29(1–3):1–19, March 2002. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp008.lwwonline.com/content/
getfile/5058/38/1/abstract.htm;   http://ipsapp008.
lwwonline.com/content/getfile/5058/38/1/fulltext.pdf.

Abdaoui:2020:SBC

[AEH20] Ilias Abdaoui, Lakhdar Elbouyahyaoui, and Mohammed Heyouni. The simpler block CMRH method for linear systems. Numerical Algorithms, 84(4):1265–1293, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00814-7.

Arge:1994:ASD

[AF94] Erlend Arge and Michael Floater. Approximating scattered data with discontinuities. Numerical Algorithms, 8(2–4):149– 166, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ala:2013:MTM

[AF13] Guido Ala and Elisa Francomano. A marching-on in time meshless kernel based solver for full-wave electromagnetic simulation. Numerical Algorithms, 62(4):541–558, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9635-1.

Abdollahi:2023:MCG

[AF23] Fahimeh Abdollahi and Masoud Fatemi. A modified conjugate gradient method for general convex functions. Numerical Algorithms, 92(3):1485–1502, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01349-0.

Alamdari:2026:NNS

[AFG26] Mohammad Saeid Alamdari, Masoud Fatemi, and Aboozar Ghaffari. Non-negative sparse signal recovery with a feasible based sequential quadratic programming. Numerical Algorithms, 102(1):199–219, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02127-
4.

Axelsson:2016:CPK

[AFN16] Owe Axelsson, Shiraz Farouq, and Maya Neytcheva. Comparison of preconditioned Krylov subspace iteration methods for PDE-constrained optimization problems. Numerical Algorithms, 73(3):631–663, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0111-1.

Axelsson:2017:CPK

[AFN17] Owe Axelsson, Shiraz Farouq, and Maya Neytcheva. Comparison of preconditioned Krylov subspace iteration methods for PDE-constrained optimization problems. Numerical Algorithms, 74(1):19–37, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0136-5.

AlvarezdeMorales:2007:DPB

[ÁFP07] María Álvarez de Morales, Lidia Fernández, and Teresa E. Pérez. On differential properties for bivariate orthogonal polynomials. Numerical Algorithms, 45(1–4):153–166, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-007-9113-3.

Apprato:2000:RAS

[AG00] Dominique Apprato and Christian Gout. A result about scale transformation families in approximation: application to surface fitting from rapidly varying data. Numerical Algorithms, 23(2–3):263–279, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
25/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/25/5/fulltext.pdf.

Akhmouch:2003:HOA

[AG03] M. Akhmouch and N. Guessous. High-order analytical nodal method for the multigroup diffusion equations.

Numerical Algorithms, 34(2–4):137–146, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/13/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
13/fulltext.pdf.

Argyros:2013:UMS

[AG13] I. K. Argyros and D. González. Unified majorizing sequences for Traub-type multipoint iterative procedures. Numerical Algorithms, 64(3):549–565, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9678-3.

Achache:2015:CAN

[AG15] Mohamed Achache and Moufida Goutali. Complexity analysis and numerical implementation of a full-Newton step interiorpoint algorithm for LCCO. Numerical Algorithms, 70(2):393– 405, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9955-4.

Argyros:2017:CNL

[AG17] Ioannis K. Argyros and Santhosh George. On the convergence of Newton-like methods using restricted domains. Numerical Algorithms, 75(3):553–567, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Alves:2019:ICI

[AG19] M. Marques Alves and Marina Geremia. Iteration complexity of an inexact Douglas–Rachford method and of a Douglas– Rachford–Tseng’s F-B four-operator splitting method for solving monotone inclusions. Numerical Algorithms, 82(1):263– 295, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Adona:2023:IVS

[AG23a] Vando A. Adona and Max L. N. Gonçalves. An inexact version of the symmetric proximal ADMM for solving separable convex optimization. Numerical Algorithms, 94(1):1–28, September 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01491-9.

Arias:2023:IFD

[AG23b] C. A. Arias and C. Gómez. Inexact free derivative quasi-Newton method for large-scale nonlinear system of equations. Numerical Algorithms, 94(3):1103–1123, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01529-6.

Aceto:2025:SSM

[AG25] Lidia Aceto and Luca Gemignani. A scaling-and-squaring method for computing the inverses of matrix φ-functions. Numerical Algorithms, 100(4):1581–1598, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02049-1.

A:2024:HHM

[AGA24] Kandappan V. A, Vaishnavi Gujjula, and Sivaram Ambikasaran. HODLR3D: hierarchical matrices for N-body problems in three dimensions. Numerical Algorithms, 97(4):1635– 1672, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01765-4.

Astorino:2017:MCM

[AGG17] A. Astorino, M. Gaudioso, and E. Gorgone. A method for convex minimization based on translated first-order approximations. Numerical Algorithms, 76(3):745–760, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Auquiert:2007:CSI

[AGN07] P. Auquiert, O. Gibaru, and E. Nyiri. On the cubic L_{1} spline interpolant to the Heaviside function. Numerical Algorithms, 46(4):321–332, December 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=4&spage=321.

Apprato:2005:MAO

[AGRT05] Dominique Apprato, Christian Gout, Christophe Rabut, and Leonardo Traversoni. Multivariate approximation: An overview. Numerical Algorithms, 39(1–3):1–6, July 2005. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abad:2008:AOG

[AGS08] J. Abad, F. J. Gómez, and J. Sesma. An algorithm to obtain global solutions of the double confluent Heun equation. Numerical Algorithms, 49(1–4):33–51, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=33.

Argyros:2020:EAI

[AGS20] Ioannis K. Argyros, Santhosh George, and Kedarnath Senapati. Extending the applicability of the inexact Newton–HSS method for solving large systems of nonlinear equations. Numerical Algorithms, 83(1):333–353, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Al-Homidan:2003:HMA

[AH03] Suliman Al-Homidan. Hybrid methods for approximating Hankel matrix. Numerical Algorithms, 32(1):57–66, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/44/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
44/4/fulltext.pdf.

Argyros:2008:FDF

[AH08] Ioannis K. Argyros and Saïd Hilout. A Fréchet derivativefree cubically convergent method for set-valued maps. Numerical Algorithms, 48(4):361–371, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=361.

Argyros:2009:CNT

[AH09] Ioannis K. Argyros and Saïd Hilout. On the convergence of Newton-type methods under mild differentiability conditions. Numerical Algorithms, 52(4):701–726, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.

asp?genre=article&issn=1017-1398&volume=52&issue=4&
spage=663.

Argyros:2010:KTC

[AH10] Ioannis K. Argyros and Saïd Hilout. A Kantorovich-type convergence analysis of the Newton–Josephy method for solving variational inequalities. Numerical Algorithms, 55(4):447–466, December 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
55&issue=4&spage=447.

Abdi:2011:EGL

[AH11a] Ali Abdi and Gholamreza Hojjati. An extension of general linear methods. Numerical Algorithms, 57(2):149–167, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=2&
spage=149.

Argyros:2011:EAG

[AH11b] Ioannis K. Argyros and Saïd Hilout. Extending the applicability of the Gauss–Newton method under average Lipschitz-type conditions. Numerical Algorithms, 58(1):23–52, September 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=58&issue=1&
spage=23.

Argyros:2011:SSE

[AH11c] Ioannis K. Argyros and Saïd Hilout. On the solution of systems of equations with constant rank derivatives. Numerical Algorithms, 57(2):235–253, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=2&spage=235.

Argyros:2012:STM

[AH12] Ioannis K. Argyros and Saïd Hilout. Secant-type methods and nondiscrete induction. Numerical Algorithms, 61(3):397–412, November 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=3&spage=397.

Argyros:2013:EUB

[AH13] Ioannis K. Argyros and Saïd Hilout. Estimating upper bounds on the limit points of majorizing sequences for Newton’s method. Numerical Algorithms, 62(1):115–132, January 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9570-1/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=115-132.

Andersen:2014:GRA

[AH14] Martin S. Andersen and Per Christian Hansen. Generalized row-action methods for tomographic imaging. Numerical Algorithms, 67(1):121–144, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9778-8.

Amiri:2015:CBS

[AH15] Sadegh Amiri and S. Mohammad Hosseini. A class of balanced stochastic Runge–Kutta methods for stiff SDE systems. Numerical Algorithms, 69(3):531–552, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9911-3.

Ait-Haddou:2017:BHP

[AH17] Rachid Ait-Haddou. q-blossoming and Hermite–Padé approximants to the q-exponential function. Numerical Algorithms, 76(1):53–66, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Asgari:2018:ENS

[AH18] Zohreh Asgari and S. M. Hosseini. Efficient numerical schemes for the solution of generalized time fractional Burgers type equations. Numerical Algorithms, 77(3):763–792, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abdi:2021:SDB

[AH21] Ali Abdi and Gholamreza Hojjati. Second derivative backward differentiation formulae for ODEs based on barycentric rational interpolants. Numerical Algorithms, 87(4):1577–1591, August 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01020-6.

Ahmed:2023:IGS

[AH23] Hoda F. Ahmed and W. A. Hashem. Improved Gegenbauer spectral tau algorithms for distributed-order time-fractional telegraph models in multi-dimensions. Numerical Algorithms, 93(3):1013–1043, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01452-2.

Al-Hayani:2005:AAS

[AHC05] Waleed Al-Hayani and Luis Casasús. Approximate analytical solution of fourth order boundary value problems. Numerical Algorithms, 40(1):67–78, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=1&spage=67.

Atkinson:2013:SMP

[AHC13] Kendall Atkinson, Olaf Hansen, and David Chien. A spectral method for parabolic differential equations. Numerical Algorithms, 63(2):213–237, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9620-8.

Abdi:2022:GEEa

[AHIJ22] Ali Abdi, Gholamreza Hojjati, Giuseppe Izzo, and Zdzislaw Jackiewicz. Global error estimation for explicit general linear methods. Numerical Algorithms, 89(3):1075–1093, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01146-1.

Abidi:2017:SPE

[AHJ17] O. Abidi, M. Heyouni, and K. Jbilou. On some properties of the extended block and global Arnoldi methods with applications to model reduction. Numerical Algorithms, 75(1):285– 304, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abdi:2022:GEEb

[AHJ22] Ali Abdi, Gholamreza Hojjati, and Zdzislaw Jackiewicz. Global error estimation for explicit second derivative general linear methods. Numerical Algorithms, 90(2):833–850, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01211-9.

Auzinger:2004:MDC

[AHKW04] W. Auzinger, H. Hofstätter, W. Kreuzer, and E. Weinmüller. Modified defect correction algorithms for ODEs. Part I: General theory. Numerical Algorithms, 36(2):135–155, June 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/54/A/3/abstract.htm.

Auzinger:2005:MDC

[AHKW05] W. Auzinger, H. Hofstätter, W. Kreuzer, and E. Weinmüller. Modified defect correction algorithms for ODEs. Part II: Stiff initial value problems. Numerical Algorithms, 40(3):285–303, November 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
40&issue=3&spage=285.

An:2020:EIG

[AHL20] Phan Thanh An, Nam Dũng Hoang, and Nguyen Kieu Linh. An efficient improvement of gift wrapping algorithm for computing the convex hull of a finite set of points in R^{n}. Numerical Algorithms, 85(4):1499–1518, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00873-1.

Ait-Haddou:2021:BHP

[AHM21] Rachid Ait-Haddou and Marie-Laurence Mazure. Blossoming and Hermite–Padé approximation for hypergeometric series. Numerical Algorithms, 88(3):1183–1214, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01071-3.

Abdi:2020:NMB

[AHP20] Ali Abdi, Seyyed Ahmad Hosseini, and Helmut Podhaisky. Numerical methods based on the Floater–Hormann interpolants for stiff VIEs. Numerical Algorithms, 85(3):867–886, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00841-4.

Abdi:2025:LBR

[AHP25] Ali Abdi, Seyyed Ahmad Hosseini, and Helmut Podhaisky. The linear barycentric rational backward differentiation formulae for stiff ODEs on nonuniform grids. Numerical Algorithms, 98 (2):877–902, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01818-8.

Ait-Haddou:2021:AOQ

[AHR21] Rachid Ait-Haddou and Helmut Ruhland. Asymptotically optimal quadrature rules for uniform splines over the real line. Numerical Algorithms, 86(3):1189–1223, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00929-2.

Andreani:2022:CBA

[AHS22a] Roberto Andreani, Gabriel Haeser, and Leonardo D. Secchin. Correction to: On the best achievable quality of limit points of augmented Lagrangian schemes. Numerical Algorithms, 90 (2):879–880, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01241-3. See

[AHS22b].

Andreani:2022:BAQ

[AHS22b] Roberto Andreani, Gabriel Haeser, and Leonardo D. Secchin. On the best achievable quality of limit points of augmented Lagrangian schemes. Numerical Algorithms, 90(2):851–877, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01212-8. See correction [AHS22a].

Ahusborde:2009:LSM

[Ahu09] E. Ahusborde. Legendre spectral methods for the grad- (div) operator with free boundary conditions. Numerical Algorithms, 52(2):151–171, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=2&spage=151.

Argyros:2017:CSM

[AHVR17] I. K. Argyros, M. A. Hernández-Verón, and M. J. Rubio. Convergence of Steffensen’s method for non-differentiable operators. Numerical Algorithms, 75(1):229–244, May 2017. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abubakar:2024:SDL

[AIA^{+}24] A. B. Abubakar, A. H. Ibrahim, M. Abdullahi, M. Aphane, and Jiawei Chen. A sufficient descent LS–PRP-BFGS-like method for solving nonlinear monotone equations with application to image restoration. Numerical Algorithms, 96(4):1423–1464, August 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01673-z.

Aihara:2017:VGU

[Aih17] Kensuke Aihara. Variants of the groupwise update strategy for short-recurrence Krylov subspace methods. Numerical Algorithms, 75(2):397–412, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abbasbandy:2013:DOC

[AJ13] S. Abbasbandy and M. Jalili. Determination of optimal convergence-control parameter value in homotopy analysis method. Numerical Algorithms, 64(4):593–605, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9680-9.

Andreani:2011:PGI

[AJMP11] Roberto Andreani, Joaquim J. Júdice, José Mario Martínez, and Joao Patrício. A projected-gradient interior-point algorithm for complementarity problems. Numerical Algorithms, 57(4):457–485, August 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=4&spage=457.

Axelsson:2000:SSR

[AK00] Owe Axelsson and Igor Kaporin. On the sublinear and superlinear rate of convergence of conjugate gradient methods. Numerical Algorithms, 25(1–4):1–22, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/3/22/abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/5058/3/22/fulltext.

pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Axelsson:2009:EOP

[AK09] O. Axelsson and J. Karátson. Equivalent operator preconditioning for elliptic problems. Numerical Algorithms, 50(3):297– 380, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
50&issue=3&spage=297.

Akbari:2012:CGM

[AK12] Hani Akbari and Asghar Kerayechian. Coiflet–Galerkin method for solving second order BVPs with variable coefficients in three dimensions. Numerical Algorithms, 61(4): 681–698, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9558-x/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=61&issue=4&spage=681-698.

Amini:2015:NLS

[AK15] Keyvan Amini and Ahmad Kamandi. A new line search strategy for finding separating hyperplane in projection-based methods. Numerical Algorithms, 70(3):559–570, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9961-1.

Asai:2016:NSR

[AK16] Y. Asai and P. E. Kloeden. Numerical schemes for random ODEs with affine noise. Numerical Algorithms, 72(1):155– 171, May 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0038-y.

Abubakar:2019:DDL

[AK19] Auwal Bala Abubakar and Poom Kumam. A descent Dai–Liao conjugate gradient method for nonlinear equations. Numerical Algorithms, 81(1):197–210, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ahmad:2026:PCN

[AK26a] Sk. Safique Ahmad and Pinki Khatun. Partial condition numbers for double saddle point problems. Numerical Algorithms, 102(2):923–952, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02161-2.

Anupindi:2026:CRD

[AK26b] Vishnupriya Anupindi and Peter Kritzer. Column reduced digital nets. Numerical Algorithms, 101(3):1451–1473, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02050-8.

Asharabi:2026:GRS

[AK26c] Rashad M. Asharabi and Mustafa Q. Khirallah. Gaussian regularization sampling for approximating derivatives of holomorphic functions in the linear canonical transform domain. Numerical Algorithms, 101(4):2911–2938, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02118-5.

Amini:2015:DPB

[AKB15] Keyvan Amini, Ahmad Kamandi, and Somayeh Bahrami. A double-projection-based algorithm for large-scale nonlinear systems of monotone equations. Numerical Algorithms, 68(2): 213–228, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9841-0.

Auzinger:2016:ASM

[AKKT16] Winfried Auzinger, Thomas Kassebacher, Othmar Koch, and Mechthild Thalhammer. Adaptive splitting methods for nonlinear Schrödinger equations in the semiclassical regime. Numerical Algorithms, 72(1):1–35, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0032-4.

Auzinger:2003:CCS

[AKKW03] Winfried Auzinger, Günter Kneisl, Othmar Koch, and Ewa Weinmüller. A collocation code for singular boundary value problems in ordinary differential equations. Numerical Algorithms, 33(1–4):27–39, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/3/fulltext.pdf.

Auzinger:2005:NPE

[AKPW05] Winfried Auzinger, Othmar Koch, Dirk Praetorius, and Ewa Weinmüller. New a posteriori error estimates for singular boundary value problems. Numerical Algorithms, 40(1): 79–100, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=40&issue=1&spage=79.

Auzinger:2017:AHO

[AKQ17] Winfried Auzinger, Othmar Koch, and Michael Quell. Adaptive high-order splitting methods for systems of nonlinear evolution equations with periodic boundary conditions. Numerical Algorithms, 75(1):261–283, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-016-
0206-8.pdf.

Auzinger:2015:DBL

[AKT15] Winfried Auzinger, Othmar Koch, and Mechthild Thalhammer. Defect-based local error estimators for high-order splitting methods involving three linear operators. Numerical Algorithms, 70(1):61–91, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9935-8.

Auzinger:2002:ECS

[AKW02] Winfried Auzinger, Othmar Koch, and Ewa Weinmüller. Efficient collocation schemes for singular boundary value problems. Numerical Algorithms, 31(1–4):5–25, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/43/8/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
43/8/fulltext.pdf.

Alouges:1997:MPP

[AL97] François Alouges and Philippe Loreaux. Massively parallel preconditioners for symmetric positive definite linear systems. Numerical Algorithms, 14(4):361–375, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/6/5/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/6/
5/fulltext.pdf.

Alt:2004:HJV

[AL04] R. Alt and J.-L. Lamotte. In honour of Jean Vignes and Ulrich Kulisch. Numerical Algorithms, 37(1–4):1–2, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/39/abstract.htm.

Amiraslani:2009:RQA

[AL09] A. Amiraslani and P. Lancaster. Rayleigh quotient algorithms for nonsymmetric matrix pencils. Numerical Algorithms, 51(1): 5–22, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=

51&issue=1&spage=5. Tributes to Gene H. Golub, Part I.

Abdalkhani:2015:CTC

[AL15] Javad Abdalkhani and David Levin. On the choice of β in the u-transformation for convergence acceleration. Numerical Algorithms, 70(1):205–213, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9942-9.

Amir:2018:QIO

[AL18] Anat Amir and David Levin. Quasi-interpolation and outliers removal. Numerical Algorithms, 78(3):805–825, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Aulisa:2023:ESE

[AL23] Eugenio Aulisa and Jonathon Loftin. Exact subdomain and embedded interface polynomial integration in finite elements with planar cuts. Numerical Algorithms, 94(1):315–350, September 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01502-3.

Abdullah:2018:NAN

[ALB^{+}18] F. A. Abdullah, F. Liu, P. Burrage, K. Burrage, and T. Li. Novel analytical and numerical techniques for fractional temporal SEIR measles model. Numerical Algorithms, 79(1):19–40, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Aliyev:2023:SME

[Ali23] Nicat Aliyev. Subspace method for the estimation of large-scale structured real stability radius. Numerical Algorithms, 92(2): 1289–1310, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01340-9.

Ai-Li:2024:MPD

[ALJLYJ24] Yang Ai-Li, Zhu Jun-Li, and Wu Yu-Jiang. Multi-parameter dimensional split preconditioner for three-by-three block system of linear equations. Numerical Algorithms, 95(2):721–745, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01587-w.

Allasia:2003:SIA

[All03] Giampietro Allasia. Simultaneous interpolation and approximation by a class of multivariate positive operators. Numerical Algorithms, 34(2–4):147–158, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/14/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
14/fulltext.pdf.

Allasia:2008:BNL

[All08a] Giampietro Allasia. Biographic notes on Luigi Gatteschi. Numerical Algorithms, 49(1–4):1–4, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=1.

Allasia:2008:LGL

[All08b] Giampietro Allasia. Luigi Gatteschi — list of publications. Numerical Algorithms, 49(1–4):5–9, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=5.

Allasia:2018:RHS

[All18] Giampietro Allasia. Remarkable Haar spaces of multivariate piecewise polynomials. Numerical Algorithms, 78(2):661–672, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

An:2025:REA

[ALL25] Liyuan An, Kun Liang, and Qilong Liu. Randomized extended average block Kaczmarz method for inconsistent tensor equations under t-product. Numerical Algorithms, 100(3): 1123–1144, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01982-x.

A:2017:NEM

[ALQ17] Dang Quang A, Dang Quang Long, and Ngo Thi Kim Quy. A novel efficient method for nonlinear boundary value problems. Numerical Algorithms, 76(2):427–439, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amat:2025:NLW

[ALRÁY25] Sergio Amat, David Levin, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. Non-linear WENO B-spline based approximation method. Numerical Algorithms, 98(3):1141–1169, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01829-5.

Amat:2016:NMA

[ALRT16] S. Amat, J. Liandrat, J. Ruiz, and J. C. Trillo. On a nonlinear mean and its application to image compression using multiresolution schemes. Numerical Algorithms, 71(4):729–752, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0019-1.

Altuntac:2021:CPP

[Alt21] Erdem Altuntac. Choice of the parameters in a primal– dual algorithm for Bregman iterated variational regularization. Numerical Algorithms, 86(2):729–759, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00909-6.

Alecsa:2020:GTA

[ALV20] Cristian Daniel Alecsa, Szilárd Csaba László, and Adrian Viorel. A gradient-type algorithm with backward inertial steps associated to a nonconvex minimization problem. Numerical Algorithms, 84(2):485–512, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00765-z.

Agrawal:1998:NCC

[ALW98] Vishwani D. Agrawal, David Lee, and Henryk Woźniakowski. Numerical computation of characteristic polynomials of Boolean functions and its applications. Numerical Algorithms, 17(3–4):261–278, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
12/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/12/2/fulltext.pdf.

Amat:2022:CEC

[ALY22] Sergio Amat, David Levin, and Dionisio F. Yáñez. A class of C^{2} quasi-interpolating splines free of Gibbs phenomenon.

Numerical Algorithms, 91(1):51–79, September 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01254-6.

Alzer:2008:GFI

[Alz08] Horst Alzer. Gamma function inequalities. Numerical Algorithms, 49(1–4):53–84, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=53.

An:2020:SGA

[ALZ20] Jing An, Huiyuan Li, and Zhimin Zhang. Spectral-Galerkin approximation and optimal error estimate for biharmonic eigenvalue problems in circular/spherical/elliptical domains. Numerical Algorithms, 84(2):427–455, June 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-019-00760-4.

Antoine:2021:DAC

[ALZ21] X. Antoine, E. Lorin, and Y. Zhang. Derivation and analysis of computational methods for fractional Laplacian equations with absorbing layers. Numerical Algorithms, 87(1):409–444, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00972-z.

Amodio:1998:ACC

[AM98a] Pierluigi Amodio and Francesca Mazzia. An algorithm for the computation of consistent initial values for differentialalgebraic equations. Numerical Algorithms, 19(1–4):13–23, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/16/18/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/

5058/16/18/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Arnold:1998:NSI

[AM98b] Martin Arnold and Ander Murua. Non-stiff integrators for differential-algebraic systems of index 2. Numerical Algorithms, 19(1–4):25–41, September 1998. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
17/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/17/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Akesbi:2001:MRM

[AM01] S. Akesbi and E. Maitre. Minimal residual method applied to the transport equation. Numerical Algorithms, 26(3):235–249, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
2/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/2/3/fulltext.pdf.

Al-Mohy:2012:MAB

[AM12] Awad H. Al-Mohy. A more accurate Briggs method for the logarithm. Numerical Algorithms, 59(3):393–402, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/content/
4110609h521kg66m/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=3&
spage=393.

Asadi:2013:PIP

[AM13] S. Asadi and H. Mansouri. Polynomial interior-point algorithm for P_{∗}(κ)(κ) horizontal linear complementarity problems. Numerical Algorithms, 63(2):385–398, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9628-0.

Argyros:2016:SLC

[AM16] Ioannis K. Argyros and Á. Alberto Magreñán. A study on the local convergence and the dynamics of Chebyshev–Halley-type methods free from second derivative. Numerical Algorithms, 71 (1):1–23, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9981-x.

Arias:2018:FCI

[AM18] C. A. Arias and J. M. Martínez. Fast convergence of an inexact interior point method for horizontal complementarity problems. Numerical Algorithms, 79(4):1187–1210, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Aljawi:2021:ESG

[AM21] Salma Aljawi and Marco Marletta. On the eigenvalues of spectral gaps of matrix-valued Schrödinger operators. Numerical Algorithms, 86(2):637–657, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00904-x;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00904-x.pdf.

Adams:2024:DES

[AM24] Mark Adams and Paul Muir. Differential equation software for the computation of error-controlled continuous approximate solutions. Numerical Algorithms, 96(3):1021–1044, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01784-1.

Al-Mohy:2021:CSA

[AMA21] Awad H. Al-Mohy and Bahar Arslan. The complex step approximation to the higher order Fréchet derivatives of a matrix function. Numerical Algorithms, 87(3):1061–1074, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00998-3.

Alonso-Mallo:2006:SRK

[AMCM06] I. Alonso-Mallo, B. Cano, and M. J. Moreta. Stable Runge– Kutta–Nyström methods for dissipative stiff problems. Numerical Algorithms, 42(2):193–203, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=2&spage=193.

Al-Mohy:2010:CSA

[AMH10] Awad H. Al-Mohy and Nicholas J. Higham. The complex step approximation to the Fréchet derivative of a matrix function. Numerical Algorithms, 53(1):133–148, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=1&spage=133.

Arioli:1996:CEA

[AMKV96] M. Arioli, H. Munthe-Kaas, and L. Valdettaro. Componentwise error analysis for FFTs with applications to fast Helmholtz solvers. Numerical Algorithms, 12(1–2):65–88, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Aceto:2011:SMP

[AMM11] Lidia Aceto, Cecilia Magherini, and Marco Marletta. Shooting methods for a PT-symmetric periodic eigenvalue problem. Numerical Algorithms, 57(4):513–536, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=4&spage=513.

Antonietti:2016:HOD

[AMM16] Paola F. Antonietti, Carlo Marcati, and Ilario Mazzieri. High order discontinuous Galerkin methods on simplicial elements for the elastodynamics equation. Numerical Algorithms, 71(1): 181–206, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0021-7.

Antonana:2017:RMR

[AMM17] Mikel Antoñana, Joseba Makazaga, and Ander Murua. Reducing and monitoring round-off error propagation for symplectic implicit Runge–Kutta schemes. Numerical Algorithms, 76(4): 861–880, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Antonana:2018:EIS

[AMM18] Mikel Antoñana, Joseba Makazaga, and Ander Murua. Efficient implementation of symplectic implicit Runge–Kutta schemes with simplified Newton iterations. Numerical Algorithms, 78(1):63–86, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amodei:2002:ODH

[Amo02] L. Amodei. Orthogonal decompositions in Hilbertian subspaces, error functions and optimal extensions of interpolation systems. Numerical Algorithms, 30(2):157–184, June 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/41/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
41/4/fulltext.pdf.

Arandiga:2015:CAI

[AMR15] Francesc Aràndiga, Pep Mulet, and Vicent Renau. Cell average image transform algorithms with exact error control. Numerical Algorithms, 69(1):75–93, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9882-4.

Alqahtani:2023:SIP

[AMR23] A. Alqahtani, T. Mach, and L. Reichel. Solution of ill-posed problems with Chebfun. Numerical Algorithms, 92(4):2341– 2364, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01390-z.

Ahookhosh:2017:OSA

[AN17] Masoud Ahookhosh and Arnold Neumaier. Optimal subgradient algorithms for large-scale convex optimization in simple domains. Numerical Algorithms, 76(4):1071–1097, December 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-017-0297-x.pdf.

Axelsson:2014:CIM

[ANA14] Owe Axelsson, Maya Neytcheva, and Bashir Ahmad. A comparison of iterative methods to solve complex valued linear algebraic systems. Numerical Algorithms, 66(4):811–841, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9764-1.

Anderson:1997:PAP

[And97] I. J. Anderson. A piecewise approach to piecewise approximation. Numerical Algorithms, 15(2):139–152, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/7/1/abstract.htm;   http:/

/ipsapp007.kluweronline.com/content/getfile/5058/7/
1/fulltext.pdf.

Andrei:2006:AGD

[And06] Neculai Andrei. An acceleration of gradient descent algorithm with backtracking for unconstrained optimization. Numerical Algorithms, 42(1):63–73, May 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=1&spage=63.

Andrei:2008:AHC

[And08] Neculai Andrei. Another hybrid conjugate gradient algorithm for unconstrained optimization. Numerical Algorithms, 47(2):143–156, February 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=143.

Andrei:2010:AHC

[And10] Neculai Andrei. Accelerated hybrid conjugate gradient algorithm with modified secant condition for unconstrained optimization. Numerical Algorithms, 54(1):23–46, May 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=54&issue=1&
spage=23.

Andreev:2014:STD

[And14a] Roman Andreev. Space–time discretization of the heat equation. Numerical Algorithms, 67(4):713–731, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9818-4.

Andrei:2014:ASM

[And14b] Neculai Andrei. An accelerated subspace minimization threeterm conjugate gradient algorithm for unconstrained optimization. Numerical Algorithms, 65(4):859–874, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9718-7.

Andrei:2015:NTT

[And15] Neculai Andrei. A new three-term conjugate gradient algorithm for unconstrained optimization. Numerical Algorithms, 68(2):305–321, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9845-9. See

comment [Don16].

Andrei:2018:ASB

[And18a] Neculai Andrei. An adaptive scaled BFGS method for unconstrained optimization. Numerical Algorithms, 77(2):413–432, February 2018. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-017-0321-1.

Andrei:2018:DLC

[And18b] Neculai Andrei. A Dai–Liao conjugate gradient algorithm with clustering of eigenvalues. Numerical Algorithms, 77(4):1273– 1282, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anderson:2019:CAA

[And19a] Donald G. M. Anderson. Comments on “Anderson acceleration, mixing and extrapolation”. Numerical Algorithms, 80(1): 135–234, January 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Andrei:2019:DQN

[And19b] Neculai Andrei. A diagonal quasi-Newton updating method for unconstrained optimization. Numerical Algorithms, 81(2):575– 590, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Andrei:2022:NML

[And22] Neculai Andrei. A note on memory-less SR1 and memoryless BFGS methods for large-scale unconstrained optimization. Numerical Algorithms, 90(1):223–240, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01185-8.

Anh:2019:LMB

[Anh19] Tran Viet Anh. Linesearch methods for bilevel split pseudomonotone variational inequality problems. Numerical Algorithms, 81(3):1067–1087, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Akaiwa:2015:FSC

[ANI15] Kanae Akaiwa, Yoshimasa Nakamura, and Masashi Iwasaki. A finite-step construction of totally nonnegative matrices with specified eigenvalues. Numerical Algorithms, 70(3):469–484, November 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9957-x.

Akaiwa:2017:ABS

[ANI^{+}17] Kanae Akaiwa, Yoshimasa Nakamura, Masashi Iwasaki, Akira Yoshida, and Koichi Kondo. An arbitrary band structure construction of totally nonnegative matrices with prescribed eigenvalues. Numerical Algorithms, 75(4):1079–1101, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:1992:IMC

[Ano92] Anonymous, editor. International Mathematical Congress on Extrapolation and Rational Approximation, volume 3(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, December 1992. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Anonymous:1993:NAW

[Ano93] Anonymous, editor. NATO Advanced Workshop on Algorithms for Approximation, volume 5(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, November 1993. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Anonymous:1995:BR

[Ano95a] Anonymous. Book reviews. Numerical Algorithms, 9(3–4):399– 406, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:1995:IJS

[Ano95b] Anonymous, editor. International Joint Symposium on Special Functions and Artificial Intelligence, volume 10(1–2) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, July 1995. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Anonymous:1995:WAC

[Ano95c] Anonymous, editor. Workshop on Algorithms for Constrained Approximation and Optimization, volume 9(1–2) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, May 1995. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Anonymous:1997:BR

[Ano97] Anonymous. Book reviews. Numerical Algorithms, 14 (4):401–403, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
6/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/6/7/fulltext.pdf.

Anonymous:1998:BRa

[Ano98a] Anonymous. Book reviews. Numerical Algorithms, 17(3– 4):357–358, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/12/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/12/10/fulltext.pdf.

Anonymous:1998:BRb

[Ano98b] Anonymous. Book reviews. Numerical Algorithms, 18 (1):109–112, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
13/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/13/6/fulltext.pdf.

Anonymous:1998:BRc

[Ano98c] Anonymous. Book reviews. Numerical Algorithms, 18(3–4): 385–388, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:1998:DP

[Ano98d] Anonymous. DAE ’97 participants. Numerical Algorithms, 19(1–4):261–264, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
16/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/2/fulltext.pdf.

Anonymous:1998:P

[Ano98e] Anonymous. Preface. Numerical Algorithms, 19(1–4):0, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/16/1/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/16/1/fulltext.pdf.

Anonymous:1999:BRa

[Ano99a] Anonymous. Book reviews. Numerical Algorithms, 20 (1):101–105, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
17/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/17/6/fulltext.pdf.

Anonymous:1999:BRb

[Ano99b] Anonymous. Book reviews. Numerical Algorithms, 20 (4):379–382, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
19/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/19/9/fulltext.pdf.

Anonymous:1999:BRc

[Ano99c] Anonymous. Book reviews. Numerical Algorithms, 22 (2):231–235, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
22/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/22/8/fulltext.pdf.

Anonymous:1999:BRd

[Ano99d] Anonymous. Book reviews. Numerical Algorithms, 22(3– 4):385–388, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/8/fulltext.pdf.

Anonymous:1999:LP

[Ano99e] Anonymous. List of participants. Numerical Algorithms, 21 (1–4):393–396, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
28/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/28/fulltext.pdf.

Anonymous:1999:P

[Ano99f] Anonymous. Preface. Numerical Algorithms, 21(1–4): 0, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/20/1/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/20/1/fulltext.pdf.

Anonymous:2000:BR

[Ano00a] Anonymous. Book reviews. Numerical Algorithms, 23 (4):407–408, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
26/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/26/7/fulltext.pdf.

Anonymous:2000:P

[Ano00b] Anonymous. Preface. Numerical Algorithms, 25(1–4): 11–14, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/28/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/28/fulltext.pdf.

Anonymous:2000:PA

[Ano00c] Anonymous. Publisher’s announcement. Numerical Algorithms, 25(1–4):1, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

3/29/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/29/fulltext.pdf.

Anonymous:2000:RRV

[Ano00d] Anonymous. Reminiscences of Richard S. Varga. Numerical Algorithms, 25(1–4):15–16, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/1/fulltext.pdf.

Anonymous:2001:AIa

[Ano01a] Anonymous. Author index. Numerical Algorithms, 26 (4):385–386, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
32/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/32/7/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/32/
7/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/32/7/fulltext.pdf.

Anonymous:2001:AIb

[Ano01b] Anonymous. Author index. Numerical Algorithms, 27(4):391– 392, August 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/6/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
6/fulltext.pdf.

Anonymous:2001:BRa

[Ano01c] Anonymous. Book reviews. Numerical Algorithms, 26 (2):197–200, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
1/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/1/8/fulltext.pdf.

Anonymous:2001:BRb

[Ano01d] Anonymous. Book reviews. Numerical Algorithms, 26 (3):301–303, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

2/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/2/6/fulltext.pdf.

Anonymous:2001:BRc

[Ano01e] Anonymous. Book reviews. Numerical Algorithms, 26 (4):381–384, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
32/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/32/6/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/32/
6/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/32/6/fulltext.pdf.

Anonymous:2001:BRd

[Ano01f] Anonymous. Book reviews. Numerical Algorithms, 27(1):115– 117, May 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/33/6/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/33/
6/fulltext.pdf.

Anonymous:2001:BRe

[Ano01g] Anonymous. Book reviews. Numerical Algorithms, 27(4):387– 390, August 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/7/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
7/fulltext.pdf.

Anonymous:2001:P

[Ano01h] Anonymous. Preface. Numerical Algorithms, 28(1–4):1– 2, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/37/21/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/37/21/fulltext.pdf.

Anonymous:2002:AIa

[Ano02a] Anonymous. Author index. Numerical Algorithms, 29 (4):379–380, April 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

39/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/39/4/fulltext.pdf.

Anonymous:2002:AIb

[Ano02b] Anonymous. Author index. Numerical Algorithms, 30 (3–4):353–354, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
42/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/42/7/fulltext.pdf.

Anonymous:2002:BR

[Ano02c] Anonymous. Book reviews. Numerical Algorithms, 30(1): 91–92, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/40/7/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/40/7/fulltext.pdf.

Anonymous:2002:P

[Ano02d] Anonymous. Preface. Numerical Algorithms, 29(4):281, April 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/39/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
39/3/fulltext.pdf.

Anonymous:2003:BR

[Ano03a] Anonymous. Book reviews. Numerical Algorithms, 34 (1):103–106, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.kluweronline.com/content/getfile/5058/
47/7/abstract.htm;   http://ipsapp008.kluweronline.
com/content/getfile/5058/47/7/fulltext.pdf.

Anonymous:2003:LP

[Ano03b] Anonymous. List of participants. Numerical Algorithms, 34 (2–4):441–452, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/10/fulltext.pdf.

Anonymous:2003:P

[Ano03c] Anonymous. Preface. Numerical Algorithms, 33(1–4):1–2, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/43/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
46/43/fulltext.pdf.

Anonymous:2004:AIV

[Ano04a] Anonymous. Author index, volume 36 (2004). Numerical Algorithms, 36(4):381, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:2004:BR

[Ano04b] Anonymous. Book reviews. Numerical Algorithms, 35(1): 131–136, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/49/A/9/
abstract.htm.

Anonymous:2004:CV

[Ano04c] Anonymous. Contents of volume 36 (2004). Numerical Algorithms, 36(4):383–384, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:2004:LP

[Ano04d] Anonymous. List of participants. Numerical Algorithms, 35 (2–4):389–395, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/50/A/11/
abstract.htm.

Anonymous:2005:BR

[Ano05a] Anonymous. Book reviews. Numerical Algorithms, 38(4):327– 331, April 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:2005:CPS

[Ano05b] Anonymous. Chebyshev polynomials and spectral methods. Numerical Algorithms, 38(1–3):1–2, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anonymous:2005:P

[Ano05c] Anonymous. Preface. Numerical Algorithms, 40(2):101, October 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=40&
issue=2&spage=101.

Anonymous:2006:BR

[Ano06] Anonymous. Book reviews. Numerical Algorithms, 42(1):91– 94, May 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
42&issue=1&spage=91.

Anonymous:2017:SIR

[Ano17] Anonymous. Special issue: recent progress on iterative methods for large systems of equations. Numerical Algorithms, 75 (2):337–338, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Antunes:2018:NAR

[Ant18] Pedro R. S. Antunes. A numerical algorithm to reduce illconditioning in meshless methods for the Helmholtz equation. Numerical Algorithms, 79(3):879–897, November 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Antunes:2022:WCM

[Ant22] Pedro R. S. Antunes. A well-conditioned method of fundamental solutions for Laplace equation. Numerical Algorithms, 91 (3):1381–1405, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01306-
x.

Asharabi:2021:ASF

[AP21] R. M. Asharabi and J. Prestin. Accurate sampling formula for approximating the partial derivatives of bivariate analytic functions. Numerical Algorithms, 86(4):1421–1441, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00939-0.

Abdi:2026:VSG

[AP26] Ali Abdi and Helmut Podhaisky. Variable stepsize general linear methods for ODEs. Numerical Algorithms, 101(1):243–271, January 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-02004-6.

Alfaro:2014:LRO

[APPR14] Manuel Alfaro, Ana Peña, Teresa E. Pérez, and M. Luisa Rezola. On linearly related orthogonal polynomials in several variables. Numerical Algorithms, 66(3):525–553, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9747-2.

Aldhafeeri:2021:RGI

[APST21] Nahier Aldhafeeri, Dimitrios Pappas, Ivan P. Stanimirović, and Milan Tasić. Representations of generalized inverses via full-rank QDR decomposition. Numerical Algorithms, 86(3): 1327–1337, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00935-4.

Amraoui:1999:RFD

[AR99] S. Amraoui and S. Lalaoui Rhali. Retarded functional differential equations with nondense domain operators. Numerical Algorithms, 21(1–4):1–8, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
27/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/27/fulltext.pdf.

Argyros:2009:ILC

[AR09] Ioannis K. Argyros and Hongmin Ren. On an improved local convergence analysis for the Secant method. Numerical Algorithms, 52(2):257–271, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=2&spage=257.

Arico:2010:FSL

[AR10] Antonio Aricò and Giuseppe Rodriguez. A fast solver for linear systems with displacement structure. Numerical Algorithms, 55(4):529–556, December 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=4&spage=529.

Akram:2013:NSE

[AR13] Ghazala Akram and Hamood Ur Rehman. Numerical solution of eighth order boundary value problems in reproducing kernel space. Numerical Algorithms, 62(3):527–540, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9608-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=527-540.

Acu:2016:NED

[AR16] Ana Maria Acu and Ioan Rasa. New estimates for the differences of positive linear operators. Numerical Algorithms, 73(3): 775–789, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0117-8.

Alqahtani:2018:SAG

[AR18] Hessah Alqahtani and Lothar Reichel. Simplified anti-Gauss quadrature rules with applications in linear algebra. Numerical Algorithms, 77(2):577–602, February 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-017-0329-6.

Acu:2020:EDP

[AR20] Ana-Maria Acu and Ioan Rasa. Estimates for the differences of positive linear operators and their derivatives. Numerical Algorithms, 85(1):191–208, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00809-
4.

Abdennebi:2024:ASS

[AR24] Issam Abdennebi and Azedine Rahmoune. Adaptive spectral solution method for Fredholm integral equations of the second kind. Numerical Algorithms, 96(2):975–998, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01671-1.

Alsubhi:2026:EGR

[AR26] Abdulmajeed Alsubhi and Rosemary A. Renaut. Enlarged GKB and RSVD for KP approximations illustrated for mixed precision ℓ_{1} image restoration. Numerical Algorithms, 101 (4):2245–2267, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02086-w.

Argyros:2009:UMU

[Arg09] Ioannis K. Argyros. On Ulm’s method using divided differences of order one. Numerical Algorithms, 52(3):295–320, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=3&
spage=295.

Argyros:2010:CSL

[Arg10] Ioannis K. Argyros. On a class of secant-like methods for solving nonlinear equations. Numerical Algorithms, 54(4):485–501, August 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=4&spage=485.

Ariyawansa:1998:LCC

[Ari98] K. A. Ariyawansa. Local convergence of collinear scaling algorithms related to direct least-change secant update methods for minimization. Numerical Algorithms, 18(3– 4):293–320, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/5/fulltext.pdf.

Achachi:2003:CAB

[ARJ03] Nisrine Achachi and Françoise Richard-Jung. Computer algebra and bifurcations. Numerical Algorithms, 34(2– 4):107–115, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/

11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/11/fulltext.pdf.

Arnold:1997:FCF

[Arn97] C. Arnold. Formal continued fractions solutions of the generalized second order Riccati equations, applications. Numerical Algorithms, 15(1):111–134, December 1997. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/30/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/7/fulltext.pdf.

Alam:2026:ARI

[ARS26] Khairul Habib Alam, Yumnam Rohen, and S. Surendra Singh. Analysis of a refined iterative method with a new setting and applications to various models of partial differential equations. Numerical Algorithms, 102(2):875–900, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02159-w.

Andreani:2019:BOM

[ARSS19] Roberto Andreani, Viviana A. Ramirez, Sandra A. Santos, and Leonardo D. Secchin. Bilevel optimization with a multiobjective problem in the lower level. Numerical Algorithms, 81 (3):915–946, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amat:2014:ICR

[ART14] Sergio Amat, Juan Ruiz, and J. Carlos Trillo. Improving the compression rate versus L^{1} error ratio in cell-average error control algorithms. Numerical Algorithms, 67(1):145–162, September 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9780-1.

Amat:2019:AAS

[ART19] Sergio Amat, Juan Ruiz, and Juan C. Trillo. On an algorithm to adapt spline approximations to the presence of discontinuities. Numerical Algorithms, 80(3):903–936, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Amat:2020:FNO

[ARTY20] Sergio Amat, Juan Ruiz, Juan C. Trillo, and Dionisio F. Yáñez. On a family of non-oscillatory subdivision schemes having regularity C^{r} with r > 1. Numerical Algorithms, 85(2):543–569, October 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00826-3.

Alkilayh:2017:NZF

[ARY17] Maged Alkilayh, Lothar Reichel, and Jin Yun Yuan. New zerofinders for trust-region computations. Numerical Algorithms, 76(2):361–375, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Al-Subaihi:2008:AFC

[AS08] I. A. Al-Subaihi. An algorithm for fitting circular arcs to data using the l_{1} norm. Numerical Algorithms, 47(1):1–14, January 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=1&
spage=1.

Abbasbandy:2010:HAM

[AS10] Saeid Abbasbandy and A. Shirzadi. Homotopy analysis method for multiple solutions of the fractional Sturm–Liouville problems. Numerical Algorithms, 54(4):521–532, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=54&issue=4&
spage=521.

Abbasbandy:2011:NAT

[AS11] Saeid Abbasbandy and E. Shivanian. A new analytical technique to solve Fredholm’s integral equations. Numerical Algorithms, 56(1):27–43, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=27.

Arioli:2014:CAI

[AS14] M. Arioli and J. Scott. Chebyshev acceleration of iterative refinement. Numerical Algorithms, 66(3):591–608, July 2014.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-013-9750-7.pdf.

Ascher:1997:SID

[Asc97] Uri M. Ascher. Stabilization of invariants of discretized differential systems. Numerical Algorithms, 14(1–3):1–24, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/5/2/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/5/

2/fulltext.pdf. Dynamical numerical analysis (Atlanta, GA, 1995).

Acosta-Soba:2023:UDS

[ASGGRG23] Daniel Acosta-Soba, Francisco Guillén-González, and J. Rafael Rodríguez-Galván. An upwind DG scheme preserving the maximum principle for the convective Cahn–Hilliard model. Numerical Algorithms, 92(3):1589–1619, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01355-2.

Anciaux-Sedrakian:2020:ASL

[ASGJ^{+}20] A. Anciaux-Sedrakian, L. Grigori, Z. Jorti, J. Papez, and S. Yousef. Adaptive solution of linear systems of equations based on a posteriori error estimators. Numerical Algorithms, 84(1):331–364, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Asharabi:2016:GSG

[Ash16] R. M. Asharabi. Generalized sinc-Gaussian sampling involving derivatives. Numerical Algorithms, 73(4):1055–1072, December 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0129-4.

Asharabi:2019:USG

[Ash19] Rashad M. Asharabi. The use of the sinc-Gaussian sampling formula for approximating the derivatives of analytic functions. Numerical Algorithms, 81(1):293–312, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Arevalo:2021:LEE

[ASHF21] Carmen Arévalo, Gustaf Söderlind, Yiannis Hadjimichael, and Imre Fekete. Local error estimation and step size control in adaptive linear multistep methods. Numerical Algorithms, 86(2):537–563, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00900-
1;   http://link.springer.com/content/pdf/10.1007/
s11075-020-00900-1.pdf.

Ameur:2003:GMM

[ASS03] El Bachir Ameur, Driss Sbibih, and Paul Sablonniere. A general multiresolution method for fitting functions on the sphere. Numerical Algorithms, 34(2–4):159–171, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/15/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
15/fulltext.pdf.

Akkas:2004:ICS

[ASS04] Ahmet Akkaş, Michael J. Schulte, and James E. Stine. Intrinsic compiler support for interval arithmetic. Numerical Algorithms, 37(1–4):13–20, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/2/abstract.htm.

Allouch:2011:SFI

[ASS11] Chafik Allouch, Paul Sablonnière, and Driss Sbibih. Solving Fredholm integral equations by approximating kernels by spline quasi-interpolants. Numerical Algorithms, 56(3):437– 453, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
56&issue=3&spage=437.

Allouch:2013:CMN

[ASS13] Chafik Allouch, Paul Sablonnière, and Driss Sbibih. A collocation method for the numerical solution of a two dimensional integral equation using a quadratic spline quasiinterpolant. Numerical Algorithms, 62(3):445–468, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9598-2/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=445-468.

Agrawal:2023:NSF

[ASV23] V. Agrawal, T. Som, and S. Verma. A note on stability and fractal dimension of bivariate α-fractal functions. Numerical Algorithms, 93(4):1811–1833, August 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01490-w.

Araujo:2021:FCB

[ASVC21a] Leonardo C. Araujo, João P. H. Sansão, and Adriano S. Vale-Cardoso. Fast computation of binomial coefficients. Numerical Algorithms, 86(2):799–812, February 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-00912-x.

Araujo:2021:FCM

[ASVC21b] Leonardo C. Araujo, João P. H. Sansão, and Adriano S. Vale-Cardoso. Fast computation of multinomial coefficients. Numerical Algorithms, 88(2):837–851, October 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-01059-5.

Al-Subaihi:2006:AMP

[ASW06] I. Al-Subaihi and G. A. Watson. An algorithm for matching point sets using the l_{1} norm. Numerical Algorithms, 41(3):203–217, March 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=3&spage=203.

AlMahbub:2023:UEG

[ASZ23] Md. Abdullah Al Mahbub, Li Shan, and Haibiao Zheng. Uncoupling evolutionary groundwater-surface water flows: stabilized mixed methods in both porous media and fluid regions. Numerical Algorithms, 92(3):1837–1874, March 2023.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01370-3.

An:2012:PAB

[AT12] Phan Thanh An and Le Hong Trang. A parallel algorithm based on convexity for the computing of Delaunay tessellation. Numerical Algorithms, 59(3):347–357, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=347.

Asharabi:2017:AED

[AT17] Rashad M. Asharabi and Mohammed M. Tharwat. Approximating eigenvalues of Dirac system with discontinuities at several points using Hermite–Gauss method. Numerical Algorithms, 76(3):655–673, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Annaby:2019:SRT

[AT19] M. H. Annaby and M. M. Tharwat. Sinc-regularized techniques to compute eigenvalues of Schrödinger operators on L^{2}(I)⊕C^{2}. Numerical Algorithms, 80(3):795–817, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anh:2021:SPM

[AT21] Pham Ngoc Anh and Ho Phi Tu. Subgradient projection methods extended to monotone bilevel equilibrium problems in Hilbert spaces. Numerical Algorithms, 86(1):55–74, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00878-w.

Ahmad:2016:PMS

[ATC16] Fayyaz Ahmad, Emran Tohidi, and Juan A. Carrasco. A parameterized multi-step Newton method for solving systems of nonlinear equations. Numerical Algorithms, 71(3):631–653, March 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0013-7.

Adams:2019:ECG

[ATM19] Mark Adams, Connor Tannahill, and Paul Muir. Error control Gaussian collocation software for boundary value ODEs and 1D time-dependent PDEs. Numerical Algorithms, 81(4): 1505–1519, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Anh:2021:HPM

[ATT21] Pham Ngoc Anh, T. V. Thang, and H. T. C. Thach. Halpern projection methods for solving pseudomonotone multivalued variational inequalities in Hilbert spaces. Numerical Algorithms, 87(1):335–363, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00968-9.

Anh:2022:SPM

[ATT22] P. N. Anh, T. V. Thang, and H. T. C. Thach. A subgradient proximal method for solving a class of monotone multivalued variational inequality problems. Numerical Algorithms, 89(1): 409–430, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01119-4.

Argyros:2008:CMM

[AU08] Ioannis K. Argyros and Livinus U. Uko. On the convergence of the midpoint method. Numerical Algorithms, 47(2):157–167, February 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
47&issue=2&spage=157.

Ahmad:2022:CWE

[AUA22] Junaid Ahmad, Kifayat Ullah, and Muhammad Arshad. Convergence, weak w^{2} stability, and data dependence results for the F iterative scheme in hyperbolic spaces. Numerical Algorithms, 91(4):1755–1778, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01321-y.

Aral:2018:NCS

[AUD18] Ali Aral, Gulsum Ulusoy, and Emre Deniz. A new construction of Szász–Mirakyan operators. Numerical Algorithms, 77 (2):313–326, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-017-0317-x.

Anh:2019:SAG

[AV19] Pham Ky Anh and Nguyen The Vinh. Self-adaptive gradient projection algorithms for variational inequalities involving non-Lipschitz continuous operators. Numerical Algorithms, 81(3): 983–1001, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Angulo:2026:CFI

[AVCO26] Wilfredo Angulo, Miguel Vivas-Cortez, and Juan Carlos Osorio. A conformable fractional iterative fixed-point method for solving one-dimensional nonlinear equations. Numerical Algorithms, 102(2):953–979, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02162-
1.

Averseng:2020:FDC

[Ave20] Martin Averseng. Fast discrete convolution in R^{2} with radial kernels using non-uniform Fast Fourier Transform with nonequispaced frequencies. Numerical Algorithms, 83(1):33– 56, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Averbuch:1997:FDE

[AVI97] A. Averbuch, L. Vozovoi, and M. Israeli. On a fast direct elliptic solver by a modified Fourier method. Numerical Algorithms, 15(3–4):287–313, January 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Arbogast:2023:DSM

[AW23] Todd Arbogast and Chuning Wang. Direct serendipity and mixed finite elements on convex polygons. Numerical Algorithms, 92(2):1451–1483, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01348-
1.

Awawdeh:2010:NIM

[Awa10] Fadi Awawdeh. On new iterative method for solving systems of nonlinear equations. Numerical Algorithms, 54(3):395–409, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=1017-1398&volume=54&
issue=3&spage=395.

Ahmed:2024:AHZ

[AWH^{+}24] Kabiru Ahmed, Mohammed Yusuf Waziri, Abubakar Sani Halilu, Salisu Murtala, and Jamilu Sabi’u. Another Hager– Zhang-type method via singular-value study for constrained monotone equations with application. Numerical Algorithms, 96(4):1583–1623, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01678-
8.

An:2024:PET

[AWL^{+}24] Xingyu An, Qingxia Wang, Fawang Liu, Vo V. Anh, and Ian W. Turner. Parameter estimation for time-fractional Black–Scholes equation with S&P 500 index option. Numerical Algorithms, 95(1):1–30, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01563-
4.

An:2026:RBB

[AX26] Congpei An and Xin Xu. Regularized Barzilai–Borwein method. Numerical Algorithms, 102(1):25–52, May 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02120-x.

Axelsson:1999:SAM

[Axe99] Owe Axelsson. Stabilization of algebraic multilevel iteration methods: additive methods. Numerical Algorithms, 21 (1–4):23–47, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/3/fulltext.pdf. Numer-

ical methods for partial differential equations (Marrakech, 1998).

Alinia:2019:NAB

[AZ19a] N. Alinia and M. Zarebnia. A numerical algorithm based on a new kind of tension B-spline function for solving Burgers– Huxley equation. Numerical Algorithms, 82(4):1121–1142, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zou:2019:NST

[aZ19b] Guang an Zou. Numerical solutions to time-fractional stochastic partial differential equations. Numerical Algorithms, 82(2): 553–571, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Abe:2004:VOM

[AZMJ04] Kuniyoshi Abe, Shao-Liang Zhang, Taketomo Mitsui, and Cheng-Hai Jin. A variant of the ORTHOMIN(2) method for singular linear systems. Numerical Algorithms, 36(3):189–202, July 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/57/A/1/abstract.htm.

Baccouch:2018:SLD

[Bac18] Mahboub Baccouch. A superconvergent local discontinuous Galerkin method for nonlinear two-point boundary-value problems. Numerical Algorithms, 79(3):697–718, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Baccouch:2020:ALD

[Bac20] Mahboub Baccouch. An adaptive local discontinuous Galerkin method for nonlinear two-point boundary-value problems. Numerical Algorithms, 84(3):1121–1153, July 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-019-00794-8.

Baccouch:2021:AOS

[Bac21] Mahboub Baccouch. Analysis of optimal superconvergence of the local discontinuous Galerkin method for nonlinear fourthorder boundary value problems. Numerical Algorithms, 86(4): 1615–1650, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00947-0.

Baccouch:2023:SUW

[Bac23] Mahboub Baccouch. A superconvergent ultra-weak local discontinuous Galerkin method for nonlinear fourth-order boundary-value problems. Numerical Algorithms, 92(4):1983– 2023, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01374-z.

Baccouch:2024:PEA

[Bac24] Mahboub Baccouch. A posteriori error analysis of an ultraweak local discontinuous Galerkin method for nonlinear fourthorder boundary-value problems. Numerical Algorithms, 97(4): 1895–1933, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01773-4.

Baccouch:2025:RBP

[Bac25] Mahboub Baccouch. Residual-based a posteriori error analysis of an ultra-weak discontinuous Galerkin method for nonlinear second-order initial-value problems. Numerical Algorithms, 98 (1):373–404, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01799-8.

Ben-Artzi:2025:NSB

[BAEF25] M. Ben-Artzi, Y. Eidelman, and D. Fishelov. Numerical solution of the boundary value problems for the biharmonic equations via quasiseparable representations. Numerical Algorithms, 98(2):625–649, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01809-9.

Baglama:2000:DLD

[Bag00] J. Baglama. Dealing with linear dependence during the iterations of the restarted block Lanczos methods. Numerical Algorithms, 25(1–4):23–36, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/23/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/23/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Bagherboum:2016:AFP

[Bag16] Mozhgan Bagherboum. Approximating fixed points of mappings satisfying condition (E) in Busemann space. Numerical Algorithms, 71(1):25–39, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9982-9.

Bai:1997:CTS

[Bai97a] Zhong-Zhi Bai. A class of two-stage iterative methods for systems of weakly nonlinear equations. Numerical Algorithms, 14(4):295–319, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
6/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/6/2/fulltext.pdf.

Bai:1997:PMT

[Bai97b] Zhong-Zhi Bai. Parallel multisplitting two-stage iterative methods for large sparse systems of weakly nonlinear equations. Numerical Algorithms, 15(3–4):347–372, January 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Barari:2025:IDP

[BAK25] Suma Barari, Keyvan Amini, and Ahmad Kamandi. An inertial-double-projection-based algorithm for large-scale nonlinear systems of monotone equations. Numerical Algorithms, 100(3):963–982, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01975-
w.

Baltensperger:2011:BRI

[Bal11] Richard Baltensperger. Barycentric rational interpolation with asymptotically monitored poles. Numerical Algorithms, 57(1): 67–81, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
57&issue=1&spage=67.

Barkatou:1991:CRS

[Bar91] M. A. Barkatou. Characterization of regular singular linear systems of difference equations. Numerical Algorithms, 1(2):

139–154, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Barry:2013:PBL

[Bar13] Phillip Barry. Properties of an h-Bernstein-like basis and curves. Numerical Algorithms, 63(3):453–481, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9632-4.

Bhadala:2025:OIR

[BAS25] Neha Bhadala, Sk. Safique Ahmad, and Predrag S. Stanimirović. Outer inverse of reduced biquaternion matrices. Numerical Algorithms, 100(3):995–1022, November 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01977-8.

Bakhtiari:2018:SDI

[BAV18] P. Bakhtiari, S. Abbasbandy, and R. A. Van Gorder. Solving the Dym initial value problem in reproducing kernel space. Numerical Algorithms, 78(2):405–421, June 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bellour:2014:TCM

[BB14a] Azzeddine Bellour and Mahmoud Bousselsal. A Taylor collocation method for solving delay integral equations. Numerical Algorithms, 65(4):843–857, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9717-8.

Burrage:2014:SPR

[BB14b] Pamela M. Burrage and Kevin Burrage. Structure-preserving Runge–Kutta methods for stochastic Hamiltonian equations with additive noise. Numerical Algorithms, 65(3):519–532, March 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9796-6.

Breuss:2006:NAT

[BBB^{+}06] Michael Breuß, Andrea Bürgel, Thomas Brox, Thomas Sonar, and Joachim Weickert. Numerical aspects of TV flow. Numerical Algorithms, 41(1):79–101, January 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=1&spage=79.

Betcke:2022:HCF

[BBB22] Timo Betcke, Michał Bosy, and Erik Burman. Hybrid coupling of finite element and boundary element methods using Nitsche’s method and the Calderon projection. Numerical Algorithms, 91(3):997–1019, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01289-9.

Bartholomew-Biggs:2020:CMT

[BBBC20] Michael Bartholomew-Biggs, Salah Beddiaf, and Bruce Christianson. A comparison of methods for traversing regions of nonconvexity in optimization problems. Numerical Algorithms, 85(1):231–253, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00811-w.

Bartholomew-Biggs:2023:GCC

[BBBC23] Michael Bartholomew-Biggs, Salah Beddiaf, and Bruce Christianson. Global convergence of a curvilinear search for nonconvex optimization. Numerical Algorithms, 92(4):2025–2043, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01375-y.

Bai:2011:PMI

[BBC11] Zhong-Zhi Bai, Michele Benzi, and Fang Chen. On preconditioned MHSS iteration methods for complex symmetric linear systems. Numerical Algorithms, 56(2):297–317, February 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=2&
spage=297.

Banert:2021:FES

[BBC21] Sebastian Banert, Radu Ioan Bot, and Ernö Robert Csetnek. Fixing and extending some recent results on the ADMM algorithm. Numerical Algorithms, 86(3):1303–1325, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00934-5;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00934-5.pdf.

Bellet:2025:SHC

[BBC25] Jean-Baptiste Bellet, Matthieu Brachet, and Jean-Pierre Croisille. Spherical harmonics collocation: a computational intercomparison of several grids. Numerical Algorithms, 100(4): 2021–2040, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02170-1.

Behling:2021:CRM

[BBCS21] Roger Behling, Yunier Bello-Cruz, and Luiz-Rafael Santos. On the circumcentered-reflection method for the convex feasibility problem. Numerical Algorithms, 86(4):1475–1494, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00941-6.

Bojanczyk:1995:SAG

[BBd95] Adam W. Bojańczyk, Richard P. Brent, and Frank R. de Hoog. Stability analysis of a general Toeplitz systems solver. Numerical Algorithms, 10(3–4):225–244, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Baltensperger:2003:LRP

[BBD03] Richard Baltensperger, Jean-Paul Berrut, and Yves Dubey. The linear rational pseudospectral method with preassigned poles. Numerical Algorithms, 33(1–4):53–63, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
46/5/fulltext.pdf.

Bergam:2003:EIM

[BBHM03] A. Bergam, C. Bernardi, F. Hecht, and Z. Mghazli. Error indicators for the mortar finite element discretization of a parabolic problem. Numerical Algorithms, 34(2–4):187–201, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/17/abstract.htm;   http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
17/fulltext.pdf.

Bentbib:2025:EMM

[BBJ25] Abdeslem Hafid Bentbib, Maryam Boubekraoui, and Khalide Jbilou. Extrapolation methods for multilinear PageRank. Numerical Algorithms, 98(2):1013–1043, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01823-x.

Beckermann:2022:RAM

[BBL22a] Bernhard Beckermann, Joanna Bisch, and Robert Luce. On the rational approximation of Markov functions, with applications to the computation of Markov functions of Toeplitz matrices. Numerical Algorithms, 91(1):109–144, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01256-4.

Burrage:2022:ENM

[BBL22b] Kevin Burrage, Pamela M. Burrage, and Grant Lythe. Effective numerical methods for simulating diffusion on a spherical surface in three dimensions. Numerical Algorithms, 91(4): 1577–1596, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01315-w.

Beddig:2023:LRU

[BBL23] Rebekka S. Beddig, Jörn Behrens, and Sabine Le Borne. A lowrank update for relaxed Schur complement preconditioners in fluid flow problems. Numerical Algorithms, 94(4):1597–1618, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01548-3.

Boggess:2008:FSL

[BBM08] A. Jason Boggess, B. Eric Bunch, and C. Charles N. Moore. Fourier series and the Lubkin W -transform. Numerical Algorithms, 47(2):133–142, February 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=133.

Bartmeyer:2021:SPU

[BBO21] Petra Maria Bartmeyer, Silvana Bocanegra, and Aurelio Ribeiro Leite Oliveira. Switching preconditioners using a hybrid approach for linear systems arising from interior point methods for linear programming. Numerical Algorithms, 86(1): 397–424, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00893-x.

Bhaya:2017:CCA

[BBP17] Amit Bhaya, Pierre-Alexandre Bliman, and Fernando Pazos. Cooperative concurrent asynchronous computation of the solution of symmetric linear systems. Numerical Algorithms, 75 (3):587–617, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beckermann:2012:CHP

[BBPV12] Bernhard Beckermann, Daniel Bessis, Luca Perotti, and Daniel Vrinceanu. Computing high precision Matrix Padé approximants. Numerical Algorithms, 61(2):189–208, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=2&
spage=189.

Barrachina:2007:EAG

[BBQO07] Sergio Barrachina, Peter Benner, and Enrique S. Quintana-Ortí. Efficient algorithms for generalized algebraic Bernoulli equations based on the matrix sign function. Numerical Algorithms, 46(4):351–368, December 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=4&spage=351.

Berti:2003:OPA

[BBR03] A. C. Berti, C. F. Bracciali, and A. Sri Ranga. Orthogonal polynomials associated with related measures and Sobolev orthogonal polynomials. Numerical Algorithms, 34 (2–4):203–216, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
18/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/18/fulltext.pdf.

Badahmane:2020:PKS

[BBS20] A. Badahmane, A. H. Bentbib, and H. Sadok. Preconditioned Krylov subspace and GMRHSS iteration methods for solving the nonsymmetric saddle point problems. Numerical Algorithms, 84(4):1295–1312, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00833-
4.

Bartholomew-Biggs:1995:UMI

[BBZ95] M. C. Bartholomew-Biggs and S. Zakovic. Using Markov’s interval arithmetic to evaluate Bessel–Riccati functions. Numerical Algorithms, 10(3–4):261–287, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beckermann:1992:RMC

[BC92] Bernhard Beckermann and Carsten Carstensen. A reliable modification of the cross rule for rational Hermite interpolation. Numerical Algorithms, 3(1–4):29–44, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bank:1994:CSB

[BC94] Randolph E. Bank and Tony F. Chan. A composite step biconjugate gradient algorithm for nonsymmetric linear systems. Numerical Algorithms, 7(1):1–16, June 1994. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Butcher:1999:EOS

[BC99] J. C. Butcher and P. Chartier. The effective order of singly-implicit Runge–Kutta methods. Numerical Algorithms, 20(4):269–284, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
19/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/19/7/fulltext.pdf.

Benouahmane:2000:MOP

[BC00] Brahim Benouahmane and Annie Cuyt. Multivariate orthogonal polynomials, homogeneous Padé approximants and Gaussian cubature. Numerical Algorithms, 24(1–2):1–15, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/2/fulltext.pdf. Computational methods from rational approximation theory (Wilrijk, 1999).

Butcher:2001:VSS

[BC01a] J. C. Butcher and T. M. H. Chan. Variable stepsize schemes for effective order methods and enhanced order composition methods. Numerical Algorithms, 26(2):131–150, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/1/4/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/1/
4/fulltext.pdf.

Butcher:2001:IEM

[BC01b] J. C. Butcher and D. J. L. Chen. On the implementation of ESIRK methods for stiff IVPs. Numerical Algorithms, 26(3):201–218, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
2/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/2/1/fulltext.pdf.

Bayro-Corrochano:2005:MRI

[BC05a] Eduardo Bayro-Corrochano. Multi-resolution image analysis using the quaternion wavelet transform. Numerical Algorithms, 39(1–3):35–55, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Blanes:2005:ROG

[BC05b] Sergio Blanes and Fernando Casas. Raising the order of geometric numerical integrators by composition and extrapolation. Numerical Algorithms, 38(4):305–326, April 2005. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Becuwe:2006:FST

[BC06] Stefan Becuwe and Annie Cuyt. On the fast solution of Toeplitz-block linear systems arising in multivariate approximation theory. Numerical Algorithms, 43(1):1–24, September 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=43&issue=1&
spage=1.

Bultheel:2009:MCR

[BC09] Adhemar Bultheel and Maria-José Cantero. A matricial computation of rational quadrature formulas on the unit circle. Numerical Algorithms, 52(1):47–68, September 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=1&spage=47.

Behloul:2014:PSN

[BC14] Djilali Behloul and Sui Sun Cheng. Polynomial solutions of a nonlinear difference equation. Numerical Algorithms, 65(2): 325–337, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9707-x.

Bot:2016:IFB

[BC16] Radu Ioan Bot and Ernö Robert Csetnek. An inertial forwardbackward-forward primal-dual splitting algorithm for solving monotone inclusion problems. Numerical Algorithms, 71(3): 519–540, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0007-5.

Bracciali:2017:QAR

[BC17] Cleonice F. Bracciali and Michael Carley. Quasi-analytical root-finding for non-polynomial functions. Numerical Algorithms, 76(3):639–653, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-017-
0274-4.pdf.

Brezinski:2022:RSV

[BC22] Claude Brezinski and Philippe G. Ciarlet. Richard Steven Varga, 1928–2022. Numerical Algorithms, 90(1):1–2, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01292-0.

Brachet:2026:HMS

[BC26] Matthieu Brachet and Jean-Paul Chehab. High modes smoothing of time schemes for nonlinear parabolic PDEs. Numerical Algorithms, 101(2):761–802, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02022-y.

Baglama:2025:HAC

[BCCP25] James Baglama, Jonathan A. Chávez-Casillas, and Vasilije Perović. A hybrid algorithm for computing a partial singular value decomposition satisfying a given threshold. Numerical Algorithms, 99(3):??, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01906-9.

Butcher:2011:PAB

[BCGVS11] John C. Butcher, Robert M. Corless, Laureano Gonzalez-Vega, and Azar Shakoori. Polynomial algebra for Birkhoff interpolants. Numerical Algorithms, 56(3):319–347, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=3&
spage=319.

Brugnano:2014:EIG

[BCI14] Luigi Brugnano, Gianluca Frasca Caccia, and Felice Iavernaro. Efficient implementation of Gauss collocation and Hamiltonian boundary value methods. Numerical Algorithms, 65(3): 633–650, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9825-0.

Butcher:1997:NRD

[BCJ97] J. C. Butcher, P. Chartier, and Z. Jackiewicz. Nordsieck representation of DIMSIMs. Numerical Algorithms, 16(2):209–230, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/8/fulltext.pdf.

Butcher:1999:EVO

[BCJ99] J. C. Butcher, P. Chartier, and Z. Jackiewicz. Experiments with a variable-order type 1 DIMSIM code. Numerical Algorithms, 22(3–4):237–261, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/9/fulltext.pdf.

Borachok:2022:MFS

[BCJ22] Ihor Borachok, Roman Chapko, and B. Tomas Johansson. A method of fundamental solutions for heat and wave propagation from lateral Cauchy data. Numerical Algorithms, 89(1): 431–449, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01120-x.

Borachok:2024:RMC

[BCJ24] Ihor Borachok, Roman Chapko, and B. Tomas Johansson. Rothe’s method in combination with a fundamental sequences method for the nonstationary Stokes problem. Numerical Algorithms, 96(1):59–73, May 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01639-1.

Bauschke:2006:EAA

[BCK06] Heinz H. Bauschke, Patrick L. Combettes, and Serge G. Kruk. Extrapolation algorithm for affine-convex feasibility problems. Numerical Algorithms, 41(3):239–274, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=3&spage=239.

Bacchelli:2000:ACB

[BCL00] S. Bacchelli, M. Cotronei, and D. Lazzaro. An algebraic construction of k-balanced multiwavelets via the lifting scheme. Numerical Algorithms, 23(4):329–356, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/26/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
26/5/fulltext.pdf.

Berriochoa:2007:NNQ

[BCM07] E. Berriochoa, A. Cachafeiro, and F. Marcellán. A new numerical quadrature formula on the unit circle. Numerical Algorithms, 44(4):391–401, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=4&spage=391.

Branquinho:2010:MIM

[BCM10] Amílcar Branquinho, Luis Cotrim, and Ana Foulquié Moreno. Matrix interpretation of multiple orthogonality. Numerical Algorithms, 55(1):19–37, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=1&spage=19.

Behl:2016:OFO

[BCM16] Ramandeep Behl, Alicia Cordero, and Sandile S. Motsa. An optimal fourth-order family of methods for multiple roots and its dynamics. Numerical Algorithms, 71(4):775–796, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0023-5.

Beccari:2019:DDN

[BCM19] Carolina Vittoria Beccari, Giulio Casciola, and Marie-Laurence Mazure. Design or not design? A numerical characterisation for piecewise Chebyshevian splines. Numerical Algorithms, 81(1):1–31, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Behl:2018:EOF

[BCMT18] Ramandeep Behl, Alicia Cordero, Sandile S. Motsa, and Juan R. Torregrosa. An eighth-order family of optimal multiple root finders and its dynamics. Numerical Algorithms, 77(4): 1249–1272, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Benzi:2006:P

[BCN06] Michele Benzi, Ljiljana Cvetković, and Michael Neumann. Preface. Numerical Algorithms, 42(3–4):205–206, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=42&issue=3&
spage=205.

Bauschke:2016:ORL

[BCN^{+}16] Heinz H. Bauschke, J. Y. Bello Cruz, Tran T. A. Nghia, Hung M. Pha, and Xianfu Wang. Optimal rates of linear convergence of relaxed alternating projections and generalized Douglas–Rachford methods for two subspaces. Numerical Algorithms, 73(1):33–76, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0085-4.

Behling:2018:CDR

[BCS18] Roger Behling, José Yunier Bello Cruz, and Luiz-Rafael Santos. Circumcentering the Douglas–Rachford method. Numerical Algorithms, 78(3):759–776, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Babajee:2014:ITS

[BCST14] Diyashvir K. R. Babajee, Alicia Cordero, Fazlollah Soleymani, and Juan R. Torregrosa. On improved three-step schemes with high efficiency index and their dynamics. Numerical Algorithms, 65(1):153–169, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9699-6.

Betts:2015:LIM

[BCT15] John T. Betts, Stephen L. Campbell, and Karmethia Thompson. Lobatto IIIA methods, direct transcription, and DAEs with delays. Numerical Algorithms, 69(2):291–300, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9896-y.

Bini:2003:STL

[BCV03] D. A. Bini, G. Codevico, and M. Van Barel. Solving Toeplitz least squares problems by means of Newton’s iteration. Numerical Algorithms, 33(1–4):93–103, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/8/abstract.htm;   http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/8/fulltext.pdf.

Bai:2013:ABD

[BCW13] Zhong-Zhi Bai, Fang Chen, and Zeng-Qi Wang. Additive block diagonal preconditioning for block two-by-two linear systems of skew-Hamiltonian coefficient matrices. Numerical Algorithms, 62(4):655–675, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9696-9.

Baszenski:1995:CAB

[BD95] Günter Baszenski and Franz-Jürgen Delvos. Computational aspects of Boolean cubature. Numerical Algorithms, 10(1–2): 1–11, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Butcher:1998:DDE

[BD98] J. C. Butcher and M. T. Diamantakis. DESIRE: diagonally extended singly implicit Runge–Kutta effective order methods. Numerical Algorithms, 17(1–2):121–145, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/31/9/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
31/9/fulltext.pdf.

Borges:2000:PVF

[BD00] Leonardo Borges and Prabir Daripa. A parallel version of a fast algorithm for singular integral transforms. Numerical Algorithms, 23(1):71–96, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
24/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/24/4/fulltext.pdf.

Badea:2002:FAT

[BD02] Lori Badea and Prabir Daripa. A fast algorithm for twodimensional elliptic problems. Numerical Algorithms, 30 (3–4):199–239, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

42/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/42/1/fulltext.pdf.

Badea:2003:FME

[BD03] Lori Badea and Prabir Daripa. On a Fourier method of embedding domains using an optimal distributed control. Numerical Algorithms, 32(2–4):261–273, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
45/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/45/7/fulltext.pdf.

Badea:2004:DEM

[BD04a] Lori Badea and Prabir Daripa. A domain embedding method using the optimal distributed control and a fast algorithm. Numerical Algorithms, 36(2):95–112, June 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/54/A/1/abstract.htm.

Boldo:2004:STQ

[BD04b] Sylvie Boldo and Marc Daumas. A simple test qualifying the accuracy of Horner’s rule for polynomials. Numerical Algorithms, 37(1–4):45–60, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/35/abstract.htm.

Bai:2006:MDN

[BD06] Zhong-Zhi Bai and Jun-Liang Dong. A modified damped Newton method for linear complementarity problems. Numerical Algorithms, 42(3–4):207–228, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=3&spage=207.

Babajee:2009:SAE

[BD09] D. K. R. Babajee and M. Z. Dauhoo. Spectral analysis of the errors of some families of multi-step Newton-like methods. Numerical Algorithms, 52(1):25–46, September 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=1&spage=25.

Babajee:2010:CSA

[BD10] Diyashvir Kreetee Rajiv Babajee and M. Z. Dauhoo. Convergence and spectral analysis of the Frontini–Sormani family of multipoint third order methods from quadrature rule. Numerical Algorithms, 53(4):467–484, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=4&spage=467.

Bastani:2017:AAS

[BD17a] Ali Foroush Bastani and Davood Damircheli. An adaptive algorithm for solving stochastic multi-point boundary value problems. Numerical Algorithms, 74(4):1119–1143, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0189-5.

Bertaccini:2017:SMC

[BD17b] Daniele Bertaccini and Fabio Durastante. Solving mixed classical and fractional partial differential equations using shortmemory principle and approximate inverses. Numerical Algorithms, 74(4):1061–1082, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0186-8.

Bihun:2020:OSC

[BD20] Oksana Bihun and Kathy Driver. Orthogonal sequences constructed from quasi-orthogonal ultraspherical polynomials. Numerical Algorithms, 85(2):503–522, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00824-5.

Bultheel:2025:PQF

[BD25] Adhemar Bultheel and Carlos Díaz Mendoza. Positive quadrature formulas on an arc of the unit circle. Numerical Algorithms, 100(2):769–802, October 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01953-2.

Buccini:2020:GFA

[BDD20] Alessandro Buccini, Pietro Dell’Acqua, and Marco Donatelli. A general framework for ADMM acceleration. Numerical Algorithms, 85(3):829–848, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00839-
y.

Burman:2025:EDT

[BDF^{+}25] Erik Burman, Rebecca Durst, Miguel A. Fernández, Johnny Guzmán, and Sijing Liu. Estimates of discrete time derivatives for the parabolic-parabolic Robin–Robin coupling method. Numerical Algorithms, 99(2):921–948, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01902-z.

Besana:2013:CDA

[BDH^{+}13] Gian Mario Besana, Sandra Di Rocco, Jonathan D. Hauenstein, Andrew J. Sommese, and Charles W. Wampler. Cell decomposition of almost smooth real algebraic surfaces. Numerical Algorithms, 63(4):645–678, August 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9646-y.

Barrera:2018:TNB

[BDIR18] D. Barrera, C. Dagnino, M. J. Ibáñez, and S. Remogna. Trivariate near-best blending spline quasi-interpolation operators. Numerical Algorithms, 78(1):217–241, May 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beardon:2011:ZPE

[BDJ11] Alan Beardon, Kathy Driver, and Kerstin Jordaan. Zeros of polynomials embedded in an orthogonal sequence. Numerical Algorithms, 57(3):399–403, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=3&spage=399.

Baldeaux:2012:ECW

[BDL^{+}12] Jan Baldeaux, Josef Dick, Gunther Leobacher, Dirk Nuyens, and Friedrich Pillichshammer. Efficient calculation of the worst-case error and (fast) component-by-component construction of higher order polynomial lattice rules. Numerical Algorithms, 59(3):403–431, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=403.

Baharev:2017:RAF

[BDN17] Ali Baharev, Ferenc Domes, and Arnold Neumaier. A robust approach for finding all well-separated solutions of sparse systems of nonlinear equations. Numerical Algorithms, 76(1):163– 189, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-016-0249-x.pdf.

Buccini:2025:FIM

[BDR25] Alessandro Buccini, Marco Donatelli, and Lothar Reichel. Flexible iterative methods for linear systems of equations with multiple right-hand sides. Numerical Algorithms, 100(4): 1477–1505, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02007-x.

Bokhari:2000:BIS

[BDS00] M. A. Bokhari, H. P. Dikshit, and A. Sharma. Birkhoff interpolation on some perturbed roots of unity: revisited. Numerical Algorithms, 25(1–4):47–62, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/25/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/25/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Barakitis:2025:BSA

[BDS25] Nikos Barakitis, Marco Donatelli, and Rosita Luisa Sormani. Blocking structures, approximation, and preconditioning. Numerical Algorithms, 100(4):1893–1927, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02157-y.

Baxter:2018:MDD

[BDV18] Mathew Baxter, Mangalagama Dewasurendra, and Kuppalapalle Vajravelu. A method of directly defining the inverse mapping for solutions of coupled systems of nonlinear differential equations. Numerical Algorithms, 77(4):1199–1211, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bristeau:1998:ACG

[BE98] M. O. Bristeau and J. Erhel. Augmented conjugate gradient. Application in an iterative process for the solution of scattering problems. Numerical Algorithms, 18(1):71–90, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/13/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
13/5/fulltext.pdf.

Brieu:2003:CNI

[BE03] Mathias Brieu and Jocelyne Erhel. On the convergence of a non-incremental homogenization method for nonlinear elastic composite materials. Numerical Algorithms, 32(2–4):141–161, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
45/2/fulltext.pdf.

Bergstrom:2017:RRS

[BE17] Per Bergström and Ove Edlund. Robust registration of surfaces using a refined iterative closest point algorithm with a trust region approach. Numerical Algorithms, 74(3):755–779, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0170-3.

Bojari:2020:TFS

[BE20] S. Bojari and M. R. Eslahchi. Two families of scaled three-term conjugate gradient methods with sufficient descent property for nonconvex optimization. Numerical Algorithms, 83(3):901– 933, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beauzamy:1996:RSK

[Bea96] Bernard Beauzamy. Reconstructing a signal from the knowledge of the norms of its multiples. Numerical Algorithms, 12 (3–4):409–414, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beardmore:1998:SBP

[Bea98] R. E. Beardmore. Stability and bifurcation properties of index-1 DAEs. Numerical Algorithms, 19(1–4):43–53, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/16/14/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/16/

14/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Beckermann:1996:SCF

[Bec96] Bernhard Beckermann. The stable computation of formal orthogonal polynomials. Numerical Algorithms, 11(1–4):1–23, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Boito:2025:NSN

[BEG25] Paola Boito, Yuli Eidelman, and Luca Gemignani. Numerical solution of nonclassical boundary value problems. Numerical Algorithms, 100(2):745–768, October 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01946-1.

Boley:1991:NLF

[BEGG91] Daniel L. Boley, Sylvan Elhay, Gene H. Golub, and Martin H. Gutknecht. Nonsymmetric Lanczos and finding orthogonal polynomials associated with indefinite weights. Numerical Algorithms, 1(1):21–43, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bouhamidi:2024:CSM

[BEH24] Abderrahman Bouhamidi, Lakhdar Elbouyahyaoui, and Mohammed Heyouni. The constant solution method for solving large-scale differential Sylvester matrix equations with time invariant coefficients. Numerical Algorithms, 96(1):449–488, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01653-3.

Bentbib:2020:KSP

[BEHS20] A. H. Bentbib, S. El-Halouy, and El M. Sadek. Krylov subspace projection method for Sylvester tensor equation with low rank right-hand side. Numerical Algorithms, 84(4):1411–1430, August 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00874-0.

Bentbib:2020:ENG

[BEJ20] A. H. Bentbib, M. El Ghomari, and K. Jbilou. Extended nonsymmetric global Lanczos method for matrix function approximation. Numerical Algorithms, 84(4):1459–1479, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00896-8.

Bentbib:2023:GGK

[BEJR23] A. H. Bentbib, M. El Ghomari, K. Jbilou, and L. Reichel. The global Golub–Kahan method and Gauss quadrature for tensor function approximation. Numerical Algorithms, 92(1):5–34, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01392-x.

Beik:2021:TEM

[BEJS21] F. P. A. Beik, A. El Ichi, K. Jbilou, and R. Sadaka. Tensor extrapolation methods with applications. Numerical Algorithms, 87(4):1421–1444, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01013-5.

Bellido:1994:CIF

[Bel94] Anne-Mercedes Bellido. Construction of iteration functions for the simultaneous computation of the solutions of equations and algebraic systems. Numerical Algorithms, 6(3–4): 317–351, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Belghith:1999:SNA

[Bel99] Safya Belghith. Symbolic and numerical analysis for studying complex nonlinear behavior. Numerical Algorithms, 20(1):51–61, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
17/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/17/1/fulltext.pdf.

Bellalij:2003:AHT

[Bel03] Mohammed Bellalij. Analysis of Henrici’s transformation for singular problems. Numerical Algorithms, 33(1–4):65–82, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
46/6/fulltext.pdf.

Belhaj:2008:FMB

[Bel08] Skander Belhaj. A fast method to block-diagonalize a Hankel matrix. Numerical Algorithms, 47(1):15–34, January 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=1&
spage=15.

Bonino:2023:DDR

[BEL23] Brigida Bonino, Claudio Estatico, and Marta Lazzaretti. Dual descent regularization algorithms in variable exponent Lebesgue spaces for imaging. Numerical Algorithms, 92(1): 149–182, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01458-w.

Baranger:1999:EEC

[BEM99] J. Baranger, D. Esslaoui, and A. Machmoum. Error estimate for convection problem with characteristics method. Numerical Algorithms, 21(1–4):49–56, December 1999. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/20/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/4/fulltext.pdf.

Benazzouz:1997:QST

[Ben97] A. Benazzouz. Quasilinear sequence transformations. Numerical Algorithms, 15(3–4):275–285, January 1997. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Benbachir:1999:SRP

[Ben99a] Saâd Benbachir. Symbolic research of the periodic solutions of the nonintegrable Hamiltonian system of Ollongren. Numerical Algorithms, 21(1–4):57–77, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/25/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
25/fulltext.pdf.

Bentbib:1999:CDM

[Ben99b] A. H. Bentbib. Conjugate directions method for solving interval linear systems. Numerical Algorithms, 21(1– 4):79–86, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/9/fulltext.pdf.

Benner:2014:FVN

[BEQOR14] Peter Benner, Pablo Ezzatti, Enrique S. Quintana-Ortí, and Alfredo Remón. A factored variant of the Newton iteration for the solution of algebraic Riccati equations via the matrix sign function. Numerical Algorithms, 66(2):363–377, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9739-2.

Berrut:1993:CFC

[Ber93] Jean-Paul Berrut. A closed formula for the čebyšev barycentric weights of optimal approximation in H^{2}. Numerical Algorithms, 5(1–4):155–163, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Berrut:2000:MDL

[Ber00] Jean-Paul Berrut. A matrix for determining lower complexity barycentric representations of rational interpolants. Numerical Algorithms, 24(1–2):17–29, December 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/27/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/27/3/fulltext.pdf. Computa-

tional methods from rational approximation theory (Wilrijk, 1999).

Berrut:2007:FEF

[Ber07] Jean-Paul Berrut. A formula for the error of finite sincinterpolation over a finite interval. Numerical Algorithms, 45 (1–4):369–374, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-007-9074-6.

Berzins:2010:NDB

[Ber10] Martin Berzins. Nonlinear data-bounded polynomial approximations and their applications in ENO methods. Numerical Algorithms, 55(2–3):171–189, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=171.

Berrut:2011:FEF

[Ber11] Jean-Paul Berrut. A formula for the error of finite sinc interpolation with an even number of nodes. Numerical Algorithms, 56(1):143–157, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=143.

Berlinet:2014:GAP

[Ber14] Alain F. Berlinet. Geometric approach to the parallel sum of vectors and application to the vector ε-algorithm. Numerical Algorithms, 65(4):783–807, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9714-y.

Bogoya:2022:FTE

[BESC22] Manuel Bogoya, Sven-Erik Ekström, and Stefano Serra-Capizzano. Fast Toeplitz eigenvalue computations, joining interpolation-extrapolation matrix-less algorithms and simpleloop theory. Numerical Algorithms, 91(4):1653–1676, December 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01318-7.

Butler:1993:ACD

[BF93] B. P. Butler and A. B. Forbes. An algorithm for combining data sets having different frames of reference. Numerical Algorithms, 5(1–4):467–477, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bini:1999:PES

[BF99a] Dario Andrea Bini and Giuseppe Fiorentino. On the parallel evaluation of a sparse polynomial at a point. Numerical Algorithms, 20(4):323–329, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
19/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/19/2/fulltext.pdf.

Broyden:1999:DCG

[BF99b] C. G. Broyden and P. Foschi. Duality in conjugate gradient methods. Numerical Algorithms, 22(2):113–128, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/22/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
22/2/fulltext.pdf.

Bini:2000:DAI

[BF00] Dario Andrea Bini and Giuseppe Fiorentino. Design, analysis, and implementation of a multiprecision polynomial rootfinder. Numerical Algorithms, 23(2–3):127–173, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/25/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
25/1/fulltext.pdf.

Birk:2014:DCG

[BF14] Sebastian Birk and Andreas Frommer. A deflated conjugate gradient method for multiple right hand sides and multiple shifts. Numerical Algorithms, 67(3):507–529, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9805-9.

Bialecki:2017:ADI

[BF17] Bernard Bialecki and Ryan I. Fernandes. Alternating direction implicit orthogonal spline collocation on some nonrectangular regions with inconsistent partitions. Numerical Algorithms, 74(4):1083–1100, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0187-7.

Belyaev:2018:MGW

[BF18] Alexander Belyaev and Pierre-Alain Fayolle. On modified Gordon–Wixom interpolation schemes and their applications to nonlinear and exterior domain problems. Numerical Algorithms, 77(3):691–708, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-017-0335-8.
pdf.

Belyaev:2020:ABS

[BF20] Alexander Belyaev and Pierre-Alain Fayolle. An ADMMbased scheme for distance function approximation. Numerical Algorithms, 84(3):983–996, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00789-
5;   http://link.springer.com/content/pdf/10.1007/
s11075-019-00789-5.pdf.

Bini:2003:EFA

[BFGM03] D. A. Bini, G. Fiorentino, L. Gemignani, and B. Meini. Effective fast algorithms for polynomial spectral factorization. Numerical Algorithms, 34(2–4):217–227, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/19/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
19/fulltext.pdf.

Bialecki:2009:MDA

[BFK^{+}09] Bernard Bialecki, Graeme Fairweather, David B. Knudson, D. Abram Lipman, Que N. Nguyen, Weiwei Sun, and Gadalia M. Weinberg. Matrix decomposition algorithms for the finite element Galerkin method with piecewise Hermite cubics. Numerical Algorithms, 52(1):1–23, September 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=1&spage=1.

Bialecki:2011:MDA

[BFK11] Bernard Bialecki, Graeme Fairweather, and Andreas Karageorghis. Matrix decomposition algorithms for elliptic boundary value problems: a survey. Numerical Algorithms, 56(2):253–295, February 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=2&spage=253.

Bialecki:2022:OTS

[BFK22] Bernard Bialecki, Graeme Fairweather, and Andreas Karageorghis. An optimal two-step quadratic spline collocation method for the Dirichlet biharmonic problem. Numerical Algorithms, 91(3):1115–1143, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01294-y.

Bialecki:2020:QSC

[BFKM20] Bernard Bialecki, Graeme Fairweather, Andreas Karageorghis, and Jonathan Maack. A quadratic spline collocation method for the Dirichlet biharmonic problem. Numerical Algorithms, 83(1):165–199, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Barry:1991:SPD

[BG91a] Phillip J. Barry and Ronald N. Goldman. Shape parameter deletion for Pólya curves. Numerical Algorithms, 1(2):121– 137, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Berry:1991:ELS

[BG91b] Michael Berry and Gene Golub. Estimating the largest singular values of large sparse matrices via modified moments. Numerical Algorithms, 1(4):353–373, November 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Beutel:2003:QIP

[BG03a] Laura Beutel and Heinz H. Gonska. Quantitative inheritance properties for simultaneous approximation by tensor product operators. Numerical Algorithms, 33(1–4):83–92, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
46/7/fulltext.pdf.

Bottcher:2003:ASD

[BG03b] A. Böttcher and S. M. Grudsky. Asymptotic spectra of dense Toeplitz matrices are unstable. Numerical Algorithms, 33(1–4):105–112, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/9/fulltext.pdf.

Bardsley:2011:RPS

[BG11] Johnathan M. Bardsley and John Goldes. Regularization parameter selection and an efficient algorithm for total variation-regularized positron emission tomography. Numerical Algorithms, 57(2):255–271, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=2&spage=255.

Borm:2013:LRA

[BG13] Steffen Börm and Jessica Gördes. Low-rank approximation of integral operators by using the Green formula and quadrature. Numerical Algorithms, 64(3):567–592, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9679-2.

Bitterlich:2024:SIM

[BG24] Sandy Bitterlich and Sorin-Mihai Grad. Stochastic incremental mirror descent algorithms with Nesterov smoothing. Numerical Algorithms, 95(1):351–382, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01574-1.

Brugnano:2025:NSF

[BGI25] Luigi Brugnano, Gianmarco Gurioli, and Felice Iavernaro. Numerical solution of FDE–IVPs by using fractional HBVMs: the fhbvm code. Numerical Algorithms, 99(1):463–489, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01884-y.

Brownlee:2007:SSF

[BGL07] R. Brownlee, A. N. Gorban, and J. Levesley. Stable simulation of fluid flow with high-Reynolds number using Ehrenfests’ steps. Numerical Algorithms, 45(1–4):389–408, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-007-9087-1.

Bao:2026:SBM

[BGL26] Wendi Bao, Zhiwei Guo, and Weiguo Li. On stochastic block methods for solving nonlinear equations. Numerical Algorithms, 101(2):915–931, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02027-7.

Bajpai:2023:PEE

[BGR23a] Saumya Bajpai, Deepjyoti Goswami, and Kallol Ray. A priori error estimates of a discontinuous Galerkin method for the Navier–Stokes equations. Numerical Algorithms, 94(2): 937–1002, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01525-w.

Barshad:2023:GMS

[BGR23b] Kay Barshad, Aviv Gibali, and Simeon Reich. The generalized modular string averaging procedure and its applications to iterative methods for solving various nonlinear operator theory problems. Numerical Algorithms, 94(4):1797–1818, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01555-4.

Buccini:2025:PIT

[BGR25] Alessandro Buccini, Silvia Gazzola, and Lothar Reichel. Projected iterated Tikhonov in general form with adaptive choice of the regularization parameter. Numerical Algorithms, 100(4): 1617–1637, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02072-2.

Bayon:2009:AAC

[BGRS09] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. An application of the algorithm of the cyclic coordinate descent in multidimensional optimization problems with constrained speed. Numerical Algorithms, 52(2):129–149, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=2&spage=129.

Bayon:2012:CCD

[BGRS12] Luis Bayón, Jose M. Grau, Maria M. Ruiz, and Pedro M. Suárez. Cyclic coordinate descent in hydrothermal nonsmooth problems. Numerical Algorithms, 59(2):227–247, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=2&
spage=227.

Bernigaud:2024:NLC

[BGS24a] Antoine Bernigaud, Serge Gratton, and Ehouarn Simon. A non-linear conjugate gradient in dual space for L_{p}-norm regularized non-linear least squares with application in data assimilation. Numerical Algorithms, 95(1):471–497, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01578-x.

Berrone:2024:PCO

[BGS24b] Stefano Berrone, Denise Grappein, and Stefano Scialò. A PDEconstrained optimization method for 3D–1D coupled problems with discontinuous solutions. Numerical Algorithms, 95 (1):499–526, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01579-w.

Bultheel:1992:COR

[BGVHN92a] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. The computation of orthogonal rational functions and their interpolating properties. Numerical Algorithms, 2(1):85–114, February 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bultheel:1992:FTO

[BGVHN92b] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. A Favard theorem for orthogonal rational functions on the unit circle. Numerical Algorithms, 3(1–4): 81–89, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bultheel:1992:MPA

[BGVHN92c] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. A moment problem associated to rational Szegő functions. Numerical Algorithms, 3(1–4):91–104, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bultheel:1992:ORF

[BGVHN92d] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. Orthogonal rational functions and quadrature on the unit circle. Numerical Algorithms, 3(1–4):105–116, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bultheel:1996:CMP

[BGVHN96a] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. On the convergence of multipoint Padé-type approximants and quadrature formulas associated with the unit circle. Numerical Algorithms, 13(3–4):321–344, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bultheel:1996:ORF

[BGVHN96b] A. Bultheel, P. González-Vera, E. Hendriksen, and O. Njåstad. Orthogonal rational functions and modified approximants. Numerical Algorithms, 11(1–4):57–69, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Burkardt:2020:HPC

[BGZ20] John Burkardt, Max Gunzburger, and Wenju Zhao. Highprecision computation of the weak Galerkin methods for the fourth-order problem. Numerical Algorithms, 84(1):181–205, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bessis:1992:MPF

[BH92a] Daniel Bessis and Carlos R. Handy. Moment problem formulation of the Schrödinger equation. Numerical Algorithms, 3 (1–4):1–16, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bischof:1992:BAC

[BH92b] Christian H. Bischof and Per Christian Hansen. A block algorithm for computing rank-revealing QR factorizations. Numerical Algorithms, 2(3–4):371–391, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Borm:2001:ATP

[BH01] S. Börm and R. Hiptmair. Analysis of tensor product multigrid. Numerical Algorithms, 26(3):219–234, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/2/2/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/2/
2/fulltext.pdf.

Butcher:2002:SNM

[BH02] J. C. Butcher and A. D. Heard. Stability of numerical methods for ordinary differential equations. Numerical Algorithms, 31(1–4):59–73, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/

11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/11/fulltext.pdf.

Butcher:2005:SDM

[BH05] J. C. Butcher and G. Hojjati. Second derivative methods with RK stability. Numerical Algorithms, 40(4):415–429, December 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=4&
spage=415.

Butcher:2009:ESG

[BH09] J. C. Butcher and L. L. Hewitt. The existence of symplectic general linear methods. Numerical Algorithms, 51(1):77–84, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=51&

issue=1&spage=77. Tributes to Gene H. Golub, Part I.

Bereziat:2011:SIP

[BH11] Dominique Béréziat and Isabelle Herlin. Solving ill-posed image processing problems using data assimilation. Numerical Algorithms, 56(2):219–252, February 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=2&spage=219.

Benner:2017:SPI

[BH17] Peter Benner and Mohammad-Sahadet Hossain. Structure preserving iterative methods for periodic projected Lyapunov equations and their application in model reduction of periodic descriptor systems. Numerical Algorithms, 76(4):881–904, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Balani:2022:GAC

[BH22] Fariba Bakrani Balani and Masoud Hajarian. On the generalized AOR and CG iteration methods for a class of block two-by-two linear systems. Numerical Algorithms, 90(2):669– 685, June 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01203-9.

Brysiewicz:2024:MCT

[BHH24] Taylor Brysiewicz, Jonathan D. Hauenstein, and Caroline Hills. Max-convolution through numerics and tropical geometry. Numerical Algorithms, 96(2):845–877, June 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01668-w.

Bonan-Hamada:1992:POL

[BHJTM92] Catherine M. Bonan-Hamada, William B. Jones, W. J. Thron, and Arne Magnus. Para-orthogonal Laurent polynomials and associated sequences of rational functions. Numerical Algorithms, 3(1–4):67–74, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bai:2021:NPA

[BHLZ21] Xueli Bai, Hongjin He, Chen Ling, and Guanglu Zhou. A nonnegativity preserving algorithm for multilinear systems with nonsingular M-tensors. Numerical Algorithms, 87(3):1301– 1320, July 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01008-2.

Bini:2005:AMR

[BHM05] Dario A. Bini, Nicholas J. Higham, and Beatrice Meini. Algorithms for the matrix pth root. Numerical Algorithms, 39(4): 349–378, August 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bates:2016:SGT

[BHNS16] Daniel J. Bates, Jonathan D. Hauenstein, Matthew E. Niemerg, and Frank Sottile. Software for the Gale transform of fewnomial systems and a Descartes rule for fewnomials. Numerical Algorithms, 73(1):281–304, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0095-2.

Bhrawy:2016:JSC

[Bhr16] A. H. Bhrawy. A Jacobi spectral collocation method for solving multi-dimensional nonlinear fractional sub-diffusion equations. Numerical Algorithms, 73(1):91–113, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0087-2.

Bates:2011:EPT

[BHS11] Daniel J. Bates, Jonathan D. Hauenstein, and Andrew J. Sommese. Efficient path tracking methods. Numerical Algorithms, 58(4):451–459, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=451.

Benner:2014:LRI

[BHS14] Peter Benner, Mohammad-Sahadet Hossain, and Tatjana Stykel. Low-rank iterative methods for periodic projected Lyapunov equations and their application in model reduction of periodic descriptor systems. Numerical Algorithms, 67(3):669– 690, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9816-6.

Boltaev:2017:COQ

[BHS17] N. D. Boltaev, A. R. Hayotov, and Kh. M. Shadimetov. Construction of optimal quadrature formulas for Fourier coefficients in Sobolev space L_{2}^{(m)}(0, 1). Numerical Algorithms, 74(2):307–336, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0150-7.

Boman:2023:OSB

[BHS23] Erik G. Boman, Andrew J. Higgins, and Daniel B. Szyld. Optimal size of the block in block GMRES on GPUs: computational model and experiments. Numerical Algorithms, 92 (1):119–147, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01439-z.

Buong:2016:NEI

[BHT16] Nguyen Buong, Nguyen Song Ha, and Nguyen Thi Thu Thuy. A new explicit iteration method for a class of variational inequalities. Numerical Algorithms, 72(2):467–481, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0056-9.

Benner:2023:LRS

[BHW23] Peter Benner, Jan Heiland, and Steffen W. R. Werner. A low-rank solution method for Riccati equations with indefinite quadratic terms. Numerical Algorithms, 92(2):1083–1103, February 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01331-w.

Ben-Israel:2000:LIN

[BI00] Adi Ben-Israel. A local inverse for nonlinear mappings. Numerical Algorithms, 25(1–4):37–46, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp007.kluweronline.com/content/getfile/
5058/3/24/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/24/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Butcher:2014:SEO

[BI14] John C. Butcher and Gulshad Imran. Symplectic effective order methods. Numerical Algorithms, 65(3):499–517, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9789-5.

Bialecki:1994:CRF

[Bia94] Bernard Bialecki. Cyclic reduction and FACR methods for piecewise Hermite bicubic orthogonal spline collocation. Numerical Algorithms, 8(2–4):167–184, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bialecki:2012:FOF

[Bia12] Bernard Bialecki. A fourth order finite difference method for the Dirichlet biharmonic problem. Numerical Algorithms, 61(3):351–375, November 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=351.

Bica:2011:NMS

[Bic11] Alexandru Mihai Bica. The numerical method of successive interpolations for Fredholm functional integral equations. Numerical Algorithms, 58(3):351–377, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=3&spage=351.

Bica:2024:MOC

[Bic24] Alexandru Mihai Bica. On the maximal order of convergence of Green’s function method for solving two-point boundary value problems with deviating argument. Numerical Algorithms, 95 (2):943–978, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01595-w.

Bini:2023:CES

[BIM^{+}23] D. A. Bini, B. Iannazzo, B. Meini, J. Meng, and L. Robol. Computing eigenvalues of semi-infinite quasi-Toeplitz matrices. Numerical Algorithms, 92(1):89–118, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01381-0.

Brugnano:2019:SAS

[BIMR19] Luigi Brugnano, Felice Iavernaro, Juan I. Montijano, and Luis Rández. Spectrally accurate space-time solution of Hamiltonian PDEs. Numerical Algorithms, 81(4):1183–1202, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bini:1996:NCP

[Bin96] Dario Andrea Bini. Numerical computation of polynomial zeros by means of Aberth’s method. Numerical Algorithms, 13 (3–4):179–200, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bartoszewski:1998:CTS

[BJ98] Z. Bartoszewski and Z. Jackiewicz. Construction of two-step Runge–Kutta methods of high order for ordinary differential equations. Numerical Algorithms, 18(1):51–70, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/13/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
13/4/fulltext.pdf.

Butcher:2002:EEN

[BJ02] J. C. Butcher and Z. Jackiewicz. Error estimation for Nordsieck methods. Numerical Algorithms, 31(1–4):75–85, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/18/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/43/
18/fulltext.pdf.

Butcher:2004:CGL

[BJ04] J. C. Butcher and Z. Jackiewicz. Construction of general linear methods with Runge–Kutta stability properties. Numerical Algorithms, 36(1):53–72, May 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/52/A/4/abstract.htm.

Burrage:2024:AHJ

[BJK^{+}24] Kevin Burrage, Zdzisław Jackiewicz, Bernd Krauskopf, Yuto Miyatake, Helmut Podhaisky, and Mayya Tokman. AN-ODE 2023 in honour of John Butcher’s 90th birthday. Numerical Algorithms, 96(3):1019–1020, July 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01849-1.

Beik:2020:GKB

[BJNKR20] Fatemeh P. A. Beik, Khalide Jbilou, Mehdi Najafi-Kalyani, and Lothar Reichel. Golub–Kahan bidiagonalization for illconditioned tensor equations with applications. Numerical Algorithms, 84(4):1535–1563, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00911-
y.

Brugnano:2019:P

[BJP^{+}19] Luigi Brugnano, Zdzisław Jackiewicz, Helmut Podhaisky, Yajuan Sun, and Caren Tischendorf. Preface. Numerical Algorithms, 81(4):1181–1182, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-019-
00766-y.pdf.

Butcher:2010:P

[BJPW10] John C. Butcher, Zdzisław Jackiewicz, Helmut Podhaisky, and Rüdiger Weiner. Preface. Numerical Algorithms, 53(2–3):149– 151, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
53&issue=2&spage=149.

Bentbib:2024:MMT

[BJT24] Abdeslem Hafid Bentbib, Khalide Jbilou, and Ridwane Tahiri. N-mode minimal tensor extrapolation methods. Numerical Algorithms, 95(2):665–691, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01585-
y.

Bentbib:2026:EME

[BJT26] Abdeslem Hafid Bentbib, Khalide Jbilou, and Ridwane Tahiri. Einstein-multidimensional extrapolation methods. Numerical Algorithms, 101(2):1211–1240, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02039-3.

Beyn:1997:NAH

[BK97] W.-J. Beyn and J.-M. Kleinkauf. Numerical approximation of homoclinic chaos. Numerical Algorithms, 14 (1–3):25–53, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/3/fulltext.pdf. Dynamical

numerical analysis (Atlanta, GA, 1995).

Bialecki:2004:LGS

[BK04] Bernard Bialecki and Andreas Karageorghis. Legendre Gauss spectral collocation for the Helmholtz equation on a rectangle. Numerical Algorithms, 36(3):203–227, July 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/57/A/2/abstract.htm.

Burstedde:2008:WBN

[BK08] Carsten Burstedde and Angela Kunoth. A wavelet-based nested iteration-inexact conjugate gradient algorithm for adaptively solving elliptic PDEs. Numerical Algorithms, 48(1– 3):161–188, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=48&issue=1&spage=161.

Bialecki:2013:MNC

[BK13] Bernard Bialecki and Andreas Karageorghis. Modified nodal cubic spline collocation for three-dimensional variable coefficient second order partial differential equations. Numerical Algorithms, 64(2):349–383, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9669-4.

Babaie-Kafaki:2016:MSP

[BK16a] Saman Babaie-Kafaki. A modified scaling parameter for the memoryless BFGS updating formula. Numerical Algorithms, 72(2):425–433, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0053-z.

Bujanovic:2016:BAC

[BK16b] Zvonimir Bujanović and Daniel Kressner. A block algorithm for computing antitriangular factorizations of symmetric matrices. Numerical Algorithms, 71(1):41–57, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9983-8.

Abda:2018:MBD

[BK18a] A. Ben Abda and F. Khayat. Missing boundary data recovery using Nash games: the Stokes system. Numerical Algorithms, 78(3):777–803, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bialecki:2018:LSQ

[BK18b] Bernard Bialecki and Andreas Karageorghis. A Legendre spectral quadrature Galerkin method for the Cauchy–Navier equations of elasticity with variable coefficients. Numerical Algorithms, 77(2):491–516, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-017-0325-x.

Babaie-Kafaki:2019:TPS

[BKA19] Saman Babaie-Kafaki and Zohre Aminifard. Two-parameter scaled memoryless BFGS methods with a nonmonotone choice for the initial step length. Numerical Algorithms, 82(4):1345– 1357, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Babaie-Kafaki:2025:SRM

[BKDA25] Saman Babaie-Kafaki, Fatemeh Dargahi, and Zohre Aminifard. On solving a revised model of the nonnegative matrix factorization problem by the modified adaptive versions of the Dai–Liao method. Numerical Algorithms, 99(1):505– 519, May 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01886-w.

Bhatt:2020:EHO

[BKF20] H. P. Bhatt, A. Q. M. Khaliq, and K. M. Furati. Efficient high-order compact exponential time differencing method for space-fractional reaction-diffusion systems with nonhomogeneous boundary conditions. Numerical Algorithms, 83(4): 1373–1397, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Babaie-Kafaki:2011:TEH

[BKFMA11] Saman Babaie-Kafaki, Masoud Fatemi, and Nezam Mahdavi-Amiri. Two effective hybrid conjugate gradient algorithms based on modified BFGS updates. Numerical Algorithms, 58(3):315–331, November 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=3&spage=315.

Babaie-Kafaki:2015:HPR

[BKG15] Saman Babaie-Kafaki and Reza Ghanbari. A hybridization of the Polak–Ribière–Polyak and Fletcher–Reeves conjugate gradient methods. Numerical Algorithms, 68(3):481–495, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9856-6.

Buhmiler:2010:PQN

[BKL10] Sandra Buhmiler, Natasa Krejić, and Zorana Luzanin. Practical Quasi–Newton algorithms for singular nonlinear systems. Numerical Algorithms, 55(4):481–502, December 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=4&
spage=481.

Birgin:2003:GCI

[BKM03] Ernesto G. Birgin, Nataša Krejić, and José Mario Martínez. Globally convergent inexact quasi-Newton methods for solving nonlinear systems. Numerical Algorithms, 32(2–4):249–260, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
45/6/fulltext.pdf.

Bini:2025:CEC

[BKM25] Dario A. Bini, Steve Kirkland, and Beatrice Meini. Cut-edge centralities in an undirected graph. Numerical Algorithms, 100(4):1555–1580, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02041-
9.

Barrodale:1993:SSS

[BKPS93] I. Barrodale, R. Kuwahara, R. Poeckert, and D. Skea. Sidescan sonar image processing using thin plate splines and control point matching. Numerical Algorithms, 5(1–4):85–98, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Babaie-Kafaki:2018:TAN

[BKR18] Saman Babaie-Kafaki and Saeed Rezaee. Two accelerated nonmonotone adaptive trust region line search methods. Numerical Algorithms, 78(3):911–928, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Benner:2013:EHC

[BKS13] Peter Benner, Patrick Kürschner, and Jens Saak. Efficient handling of complex shift parameters in the low-rank Cholesky factor ADI method. Numerical Algorithms, 62(2): 225–251, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9569-7/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=2&spage=225-251.

Bujanovic:2023:IRS

[BKS23] Zvonimir Bujanović, Daniel Kressner, and Christian Schröder. Iterative refinement of Schur decompositions. Numerical Algorithms, 92(1):247–267, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01327-
6.

Beckermann:2025:CPL

[BKW25] Bernhard Beckermann, Daniel Kressner, and Heather Wilber. Compression properties for large Toeplitz-like matrices. Numerical Algorithms, 100(4):2041–2067, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02185-8.

Beckermann:1992:UAH

[BL92] Bernhard Beckermann and George Labahn. A uniform approach for Hermite Padé and simultaneous Padé approximants and their matrix-type generalizations. Numerical Algorithms, 3(1–4):45–54, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bosworth:1993:SSA

[BL93] Ken W. Bosworth and Upmanu Lall. An L_{1} smoothing spline algorithm with cross validation. Numerical Algorithms, 5(1–4):

407–417, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Bosworth:1995:LLW

[BL95] Ken W. Bosworth and Upmanu Lall. LOWLAD: a locally weighted L_{1} smoothing spline algorithm with cross validated choice of smoothing parameters. Numerical Algorithms, 9(1– 2):85–106, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Bond:1998:TTR

[BL98] Stephen D. Bond and Benedict J. Leimkuhler. Timetransformations for reversible variable stepsize integration. Numerical Algorithms, 19(1–4):55–71, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/16/19/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/16/

19/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Boese:2004:EZS

[BL04] Fritz G. Boese and Wolfram J. Luther. Enclosure of the zero set of multivariate exponential interval polynomials. Numerical Algorithms, 37(1–4):35–44, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/34/abstract.htm.

Bai:2023:EIO

[BL23] Zhong-Zhi Bai and Kang-Ya Lu. An economic implementation of the optimal rotated block-diagonal preconditioning method. Numerical Algorithms, 93(1):85–101, May 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01404-w.

Blanes:2015:HOS

[Bla15] Sergio Blanes. High order structure preserving explicit methods for solving linear-quadratic optimal control problems. Numerical Algorithms, 69(2):271–290, June 2015.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9894-0.

Bloom:1992:IIR

[BLS92] T. Bloom, D. S. Lubinsky, and H. Stahl. Interpolatory integration rules and orthogonal polynomials with varying weights. Numerical Algorithms, 3(1–4):55–65, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Bozzini:2006:KBS

[BLS06] M. Bozzini, L. Lenarduzzi, and R. Schaback. Kernel B-splines and interpolation. Numerical Algorithms, 41(1):1–16, January 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=41&issue=1&
spage=1.

Bao:2009:RGK

[BLW09] Liang Bao, Yiqin Lin, and Yimin Wei. Restarted generalized Krylov subspace methods for solving large-scale polynomial eigenvalue problems. Numerical Algorithms, 50(1):17–32, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=50&
issue=1&spage=17.

Baxter:1994:NEI

[BM94] B. J. C. Baxter and C. A. E. Micchelli. Norm estimates for the l^{2}-inverses of multivariate Toeplitz matrices. Numerical Algorithms, 6(1–2):103–117, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1996:ALF

[BM96] C. Brezinski and P. Maroni. The algebra of linear functionals on polynomials, with applications to Padé approximation. Numerical Algorithms, 11(1–4):25–33, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Bini:1997:ICR

[BM97] Dario Andrea Bini and Beatrice Meini. Improved cyclic reduction for solving queueing problems. Numerical Algorithms, 15(1):57–74, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
30/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/4/fulltext.pdf.

Berrut:2000:RIT

[BM00] Jean-Paul Berrut and Hans D. Mittelmann. Rational interpolation through the optimal attachment of poles to the interpolating polynomial. Numerical Algorithms, 23(4):315–328, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/26/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
26/4/fulltext.pdf.

Butcher:2003:ENF

[BM03] John C. Butcher and Nicolette Moir. Experiments with a new fifth order method. Numerical Algorithms, 33 (1–4):137–151, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
12/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/12/fulltext.pdf.

Bini:2006:CCI

[BM06] Dario A. Bini and Ana Marco. Computing curve intersection by means of simultaneous iterations. Numerical Algorithms, 43(2):151–175, October 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=2&spage=151.

Bini:2009:CRA

[BM09] Dario A. Bini and Beatrice Meini. The cyclic reduction algorithm: from Poisson equation to stochastic processes and beyond. Numerical Algorithms, 51(1):23–60, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=1&

spage=23. Tributes to Gene H. Golub, Part I.

Bergamaschi:2012:BTM

[BM12] Luca Bergamaschi and Ángeles Martínez. Banded target matrices and recursive FSAI for parallel preconditioning. Numerical Algorithms, 61(2):223–241, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=2&spage=223.

Brilleaud:2014:DS

[BM14] Martine Brilleaud and Marie-Laurence Mazure. Design with L-splines. Numerical Algorithms, 65(1):91–124, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9697-8.

Bai:2019:CEH

[BM19] Zhong-Zhi Bai and Cun-Qiang Miao. Computing eigenpairs of Hermitian matrices in perfect Krylov subspaces. Numerical Algorithms, 82(4):1251–1277, December 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Becher:2022:UAV

[BM22] Simon Becher and Gunar Matthies. Unified analysis for variational time discretizations of higher order and higher regularity applied to non-stiff ODEs. Numerical Algorithms, 89(4):1533– 1565, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01164-z.

Bultheel:2023:TPP

[BM23] Adhemar Bultheel and Carlos Díaz Mendoza. Two-point Padé approximation to Herglotz–Riesz transforms. Numerical Algorithms, 92(1):269–299, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01467-9.

Birgin:2024:PIO

[BM24a] E. G. Birgin and J. M. Martínez. A PDE-informed optimization algorithm for river flow predictions. Numerical Algorithms, 96(1):289–304, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01647-
1.

Bucataru:2024:AIA

[BM24b] Mihai Bucataru and Liviu Marin. Accelerated iterative algorithms for the Cauchy problem in steady-state anisotropic heat conduction. Numerical Algorithms, 95(2):605–636, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01583-0.

Bini:2025:CMA

[BM25] Dario A. Bini and Beatrice Meini. On certain matrix algebras related to quasi-Toeplitz matrices. Numerical Algorithms, 100 (2):685–705, October 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01855-3.

Bagherpour:2016:NEV

[BMA16] Negin Bagherpour and Nezam Mahdavi-Amiri. A new error in variables model for solving positive definite linear system using orthogonal matrix decompositions. Numerical Algorithms, 72 (1):211–241, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0042-2.

Bahi:1997:AMM

[BMR97] J. Bahi, J. C. Miellou, and K. Rhofir. Asynchronous multisplitting methods for nonlinear fixed point problems. Numerical Algorithms, 15(3–4):315–345, January 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bini:2019:QTM

[BMR19a] Dario A. Bini, Stefano Massei, and Leonardo Robol. Quasi-Toeplitz matrix arithmetic: a MATLAB toolbox. Numerical Algorithms, 81(2):741–769, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brugnano:2019:ESM

[BMR19b] Luigi Brugnano, Juan I. Montijano, and Luis Rández. On the effectiveness of spectral methods for the numerical solution of multi-frequency highly oscillatory Hamiltonian problems. Numerical Algorithms, 81(1):345–376, May 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Birgin:2021:CON

[BMR21] E. G. Birgin, J. M. Martínez, and A. Ramos. On constrained optimization with nonconvex regularization. Numerical Algorithms, 86(3):1165–1188, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00928-
3.

Bouyghf:2024:UAK

[BMS24] F. Bouyghf, A. Messaoudi, and H. Sadok. A unified approach to Krylov subspace methods for solving linear systems. Numerical Algorithms, 96(1):305–332, May 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01648-0.

Breda:2009:AAE

[BMV09] Dimitri Breda, Stefano Maset, and Rossana Vermiglio. An adaptive algorithm for efficient computation of level curves of surfaces. Numerical Algorithms, 52(4):605–628, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=4&
spage=605.

Bulthel:2025:P

[BMV25] Adhemar Bulthel, Nicola Mastronardi, and Raf Vandebril. Preface. Numerical Algorithms, 100(2):597–598, October 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02219-1.

Brown:2018:ICD

[BN18] David A. Brown and Siva Nadarajah. Inexactly constrained discrete adjoint approach for steepest descent-based optimization algorithms. Numerical Algorithms, 78(3):983–1000, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bansal:2025:NHO

[BN25] Saurabh Bansal and Srinivasan Natesan. A novel higherorder efficient computational method for pricing European and Asian options. Numerical Algorithms, 99(3):1111–1125, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01909-6.

Bates:2016:BMI

[BNN16] Daniel J. Bates, Andrew J. Newell, and Matthew Niemerg. BertiniLab: A MATLAB interface for solving systems of polynomial equations. Numerical Algorithms, 71(1):229–244, January 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0014-6.

Bnouhachem:2021:SAD

[Bno21] Abdellah Bnouhachem. A self-adaptive descent LQP alternating direction method for the structured variational inequalities. Numerical Algorithms, 86(1):303–324, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00890-0.

Butcher:2002:NMS

[BO02] J. C. Butcher and A. E. O’Sullivan. Nordsieck methods with an off-step point. Numerical Algorithms, 31(1– 4):87–101, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
19/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/19/fulltext.pdf.

Beckermann:2003:SSP

[BO03] B. Beckermann and A. Osipov. Some spectral properties of infinite band matrices. Numerical Algorithms, 34 (2–4):173–185, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
16/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/16/fulltext.pdf.

Boglaev:2002:SSP

[Bog02] Igor P. Boglaev. The solution of a singularly perturbed convection–diffusion problem by an iterative domain decomposition method. Numerical Algorithms, 31(1–4):27–46, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/43/9/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
43/9/fulltext.pdf.

Boglaev:2013:UQC

[Bog13] Igor Boglaev. Uniform quadratic convergence of monotone iterates for nonlinear singularly perturbed parabolic problems. Numerical Algorithms, 64(4):607–631, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9682-7.

Bar-On:1998:SST

[BOL98] Ilan Bar-On and Mauro Leoncini. Stable solution of tridiagonal systems. Numerical Algorithms, 18(3–4):361–388, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/15/8/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
15/8/fulltext.pdf.

Bar-On:1998:HPS

[BOP98] Ilan Bar-On and Marcin Paprzycki. High performance solution of the complex symmetric eigenproblem. Numerical Algorithms, 18(2):195–208, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
14/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/14/5/fulltext.pdf.

Buccini:2023:ABP

[BOR23] Alessandro Buccini, Lucas Onisk, and Lothar Reichel. An Arnoldi-based preconditioner for iterated Tikhonov regularization. Numerical Algorithms, 92(1):223–245, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01407-7.

Borm:2026:AMM

[Bör26] Steffen Börm. Adaptive multiplication of H^{∈}-matrices with block-relative error control. Numerical Algorithms, 101(1):65– 96, January 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01996-5.

Bosner:2021:PRF

[Bos21] Nela Bosner. Parallel reduction of four matrices to condensed form for a generalized matrix eigenvalue algorithm. Numerical Algorithms, 86(1):153–178, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00883-
z.

Boutry:2003:AGP

[Bou03] Gregory Boutry. Anti-Gaussian Padé approximants. Numerical Algorithms, 33(1–4):113–122, August 2003. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/46/10/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/10/fulltext.pdf.

Boumenir:2006:PSS

[Bou06] Amin Boumenir. Power series solutions for the KPP equation. Numerical Algorithms, 43(2):177–187, October 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=2&spage=177.

Bourquin:2017:ACH

[Bou17] Raoul Bourquin. Algorithms for the construction of highorder Kronrod rule extensions with application to sparse-grid integration. Numerical Algorithms, 76(3):617–637, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Boumenir:2024:RSD

[Bou24] A. Boumenir. The reconstruction of the space-dependent thermal conductivity coefficient. Numerical Algorithms, 97(2): 801–817, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01724-5.

Bauschke:2021:LCC

[BOW21] Heinz H. Bauschke, Hui Ouyang, and Xianfu Wang. On the linear convergence of circumcentered isometry methods. Numerical Algorithms, 87(1):263–297, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00966-
x.

Boyd:2005:CSN

[Boy05] John P. Boyd. Chebyshev solution of the nearly-singular onedimensional Helmholtz equation and related singular perturbation equations: Multiple scale series and the boundary layer rule-of-thumb. Numerical Algorithms, 38(1–3):197–207, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bozzini:1993:ACS

[BP93] M. T. Bozzini and C. Paracelli. An algorithm for constrained smoothing functions. Numerical Algorithms, 5(1–4):341–351, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Bacchelli:2003:MTM

[BP03] Silvia Bacchelli and Serena Papi. Matrix thresholding for multiwavelet image denoising. Numerical Algorithms, 33 (1–4):41–52, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/4/fulltext.pdf.

Bittens:2019:SFD

[BP19] Sina Bittens and Gerlind Plonka. Sparse fast DCT for vectors with one-block support. Numerical Algorithms, 82(2):663–697, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bochacik:2022:RES

[BP22] Tomasz Bochacik and Paweł Przybyłowicz. On the randomized Euler schemes for ODEs under inexact information. Numerical Algorithms, 91(3):1205–1229, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01299-7.

Bracciali:2023:MOP

[BP23] Cleonice F. Bracciali and Miguel A. Piñar. On multivariate orthogonal polynomials and elementary symmetric functions. Numerical Algorithms, 92(1):183–206, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01434-4.

Baglama:2023:HIR

[BPP23] James Baglama, Vasilije Perović, and Jennifer Picucci. Hybrid iterative refined restarted Lanczos bidiagonalization methods. Numerical Algorithms, 92(2):1183–1212, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01336-5.

Bracciali:2020:QRT

[BPR20] Cleonice F. Bracciali, Junior A. Pereira, and A. Sri Ranga. Quadrature rules from a R_{II} type recurrence relation and associated quadrature rules on the unit circle. Numerical Algorithms, 83(3):1029–1061, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Buccini:2021:LKS

[BPR21] Alessandro Buccini, Mirjeta Pasha, and Lothar Reichel. Linearized Krylov subspace Bregman iteration with nonnegativity constraint. Numerical Algorithms, 87(3):1177–1200, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01004-6;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01004-6.pdf.

Babu:2022:RNA

[BPR22] Gajendra Babu, M. Prithvi, and V. P. Ramesh. A robust numerical algorithm on harmonic mesh for parabolic singularly perturbed convection-diffusion problems with time delay. Numerical Algorithms, 91(2):615–634, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01275-1.

Benner:2023:IKA

[BPS23] Peter Benner, Davide Palitta, and Jens Saak. On an integrated Krylov–ADI solver for large-scale Lyapunov equations. Numerical Algorithms, 92(1):35–63, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01409-5.

Bessis:2013:NCP

[BPV13] Daniel Bessis, Luca Perotti, and Daniel Vrinceanu. Noise in the complex plane: open problems. Numerical Algorithms, 62 (4):559–569, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9640-4.

Bazan:2019:GAE

[BQ19] Fermin S. V. Bazán and J. R. Quiroz. Galerkin approach for estimating boundary data in Poisson equation on annular domain with application to heat transfer coefficient estimation in coiled tubes. Numerical Algorithms, 81(1):79–98, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Benner:1999:SSG

[BQO99] Peter Benner and Enrique S. Quintana-Ortí. Solving stable generalized Lyapunov equations with the matrix sign function. Numerical Algorithms, 20(1):75–100, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/17/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
17/5/fulltext.pdf.

Baglama:2006:RBL

[BR06] James Baglama and Lothar Reichel. Restarted block Lanczos bidiagonalization methods. Numerical Algorithms, 43 (3):251–272, November 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=43&issue=3&spage=251.

Bertelle:2007:AGM

[BR07a] Roberto Bertelle and Maria Rosaria Russo. An approach to the Gummel map by vector extrapolation methods. Numerical Algorithms, 45(1–4):331–343, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-007-9101-7.

Bosner:2007:NUE

[BR07b] Tina Bosner and Mladen Rogina. Non-uniform exponential tension splines. Numerical Algorithms, 46(3):265–294, November 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=46&
issue=3&spage=265.

Bonazzoli:2017:HOF

[BR17] Marcella Bonazzoli and Francesca Rapetti. High-order finite elements in numerical electromagnetism: degrees of freedom and generators in duality. Numerical Algorithms, 74(1):111–136, January 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0141-8.

Buccini:2021:GCV

[BR21] Alessandro Buccini and Lothar Reichel. Generalized cross validation for l^{p}–l^{q} minimization. Numerical Algorithms, 88 (4):1595–1616, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01087-
9;   http://link.springer.com/content/pdf/10.1007/
s11075-021-01087-9.pdf.

Brass:1996:QAB

[Bra96] Helmut Brass. On the quality of algorithms based on spline interpolation. Numerical Algorithms, 13(3–4):159–177, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bratsos:2006:MPC

[Bra06] A. G. Bratsos. A modified predictor–corrector scheme for the two-dimensional sine-Gordon equation. Numerical Algorithms, 43(4):295–308, December 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=4&spage=295.

Bratsos:2007:SON

[Bra07] Athanassios G. Bratsos. A second order numerical scheme for the solution of the one-dimensional Boussinesq equation. Numerical Algorithms, 46(1):45–58, September 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=1&spage=45.

Brezinski:1996:BR

[Bre96] C. Brezinski. Book reviews. Numerical Algorithms, 13(3–4): 399–404, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1997:BRb

[Bre97a] C. Brezinski. Book reviews. Numerical Algorithms, 15 (2):243–246, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
7/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/7/7/fulltext.pdf.

Brezinski:1997:BRa

[Bre97b] Claude Brezinski. Book reviews. Numerical Algorithms, 15 (1):135–137, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
30/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/8/fulltext.pdf.

Brezinski:1997:BRc

[Bre97c] Claude Brezinski. Book reviews. Numerical Algorithms, 15 (3–4):385–388, January 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1997:BRd

[Bre97d] Claude Brezinski. Book reviews. Numerical Algorithms, 16 (3–4):381–382, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
11/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/11/8/fulltext.pdf.

Brezinski:1999:KSM

[Bre99a] C. Brezinski. Krylov subspace methods, biorthogonal polynomials and Padé-type approximants. Numerical Algorithms, 21(1–4):97–107, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/5/fulltext.pdf. Numer-

ical methods for partial differential equations (Marrakech, 1998).

Brezinski:1999:NMP

[Bre99b] C. Brezinski, editor. Numerical methods for partial differential equations, volume 21(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 1999. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ???? Papers from the International Congress held in Marrakech, September 14–18, 1998.

Brezinski:2000:APM

[Bre00a] C. Brezinski. Acceleration procedures for matrix iterative methods. Numerical Algorithms, 25(1–4):63–73, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/9/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/3/

9/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Brezinski:2000:BR

[Bre00b] Claude Brezinski. Book reviews. Numerical Algorithms, 23(2–3):281–284, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
25/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/25/6/fulltext.pdf.

Brezinski:2002:BDM

[Bre02] C. Brezinski. Block descent methods and hybrid procedures for linear systems. Numerical Algorithms, 29(1–3):21–32, March 2002. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp008.lwwonline.
com/content/getfile/5058/38/2/abstract.htm;   http:
//ipsapp008.lwwonline.com/content/getfile/5058/38/
2/fulltext.pdf.

Brezinski:2003:CQN

[Bre03a] C. Brezinski. A classification of quasi-Newton methods. Numerical Algorithms, 33(1–4):123–135, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/11/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
11/fulltext.pdf.

Brezinski:2003:BR

[Bre03b] Claude Brezinski. Book reviews. Numerical Algorithms, 32(1):99–103, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
44/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/44/7/fulltext.pdf.

Brezinski:2003:T

[Bre03c] Claude Brezinski. A tribute. Numerical Algorithms, 33(1–4):3–9, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
46/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/1/fulltext.pdf.

Brezinski:2004:EAF

[Bre04] C. Brezinski. Extrapolation algorithms for filtering series of functions, and treating the Gibbs phenomenon. Numerical Algorithms, 36(4):309–329, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:2006:LWA

[Bre06a] C. Brezinski. The life and work of André Cholesky. Numerical Algorithms, 43(3):279–288, November 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=3&spage=279.

Brezinski:2006:BR

[Bre06b] Claude Brezinski. Book reviews. Numerical Algorithms, 43 (3):289–294, November 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=43&issue=3&spage=289.

Brezinski:2019:RPW

[Bre19] C. Brezinski. Reminiscences of Peter Wynn. Numerical Algorithms, 80(1):5–10, January 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-018-0542-y.
pdf.

Brezinski:2025:P

[Bre25] Claude Brezinski. Preface. Numerical Algorithms, 100(4): 1435–1438, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02125-6.

Buhmiler:2018:FDM

[BRMG18] Sandra Buhmiler, Sanja Rapajić, Slavica Medić, and Tatjana Grbić. Finite-difference method for singular nonlinear systems. Numerical Algorithms, 79(1):65–86, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brownlee:2005:EEI

[Bro05] R. A. Brownlee. Error estimates for interpolation of rough data using the scattered shifts of a radial basis function. Numerical Algorithms, 39(1–3):57–68, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Baglama:2013:ALM

[BRR13] J. Baglama, L. Reichel, and D. Richmond. An augmented LSQR method. Numerical Algorithms, 64(2):263–293, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9665-8.

Brezinski:1991:ABN

[BRS91] C. Brezinski, M. Redivo Zaglia, and H. Sadok. Avoiding breakdown and near-breakdown in Lanczos type algorithms. Numerical Algorithms, 1(3):261–284, September 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See erratum [BRS92].

Brezinski:1992:AAB

[BRS92] C. Brezinski, M. Redivo Zaglia, and H. Sadok. Addendum to: “Avoiding breakdown and near-breakdown in Lanczos type algorithms” [Numer. Algorithms 1 (1991), no. 3, 261–284, MR 92j:65041]. Numerical Algorithms, 2(2):133–136, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See [BRS91].

Brezinski:2008:EEL

[BRS08] C. Brezinski, G. Rodriguez, and S. Seatzu. Error estimates for linear systems with applications to regularization. Numerical Algorithms, 49(1–4):85–104, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=85.

Brezinski:2009:EER

[BRS09] C. Brezinski, G. Rodriguez, and S. Seatzu. Error estimates for the regularization of least squares problems. Numerical Algorithms, 51(1):61–76, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=1&spage=61. Trib-

utes to Gene H. Golub, Part I.

Bujok:2026:NSS

[BRS26] Maksymilian Bujok, Miroslav Rozlozník, and Alicja Smoktunowicz. Numerical stability of the symplectic LL^{T} factorization. Numerical Algorithms, 101(3):2017–2042, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02075-z.

Brysiewicz:2021:SDS

[BRSY21] Taylor Brysiewicz, Jose Israel Rodriguez, Frank Sottile, and Thomas Yahl. Solving decomposable sparse systems. Numerical Algorithms, 88(1):453–474, September 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-01045-x.

Brutman:1993:AGA

[Bru93] L. Brutman. An application of the generalized alternating polynomials to the numerical solution of Fredholm integral equations. Numerical Algorithms, 5(1–4):437–442, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Bi:2011:CMH

[BRW11] Weihong Bi, Hongmin Ren, and Qingbiao Wu. Convergence of the modified Halley’s method for multiple zeros under Hölder continuous derivative. Numerical Algorithms, 58(4):497–512, December 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=497.

Bai:2014:AIU

[BRY14] Zhong-Zhi Bai, Yu-Hong Ran, and Li Yuan. On approximated ILU and UGS preconditioning methods for linearized discretized steady incompressible Navier–Stokes equations. Numerical Algorithms, 65(1):43–68, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9694-y.

Brezinski:1996:LAS

[BRZ96] C. Brezinski and M. Redivo-Zaglia. A look-ahead strategy for the implementation of some old and new extrapolation methods. Numerical Algorithms, 11(1–4):35–55, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Brezinski:1998:TFL

[BRZ98] C. Brezinski and M. Redivo-Zaglia. Transpose-free Lanczostype algorithms for nonsymmetric linear systems. Numerical Algorithms, 17(1–2):67–103, March 1998. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
31/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/31/5/fulltext.pdf.

Brezinski:2017:STA

[BRZ17] Claude Brezinski and Michela Redivo-Zaglia. The simplified topological ϵ-algorithms: software and applications. Numerical Algorithms, 74(4):1237–1260, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0238-0.

Bargetz:2018:CPD

[BRZ18] Christian Bargetz, Simeon Reich, and Rafal Zalas. Convergence properties of dynamic string-averaging projection methods in the presence of perturbations. Numerical Algorithms, 77(1):185–209, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-017-0310-4.
pdf.

Brezinski:2019:GED

[BRZ19] Claude Brezinski and Michela Redivo-Zaglia. The genesis and early developments of Aitken’s process, Shanks’ transformation, the ϵ-algorithm, and related fixed point methods. Numerical Algorithms, 80(1):11–133, January 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:2020:SUR

[BRZ20] Claude Brezinski and Michela Redivo-Zaglia. Some unusual results on extrapolation methods. Numerical Algorithms, 84(4): 1241–1264, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00782-y.

Brezinski:2023:KVA

[BRZS23] Claude Brezinski, Michela Redivo-Zaglia, and Ahmed Salam. On the kernel of vector ϵ-algorithm and related topics. Numerical Algorithms, 92(1):207–221, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01358-z.

Brezinski:1991:ABC

[BS91] Claude Brezinski and Hassane Sadok. Avoiding breakdown in the CGS algorithm. Numerical Algorithms, 1(2):199–206, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1992:SVS

[BS92] C. Brezinski and H. Sadok. Some vector sequence transformations with applications to systems of equations. Numerical Algorithms, 3(1–4):75–80, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Branden:2004:ACF

[BS04] Henrik Brandén and Per Sundqvist. An algorithm for computing fundamental solutions of difference operators. Numerical Algorithms, 36(4):331–343, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brugnano:2014:MIC

[BS14] Luigi Brugnano and Yajuan Sun. Multiple invariants conserving Runge–Kutta type methods for Hamiltonian problems. Numerical Algorithms, 65(3):611–632, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9769-9.

Bansal:2017:PUN

[BS17] Komal Bansal and Kapil K. Sharma. Parameter uniform numerical scheme for time dependent singularly perturbed convection-diffusion-reaction problems with general shift arguments. Numerical Algorithms, 75(1):113–145, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bohmer:2019:MMS

[BS19] Klaus Böhmer and Robert Schaback. A meshfree method for solving the Monge–Ampère equation. Numerical Algorithms, 82(2):539–551, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Baranczuk:2021:TAE

[BS21] Stefan Barańczuk and Bogdan Szydło. Two algorithms for the exchange lemma. Numerical Algorithms, 86(3):1041–1050, March 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00921-w;   http://link.
springer.com/content/pdf/10.1007/s11075-020-00921-
w.pdf.

Baggag:2025:DLT

[BS25] Abdelkader Baggag and Yousef Saad. Deep learning, transformers and graph neural networks: a linear algebra perspective. Numerical Algorithms, 100(4):2095–2134, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02218-2.

Bujok:2023:CSE

[BSB23] Maksymilian Bujok, Alicja Smoktunowicz, and Grzegorz Borowik. On computing the symplectic LL^{T} factorization. Numerical Algorithms, 93(3):1401–1416, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01472-y.

Borhanifar:2017:MFA

[BSF17] A. Borhanifar, S. Shahmorad, and E. Feizi. A matrix formulated algorithm for solving parabolic equations with nonlocal boundary conditions. Numerical Algorithms, 74(4):1203– 1221, April 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0192-x.

Bultheel:2018:PIQ

[BSL18] Adhemar Bultheel and Juan Carlos Santos-León. Properties of interpolatory quadrature with equidistant nodes on the unit circle. Numerical Algorithms, 77(2):327–359, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-017-0318-9.

Burrage:2014:P

[BSSJ14] Kevin Burrage, Chus Sanz-Serna, and Zdzisław Jackiewicz. Preface. Numerical Algorithms, 65(3):375–376, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9833-0;   http://link.springer.com/
content/pdf/10.1007/s11075-014-9833-0.pdf.

Benzi:2014:MPS

[BT14] Michele Benzi and Elena Toscano. Mauro Picone, Sandro Faedo, and the numerical solution of partial differential equations in Italy (1928–1953). Numerical Algorithms, 66(1):105– 145, May 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9727-6.

Berrut:2023:EQE

[BT23] Jean-Paul Berrut and Manfred R. Trummer. Extrapolation quadrature from equispaced samples of functions with jumps. Numerical Algorithms, 92(1):65–88, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01462-0.

Bunger:2018:SWT

[Bün18] Florian Bünger. Shrink wrapping for Taylor models revisited. Numerical Algorithms, 78(4):1001–1017, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bunger:2024:RTE

[Bün24] Florian Bünger. Reducing the truncation error in Taylor model multiplication. Numerical Algorithms, 97(2):819–841, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01725-4.

Buong:2017:IAM

[Buo17] Nguyen Buong. Iterative algorithms for the multiple-sets split feasibility problem in Hilbert spaces. Numerical Algorithms, 76 (3):783–798, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Bibi:2024:NTB

[BuR24] Amna Bibi and Mujeeb ur Rehman. A numerical technique based on multiplicative integration strategy for fractional Darboux problem. Numerical Algorithms, 97(2):617–643, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01718-3.

Butcher:1996:OPP

[But96] J. C. Butcher. Orthogonal polynomials, Padé approximations and A-stability. Numerical Algorithms, 11(1–4):71–78, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Butcher:1998:AMO

[But98] J. C. Butcher. ARK methods up to order five. Numerical Algorithms, 17(3–4):193–221, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
12/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/12/6/fulltext.pdf.

Butcher:2002:SMP

[But02a] J. C. Butcher. The A-stability of methods with Padé and generalized Padé stability functions. Numerical Algorithms, 31(1–4):47–58, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/10/fulltext.pdf.

Butcher:2002:P

[But02b]  J. C. Butcher.  Preface.  Numerical Algorithms, 31

(1–4):1–3, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
43/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/43/7/fulltext.pdf.

Butcher:2002:SIO

[But02c] J. C. Butcher. Software issues for ordinary differential equations. Numerical Algorithms, 31(1–4):401–418, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/22/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/43/
22/fulltext.pdf.

Butcher:2010:TNM

[But10] J. C. Butcher. Trees and numerical methods for ordinary differential equations. Numerical Algorithms, 53(2–3):153–170, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
53&issue=2&spage=153.

Butcher:2015:CGS

[But15] J. C. Butcher. The cohesiveness of G-symplectic methods. Numerical Algorithms, 70(3):607–624, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9964-y.

Butcher:2019:TBS

[But19] J. C. Butcher. Trees and B-series. Numerical Algorithms, 81 (4):1311–1325, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Barone:1993:NEC

[BV93] Claudio Barone and Ezio Venturino. On the numerical evaluation of Cauchy transforms. Numerical Algorithms, 5(1–4): 429–436, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Bultheel:1995:FOP

[BV95] Adhemar Bultheel and Marc Van Barel. Formal orthogonal polynomials and Hankel/Toeplitz duality. Numerical Algorithms, 10(3–4):289–335, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1996:OPN

[BV96] C. Brezinski and R. S. Varga, editors. Orthogonal Polynomials and Numerical Analysis. Proceedings of the international conference, Luminy, France, September 5–10, 1994, volume 11(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, March 1996. CO-DEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ????

Biro:1999:NIC

[BV99] Zsolt Biro and Maria Vicsek. Numerical interface curves for some nonlinear diffusion equations. Numerical Algorithms, 21(1–4):87–96, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/11/fulltext.pdf.

Bulatov:2009:TSF

[BV09] Mikhail V. Bulatov and Guido Vanden Berghe. Two-step fourth order methods for linear ODEs of the second order. Numerical Algorithms, 51(4):449–460, August 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=4&spage=449.

Bogdanov:2021:NOT

[BV21] Vladimir V. Bogdanov and Yuriy S. Volkov. Near-optimal tension parameters in convexity preserving interpolation by generalized cubic splines. Numerical Algorithms, 86(2):833– 861, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00914-9.

Bleichrodt:2016:EIA

[BvLP16] Folkert Bleichrodt, Tristan van Leeuwen, and Willem Jan Palenstijn. Easy implementation of advanced tomography algorithms using the ASTRA toolbox with spot operators. Numerical Algorithms, 71(3):673–697, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0016-4.

Baxter:2014:CAL

[BVV14] Mathew Baxter, Robert A. Van Gorder, and Kuppalapalle Vajravelu. On the choice of auxiliary linear operator in the optimal homotopy analysis of the Cahn–Hilliard initial value problem. Numerical Algorithms, 66(2):269–298, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9733-8.

Barb:1993:MRT

[BW93] Florin Dan Barb and Martin Weiss. Model reduction techniques for sampled-data systems. Numerical Algorithms, 4(1– 2):47–64, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Benzi:2013:PIM

[BW13] Michele Benzi and Zhen Wang. A parallel implementation of the modified augmented Lagrangian preconditioner for the incompressible Navier–Stokes equations. Numerical Algorithms, 64(1):73–84, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9655-x.

Bialecki:2015:FDS

[BW15] Bernard Bialecki and Lyndsey Wright. A fast direct solver for a fourth order finite difference scheme for Poisson’s equation on the unit disc in polar coordinates. Numerical Algorithms, 70(4):727–751, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9971-z.

Burrage:2024:ULC

[BWB24] Pamela M. Burrage, Hasitha N. Weerasinghe, and Kevin Burrage. Using a library of chemical reactions to fit systems of ordinary differential equations to agent-based models: a machine learning approach. Numerical Algorithms, 96(3):1063– 1077, July 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01737-0.

Bo:2022:AHO

[BWC22] Yonghui Bo, Yushun Wang, and Wenjun Cai. Arbitrary highorder linearly implicit energy-preserving algorithms for Hamiltonian PDEs. Numerical Algorithms, 90(4):1519–1546, August 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01239-x.

Besse:2017:SWR

[BX17] Christophe Besse and Feng Xing. Schwarz waveform relaxation method for one-dimensional Schrödinger equation with general potential. Numerical Algorithms, 74(2):393–426, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0153-4.

Bu:2019:VCP

[BX19] Weiping Bu and Aiguo Xiao. An h–p version of the continuous Petrov–Galerkin finite element method for Riemann–Liouville fractional differential equation with novel test basis functions. Numerical Algorithms, 81(2):529–545, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1991:NPO

[BZ91] C. Brezinski and M. Redivo Zaglia. A new presentation of orthogonal polynomials with applications to their computation. Numerical Algorithms, 1(2):207–221, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:1994:TNB

[BZ94] C. Brezinski and M. Redivo Zaglia. Treatment of nearbreakdown in the CGS algorithm. Numerical Algorithms, 7 (1):33–73, June 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Brezinski:2002:BPM

[BZ02] C. Brezinski and M. Redivo Zaglia. Block projection methods for linear systems. Numerical Algorithms, 29(1–3):33–43, March 2002. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp008.lwwonline.
com/content/getfile/5058/38/3/abstract.htm;   http:
//ipsapp008.lwwonline.com/content/getfile/5058/38/
3/fulltext.pdf.

Bai:2013:MBS

[BZ13] Zhong-Zhi Bai and Li-Li Zhang. Modulus-based synchronous two-stage multisplitting iteration methods for linear complementarity problems. Numerical Algorithms, 62 (1):59–77, January 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9566-x/;

http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=1&spage=59-77.

Brown:2018:MFM

[BZ18] David A. Brown and David W. Zingg. Matrix-free monolithic homotopy continuation with application to computational aerodynamics. Numerical Algorithms, 78(4):1303–1320, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Boukhelkhal:2024:LIP

[BZ24] Ikram Boukhelkhal and Rebiha Zeghdane. Lagrange interpolation polynomials for solving nonlinear stochastic integral equations. Numerical Algorithms, 96(2):583–618, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01659-x. See correction [IZ24].

Bodaghi:2022:DMB

[BZS22] S. Bodaghi, A. Zakeri, and A. H. Salehi Shayegan. Discrete mollification in Bernstein basis and space marching scheme for numerical solution of an inverse two-phase one-dimensional Stefan problem. Numerical Algorithms, 90(4):1569–1592, August 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01242-2.

Bhrawy:2016:STL

[BZV16] A. H. Bhrawy, M. A. Zaky, and R. A. Van Gorder. A space– time Legendre spectral tau method for the two-sided space– time Caputo fractional diffusion-wave equation. Numerical Algorithms, 71(1):151–180, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9990-9.

Caraus:2007:ASS

[CA07] Iurie Nicolae Caraus and Feras M. Al Faqih. Approximate solution of singular integro-differential equations in generalized Hölder spaces. Numerical Algorithms, 45(1–4):205–215, August 2007. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-007-9079-1.

Chandra:2022:AFD

[CA22] Subhash Chandra and Syed Abbas. Analysis of fractal dimension of mixed Riemann–Liouville integral. Numerical Algorithms, 91(3):1021–1046, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01290-2.

Chakir:2022:TDL

[CAB22] Y. Chakir, J. Abouir, and B. Benouahmane. Two-dimensional Laplace transform inversion using bivariate homogeneous twopoint Padé approximants. Numerical Algorithms, 90(3):1153– 1174, July 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01225-3.

Cacao:2010:COS

[Caç10] Isabel Cação. Complete orthonormal sets of polynomial solutions of the Riesz and Moisil–Teodorescu systems in R^{3}. Numerical Algorithms, 55(2–3):191–203, November 2010. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=2&
spage=191.

Cai:2022:OPL

[Cai22] Haotao Cai. Oscillation-preserving Legendre–Galerkin methods for second kind integral equations with highly oscillatory kernels. Numerical Algorithms, 90(3):1091–1115, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01223-5.

Calvo:2020:NCS

[Cal20] Juan G. Calvo. A new coarse space for overlapping Schwarz algorithms for H(curl) problems in three dimensions with irregular subdomains. Numerical Algorithms, 83(3):885–899, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Campos:1995:QFH

[Cam95] R. G. Campos. A quadrature formula for the Hankel transform. Numerical Algorithms, 9(3–4):343–354, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cameron:2019:EIM

[Cam19] Thomas R. Cameron. An effective implementation of a modified Laguerre method for the roots of a polynomial. Numerical Algorithms, 82(3):1065–1084, November 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cao:2012:NSA

[Cao12] Zhi-Hao Cao. A note on spectrum analysis of augmentation block Schur complement preconditioners. Numerical Algorithms, 59(4):561–569, April 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=561.

Carnicer:1991:BCI

[Car91] J. M. Carnicer. On best constrained interpolation. Numerical Algorithms, 1(2):155–176, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Carnicer:1992:RIS

[Car92] J. M. Carnicer. Rational interpolation with a single variable pole. Numerical Algorithms, 3(1–4):125–132, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Carlson:1995:NCR

[Car95] B. C. Carlson. Numerical computation of real or complex elliptic integrals. Numerical Algorithms, 10(1–2):13–26, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Carletti:2001:NDI

[Car01] M. Carletti. Numerical determination of the instability region for a delay model of phage–bacteria interaction. Numerical Algorithms, 28(1–4):27–44, December 2001. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
37/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/3/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
3/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/3/fulltext.pdf.

Carnicer:2010:WIE

[Car10] Jesus M. Carnicer. Weighted interpolation for equidistant nodes. Numerical Algorithms, 55(2–3):223–232, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=2&
spage=223.

Castellini:2017:KIM

[Cas17] Jacopo Castellini. Krylov iterative methods for the geometric mean of two matrices times a vector. Numerical Algorithms, 74(2):561–571, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0161-4.

Catinas:2024:CSS

[Cat24a] Emil Catinas. Correction to: The strict superlinear order can be faster than the infinite order. Numerical Algorithms, 95(3):1187, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01658-y. See

[Cat24b].

Catinas:2024:SSO

[Cat24b] Emil Catinas. The strict superlinear order can be faster than the infinite order. Numerical Algorithms, 95(3):1177– 1186, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01604-y. See

correction [Cat24a].

Causley:2022:GFI

[Cau22] Matthew F. Causley. The gamma function via interpolation. Numerical Algorithms, 90(2):687–707, June 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01204-8.

Chandra:2023:BSF

[CAV23] Subhash Chandra, Syed Abbas, and Saurabh Verma. Bernstein super fractal interpolation function for countable data systems. Numerical Algorithms, 92(4):2457–2481, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01398-5.

Coulaud:2000:PCM

[CB00] Olivier Coulaud and Pierre-Eric Bernard. Parallel constrained molecular dynamics. Numerical Algorithms, 24 (4):393–405, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
29/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/29/6/fulltext.pdf. Sparse

matrices and differential equations (Rennes, 1999).

Chang:2013:ESA

[CB13] Chia-Tche Chang and Vincent D. Blondel. An experimental study of approximation algorithms for the joint spectral radius. Numerical Algorithms, 64(1):181–202, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9661-z.

Campbell:2016:CDT

[CB16] Stephen L. Campbell and John T. Betts. Comments on direct transcription solution of DAE constrained optimal control problems with two discretization approaches. Numerical Algorithms, 73(3):807–838, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0119-6.

Cruz-Barroso:2007:BOS

[CBGVN07] Ruymán Cruz-Barroso, Pablo González-Vera, and Olav Njåstad. On bi-orthogonal systems of trigonometric functions and quadrature formulas for periodic integrands. Numerical Algorithms, 44(4):309–333, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=4&spage=309.

Cruz-Barroso:2009:ASP

[CBGVPP09] R. Cruz-Barroso, P. González-Vera, and F. Perdomo-Pío. An application of Szegő polynomials to the computation of certain weighted integrals on the real line. Numerical Algorithms, 52(3):273–293, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=273.

Chehab:2003:MSE

[CC03] J. P. Chehab and B. Costa. Multiparameter schemes for evolutionary equations. Numerical Algorithms, 34(2– 4):245–257, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
24/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/24/fulltext.pdf.

Chan:2006:NMA

[CC06] Tony F. Chan and Ke Chen. On a nonlinear multigrid algorithm with primal relaxation for the image total variation minimisation. Numerical Algorithms, 41(4):387–411, April 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=41&issue=4&
spage=387.

Capobianco:2007:SRN

[CC07] M. R. Capobianco and G. Criscuolo. Some remarks on the numerical computation of integrals on an unbounded interval. Numerical Algorithms, 45(1–4):37–48, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=37.

Capobianco:2012:CSN

[CC12] M. R. Capobianco and G. Criscuolo. Convergence and stability of a new quadrature rule for evaluating Hilbert transform. Numerical Algorithms, 60(4):579–592, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=579.

Cheng:2013:NSM

[CC13] Wanyou Cheng and Zixin Chen. Nonmonotone spectral method for large-scale symmetric nonlinear equations. Numerical Algorithms, 62(1):149–162, January 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9572-z/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=149-162.

Chao:2015:NSC

[CC15] Zhen Chao and Guoliang Chen. A note on semi-convergence of generalized parameterized inexact Uzawa method for singular saddle point problems. Numerical Algorithms, 68(1): 95–105, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9840-1.

Chao:2016:GMH

[CC16a] Zhen Chao and Guo-Liang Chen. A generalized modified HSS method for singular complex symmetric linear systems. Numerical Algorithms, 73(1):77–89, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0086-3.

Chen:2016:NMH

[CC16b] Y. Chen and Chang-Ming Chen. Numerical method with high order accuracy for solving a anomalous subdiffusion equation. Numerical Algorithms, 72(3):687–703, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0062-y.

Chen:2018:NAS

[CC18] Y. Chen and Chang-Ming Chen. Numerical algorithm for solving the Stokes’ first problem for a heated generalized second grade fluid with fractional derivative. Numerical Algorithms, 77(3):939–953, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2025:STM

[CCC25] Ziheng Chen, Liangmin Cao, and Lin Chen. Stochastic theta methods for random periodic solution of stochastic differential equations under non-globally Lipschitz conditions. Numerical Algorithms, 99(2):651–681, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01892-y.

Christara:2010:QSC

[CCD10] Christina C. Christara, Tong Chen, and Duy Minh Dang. Quadratic spline collocation for one-dimensional linear parabolic partial differential equations. Numerical Algorithms, 53(4): 511–553, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=53&issue=4&spage=511.

Capobianco:2001:SCA

[CCG01] M. R. Capobianco, G. Criscuolo, and R. Giova. A stable and convergent algorithm to evaluate Hilbert transform. Numerical Algorithms, 28(1–4):11–26, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/37/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
37/2/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/2/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/37/2/fulltext.pdf.

Chen:2023:GDL

[CCHH23] Xiao-Min Chen, Xiang-Ke Chang, Yi He, and Xing-Biao Hu. Generalized discrete Lotka–Volterra equation, orthogonal polynomials and generalized epsilon algorithm. Numerical Algorithms, 92(1):335–375, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01365-0.

Chien:1999:CDD

[CCJ99] C.-S. Chien, H.-S. Chou, and B.-W. Jeng. A continuation– domain decomposition algorithm for bifurcation problems.

Numerical Algorithms, 22(3–4):367–383, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/23/1/abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/5058/23/1/fulltext.
pdf.

Capobianco:2010:NMC

[CCJ10] Maria Rosaria Capobianco, Giuliana Criscuolo, and Peter Junghanns. Newton methods for a class of nonlinear hypersingular integral equations. Numerical Algorithms, 55(2–3):205–221, November 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=205.

Chen:2018:CSA

[CCJC18] H.-S. Chen, S.-L. Chang, B.-W. Jeng, and C.-S. Chien. Continuation and stability analysis for Bloch waves of the Gross– Pitaevskii equation. Numerical Algorithms, 77(3):709–726, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Csallner:2004:RNC

[CCK04] András Erik Csallner, Tibor Csendes, and András Balázs Kocsis. Reliable numerical computation in civil engineering. Numerical Algorithms, 37(1–4):85–91, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/36/abstract.htm.

Chen:2016:PBO

[CCL16] Yan Mei Chen, Xiao Shan Chen, and Wen Li. On perturbation bounds for orthogonal projections. Numerical Algorithms, 73(2):433–444, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0102-2.

Che:2018:NSI

[CCL18] Haitao Che, Haibin Chen, and Meixia Li. A new simultaneous iterative method with a parameter for solving the extended split equality problem and the extended split equality fixed point problem. Numerical Algorithms, 79(4):1231–1256, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cuyt:2016:AMP

[CCLi16] Annie Cuyt, Oliver Salazar Celis, Maryna Lukach, and Karel in ’t Hout. Analytic models for parameter dependency in option price modelling. Numerical Algorithms, 73(1):15–31, September 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0084-5.

Cai:2025:CRE

[CCN25] Jingjing Cai, Ziheng Chen, and Yuanling Niu. Convergence rate and exponential stability of backward Euler method for neutral stochastic delay differential equations under generalized monotonicity conditions. Numerical Algorithms, 98(4): 2005–2035, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01862-4.

Charina:2005:RMV

[CCS05] Maria Charina, Costanza Conti, and Thomas Sauer. Regularity of multivariate vector subdivision schemes. Numerical Algorithms, 39(1–3):97–113, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Campos:2016:OPT

[CCTV16] Beatriz Campos, Alicia Cordero, Juan R. Torregrosa, and Pura Vindel. Orbits of period two in the family of a multipoint variant of Chebyshev–Halley family. Numerical Algorithms, 73(1):141–156, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0089-0.

Candelario:2023:GCF

[CCTV23] Giro Candelario, Alicia Cordero, Juan R. Torregrosa, and María P. Vassileva. Generalized conformable fractional Newton-type method for solving nonlinear systems. Numerical Algorithms, 93(3):1171–1208, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01463-
z.

Celis:2007:RAV

[CCV07] Oliver Salazar Celis, Annie Cuyt, and Brigitte Verdonk. Rational approximation of vertical segments. Numerical Algorithms, 45(1–4):375–388, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-007-9077-3.

Campos:2023:DNL

[CCV23] B. Campos, J. Canela, and P. Vindel. Dynamics of Newton-like root finding methods. Numerical Algorithms, 93(4):1453–1480, August 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01474-w.

Choi:2021:HDS

[CCW21] Bosu Choi, Andrew Christlieb, and Yang Wang. Highdimensional sparse Fourier algorithms. Numerical Algorithms, 87(1):161–186, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00962-1.

Caliari:2023:MBA

[CCZ23] Marco Caliari, Fabio Cassini, and Franco Zivcovich. A μmode BLAS approach for multidimensional tensor-structured problems. Numerical Algorithms, 92(4):2483–2508, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01399-4.

Cecchi:1996:MBM

[CD96] M. Morandi Cecchi and E. Di Nardo. The modified bordering method to evaluate eigenvalues and eigenvectors of normal matrices. Numerical Algorithms, 11(1–4):285–309, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Chen:1999:NFD

[CD99] Guoting Chen and Jean Della Dora. Normal forms for differentiable maps near a fixed point. Numerical Algorithms, 22(2):213–230, December 1999. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
22/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/22/7/fulltext.pdf.

Csordas:2000:CTT

[CD00] George Csordas and Dimitar K. Dimitrov. Conjectures and theorems in the theory of entire functions. Numerical Algorithms, 25(1–4):109–122, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/5/fulltext.pdf. Mathemat-

ical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Costabile:2001:ESR

[CD01] F. Costabile and F. Dell’Accio. Expansions over a simplex of real functions by means of Bernoulli polynomials. Numerical Algorithms, 28(1–4):63–86, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/37/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
37/5/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/5/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/37/5/fulltext.pdf.

Costabile:2007:NEB

[CD07] F. A. Costabile and F. Dell’Accio. New embedded boundarytype quadrature formulas for the simplex. Numerical Algorithms, 45(1–4):253–267, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-007-9088-0.

Chen:2015:SMM

[CD15] Shaohua Chen and Lauren DeDieu. A simple moving mesh method for blowup problems. Numerical Algorithms, 69(2): 343–356, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9901-5.

Cetin:2025:AMF

[ÇD25a] Coskun Çetin and Jasmina Dordević. Approximate moment functions for logistic stochastic differential equations. Numerical Algorithms, 99(3):1161–1202, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01911-y.

Chen:2025:ISA

[CD25b] Xin Chen and Peichao Duan. Inertial self-adaptive algorithms for solving non-smooth convex optimization problems. Numerical Algorithms, 98(1):133–163, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01788-x.

Costabile:2013:CLI

[CDD13] F. A. Costabile, F. Dell’Accio, and F. Di Tommaso. Complementary Lidstone interpolation on scattered data sets. Numerical Algorithms, 64(1):157–180, September 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9659-6.

Cipolla:2021:RIP

[CDD21] Stefano Cipolla, Marco Donatelli, and Fabio Durastante. Regularization of inverse problems by an approximate matrixfunction technique. Numerical Algorithms, 88(3):1275–1308, November 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01076-y;   http://link.
springer.com/content/pdf/10.1007/s11075-021-01076-
y.pdf.

Cecchi:1999:PRC

[CDF99] M. Morandi Cecchi, S. De Marchi, and D. Fasoli. A package for representing C1 interpolating surfaces: Application to the lagoon of venice’s bed. Numerical Algorithms, 20(2–3):197–215, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/5/fulltext.pdf.

Chaudet-Dumas:2023:CPD

[CDG23] Bastien Chaudet-Dumas and Martin J. Gander. Cross-points in the Dirichlet–Neumann method I: well-posedness and convergence issues. Numerical Algorithms, 92(1):301–334, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01445-1.

Chueca-Diez:2025:FNA

[CDGC25] Fernando Chueca-Díez and Alfonso M. Gañán-Calvo. A fast numerical algorithm for finding all real solutions to a system of N nonlinear equations in a finite domain. Numerical Algorithms, 99(3):1081–1110, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01908-7.

Conte:2014:NVR

[CDI14] Dajana Conte, Raffaele D’Ambrosio, and Giuseppe Izzo. Natural Volterra Runge–Kutta methods. Numerical Algorithms, 65 (3):421–445, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9790-z.

Chen:2021:IEP

[CDLW21] Xu Chen, Deng Ding, Siu-Long Lei, and Wenfei Wang. An implicit–explicit preconditioned direct method for pricing options under regime-switching tempered fractional partial differential models. Numerical Algorithms, 87(3):939–965, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00994-7.

Conte:2016:GAW

[CDP16] Dajana Conte, Raffaele D’Ambrosio, and Beatrice Paternoster. GPU-acceleration of waveform relaxation methods for large differential systems. Numerical Algorithms, 71(2):293– 310, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9993-6.

CortesGarcia:2020:SAF

[CDS20] Idoia Cortes Garcia, Herbert De Gersem, and Sebastian Schöps. A structural analysis of field/circuit coupled problems based on a generalised circuit element. Numerical Algorithms, 83(1):373–394, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Caliari:2011:PFI

[CDSV11] Marco Caliari, Stefano De Marchi, Alvise Sommariva, and Marco Vianello. Padua2DM: fast interpolation and cubature at the Padua points in Matlab/Octave. Numerical Algorithms, 56(1):45–60, January 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=45.

Cachadina:2010:NAC

[CDT10] Maria Isabel Bueno Cachadina, Alfredo Deaño, and Edward Tavernetti. A new algorithm for computing the Geronimus transformation with large shifts. Numerical Algorithms, 54(1):101–139, May 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=101.

Chan:1998:TFF

[Cdv98] Tony F. Chan, Lisette de Pillis, and Henk van der Vorst. Transpose-free formulations of Lanczos-type methods for nonsymmetric linear systems. Numerical Algorithms, 17 (1–2):51–66, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/31/
12/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/31/12/fulltext.pdf.

Choi:1995:DPD

[CDW95] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. The design of a parallel dense linear algebra software library: reduction to Hessenberg, tridiagonal, and bidiagonal form. Numerical Algorithms, 10(3–4):379–399, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Censor:1994:MAU

[CE94] Yair Censor and Tommy Elfving. A multiprojection algorithm using Bregman projections in a product space. Numerical Algorithms, 8(2–4):221–239, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Calver:2017:NMC

[CE17] Jonathan Calver and Wayne Enright. Numerical methods for computing sensitivities for ODEs and DDEs. Numerical Algorithms, 74(4):1101–1117, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0188-6.

Cegielski:2024:SPD

[Ceg24] Andrzej Cegielski. Strict pseudocontractions and demicontractions, their properties, and applications. Numerical Algorithms, 95(4):1611–1642, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01623-
9.

Cegielski:2025:LCL

[Ceg25] Andrzej Cegielski. On linear convergence of Landweber-type methods. Numerical Algorithms, 100(4):1929–1964, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02163-0.

Cibik:2021:PCS

[CEK21] Aytekin Cibik, Fatma G. Eroglu, and Songül Kaya. On the performance of curvature stabilization time stepping methods for double-diffusive natural convection flows in the presence of magnetic field. Numerical Algorithms, 88(1):475–498, September 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01046-w.

Cen:2014:ASO

[CEX14] Zhongdi Cen, Fevzi Erdogan, and Aimin Xu. An almost second order uniformly convergent scheme for a singularly perturbed initial value problem. Numerical Algorithms, 67(2): 457–476, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9801-0.

Calloo:2026:CFP

[CEZ26] Ansar Calloo, Matthew Evans, and Luca Zanetti. Cyclefree polytopal mesh sweeping for Boltzmann transport. Numerical Algorithms, 101(4):2157–2180, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02081-1.

Carnicer:1996:PLI

[CF96] J. M. Carnicer and M. S. Floater. Piecewise linear interpolants to Lagrange and Hermite convex scattered data. Numerical Algorithms, 13(3–4):345–364, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cecchi:2005:FHA

[CF05] Maria Morandi Cecchi and Massimo Fornasier. Fast homogenization algorithm based on asymptotic theory and multiscale schemes. Numerical Algorithms, 40(2):171–186, October 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=2&
spage=171.

Crespo:2020:MSM

[CFK^{+}20] S. Crespo, M. Fasondini, C. Klein, N. Stoilov, and C. Vallée. Multidomain spectral method for the Gauss hypergeometric function. Numerical Algorithms, 84(1):1–35, May 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2019:DMA

[CFL19] Xiaohong Chen, Xiufang Feng, and Zhilin Li. A direct method for accurate solution and gradient computations for elliptic interface problems. Numerical Algorithms, 80(3):709–740, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cavoretto:2015:IHS

[CFM15] Roberto Cavoretto, Gregory E. Fasshauer, and Michael Mc-Court. An introduction to the Hilbert–Schmidt SVD using iterated Brownian bridge kernels. Numerical Algorithms, 68(2): 393–422, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9850-z.

Casciola:2006:MDD

[CFR06] G. Casciola, E. Franchini, and L. Romani. The mixed directional difference-summation algorithm for generating the Bézier net of a trivariate four-direction Box-spline. Numerical Algorithms, 43(1):75–98, September 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=1&spage=75.

Concas:2019:PMP

[CFR19] Anna Concas, Caterina Fenu, and Giuseppe Rodriguez. PQser: a Matlab package for spectral seriation. Numerical Algorithms, 80(3):879–902, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Concas:2023:SPP

[CFRV23] Anna Concas, Caterina Fenu, Giuseppe Rodriguez, and Raf Vandebril. The seriation problem in the presence of a double Fiedler value. Numerical Algorithms, 92(1):407–435, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01461-1.

Chow:2021:ARI

[CFS21] Edmond Chow, Andreas Frommer, and Daniel B. Szyld. Asynchronous Richardson iterations: theory and practice. Numerical Algorithms, 87(4):1635–1651, August 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-01023-3.

Carnicer:2003:ACD

[CG03] J. M. Carnicer and M. Gasca. Asymptotic conditions for degree diminution along prescribed lines. Numerical Algorithms, 33(1–4):183–192, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
15/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/15/fulltext.pdf.

Carnicer:2005:GLP

[CG05] J. M. Carnicer and M. Gasca. Generation of lattices of points for bivariate interpolation. Numerical Algorithms, 39(1–3):69– 79, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Carnicer:2007:GPL

[CG07a] J. M. Carnicer and C. Godés. Generalized principal lattices and cubic pencils. Numerical Algorithms, 44(2):133–145, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=44&issue=2&
spage=133.

Cates:2007:DDD

[CG07b] Dennis Cates and Anne Gelb. Detecting derivative discontinuity locations in piecewise continuous functions from Fourier spectral data. Numerical Algorithms, 46(1):59–84, September 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=46&issue=1&
spage=59.

Carnicer:2009:CSS

[CG09] J. M. Carnicer and C. Godés. Classification of sets satisfying the geometric characterization. Numerical Algorithms, 50(2):145–154, February 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=145.

Courvoisier:2013:OSW

[CG13] Yves Courvoisier and Martin J. Gander. Optimization of Schwarz waveform relaxation over short time windows. Numerical Algorithms, 64(2):221–243, October 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9662-y.

Carnicer:2014:ID

[CG14] J. M. Carnicer and C. Godés. Interpolation on the disk. Numerical Algorithms, 66(1):1–16, May 2014. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9720-0.

Carnicer:2017:ISP

[CG17] J. M. Carnicer and C. Godés. Interpolation with symmetric polynomials. Numerical Algorithms, 74(1):1–18, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0135-6.

Clavero:2019:UCA

[CG19] C. Clavero and J. L. Gracia. Uniformly convergent additive schemes for 2 d singularly perturbed parabolic systems of reaction–diffusion type. Numerical Algorithms, 80(4):1097– 1120, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Castillo:2020:CLD

[CG20] P. Castillo and S. Gómez. Conservative local discontinuous Galerkin method for the fractional Klein–Gordon–Schrödinger system with generalized Yukawa interaction. Numerical Algorithms, 84(1):407–425, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Censor:2021:DFS

[CGHH21] Yair Censor, Edgar Garduño, Elias S. Helou, and Gabor T. Herman. Derivative-free superiorization: principle and algorithm. Numerical Algorithms, 88(1):227–248, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01038-w.

Clavero:1999:HOM

[CGL99] C. Clavero, J. L. Gracia, and F. Lisbona. High order methods on Shishkin meshes for singular perturbation problems of convection–diffusion type. Numerical Algorithms, 22(1):73–97, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
21/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/21/6/fulltext.pdf.

Costabile:2001:NIM

[CGL01] F. Costabile, M. I. Gualtieri, and R. Luceri. A new iterative method for the computation of the solutions of nonlinear equations. Numerical Algorithms, 28(1–4): 87–100, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
37/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/6/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
6/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/6/fulltext.pdf.

Coope:1993:RFV

[CGM93] I. D. Coope and P. R. Graves-Morris. The rise and fall of the vector ϵ-algorithm. Numerical Algorithms, 5(1–4):275–286, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Castillo:2012:ZSO

[CGM12] K. Castillo, L. E. Garza, and F. Marcellán. Zeros of Sobolev orthogonal polynomials on the unit circle. Numerical Algorithms, 60(4):669–681, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=669.

Chouly:2025:OSW

[CGM25] Franz Chouly, Martin J. Gander, and Véronique Martin. Optimized Schwarz waveform relaxation methods for waveheat coupling in one dimensional bounded domains. Numerical Algorithms, 100(4):1739–1763, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02100-1.

Chakib:2003:SNA

[CGN03] A. Chakib, T. Ghemires, and A. Nachaoui. Study of a numerical approach for a transient flow problem in porous media. Numerical Algorithms, 34(2–4):229–243, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/26/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
26/fulltext.pdf.

Costabile:2022:BGA

[CGN22] F. A. Costabile, M. I. Gualtieri, and A. Napoli. Bivariate general Appell interpolation problem. Numerical Algorithms, 91 (2):531–556, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01272-4.

Calvo:2000:ISA

[CGPM00] M. Calvo, S. González-Pinto, and J. I. Montijano. On the iterative solution of the algebraic equations in fully implicit Runge– Kutta methods. Numerical Algorithms, 23(1):97–113, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/24/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
24/5/fulltext.pdf.

Censor:2012:ASV

[CGR12] Yair Censor, Aviv Gibali, and Simeon Reich. Algorithms for the Split Variational Inequality Problem. Numerical Algorithms, 59(2):301–323, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=2&spage=301.

Chen:2024:NBP

[CGSS24] Xu Chen, Xin-Xin Gong, Zi-Run She, and Zi-Hang She. A novel banded preconditioner for coupled tempered fractional diffusion equation generated from the regime-switching CGMY model. Numerical Algorithms, 97(4):1791–1822, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01769-0.

Camacho:1992:RAS

[CGV92] M. Camacho and P. González-Vera. Rational approximants to symmetric formal Laurent series. Numerical Algorithms, 3(1–4):117–124, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Chaouqui:2022:LNS

[CGV22] F. Chaouqui, M. J. Gander, and T. Vanzan. Linear and nonlinear substructured Restricted Additive Schwarz iterations and preconditioning. Numerical Algorithms, 91(1):81–107, September 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01255-5.

Cocquet:2026:DTO

[CGV26] Pierre-Henri Cocquet, Martin J. Gander, and Alexandre Vieira. Decompose-then-optimize: a new approach to design domain decomposition methods for optimal control problems. Numerical Algorithms, 101(2):1395–1423, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02046-4.

Chen:2025:OPA

[CGW25] Tyler Chen, Anne Greenbaum, and Natalie Wellen. Optimal polynomial approximation to rational matrix functions using the Arnoldi algorithm. Numerical Algorithms, 100(4): 1657–1678, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02085-x.

Chen:2019:IOM

[CGYZ19] Ke Chen, Geovani Nunes Grapiglia, Jinyun Yuan, and Daoping Zhang. Improved optimization methods for image registration problems. Numerical Algorithms, 80(2):305–336, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chui:1995:BIR

[CH95] C. K. Chui and T. X. He. Bivariate interpolatory rational splines. Numerical Algorithms, 9(3–4):277–291, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cheng:2011:EAS

[CH11] Pan Cheng and Jin Huang. Extrapolation algorithms for solving nonlinear boundary integral equations by mechanical quadrature methods. Numerical Algorithms, 58(4):545–554, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
58&issue=4&spage=545.

Chen:2022:FDB

[CH22] Jia-Qi Chen and Zheng-Da Huang. On a fast deterministic block Kaczmarz method for solving large-scale linear systems. Numerical Algorithms, 89(3):1007–1029, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01143-4.

Carson:2026:DCS

[CH26] Erin Claire Carson and Jakub Hercík. The detection and correction of silent errors in pipelined Krylov subspace methods. Numerical Algorithms, 101(2):1175–1210, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02037-5.

Chausova:2004:DNI

[Cha04] Elena V. Chausova. Dynamic network inventory control model with interval nonstationary demand uncertainty. Numerical Algorithms, 37(1–4):71–84, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/5/abstract.htm.

Chan:2014:RBH

[Cha14] T. M. H. Chan. Relations between the Hopf algebra and the B-series with new concepts. Numerical Algorithms, 65(3): 669–685, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9812-x.

Chesneaux:1994:ERS

[Che94] Jean-Marie Chesneaux. The equality relations in scientific computing. Numerical Algorithms, 7(2–4):129–143, July 1994.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:1999:GDQ

[Che99] Chang-New Chen. Generalization of differential quadrature discretization. Numerical Algorithms, 22(2):167–182, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/22/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
22/4/fulltext.pdf.

Chen:2001:FRN

[Che01] Guoting Chen. Further reduction of normal forms for vector fields. Numerical Algorithms, 27(1):1–33, May 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/33/1/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/33/
1/fulltext.pdf.

Chen:2002:VOE

[Che02] D. J. L. Chen. Variable-order ESIRK methods for stiff differential equations. Numerical Algorithms, 31(1–4): 103–114, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/20/fulltext.pdf.

Chen:2004:URH

[Che04] Guizhi Chen. The use of refined harmonic shifts in the implicitly restarted refined harmonic Arnoldi method. Numerical Algorithms, 35(1):91–96, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/49/A/6/abstract.htm.

Chen:2013:UTS

[Che13] Chao-Ping Chen. Unified treatment of several asymptotic formulas for the gamma function. Numerical Algorithms, 64(2): 311–319, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9667-6.

Chen:2014:ESI

[Che14] D. J. L. Chen. The efficiency of singly-implicit Runge–Kutta methods for stiff differential equations. Numerical Algorithms, 65(3):533–554, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9797-5.

Chen:2016:IAE

[Che16a] Chao-Ping Chen. Inequalities and asymptotics for the Euler– Mascheroni constant based on DeTemple’s result. Numerical Algorithms, 73(3):761–774, November 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-016-0116-9.

Chen:2016:SPI

[Che16b] Hao Chen. A splitting preconditioner for implicit Runge– Kutta discretizations of a partial differential-algebraic equation. Numerical Algorithms, 73(4):1037–1054, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0128-5.

Cheng:2019:OEE

[Che19] Yao Cheng. Optimal error estimate of the local discontinuous Galerkin methods based on the generalized alternating numerical fluxes for nonlinear convection-diffusion equations. Numerical Algorithms, 80(4):1329–1359, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2022:SOI

[Che22] Yong Chen. Second-order IMEX scheme for a system of partial integro-differential equations from Asian option pricing under regime-switching jump-diffusion models. Numerical Algorithms, 89(4):1823–1843, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01174-

Chen:2024:CIS

[Che24] Yong Chen. Compact IMEX scheme for a moving boundary PIDE system of the regime-switching jump-diffusion Asian option pricing. Numerical Algorithms, 95(3):1055–1077, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01599-6.

Cox:1993:ARN

[CHH93] M. G. Cox, P. M. Harris, and D. A. Humphreys. An algorithm for the removal of noise and jitter in signals and its application to picosecond electrical measurement. Numerical Algorithms, 5(1–4):491–508, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chang:2020:CNG

[CHH^{+}20] Xiang-Ke Chang, Yi He, Xing-Biao Hu, Jian-Qing Sun, and Ernst Joachim Weniger. Construction of new generalizations of Wynn’s epsilon and rho algorithm by solving finite difference equations in the transformation order. Numerical Algorithms, 83(2):593–627, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chang:2018:NIC

[CHHL18] Xiang-Ke Chang, Yi He, Xing-Biao Hu, and Shi-Hao Li. A new integrable convergence acceleration algorithm for computing Brezinski–Durbin–Redivo–Zaglia’s sequence transformation via pfaffians. Numerical Algorithms, 78(1):87–106, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cvetkovic:2014:CPM

[CHK14] Lj. Cvetković, A. Hadjidimos, and V. Kostić. On the choice of parameters in MAOR type splitting methods for the linear complementarity problem. Numerical Algorithms, 67(4):793– 806, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9824-1.

Chkifa:2020:LFL

[Chk20] Moulay Abdellah Chkifa. Leja, Fejér–Leja and R–Leja sequences for Richardson iteration. Numerical Algorithms, 84(4):

1481–1505, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00903-y.

Cordero:2010:ETS

[CHMT10a] Alicia Cordero, José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. Efficient three-step iterative methods with sixth order convergence for nonlinear equations. Numerical Algorithms, 53(4):485–495, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=4&spage=485.

Cordero:2010:MNJ

[CHMT10b] Alicia Cordero, José L. Hueso, Eulalia Martínez, and Juan R. Torregrosa. A modified Newton–Jarratt’s composition. Numerical Algorithms, 55(1):87–99, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=1&spage=87.

Cholamjiak:2016:GFB

[Cho16] Prasit Cholamjiak. A generalized forward-backward splitting method for solving quasi inclusion problems in Banach spaces. Numerical Algorithms, 71(4):915–932, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0030-6.

Cho:2017:NAB

[Cho17] C.-H. Cho. A numerical algorithm for blow-up problems revisited. Numerical Algorithms, 75(3):675–697, July 2017. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Censor:2019:DFS

[CHS19] Yair Censor, Howard Heaton, and Reinhard Schulte. Derivativefree superiorization with component-wise perturbations. Numerical Algorithms, 80(4):1219–1240, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cummings:2025:SCR

[CHS25] Joseph Cummings, Jonathan D. Hauenstein, and Clifford D. Smyth. Smooth connectivity in real algebraic varieties. Numerical Algorithms, 100(1):35–62, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01952-3.

Cao:2019:SAT

[CHY19] Mingyuan Cao, Qingdao Huang, and Yueting Yang. A selfadaptive trust region method for extreme B-eigenvalues of symmetric tensors. Numerical Algorithms, 81(2):407–420, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2024:OPS

[CHYH24] Cairong Chen, Bo Huang, Dongmei Yu, and Deren Han. Optimal parameter of the SOR-like iteration method for solving absolute value equations. Numerical Algorithms, 96(2):799– 826, June 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01666-y.

Campbell:1998:MSNa

[CHYZ98] S. L. Campbell, R. Hollenbeck, K. Yeomans, and Y. Zhong. Mixed symbolic-numerical computations with general DAEs. I. System properties. Numerical Algorithms, 19(1–4): 73–83, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
16/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/8/fulltext.pdf. Differen-

tial algebraic equations (Grenoble, 1997).

Ciarlet:1994:RRB

[Cia94] P. Ciarlet, Jr. Repeated Red-Black ordering: a new approach. Numerical Algorithms, 7(2–4):295–324, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cicone:2020:IFD

[Cic20] Antonio Cicone. Iterative filtering as a direct method for the decomposition of nonstationary signals. Numerical Algorithms, 85(3):811–827, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00838-z.

Cardone:2010:EFD

[CIP10] Angelamaria Cardone, Liviu Gr. Ixaru, and Beatrice Paternoster. Exponential fitting Direct Quadrature methods for Volterra integral equations. Numerical Algorithms, 55(4):467–480, December 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=4&spage=467.

Conti:2004:COC

[CJ04] C. Conti and K. Jetter. Concerning order of convergence for subdivision. Numerical Algorithms, 36(4):345–363, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cardone:2012:ENM

[CJ12] Angelamaria Cardone and Zdzisław Jackiewicz. Explicit Nordsieck methods with quadratic stability. Numerical Algorithms, 60(1):1–25, May 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=1&spage=1.

Clavero:2017:FSM

[CJ17] C. Clavero and J. C. Jorge. A fractional step method for 2D parabolic convection-diffusion singularly perturbed problems: uniform convergence and order reduction. Numerical Algorithms, 75(3):809–826, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Clavero:2020:EUC

[CJ20] C. Clavero and J. C. Jorge. An efficient and uniformly convergent scheme for one-dimensional parabolic singularly perturbed semilinear systems of reaction-diffusion type. Numerical Algorithms, 85(3):1005–1027, November 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-019-00850-3.

Chen:2022:RDA

[CJK22] C. S. Chen, Malgorzata A. Jankowska, and Andreas Karageorghis. RBF-DQ algorithms for elliptic problems in axisymmetric domains. Numerical Algorithms, 89(1):33–63, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01105-w.

Choi:2023:LCD

[CJKL23] Yoonjeong Choi, Gwanghyun Jo, Do Y. Kwak, and Young Ju Lee. Locally conservative discontinuous bubble scheme for Darcy flow and its application to Hele–Shaw equation based on structured grids. Numerical Algorithms, 92(2):1127–1152, February 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01333-8.

Cardone:2014:EBI

[CJSZ14] Angelamaria Cardone, Zdzisław Jackiewicz, Adrian Sandu, and Hong Zhang. Extrapolation-based implicit-explicit general linear methods. Numerical Algorithms, 65(3):377–399, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9759-y.

Craviotto:1996:CSF

[CJTW96] Cathleen M. Craviotto, William B. Jones, W. J. Thron, and Nancy J. Wyshinski. Computation of special functions by Padé approximants with orthogonal polynomial denominators. Numerical Algorithms, 11(1–4):117–141, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Castano:2005:MRW

[CK05] Daniel Castaño and Angela Kunoth. Multilevel regularization of wavelet based fitting of scattered data — some experiments. Numerical Algorithms, 39(1–3):81–96, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cvetkovic:2006:NSB

[CK06] Ljiljana Cvetković and Vladimir Kostić. New subclasses of block H-matrices with applications to parallel decompositiontype relaxation methods. Numerical Algorithms, 42(3–4):325– 334, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
42&issue=3&spage=325.

Chen:2020:RBA

[CK20] Minhong Chen and Daniel Kressner. Recursive blocked algorithms for linear systems with Kronecker product structure. Numerical Algorithms, 84(3):1199–1216, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00797-5.

Chouly:2022:WHC

[CK22] Franz Chouly and Pauline Klein. Wave-heat coupling in onedimensional unbounded domains: artificial boundary conditions and an optimized Schwarz method. Numerical Algorithms, 90(2):631–668, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01201-
x.

Castillo:2024:RRS

[CK24] S. J. Castillo and J. Karátson. Rates of robust superlinear convergence of preconditioned Krylov methods for elliptic FEM problems. Numerical Algorithms, 96(2):719–738, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01663-1.

Cordero:2016:SCI

[CKKT16] Alicia Cordero, Munish Kansal, Vinay Kanwar, and Juan R. Torregrosa. A stable class of improved second-derivative free Chebyshev–Halley type methods with optimal eighth order convergence. Numerical Algorithms, 72(4):937–958, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0075-6.

Chen:2016:CLS

[CKL16] C. S. Chen, A. Karageorghis, and Yan Li. On choosing the location of the sources in the MFS. Numerical Algorithms, 72 (1):107–130, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0036-0.

Corless:2019:ORD

[CKM19] Robert M. Corless, C. Yalçin Kaya, and Robert H. C. Moir. Optimal residuals and the Dahlquist test problem. Numerical Algorithms, 81(4):1253–1274, August 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cai:2025:LNM

[CKM^{+}25] Xiao-Min Cai, Yi-Fen Ke, Chang-Feng Ma, Ya-Jun Xie, and Ri-Wei Liao. Logarithmic norm minimization of quaternion matrix decomposition for color image sparse representation. Numerical Algorithms, 99(2):521–547, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01887-9.

CJulio:2024:CPL

[CKO24] Dana C’Julio, Bernd Krauskopf, and Hinke M. Osinga. Computing parametrised large intersection sets of 1D invariant manifolds: a tool for blender detection. Numerical Algorithms, 96(3):1079–1108, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01812-0.

Coomes:1997:LPS

[CKP97] Brian A. Coomes, Hüseyin Koçak, and Kenneth J. Palmer. Long periodic shadowing. Numerical Algorithms, 14(1–3): 55–78, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/5/4/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/

5058/5/4/fulltext.pdf. Dynamical numerical analysis (Atlanta, GA, 1995).

Carnicer:2019:OIL

[CKP19] J. M. Carnicer, Y. Khiar, and J. M. Peña. Optimal interval length for the collocation of the Newton interpolation basis. Numerical Algorithms, 82(3):895–908, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Carnicer:2022:ICO

[CKP22] J. M. Carnicer, Y. Khiar, and J. M. Peña. Inverse central ordering for the Newton interpolation formula. Numerical Algorithms, 90(4):1691–1713, August 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01247-x.

Cvetkovic:2016:WCA

[CKS16] Ljiljana Cvetković, Vladimir Kostić, and Ernest Sanca. A wider convergence area for the MSTMAOR iteration methods for LCP. Numerical Algorithms, 71(1):77–88, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9985-6.

Chakib:2024:INA

[CKS24] Abdelkrim Chakib, Ibrahim Khalil, and Azeddine Sadik. An improved numerical approach for solving shape optimization problems on convex domains. Numerical Algorithms, 96 (2):621–663, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01660-4.

Chen:1999:GMG

[CKY99] Jen-Yuan Chen, David R. Kincaid, and David M. Young. Generalizations and modifications of the GMRES iterative method. Numerical Algorithms, 21(1–4):119–146, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/12/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
12/fulltext.pdf.

Cabay:1992:SAM

[CL92] Stan Cabay and George Labahn. A superfast algorithm for multidimensional Padé systems. Numerical Algorithms, 2(2): 201–224, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Clavero:1993:UCF

[CL93] Carmelo Clavero and Francisco Lisbona. Uniformly convergent finite difference methods for singularly perturbed problems with turning points. Numerical Algorithms, 4(4):339–359, May 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ciarlet:1996:DCP

[CL96a] P. Ciarlet, Jr. and F. Lamour. Does contraction preserve triangular meshes? Numerical Algorithms, 13(3–4):201–223, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ciarlet:1996:VFO

[CL96b] P. Ciarlet, Jr. and F. Lamour. On the validity of a frontoriented approach to partitioning large sparse graphs with a connectivity constraint. Numerical Algorithms, 12(1–2):193– 214, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cohen:1999:AIP

[CL99] Alan M. Cohen and David Levin. Accelerating infinite products. Numerical Algorithms, 22(2):157–165, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/22/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
22/3/fulltext.pdf.

Carre:2000:PLM

[CL00] G. Carré and S. Lanteri. Parallel linear multigrid by agglomeration for the acceleration of 3D compressible flow calculations on unstructured meshes. Numerical Algorithms, 24(4):309–332, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
29/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/29/2/fulltext.pdf. Sparse

matrices and differential equations (Rennes, 1999).

Chehab:2005:DES

[CL05] Jean-Paul Chehab and Jacques Laminie. Differential equations and solution of linear systems. Numerical Algorithms, 40(2):103–124, October 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=103.

Chen:2006:ERF

[CL06] Jinhai Chen and Weiguo Li. An exponential regula falsi method for solving nonlinear equations. Numerical Algorithms, 41(4):327–338, April 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=4&spage=327.

Cen:2010:RFD

[CL10a] Zhongdi Cen and Anbo Le. A robust finite difference scheme for pricing American put options with Singularity–Separating method. Numerical Algorithms, 53(4):497–510, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=4&
spage=497.

Cheng:2010:SDN

[CL10b] Wanyou Cheng and Qunfeng Liu. Sufficient descent nonlinear conjugate gradient methods with conjugacy condition. Numerical Algorithms, 53(1):113–131, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=1&spage=113.

Cardoso:2011:IFP

[CL11] João Ribeiro Cardoso and Ana F. Loureiro. Iteration functions for pth roots of complex numbers. Numerical Algorithms, 57(3):329–356, July 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=3&spage=329.

Chen:2013:SCM

[CL13a] Fang Chen and Qing-Quan Liu. On semi-convergence of modified HSS iteration methods. Numerical Algorithms, 64(3):507– 518, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9676-5.

Costabile:2013:ASR

[CL13b] F. A. Costabile and E. Longo. Δ_{h}-Appell sequences and related interpolation problem. Numerical Algorithms, 63(1):165– 186, May 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9619-1.

Chen:2019:SSC

[CL19] Jixin Chen and Ignace Loris. On starting and stopping criteria for nested primal-dual iterations. Numerical Algorithms, 82(2): 605–621, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2021:FIS

[CL21] Fang Chen and Tian-Yi Li. Fast and improved scaled HSS preconditioner for steady-state space-fractional diffusion equations. Numerical Algorithms, 87(2):651–665, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00982-x.

Chen:2024:EOR

[CL24a] Meng Chen and Leevan Ling. Exploring oversampling in RBF least-squares collocation method of lines for surface diffusion. Numerical Algorithms, 97(3):1067–1087, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01741-4.

Chu:2024:ESW

[CL24b] Moody T. Chu and Matthew M. Lin. On the enumeration of subcells within hypercubes and its application to the Borsuk– Ulam theorem. Numerical Algorithms, 97(2):595–615, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01716-5.

Chen:2011:NIF

[CLA11] S. Chen, F. Liu, and V. Anh. A novel implicit finite difference method for the one-dimensional fractional percolation equation. Numerical Algorithms, 56(4):517–535, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=517.

Chui:2000:ACM

[CLaL00] Charles K. Chui, Ming-Jun Lai, and Jian ao Lian. Algorithms for G^{1} connection of multiple parametric bicubic NURBS surfaces. Numerical Algorithms, 23(4):285–313, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/26/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
26/6/fulltext.pdf.

Cordero:2015:ETP

[CLBT15] Alicia Cordero, Taher Lotfi, Parisa Bakhtiari, and Juan R. Torregrosa. An efficient two-parametric family with memory for nonlinear equations. Numerical Algorithms, 68(2):323–335, February 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9846-8.

Chau:2017:GSP

[CLGS17] M. Chau, A. Laouar, T. Garcia, and P. Spiteri. Grid solution of problem with unilateral constraints. Numerical Algorithms, 75(4):879–908, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Casciola:2005:FSR

[CLMM05] G. Casciola, D. Lazzaro, L. B. Montefusco, and S. Morigi. Fast surface reconstruction and hole filling using positive definite radial basis functions. Numerical Algorithms, 39(1–3):289–305, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2023:ESO

[CLPY23] Rui Chen, Yaxiang Li, Kejia Pan, and Xiaofeng Yang. Efficient second-order, linear, decoupled and unconditionally energy stable schemes of the Cahn–Hilliard-Darcy equations for the Hele–Shaw flow. Numerical Algorithms, 92(4):2275–2306, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01388-7.

Chen:2026:RVE

[CLQ26] Yanping Chen, Jian Lei, and Fangfang Qin. β-robust virtual element method for distributed order time-fractional reaction–diffusion equation with variable coefficients. Numerical Algorithms, 101(3):1475–1503, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02052-6.

Chen:2013:NAV

[CLT^{+}13] Chang-Ming Chen, F. Liu, I. Turner, V. Anh, and Y. Chen. Numerical approximation for a variable-order nonlinear reaction-subdiffusion equation. Numerical Algorithms, 63(2): 265–290, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9622-6.

Chen:2010:NSM

[CLTA10] Chang-Ming Chen, Fawang Liu, Ian Turner, and Vo Anh. Numerical schemes and multivariate extrapolation of a twodimensional anomalous sub-diffusion equation. Numerical Algorithms, 54(1):1–21, May 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=1.

Chen:2020:TGM

[CLWH20] Yanping Chen, Qingfeng Li, Yang Wang, and Yunqing Huang. Two-grid methods of finite element solutions for semi-linear elliptic interface problems. Numerical Algorithms, 84(1):307– 330, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2015:EBL

[CLWV15] Tingting Chen, Wen Li, Xianping Wu, and Seakweng Vong. Error bounds for linear complementarity problems of MBmatrices. Numerical Algorithms, 70(2):341–356, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9950-9.

Chen:2025:UES

[CLY25] Yaoyao Chen, Hailiang Liu, and Peimeng Yin. Unconditionally energy stable IEQ-FEMs for the Cahn–Hilliard equation and Allen–Cahn equation. Numerical Algorithms, 99(3):1127– 1159, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01910-z.

Carstensen:1992:NAF

[CM92] C. Carstensen and G. Mühlbach. The Neville–Aitken formula for rational interpolants with prescribed poles. Numerical Algorithms, 3(1–4):133–141, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Chesneaux:1996:BNB

[CM96] Jean-Marie Chesneaux and Ana C. Matos. Breakdown and near-breakdown control in the CGS algorithm using stochastic arithmetic. Numerical Algorithms, 11(1–4):99–116, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Campbell:1998:MSNb

[CM98] Stephen L. Campbell and Wiesław Marszalek. Mixed symbolic-numerical computations with general DAEs. II. An applications case study. Numerical Algorithms, 19(1– 4):85–94, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
16/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/9/fulltext.pdf. Differen-

tial algebraic equations (Grenoble, 1997).

Cecchi:1999:AWB

[CM99] M. Morandi Cecchi and F. Marcuzzi. An approximate waves-bordering algorithm for adaptive finite elements analysis. Numerical Algorithms, 21(1–4):311–322, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/10/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
10/fulltext.pdf.

Calio:2001:ASG

[CM01] F. Calio’ and E. Marchetti. On an algorithm for the solution of generalized Prandtl equations. Numerical Algorithms, 28(1–4):3–10, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
37/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/1/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
1/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/1/fulltext.pdf.

Crampton:2005:DAF

[CM05] Andrew Crampton and John C. Mason. Detecting and approximating fault lines from randomly scattered data. Numerical Algorithms, 39(1–3):115–130, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cano:2015:NMN

[CM15] Alfredo Cano and Carlos Moreno. A new method for numerical integration of singular functions on the plane. Numerical Algorithms, 68(3):547–568, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9860-x.

Chen:2016:GHB

[CM16] Cairong Chen and Changfeng Ma. A generalization of the HSSbased sequential two-stage method for solving non-Hermitian saddle point problems. Numerical Algorithms, 73(4):1073– 1090, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0130-y.

Canak:2024:EAE

[CM24a] Melih Cem Canak and Gulcin M. Muslu. Error analysis of the exponential wave integrator sine pseudo-spectral method for the higher-order Boussinesq equation. Numerical Algorithms, 97(4):1583–1606, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01763-6.

Catinas:2024:SIS

[CM24b] Teodora Catinas and Andra Malina. Spherical interpolation of scattered data using least squares thin-plate spline and inverse multiquadric functions. Numerical Algorithms, 97(3): 1397–1414, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01755-6.

Chen:2025:GPR

[CM25] Fang Chen and Jin-Feng Mao. On greedy partially randomized extended Kaczmarz method for solving large sparse inconsistent linear systems. Numerical Algorithms, 100(1):311–338, September 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01962-1.

Caille:2019:MFR

[CMD19] Laëtitia Caillé, Liviu Marin, and Franck Delvare. A meshless fading regularization algorithm for solving the Cauchy problem for the three-dimensional Helmholtz equation. Numerical Algorithms, 82(3):869–894, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cariow:2015:FAD

[CMM15] Aleksandr Cariow and Dorota Majorkowska-Mech. Fast algorithm for discrete fractional Hadamard transform. Numerical Algorithms, 68(3):585–600, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-014-
9862-8.pdf.

Cerdan:2017:LRU

[CMM17] J. Cerdán, J. Marín, and J. Mas. Low-rank updates of balanced incomplete factorization preconditioners. Numerical Algorithms, 74(2):337–370, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0151-6.

Combettes:2019:LOI

[CMMP19] Patrick L. Combettes, Andrew M. McDonald, Charles A. Micchelli, and Massimiliano Pontil. Learning with optimal interpolation norms. Numerical Algorithms, 81(2):695–717, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Carnicer:2021:SPD

[CMP21] J. M. Carnicer, E. Mainar, and J. M. Peña. Stability properties of disk polynomials. Numerical Algorithms, 87(1):119– 135, May 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00960-3.

Czyzewska:2022:ASD

[CMP22] Natalia Czyzewska, Paweł M. Morkisz, and Paweł Przybyłowicz. Approximation of solutions of DDEs under nonstandard assumptions via Euler scheme. Numerical Algorithms, 91(4): 1829–1854, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01324-9.

Calvo:1993:CSS

[CMR93] M. Calvo, J. I. Montijano, and L. Rández. On the change of step size in multistep codes. Numerical Algorithms, 4(3):283– 304, March 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Calvo:2003:SDI

[CMR03] M. Calvo, J. I. Montijano, and L. Rández. On the solution of discontinuous IVPs by adaptive Runge–Kutta codes. Numerical Algorithms, 33(1–4):163–182, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/14/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
14/fulltext.pdf.

Castillo:2016:ZPO

[CMR16] K. Castillo, F. Marcellán, and M. N. Rebocho. Zeros of paraorthogonal polynomials and linear spectral transformations on the unit circle. Numerical Algorithms, 71(3):699–714, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0017-3.

Calvetti:2000:CEB

[CMRS00a] D. Calvetti, S. Morigi, L. Reichel, and F. Sgallari. Computable error bounds and estimates for the conjugate gradient method. Numerical Algorithms, 25(1–4):75–88, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/8/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/3/

8/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Calvetti:2000:RLU

[CMRS00b] D. Calvetti, S. Morigi, L. Reichel, and F. Sgallari. An L-ribbon for large underdetermined linear discrete ill-posed problems. Numerical Algorithms, 25(1–4):89–107, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/6/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/3/

6/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Conti:2001:CSD

[CMRS01] C. Conti, R. Morandi, C. Rabut, and A. Sestini. Cubic spline data reduction choosing the knots from a third derivative criterion. Numerical Algorithms, 28(1–4):45–61, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/37/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
37/4/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/4/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/37/4/fulltext.pdf.

Chousurin:2020:PCN

[CMWP20] R. Chousurin, T. Mouktonglang, B. Wongsaijai, and K. Poochinapan. Performance of compact and non-compact structure preserving algorithms to traveling wave solutions modeled by the Kawahara equation. Numerical Algorithms, 85(2):523–541, October 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00825-4.

Costabile:2001:MGA

[CN01] F. Costabile and A. Napoli. A method for global approximation of the initial value problem. Numerical Algorithms, 27(2):119– 130, June 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/35/1/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/35/
1/fulltext.pdf.

Chun:2016:NFO

[CN16] Changbum Chun and Beny Neta. On the new family of optimal eighth order methods developed by Lotfi et al. Numerical Algorithms, 72(2):363–376, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0048-9.

Chun:2017:CSE

[CN17] Changbum Chun and Beny Neta. Comparative study of eighth-order methods for finding simple roots of nonlinear equations. Numerical Algorithms, 74(4):1169–1201, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0191-y.

Ciurte:2015:SNA

[CNR15] Anca Ciurte, Sergiu Nedevschi, and Ioan Rasa. Systems of nonlinear algebraic equations with unique solution. Numerical Algorithms, 68(2):367–376, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9849-5.

Chan:1994:CPT

[CO94] Tony F. Chan and Julia A. Olkin. Circulant preconditioners for Toeplitz-block matrices. Numerical Algorithms, 6(1–2):89–101, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Celledoni:2003:ILG

[CO03] Elena Celledoni and Brynjulf Owren. On the implementation of Lie group methods on the Stiefel manifold. Numerical Algorithms, 32(2–4):163–183, April 2003. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
45/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/45/3/fulltext.pdf.

Curry:2019:VSS

[CO19] Charles Curry and Brynjulf Owren. Variable step size commutator free Lie group integrators. Numerical Algorithms, 82(4): 1359–1376, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Coleman:1992:RAC

[Col92] John P. Coleman. Rational approximations for the cosine function: p-acceptability and order. Numerical Algorithms, 3(1–4): 143–158, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Cong:1993:SDI

[Con93] Nguyen Huu Cong. A-stable diagonally implicit Runge–Kutta– Nyström methods for parallel computers. Numerical Algorithms, 4(3):263–281, March 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cools:2003:EAS

[Coo03] Ronald Cools. Extrapolation and adaptivity in software for automatic numerical integration on a cube. Numerical Algorithms, 34(2–4):259–269, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/20/fulltext.pdf.

Coope:2009:CGC

[Coo09] I. D. Coope. On calculating generalized Cayley transforms of bounded norm. Numerical Algorithms, 52(4):575–583, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=52&
issue=4&spage=575.

Cordellier:1991:UKA

[Cor91] Florent Cordellier. On the use of Kronecker’s algorithm in the generalized rational interpolation problem. Numerical Algorithms, 1(4):401–413, November 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Corless:2002:NVC

[Cor02] Robert M. Corless. A new view of the computational complexity of IVP for ODE. Numerical Algorithms, 31(1– 4):115–124, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
21/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/21/fulltext.pdf.

Corless:2024:HOM

[Cor24] Robert M. Corless. An Hermite–Obreshkov method for 2ndorder linear initial-value problems for ODE. Numerical Algorithms, 96(3):1109–1141, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01738-
z.

Chalhoub:2022:PEE

[COSE22] Nancy Chalhoub, Pascal Omnes, Toni Sayah, and Rebecca El Zahlaniyeh. A posteriori error estimates for the timedependent convection-diffusion-reaction equation coupled with the Darcy system. Numerical Algorithms, 89(3):1247–1286, March 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01152-3.

Courreges:2015:DDA

[Cou15a] Fabien Courreges. Design of a discrete algebraic robust differentiation FIR filter using an annihilator of the Ztransform; frequency response analysis and parameter tuning. Numerical Algorithms, 68(4):867–901, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9875-3. See erratum [Cou15b].

Courreges:2015:EDD

[Cou15b] Fabien Courreges. Erratum to: Design of a discrete algebraic robust differentiation FIR filter using an annihilator of the Z-transform; frequency response analysis and parameter tuning. Numerical Algorithms, 69(4):913, August 2015. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-015-0020-8.pdf. See [Cou15a].

Cox:1993:RDI

[Cox93] M. G. Cox. Reliable determination of interpolating polynomials. Numerical Algorithms, 5(1–4):133–154, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Ciarlini:1993:ASR

[CP93] P. Ciarlini and F. Pavese. Application of special reduction procedures to metrological data. Numerical Algorithms, 5(1– 4):479–489, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See erratum [CP95b].

Cecchi:1995:RAM

[CP95a] M. Morandi Cecchi and E. Pirozzi. A recursive algorithm by the moments method to evaluate a class of numerical integrals over an infinite interval. Numerical Algorithms, 10(1–2):155– 165, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Ciarlini:1995:EAS

[CP95b] P. Ciarlini and F. Pavese. Erratum to “Application of special reduction procedures to metrological data” [Numer. Algorithms 5 (1993), no. 1–4, 479–489]. Numerical Algorithms, 10 (3–4):421, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See [CP93].

Chartier:2000:P

[CP00a] P. Chartier and B. Philippe. Preface. Numerical Algorithms, 24(4):0, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
29/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/29/1/fulltext.pdf.

Chartier:2000:SMD

[CP00b] P. Chartier and B. Philippe, editors. Sparse matrices and differential equations, volume 24(4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 2000. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ???? Papers from the 3rd Seminar on Computational Techniques Applied to Industrial Problems (Sparse Matrices and Differential Equations) held in Rennes, March 11–12, 1999.

Capizzano:2001:FEM

[CP01a] Stefano Serra Capizzano and Cristina Tablino Possio. Finite element matrix sequences: the case of rectangular domains. Numerical Algorithms, 28(1–4):309–327, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/37/18/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/37/
18/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/18/abstract.htm;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/18/
fulltext.pdf.

Costantini:2001:SPA

[CP01b] Paolo Costantini and Francesca Pelosi. Shape-preserving approximation by space curves. Numerical Algorithms, 27(3):237–264, July 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.lwwonline.com/content/getfile/5058/34/4/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/34/4/fulltext.pdf.

Cornelio:2014:CNO

[CPN14] Anastasia Cornelio, Elena Loli Piccolomini, and James G. Nagy. Constrained numerical optimization methods for blind deconvolution. Numerical Algorithms, 65(1):23–42, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9693-z.

Censor:2014:FCI

[CPP14] Yair Censor, Ioana Pantelimon, and Constantin Popa. Family constraining of iterative algorithms. Numerical Algorithms, 66 (2):323–338, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9736-5.

Cipolla:2023:LTP

[CPRZV23] Stefano Cipolla, Stefano Pozza, Michela Redivo-Zaglia, and Niel Van Buggenhout. A Lanczos-type procedure for tensors. Numerical Algorithms, 92(1):377–406, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01351-6.

Cravero:2012:ASL

[CPS12] Isabella Cravero, Giovanna Pittaluga, and Laura Sacripante. An algorithm for solving linear Volterra integro-differential equations. Numerical Algorithms, 60(1):101–114, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=1&
spage=101.

Codevico:2004:NLI

[CPV04] Gianni Codevico, Victor Y. Pan, and Marc Van Barel. Newtonlike iteration based on a cubic polynomial for structured matrices. Numerical Algorithms, 36(4):365–380, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2014:ARP

[CPZ14] K. Chen, E. Loli Piccolomini, and F. Zama. An automatic regularization parameter selection algorithm in the total variation model for image deblurring. Numerical Algorithms, 67(1): 73–92, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9775-y.

Cai:2016:LGM

[CQ16] Haotao Cai and Jiafei Qi. A Legendre–Galerkin method for solving general Volterra functional integral equations. Numerical Algorithms, 73(4):1159–1180, December 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-016-0134-7.

Chen:2015:TDR

[CQLY15] Zhaoxia Chen, Zeyu Qiu, Juan Li, and Xiong You. Twoderivative Runge–Kutta-Nyström methods for second-order ordinary differential equations. Numerical Algorithms, 70(4): 897–927, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9979-4.

Calvetti:1996:ARI

[CR96a] D. Calvetti and L. Reichel. An adaptive Richardson iteration method for indefinite linear systems. Numerical Algorithms, 12(1–2):125–149, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Calvetti:1996:HIM

[CR96b] D. Calvetti and L. Reichel. A hybrid iterative method for symmetric positive definite linear systems. Numerical Algorithms, 11(1–4):79–98, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Calvetti:1999:BLM

[CR99] D. Calvetti and L. Reichel. A block-Lanczos method for large continuation problems. Numerical Algorithms, 21(1– 4):109–118, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
15/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/15/fulltext.pdf.

Calvetti:2000:MJT

[CR00] Daniela Calvetti and Lothar Reichel, editors. Mathematical journey through analysis, matrix theory and scientific computation, volume 25(1–4) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 2000. CODEN NUALEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (electronic). LCCN ???? Papers from the conference held in honor of Richard S. Varga on the occasion of his 70th birthday at Kent State University, Kent, OH, March 25–27,

Calvetti:2002:LBE

[CR02] D. Calvetti and L. Reichel. Lanczos-based exponential filtering for discrete ill-posed problems. Numerical Algorithms, 29(1–3):45–65, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp008.lwwonline.com/content/getfile/5058/38/4/
abstract.htm;   http://ipsapp008.lwwonline.com/content/
getfile/5058/38/4/fulltext.pdf.

Calvetti:2003:GQA

[CR03a] Daniela Calvetti and Lothar Reichel. Gauss quadrature applied to trust region computations. Numerical Algorithms, 34(1):85–102, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.kluweronline.com/content/getfile/5058/
47/6/abstract.htm;   http://ipsapp008.kluweronline.
com/content/getfile/5058/47/6/fulltext.pdf.

Calvetti:2003:EPC

[CR03b] Daniela Calvetti and Lothar Reichel. On the evaluation of polynomial coefficients. Numerical Algorithms, 33 (1–4):153–161, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
13/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/13/fulltext.pdf.

Campos:2012:SSS

[CR12] Carmen Campos and Jose E. Roman. Strategies for spectrum slicing based on restarted Lanczos methods. Numerical Algorithms, 60(2):279–295, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=2&spage=279.

Chan:2014:SEI

[CR14] R. P. K. Chan and N. Razali. Smoothing effects on the IMR and ITR. Numerical Algorithms, 65(3):401–420, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9779-7.

Chehab:2020:GIM

[CR20] Jean-Paul Chehab and Marcos Raydan. Geometrical inverse matrix approximation for least-squares problems and acceleration strategies. Numerical Algorithms, 85(4):1213–1231, December 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00862-z.

Chenouf:2023:HGO

[CR23] Chahinaz Chenouf and Mohamed Rahal. On Hölder global optimization method using piecewise affine bounding functions. Numerical Algorithms, 94(2):905–935, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01524-x.

Cordero:2024:EPF

[CRHTTN24] Alicia Cordero, Renso V. Rojas-Hiciano, Juan R. Torregrosa, and Paula Triguero-Navarro. Efficient parametric family of fourth-order Jacobian-free iterative vectorial schemes. Numerical Algorithms, 97(4):2011–2029, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01776-1.

Cordero:2024:HEC

[CRHTV24] Alicia Cordero, Renso V. Rojas-Hiciano, Juan R. Torregrosa, and Maria P. Vassileva. A highly efficient class of optimal fourth-order methods for solving nonlinear systems. Numerical Algorithms, 95(4):1879–1904, April 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01631-9.

Chidume:2019:IAS

[CRN19] C. E. Chidume, O. M. Romanus, and U. V. Nnyaba. An iterative algorithm for solving split equality fixed point problems for a class of nonexpansive-type mappings in Banach spaces. Numerical Algorithms, 82(3):987–1007, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Croft:1992:ACA

[Cro92] A. Croft. An application of convergence acceleration techniques to a class of two-point boundary value problems on a semi-infinite domain. Numerical Algorithms, 2(3–4):307–320, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Crouzeix:2003:CAS

[Cro03] Michel Crouzeix. Contractivity and analyticity in l^{p} of some approximation of the heat equation. Numerical Algorithms, 33(1–4):193–201, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
16/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/16/fulltext.pdf.

Calvetti:2004:CCB

[CRS04] D. Calvetti, L. Reichel, and A. Shuibi. L-curve and curvature bounds for Tikhonov regularization. Numerical Algorithms, 35(2–4):301–314, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/50/A/7/
abstract.htm.

Csordas:1991:LIA

[CRV91] G. Csordas, A. Ruttan, and R. S. Varga. The Laguerre inequalities with applications to a problem associated with the Riemann hypothesis. Numerical Algorithms, 1(3):305–329, September 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cegielski:2026:RPT

[CRZ26] Andrzej Cegielski, Simeon Reich, and Rafał Zalas. Regularity of the product of two relaxed cutters with relaxation parameters beyond two. Numerical Algorithms, 101(3):1809–1831, March 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02064-2.

Chan:1994:CSC

[CS94] Tony F. Chan and Tedd Szeto. A composite step conjugate gradients squared algorithm for solving nonsymmetric linear systems. Numerical Algorithms, 7(1):17–32, June 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Celorrio:1999:FCM

[CS99] Ricardo Celorrio and Francisco-Javier Sayas. Full collocation methods for some boundary integral equations. Numerical Algorithms, 22(3–4):327–351, December 1999. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/23/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/5/fulltext.pdf.

Cvijovic:2008:CFE

[CS08] Djurdje Cvijović and H. M. Srivastava. Closed-form evaluations of certain definite integrals by employing the Cauchy integral theorem. Numerical Algorithms, 49(1–4):129–141, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=49&
issue=1&spage=129.

Chernogorova:2012:CAA

[CS12] Tatiana Paraskevova Chernogorova and Beata Stehlíková. A comparison of asymptotic analytical formulae with finitedifference approximations for pricing zero coupon bond. Numerical Algorithms, 59(4):571–588, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=571.

Cardoso:2018:CMM

[CS18] João R. Cardoso and Amir Sadeghi. Conditioning of the matrix–matrix exponentiation. Numerical Algorithms, 79(2): 457–477, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2022:SOM

[CS22] Hao Chen and Hai-Wei Sun. Second-order maximum principle preserving Strang’s splitting schemes for anisotropic fractional Allen–Cahn equations. Numerical Algorithms, 90(2):749–771, June 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01207-5.

Chen:2024:TMV

[CS24] Qian Chen and Zhengwei Shen. A two-metric variable scaled forward-backward algorithm for ℓ_{0} optimization problem and its applications. Numerical Algorithms, 97(1):191–221, September 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01700-z.

Chakir:2025:SWS

[CS25] Y. Chakir and H. Safouhi. On solving 2D weakly singular Volterra integral equations of the second kind. Numerical Algorithms, 98(4):1827–1847, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01854-
4.

Chen:2026:EIA

[CSC26] Zebin Chen, Hui-Jie Sun, and Xuesong Chen. Explicit iterative algorithm with optimal tuning parameter for stochastic Lyapunov matrix equation. Numerical Algorithms, 101 (1):679–702, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02019-7.

Csendes:2004:GSS

[Cse04] Tibor Csendes. Generalized subinterval selection criteria for interval global optimization. Numerical Algorithms, 37(1–4):93–100, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/37/abstract.htm.

Correa:2004:GAS

[CSFC04] Elon Santos Correa, Maria Teresinha A. Steiner, Alex A. Freitas, and Celso Carnieri. A genetic algorithm for solving a capacitated p-median problem. Numerical Algorithms, 35(2–4):373–388, April 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/50/A/10/abstract.htm.

Cai:2016:IAS

[CSI16] Gang Cai, Yekini Shehu, and Olaniyi Samuel Iyiola. Iterative algorithms for solving variational inequalities and fixed point problems for asymptotically nonexpansive mappings in Banach spaces. Numerical Algorithms, 73(3):869–906, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0121-z.

Cai:2017:VIA

[CSI17] Gang Cai, Yekini Shehu, and Olaniyi Samuel Iyiola. Viscosity iterative algorithms for fixed point problems of asymptotically nonexpansive mappings in the intermediate sense and variational inequality problems in Banach spaces. Numerical Algorithms, 76(2):521–553, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cai:2018:SCR

[CSI18] Gang Cai, Yekini Shehu, and Olaniyi Samuel Iyiola. Strong convergence results for variational inequalities and fixed point problems using modified viscosity implicit rules. Numerical Algorithms, 77(2):535–558, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-017-0327-8.

Cuyt:2024:MMP

[CsL24] Annie Cuyt and Wen shin Lee. Multiscale matrix pencils for separable reconstruction problems. Numerical Algorithms, 95(1):31–72, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01564-3.

Chen:2022:SEE

[CSZ22] Hu Chen, Yanhua Shi, and Yanmin Zhao. Sharp error estimate of a Grünwald–Letnikov scheme for reaction–subdiffusion equations. Numerical Algorithms, 89(4):1465–1477, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01161-2.

Chan:1993:CMA

[CT93] Raymond H. Chan and Ping Tak Peter Tang. Constrained minimax approximation and optimal preconditioners for Toeplitz matrices. Numerical Algorithms, 5(1–4):353–364, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Cvetkovic:2006:MOE

[CT06] Ljiljana Cvetković and Djurdjica Takaci. On matrices with operator entries. Numerical Algorithms, 42(3–4):335–344, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=42&issue=3&
spage=335.

Chan:2010:ETD

[CT10] Robert P. K. Chan and Angela Y. J. Tsai. On explicit two-derivative Runge–Kutta methods. Numerical Algorithms, 53(2–3):171–194, March 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=171.

Coope:2021:GDH

[CT21] Ian D. Coope and Rachael Tappenden. Gradient and diagonal Hessian approximations using quadratic interpolation models and aligned regular bases. Numerical Algorithms, 88(2): 767–791, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01056-8.

Chau:2009:PSA

[CTS09] M. Chau, C. Tauber, and P. Spiteri. Parallel Schwarz alternating methods for anisotropic diffusion of speckled medical images. Numerical Algorithms, 51(1):85–114, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=1&

spage=85. Tributes to Gene H. Golub, Part I.

Chen:2025:ISD

[CTY25] Wang Chen, Liping Tang, and Xinmin Yang. Improvements to steepest descent method for multi-objective optimization. Numerical Algorithms, 100(3):1407–1433, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01993-8.

Cui:2013:CAH

[Cui13] Mingrong Cui. Convergence analysis of high-order compact alternating direction implicit schemes for the two-dimensional time fractional diffusion equation. Numerical Algorithms, 62 (3):383–409, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9589-3/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=3&spage=383-409.

Cui:2026:GCS

[Cui26] Haijuan Cui. Global convergence of a spectral conjugate gradient method for unconstrained optimization. Numerical Algorithms, 101(2):1049–1072, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02032-w.

Cuyt:2000:CMR

[Cuy00] Annie Cuyt, editor. Computational methods from rational approximation theory, volume 24(1–2) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, December 2000. CODEN NUA-LEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). LCCN ???? URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/1/fulltext.pdf. Papers from the International Conference on Rational Approximation (ICRA99) held at the University of Antwerp, Wilrijk, June 6–11, 1999.

Chen:2024:CGF

[CUY24a] Qiyuan Chen, Jeffrey Uhlmann, and Ke Ye. Correction to: Generalizing Frobenius inversion to quaternion matrices. Numerical Algorithms, 97(1):117, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01733-4. See [CUY24b].

Chen:2024:GFI

[CUY24b] Qiyuan Chen, Jeffrey Uhlmann, and Ke Ye. Generalizing Frobenius inversion to quaternion matrices. Numerical Algorithms, 97(1):101–116, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01694-8. See correction [CUY24a].

Cuyt:1992:MRD

[CV92] Annie Cuyt and Brigitte Verdonk. Multivariate rational data fitting: general data structure, maximal accuracy and object orientation. Numerical Algorithms, 3(1–4):159–172, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Chernogorova:2015:FFV

[CV15] Tatiana Chernogorova and Lubin Lubin Vulkov. Fitted finite volume positive difference scheme for a stationary model of air pollution. Numerical Algorithms, 70(1):171–189, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9940-y.

Ciaramella:2022:STG

[CV22] G. Ciaramella and T. Vanzan. Substructured two-grid and multi-grid domain decomposition methods. Numerical Algorithms, 91(1):413–448, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01268-
0.

Calvo:2001:NSS

[CVA01] Manuel Calvo and Jesús Vigo-Aguiar. A note on the step size selection in Adams multistep methods. Numerical Algorithms, 27(4):359–366, August 2001. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp007.lwwonline.com/content/
getfile/5058/36/3/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/36/3/fulltext.pdf.

Cvetkovic:2006:MTV

[Cve06] Ljiljana Cvetković. H-matrix theory vs. eigenvalue localization. Numerical Algorithms, 42(3–4):229–245, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=42&issue=3&
spage=229.

Chen:2019:NIG

[CVLX19] Xiao Shan Chen, Seak-Weng Vong, Wen Li, and Hongguo Xu. Noda iterations for generalized eigenproblems following Perron–Frobenius theory. Numerical Algorithms, 80(3):937– 955, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chernov:2015:QAH

[CvPS15] Alexey Chernov, Tobias von Petersdorff, and Christoph Schwab. Quadrature algorithms for high dimensional singular integrands on simplices. Numerical Algorithms, 70(4):847– 874, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9977-6.

Constantinou:2016:FEM

[CVX16] Philippos Constantinou, Charalambia Varnava, and Christos Xenophontos. An hp finite element method for 4-th order singularly perturbed problems. Numerical Algorithms, 73(2): 567–590, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0108-9.

Chu:2008:ALS

[CW08] Wenchang Chu and Xiaoyuan Wang. Abel’s lemma on summation by parts and terminating q-series identities. Numerical Algorithms, 49(1–4):105–128, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=105.

Chen:2014:DDC

[CW14] Fengmin Chen and Patricia J. Y. Wong. Deficient discrete cubic spline solution for a system of second order boundary value problems. Numerical Algorithms, 66(4):793–809, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9763-2.

Chow:2017:SLM

[CW17] S.-S. Chow and A. Washburn. A shooting like method for non-Darcian seepage flow problems. Numerical Algorithms, 74 (4):951–966, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0179-7.

Cao:2019:TSM

[CW19a] Yang Cao and An Wang. Two-step modulus-based matrix splitting iteration methods for implicit complementarity problems. Numerical Algorithms, 82(4):1377–1394, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2019:MNM

[CW19b] Min-Hong Chen and Qing-Biao Wu. Modified Newton– MDPMHSS method for solving nonlinear systems with block two-by-two complex symmetric Jacobian matrices. Numerical Algorithms, 80(2):355–375, February 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cao:2021:PIM

[CW21a] Shan-Mou Cao and Zeng-Qi Wang. PMHSS iteration method and preconditioners for Stokes control PDE-constrained optimization problems. Numerical Algorithms, 87(1):365–380, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00970-1.

Chudy:2021:FAE

[CW21b] Filip Chudy and Paweł Woźny. Fast and accurate evaluation of dual Bernstein polynomials. Numerical Algorithms, 87(3): 1001–1015, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00996-5.

Cho:2025:CNB

[CW25] Chien-Hong Cho and Jhih-Sin Wu. On the convergence of the numerical blow-up time for a rescaling algorithm.

Numerical Algorithms, 98(1):191–217, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01791-2.

Chen:2020:CAP

[CWHL20] Jiawei Chen, Yiyun Wang, Hongjin He, and Yibing Lv. Convergence analysis of positive-indefinite proximal ADMM with a Glowinski’s relaxation factor. Numerical Algorithms, 83(4): 1415–1440, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2016:SCA

[CWL16] Minhong Chen, Qingbiao Wu, and Rongfei Lin. Semilocal convergence analysis for the modified Newton–HSS method under the Hölder condition. Numerical Algorithms, 72(3):667– 685, July 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0061-z.

Cai:2013:CAN

[CWZ13] Xinzhong Cai, Guoqiang Wang, and Zihou Zhang. Complexity analysis and numerical implementation of primal-dual interiorpoint methods for convex quadratic optimization based on a finite barrier. Numerical Algorithms, 62(2):289–306, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9581-y/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=289-306.

Chen:2020:EPR

[CX20] Hao Chen and Dongping Xu. Efficient preconditioners for Radau–IIA time discretization of space fractional diffusion equations. Numerical Algorithms, 83(4):1349–1372, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cen:2025:GSF

[CX25] Xiaoli Cen and Yong Xia. Globally solving the fractional squared least squares model for GPS localization. Numerical Algorithms, 99(4):1829–1853, August 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01935-4.

Chang:2025:SPP

[CXC25] Xiaokai Chang, Long Xu, and Jianxiong Cao. A splitting preconditioned primal-dual algorithm with interpolation and extrapolation for bilinear saddle point problem. Numerical Algorithms, 100(3):937–962, November 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01974-x.

Cen:2016:HFD

[CXL16] Zhongdi Cen, Aimin Xu, and Anbo Le. On the hybrid finite difference scheme for a singularly perturbed Riccati equation. Numerical Algorithms, 71(2):417–436, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0001-y.

Cuyt:2010:PEF

[CY10] Annie Cuyt and Xianglan Yang. A practical error formula for multivariate rational interpolation and approximation. Numerical Algorithms, 55(2–3):233–243, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=233.

Chen:2019:EET

[CY19] Minghua Chen and Wenshan Yu. Energy estimates for twodimensional space-Riesz fractional wave equation. Numerical Algorithms, 80(3):989–1014, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cuyt:2012:ROL

[CYIB12] Annie Cuyt, Irem Yaman, Bayram Ali Ibrahimoglu, and Brahim Benouahmane. Radial orthogonality and Lebesgue constants on the disk. Numerical Algorithms, 61(2):291–313, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=2&spage=291.

Cheng:2022:BNE

[CYM22] Yao Cheng, Li Yan, and Yanjie Mei. Balanced-norm error estimate of the local discontinuous Galerkin method on layer-adapted meshes for reaction-diffusion problems. Numerical Algorithms, 91(4):1597–1626, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01316-9.

Cui:2025:TAP

[CYWH25] Zixuan Cui, Lei Yang, Jing Wu, and Guanghui Hu. A treecode algorithm for the Poisson equation in a general domain with unstructured grids. Numerical Algorithms, 99(2):549–571, June 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01888-8.

Chan:1994:ASD

[CZ94] Tony F. Chan and Jun Zou. Additive Schwarz domain decomposition methods for elliptic problems on unstructured meshes. Numerical Algorithms, 8(2–4):329–346, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See erratum [CZ95].

Chan:1995:EAS

[CZ95] Tony F. Chan and Jun Zou. Erratum to: “Additive Schwarz domain decomposition methods for elliptic problems on unstructured meshes” [Numer. Algorithms 8 (1994), no. 2–4, 329– 346, MR 95i:65174]. Numerical Algorithms, 9(3–4):397, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See [CZ94].

Chan:1996:CTM

[CZ96] Tony F. Chan and Jun Zou. A convergence theory of multilevel additive Schwarz methods on unstructured meshes. Numerical Algorithms, 13(3–4):365–398, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chao:2014:GPH

[CZ14] Zhen Chao and Naimin Zhang. A generalized preconditioned HSS method for singular saddle point problems. Numerical Algorithms, 66(2):203–221, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9730-y.

Chen:2020:CDZ

[CZ20] Jianrong Chen and Yunong Zhang. Continuous and discrete zeroing neural dynamics handling future unknown-transpose matrix inequality as well as scalar inequality of linear class. Numerical Algorithms, 83(2):529–547, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cai:2023:UAN

[CZ23] Yongyong Cai and Xuanxuan Zhou. Uniformly accurate nested Picard iterative integrators for the Klein–Gordon– Schrödinger equation in the nonrelativistic regime. Numerical Algorithms, 94(1):371–396, September 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01505-0.

Cao:2022:AEB

[CZH22] Huanxin Cao, Hongchan Zheng, and Gang Hu. Adjusting the energy of Ball surfaces by modifying unfixed control balls. Numerical Algorithms, 89(2):749–768, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01132-7.

Cai:2026:QMC

[CZL26] Xiaomin Cai, Wenxin Zhuo, and Riwei Liao. Quaternion modified conjugate gradient algorithm for η (-anti)-hermitian solutions of generalized coupled Sylvester-type quaternion matrix equations. Numerical Algorithms, 101(4):2357–2379, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02090-0.

Cui:2018:LCP

[CZLS18] Jifeng Cui, Wenyu Zhang, Zeng Liu, and Jianglong Sun. On the limit cycles, period-doubling, and quasi-periodic solutions of the forced Van der Pol–Duffing oscillator. Numerical Algorithms, 78(4):1217–1231, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Chen:2021:TSM

[CZM21] Fang Chen, Yu Zhu, and Galina V. Muratova. Two-step modulus-based matrix splitting iteration methods for retinex problem. Numerical Algorithms, 88(4):1989–2005, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01103-y.

Cui:2026:LFM

[CZX26] Angang Cui, Qingyong Zhu, and Shengli Xue. Logarithmic function minimization to compressed sensing with application to magnetic resonance imaging. Numerical Algorithms, 101 (3):2043–2067, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02078-w.

Chen:2025:STG

[CZZ^{+}25] Hao Chen, Mahmoud A. Zaky, Xiangcheng Zheng, Ahmed S. Hendy, and Wenlin Qiu. Spatial two-grid compact difference method for nonlinear Volterra integro-differential equation with Abel kernel. Numerical Algorithms, 98(2):677–718, February 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01811-1.

DaRocha:1992:IRR

[Da 92] Z. Da Rocha. Implementation of the recurrence relations of biorthogonality. Numerical Algorithms, 3(1–4):173–183, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Dabrowska:2004:JME

[Dab04] Dorota Dabrowska. Jitter and measurement errors in approximation and integration of Lipschitz functions. Numerical Algorithms, 35(1):45–60, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/49/A/3/abstract.htm.

deAndrade:2001:GQR

[dABR01] E. X. L. de Andrade, C. F. Bracciali, and A. Sri Ranga. Gaussian quadrature rules with simple node-weight relations. Numerical Algorithms, 27(1):61–76, May 2001. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp007.lwwonline.com/content/
getfile/5058/33/3/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/33/3/fulltext.pdf.

deAguiar:2004:IIT

[dADdRC04] Marilton Sanchotene de Aguiar, Graçaliz Pereira Dimuro, and Antônio Carlos da Rocha Costa. ICTM: An interval tessellation-based model for reliable topographic segmentation. Numerical Algorithms, 37(1–4):3–11, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/1/abstract.htm.

deAlba:2023:RMN

[dAFPR23] Patricia Díaz de Alba, Luisa Fermo, Federica Pes, and Giuseppe Rodriguez. Regularized minimal-norm solution of an overdetermined system of first kind integral equations. Numerical Algorithms, 92(1):471–502, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01282-2.

Dahmen:1993:DRS

[Dah93] Wolfgang Dahmen. Decomposition of refinable spaces and applications to operator equations. Numerical Algorithms, 5(1– 4):229–245, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Damelin:2008:WTE

[Dam08] S. B. Damelin. A walk through energy, discrepancy, numerical integration and group invariant measures on measurable subsets of Euclidean space. Numerical Algorithms, 48(1–3):213–235, July 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=213.

Dehghan:2016:AMM

[DAM16] Mehdi Dehghan, Mostafa Abbaszadeh, and Akbar Mohebbi. Analysis of a meshless method for the time fractional diffusionwave equation. Numerical Algorithms, 73(2):445–476, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0103-1.

Deka:2026:GFG

[DAN26] Mandeep Deka, Ashwani Assam, and Ganesh Natarajan. A generalized formulation for gradient reconstruction schemes in unstructured finite volume method. Numerical Algorithms, 102 (2):1403–1431, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02181-y.

Daras:1999:RAH

[Dar99] Nicholas J. Daras. Rational approximation to harmonic functions. Numerical Algorithms, 20(4):285–301, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/19/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
19/5/fulltext.pdf.

deAndrade:2006:MCA

[dAR06] E. X. L. de Andrade and A. Sri Ranga. Modified Chebyshev algorithm: some applications. Numerical Algorithms, 43(3):215–233, November 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=3&spage=215.

Das:2019:PBC

[Das19] Pratibhamoy Das. An a posteriori based convergence analysis for a nonlinear singularly perturbed system of delay differential equations on an adaptive mesh. Numerical Algorithms, 81 (2):465–487, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dassi:2026:VCL

[Das26] Franco Dassi. Vem++, a C++ library to handle and play with the virtual element method. Numerical Algorithms, 101 (3):1633–1675, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02059-z.

Dax:2009:HAS

[Dax09] Achiya Dax. A hybrid algorithm for solving linear inequalities in a least squares sense. Numerical Algorithms, 50(2):97–114, February 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=97.

Dax:2017:GRI

[Dax17] Achiya Dax. A gradual rank increasing process for matrix completion. Numerical Algorithms, 76(4):959–976, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DeSamblanx:1998:NLI

[DB98] Gorik De Samblanx and Adhemar Bultheel. Nested Lanczos: Implicitly restarting an unsymmetric Lanczos algorithm. Numerical Algorithms, 18(1):31–50, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/13/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
13/3/fulltext.pdf.

Doha:2006:ESG

[DB06] E. H. Doha and A. H. Bhrawy. Efficient spectral-Galerkin algorithms for direct solution for second-order differential equations using Jacobi polynomials. Numerical Algorithms, 42 (2):137–164, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=42&issue=2&spage=137.

deBoor:2007:MPI

[dB07a] Carl de Boor. Multivariate polynomial interpolation: conjectures concerning GC-sets. Numerical Algorithms, 45(1–4):113– 125, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
45&issue=1&spage=113.

deBruin:2007:CCT

[dB07b] Marcel G. de Bruin. ‘classical’ convergence theorems for generalized continued fractions. Numerical Algorithms, 44(4):367– 380, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
44&issue=4&spage=367.

Doha:2009:JSG

[DBAE09] E. H. Doha, A. H. Bhrawy, and W. M. Abd-Elhameed. Jacobi spectral Galerkin method for elliptic Neumann problems. Numerical Algorithms, 50(1):67–91, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=1&spage=67.

deBruin:2005:PTB

[dBD05] M. G. de Bruin and H. P. Dikshit. Pál-type Birkhoff interpolation on nonuniformly distributed points. Numerical Algorithms, 40(1):1–16, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=1&spage=1.

Dookhitram:2011:PRB

[DBGB11] Kumar Dookhitram, Ravindra Boojhawon, Ashvin Gopaul, and Muddun Bhuruth. A-posteriori residual bounds for Arnoldi’s methods for nonsymmetric eigenvalue problems. Numerical Algorithms, 56(4):481–495, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=4&spage=481.

deBoor:2008:MAT

[dBGKR08] Carl de Boor, Christian Gout, Angela Kunoth, and Christophe Rabut. Multivariate approximation: theory and applications. An overview. Numerical Algorithms, 48(1–3):1–9, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=1&
spage=1.

Dehghani:2021:UNF

[DBH21] R. Dehghani, N. Bidabadi, and M. M. Hosseini. Using nonlinear functions to approximate a new quasi-Newton method for unconstrained optimization problems. Numerical Algorithms, 87(2):755–777, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00986-7.

Dallerit:2024:SOR

[DBTG24] Valentin Dallerit, Tommaso Buvoli, Mayya Tokman, and Stéphane Gaudreault. Second-order Rosenbrock-exponential (ROSEXP) methods for partitioned differential equations. Numerical Algorithms, 96(3):1143–1161, July 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01698-4.

Dewaele:2023:CNM

[DBV23] Nick Dewaele, Paul Breiding, and Nick Vannieuwenhoven. The condition number of many tensor decompositions is invariant under Tucker compression. Numerical Algorithms, 94(2): 1003–1029, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01526-9.

deCamargo:2016:ENS

[dC16a] André Pierro de Camargo. Erratum: “On the numerical stability of Floater–Hormann’s rational interpolant”. Numerical Algorithms, 72(1):153–154, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-015-

0071-x.pdf. See [dC16b].

deCamargo:2016:NSF

[dC16b] André Pierro de Camargo. On the numerical stability of Floater–Hormann’s rational interpolant. Numerical Algorithms, 72(1):131–152, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0037-z. See

erratum [dC16a].

Delvare:2017:UDH

[DC17] Franck Delvare and Alain Cimetière. Unique discrete harmonic continuation and data completion problems using the fading regularization method. Numerical Algorithms, 75(3):731–751, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

deCamargo:2020:REA

[dC20] André Pierro de Camargo. Rounding error analysis of divided differences schemes: Newton’s divided differences; Neville’s algorithm; Richardson extrapolation; Romberg quadrature; etc. Numerical Algorithms, 85(2):591–606, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00828-1.

deCamargo:2022:BFS

[dC22] André Pierro de Camargo. Backward and forward stability analysis of Neville’s algorithm for interpolation and a pyramid algorithm for the computation of Lebesgue functions. Numerical Algorithms, 89(4):1521–1531, April 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01163-0.

deCamargo:2023:CMS

[dC23] André Pierro de Camargo. A Chebyshev–Markov–Stieltjes separation type theorem for classical Romberg quadrature. Numerical Algorithms, 92(4):2365–2376, April 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01393-w.

Dang:2025:MBR

[DCG25] Yazheng Dang, Liyuan Chen, and Yan Gao. Multi-block relaxed-dual linear inertial ADMM algorithm for nonconvex and nonsmooth problems with nonseparable structures. Numerical Algorithms, 98(1):251–285, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01793-0.

DAmore:2013:ERA

[DCM^{+}13] Luisa D’Amore, Rosanna Campagna, Valeria Mele, Almerico Murli, and Mariarosaria Rizzardi. Erratum to: ReLaTIve. An ANSI C90 software package for the Real Laplace Transform Inversion. Numerical Algorithms, 63(3):571, July 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-013-9725-8.pdf. See [DCMM13].

DAmore:2013:RAC

[DCMM13] Luisa D’Amore, Rosanna Campagna, Valeria Mele, and Almerico Murli. ReLaTIve. an ANSI C90 software package for the Real Laplace Transform Inversion. Numerical Algorithms, 63(1):187–211, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9636-0. See

erratum [DCM^{+}13].

daCunha:2021:MAG

[dCOS21] Rudnei D. da Cunha, Elismar R. Oliveira, and Filip Strobin. A multiresolution algorithm to generate images of generalized fuzzy fractal attractors. Numerical Algorithms, 86(1): 223–256, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00886-w.

Deng:2023:EPD

[DCW23] Dingwen Deng, Jingliang Chen, and Qihong Wang. Energypreserving Du Fort–Frankel difference schemes for solving sine-Gordon equation and coupled sine-Gordon equations. Numerical Algorithms, 93(3):1045–1081, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01453-1.

Driver:1999:AZD

[DD99] Kathy Driver and Peter Duren. Asymptotic zero distribution of hypergeometric polynomials. Numerical Algorithms, 21(1–4):147–156, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/8/fulltext.pdf.

Dang:2020:ERI

[DD20] Quang A Dang and Quang Long Dang. Existence results and iterative method for a fully fourth-order nonlinear integral boundary value problem. Numerical Algorithms, 85(3): 887–907, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00842-3.

Dang:2021:SNM

[DD21] Quang A Dang and Quang Long Dang. Simple numerical methods of second- and third-order convergence for solving a fully third-order nonlinear boundary value problem. Numerical Algorithms, 87(4):1479–1499, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01016-2.

Deidda:2020:FMP

[DdAF^{+}20] G. P. Deidda, P. Díaz de Alba, C. Fenu, G. Lovicu, and G. Rodriguez. FDEMtools: a MATLAB package for FDEM data inversion. Numerical Algorithms, 84(4):1313–1327, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00843-2.

DiazCalle:2005:WAG

[DDG05] Jorge L. Díaz Calle, Philippe R. B. Devloo, and Sónia M. Gomes. Wavelets and adaptive grids for the discontinuous Galerkin method. Numerical Algorithms, 39(1–3):143–154, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

deBruin:1992:CSH

[dDL92] M. G. de Bruin, K. A. Driver, and D. S. Lubinsky. Convergence of simultaneous Hermite–Padé approximants to the n-tuple of q-hypergeometric series \{_{1}Φ_{1}(\binom{1,1}{c,γ_{j}}andz)\}_{j=1}^{n}. Numerical Algorithms, 3(1–4):185–192, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

DAmbrosio:2014:OCG

[DDP14] Raffaele D’Ambrosio, Giuseppe De Martino, and Beatrice Paternoster. Order conditions for General Linear Nyström methods. Numerical Algorithms, 65(3):579–595, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9819-3.

Djukic:2023:IGA

[DDRS23] D. Lj. Djukić, R. M. Mutavdžić Djukić, L. Reichel, and M. M. Spalević. Internality of generalized averaged Gauss quadrature rules and truncated variants for modified Chebyshev measures of the third and fourth kinds. Numerical Algorithms, 92 (1):523–544, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01385-w.

Dumitrescu:1997:TDB

[DDRT97] B. Dumitrescu, M. Doreille, J.-L. Roch, and D. Trystram. Two-dimensional block partitionings for the parallel sparse Cholesky factorization. Numerical Algorithms, 16(1):17–38, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
9/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/9/3/fulltext.pdf. Sparse ma-

trices in industry (Lille, 1997).

Dagnino:1993:ANI

[DDS93] C. Dagnino, V. Demichelis, and E. Santi. An algorithm for numerical integration based on quasi-interpolating splines. Numerical Algorithms, 5(1–4):443–452, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

deBruin:2000:BIU

[dDS00] M. G. de Bruin, H. P. Dikshit, and A. Sharma. Birkhoff interpolation on unity and on Möbius transform of the roots of unity. Numerical Algorithms, 23(1):115–125, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/24/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
24/6/fulltext.pdf.

delaCalleYsern:2024:LIP

[ddS24] Bernardo de la Calle Ysern and Pedro Galán del Sastre. A Lagrange interpolation with preprocessing to nearly eliminate oscillations. Numerical Algorithms, 97(4):2051–2082, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01778-z.

DAmbra:2025:OPS

[DDT25] Pasqua D’Ambra, Fabio Durastante, and Stephen Thomas. Optimal polynomial smoothers for parallel AMG. Numerical Algorithms, 100(4):1783–1812, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02117-6.

deBoor:1993:EBS

[de 93] Carl de Boor. On the evaluation of box splines. Numerical Algorithms, 5(1–4):5–23, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

DeRoeck:2002:SLA

[De 02] Yann-Hervé De Roeck. Sparse linear algebra and geophysical migration: a review of direct and iterative methods. Numerical Algorithms, 29(4):283–322, April 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
39/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/39/1/fulltext.pdf.

Deadman:2015:ECN

[Dea15] Edvin Deadman. Estimating the condition number of f(A)b. Numerical Algorithms, 70(2):287–308, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9947-4.

Dajana:2024:NSW

[DEC24] Conte Dajana, Cuesta Eduardo, and Valentino Carmine. Nonstationary wave relaxation methods for general linear systems of Volterra equations: convergence and parallel GPU implementation. Numerical Algorithms, 95(1):149–180, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01567-0.

Dehmiry:2020:GCB

[Deh20] Alireza Hosseini Dehmiry. The global convergence of the BFGS method under a modified Yuan–Wei–Lu line search technique. Numerical Algorithms, 84(2):781–793, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00779-7.

Diniz-Ehrhardt:1993:PPM

[DEM93] Maria A. Diniz-Ehrhardt and José Mario Martínez. A parallel projection method for overdetermined nonlinear systems of equations. Numerical Algorithms, 4(3):241–262, March 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dodu:1994:IPA

[DEM94] J. C. Dodu, T. Eve, and M. Minoux. Implementation of a proximal algorithm for linearly constrained nonsmooth optimization problems and computational results. Numerical Algorithms, 6(3–4):245–273, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Deng:2014:NIA

[Den14a] Wei-Qi Deng. A new iterative algorithm for solving a system of generalized mixed equilibrium problems for a kind of multivalued nonlinear mappings. Numerical Algorithms, 66(4):779– 791, August 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9762-3.

Deng:2014:SCF

[Den14b] Wei-Qi Deng. The split common fixed point problem for infinite families of total quasi-asymptotically nonexpansive operators. Numerical Algorithms, 67(2):243–256, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9785-9.

DAmbrosio:2012:GLM

[DEP12] R. D’Ambrosio, E. Esposito, and B. Paternoster. General linear methods for y′′ = f(y(t)). Numerical Algorithms, 61(2):331–349, October 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=2&spage=331.

Despres:2017:PBN

[Des17] Bruno Després. Polynomials with bounds and numerical approximation. Numerical Algorithms, 76(3):829–859, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See correction [DH18].

Dieci:1997:P

[DEV97] Luca Dieci, Don Estep, and Eric Van de Velde. Preface. Numerical Algorithms, 14(1–3):1, March 1997. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/5/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/1/fulltext.pdf.

Dey:2023:HIC

[Dey23] Soumitra Dey. A hybrid inertial and contraction proximal point algorithm for monotone variational inclusions. Numerical Algorithms, 93(1):1–25, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01400-
0.

Daehlen:1993:IPI

[DF93] Morten Dæhlen and Michael Floater. Iterative polynomial interpolation and data compression. Numerical Algorithms, 5 (1–4):165–177, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Dedieu:1994:AOU

[DF94] J.-P. Dedieu and Ch. Favardin. Algorithms for ordering unorganized points along parametrized curves. Numerical Algorithms, 6(1–2):169–200, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Driscoll:2001:PBA

[DF01] Tobin A. Driscoll and Bengt Fornberg. A Padé-based algorithm for overcoming the Gibbs phenomenon. Numerical Algorithms, 26(1):77–92, January 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
4/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/4/5/fulltext.pdf.

Djukic:2023:ACS

[DFD23] Dusan Lj Djukić, Luisa Fermo, and Rada M. Mutavdzić Djukić. Averaged cubature schemes on the real positive semiaxis. Numerical Algorithms, 92(1):545–569, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01408-6.

Diethelm:2004:DEA

[DFF04] Kai Diethelm, Neville J. Ford, and Alan D. Freed. Detailed error analysis for a fractional Adams method. Numerical Algorithms, 36(1):31–52, May 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/52/A/3/abstract.htm.

deFalco:1993:UAR

[dFG93] D. de Falco, K. M. Frontini, and L. Gotusso. A unifying approach to the regularization of Fourier polynomials. Numerical Algorithms, 5(1–4):419–424, November 1993. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

DAmbrosio:2010:TSA

[DFJP10] R. DÁmbrosio, M. Ferro, Z. Jackiewicz, and B. Paternoster. Two-step almost collocation methods for ordinary differential equations. Numerical Algorithms, 53(2–3):195–217, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=195.

Doedel:1997:SCL

[DFK97] E. J. Doedel, M. J. Friedman, and B. I. Kunin. Successive continuation for locating connecting orbits. Numerical Algorithms, 14(1–3):103–124, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/6/fulltext.pdf. Dynamical

numerical analysis (Atlanta, GA, 1995).

deFalco:2011:PRP

[dFO11] Carlo de Falco and Eugene O’Riordan. A parameter robust Petrov–Galerkin scheme for advection-diffusion-reaction equations. Numerical Algorithms, 56(1):107–127, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=1&
spage=107.

Delgado:2010:OPS

[DFP^{+}10] Antonia M. Delgado, Lidia Fernández, Teresa E. Pérez, Miguel A. Piñar, and Yuan Xu. Orthogonal polynomials in several variables for measures with mass points. Numerical Algorithms, 55(2–3):245–264, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=245.

DeGiosa:2026:RVC

[DFR26] Marcello De Giosa, Antonella Falini, and Giorgia Rubino. A robust variant of cubic smoothing spline approximation. Numerical Algorithms, 101(1):221–241, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-02003-7.

deFigueiredo:2004:AAC

[dFS04] Luiz Henrique de Figueiredo and Jorge Stolfi. Affine arithmetic: Concepts and applications. Numerical Algorithms, 37(1–4):147–158, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/10/abstract.htm.

Dubrulle:1994:MQI

[DG94] Augustin A. Dubrulle and Gene H. Golub. A multishift QR iteration without computation of the shifts. Numerical Algorithms, 7(2–4):173–181, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DeMaerschalck:2005:UCP

[DG05] Bart De Maerschalck and Marc I. Gerritsma. The use of Chebyshev polynomials in the space–time least-squares spectral element method. Numerical Algorithms, 38(1–3):173–196, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Daripa:2017:FBF

[DG17] Prabir Daripa and Aditi Ghosh. The FFTRR-based fast direct algorithms for complex inhomogeneous biharmonic problems with applications to incompressible flows. Numerical Algorithms, 75(4):937–971, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dominguez:2006:PDI

[DGL06] Juan Dominguez and María D. González-Lima. A primaldual interior-point algorithm for quadratic programming. Numerical Algorithms, 42(1):1–30, May 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=1&spage=1.

Dick:2026:TFD

[DGM26] Josef Dick, Hecong Gao, and Kassem Mustapha. Timefractional diffusion equations with randomness, and efficient numerical estimations of expected values. Numerical Algorithms, 101(3):1865–1883, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02066-
0.

Duenas:2015:HOS

[DGP15] Herbert Dueñas, Luis E. Garza, and Miguel Piñar. A higher order Sobolev-type inner product for orthogonal polynomials in several variables. Numerical Algorithms, 68(1):35–46, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9836-x.

Dunster:2015:CNS

[DGST15] T. M. Dunster, A. Gil, J. Segura, and N. M. Temme. Computation of a numerically satisfactory pair of solutions of the differential equation for conical functions of non-negative integer orders. Numerical Algorithms, 68(3):497–509, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9857-5.

Demmel:2004:FAF

[DH04] James Demmel and Yozo Hida. Fast and accurate floating point summation with application to computational geometry. Numerical Algorithms, 37(1–4):101–112, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/6/abstract.htm.

Despres:2018:CPB

[DH18] B. Despres and M. Herda. Correction to: Polynomials with bounds and numerical approximation. Numerical Algorithms, 77(1):309–311, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-017-0441-7.

pdf. See [Des17].

Ding:2026:QRG

[DH26] Xiaojian Ding and Baohua Huang. The quaternion relaxed greedy randomized Kaczmarz method with adaptive parameters for solving quaternion matrix equation. Numerical Algorithms, 101(3):1989–2016, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02073-1.

Deng:2021:PEA

[DHF21] Lanmei Deng, Rong Hu, and Yaping Fang. Projection extragradient algorithms for solving nonmonotone and non-Lipschitzian equilibrium problems in Hilbert spaces. Numerical Algorithms, 86(1):191–221, January 2021. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00885-
x.

Dahl:2010:AST

[DHJJ10] Joachim Dahl, Per Christian Hansen, Søren Holdt Jensen, and Tobias Lindstrøm Jensen. Algorithms and software for total variation image reconstruction via first-order methods. Numerical Algorithms, 53(1):67–92, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=1&spage=67.

Defour:2004:PSM

[DHL^{+}04] David Defour, Guillaume Hanrot, Vincent Lefèvre, Jean-Michel Muller, Nathalie Revol, and Paul Zimmermann. Proposal for a standardization of mathematical function implementation in floating-point arithmetic. Numerical Algorithms, 37(1–4):367–375, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp009.kluweronline.com/IPS/content/ext/
x/J/5058/I/58/A/30/abstract.htm;   http://perso.ens-
lyon.fr/jean-michel.muller/NumAlg04.pdf;   http://
www.loria.fr/~zimmerma/papers/PropStandFunctions.pdf.

Duenas:2012:APL

[DHM12] Herbert Dueñas, Edmundo J. Huertas, and Francisco Marcellán. Asymptotic properties of Laguerre–Sobolev type orthogonal polynomials. Numerical Algorithms, 60(1):51–73, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=60&
issue=1&spage=51.

Donatelli:2016:IDS

[DHMS16a] Marco Donatelli, Thomas Huckle, Mariarosa Mazza, and Debora Sesana. Image deblurring by sparsity constraint on the Fourier coefficients. Numerical Algorithms, 72(2):341–361, June 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0047-x.

Duminil:2016:ACM

[DHMS16b] Sébastien Duminil, Mohammed Heyouni, Philippe Marion, and Hassane Sadok. Algorithms for the CMRH method for dense linear systems. Numerical Algorithms, 71(2):383–394, February 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9997-2.

Dragan:1997:WCC

[DHS97] Vasile Dragan, Aristide Halanay, and Adrian Stoica. Wellconditioned computation for H^{∞} controller near the optimum. Numerical Algorithms, 15(2):193–206, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/7/4/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/7/
4/fulltext.pdf.

Dimulyo:2009:FCT

[DHS09] Sarpono Dimulyo, Zulfiqar Habib, and Manabu Sakai. Fair cubic transition between two circles with one circle inside or tangent to the other. Numerical Algorithms, 51(4):461–476, August 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=51&
issue=4&spage=461.

Dostal:2022:HSH

[DHV22] Zdenek Dostál, David Horák, and Oldrich Vlach. Highly scalable hybrid domain decomposition method for the solution of huge scalar variational inequalities. Numerical Algorithms, 91 (2):773–801, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01281-3.

Deng:2025:GFB

[DHWL25] Ting Deng, Jin Huang, Yifei Wang, and Hu Li. A generalized fuzzy barycentric Lagrange interpolation method for solving two-dimensional fuzzy fractional Volterra integral equations. Numerical Algorithms, 98(2):743–766, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01814-y.

Dragan:2011:CSS

[DI11] Vasile Dragan and Ivan G. Ivanov. Computation of the stabilizing solution of game theoretic Riccati equation arising in stochastic H^{∞} control problems. Numerical Algorithms, 57(3):357–375, July 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=3&spage=357.

Diethelm:2008:ISN

[Die08] Kai Diethelm. An investigation of some nonclassical methods for the numerical approximation of Caputo-type fractional derivatives. Numerical Algorithms, 47(4):361–390, April 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=4&
spage=361.

DeMarinis:2022:MTO

[DIM22] Arturo De Marinis, Felice Iavernaro, and Francesca Mazzia. A minimum-time obstacle-avoidance path planning algorithm for unmanned aerial vehicles. Numerical Algorithms, 89(4):1639– 1661, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01167-w.

Driver:2002:ZP

[DJ02] K. Driver and K. Jordaan. Zeros of _{3}F_{2}( −n, b, c,.e; z) polynomials. Numerical Algorithms, 30(3–4):323–333, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/42/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
42/5/fulltext.pdf.

DAmbrosio:2010:CTS

[DJ10] Raffaele DÁmbrosio and Zdzisław Jackiewicz. Continuous twostep Runge–Kutta methods for ordinary differential equations. Numerical Algorithms, 54(2):169–193, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=2&spage=169.

Driver:2018:ZJP

[DJ18] Kathy Driver and Kerstin Jordaan. Zeros of Jacobi polynomials P_{n}^{(α,β)}, −2 < α, β < −1. Numerical Algorithms, 79(4): 1075–1085, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dong:2018:MSE

[DJG18] Qiao-Li Dong, Dan Jiang, and Aviv Gibali. A modified subgradient extragradient method for solving the variational inequality problem. Numerical Algorithms, 79(3):927–940, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Driver:2008:IZJ

[DJM08] Kathy Driver, Kerstin Jordaan, and Norbert Mbuyi. Interlacing of the zeros of Jacobi polynomials with different parameters. Numerical Algorithms, 49(1–4):143–152, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=49&issue=1&
spage=143.

Duarte:2018:OHA

[DJM^{+}18] Jorge Duarte, Cristina Januário, Nuno Martins, C. Correia Ramos, Carla Rodrigues, and Josep Sardanyés. Optimal homotopy analysis of a chaotic HIV-1 model incorporating AIDSrelated cancer cells. Numerical Algorithms, 77(1):261–288, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dopico:2020:SVM

[DJS20] Froilan Dopico, Khalide Jbilou, and Mohammed Seaid. Special volume on mathematical modeling with applications. Numerical Algorithms, 84(4):1239–1240, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00969-8;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00969-8.pdf.

Druskin:2000:GSR

[DK00] Vladimir Druskin and Leonid Knizhnerman. Gaussian spectral rules for second order finite-difference schemes. Numerical Algorithms, 25(1–4):139–159, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/3/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/3/fulltext.pdf. Mathemat-

ical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Datsko:2025:ENM

[DK25] B. Y. Datsko and M. V. Kutniv. Explicit numerical methods for solving singular initial value problems for systems of secondorder nonlinear ODEs. Numerical Algorithms, 98(2):929–942, February 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01820-0.

Dick:2015:NIL

[DKL15] Josef Dick, Peter Kritzer, and Gunther Leobacher. Numerical integration in log-Korobov and log-cosine spaces. Numerical Algorithms, 70(4):753–775, December 2015. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-015-9972-y.

Dieci:1997:LTN

[DL97] Luca Dieci and Jens Lorenz. Lyapunov-type numbers and torus breakdown: numerical aspects and a case study. Numerical Algorithms, 14(1–3):79–102, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/5/fulltext.pdf. Dynamical

numerical analysis (Atlanta, GA, 1995).

Driver:2001:PHF

[DL01] K. A. Driver and A. D. Love. Products of hypergeometric functions and the zeros of _{4}F_{3} polynomials. Numerical Algorithms, 26(1):1–9, January 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
4/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/4/1/fulltext.pdf.

DelBuono:2003:NIC

[DL03] N. Del Buono and L. Lopez. Numerical integration of a class of ordinary differential equations on the general linear group of matrices. Numerical Algorithms, 34(2–4):271–281, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/27/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
27/fulltext.pdf.

Dyllong:2004:GDA

[DL04] Eva Dyllong and Wolfram Luther. The GJK distance algorithm: An interval version for incremental motions. Numerical Algorithms, 37(1–4):127–136, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/8/abstract.htm.

Dadourian:2008:ASB

[DL08] Karine Dadourian and Jacques Liandrat. Analysis of some bivariate non-linear interpolatory subdivision schemes. Numerical Algorithms, 48(1–3):261–278, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=261.

Dubois:2009:CEO

[DL09] O. Dubois and S. H. Lui. Convergence estimates for an optimized Schwarz method for PDEs with discontinuous coefficients. Numerical Algorithms, 51(1):115–131, May 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=1&

spage=115. Tributes to Gene H. Golub, Part I.

DeBonis:2018:SMN

[DL18] M. C. De Bonis and C. Laurita. On the stability of a modified Nyström method for Mellin convolution equations in weighted spaces. Numerical Algorithms, 79(2):611–631, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dastour:2021:OPF

[DL21] Hatef Dastour and Wenyuan Liao. An optimal 13-point finite difference scheme for a 2D Helmholtz equation with a perfectly matched layer boundary condition. Numerical Algorithms, 86(3):1109–1141, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00926-5.

Dong:2014:NRF

[DLC14] Bo Dong, Matthew M. Lin, and Moody T. Chu. Nonnegative rank factorization — a heuristic approach via rank reduction. Numerical Algorithms, 65(2):251–274, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9704-0.

Duan:2021:NMG

[DLDW21] Xue-Feng Duan, Juan Li, Shan-Qi Duan, and Qing-Wen Wang. Numerical method for the generalized nonnegative tensor factorization problem. Numerical Algorithms, 87(2):499–510, June 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00975-w.

Dai:2004:RPC

[DLL04] Yu-Hong Dai, Li-Zhi Liao, and Duan Li. On restart procedures for the conjugate gradient method. Numerical Algorithms, 35 (2–4):249–260, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/50/A/4/
abstract.htm.

Dai:2012:EEB

[DLL12a] Ping-Fan Dai, Yao-Tang Li, and Chang-Jing Lu. Erratum to: Error bounds for linear complementarity problems for SB-matrices. Numerical Algorithms, 61(1):187, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=1&

spage=187. See [DLL12b].

Dai:2012:EBL

[DLL12b] Ping-Fan Dai, Yao-Tang Li, and Chang-Jing Lu. Error bounds for linear complementarity problems for SB-matrices. Numerical Algorithms, 61(1):121–139, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=1&spage=121.

See erratum [DLL12a].

Dai:2013:NEB

[DLL13] Ping-Fan Dai, Chang-Jing Lu, and Yao-Tang Li. New error bounds for the linear complementarity problem with an SB-matrix. Numerical Algorithms, 64(4):741–757, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9691-6.

Ding:2024:NSP

[DLL^{+}24] Wenxv Ding, Zhihong Liu, Ying Li, Anli Wei, and Mingcui Zhang. New structure-preserving algorithms of Gauss–Seidel and successive over-relaxation iteration methods for quaternion linear systems. Numerical Algorithms, 95(3):1309–1323, March 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01609-7.

Duan:2017:NMS

[DLLD17] Xue-Feng Duan, Chun-Mei Li, Jiao-Fen Li, and Yong Ding. Numerical methods for solving some matrix feasibility problems. Numerical Algorithms, 74(2):461–479, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0155-2.

Dagnino:2012:BSB

[DLR12] Catterina Dagnino, Paola Lamberti, and Sara Remogna. Bspline bases for unequally smooth quadratic spline spaces on non-uniform criss-cross triangulations. Numerical Algorithms, 61(2):209–222, October 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=2&spage=209.

Duff:2024:GSS

[DLR24] Timothy Duff, Anton Leykin, and Jose Israel Rodriguez. u-generation: solving systems of polynomials equation-by-equation. Numerical Algorithms, 95(2):813–838, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01590-1.

Dong:2017:ASM

[DLYH17] Qiao-Li Dong, Yan-Yan Lu, Jinfeng Yang, and Songnian He. Approximately solving multi-valued variational inequalities by using a projection and contraction algorithm. Numerical Algorithms, 76(3):799–812, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dubuc:1992:CHP

[DM92] Serge Dubuc and Abdul Malik. Convex hull of powers of a complex number, trinomial equations and the Farey sequence. Numerical Algorithms, 2(1):1–32, February 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Draux:1997:RMP

[DM97] André Draux and Borhane Moalla. Rectangular matrix Padé approximants and square matrix orthogonal polynomials. Numerical Algorithms, 14(4):321–341, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/6/3/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/6/
3/fulltext.pdf.

Daripa:1998:ENN

[DM98a] Prabir Daripa and Daoud Mashat. An efficient and novel numerical method for quasiconformal mappings of doubly connected domains. Numerical Algorithms, 18(2): 159–175, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
14/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/14/4/fulltext.pdf.

Daripa:1998:SIT

[DM98b] Prabir Daripa and Daoud Mashat. Singular integral transforms and fast numerical algorithms. Numerical Algorithms, 18(2):133–157, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
14/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/14/3/fulltext.pdf.

Dubois:2003:MCF

[DM03] François Dubois and Stéphanie Mengué. Mixed collocation for fractional differential equations. Numerical Algorithms, 34 (2–4):303–311, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
22/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/22/fulltext.pdf.

Derevyagin:2014:NGT

[DM14] Maxim Derevyagin and Francisco Marcellán. A note on the Geronimus transformation and Sobolev orthogonal polynomials. Numerical Algorithms, 67(2):271–287, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9788-6.

Dessole:2021:MPA

[DM21] M. Dessole and F. Marcuzzi. A massively parallel algorithm for Bordered Almost Block Diagonal Systems on GPUs. Numerical Algorithms, 86(3):1243–1263, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00931-8.

Dessole:2022:DMR

[DM22] Monica Dessole and Fabio Marcuzzi. Deviation maximization for rank-revealing QR factorizations. Numerical Algorithms, 91(3):1047–1079, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01291-1.

Dehghan:2009:FOC

[DMA09] Mehdi Dehghan, Akbar Mohebbi, and Zohreh Asgari. Fourthorder compact solution of the nonlinear Klein–Gordon equation. Numerical Algorithms, 52(4):523–540, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=4&
spage=523.

Dehghani:2019:SNC

[DMA19] R. Dehghani and N. Mahdavi-Amiri. Scaled nonlinear conjugate gradient methods for nonlinear least squares problems. Numerical Algorithms, 82(1):1–20, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dansou:2020:OFM

[DMC20] Anicet Dansou, Saïda Mouhoubi, and Cyrille Chazallon. Optimizations of a fast multipole symmetric Galerkin boundary element method code. Numerical Algorithms, 84(3):825–846, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00781-z.

Dehghani-Madiseh:2016:PAS

[DMD16] Marzieh Dehghani-Madiseh and Mehdi Dehghan. Parametric AE-solution sets to the parametric linear systems with multiple right-hand sides and parametric matrix equation A(p)X = B(p). Numerical Algorithms, 73(1):245–279, September 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0094-3.

DelaCruzCabrera:2020:EIN

[DMR20] Omar De la Cruz Cabrera, Mona Matar, and Lothar Reichel. Edge importance in a network via line graphs and the matrix exponential. Numerical Algorithms, 83(2):807–832, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DelaCruzCabrera:2021:CMN

[DMR21] Omar De la Cruz Cabrera, Mona Matar, and Lothar Reichel. Centrality measures for node-weighted networks via line graphs and the matrix exponential. Numerical Algorithms, 88(2): 583–614, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01050-0.

Dora:2003:HSH

[DMRT03] Jean Della Dora, Mihaela Mirica-Ruse, and Evelyne Tournier. Hybrid systems and hybrid computation 1st part: Hybrid systems. Numerical Algorithms, 33(1–4):203–213, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/17/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
17/fulltext.pdf.

Draux:2009:GQA

[DMS09] André Draux, Borhane Moalla, and Mohamed Sadik. Generalized qd algorithm and Markov–Bernstein inequalities for Jacobi weight. Numerical Algorithms, 51(4):429–447, August 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=4&
spage=429.

Danelakis:2013:BID

[DMT13] Antonios Danelakis, Marilena Mitrouli, and Dimitrios Triantafyllou. Blind image deconvolution using a banded matrix method. Numerical Algorithms, 64(1):43–72, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9654-y.

Dinh:2022:MSS

[DMT22] Bui Van Dinh, Hy Duc Manh, and Tran Thi Huyen Thanh. A modified Solodov–Svaiter method for solving nonmonotone variational inequality problems. Numerical Algorithms, 90(4): 1715–1734, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01248-w.

Du:2023:LPF

[DMW23] Guangzhi Du, Shilin Mi, and Xinhui Wang. Local and parallel finite element methods based on two-grid discretizations for the unsteady mixed Stokes–Darcy model with the Beavers–Joseph interface condition. Numerical Algorithms, 94(4):1883–1918, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01558-1.

DiNunno:2023:DIE

[DMYT23] Giulia Di Nunno, Yuliya Mishura, and Anton Yurchenko-Tytarenko. Drift-implicit Euler scheme for sandwiched processes driven by Hölder noises. Numerical Algorithms, 93 (2):459–491, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01424-6.

Dieci:2020:DDI

[DMZ20] Luca Dieci, Manuela Manetta, and Haomin Zhou. Double descent and intermittent color diffusion for landscape exploration. Numerical Algorithms, 85(1):145–169, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00807-6.

Dhanush:2019:IVE

[DN19] V. Dhanush and S. Natarajan. Implementation of the virtual element method for coupled thermo-elasticity in Abaqus. Numerical Algorithms, 80(3):1037–1058, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Denich:2024:FSA

[DN24a] Eleonora Denich and Paolo Novati. A fast and simple algorithm for the computation of the Lerch transcendent. Numerical Algorithms, 96(1):13–32, May 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01637-3.

Duc:2024:RMC

[DN24b] Nguyen Van Duc and Thi-Phong Nguyen. A regularization method for Caputo fractional derivatives in the Banach space L^{∞}[0, T]. Numerical Algorithms, 95(2):1033–1053, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01598-7.

Duc:2025:TTR

[DNH^{+}25] Nguyen Van Duc, Thi-Phong Nguyen, Nguyen Phuong Ha, Nguyen The Anh, Luu Duc Manh, and Hoang Cong Gia Bao. A Tikhonov-type regularization method for Caputo fractional derivative. Numerical Algorithms, 99(1):441–462, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01883-z.

Dykes:2015:RGA

[DNR15] L. Dykes, S. Noschese, and L. Reichel. Rescaling the GSVD with application to ill-posed problems. Numerical Algorithms, 68(3):531–545, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9859-3.

Dykes:2017:CPD

[DNR17] L. Dykes, S. Noschese, and L. Reichel. Circulant preconditioners for discrete ill-posed Toeplitz systems. Numerical Algorithms, 75(2):477–490, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dang:2025:CHO

[DNV25] Quang A Dang, Thanh Huong Nguyen, and Vinh Quang Vu. Construction of high order numerical methods for solving fourth order nonlinear boundary value problems. Numerical Algorithms, 99(1):323–354, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01879-
9.

Dieci:2025:SSD

[DO25] Luca Dieci and Daniyar Omarov. Solving semi-discrete optimal transport problems: star shapedeness and Newton’s method. Numerical Algorithms, 99(2):949–1004, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01903-y.

Dong:2010:SSS

[Don10] Jun-Liang Dong. Shifted skew-symmetric iteration methods for nonsymmetric linear complementarity problems. Numerical Algorithms, 54(3):343–357, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=343.

Donatelli:2012:NSR

[Don12] Marco Donatelli. On nondecreasing sequences of regularization parameters for nonstationary iterated Tikhonov. Numerical Algorithms, 60(4):651–668, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=651.

Dong:2013:TIP

[Don13] Xuanchun Dong. A trigonometric integrator pseudospectral discretization for the N-coupled nonlinear Klein–Gordon equations. Numerical Algorithms, 62(2):325–336, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9586-6/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=325-336.

Dong:2016:CNT

[Don16] XiaoLiang Dong. Comment on “A new three-term conjugate gradient method for unconstrained problem”. Numerical Algorithms, 72(1):173–179, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0039-x. See

[And15].

Dehestani:2026:PNA

[DOR26] Haniye Dehestani, Yadollah Ordokhani, and Mohsen Razzaghi. Piecewise numerical approach for solving piecewise fractional optimal control and variational problems. Numerical Algorithms, 101(2):823–846, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02024-w.

Dostal:2003:PBA

[Dos03a] Z. Dostál. A proportioning based algorithm with rate of convergence for bound constrained quadratic programming. Numerical Algorithms, 34(2–4):293–302, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/48/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
48/1/fulltext.pdf.

Dubeau:2003:GMN

[DOS03b] F. Dubeau, A. Ouansafi, and A. Sakat. Galerkin methods for nonlinear ordinary differential equation with impulses. Numerical Algorithms, 33(1–4):215–225, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/18/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
18/fulltext.pdf.

deOliveira:2007:UFT

[dOS07] Rodrigo M. S. de Oliveira and Carlos Leonidas S. S. Sobrinho. UPML formulation for truncating conductive media in curvilinear coordinates. Numerical Algorithms, 46(4):295–319, December 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=4&spage=295.

Bonis:2021:FIS

[DOT21] M. C. De Bonis, D. Occorsio, and W. Themistoclakis. Filtered interpolation for solving Prandtl’s integro-differential equations. Numerical Algorithms, 88(2):679–709, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01053-x.

Dieci:2001:CEB

[DP01] Luca Dieci and Alessandra Papini. Conditioning of the exponential of a block triangular matrix. Numerical Algorithms, 28(1–4):137–150, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/37/
20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/37/20/fulltext.pdf;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/20/

abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/20/fulltext.pdf.

Dopico:2016:SCN

[DP16] Froilán M. Dopico and Kenet Pomés. Structured condition numbers for linear systems with parameterized quasiseparable coefficient matrices. Numerical Algorithms, 73(4):1131– 1158, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0133-8.

Delkhosh:2021:NCM

[DP21] Mehdi Delkhosh and Kourosh Parand. A new computational method based on fractional Lagrange functions to solve multiterm fractional differential equations. Numerical Algorithms, 88(2):729–766, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01055-9.

Dieci:2019:CPE

[DPP19] Luca Dieci, Alessandra Papini, and Alessandro Pugliese. Coalescing points for eigenvalues of banded matrices depending on parameters with application to banded random matrix functions. Numerical Algorithms, 80(4):1241–1266, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dieci:2022:DCE

[DPP22] Luca Dieci, Alessandra Papini, and Alessandro Pugliese. Decompositions and coalescing eigenvalues of symmetric definite pencils depending on parameters. Numerical Algorithms, 91 (4):1879–1910, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01326-
7.

Derevianko:2023:EVE

[DPR23] Nadiia Derevianko, Gerlind Plonka, and Raha Razavi. ES-PRIT versus ESPIRA for reconstruction of short cosine sums and its application. Numerical Algorithms, 92(1):437–470, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01432-6.

Djukic:2018:EBG

[DPS18] Dusan Lj. Djukić, Aleksandar V. Pejcev, and Miodrag M. Spalević. The error bounds of Gauss–Kronrod quadrature formulae for weight functions of Bernstein–Szegő type. Numerical Algorithms, 77(4):1003–1028, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

daRocha:1999:SFC

[dR99] Zélia da Rocha. Shohat–Favard and Chebyshev’s methods in d-orthogonality. Numerical Algorithms, 20(2–3):139–164, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/4/fulltext.pdf.

DelCorso:2001:HST

[DR01] Gianna M. Del Corso and Francesco Romani. Heuristic spectral techniques for the reduction of bandwidth and workbound of sparse matrices. Numerical Algorithms, 28(1– 4):117–136, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
37/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/8/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
8/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/8/fulltext.pdf.

Didier:2004:HRB

[DR04] Laurent-Stéphane Didier and Fabien Rico. High radix BKM algorithm. Numerical Algorithms, 37(1–4):113–125, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/7/abstract.htm.

Drioli:2007:UKB

[DR07] Carlo Drioli and Davide Rocchesso. On the use of kernelbased methods in sound synthesis by physical modeling. Numerical Algorithms, 45(1–4):315–329, August 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-007-9117-z.

Dykes:2012:RTM

[DR12] L. Dykes and L. Reichel. On the reduction of Tikhonov minimization problems and the construction of regularization matrices. Numerical Algorithms, 60(4):683–696, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=4&
spage=683.

Draux:1996:FOP

[Dra96] André Draux. Formal orthogonal polynomials revisited. Applications. Numerical Algorithms, 11(1–4):143–158, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Draux:2000:IFN

[Dra00] André Draux. Improvement of the formal and numerical estimation of the constant in some Markov–Bernstein inequalities. Numerical Algorithms, 24(1–2):31–58, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/4/fulltext.pdf. Computational methods from rational approximation theory (Wilrijk, 1999).

Draux:2002:FOP

[Dra02] André Draux. Formal orthogonal polynomials and Newton– Padé approximants. Numerical Algorithms, 29(1–3):67–74, March 2002. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp008.lwwonline.
com/content/getfile/5058/38/5/abstract.htm;   http:
//ipsapp008.lwwonline.com/content/getfile/5058/38/
5/fulltext.pdf.

Drieschner:1993:CAD

[Dri93] Rudolf Drieschner. Chebyshev approximation to data by geometric elements. Numerical Algorithms, 5(1–4):509–522, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Marchi:2007:P

[DRZV07] Stefano De Marchi, Michela Redivo-Zaglia, and Marco Vianello. Preface. Numerical Algorithms, 45(1–4):1–9, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=45&issue=1&
spage=1.

deBruin:2000:BIN

[dS00] Marcel G. de Bruin and A. Sharma. Birkhoff interpolation on non-uniformly distributed roots of unity. Numerical Algorithms, 25(1–4):123–138, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/4/fulltext.pdf. Mathemat-

ical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Dehghan:2009:MLS

[DS09a] Mehdi Dehghan and Fatemeh Shakeri. Method of lines solutions of the parabolic inverse problem with an overspecification at a point. Numerical Algorithms, 50(4):417–437, April 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=50&issue=4&
spage=417.

Dehghan:2009:MMN

[DS09b] Mehdi Dehghan and Ali Shokri. A meshless method for numerical solution of the one-dimensional wave equation with an integral condition using radial basis functions. Numerical Algorithms, 52(3):461–477, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=461.

Dosso:2009:STM

[DS09c] M. Dosso and M. Sadkane. A spectral trichotomy method for symplectic matrices. Numerical Algorithms, 52(2):187–212, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
52&issue=2&spage=187.

Draux:2012:GAB

[DS12] André Draux and Mohamed Sadik. Generalized qd algorithm for block band matrices. Numerical Algorithms, 61(3):377–396, November 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=3&spage=377.

Demichelis:2015:MSF

[DS15] Vittoria Demichelis and Matteo Sciarra. Martensen splines and finite-part integrals. Numerical Algorithms, 69(4):693–712, August 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9921-1.

Dehghan:2020:ADS

[DS20] Mehdi Dehghan and Akbar Shirilord. Accelerated double-step scale splitting iteration method for solving a class of complex symmetric linear systems. Numerical Algorithms, 83(1): 281–304, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DAmbrosio:2021:NSP

[DS21] Raffaele D’Ambrosio and Carmela Scalone. On the numerical structure preservation of nonlinear damped stochastic oscillators. Numerical Algorithms, 86(3):933–952, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00918-5.

Doliwa:2023:IGW

[DS23] Adam Doliwa and Artur Siemaszko. Integrability and geometry of the Wynn recurrence. Numerical Algorithms, 92 (1):571–596, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01344-5.

ditSandretto:2024:VSR

[dS24] Julien Alexandre dit Sandretto. Validated B-series and Runge– Kutta pairs. Numerical Algorithms, 96(3):1045–1062, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01697-5.

Deif:2026:ISS

[DS26] Sarah Deif and Sedaghat Shahmorad. An IRM for solving a system of fractional Fredholm–Hammerstein integrodifferential equations of the second kind. Numerical Algorithms, 101(4):2709–2735, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02107-
8.

deSaboia:2004:CSG

[dSCS04] Carlos Henrique Medeiros de Sabóia, Manoel Campêlo, and Susana Scheimberg. A computational study of global algorithms for linear bilevel programming. Numerical Algorithms, 35(2–4):155–173, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/50/A/3/
abstract.htm.

Darania:2011:NCM

[DSI11] Parviz Darania, Jafar Ahmadi Shali, and Karim Ivaz. New computational method for solving some 2-dimensional nonlinear Volterra integro-differential equations. Numerical Algorithms, 57(1):125–147, May 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=1&spage=125.

Duminil:2014:FSD

[DSS14] Sebastien Duminil, Hassane Sadok, and David Silvester. Fast solvers for discretized Navier–Stokes problems using vector extrapolation. Numerical Algorithms, 66(1):89–104, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9726-7.

Delgado-Tellez:2009:NAS

[DTI09] Marina Delgado-Téllez and Alberto Ibort. A numerical algorithm for singular optimal LQ control systems. Numerical Algorithms, 51(4):477–500, August 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=4&spage=477.

Dumont:2003:SRV

[Dum03] Yves Dumont. Some remarks on a vibro-impact scheme. Numerical Algorithms, 33(1–4):227–240, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp007.kluweronline.com/content/getfile/
5058/46/19/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/19/fulltext.pdf.

Duminil:2013:PIC

[Dum13] Sébastien Duminil. A parallel implementation of the CMRH method for dense linear systems. Numerical Algorithms, 63 (1):127–142, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9616-4.

Dunham:1994:UAM

[Dun94] Charles B. Dunham. Uniform approximation by McCallig rationals. Numerical Algorithms, 7(2–4):201–204, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Durand:1993:CCA

[Dur93] Sylvain Durand. Convergence of cascade algorithms introduced by I. Daubechies. Numerical Algorithms, 4(4):307–322, May 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

DeMarchi:2001:AAN

[DV01] Stefano De Marchi and Marco Vianello. Approximating the approximant: a numerical code for polynomial compression of discrete integral operators. Numerical Algorithms, 28(1–4):101–116, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

37/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/7/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
7/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/7/fulltext.pdf.

Denis-Vidal:2003:SIS

[DVJBN03] Lilianne Denis-Vidal, Ghislaine Joly-Blanchard, and Céline Noiret. System identifiability (symbolic computation) and parameter estimation (numerical computation). Numerical Algorithms, 34(2–4):283–292, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
21/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/21/fulltext.pdf.

Diethelm:1997:NSF

[DW97] Kai Diethelm and Guido Walz. Numerical solution of fractional order differential equations by extrapolation. Numerical Algorithms, 16(3–4):231–253, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
11/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/11/1/fulltext.pdf.

Dai:2012:GCM

[DW12] Zhifeng Dai and Fenghua Wen. Global convergence of a modified Hestenes–Stiefel nonlinear conjugate gradient method with Armijo line search. Numerical Algorithms, 59(1):79–93, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=59&
issue=1&spage=79.

Dai:2015:CAH

[DW15a] Zhifeng Dai and Fenghua Wen. Comments on another hybrid conjugate gradient algorithm for unconstrained optimization by Andrei. Numerical Algorithms, 69(2):337–341, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9899-8. See erratum [DW15b].

Dai:2015:ECA

[DW15b] Zhifeng Dai and Fenghua Wen. Erratum to: “Comments on another hybrid conjugate gradient algorithm for unconstrained optimization by Andrei”. Numerical Algorithms, 70(2):449–450, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/link.springer.com/article/10.1007/s11075-015-0049-
8;   http://link.springer.com/content/pdf/10.1007/
s11075-015-0049-8.pdf. See [DW15a].

Dingfelder:2015:ITM

[DW15c] Benedict Dingfelder and J. A. C. Weideman. An improved Talbot method for numerical Laplace transform inversion. Numerical Algorithms, 68(1):167–183, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9895-z.

Deng:2021:SLM

[DW21] Dingwen Deng and Qiang Wu. The studies of the linearly modified energy-preserving finite difference methods applied to solve two-dimensional nonlinear coupled wave equations. Numerical Algorithms, 88(4):1875–1914, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01099-5.

Deng:2022:EET

[DW22] Dingwen Deng and Qiang Wu. The error estimations of a twolevel linearized compact ADI method for solving the nonlinear coupled wave equations. Numerical Algorithms, 89(4):1663– 1693, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01168-9.

Dai:2024:MBB

[DW24] Pingfei Dai and Qingbiao Wu. Modulus-based block triangular splitting iteration method for solving the generalized absolute value equations. Numerical Algorithms, 96(2):537– 555, June 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01656-0.

Dai:2018:MNN

[DWC18] Ping-Fei Dai, Qing-Biao Wu, and Min-Hong Chen. Modified Newton–NSS method for solving systems of nonlinear equations. Numerical Algorithms, 77(1):1–21, January 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Diao:2017:SSS

[DWX17] Huai-An Diao, Yimin Wei, and Pengpeng Xie. Small sample statistical condition estimation for the total least squares problem. Numerical Algorithms, 75(2):435–455, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dai:2014:NAB

[DWZ14] Huiya Dai, Leilei Wei, and Xindong Zhang. Numerical algorithm based on an implicit fully discrete local discontinuous Galerkin method for the fractional diffusion-wave equation. Numerical Algorithms, 67(4):845–862, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9827-y.

Duchamp:2018:SNS

[DXY18] Tom Duchamp, Gang Xie, and Thomas Yu. Smoothing nonlinear subdivision schemes by averaging. Numerical Algorithms, 77(2):361–379, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-017-0319-8.

Dedieu:1993:CRR

[DY93] Jean-Pierre Dedieu and Jean-Claude Yakoubsohn. Computing the real roots of a polynomial by the exclusion algorithm. Numerical Algorithms, 4(1–2):1–24, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Dou:2016:MPI

[DYW16] Yan Dou, Ai-Li Yang, and Yu-Jiang Wu. Modified parameterized inexact Uzawa method for singular saddle-point problems. Numerical Algorithms, 72(2):325–339, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0046-y.

Dai:2001:ATP

[DZ01] Yu-Hong Dai and Hongchao Zhang. Adaptive two-point stepsize gradient algorithm. Numerical Algorithms, 27(4):377–385, August 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/5/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
5/fulltext.pdf.

Deng:2013:AFO

[DZ13] Dingwen Deng and Chengjian Zhang. Analysis of a fourthorder compact ADI method for a linear hyperbolic equation with three spatial variables. Numerical Algorithms, 63(1):1– 26, May 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9604-8.

Du:2019:TGP

[DZ19] Guangzhi Du and Liyun Zuo. A two-grid parallel partition of unity finite element scheme. Numerical Algorithms, 80(2): 429–445, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Du:2021:LPF

[DZ21] Guangzhi Du and Liyun Zuo. Local and parallel finite element methods for the coupled Stokes/ Darcy model. Numerical Algorithms, 87(4):1593–1611, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01021-
5.

Du:2022:LPP

[DZ22] Guangzhi Du and Liyun Zuo. Local and parallel partition of unity scheme for the mixed Navier–Stokes–Darcy problem. Numerical Algorithms, 91(2):635–650, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01276-0.

Du:2024:IPF

[DZ24] Guangzhi Du and Liyun Zuo. Improved parallel finite element methods for the stationary Navier–Stokes problem. Numerical Algorithms, 97(3):1383–1396, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01754-7.

Dzetkulic:2015:RIN

[Dze15] Tomás Dzetkulic. Rigorous integration of non-linear ordinary differential equations in Chebyshev basis. Numerical Algorithms, 69(1):183–205, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9889-x.

Du:2023:FAI

[DZH23] Jiarui Du, Yuanpeng Zhu, and Xuli Han. Fast algorithms for interpolation and smoothing for a general class of fourth order exponential splines. Numerical Algorithms, 94(4):1849–1881, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01557-2.

Ding:2021:LPF

[DZS21] Qi Ding, Bo Zheng, and Yueqiang Shang. Local and parallel finite element algorithms for the time-dependent Oseen equations. Numerical Algorithms, 87(4):1653–1677, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01024-2.

Dzunic:2013:ETP

[Dzu13] Jovana Dzunić. On efficient two-parameter methods for solving nonlinear equations. Numerical Algorithms, 63(3):549– 569, July 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9641-3.

Deng:2017:FPC

[DZW17] Jingwei Deng, Lijing Zhao, and Yujiang Wu. Fast predictor– corrector approach for the tempered fractional differential equations. Numerical Algorithms, 74(3):717–754, March 2017.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0169-9.

Erdogan:2012:FFD

[EA12] Fevzi Erdogan and Gabil M. Amiraliyev. Fitted finite difference method for singularly perturbed delay differential equations. Numerical Algorithms, 59(1):131–145, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=1&
spage=131.

Elidrissi:2020:CPM

[EAB20] A. Elidrissi, J. Abouir, and B. Benouahmane. On the calculation of the poles of multivariate meromorphic functions using the symbolic–numeric two-point qd-algorithm. Numerical Algorithms, 84(4):1443–1458, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00887-
9.

Essaouini:2020:CEE

[EAGS20] M. Essaouini, B. Abouzaid, P. Gaudreau, and H. Safouhi. Computation of energy eigenvalues of the anharmonic Coulombic potential with irregular singularities. Numerical Algorithms, 84(4):1397–1409, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00853-
0.

Ebadi:2018:NCH

[Eba18] Moosa Ebadi. New class of hybrid BDF methods for the computation of numerical solutions of IVPs. Numerical Algorithms, 79(1):179–193, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

El-Daou:2005:TMP

[ED05] Mohamed K. El-Daou. A tau method with perturbed boundary conditions for certain ordinary differential equations. Numerical Algorithms, 38(1–3):31–45, March 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Estatico:2013:SIA

[ED13] Claudio Estatico and Fabio Di Benedetto. Shift-invariant approximations of structured shift-variant blurring matrices. Numerical Algorithms, 62(4):615–635, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9686-3.

Elden:2022:KSL

[ED22] Lars Eldén and Maryam Dehghan. A Krylov–Schur-like method for computing the best rank-(r_{1}, r_{2}, r_{3}) approximation of large and sparse tensors. Numerical Algorithms, 91(3): 1315–1347, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01303-0.

El-Daou:2012:CCP

[EDAH12] Mohamed K. El-Daou and Khaled M. Al-Hamad. Computation of the canonical polynomials and applications to some optimal control problems. Numerical Algorithms, 61(4): 545–566, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9550-5/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=61&issue=4&spage=545-566.

El-Daou:2013:CEE

[EDAM13] Mohamed K. El-Daou, Suad Sh. Al Enezi, and Mona M. Mekkaoui. Correction of eigenvalues estimated by the Legendre–Gauss Tau method. Numerical Algorithms, 64(2): 203–220, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9660-0.

ElMoudden:2018:MRG

[EE18] M. El Moudden and A. El Ghali. Multiple reduced gradient method for multiobjective optimization problems. Numerical Algorithms, 79(4):1257–1282, December 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ezz-Eldien:2019:FPS

[EED19] S. S. Ezz-Eldien and E. H. Doha. Fast and precise spectral method for solving pantograph type Volterra integrodifferential equations. Numerical Algorithms, 81(1):57–77, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Errachid:2020:RNA

[EEM20] M. Errachid, A. Essanhaji, and A. Messaoudi. RMVPIA: a new algorithm for computing the Lagrange multivariate polynomial interpolation. Numerical Algorithms, 84(4):1507–1534, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00907-8.

Eftekhari:2015:ENP

[Eft15a] Tahereh Eftekhari. Erratum to: “A new proof of interval extension of the classic Ostrowski’s method and its modified method for computing the enclosure solutions of nonlinear equations” by Tahereh Eftekhari. Numerical Algorithms, 69(1):167, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11075-014-9920-2.pdf. See [Eft15b].

Eftekhari:2015:NPI

[Eft15b] Tahereh Eftekhari. A new proof of interval extension of the classic Ostrowski’s method and its modified method for computing the enclosure solutions of nonlinear equations. Numerical Algorithms, 69(1):157–165, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9887-z. See erratum [Eft15a].

Espelid:1994:DAAb

[EG94] Terje O. Espelid and Alan Genz. DECUHR: an algorithm for automatic integration of singular functions over a hyperrectangular region. Numerical Algorithms, 8(2–4):201–220, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ebadi:2010:HBM

[EG10] Moosa Ebadi and M. Y. Gokhale. Hybrid BDF methods for the numerical solutions of ordinary differential equations. Numerical Algorithms, 55(1):1–17, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=1&spage=1.

El-Gamel:2018:EAS

[EG18] Mohamed El-Gamel. Error analysis of sinc-Galerkin method for time-dependent partial differential equations. Numerical Algorithms, 77(2):517–533, February 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-017-0326-9.

Ekstrom:2019:MLP

[EG19] Sven-Erik Ekström and Carlo Garoni. A matrix-less and parallel interpolation-extrapolation algorithm for computing the eigenvalues of preconditioned banded symmetric Toeplitz matrices. Numerical Algorithms, 80(3):819–848, March 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-018-0508-0.pdf.

Eidelman:2008:EEC

[EGG08] Yuli Eidelman, Luca Gemignani, and Israel Gohberg. Efficient eigenvalue computation for quasiseparable Hermitian matrices under low rank perturbations. Numerical Algorithms, 47(3):253–273, March 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=3&spage=253.

Ezquerro:2013:ETV

[EGGSH13] José Antonio Ezquerro, Angela Grau, Miquel Grau-Sánchez, and Miguel Ángel Hernández. On the efficiency of two variants of Kurchatov’s method for solving nonlinear systems. Numerical Algorithms, 64(4):685–698, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9685-4.

Ezquerro:2015:FIM

[EGSHVN15] J. A. Ezquerro, M. Grau-Sánchez, M. A. Hernández-Verón, and M. Noguera. A family of iterative methods that uses divided differences of first and second orders. Numerical Algorithms, 70(3):571–589, November 2015. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9962-0.

Echebest:2004:ASS

[EGSV04] N. Echebest, M. T. Guardarucci, H. Scolnik, and M. C. Vacchino. An acceleration scheme for solving convex feasibility problems using incomplete projection algorithms. Numerical Algorithms, 35(2–4):331–350, April 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/16/abstract.htm.

Egorova:2025:NAC

[EGT25] Vera Egorova, Amparo Gil, and Nico M. Temme. A numerical algorithm for the computation of the noncentral beta distribution function. Numerical Algorithms, 99(4):1763–1790, August 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01931-8.

Enright:1997:DDE

[EH97a] W. H. Enright and H. Hayashi. A delay differential equation solver based on a continuous Runge–Kutta method with defect control. Numerical Algorithms, 16(3–4):349–364, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/11/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
11/6/fulltext.pdf.

Enright:1997:IPW

[EH97b] W. H. Enright and Min Hu. Improving performance when solving high-order and mixed-order boundary value problems in ODEs. Numerical Algorithms, 16(2):107–116, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/10/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
10/3/fulltext.pdf.

Erfanifar:2024:NMS

[EH24a] Raziyeh Erfanifar and Masoud Hajarian. A new multistep method for solving nonlinear systems with high efficiency indices. Numerical Algorithms, 97(2):959–984, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01735-2.

Erfanifar:2024:SIM

[EH24b] Raziyeh Erfanifar and Masoud Hajarian. Splitting iteration methods to solve non-symmetric algebraic Riccati matrix equation Y AY − Y B − CY + D = 0. Numerical Algorithms, 97 (2):753–778, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01722-7.

Elfving:2017:CAC

[EHN17a] Tommy Elfving, Per Christian Hansen, and Touraj Nikazad. Convergence analysis for column-action methods in image reconstruction. Numerical Algorithms, 74(3):905–924, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0176-x. See erratum [EHN17b].

Elfving:2017:ECA

[EHN17b] Tommy Elfving, Per Christian Hansen, and Touraj Nikazad. Erratum to: Convergence analysis for column-action methods in image reconstruction. Numerical Algorithms, 74(3): 925, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-016-0232-6.pdf. See

[EHN17a].

El-Halouy:2023:PCS

[EHNR23] Smahane El-Halouy, Silvia Noschese, and Lothar Reichel. Perron communicability and sensitivity of multilayer networks. Numerical Algorithms, 92(1):597–617, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01418-4.

Ehrich:1997:SOP

[Ehr97] Sven Ehrich. Sard-optimal prefilters for the fast wavelet transform. Numerical Algorithms, 16(3–4):303–319, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.

com/content/getfile/5058/11/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
11/4/fulltext.pdf.

Elbouyahyaoui:2021:RDB

[EHTSM21] Lakhdar Elbouyahyaoui, Mohammed Heyouni, Azita Tajaddini, and Farid Saberi-Movahed. On restarted and deflated block FOM and GMRES methods for sequences of shifted linear systems. Numerical Algorithms, 87(3):1257–1299, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01007-3.

Ezquerro:2019:NFI

[EHV19] J. A. Ezquerro and M. A. Hernández-Verón. Nonlinear Fredholm integral equations and majorant functions. Numerical Algorithms, 82(4):1303–1323, December 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ezquerro:2014:HMI

[EHVRV14] J. A. Ezquerro, M. A. Hernández-Verón, M. J. Rubio, and A. I. Velasco. An hybrid method that improves the accessibility of Steffensen’s method. Numerical Algorithms, 66(2): 241–267, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9732-9.

ElGuennouni:2002:BKS

[EJR02] A. El Guennouni, K. Jbilou, and A. J. Riquet. Block Krylov subspace methods for solving large Sylvester equations. Numerical Algorithms, 29(1–3):75–96, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp008.lwwonline.com/content/
getfile/5058/38/6/abstract.htm;   http://ipsapp008.
lwwonline.com/content/getfile/5058/38/6/fulltext.pdf.

ElHachimi:2026:NNE

[EJR26] Anas El Hachimi, Khalide Jbilou, and Ahmed Ratnani. Nonnegative Einstein tensor factorization for unmixing hyperspectral images. Numerical Algorithms, 101(1):423–453, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02011-1.

Elhay:1994:JMM

[EK94] Sylvan Elhay and Jaroslav Kautsky. Jacobi matrices for measures modified by a rational factor. Numerical Algorithms, 6 (3–4):205–227, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Emmel:2003:PJF

[EKM03] L. Emmel, S. M. Kaber, and Y. Maday. Padé–Jacobi filtering for spectral approximations of discontinuous solutions. Numerical Algorithms, 33(1–4):251–264, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/21/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
21/fulltext.pdf.

Eisenmann:2023:RTM

[EKPU23] Henrik Eisenmann, Felix Krahmer, Max Pfeffer, and André Uschmajew. Riemannian thresholding methods for row-sparse and low-rank matrix recovery. Numerical Algorithms, 93 (2):669–693, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01433-5.

ElMouatasim:2018:IRG

[El 18] Abdelkrim El Mouatasim. Implementation of reduced gradient with bisection algorithms for non-convex optimization problem via stochastic perturbation. Numerical Algorithms, 78(1):41– 62, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

EstevezSchwarz:2001:CCI

[EL01] Diana Estévez Schwarz and René Lamour. The computation of consistent initial values for nonlinear index-2 differentialalgebraic equations. Numerical Algorithms, 26(1):49–75, January 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/4/4/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/4/
4/fulltext.pdf.

Elbert:2008:ZCE

[EL08] Árpád Elbert and Andrea Laforgia. The zeros of the complementary error function. Numerical Algorithms, 49(1–4): 153–157, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=49&issue=1&spage=153.

Elfving:2025:NPL

[Elf25] Tommy Elfving. Noise propagation in linear stationary iterations. Numerical Algorithms, 99(2):603–625, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01890-0.

Elliott:1993:LSD

[Ell93] G. H. Elliott. Least squares data fitting using shape preserving piecewise approximations. Numerical Algorithms, 5(1–4):365– 371, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Erturk:2007:RAS

[EM07] Vedat Suat Erturk and Shaher Momani. A reliable algorithm for solving tenth-order boundary value problems. Numerical Algorithms, 44(2):147–158, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=2&spage=147.

Enright:2010:NIA

[EM10] W. H. Enright and P. H. Muir. New interpolants for asymptotically correct defect control of BVODEs. Numerical Algorithms, 53(2–3):219–238, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=219.

Emad:1996:PRR

[Ema96] Nahid Emad. The Padé-Rayleigh–Ritz method for solving large Hermitian eigenproblems. Numerical Algorithms, 11(1– 4):159–179, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Ebadi:2025:UMS

[EMS25] Ghodrat Ebadi, Khosro Mehrabi, and Predrag S. Stanimirović. An Uzawa–DOS method for solving saddle-point problems. Numerical Algorithms, 99(1):165–182, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01873-1.

Eldar:2011:ARK

[EN11] Yonina C. Eldar and Deanna Needell. Acceleration of randomized Kaczmarz method via the Johnson–Lindenstrauss Lemma. Numerical Algorithms, 58(2):163–177, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=2&spage=163.

Enright:2002:DIU

[Enr02] W. H. Enright. The design and implementation of usable ODE software. Numerical Algorithms, 31(1–4):125–137, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/23/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/43/
23/fulltext.pdf.

Espelid:1994:DAAa

[EO94] T. O. Espelid and K. J. Overholt. DQAINF: an algorithm for automatic integration of infinite oscillating tails. Numerical Algorithms, 8(1):83–101, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Erra:1997:SSI

[EP97] Robert Erra and Bernard Philippe. On some structured inverse eigenvalue problems. Numerical Algorithms, 15 (1):15–35, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
30/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/2/fulltext.pdf.

Eskandani:2019:ZPP

[ER19] G. Zamani Eskandani and M. Raeisi. On the zero point problem of monotone operators in Hadamard spaces. Numerical Algorithms, 80(4):1155–1179, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Erb:2015:ALM

[Erb15] Wolfgang Erb. Accelerated Landweber methods based on codilated orthogonal polynomials. Numerical Algorithms, 68(2): 229–260, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9842-z.

Ernst:2000:EIM

[Ern00] Oliver G. Ernst. Equivalent iterative methods for p-cyclic matrices. Numerical Algorithms, 25(1–4):161–180, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/27/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/27/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Ernst:2019:LBC

[ES19] Philip A. Ernst and Fazlollah Soleymani. A Legendre-based computational method for solving a class of Itô stochastic delay differential equations. Numerical Algorithms, 80(4):1267–1282, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Espelid:2005:CII

[Esp05] T. O. Espelid. Computation of an infinite integral using series acceleration. Numerical Algorithms, 40(1):17–22, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=1&
spage=17.

Essai:1998:WFG

[Ess98] Azeddine Essai. Weighted FOM and GMRES for solving nonsymmetric linear systems. Numerical Algorithms, 18(3– 4):277–292, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/3/fulltext.pdf.

Espedal:1998:HND

[ETY98] Magne S. Espedal, Xue-Cheng Tai, and Ningning Yan. A hybrid nonoverlapping domain decomposition scheme for advection dominated advection-diffusion problems. Numerical Algorithms, 18(3–4):321–336, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/6/fulltext.pdf.

Ekstrom:2022:MLM

[EV22] Sven-Erik Ekström and Paris Vassalos. A matrix-less method to approximate the spectrum and the spectral function of Toeplitz matrices with real eigenvalues. Numerical Algorithms, 89(2):701–720, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01130-
9;   http://link.springer.com/content/pdf/10.1007/
s11075-021-01130-9.pdf.

Enright:2010:RCT

[EY10] W. H. Enright and Li Yan. The reliability/cost trade-off for a class of ODE solvers. Numerical Algorithms, 53(2–3):239–260, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
53&issue=2&spage=239.

Fabien:2016:PCS

[Fab16] Brian C. Fabien. Parallel collocation solution of index-1 BVP–DAEs arising from constrained optimal control problems. Numerical Algorithms, 71(2):311–335, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9994-5.

Ferreira:2025:CSC

[FAM25a] Inês A. Ferreira, Juan A. Acebrón, and José Monteiro. Correction to: Survey of a class of iterative row-action methods: The Kaczmarz method. Numerical Algorithms, 100(1):553–593, September 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01954-1. See [FAM25b].

Ferreira:2025:SCI

[FAM25b] Inês A. Ferreira, Juan A. Acebrón, and José Monteiro. Survey of a class of iterative row-action methods: The Kaczmarz method. Numerical Algorithms, 99(4):2063–2095, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01945-2. See correction [FAM25a].

Fairag:2020:TLM

[FAMA20] Faisal Fairag, Adel Al-Mahdi, and Shahbaz Ahmad. Twolevel method for the total fractional-order variation model in image deblurring problem. Numerical Algorithms, 85(3):931– 950, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00845-0.

Fan:2015:HIM

[Fan15] Bin Fan. A hybrid iterative method with averaged mappings for hierarchical fixed point problems and variational inequalities. Numerical Algorithms, 70(3):451–467, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9956-3.

Fan:2019:CDT

[Fan19] Zhencheng Fan. Convergence of discrete time waveform relaxation methods. Numerical Algorithms, 80(2):469–483, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fang:2022:CMB

[Fan22] Ximing Fang. Convergence of modulus-based matrix splitting iteration method for a class of nonlinear complementarity problems. Numerical Algorithms, 90(3):931–950, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01215-5.

Fang:2026:NAP

[Fan26] Rui Fang. Numerical analysis of penalty-based ensemble methods. Numerical Algorithms, 101(1):509–542, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02014-y.

Farhloul:2020:MFE

[Far20] Mohamed Farhloul. Mixed finite element methods for the Oseen problem. Numerical Algorithms, 84(4):1431–1442, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00879-9.

Fasino:2023:OCL

[Fas23] Dario Fasino. Orthogonal Cauchy-like matrices. Numerical Algorithms, 92(1):619–637, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01391-y.

Fasino:2025:REM

[Fas25] Dario Fasino. Role extraction by matrix equations and generalized random walks. Numerical Algorithms, 100(4):1867–1891, December 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02155-0.

Fazzi:2023:SLR

[Faz23] Antonio Fazzi. Structured low-rank approximation for nonlinear matrices. Numerical Algorithms, 93(4):1561–1580, August 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01479-5.

Ford:2001:AAC

[FC01] Judith Ford and Ke Chen. An algorithm for accelerated computation of DWTPer-based band preconditioners. Numerical Algorithms, 26(2):167–172, February 2001. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/1/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/1/6/fulltext.pdf.

Fuentes:2023:ETM

[FDFM23] Rafael Díaz Fuentes, Marco Donatelli, Caterina Fenu, and Giorgio Mantica. Estimating the trace of matrix functions with application to complex networks. Numerical Algorithms, 92(1):503–522, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01417-5.

Frederix:2013:ACE

[FDV13] Katrijn Frederix, Steven Delvaux, and Marc Van Barel. An algorithm for computing the eigenvalues of block companion matrices. Numerical Algorithms, 62(2):261–287, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9579-5/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=261-287.

Fucik:2023:LBM

[FEK^{+}23] Radek Fučík, Pavel Eichler, Jakub Klinkovský, Robert Straka, and Tomáš Oberhuber. Lattice Boltzmann Method Analysis Tool (LBMAT). Numerical Algorithms, 93(4):1509–1525, August 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01476-8.

Fazeli:2015:KCS

[FEL15] S. A. Shahzadeh Fazeli, Nahid Emad, and Zifan Liu. A key to choose subspace size in implicitly restarted Arnoldi method. Numerical Algorithms, 70(2):407–426, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9954-5.

Ferrari:2025:FSM

[FFSC25] Paola Ferrari, Isabella Furci, and Stefano Serra-Capizzano. Flipped structured matrix-sequences in image deblurring with reflective and anti-reflective boundary conditions. Numerical Algorithms, 100(1):227–275, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01960-3.

Fasino:2003:DIE

[FG03a] Dario Fasino and Luca Gemignani. Direct and inverse eigenvalue problems for diagonal-plus-semiseparable matrices. Numerical Algorithms, 34(2–4):313–324, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/48/25/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/48/
25/fulltext.pdf.

Franco:2003:FOS

[FG03b] J. M. Franco and I. Gómez. Fourth-order symmetric DIRK methods for periodic stiff problems. Numerical Algorithms, 32(2–4):317–336, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
45/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/45/9/fulltext.pdf.

Fasino:2007:SEP

[FG07] Dario Fasino and Luca Gemignani. Structured eigenvalue problems for rational Gauss quadrature. Numerical Algorithms, 45 (1–4):195–204, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-007-9082-6.

Francisco:2021:NIR

[FGBP21] Juliano B. Francisco, Douglas S. Gonçalves, Fermín S. V. Bazán, and Lila L. T. Paredes. Nonmonotone inexact restoration approach for minimization with orthogonality constraints. Numerical Algorithms, 86(4):1651–1684, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00948-z.

Feng:2019:MAL

[FGC19] Ting-Ting Feng, Xue-Ping Guo, and Guo-Liang Chen. A modified ASOR-like method for augmented linear systems. Numerical Algorithms, 82(3):1097–1115, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fasshauer:2000:NIP

[FGJ00] Gregory E. Fasshauer, Eugene C. Gartland, Jr., and Joseph W. Jerome. Newton iteration for partial differential equations and the approximation of the identity. Numerical Algorithms, 25(1–4):181–195, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/7/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/3/

7/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Fillion-Gourdeau:2019:SDQ

[FGL19] F. Fillion-Gourdeau and E. Lorin. Simple digital quantum algorithm for symmetric first-order linear hyperbolic systems. Numerical Algorithms, 82(3):1009–1045, November 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fazzi:2019:OBM

[FGM19] Antonio Fazzi, Nicola Guglielmi, and Ivan Markovsky. An ODE-based method for computing the approximate greatest common divisor of polynomials. Numerical Algorithms, 81 (2):719–740, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ferng:1991:ALM

[FGP91] William R. Ferng, Gene H. Golub, and Robert J. Plemmons. Adaptive Lanczos methods for recursive condition estimation. Numerical Algorithms, 1(1):1–19, ???? 1991. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Franco:2001:FSS

[FGR01] J. M. Franco, I. Gómez, and L. Rández. Four-stage symplectic and P-stable SDIRKN methods with dispersion of high order. Numerical Algorithms, 26(4):347–363, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/32/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

32/4/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/32/4/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/32/4/fulltext.pdf.

Fowkes:2024:ASH

[FGS24] Jaroslav M. Fowkes, Nicholas I. M. Gould, and Jennifer A. Scott. Approximating sparse Hessian matrices using large-scale linear least squares. Numerical Algorithms, 96(4):1675–1698, August 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01681-z.

Fierro:1997:LRR

[FH97] Ricardo D. Fierro and Per Christian Hansen. Low-rank revealing UTV decompositions. Numerical Algorithms, 15 (1):37–55, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
30/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/3/fulltext.pdf.

Frankenberger:2000:KPS

[FH00] Harald Frankenberger and Martin Hanke. Kernel polynomials for the solution of indefinite and ill-posed problems. Numerical Algorithms, 25(1–4):197–212, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/21/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/21/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Farouki:2004:CMV

[FH04] Rida T. Farouki and Chang Yong Han. Computation of Minkowski values of polynomials over complex sets. Numerical Algorithms, 36(1):13–29, May 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/52/A/1/abstract.htm.

Fierro:2005:UEP

[FH05] Ricardo D. Fierro and Per Christian Hansen. UTV expansion pack: Special-purpose rank-revealing algorithms. Numerical Algorithms, 40(1):47–66, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=1&spage=47.

Fazeli:2015:NSV

[FH15] Somayyeh Fazeli and Gholamreza Hojjati. Numerical solution of Volterra integro-differential equations by superimplicit multistep collocation methods. Numerical Algorithms, 68(4): 741–768, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9870-8.

Ferrandi:2025:LMG

[FH25] Giulia Ferrandi and Michiel E. Hochstenbach. Limited memory gradient methods for unconstrained optimization. Numerical Algorithms, 99(2):735–764, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01895-9.

Farooq:2015:SSN

[FHAL15] U. Farooq, T. Hayat, A. Alsaedi, and S. J. Liao. Series solutions of non-similarity boundary layer flows of nano-fluids over stretching surfaces. Numerical Algorithms, 70(1):43–59, September 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9934-9.

Feng:2021:SND

[FHC21] Xinlong Feng, Ruijian He, and Zhangxin Chen. Superconvergence in H^{1}-norm of a difference finite element method for the heat equation in a 3D spatial domain with almostuniform mesh. Numerical Algorithms, 86(1):357–395, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00892-y.

Fischer:1996:NCG

[FHH96] Bernd Fischer, Martin Hanke, and Marlis Hochbruck. A note on conjugate-gradient type methods for indefinite and/or inconsistent linear systems. Numerical Algorithms, 11(1–4):181– 187, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Fierro:1999:UTM

[FHH99] Ricardo D. Fierro, Per Christian Hansen, and Peter Søren Kirk Hansen. UTV Tools: Matlab templates for rank-revealing UTV decompositions. Numerical Algorithms, 20(2–3):165–194, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/9/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/9/fulltext.pdf.

Farouki:2005:BEA

[FHH05] Rida T. Farouki, Chang Yong Han, and Joel Hass. Boundary evaluation algorithms for Minkowski combinations of complex sets using topological analysis of implicit curves. Numerical Algorithms, 40(3):251–283, November 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=3&spage=251.

Fang:2021:EPT

[FHL21] Yonglei Fang, Xianfa Hu, and Jiyong Li. Explicit pseudo twostep exponential Runge–Kutta methods for the numerical integration of first-order differential equations. Numerical Algorithms, 86(3):1143–1163, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00927-
4.

Fazeli:2012:MHC

[FHS12] Somayyeh Fazeli, Gholamreza Hojjati, and Sedaghat Shahmorad. Multistep Hermite collocation methods for solving Volterra Integral Equations. Numerical Algorithms, 60(1):27– 50, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
60&issue=1&spage=27.

Filipuk:2015:CRC

[FHV15] Galina Filipuk, Maciej Haneczok, and Walter Van Assche. Computing recurrence coefficients of multiple orthogonal polynomials. Numerical Algorithms, 70(3):519–543, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9959-8.

Fiorot:1996:BAA

[FJ96] J.-C. Fiorot and P. Jeannin. A blossoming approach to accuracy of the degree elevation process. Numerical Algorithms, 13 (3–4):265–306, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fan:2025:ITB

[FJ25] Xin Fan and Ji-Teng Jia. An incomplete tridiagonalizationbased determinant evaluation for a generalized periodic tridiagonal matrix. Numerical Algorithms, 100(3):1023–1043, November 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01978-7.

Fiorot:1994:NCM

[FJT94] Jean-Charles Fiorot, Pierre Jeannin, and Salim Taleb. New control massic polygon of a B-rational curve resulting from a homographic change of parameter. Numerical Algorithms, 6(3– 4):379–418, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Farouki:2001:ESC

[FKMS01] Rida T. Farouki, Bethany K. Kuspa, Carla Manni, and Alessandra Sestini. Efficient solution of the complex quadratic tridiagonal system for C2 PH quintic splines. Numerical Algorithms, 27(1):35–60, May 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.lwwonline.com/content/getfile/5058/33/2/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/33/2/fulltext.pdf.

Fenn:2006:FET

[FKP06] Markus Fenn, Stefan Kunis, and Daniel Potts. Fast evaluation of trigonometric polynomials from hyperbolic crosses. Numerical Algorithms, 41(4):339–352, April 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=4&spage=339.

Forcadel:2008:GFM

[FLG08] Nicolas Forcadel, Carole Le Guyader, and Christian Gout. Generalized fast marching method: applications to image segmentation. Numerical Algorithms, 48(1–3):189–211, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=1&
spage=189.

Fausten:2004:ACT

[FLH04] Daniela Fausten, Wolfram Luther, and Gerhard Haßlinger. Accurate computation of traffic workload distributions. Numerical Algorithms, 37(1–4):137–146, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/9/abstract.htm.

Feng:2025:OEE

[FLH25] Xinlong Feng, Xiaoli Lu, and Yinnian He. Optimal error estimates of penalty difference finite element method for the 3D steady Navier–Stokes equations. Numerical Algorithms, 98 (3):1395–1427, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01838-4.

Favati:1999:IPB

[FLMR99] P. Favati, G. Lotti, O. Menchi, and F. Romani. An infinite precision bracketing algorithm with guaranteed convergence. Numerical Algorithms, 20(1):63–73, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/17/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
17/2/fulltext.pdf.

Favati:2000:SAC

[FLMR00] P. Favati, G. Lotti, O. Menchi, and F. Romani. Separable asymptotic cost of evaluating elementary functions. Numerical Algorithms, 24(3):255–274, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Floater:2003:ACR

[Flo03]  Michael S. Floater.  Analysis of curve reconstruction

by meshless parameterization. Numerical Algorithms, 32 (1):87–98, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
44/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/44/6/fulltext.pdf.

Floater:2016:GMV

[Flo16] Michael S. Floater. A generalized mean value property for polyharmonic functions. Numerical Algorithms, 73(1):157– 165, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0090-7.

Frohner:2001:DCS

[FLR01] Anja Fröhner, Torsten Linß, and Hans-Görg Roos. Defect correction on Shishkin-type meshes. Numerical Algorithms, 26(3):281–299, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
2/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/2/5/fulltext.pdf.

Faber:2009:ORH

[FLT09] V. Faber, J. Liesen, and P. Tichý. On orthogonal reduction to Hessenberg form with small bandwidth. Numerical Algorithms, 51(2):133–142, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=2&spage=133.

Tributes to Gene H. Golub, Part II.

Ferranti:2014:CBC

[FLV14] Micol Ferranti, Thanh Hieu Le, and Raf Vandebril. A comparison between the complex symmetric based and classical computation of the singular value decomposition of normal matrices. Numerical Algorithms, 67(1):109–120, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9777-9.

Ling:2012:NDF

[fLxX12] Ai fan Ling and Cheng xian Xu. A new discrete filled function method for solving large scale max-cut problems. Numerical Algorithms, 60(3):435–461, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=3&spage=435.

Flynn:2022:PAM

[Fly22] Thomas Flynn. A persistent adjoint method with dynamic time-scaling and an application to mass action kinetics. Numerical Algorithms, 89(1):87–113, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01107-8.

Fang:2025:EEP

[FLY25] Yonglei Fang, Changying Liu, and Xiong You. Extended explicit pseudo two-step Runge–Kutta–Nyström methods for general second-order oscillatory systems. Numerical Algorithms, 99(2):765–789, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01896-
8.

Fischer:1993:ACL

[FM93] Bernd Fischer and Jan Modersitzki. An algorithm for complex linear approximation based on semi-infinite programming. Numerical Algorithms, 5(1–4):287–297, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Fischer:1999:FIM

[FM99] Bernd Fischer and Jan Modersitzki. Fast inversion of matrices arising in image processing. Numerical Algorithms, 22(1):1–11, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
21/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/21/3/fulltext.pdf.

Fampa:2004:UCF

[FM04] Marcia Fampa and Nelson Maculan. Using a conic formulation for finding Steiner minimal trees. Numerical Algorithms, 35(2–4):315–330, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/50/A/8/
abstract.htm.

Fassino:2016:MPI

[FM16] Claudia Fassino and Hans Michael Möller. Multivariate polynomial interpolation with perturbed data. Numerical Algorithms, 71(2):273–292, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9992-7.

Fenzi:2019:PCA

[FM19] Luca Fenzi and Wim Michiels. Polynomial (chaos) approximation of maximum eigenvalue functions. Numerical Algorithms, 82(4):1143–1169, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fidalgo:2018:CIQ

[FMD18] Ulises Fidalgo and Erwin Miña-Díaz. Convergent interpolatory quadrature rules and orthogonal polynomials of varying measures. Numerical Algorithms, 79(2):423–435, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fouladi:2023:OSL

[FMD23] Somayeh Fouladi, Reza Mokhtari, and Mohammad Shafi Dahaghin. Operator-splitting local discontinuous Galerkin method for multi-dimensional linear convection-diffusion equations. Numerical Algorithms, 92(2):1425–1449, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01347-2.

Frittelli:2025:VVE

[FMS25] Massimo Frittelli, Anotida Madzvamuse, and Ivonne Sgura. VEMcomp: a virtual elements MATLAB package for bulksurface PDEs in 2D and 3D. Numerical Algorithms, 99 (3):1365–1392, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01919-4.

Fang:2019:CPK

[FNS19] Zhi-Wei Fang, Michael K. Ng, and Hai-Wei Sun. Circulant preconditioners for a kind of spatial fractional diffusion equations. Numerical Algorithms, 82(2):729–747, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Forbes:1993:GRP

[For93] A. B. Forbes. Generalized regression problems in metrology. Numerical Algorithms, 5(1–4):523–533, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fornberg:2021:GTR

[For21] Bengt Fornberg. Generalizing the trapezoidal rule in the complex plane. Numerical Algorithms, 87(1):187–202, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00963-0.

Fornberg:2022:FDF

[For22] Bengt Fornberg. Finite difference formulas in the complex plane. Numerical Algorithms, 90(3):1305–1326, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01231-5.

Ferronato:2018:SBF

[FP18] Massimiliano Ferronato and Giorgio Pini. A supernodal block factorized sparse approximate inverse for non-symmetric linear systems. Numerical Algorithms, 78(1):333–354, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Floater:2020:TMV

[FP20] Michael S. Floater and Francesco Patrizi. Transfinite mean value interpolation over polygons. Numerical Algorithms, 85 (3):995–1003, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00849-w.

FernAndez:2005:COP

[FPP05] Lidia Fernández, Teresa E. Pérez, and Miguel A. Piñar. Classical orthogonal polynomials in two variables: a matrix approach. Numerical Algorithms, 39(1–3):131–142, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fuhry:2012:NTR

[FR12] Martin Fuhry and Lothar Reichel. A new Tikhonov regularization method. Numerical Algorithms, 59(3):433–445, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=3&
spage=433.

Franco:2018:EOE

[FR18] J. M. Franco and L. Rández. An eighth-order exponentially fitted two-step hybrid method of explicit type for solving orbital and oscillatory problems. Numerical Algorithms, 78(1):243– 262, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Francisco:2026:QNI

[Fra26] Juliano B. Francisco. Quasi-Newton interior point method for solving nonlinear system of equations with box constraints. Numerical Algorithms, 101(4):2301–2331, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02088-8.

Fauvet:2009:ACS

[FRJT09] F. Fauvet, F. Richard-Jung, and J. Thomann. Automatic computation of Stokes matrices. Numerical Algorithms, 50(2):179–213, February 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=179.

Floater:2007:EMA

[FRR07] Michael S. Floater, Atgeirr F. Rasmussen, and Ulrich Reif. Extrapolation methods for approximating arc length and surface area. Numerical Algorithms, 44(3):235–248, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=44&issue=3&
spage=235.

Fermo:2021:NTG

[FRS21] Luisa Fermo, Maria Grazia Russo, and Giada Serafini. Numerical treatment of the generalized Love integral equation. Numerical Algorithms, 86(4):1769–1789, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00953-2;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00953-2.pdf.

Ford:2001:NSF

[FS01] Neville J. Ford and A. Charles Simpson. The numerical solution of fractional differential equations: Speed versus accuracy. Numerical Algorithms, 26(4):333–346, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/32/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
32/3/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/32/3/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/32/3/fulltext.pdf.

Ferreira:2016:GSB

[FS16] E. M. Ferreira and J. Sesma. Global solutions of the biconfluent Heun equation. Numerical Algorithms, 71(4):797–809, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0024-4.

Farouki:2020:CRS

[FS20] Rida T. Farouki and Jeffrey A. Strom. Computing the roots of sparse high-degree polynomials that arise from the study of random simplicial complexes. Numerical Algorithms, 83(4): 1653–1670, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Funken:2021:CAA

[FS21] Stefan A. Funken and Anja Schmidt. A coarsening algorithm on adaptive red–green–blue refined meshes. Numerical Algorithms, 87(3):1147–1176, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01003-
7;   http://link.springer.com/content/pdf/10.1007/
s11075-020-01003-7.pdf.

Frommer:2023:CRG

[FS23] Andreas Frommer and Daniel B. Szyld. On the convergence of randomized and greedy relaxation schemes for solving nonsingular linear systems of equations. Numerical Algorithms, 92 (1):639–664, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01431-7.

Feng:2023:ISD

[FSY23] Bao-Feng Feng, Han-Han Sheng, and Guo-Fu Yu. Integrable semi-discretizations and self-adaptive moving mesh method for a generalized sine-Gordon equation. Numerical Algorithms, 94 (1):351–370, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01504-1.

Fischer:2002:OPW

[FT02] Bernd Fischer and Woula Themistoclakis. Orthogonal polynomial wavelets. Numerical Algorithms, 30(1):37–58, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/40/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
40/4/fulltext.pdf.

Fauvet:2005:FNC

[FT05a] F. Fauvet and J. Thomann. Formal and numerical computations with resurgent functions. Numerical Algorithms, 40(4):323–353, December 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=40&issue=4&spage=323.

Ford:2005:SLS

[FT05b] Judith M. Ford and Eugene E. Tyrtyshnikov. Solving linear systems using wavelet compression combined with Kronecker product approximation. Numerical Algorithms, 40(2):125–135, October 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=125.

Fernandez-Torres:2014:DFI

[FT14] Gustavo Fernández-Torres. Derivative free iterative methods with memory of arbitrary high convergence order. Numerical Algorithms, 67(3):565–580, November 2014. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-013-9808-6.

Fang:2024:SPA

[FT24] Rui Fang and Richard Tsai. Stabilization of parareal algorithms for long-time computation of a class of highly oscillatory Hamiltonian flows using data. Numerical Algorithms, 96 (3):1163–1187, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01826-8.

Fuentes:2007:DSD

[Fue07] Marc Fuentes. Diagonalization of the symmetrized discrete i-th right shift operator: an elementary proof. Numerical Algorithms, 44(1):29–43, January 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=1&spage=29.

Funaro:2001:SCC

[Fun01] Daniele Funaro. A superconsistent Chebyshev collocation method for second-order differential operators. Numerical Algorithms, 28(1–4):151–157, December 2001. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/37/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/9/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
9/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/9/fulltext.pdf.

Faghih:2025:CSO

[FVV25] Amin Faghih, Marc Van Barel, and Raf Vandebril. Constructing Sobolev orthonormal rational functions via an updating procedure. Numerical Algorithms, 100(4):1837–1865, December 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02149-y.

Feng:2013:TFA

[FW13] Lei Feng and Weiping Wang. Two families of approximations for the gamma function. Numerical Algorithms, 64(3):403– 416, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9671-x.

Fan:2025:NSE

[FW25] Guangyu Fan and Beibei Wu. Numerical solutions of the EW and MEW equations using a fourth-order improvised Bspline collocation method. Numerical Algorithms, 98(4):1799– 1825, April 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01853-5.

Fan:2016:NSG

[FWC16] Hung-Yuan Fan, Peter Chang-Yi Weng, and Eric King-Wah Chu. Numerical solution to generalized Lyapunov/Stein and rational Riccati equations in stochastic control. Numerical Algorithms, 71(2):245–272, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9991-8.

Fan:2013:OHA

[FY13] Tao Fan and Xiangcheng You. Optimal homotopy analysis method for nonlinear differential equations in the boundary layer. Numerical Algorithms, 62(2):337–354, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9587-5/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=337-354.

Fan:2019:SSI

[FY19] Xiaona Fan and Qinglun Yan. Solving system of inequalities via a smoothing homotopy method. Numerical Algorithms, 82 (2):719–728, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fukuda:2012:EAM

[FYI^{+}12] Akiko Fukuda, Yusaku Yamamoto, Masashi Iwasaki, Emiko Ishiwata, and Yoshimasa Nakamura. Error analysis for matrix eigenvalue algorithm based on the discrete hungry Toda equation. Numerical Algorithms, 61(2):243–260, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=2&
spage=243.

Fang:2014:TFT

[FYM14] Yonglei Fang, Xiong You, and Qinghe Ming. Trigonometrically fitted two-derivative Runge–Kutta methods for solving oscillatory differential equations. Numerical Algorithms, 65 (3):651–667, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9802-z.

Fang:2019:NFS

[FYYW19] Yonglei Fang, Yanping Yang, Xiong You, and Bin Wang. A new family of A-stable Runge–Kutta methods with equationdependent coefficients for stiff problems. Numerical Algorithms, 81(4):1235–1251, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fasshauer:2007:COS

[FZ07] Gregory E. Fasshauer and Jack G. Zhang. On choosing “optimal” shape parameters for RBF approximation. Numerical Algorithms, 45(1–4):345–368, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-007-9072-8.

Fei:2026:SME

[FZ26] Zetao Fei and Hai Zhang. SCAN-MUSIC: An efficient superresolution algorithm for large-scale single snapshot line spectral estimation. Numerical Algorithms, 101(3):2069–2100, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02079-9.

Feng:2016:FEM

[FZL^{+}16] L. B. Feng, P. Zhuang, F. Liu, I. Turner, and Y. T. Gu. Finite element method for space-time fractional diffusion equation. Numerical Algorithms, 72(3):749–767, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0065-8.

Fang:2023:FTT

[FZLL23] Zhichao Fang, Jie Zhao, Hong Li, and Yang Liu. A fast time two-mesh finite volume element algorithm for the nonlinear time-fractional coupled diffusion model. Numerical Algorithms, 93(2):863–898, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01444-2.

Gabutti:2008:EQG

[GA08] Bruno Gabutti and Giampietro Allasia. Evaluation of qgamma function and q-analogues by iterative algorithms. Numerical Algorithms, 49(1–4):159–168, December 2008. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=49&issue=1&
spage=159.

Galantai:2015:ACI

[GA15] A. Galántai and J. Abaffy. Always convergent iteration methods for nonlinear equations of Lipschitz functions. Numerical Algorithms, 69(2):443–453, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9905-1.

Ghafli:2020:OEB

[GA20] Ahmed A. Al Ghafli and Hassan J. Al Salman. An optimal error bound for a finite element approximation of spatially extended predator–prey interaction model. Numerical Algorithms, 85(1):209–229, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00810-
x.

Gajda:2005:SAW

[Gaj05] Piotr Gajda. Smolyak’s algorithm for weighted L_{1}-approximation of multivariate functions with bounded rth mixed derivatives over R^{d}. Numerical Algorithms, 40(4):401–414, December 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=4&
spage=401.

Galligani:1993:SIC

[Gal93] Emanuele Galligani. C^{1} surface interpolation with constraints. Numerical Algorithms, 5(1–4):549–555, November 1993. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Galantai:2018:ACM

[Gal18] A. Galántai. Always convergent methods for nonlinear equations of several variables. Numerical Algorithms, 78(2):625– 641, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Galantai:2022:CNM

[Gal22] Aurél Galántai. Convergence of the Nelder–Mead method. Numerical Algorithms, 90(3):1043–1072, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01221-7.

Garcia-Ardila:2024:APS

[GAM24] Juan C. García-Ardila and Misael E. Marriaga. Approximation by polynomials in Sobolev spaces associated with classical moment functionals. Numerical Algorithms, 95(1):285– 318, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01572-3.

Garcia:2019:ARN

[Gar19] G. García. Approximating roots of nonlinear systems by αdense curves. Numerical Algorithms, 82(3):749–760, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Garg:2020:CUM

[Gar20] Naveen Kumar Garg. A class of upwind methods based on generalized eigenvectors for weakly hyperbolic systems. Numerical Algorithms, 83(3):1091–1121, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Garcia:2024:NMA

[Gar24] G. García. A numerical method to approximate the solutions of nonlinear systems in densifiable sets. Numerical Algorithms, 96 (4):1925–1943, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01690-y.

Gaspar:1999:MTB

[Gás99] C. Gáspár. Multigrid technique for biharmonic interpolation with application to dual and multiple reciprocity method. Numerical Algorithms, 21(1–4):165–183, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/21/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
21/fulltext.pdf.

Gautschi:1995:WPR

[Gau95] Walter Gautschi. The work of Philip Rabinowitz on numerical integration. Numerical Algorithms, 9(3–4):199–222, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gautschi:2000:HOG

[Gau00]  Walter Gautschi.  High-order Gauss–Lobatto formulae.

Numerical Algorithms, 25(1–4):213–222, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/3/20/abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/5058/3/20/fulltext.

pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Gautschi:2007:GTT

[Gau07] Walter Gautschi. A guided tour through my bibliography. Numerical Algorithms, 45(1–4):11–35, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=11.

Gautschi:2008:NEC

[Gau08a] Walter Gautschi. The numerical evaluation of a challenging integral. Numerical Algorithms, 49(1–4):187–194, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=49&issue=1&
spage=187.

Gautschi:2008:CIL

[Gau08b] Walter Gautschi. On a conjectured inequality for the largest zero of Jacobi polynomials. Numerical Algorithms, 49(1–4):195–198, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=195.

Gautschi:2009:HOG

[Gau09a] Walter Gautschi. High-order generalized Gauss–Radau and Gauss–Lobatto formulae for Jacobi and Laguerre weight functions. Numerical Algorithms, 51(2):143–149, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=2&

spage=143. Tributes to Gene H. Golub, Part II.

Gautschi:2009:NCI

[Gau09b] Walter Gautschi. New conjectured inequalities for zeros of Jacobi polynomials. Numerical Algorithms, 50(3):293–296, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
50&issue=3&spage=293. See remark [Gau11b].

Gautschi:2009:CIZ

[Gau09c] Walter Gautschi. On conjectured inequalities for zeros of Jacobi polynomials. Numerical Algorithms, 50(1):93–96, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.

asp?genre=article&issn=1017-1398&volume=50&issue=1&
spage=93.

Gautschi:2009:VPR

[Gau09d] Walter Gautschi. Variable-precision recurrence coefficients for nonstandard orthogonal polynomials. Numerical Algorithms, 52(3):409–418, November 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=409.

Gautschi:2010:GQR

[Gau10] Walter Gautschi. Gauss quadrature routines for two classes of logarithmic weight functions. Numerical Algorithms, 55(2–3):265–277, November 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=265.

Gautschi:2011:LWF

[Gau11a] Walter Gautschi. The Lambert W -functions and some of their integrals: a case study of high-precision computation. Numerical Algorithms, 57(1):27–34, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=1&spage=27.

Gautschi:2011:RSC

[Gau11b] Walter Gautschi. Remark on “New conjectured inequalities for zeros of Jacobi polynomials” by Walter Gautschi, Numer. Algorithms 50:293–296 (2009). Numerical Algorithms, 57(4):511, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
57&issue=4&spage=511. See [Gau09b].

Gautschi:2012:NIS

[Gau12a] Walter Gautschi. Numerical integration over the square in the presence of algebraic/logarithmic singularities with an application to aerodynamics. Numerical Algorithms, 61(2):275–290, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=2&spage=275. See erratum [Gau13a].

Gautschi:2012:SRJ

[Gau12b] Walter Gautschi. Sub-range Jacobi polynomials. Numerical Algorithms, 61(4):649–657, December 2012. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9556-z/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=4&spage=649-657. See erratum [Gau17a].

Gautschi:2013:ENI

[Gau13a] Walter Gautschi. Erratum to: Numerical integration over the square in the presence of algebraic/logarithmic singularities with an application to aerodynamics. Numerical Algorithms, 64(4):759, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/link.springer.com/article/10.1007/s11075-013-9787-
7;   http://link.springer.com/content/pdf/10.1007/
s11075-013-9787-7.pdf. See [Gau12a].

Gautschi:2013:NNS

[Gau13b] Walter Gautschi. Neutralizing nearby singularities in numerical quadrature. Numerical Algorithms, 64(3):417–425, November 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9672-9.

Gautschi:2013:RMO

[Gau13c] Walter Gautschi. Repeated modifications of orthogonal polynomials by linear divisors. Numerical Algorithms, 63(2):369– 383, June 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9627-1.

Gautschi:2014:HPG

[Gau14] Walter Gautschi. High-precision Gauss–Turán quadrature rules for Laguerre and Hermite weight functions. Numerical Algorithms, 67(1):59–72, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9774-z.

Gautschi:2015:POR

[Gau15] Walter Gautschi. Polynomials orthogonal with respect to exponential integrals. Numerical Algorithms, 70(1):215–226, September 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9943-8.

Gautschi:2017:ESR

[Gau17a] Walter Gautschi. Erratum to: Sub-range Jacobi polynomials. Numerical Algorithms, 74(2):637, February 2017. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-016-0257-x.pdf. See [Gau12b].

Gautschi:2017:POR

[Gau17b] Walter Gautschi. Polynomials orthogonal with respect to cardinal B-spline weight functions. Numerical Algorithms, 76(4): 1099–1107, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gautschi:2018:ZSJ

[Gau18] Walter Gautschi. On the zeros of subrange Jacobi polynomials. Numerical Algorithms, 79(3):759–768, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See correction [Gau19].

Gautschi:2019:CSR

[Gau19] Walter Gautschi. Correction to: Sub-range Jacobi polynomials. Numerical Algorithms, 81(2):771, June 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-019-00715-9.pdf. See [Gau18].

Gautschi:2022:ALP

[Gau22] Walter Gautschi. Another look at polynomials orthogonal relative to exponential integral weight functions. Numerical Algorithms, 91(4):1547–1557, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01313-y.

Gande:2021:MTF

[GB21] Naga Raju Gande and Ashlesha A. Bhise. Modified third and fifth order WENO schemes for inviscid compressible flows. Numerical Algorithms, 88(1):249–279, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01039-9.

Gonzaga:2004:PCP

[GC04] Clovis C. Gonzaga and Marli Cardia. Properties of the central points in linear programming problems. Numerical Algorithms, 35(2–4):185–204, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/50/A/15/abstract.htm.

Gonzalez-Concepcion:1995:AIT

[GCFF95] C. González-Concepción, V. Cano Fernández, and C. Gil Fariña. The ϵ-algorithm for the identification of a transferfunction model: some applications. Numerical Algorithms, 9 (3–4):379–395, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gonzalez-Concepcion:2003:PAE

[GCGF03] Concepción González-Concepción and María Candelaria Gil-Fariña. Padé approximation in economics. Numerical Algorithms, 33(1–4):277–292, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
23/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/23/fulltext.pdf.

Gonzalez-Concepcion:1992:MOL

[GCGVH92] C. González-Concepción, P. González-Vera, and E. Hendriksen. Matrix orthogonal Laurent polynomials and two-point Padé approximants. Numerical Algorithms, 3(1–4):201–209, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Gu:2026:ERB

[GCJ26] Jiaxi Gu, Xinjuan Chen, and Jae-Hun Jung. Explicit radial basis function Runge–Kutta methods. Numerical Algorithms, 102(1):53–84, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02121-w.

Gonzalez-Concepcion:1999:ASR

[GCPG99] Concepción González-Concepción and Celina Pestano-Gabino. Approximated solutions in rational form for systems of differential equations. Numerical Algorithms, 21(1–4): 185–203, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
14/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/14/fulltext.pdf.

Gu:2024:FLV

[GCZH24] Qiling Gu, Yanping Chen, Jianwei Zhou, and Jian Huang. A fast linearized virtual element method on graded meshes for nonlinear time-fractional diffusion equations. Numerical Algorithms, 97(3):1141–1177, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01744-1.

Gander:2015:OSM

[GD15a] Martin J. Gander and Olivier Dubois. Optimized Schwarz methods for a diffusion problem with discontinuous coefficient. Numerical Algorithms, 69(1):109–144, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9884-2.

Garshasbi:2015:EUB

[GD15b] M. Garshasbi and H. Dastour. Estimation of unknown boundary functions in an inverse heat conduction problem using a mollified marching scheme. Numerical Algorithms, 68(4): 769–790, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9871-7.

Gursoy:2020:PTI

[GEA20] Faik Gürsoy, Müzeyyen Ertürk, and Mujahid Abbas. A Picardtype iterative algorithm for general variational inequalities and nonexpansive mappings. Numerical Algorithms, 83(3):867– 883, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gemignani:1997:CRI

[Gem97] Luca Gemignani. Chebyshev rational interpolation. Numerical Algorithms, 15(1):91–110, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
30/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/30/6/fulltext.pdf.

Gensane:2012:IHT

[Gen12] Thierry Gensane. Interpolation on the hypersphere with Thiele type rational interpolants. Numerical Algorithms, 60(3):523– 529, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
60&issue=3&spage=523.

Garcia-Esnaola:2014:EBL

[GEP14] Marta García-Esnaola and Juan Manuel Peña. Error bounds for linear complementarity problems of Nekrasov matrices. Numerical Algorithms, 67(3):655–667, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9815-7.

Garcia-Esnaola:2016:BNM

[GEP16] M. García-Esnaola and J. M. Peña. B-Nekrasov matrices and error bounds for linear complementarity problems. Numerical Algorithms, 72(2):435–445, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0054-y.

Garcia-Esnaola:2019:AOE

[GEP19] M. García-Esnaola and J. M. Peña. On the asymptotic optimality of error bounds for some linear complementarity problems. Numerical Algorithms, 80(2):521–532, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Griewank:2002:RFG

[GF02] Andreas Griewank and Christèle Faure. Reduced functions, gradients and Hessians from fixed-point iterations for state equations. Numerical Algorithms, 30(2):113–139, June 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/41/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
41/3/fulltext.pdf.

Gerstner:1998:NIU

[GG98] Thomas Gerstner and Michael Griebel. Numerical integration using sparse grids. Numerical Algorithms, 18(3– 4):209–232, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/4/fulltext.pdf.

Gabutti:2001:NAZ

[GG01] Bruno Gabutti and Luigi Gatteschi. New asymptotics for the zeros of Whittaker’s functions. Numerical Algorithms, 28 (1–4):159–170, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/37/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/37/10/fulltext.pdf;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/10/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/10/fulltext.pdf.

Gautschi:2008:LGW

[GG08] Walter Gautschi and Carla Giordano. Luigi Gatteschi’s work on asymptotics of special functions and their zeros. Numerical Algorithms, 49(1–4):11–31, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=11.

Garg:2022:GLP

[GG22] Deepika Garg and Sashikumaar Ganesan. Generalized local projection stabilized nonconforming finite element methods for Darcy equations. Numerical Algorithms, 89(1):341–369, January 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01117-6.

Gao:2024:UTB

[GGJL24] Lei Gao, Xian-Ming Gu, Xiudan Jia, and Chaoqian Li. Upper triangulation-based infinity norm bounds for the inverse of Nekrasov matrices with applications. Numerical Algorithms, 97(4):1453–1479, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01758-3.

Gajny:2014:PSA

[GGN14] Laurent Gajny, Olivier Gibaru, and Eric Nyiri. L_{1}C^{1} polynomial spline approximation algorithms for large data sets. Numerical Algorithms, 67(4):807–826, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9828-x.

Gajny:2018:SFS

[GGN18] Laurent Gajny, Olivier Gibaru, and Eric Nyiri. L_{1} spline fits via sliding window process: continuous and discrete cases. Numerical Algorithms, 78(2):449–464, June 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gajny:2017:BAH

[GGNF17] Laurent Gajny, Olivier Gibaru, Eric Nyiri, and Shu-Cherng Fang. Best L_{1} approximation of Heaviside-type functions from Chebyshev and weak-Chebyshev spaces. Numerical Algorithms, 75(3):827–843, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Grapiglia:2022:CRN

[GGS22] G. N. Grapiglia, M. L. N. Gonçalves, and G. N. Silva. A cubic regularization of Newton’s method with finite difference Hessian approximations. Numerical Algorithms, 90(2):607– 630, June 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01200-y.

Gallivan:1996:RLA

[GGV96] K. Gallivan, E. Grimme, and P. Van Dooren. A rational Lanczos algorithm for model reduction. Numerical Algorithms, 12 (1–2):33–63, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Golub:2002:BOT

[GGV02] Gene H. Golub, Chen Greif, and James M. Varah. Block orderings for tensor-product grids in two and three dimensions. Numerical Algorithms, 30(2):93–111, June 2002. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/41/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/41/1/fulltext.pdf.

Gifford:1995:BAC

[GH95] David Gifford and Robert Huotari. Best L_{1} approximation by convex functions. Numerical Algorithms, 9(1–2):107–111, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Ganesh:2006:SAA

[GH06] M. Ganesh and S. C. Hawkins. A spectrally accurate algorithm for electromagnetic scattering in three dimensions. Numerical Algorithms, 43(1):25–60, September 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=1&spage=25.

Ganesh:2009:HOA

[GH09a] M. Ganesh and S. C. Hawkins. A high-order algorithm for multiple electromagnetic scattering in three dimensions. Numerical Algorithms, 50(4):469–510, April 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=4&spage=469.

Ghoreishi:2009:NCT

[GH09b] F. Ghoreishi and M. Hadizadeh. Numerical computation of the Tau approximation for the Volterra–Hammerstein integral equations. Numerical Algorithms, 52(4):541–559, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=4&
spage=541.

Giraud:2009:PAH

[GH09c]  L. Giraud and A. Haidar.  Parallel algebraic hybrid

solvers for large 3D convection-diffusion problems. Numerical Algorithms, 51(2):151–177, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=2&spage=151.

Tributes to Gene H. Golub, Part II.

Galantai:2010:SAN

[GH10] Aurél Galántai and Csaba J. Hegedüs. A study of accelerated Newton methods for multiple polynomial roots. Numerical Algorithms, 54(2):219–243, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=2&spage=219.

Guglielmi:2022:EAS

[GH22] Nicola Guglielmi and Ernst Hairer. An efficient algorithm for solving piecewise-smooth dynamical systems. Numerical Algorithms, 89(3):1311–1334, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01154-
1;   http://link.springer.com/content/pdf/10.1007/
s11075-021-01154-1.pdf.

Georgieva:2023:NMB

[GH23a] Irina Georgieva and Clemens Hofreither. A Newton method for best uniform rational approximation. Numerical Algorithms, 93(4):1741–1758, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01487-
5.

Guglielmi:2023:SMH

[GH23b] Nicola Guglielmi and Ernst Hairer. Sliding modes of high codimension in piecewise-smooth dynamical systems. Numerical Algorithms, 94(1):257–273, September 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01499-9.

Georgieva:2025:CPR

[GH25] Irina Georgieva and Clemens Hofreither. Computation of polynomial and rational approximations in complex domains by the τ-method. Numerical Algorithms, 99(2):791–808, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01897-7.

Ghasemi:2016:UCS

[Gha16] M. Ghasemi. On using cubic spline for the solution of problems in calculus of variations. Numerical Algorithms, 73(3):685– 710, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0113-z.

Ghasemi:2018:NED

[Gha18] M. Ghasemi. A new efficient DQ algorithm for the solution of elliptic problems in higher dimensions. Numerical Algorithms, 77(3):809–829, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gharibi:2024:WGP

[Gha24] Zeinab Gharibi. A weak Galerkin pseudostress-based mixed finite element method on polygonal meshes: application to the Brinkman problem appearing in porous media. Numerical Algorithms, 97(3):1341–1366, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01752-9.

Gao:2015:FDM

[GHC15] Qin Gao, Zhengda Huang, and Xiaoliang Cheng. A finite difference method for an inverse Sturm–Liouville problem in impedance form. Numerical Algorithms, 70(3):669– 690, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9968-7.

Gheorghiu:2013:LCS

[Ghe13] Calin-Ioan Gheorghiu. Laguerre collocation solutions to boundary layer type problems. Numerical Algorithms, 64(2): 385–401, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9670-y.

Gheorghiu:2015:PSS

[Ghe15] Calin-Ioan Gheorghiu. Pseudospectral solutions to some singular nonlinear BVPs. Numerical Algorithms, 68(1):1–14, January 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9834-z.

Gheorghiu:2016:SCS

[Ghe16] Calin-Ioan Gheorghiu. Spectral collocation solutions to systems of boundary layer type. Numerical Algorithms, 73(1): 1–14, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0083-6.

Gheorghiu:2018:NTE

[Ghe18] Calin-Ioan Gheorghiu. On the numerical treatment of the eigenparameter dependent boundary conditions. Numerical Algorithms, 77(1):77–93, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Garza:2008:STM

[GHM08] Luis Garza, Javier Hernández, and Francisco Marcellán. Spectral transformations of measures supported on the unit circle and the Szegő transformation. Numerical Algorithms, 49(1–4):169–185, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=169.

Gander:2016:NAB

[GHM16] Martin J. Gander, Laurence Halpern, and Véronique Martin. A new algorithm based on factorization for heterogeneous domain decomposition. Numerical Algorithms, 73(1): 167–195, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0091-6.

Gibbs:2023:NQS

[GHM23] Andrew Gibbs, David Hewett, and Andrea Moiola. Numerical quadrature for singular integrals on fractals. Numerical Algorithms, 92(4):2071–2124, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01378-
9.

Gibbs:2024:NES

[GHM24] A. Gibbs, D. P. Hewett, and B. Major. Numerical evaluation of singular integrals on non-disjoint self-similar fractal sets. Numerical Algorithms, 97(1):311–343, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01705-8.

Gazzola:2019:ITM

[GHN19] Silvia Gazzola, Per Christian Hansen, and James G. Nagy. IR tools: a MATLAB package of iterative regularization methods and large-scale test problems. Numerical Algorithms, 81 (3):773–811, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Goudin:2000:PSL

[GHP^{+}00] David Goudin, Pascal Hénon, François Pellegrini, Pierre Ramet, Jean Roman, and Jean-Jacques Pesqué. Parallel sparse linear algebra and application to structural mechanics. Numerical Algorithms, 24(4):371–391, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/29/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
29/5/fulltext.pdf.

Gutierrez:2014:IMR

[GHPMGRR14] José M. Gutiérrez, Luis J. Hernández-Paricio, Miguel Marañón-Grandes, and M. Teresa Rivas-Rodríguez. Influence of the multiplicity of the roots on the basins of attraction of Newton’s method. Numerical Algorithms, 66(3):431–455, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9742-7.

Guo:2025:FEM

[GHZ25] Huijiao Guo, Jin Huang, and Xueli Zhang. The fast Euler– Maruyama method for solving multiterm Caputo fractional stochastic delay integro-differential equations. Numerical Algorithms, 99(3):1551–1600, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01925-
6.

Gutzmer:1997:DDS

[GI97] Tim Gutzmer and Armin Iske. Detection of discontinuities in scattered data approximation. Numerical Algorithms, 16(2):155–170, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/5/fulltext.pdf.

Greco:2010:BPS

[GI10] Federico Greco and Bruno Iannazzo. A binary powering Schur algorithm for computing primary matrix roots. Numerical Algorithms, 55(1):59–78, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=1&spage=59.

Ghili:2015:RCP

[GI15] Saman Ghili and Gianluca Iaccarino. Reusing Chebyshev points for polynomial interpolation. Numerical Algorithms, 70(2):249–267, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9945-6.

Garijo:2017:EAC

[GJV17] Antonio Garijo, Xavier Jarque, and Jordi Villadelprat. An effective algorithm to compute Mandelbrot sets in parameter planes. Numerical Algorithms, 76(2):555–571, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Guo:2024:CSP

[GJVW24] Zhenwei Guo, Tongsong Jiang, V. I. Vasil’ev, and Gang Wang. Complex structure-preserving method for Schrödinger equations in quaternionic quantum mechanics. Numerical Algorithms, 97(1):271–287, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01703-
w.

Gholampoor:2020:DNM

[GK20] I. Gholampoor and M. Tavassoli Kajani. A direct numerical method for approximate solution of inverse reaction diffusion equation via two-dimensional Legendre hybrid functions. Numerical Algorithms, 83(2):511–528, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gu:2021:SCM

[GK21] Zhendong Gu and Yinying Kong. Spectral collocation method for Caputo fractional terminal value problems. Numerical Algorithms, 88(1):93–111, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01031-
3.

Grajewski:2024:DAD

[GK24] Matthias Grajewski and Andreas Kleefeld. Detecting and approximating decision boundaries in low-dimensional spaces. Numerical Algorithms, 95(4):1503–1537, April 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01618-6.

Kang:2025:BKT

[gK25] Chuan gang Kang. On the block Kaczmarz–Tanabe methods with relaxation parameters for solving linear systems. Numerical Algorithms, 99(3):1063–1079, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01907-8.

Gdawiec:2021:NMF

[GKL21] Krzysztof Gdawiec, Wiesław Kotarski, and Agnieszka Lisowska. Newton’s method with fractional derivatives and various iteration processes via visual analysis. Numerical Algorithms, 86(3):953–1010, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00919-
4;   http://link.springer.com/content/pdf/10.1007/
s11075-020-00919-4.pdf.

Grimm:2024:DGS

[GKQ24] Volker Grimm, Tobias Kliesch, and G. R. W. Quispel. Discrete gradients in short-range molecular dynamics simulations. Numerical Algorithms, 96(3):1189–1220, July 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01717-4.

Garcia:2022:PIT

[GKRS22] Idoia Cortes Garcia, Iryna Kulchytska-Ruchka, and Sebastian Schöps. Parareal for index two differential algebraic equations. Numerical Algorithms, 91(1):389–412, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01267-1.

Gil:2004:CRZ

[GKS04] Amparo Gil, Wolfram Koepf, and Javier Segura. Computing the real zeros of hypergeometric functions. Numerical Algorithms, 36(2):113–134, June 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/54/A/2/abstract.htm.

Grammont:2023:FAS

[GKV23] Laurence Grammont, Rekha P. Kulkarni, and Paulo B. Vasconcelos. Fast and accurate solvers for weakly singular integral equations. Numerical Algorithms, 92(4):2045–2070, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01376-x.

Gemignani:2004:REA

[GL04] Luca Gemignani and Grazia Lotti. Rounding error analysis in solving M-matrix linear systems of block Hessenberg form. Numerical Algorithms, 36(2):157–168, June 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/54/A/4/abstract.htm.

Gautschi:2007:CIJ

[GL07] Walter Gautschi and Paul Leopardi. Conjectured inequalities for Jacobi polynomials and their largest zeros. Numerical Algorithms, 45(1–4):217–230, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-007-9067-5.

Gaviano:2012:PNT

[GL12] Marco Gaviano and Daniela Lera. Properties and numerical testing of a parallel global optimization algorithm. Numerical Algorithms, 60(4):613–629, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=613.

Greer:2015:OSM

[GL15] Neil Greer and Sébastien Loisel. The optimised Schwarz method and the two-Lagrange multiplier method for heterogeneous problems in general domains with two general subdomains. Numerical Algorithms, 69(4):737–762, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9923-z.

Guo:2019:ASS

[GL19] Ping Guo and Chong-Jun Li. Almost sure stability with general decay rate of exact and numerical solutions for stochastic pantograph differential equations. Numerical Algorithms, 80 (4):1391–1411, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Guo:2020:SSM

[GL20] Chun-Hua Guo and Di Lu. A study of Schröder’s method for the matrix p-th root using power series expansions. Numerical Algorithms, 83(1):265–279, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gruberger:2021:TAP

[GL21] Nira Gruberger and David Levin. Two algorithms for periodic extension on uniform grids. Numerical Algorithms, 86(2): 475–494, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00897-7.

Guo:2023:ITS

[GL23] Zhen-Chen Guo and Xin Liang. The intrinsic Toeplitz structure and its applications in algebraic Riccati equations. Numerical Algorithms, 93(1):227–267, May 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01413-9.

Glader:2001:MRC

[Gla01] Christer Glader. A method for rational Chebyshev approximation of rational functions on the unit disk and on the unit interval. Numerical Algorithms, 26(2):151–165, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/1/5/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/1/
5/fulltext.pdf.

Glazunov:2004:ICT

[Gla04] Nikolaj Glazunov. Interval computations and their categorification. Numerical Algorithms, 37(1–4):159–164, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/11/abstract.htm.

Gu:2022:HTA

[GLC22] Xian-Ming Gu, Siu-Long Lei, and Bruno Carpentieri. A Hessenberg-type algorithm for computing PageRank problems. Numerical Algorithms, 89(4):1845–1863, April 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01175-w.

Gonzalez-Lima:2009:NLM

[GLdO09] María D. González-Lima and Flor Montes de Oca. A Newtonlike method for nonlinear system of equations. Numerical Algorithms, 52(3):479–506, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=479.

Gong:2026:BLM

[GLG26] Ziche Gong, Yuan Lei, and Guangcai Gong. A block Lanczos method for the symmetric generalized eigenvalue complementarity problem. Numerical Algorithms, 102(2):1237– 1264, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02176-9.

Guo:2019:MNI

[GLL19] Chun-Hua Guo, Wen-Wei Lin, and Ching-Sung Liu. A modified Newton iteration for finding nonnegative Z-eigenpairs of a nonnegative tensor. Numerical Algorithms, 80(2):595–616, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gao:2012:OFH

[GLLJ12] Wei Gao, Hong Li, Yang Liu, and Yong-Jun Jian. An oscillation-free high order TVD/CBC-based upwind scheme for convection discretization. Numerical Algorithms, 59(1):29– 50, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
59&issue=1&spage=29.

Graillat:2015:MRE

[GLM15] Stef Graillat, Vincent Lefèvre, and Jean-Michel Muller. On the maximum relative error when computing integer powers by iterated multiplications in floating-point arithmetic. Numerical Algorithms, 70(3):653–667, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9967-8.

Gout:2008:ASF

[GLRSG08] C. Gout, C. Le Guyader, L. Romani, and A.-G. Saint-Guirons. Approximation of surfaces with fault(s) and/or rapidly varying data, using a segmentation process, D^{m}-splines and the finite element method. Numerical Algorithms, 48(1–3):67–92, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=1&
spage=67.

Guo:2018:DAT

[GLS^{+}18] Dongsheng Guo, Xinjie Lin, Zhaozhu Su, Sibo Sun, and Zhijing Huang. Design and analysis of two discrete-time ZD algorithms for time-varying nonlinear minimization. Numerical Algorithms, 77(1):23–36, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gao:2026:AID

[GLS26] Lei Gao, Tianyuan Li, and Xinnian Song. Accurate inverses and determinants of doubly strictly diagonally dominant Z-matrices. Numerical Algorithms, 102(2):1065–1085, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02169-8.

Gout:2005:SUG

[GLV05] Christian Gout, Carole Le Guyader, and Luminita Vese. Segmentation under geometrical conditions using geodesic active contours and interpolation using level set methods. Numerical Algorithms, 39(1–3):155–173, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gu:2013:PRC

[GLW13] Ming Gu, Lek-Heng Lim, and Cinna Julie Wu. ParNes: a rapidly convergent algorithm for accurate recovery of sparse and approximately sparse signals. Numerical Algorithms, 64 (2):321–347, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9668-5.

Gospodarczyk:2016:CMD

[GLW16] Przemysław Gospodarczyk, Stanisław Lewanowicz, and Pawel Woźny. G^{k,l}-constrained multi-degree reduction of Bézier curves. Numerical Algorithms, 71(1):121–137, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9988-3.

Gasca:1992:MPI

[GM92a] M. Gasca and G. Mühlbach. Multivariate polynomial interpolation under projectivities. II. Neville–Aitken formulas. Numerical Algorithms, 2(3–4):255–277, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gasca:1992:BHB

[GM92b] Mariano Gasca and José J. Martínez. Bivariate Hermite– Birkhoff interpolation and Vandermonde determinants. Numerical Algorithms, 3(1–4):193–199, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Graves-Morris:1996:NAA

[GM96] P. R. Graves-Morris. A new approach to acceleration of convergence of a sequence of vectors. Numerical Algorithms, 11 (1–4):189–201, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Graves-Morris:1997:LAL

[GM97] P. R. Graves-Morris. A “look-around Lanczos” algorithm for solving a system of linear equations. Numerical Algorithms, 15 (3–4):247–274, January 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Graves-Morris:2002:BB

[GM02] P. R. Graves-Morris. The breakdowns of BiCGStab. Numerical Algorithms, 29(1–3):97–105, March 2002. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp008.lwwonline.com/content/
getfile/5058/38/7/abstract.htm;   http://ipsapp008.
lwwonline.com/content/getfile/5058/38/7/fulltext.pdf.

Graves-Morris:2003:VSV

[GM03] P. R. Graves-Morris. VPAStab: Stabilised vector-Padé approximation with application to linear systems. Numerical Algorithms, 33(1–4):293–304, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
24/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/24/fulltext.pdf.

Ganesh:2006:FDH

[GM06] M. Ganesh and K. Mustapha. A fully discrete H^{1}-Galerkin method with quadrature for nonlinear advection-diffusion-reaction equations. Numerical Algorithms, 43(4):355–383, December 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
43&issue=4&spage=355.

Ganesh:2020:CHM

[GM20] M. Ganesh and C. Morgenstern. A coercive heterogeneous media Helmholtz model: formulation, wavenumber-explicit analysis, and preconditioned high-order FEM. Numerical Algorithms, 83(4):1441–1487, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gautschi:2021:OPR

[GM21] Walter Gautschi and Gradimir V. Milovanović. Orthogonal polynomials relative to a generalized Marchenko–Pastur probability measure. Numerical Algorithms, 88(3):1233–1249, November 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01073-1.

Gautschi:2022:OPR

[GM22] Walter Gautschi and Gradimir V. Milovanović. Orthogonal polynomials relative to weight functions of Prudnikov type. Numerical Algorithms, 90(1):263–270, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01187-6.

Gemignani:2023:RFP

[GM23] Luca Gemignani and Beatrice Meini. Relaxed fixed point iterations for matrix equations arising in Markov chain modeling. Numerical Algorithms, 94(1):149–173, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01496-y.

Gasca:1992:AST

[GMP92] M. Gasca, Charles A. Micchelli, and J. M. Peña. Almost strictly totally positive matrices. Numerical Algorithms, 2(2): 225–236, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Graves-Morris:1999:ABV

[GMS99] P. R. Graves-Morris and A. Salam. Avoiding breakdown in Van der Vorst’s method. Numerical Algorithms, 21(1– 4):205–223, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/7/fulltext.pdf.

Graves-Morris:1992:SIE

[GMT92] P. R. Graves-Morris and R. Thukral. Solution of integral equations using function-valued Padé approximants. II. Numerical Algorithms, 3(1–4):223–234, December 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Guo:2018:TEM

[GMY18] Qian Guo, Xuerong Mao, and Rongxian Yue. The truncated Euler–Maruyama method for stochastic differential delay equations. Numerical Algorithms, 78(2):599–624, June 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0391-0.pdf.

Gao:2019:TKH

[GMZ19] Xing Hui Gao, Le Rong Ma, and Hai Yun Zhou. Three kinds of hybrid algorithms and their numerical realizations for a finite family of quasi-asymptotically pseudocontractive mappings. Numerical Algorithms, 80(3):1015–1035, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gemignani:2012:MNM

[GN12] Luca Gemignani and Vanni Noferini. Modifications of Newton’s method for even-grade palindromic polynomials and other twined polynomials. Numerical Algorithms, 61(2): 315–329, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=61&issue=2&spage=315.

Goh:2010:SCC

[GNH10] S. M. Goh, M. S. M. Noorani, and I. Hashim. On solving the chaotic Chen system: a new time marching design for the variational iteration method using Adomian’s polynomial. Numerical Algorithms, 54(2):245–260, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=2&spage=245.

Gergelits:2022:NAS

[GNS22] Tomás Gergelits, Bjørn Fredrik Nielsen, and Zdenek Strakos. Numerical approximation of the spectrum of self-adjoint operators in operator preconditioning. Numerical Algorithms, 91 (1):301–325, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01263-5.

Greif:2024:IDC

[GNT24] Chen Greif, Sarah Nataj, and Manfred Trummer. Incomplete double-cone factorizations of centrosymmetric matrices arising in spectral methods. Numerical Algorithms, 95(3):1359–1386, March 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01612-y.

Gracia:2006:DCP

[GO06] J. L. Gracia and E. O’Riordan. A defect-correction parameter-uniform numerical method for a singularly perturbed convection-diffusion problem in one dimension. Numerical Algorithms, 41(4):359–385, April 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=4&spage=359.

Goncalves:2020:GCI

[GO20] M. L. N. Gonçalves and F. R. Oliveira. On the global convergence of an inexact quasi-Newton conditional gradient method for constrained nonlinear systems. Numerical Algorithms, 84 (2):609–631, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00772-0.

Gracia:2021:NAS

[GO21] J. L. Gracia and E. O’Riordan. Numerical approximations to a singularly perturbed convection–diffusion problem with a discontinuous initial condition. Numerical Algorithms, 88 (4):1851–1873, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01098-6.

Goda:2015:FCH

[God15] Takashi Goda. Fast construction of higher order digital nets for numerical integration in weighted Sobolev spaces. Numerical Algorithms, 69(2):357–396, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9902-4.

Godeme:2026:IBP

[God26] Jean-Jacques Godeme. Inertial Bregman proximal gradient under partial smoothness. Numerical Algorithms, 102(2):1135– 1183, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02173-y.

Gomez:1999:ASA

[Góm99] Manuel A. Gómez. An O(n^{2}) active set algorithm for the solution of a parametric quadratic program. Numerical Algorithms, 22(3–4):305–316, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/3/fulltext.pdf.

Gomez:2001:NAS

[Góm01] Manuel A. Gómez. An O(n^{2}) active set algorithm for solving two related box constrained parametric quadratic programs. Numerical Algorithms, 27(4):367–375, August 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/4/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
4/fulltext.pdf.

Ou:2014:MOB

[gOM14] Yi gui Ou and Wei Ma. A modified ODE-based algorithm for unconstrained optimization problems. Numerical Algorithms, 65(2):233–250, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9703-1.

Goncalves:2016:IGN

[Gon16] M. L. N. Gonçalves. Inexact Gauss–Newton like methods for injective-overdetermined systems of equations under a majorant condition. Numerical Algorithms, 72(2):377–392, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0050-2.

Gordon:2018:DAR

[Gor18] Dan Gordon. A derandomization approach to recovering bandlimited signals across a wide range of random sampling rates. Numerical Algorithms, 77(4):1141–1157, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Garcia-Palomares:1999:PPM

[GP99] Ubaldo M. García-Palomares. Preconditioning projection methods for solving algebraic linear systems. Numerical Algorithms, 21(1–4):157–164, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/6/fulltext.pdf.

Grau:2005:IMG

[GP05] Miquel Grau and Josep M. Peris. Iterative method generated by inverse interpolation with additional evaluations. Numerical Algorithms, 40(1):33–45, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=1&spage=33.

Guttel:2014:SOW

[GP14] Stefan Güttel and Jennifer Pestana. Some observations on weighted GMRES. Numerical Algorithms, 67(4):733–752, December 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9820-x.

Gonzalez-Parra:2014:NAO

[GPAA14] Gilberto González-Parra, Luis Acedo, and Abraham J. Arenas. A novel approach to obtain analytical-numerical solutions of nonlinear Lorenz system. Numerical Algorithms, 67(1):93– 107, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9776-x.

Garcia-Palomares:1998:IPA

[GPGC98] Ubaldo M. García-Palomares and Francisco J. González-Castaño. Incomplete projection algorithms for solving the convex feasibility problem. Numerical Algorithms, 18(2): 177–193, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
14/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/14/2/fulltext.pdf.

Gonzalez-Pinto:1992:EJS

[GPGVS92] S. González-Pinto, P. González-Vera, and J. C. Santos. On extrapolation of Jagerman and Stetter rules. Numerical Algorithms, 3(1–4):211–222, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Gonzalez-Pinto:2024:BCM

[GPHAPR24] S. González-Pinto, D. Hernández-Abreu, and S. Pérez-Rodríguez. Boundary corrections on multi-dimensional PDEs. Numerical Algorithms, 96(2):507–535, June 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01655-1.

Gonzalez-Pinto:2018:PWM

[GPHHAPR18] S. González-Pinto, E. Hairer, D. Hernández-Abreu, and S. Pérez-Rodríguez. PDE-W-methods for parabolic problems with mixed derivatives. Numerical Algorithms, 78(3):957–981, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gasparo:2001:NAP

[GPP01a] Maria Grazia Gasparo, Alessandra Papini, and Aldo Pasquali. Nonmonotone algorithms for pattern search methods. Numerical Algorithms, 28(1–4):171–186, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp007.kluweronline.com/content/getfile/
5058/37/11/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/37/11/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/37/
11/abstract.htm;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/11/fulltext.pdf.

Gori:2001:GMB

[GPP01b] Laura Gori, Francesca Pitolli, and Laura Pezza. On the Galerkin method based on a particular class of scaling functions. Numerical Algorithms, 28(1–4):187–198, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/37/19/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/37/
19/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/19/abstract.htm;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/19/
fulltext.pdf.

Gori:2001:PRO

[GPS01] Laura Gori, Francesca Pitolli, and Elisabetta Santi. Positive refinable operators. Numerical Algorithms, 28(1– 4):199–213, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/37/
12/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/37/12/fulltext.pdf;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/12/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/12/fulltext.pdf.

Guo:2009:SPL

[GQ09] Bai-Ni Guo and Feng Qi. A simple proof of logarithmic convexity of extended mean values. Numerical Algorithms, 52(1):89–92, September 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=1&spage=89.

Gutknecht:2001:HMC

[GR01] Martin H. Gutknecht and Miroslav Rozložník. By how much can residual minimization accelerate the convergence of orthogonal residual methods? Numerical Algorithms, 27(2):189– 213, June 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/35/6/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/35/
6/fulltext.pdf.

Grau:2003:ICN

[Gra03] Miquel Grau. An improvement to the computing of nonlinear equation solutions. Numerical Algorithms, 34(1):1–12, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.kluweronline.
com/content/getfile/5058/47/1/abstract.htm;   http:
//ipsapp008.kluweronline.com/content/getfile/5058/
47/1/fulltext.pdf.

Gil:2023:CCH

[GRAST23] Amparo Gil, Diego Ruiz-Antolín, Javier Segura, and Nico M. Temme. Computation of the confluent hypergeometric function U(a, b, x) and its derivative for positive arguments. Numerical Algorithms, 94(2):669–679, October 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01515-y.

Gresbrand:1996:RSP

[Gre96] Anke Gresbrand. Rational B-splines with prescribed poles. Numerical Algorithms, 12(1–2):151–158, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Grosse:1993:AVS

[Gro93] Eric Grosse. Approximation in VLSI simulation. Numerical Algorithms, 5(1–4):591–601, November 1993. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Gregoire:1997:DIS

[GRT97] J. P. Gregoire, C. Rose, and B. Thomas. Direct and iterative solvers for finite-element problems. Numerical Algorithms, 16(1):39–53, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
9/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/9/4/fulltext.pdf. Sparse ma-

trices in industry (Lille, 1997).

Golub:1994:EQF

[GS94] Gene H. Golub and Zdeněk Strakoš. Estimates in quadratic formulas. Numerical Algorithms, 8(2–4):241–268, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gori:1995:EHT

[GS95] Laura Gori and Elisabetta Santi. On the evaluation of Hilbert transforms by means of a particular class of Turán quadrature rules. Numerical Algorithms, 10(1–2):27–39, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Gergelits:2014:CCB

[GS14] Tomás Gergelits and Zdenek Strakos. Composite convergence bounds based on Chebyshev polynomials and finite precision conjugate gradient computations. Numerical Algorithms, 65 (4):759–782, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9713-z.

Goldman:2016:NPB

[GS16a] Ron Goldman and Plamen Simeonov. Novel polynomial Bernstein bases and Bézier curves based on a general notion of polynomial blossoming. Numerical Algorithms, 72(3):605–634, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0059-6.

Goncalves:2016:LAS

[GS16b] Douglas S. Gonçalves and Sandra A. Santos. Local analysis of a spectral correction for the Gauss–Newton model applied to quadratic residual problems. Numerical Algorithms, 73(2): 407–431, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0101-3.

Gander:2019:MCA

[GS19a] Walter Gander and Qiquan Shi. Matrix completion with e-algorithm. Numerical Algorithms, 80(1):279–301, January 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Goldman:2019:BAF

[GS19b] Ron Goldman and Plamen Simeonov. q-blossoming for analytic functions. Numerical Algorithms, 82(1):107–121, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Green:2019:GEI

[GS19c] Kevin R. Green and Raymond J. Spiteri. Gating-enhanced IMEX splitting methods for cardiac monodomain simulation. Numerical Algorithms, 81(4):1443–1457, August 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Grapiglia:2021:GWC

[GS21] Geovani N. Grapiglia and Ekkehard W. Sachs. A generalized worst-case complexity analysis for non-monotone line searches. Numerical Algorithms, 87(2):779–796, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00987-6.

Ganesan:2025:OSF

[GS25] Sashikumaar Ganesan and Maneesh Kumar Singh. Operatorsplitting finite element method for solving the radiative transfer equation. Numerical Algorithms, 98(4):1725–1753, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01850-8.

Guo:2026:LPT

[GS26] Xiaoyun Guo and Yueqiang Shang. Local and parallel twogrid discretization algorithms for natural convection flows.

Numerical Algorithms, 101(4):2413–2447, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02094-w.

Gugercin:2003:MLR

[GSA03] S. Gugercin, D. C. Sorensen, and A. C. Antoulas. A modified low-rank Smith method for large-scale Lyapunov equations. Numerical Algorithms, 32(1):27–55, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/44/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
44/3/fulltext.pdf.

Guzman:2026:MPA

[GSF26] Grover Enrique Castro Guzman, Peter Florian Stadler, and Andre Fujita. A message-passing approach to obtain the trace of matrix functions with applications to network analysis. Numerical Algorithms, 101(1):455–476, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02012-0.

Gil:2002:CCA

[GST02] Amparo Gil, Javier Segura, and Nico M. Temme. Computing complex Airy functions by numerical quadrature. Numerical Algorithms, 30(1):11–23, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
40/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/40/2/fulltext.pdf.

Gil:2003:CSF

[GST03] Amparo Gil, Javier Segura, and Nico M. Temme. Computing special functions by using quadrature rules. Numerical Algorithms, 33(1–4):265–275, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
22/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/22/fulltext.pdf.

Gil:2017:EAI

[GST17] A. Gil, J. Segura, and N. M. Temme. Efficient algorithms for the inversion of the cumulative central beta distribution. Numerical Algorithms, 74(1):77–91, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0139-2.

Gil:2021:FRH

[GST21] Amparo Gil, Javier Segura, and Nico M. Temme. Fast and reliable high-accuracy computation of Gauss–Jacobi quadrature. Numerical Algorithms, 87(4):1391–1419, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01012-6.

Grimme:1996:MRS

[GSV96] E. J. Grimme, D. C. Sorensen, and P. Van Dooren. Model reduction of state space systems via an implicitly restarted Lanczos method. Numerical Algorithms, 12(1–2):1–31, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gao:2022:MQR

[GSZ22] Qinjiao Gao, Xingping Sun, and Shenggang Zhang. Multivariate quadrature rules on crosslet sparse grids. Numerical Algorithms, 90(3):951–962, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01217-
3.

Gupta:2019:MKO

[GTA19] Vijay Gupta, Gancho Tachev, and Ana-Maria Acu. Modified Kantorovich operators with better approximation properties. Numerical Algorithms, 81(1):125–149, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gu:2020:CSC

[Gu20] Zhendong Gu. Chebyshev spectral collocation method for system of nonlinear Volterra integral equations. Numerical Algorithms, 83(1):243–263, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gugat:1996:NDC

[Gug96] M. Gugat. The Newton differential correction algorithm for rational Chebyshev approximation with constrained denominators. Numerical Algorithms, 13(1–2):107–122, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Guo:2013:MCN

[Guo13] Chun-Hua Guo. Monotone convergence of Newton-like methods for M-matrix algebraic Riccati equations. Numerical Algorithms, 64(2):295–309, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9666-7.

Guo:2016:MLS

[Guo16] Pei-Chang Guo. A modified large-scale structure-preserving doubling algorithm for a large-scale Riccati equation from transport theory. Numerical Algorithms, 71(3):541–552, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0008-4.

Gutknecht:2015:RSM

[Gut15] Martin H. Gutknecht. Revisiting (k, l)-step methods. Numerical Algorithms, 69(2):455–469, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9906-0.

Gallivan:1999:RAP

[GV99] Kyle Gallivan and Paul Van Dooren. Rational approximations of pre-filtered transfer functions via the Lanczos algorithm. Numerical Algorithms, 20(4):331–342, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/19/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
19/3/fulltext.pdf.

Golub:2000:PMS

[GV00] Gene H. Golub and Denis Vanderstraeten. On the preconditioning of matrices with skew-symmetric splittings. Numerical Algorithms, 25(1–4):223–239, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/3/19/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/19/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Guan:2024:GAI

[GW24] Jinrui Guan and Zhixin Wang. A generalized ALI iteration method for nonsymmetric algebraic Riccati equations. Numerical Algorithms, 97(1):119–134, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01695-7.

Gallois-Wong:2020:OIP

[GWBC20] Diane Gallois-Wong, Sylvie Boldo, and Pascal Cuoq. Optimal inverse projection of floating-point addition. Numerical Algorithms, 83(3):957–986, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gao:2018:NEB

[GWL18] Lei Gao, Yaqiang Wang, and Chaoqian Li. New error bounds for the linear complementarity problem of QN-matrices. Numerical Algorithms, 77(1):229–242, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gu:2020:RDG

[GWL20] Chuanye Gu, Zhiyou Wu, and Jueyou Li. Regularized dual gradient distributed method for constrained convex optimization over unbalanced directed graphs. Numerical Algorithms, 84(1):91–115, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Geng:2015:NCA

[GWW15] Juan Geng, Laisheng Wang, and Yanfei Wang. A non-convex algorithm framework based on DC programming and DCA for matrix completion. Numerical Algorithms, 68(4):903–921, April 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9876-2.

Gurioli:2026:HBV

[GWY26] Gianmarco Gurioli, Weijie Wang, and Xiaoqiang Yan. Hamiltonian Boundary Value Methods (HBVMs) for functional differential equations with piecewise continuous arguments. Numerical Algorithms, 101(1):1–31, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01994-7.

Geng:2019:CBI

[GX19] Hongrui Geng and Zhenhua Xu. Coupling of boundary integral equation and finite element methods for transmission problems in acoustics. Numerical Algorithms, 82(2):479–501, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gu:2011:NML

[GZ11] Chao Gu and Detong Zhu. A non-monotone line search multidimensional filter-SQP method for general nonlinear programming. Numerical Algorithms, 56(4):537–559, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=537.

Gao:2018:MTS

[GZ18] Wenwu Gao and Ran Zhang. Multiquadric trigonometric spline quasi-interpolation for numerical differentiation of noisy data: a stochastic perspective. Numerical Algorithms, 77(1):243–259, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Gao:2020:MRK

[GZ20] Wenwu Gao and Xuan Zhou. Multiscale radial kernels with high-order generalized Strang–Fix conditions. Numerical Algorithms, 85(2):427–448, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00820-
9.

Guo:2024:STS

[GZD24] Chenzheng Guo, Jing Zhao, and Qiao-Li Dong. A stochastic two-step inertial Bregman proximal alternating linearized minimization algorithm for nonconvex and nonsmooth problems. Numerical Algorithms, 97(1):51–100, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01693-9.

Gan:2024:PTA

[GZL24] Di Gan, Guo-Feng Zhang, and Zhao-Zheng Liang. Preconditioning techniques of all-at-once systems for multi-term timefractional diffusion equations. Numerical Algorithms, 96(4): 1499–1531, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01675-x.

Guo:2025:RTS

[GZL^{+}25] Wenxiu Guo, Hua Zheng, Xiaoping Lu, Yongxiong Zhang, and Seakweng Vong. A relaxation two-step parallel modulus method without auxiliary variable for solving large sparse vertical linear complementarity problems. Numerical Algorithms, 98(1):405–429, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01800-4.

Gu:2018:NIS

[GZP18] Chao Gu, Detong Zhu, and Yonggang Pei. A new inexact SQP algorithm for nonlinear systems of mixed equalities and inequalities. Numerical Algorithms, 78(4):1233–1253, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hesameddini:2016:NRA

[HA16] Esmail Hesameddini and Elham Asadollahifard. A new reliable algorithm based on the sinc function for the time fractional diffusion equation. Numerical Algorithms, 72(4):893–913, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0073-8.

Hagler:2013:ASG

[Hag13] Brian A. Hagler. Anchoring and steering Gaussian quadrature for positive definite strong moment functionals. Numerical Algorithms, 62(1):45–58, January 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9565-y/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=45-58.

Haghighi:2025:CPT

[Hag25] Amir Haghighi. Convergence of a partially truncated Euler– Maruyama method for SDEs with super-linear piecewise continuous drift and Hölder diffusion coefficients. Numerical Algorithms, 99(4):1649–1695, August 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01928-3.

Haijun:2008:NTO

[Hai08] Wang Haijun. On new third-order convergent iterative formulas. Numerical Algorithms, 48(4):317–325, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=4&
spage=317.

Haijun:2009:NTO

[Hai09] Wang Haijun. New third-order method for solving systems of nonlinear equations. Numerical Algorithms, 50(3):271–282, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
50&issue=3&spage=271.

Hajarian:2016:GCD

[Haj16a] Masoud Hajarian. Generalized conjugate direction algorithm for solving the general coupled matrix equations over symmetric matrices. Numerical Algorithms, 73(3):591–609, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0109-8.

Hajej:2016:EEA

[Haj16b] A. Hajej. Error estimates for approximation schemes of effective Hamiltonians arising in stochastic homogenization of Hamilton–Jacobi equations. Numerical Algorithms, 73(3):839– 868, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0120-0.

Halidias:2014:NAC

[Hal14] Nikolaos Halidias. A novel approach to construct numerical methods for stochastic differential equations. Numerical Algorithms, 66(1):79–87, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9724-9.

Handscomb:1993:LRS

[Han93] David Handscomb. Local recovery of a solenoidal vector field by an extension of the thin-plate spline technique. Numerical Algorithms, 5(1–4):121–129, November 1993. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Hansen:1994:RTM

[Han94] Per Christian Hansen. REGULARIZATION TOOLS: A Matlab package for analysis and solution of discrete ill-posed problems. Numerical Algorithms, 6(1–2):1–35, January 1994. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hanke:1996:AOP

[Han96] Martin Hanke. Asymptotics of orthogonal polynomials and the numerical solution of ill-posed problems. Numerical Algorithms, 11(1–4):203–214, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Hansen:1999:RTV

[Han99] Per Christian Hansen. Regularization Tools Version 3.0 for Matlab 5.2. Numerical Algorithms, 20(2–3):195–196, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/8/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/8/fulltext.pdf.

Hansen:2002:DRT

[Han02] Per Christian Hansen. Deconvolution and regularization with Toeplitz matrices. Numerical Algorithms, 29(4):323–378, April 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/39/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
39/2/fulltext.pdf.

Hansen:2007:RTV

[Han07] Per Christian Hansen. Regularization Tools version 4.0 for Matlab 7.3. Numerical Algorithms, 46(2):189–194, October 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=46&issue=2&
spage=189.

Hanke:2022:SIL

[Han22] Michael Hanke. On the sensitivity of implementations of a least-squares collocation method for linear higher-index differential-algebraic equations. Numerical Algorithms, 91(4): 1721–1754, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01320-z.

Hossain:2024:FIR

[HAN24] Alamgir Hossain, Md. Nasim Akhtar, and Maria A. Navascués. Fractal interpolation on the real projective plane. Numerical Algorithms, 96(2):557–582, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01657-
z.

Hari:2018:GCJ

[Har18] Vjeran Hari. Globally convergent Jacobi methods for positive definite matrix pairs. Numerical Algorithms, 79(1):221–249, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hari:2020:CFL

[Har20] Vjeran Hari. On the complex Falk–Langemeyer method. Numerical Algorithms, 83(2):451–483, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Horiuchi:2023:GGE

[HASI23] Itsuki Horiuchi, Kensuke Aihara, Toshio Suzuki, and Emiko Ishiwata. Global GPBiCGstab(L) method for solving linear matrix equations. Numerical Algorithms, 93(1):295–319, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01415-7.

Han:2013:NFS

[HBP13] Le Han, Shujun Bi, and Shaohua Pan. Nonsingularity of FB system and constraint nondegeneracy in semidefinite programming. Numerical Algorithms, 62(1):79–113, January 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9567-9/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=79-113.

Huckaby:2003:CSQ

[HC03] David A. Huckaby and Tony F. Chan. On the convergence of Stewart’s QLP algorithm for approximating the SVD. Numerical Algorithms, 32(2–4):287–316, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/45/10/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/45/
10/fulltext.pdf.

Huang:2024:MAR

[HC24] Zhengge Huang and Jingjing Cui. Modified and accelerated relaxed gradient-based iterative algorithms for the complex conjugate and transpose matrix equations. Numerical Algorithms, 97(4):1955–2009, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01775-2.

Herrera:2023:NIH

[HCBAEC23] César Herrera, Ricardo Corrales-Barquero, Jorge Arroyo-Esquivel, and Juan G. Calvo. A numerical implementation for the high-order 2D virtual element method in MATLAB. Numerical Algorithms, 92(3):1707–1721, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01361-4.

Huang:2026:NMM

[HCD26] Shenghua Huang, Gang Cai, and Qiao-Li Dong. A new modified Mann-type method for solving the monotone inclusion problem and the fixed point problem in reflexive Banach spaces. Numerical Algorithms, 102(1):247–266, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02129-2.

Hu:2018:DDS

[HCH18] Xiulin Hu, Yuhao Cong, and Guang-Da Hu. Delay-dependent stability of linear multistep methods for DAEs with multiple delays. Numerical Algorithms, 79(3):719–739, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hacat:2026:OCP

[HÇK26] Gülnur Haçat, Aytekin Çibik, and Songül Kaya. Optimal control problems of NS-α and NS-ω turbulence models: analysis and numerical tests. Numerical Algorithms, 101(1):477–507, January 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02013-z.

Hao:2021:NCF

[HCL21] Zhaopeng Hao, Wanrong Cao, and Shengyue Li. Numerical correction of finite difference solution for two-dimensional space-fractional diffusion equations with boundary singularity. Numerical Algorithms, 86(3):1071–1087, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00923-8.

Huang:2020:PEE

[HCXL20] Jian Huang, Zhongdi Cen, Aimin Xu, and Li-Bin Liu. A posteriori error estimation for a singularly perturbed Volterra integro-differential equation. Numerical Algorithms, 83(2): 549–563, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hook:2018:PAA

[HD18] James Hook and Nicholas Dingle. Performance analysis of asynchronous parallel Jacobi. Numerical Algorithms, 77(3): 831–866, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

He:2023:RKA

[HDL23] Songnian He, Qiao-Li Dong, and Xiaoxiao Li. The randomized Kaczmarz algorithm with the probability distribution depending on the angle. Numerical Algorithms, 93(1):415–440, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01422-8.

Hendy:2018:SLD

[HDP18] A. S. Hendy, R. H. De Staelen, and V. G. Pimenov. A semilinear delayed diffusion-wave system with distributed order in time. Numerical Algorithms, 77(3):885–903, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Heyouni:2005:MKS

[HE05] M. Heyouni and A. Essai. Matrix Krylov subspace methods for linear systems with multiple right-hand sides. Numerical Algorithms, 40(2):137–156, October 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=137.

He:2016:USS

[He16] Dongdong He. An unconditionally stable spatial sixth-order CCD–ADI method for the two-dimensional linear telegraph equation. Numerical Algorithms, 72(4):1103–1117, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0082-7.

Heard:2010:SMS

[Hea10] Allison Heard. Scale and modify for the second and third order BDF methods. Numerical Algorithms, 53(2–3):261–280, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=261.

Heinrichs:2006:LSS

[Hei06] Wilhelm Heinrichs. Least-squares spectral collocation with the overlapping Schwarz method for the incompressible Navier– Stokes equations. Numerical Algorithms, 43(1):61–73, September 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=43&
issue=1&spage=61.

Heinrichs:2007:ASL

[Hei07] Wilhelm Heinrichs. An adaptive spectral least-squares scheme for the Burgers equation. Numerical Algorithms, 44(1):1–10, January 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
44&issue=1&spage=1.

Hemmingsson:1994:FMS

[Hem94] Lina Hemmingsson. A fast modified sine transform for solving block-tridiagonal systems with Toeplitz blocks. Numerical Algorithms, 7(2–4):375–389, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hempel:1996:IRR

[Hem96] Daniel Hempel. Isotropic refinement and recoarsening in two dimensions. Numerical Algorithms, 13(1–2):33–43, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hamdi:2004:ESE

[HEOS04] S. Hamdi, W. H. Enright, Y. Ouellet, and W. E. Schiesser. Exact solutions of extended Boussinesq equations. Numerical Algorithms, 37(1–4):165–175, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/12/abstract.htm.

Hermann:1996:SPT

[Her96] Thomas Hermann. On the stability of polynomial transformations between Taylor, Bernstein and Hermite forms. Numerical Algorithms, 13(3–4):307–320, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Heyouni:1999:NGH

[Hey99] M. Heyouni. Newton generalized Hessenberg method for solving nonlinear systems of equations. Numerical Algorithms, 21 (1–4):225–246, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
24/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/24/fulltext.pdf.

Heyouni:2001:GHC

[Hey01] M. Heyouni. The global Hessenberg and CMRH methods for linear systems with multiple right-hand sides. Numerical Algorithms, 26(4):317–332, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
32/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/32/2/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/32/
2/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/32/2/fulltext.pdf.

Hon:2024:STB

[HFDSC24] Sean Hon, Po Yin Fung, Jiamei Dong, and Stefano Serra-Capizzano. A sine transform based preconditioned MINRES method for all-at-once systems from constant and variablecoefficient evolutionary PDEs. Numerical Algorithms, 95(4): 1769–1799, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01627-5.

Huang:2021:MNI

[HFW^{+}21] Haoen Huang, Dongyang Fu, Guancheng Wang, Long Jin, Shan Liao, and Huan Wang. Modified Newton integration algorithm with noise suppression for online dynamic nonlinear optimization. Numerical Algorithms, 87(2):575–599, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00979-6.

Hu:2019:NSM

[HFZ19] Hanzhang Hu, Yiping Fu, and Jie Zhou. Numerical solution of a miscible displacement problem with dispersion term using a two-grid mixed finite element approach. Numerical Algorithms, 81(3):879–914, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hansen:1993:FOD

[HG93] Per Christian Hansen and Henrik Gesmar. Fast orthogonal decomposition of rank deficient Toeplitz matrices. Numerical Algorithms, 4(1–2):151–166, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hussain:2025:LMM

[HG25] Manzoor Hussain and Abdul Ghafoor. The localized meshless method of lines for the approximation of two-dimensional reaction–diffusion system. Numerical Algorithms, 98(3): 1507–1551, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01842-8.

Hayek:1992:RAN

[HGVPA92] N. Hayek, P. González-Vera, and F. Pérez-Acosta. Rational approximation to Neumann series of Bessel functions. Numerical Algorithms, 3(1–4):235–244, December 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Gao:2017:TDS

[hGzS17] Guang hua Gao and Zhi zhong Sun. Two difference schemes for solving the one-dimensional time distributed-order fractional wave equations. Numerical Algorithms, 74(3):675–697, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0167-y.

Hargreaves:2005:EAM

[HH05] Gareth I. Hargreaves and Nicholas J. Higham. Efficient algorithms for the matrix cosine and sine. Numerical Algorithms, 40(4):383–400, December 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=4&spage=383.

Haghighi:2011:SSS

[HH11] Amir Haghighi and S. Mohammad Hosseini. On the stability of some second order numerical methods for weak approximation of Itô SDEs. Numerical Algorithms, 57(1):101–124, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=1&
spage=101.

Haghighi:2012:CSS

[HH12] Amir Haghighi and S. Mohammad Hosseini. A class of split-step balanced methods for stiff stochastic differential equations. Numerical Algorithms, 61(1):141–162, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=1&
spage=141.

He:2022:IAP

[HHF22] Xin He, Rong Hu, and Ya-Ping Fang. Inertial accelerated primal-dual methods for linear equality constrained convex optimization problems. Numerical Algorithms, 90(4):1669–1690, August 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01246-y.

Hasegawa:2007:EDA

[HHHN07] Takemitsu Hasegawa, Susumu Hibino, Yohsuke Hosoda, and Ichizo Ninomiya. An extended doubly-adaptive quadrature method based on the combination of the Ninomiya and the FLR schemes. Numerical Algorithms, 45(1–4):101–112, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=45&issue=1&
spage=101.

Hermoso:2023:HOR

[HHLM23] Carlos Hermoso, Edmundo J. Huertas, Alberto Lastra, and Francisco Marcellán. Higher-order recurrence relations, Sobolev-type inner products and matrix factorizations. Numerical Algorithms, 92(1):665–692, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01402-y.

Hasannasab:2021:AEA

[HHLS21a] Marzieh Hasannasab, Johannes Hertrich, Friederike Laus, and Gabriele Steidl. Alternatives to the EM algorithm for ML estimation of location, scatter matrix, and degree of freedom of the Student t distribution. Numerical Algorithms, 87(1): 77–118, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00959-w. See correction

[HHLS21b].

Hasannasab:2021:CAE

[HHLS21b] Marzieh Hasannasab, Johannes Hertrich, Friederike Laus, and Gabriele Steidl. Correction to: Alternatives to the EM algorithm for ML estimation of location, scatter matrix, and degree of freedom of the Student t distribution. Numerical Algorithms, 88(1):521–522, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01156-z;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01156-z.pdf. See

[HHLS21a].

Haddad:2019:PCI

[HHST19] Mireille Haddad, Frédéric Hecht, Toni Sayah, and Pierre Henri Tournier. Parallel computing investigations for the projection method applied to the interface transport scheme of a twophase flow by the method of characteristics. Numerical Algorithms, 80(2):447–467, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hieu:2018:CAN

[Hie18] Dang Van Hieu. Convergence analysis of a new algorithm for strongly pseudomontone equilibrium problems. Numerical Algorithms, 77(4):983–1001, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hieu:2019:NIA

[Hie19] Dang Van Hieu. New inertial algorithm for a class of equilibrium problems. Numerical Algorithms, 80(4):1413–1436, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Higham:1997:SIM

[Hig97] Nicholas J. Higham. Stable iterations for the matrix square root. Numerical Algorithms, 15(2):227–242, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/7/6/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/7/
6/fulltext.pdf.

Hu:2017:RCS

[HIK17] Xianfeng Hu, Mark Iwen, and Hyejin Kim. Rapidly computing sparse Legendre expansions via sparse Fourier transforms. Numerical Algorithms, 74(4):1029–1059, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0184-x.

Hill:2010:MAS

[Hil10] A. T. Hill. G-matrices for algebraically stable general linear methods. Numerical Algorithms, 53(2–3):281–292, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=281.

Haselby:2026:FLM

[HIS26] Cullen Haselby, Mark A. Iwen, and William Swartworth. Fast and low-memory compressive sensing algorithms for low Tucker-rank tensor approximation from streamed measurements. Numerical Algorithms, 101(4):2567–2629, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02101-0.

Hached:2018:NSL

[HJ18a] M. Hached and K. Jbilou. Numerical solutions to large-scale differential Lyapunov matrix equations. Numerical Algorithms, 79(3):741–757, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hansen:2018:ATI

[HJ18b] Per Christian Hansen and Jakob Sauer Jørgensen. AIR Tools II: algebraic iterative reconstruction methods, improved implementation. Numerical Algorithms, 79(1):107–137, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2021:CFC

[HJ21] Jinzhi Huang and Zhongxiao Jia. On choices of formulations of computing the generalized singular value decomposition of a large matrix pair. Numerical Algorithms, 87(2):689–718, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00984-9.

Huang:2025:RRH

[HJ25] Jinzhi Huang and Zhongxiao Jia. Refined and refined harmonic Jacobi–Davidson methods for computing several GSVD components of a large regular matrix pair. Numerical Algorithms, 99(2):895–920, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01901-0.

Hajipour:2018:ADH

[HJB18] Mojtaba Hajipour, Amin Jajarmi, and Dumitru Baleanu. On the accurate discretization of a highly nonlinear boundary value problem. Numerical Algorithms, 79(3):679–695, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hofstatter:2006:ADC

[HK06] Harald Hofstätter and Othmar Koch. Analysis of a defect correction method for geometric integrators. Numerical Algorithms, 41(2):103–126, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=2&spage=103.

Hofstatter:2014:AES

[HK14] Harald Hofstätter and Othmar Koch. An approximate eigensolver for self-consistent field calculations. Numerical Algorithms, 66(3):609–641, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9751-6.

Hu:2024:FDL

[HKCW24] Dongdong Hu, Linghua Kong, Wenjun Cai, and Yushun Wang. Fully decoupled, linear, and energy-preserving GSAV difference schemes for the nonlocal coupled sine-Gordon equations in multiple dimensions. Numerical Algorithms, 95(4):1953– 1980, April 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01634-6.

He:1997:CSF

[HKE97] Qiming He, Lishan Kang, and D. J. Evans. Convergence and stability of the finite difference scheme for nonlinear parabolic systems with time delay. Numerical Algorithms, 16(2):129–153, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/1/fulltext.pdf.

Hickernell:2012:WCI

[HKKN12] Fred J. Hickernell, Peter Kritzer, Frances Y. Kuo, and Dirk Nuyens. Weighted compound integration rules with higher order convergence for all N. Numerical Algorithms, 59(2):161–183, February 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=2&spage=161.

He:2025:RMC

[HKP25] Haoze He, Daniel Kressner, and Bor Plestenjak. Randomized methods for computing joint eigenvalues, with applications to multiparameter eigenvalue problems and root finding. Numerical Algorithms, 100(3):861–892, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01971-0.

Hinrichs:2019:TDL

[HKPW19] Aicke Hinrichs, Peter Kritzer, Friedrich Pillichshammer, and G. W. Wasilkowski. Truncation dimension for linear problems on multivariate function spaces. Numerical Algorithms, 80(2): 661–685, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hales:2002:EEM

[HL02] S. J. Hales and J. Levesley. Error estimates for multilevel approximation using polyharmonic splines. Numerical Algorithms, 30(1):1–10, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
40/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/40/1/fulltext.pdf.

Hanke:2003:CIN

[HL03] Michael Hanke and René Lamour. Consistent initialization for nonlinear index-2 differential–algebraic equation: Large sparse systems in MATLAB. Numerical Algorithms, 32(1):67–85, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/44/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
44/5/fulltext.pdf.

Hang:2006:ASG

[HL06] Fengbo Hang and Youming Li. On applicability of the sparse grid method in the worst case setting. Numerical Algorithms, 42(2):95–105, June 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=2&spage=95.

Hegland:2015:SGQ

[HL15a] Markus Hegland and Paul Leopardi. Sparse grid quadrature on products of spheres. Numerical Algorithms, 70(3):485– 517, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9958-9.

Huang:2015:BBT

[HL15b] Yakui Huang and Hongwei Liu. A Barzilai–Borwein type method for minimizing composite functions. Numerical Algorithms, 69(4):819–838, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9927-8.

Huang:2017:FNM

[HL17] Yun-Chi Huang and Siu-Long Lei. A fast numerical method for block lower triangular Toeplitz with dense Toeplitz blocks system with applications to time-space fractional diffusion equations. Numerical Algorithms, 76(3):605–616, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2020:FSF

[HL20] Yun-Chi Huang and Siu-Long Lei. Fast solvers for finite difference scheme of two-dimensional time-space fractional differential equations. Numerical Algorithms, 84(1):37–62, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hou:2023:AOS

[HL23a] Zhisong Hou and Sanyang Liu. An accelerating outer space algorithm for globally solving generalized linear multiplicative problems. Numerical Algorithms, 94(2):877–904, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01523-y.

Hou:2023:GAC

[HL23b] Zhisong Hou and Sanyang Liu. Global algorithm for a class of multiplicative programs using piecewise linear approximation technique. Numerical Algorithms, 92(2):1063–1082, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01330-x.

Huang:2015:NSQ

[HLC15] Yakui Huang, Hongwei Liu, and Weijie Cong. A note on the smoothing quadratic regularization method for non-Lipschitz optimization. Numerical Algorithms, 69(4):863–874, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9929-6.

Hu:2025:SPJ

[HLJ25] Zhe-Han Hu, Si-Tao Ling, and Zhi-Gang Jia. Structurepreserving joint Lanczos bidiagonalization with thick-restart for the partial quaternion GSVD. Numerical Algorithms, 99 (2):873–894, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01900-1.

Han:2022:FSC

[HLL22] Ningning Han, Jian Lu, and Shidong Li. The finite steps of convergence of the fast thresholding algorithms with f-feedbacks in compressed sensing. Numerical Algorithms, 90(3):1197– 1223, July 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01227-1.

Hou:2004:CFR

[HLM04] Qing-Hu Hou, Alain Lascoux, and Yan-Ping Mu. Continued fractions for Rogers–Szegő polynomials. Numerical Algorithms, 35(1):81–90, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/49/A/5/
abstract.htm.

Hadri:2025:RNS

[HLN25] Aissam Hadri, Amine Laghrib, and Mourad Nachaoui. On the resolution of the non-smooth inverse Cauchy problem by the primal-dual method. Numerical Algorithms, 100(1):??, September 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01950-5.

Ha:2010:PTD

[HLS10] Yonghoon Ha, Keunwha Lee, and Woojae Seong. Pseudospectral time-domain method for pressure-release rough surface scattering using a surface transformation and image method. Numerical Algorithms, 54(3):411–430, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=411.

Huertas:2025:ZBH

[HLSLSL25] Edmundo J. Huertas, Alberto Lastra, Anier Soria-Lorente, and Víctor Soto-Larrosa. On zero behavior of higherorder Sobolev-type discrete q-Hermite I orthogonal polynomials. Numerical Algorithms, 99(1):1–25, May 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01868-y.

Hu:2016:INM

[HLTA16] Xiuling Hu, F. Liu, I. Turner, and V. Anh. An implicit numerical method of a new time distributed-order and two-sided space-fractional advection-dispersion equation. Numerical Algorithms, 72(2):393–407, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0051-1.

Hu:2026:STO

[HLV26] Ting Hu, Dongdong Liu, and Seak-Weng Vong. Solving thirdorder tensor linear complementarity problems via modulusbased methods and applications. Numerical Algorithms, 101 (2):993–1021, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02030-y.

He:2026:AQN

[HLW26] Yi He, Shi-Hao Li, and Yuan-Zi Wang. Applications of quasideterminants to non-commutative multistep ε-algorithm. Numerical Algorithms, 101(2):1373–1394, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02045-5.

Huang:2014:BBT

[HLZ14] Yakui Huang, Hongwei Liu, and Sha Zhou. A Barzilai– Borwein type method for stochastic linear complementarity problems. Numerical Algorithms, 67(3):477–489, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9803-y.

Hubbert:2006:ICB

[HM06] Simon Hubbert and Stefan Müller. Interpolation with circular basis functions. Numerical Algorithms, 42(1):75–90, May 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=42&issue=1&
spage=75.

Hubbert:2007:TPS

[HM07] Simon Hubbert and Stefan Müller. Thin plate spline interpolation on the unit interval. Numerical Algorithms, 45(1–4):167– 177, August 2007. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-007-9103-5.

Hueso:2014:SCF

[HM14] José L. Hueso and Eulalia Martínez. Semilocal convergence of a family of iterative methods in Banach spaces. Numerical Algorithms, 67(2):365–384, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9795-7.

Huang:2018:SIM

[HM18a] Bao-Hua Huang and Chang-Feng Ma. Some iterative methods for the largest positive definite solution to a class of nonlinear matrix equation. Numerical Algorithms, 79(1):153–178, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2018:IAL

[HM18b] Baohua Huang and Changfeng Ma. An iterative algorithm for the least Frobenius norm Hermitian and generalized skew Hamiltonian solutions of the generalized coupled Sylvesterconjugate matrix equations. Numerical Algorithms, 78(4): 1271–1301, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2018:ESP

[HM18c] Na Huang and Chang-Feng Ma. On the eigenvalues of the saddle point matrices discretized from Navier–Stokes equations. Numerical Algorithms, 79(1):41–64, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2019:FIA

[HM19a] Baohua Huang and Changfeng Ma. Finite iterative algorithm for the symmetric periodic least squares solutions of a class of periodic Sylvester matrix equations. Numerical Algorithms, 81 (1):377–406, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2019:SAP

[HM19b] Na Huang and Chang-Feng Ma. Spectral analysis of the preconditioned system for the 3 × 3 block saddle point problem. Numerical Algorithms, 81(2):421–444, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hanke:2022:TRIa

[HM22a] Michael Hanke and Roswitha März. Towards a reliable implementation of least-squares collocation for higher index differential-algebraic equations — Part 1: basics and ansatz function choices. Numerical Algorithms, 89(3):931–963, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01140-7;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01140-7.pdf.

Hanke:2022:TRIb

[HM22b] Michael Hanke and Roswitha März. Towards a reliable implementation of least-squares collocation for higher index differential-algebraic equations — Part 2: the discrete leastsquares problem. Numerical Algorithms, 89(3):965–986, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01141-6;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01141-6.pdf.

Higham:2022:AEM

[HM22c] Nicholas J. Higham and Mantas Mikaitis. Anymatrix: an extensible MATLAB matrix collection. Numerical Algorithms, 90(3):1175–1196, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01226-2.

Hanke:2026:CAI

[HM26] Michael Hanke and Roswitha März. On the computation of accurate initial conditions for linear higher-index differentialalgebraic equations and its application in initial value solvers. Numerical Algorithms, 101(4):2849–2909, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02116-7.

Hieu:2016:PHE

[HMA16] Dang Van Hieu, Le Dung Muu, and Pham Ky Anh. Parallel hybrid extragradient methods for pseudomonotone equilibrium problems and nonexpansive mappings. Numerical Algorithms, 73(1):197–217, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0092-5.

Hernandez-Mederos:2008:IUB

[HMdÁES08] Victoria Hernández-Mederos, Pedro L. del Ángel, and Jorge Estrada-Sarlabous. Isotropic umbrella based triangulation of regular parametric surfaces. Numerical Algorithms, 48(1–3):29–47, July 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=29.

Hietel:1996:CFD

[HMS96] Dietmar Hietel, Andreas Meister, and Thomas Sonar. On the comparison of four different implementations of a third-order ENO scheme of box type for the computation of compressible flow. Numerical Algorithms, 13(1–2):77–105, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hayotov:2011:OQF

[HMS11] Abdullo Rakhmonovich Hayotov, Gradimir V. Milovanović, and Kholmat Mahkambaevich Shadimetov. On an optimal quadrature formula in the sense of Sard. Numerical Algorithms, 57(4):487–510, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=4&spage=487.

Hammouda:2017:MHS

[HMT17] Chiheb Ben Hammouda, Alvaro Moraes, and Raúl Tempone. Multilevel hybrid split-step implicit tau-leap. Numerical Algorithms, 74(2):527–560, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0158-z.

Hanke:1994:TAI

[HN94] Martin Hanke and James G. Nagy. Toeplitz approximate inverse preconditioner for banded Toeplitz matrices. Numerical Algorithms, 7(2–4):183–199, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Higham:2016:ACP

[HN16] Nicholas J. Higham and Vanni Noferini. An algorithm to compute the polar decomposition of a 3 × 3 matrix. Numerical Algorithms, 73(2):349–369, October 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-016-0098-7.pdf.

Hatano:2009:NEG

[HNSH09] Yasuyo Hatano, Ichizo Ninomiya, Hiroshi Sugiura, and Takemitsu Hasegawa. Numerical evaluation of Goursat’s infinite integral. Numerical Algorithms, 52(2):213–224, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=2&
spage=213.

Huai:2018:GIU

[HNY^{+}18] Kaizhan Huai, Mingfang Ni, Zhanke Yu, Xiang Zheng, and Feng Ma. A generalized inexact Uzawa method for stable principal component pursuit problem with nonnegative constraints. Numerical Algorithms, 77(3):653–674, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hoang:2025:HON

[Hoa25] Manh Tuan Hoang. High-order nonstandard finite difference methods preserving dynamical properties of one-dimensional dynamical systems. Numerical Algorithms, 98(1):219–249, January 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01792-1.

Hoffmann:2005:NCK

[Hof05] Miklós Hoffmann. Numerical control of Kohonen neural network for scattered data approximation. Numerical Algorithms, 39(1–3):175–186, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hoffmann:2016:HGA

[Hof16] Philipp H. W. Hoffmann. A Hitchhiker’s guide to automatic differentiation. Numerical Algorithms, 72(3):775–811, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0067-6.

Hofreither:2021:ABR

[Hof21] Clemens Hofreither. An algorithm for best rational approximation based on barycentric rational interpolation. Numerical Algorithms, 88(1):365–388, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01042-
0.

Holstad:1998:MSC

[Hol98] Astrid Holstad. Modelling subsurface coupled chemical reactions and fluid flow over long time periods. Numerical Algorithms, 19(1–4):95–110, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
16/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/7/fulltext.pdf. Differen-

tial algebraic equations (Grenoble, 1997).

Homeier:1992:LTA

[Hom92] Herbert H. H. Homeier. A Levin-type algorithm for accelerating the convergence of Fourier series. Numerical Algorithms, 3(1–4):245–254, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Homeier:1994:HCI

[Hom94] Herbert H. H. Homeier. A hierarchically consistent, iterative sequence transformation. Numerical Algorithms, 8(1):47–81, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Homeier:1998:AHC

[Hom98a] Herbert H. H. Homeier. An asymptotically hierarchyconsistent, iterative sequence transformation for convergence acceleration of Fourier series. Numerical Algorithms, 18 (1):1–30, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
13/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/13/2/fulltext.pdf.

Homeier:1998:ST

[Hom98b] Herbert H. H. Homeier. On the stability of the J transformation. Numerical Algorithms, 17(3–4):223–239, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/12/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
12/4/fulltext.pdf.

Healy:1995:RMA

[HOW95] D. Healy, T. Olsen, and J. Weaver. Reduced motion artifacts in medical imaging by adaptive spatio-temporal reconstruction. Numerical Algorithms, 9(1–2):55–84, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hartmann:2018:PBA

[HP18a] Rainer Hartmann and Christoph Pflaum. A prewavelet-based algorithm for the solution of second-order elliptic differential equations with variable coefficients on sparse grids. Numerical Algorithms, 78(3):929–956, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

He:2018:USL

[HP18b] Dongdong He and Kejia Pan. An unconditionally stable linearized difference scheme for the fractional Ginzburg–Landau equation. Numerical Algorithms, 79(3):899–925, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Heuveline:1997:PCS

[HPS97] Vincent Heuveline, Bernard Philippe, and Miloud Sadkane. Parallel computation of spectral portrait of large matrices by Davidson type methods. Numerical Algorithms, 16(1):55–75, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
9/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/9/5/fulltext.pdf. Sparse ma-

trices in industry (Lille, 1997).

Hiptmair:2013:MPE

[HPS13] R. Hiptmair, G. Phillips, and G. Sinha. Multiple point evaluation on combined tensor product supports. Numerical Algorithms, 63(2):317–337, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9624-4.

Hrncir:2020:DSP

[HPS20] Jakub Hrncír, Ivana Pultarová, and Zdenek Strakos. Decomposition into subspaces preconditioning: abstract framework. Numerical Algorithms, 83(1):57–98, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Peng:2009:GBS

[hPwL09] Zhuo hua Peng and Jin wang Liu. The generalized bisymmetric solutions of the matrix equation A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{ℓ}X_{ℓ}B_{ℓ} = C and its optimal approximation. Numerical Algorithms, 50(2):127–144, February 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=127.

Higueras:2000:SAS

[HR00] Immaculada Higueras and Teo Roldán. Starting algorithms for some DIRK methods. Numerical Algorithms, 23 (4):357–369, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
26/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/26/1/fulltext.pdf.

Heinig:2003:FAC

[HR03a] Georg Heinig and Karla Rost. Fast algorithms for centrosymmetric and centro-skewsymmetric Toeplitz-plus-Hankel matrices. Numerical Algorithms, 33(1–4):305–317, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/25/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
25/fulltext.pdf.

Hill:2003:CSA

[HR03b] Michael Hill and Ian Robinson. A comparison of strategies for the automatic computation of two-dimensional integrals over infinite domains. Numerical Algorithms, 34 (2–4):325–338, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/2/fulltext.pdf.

Hernandez-Ramos:2005:AOP

[HR05] Luis Manuel Hernández-Ramos. Alternating oblique projections for coupled linear systems. Numerical Algorithms, 38 (4):285–303, April 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hernandez:2007:EIO

[HR07] M. A. Hernández and N. Romero. On the efficiency index of one-point iterative processes. Numerical Algorithms, 46(1):35–44, September 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=1&spage=35.

Hearn:2014:ADM

[HR14] Tristan A. Hearn and Lothar Reichel. Application of denoising methods to regularization of ill-posed problems. Numerical Algorithms, 66(4):761–777, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9760-5.

Hieu:2022:NPL

[HRAH22] Dang Van Hieu, Simeon Reich, Pham Ky Anh, and Nguyen Hai Ha. A new proximal-like algorithm for solving split variational inclusion problems. Numerical Algorithms, 89(2):811– 837, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01135-4.

Huang:2016:CSS

[HRY16] Guangxin Huang, Lothar Reichel, and Feng Yin. On the choice of solution subspace for nonstationary iterated Tikhonov regularization. Numerical Algorithms, 72(4):1043–1063, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0079-2.

Huang:2019:CSL

[HRY19] Guangxin Huang, Lothar Reichel, and Feng Yin. On the choice of subspace for large-scale Tikhonov regularization problems in general form. Numerical Algorithms, 81(1):33–55, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hasegawa:1996:AIP

[HS96] Takemitsu Hasegawa and Avram Sidi. An automatic integration procedure for infinite range integrals involving oscillatory kernels. Numerical Algorithms, 13(1–2):1–19, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Higham:2003:CMC

[HS03] Nicholas J. Higham and Matthew I. Smith. Computing the matrix cosine. Numerical Algorithms, 34(1):13–26, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.kluweronline.
com/content/getfile/5058/47/2/abstract.htm;   http:
//ipsapp008.kluweronline.com/content/getfile/5058/
47/2/fulltext.pdf.

Huckle:2012:TPR

[HS12] Thomas Kilian Huckle and Matous Sedlacek. Tikhonov– Phillips regularization with operator dependent seminorms. Numerical Algorithms, 60(2):339–353, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=2&spage=339.

Hosseini:2015:NIM

[HS15] S. Mohammed Hosseini and Samira Smaeili. Numerical integration of multi-dimensional highly oscillatory integrals, based on eRPIM. Numerical Algorithms, 68(2):423–442, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9851-y.

Higham:2016:AAA

[HS16] Nicholas J. Higham and Nataša Strabić. Anderson acceleration of the alternating projections method for computing the nearest correlation matrix. Numerical Algorithms, 72(4):1021– 1042, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-015-0078-3.pdf.

Hanjing:2020:HIA

[HS20] Adisak Hanjing and Suthep Suantai. Hybrid inertial accelerated algorithms for split fixed point problems of demicontractive mappings and equilibrium problems. Numerical Algorithms, 85(3):1051–1073, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00855-
y.

Huang:2021:REA

[HS21] Chaobao Huang and Martin Stynes. α-robust error analysis of a mixed finite element method for a time-fractional biharmonic equation. Numerical Algorithms, 87(4):1749–1766, August 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01036-y.

Han:2026:RML

[HS26] Sujia Han and Caiqin Song. H-representation method for the least squares symmetric solutions of the reduced biquaternion matrix equation and its application in color image restoration. Numerical Algorithms, 101(3):1537–1566, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02054-4.

Hezari:2016:NIM

[HSE16] Davod Hezari, Davod Khojasteh Salkuyeh, and Vahid Edalatpour. A new iterative method for solving a class of complex symmetric system of linear equations. Numerical Algorithms, 73(4):927–955, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0123-x.

Hoseini:2020:PHS

[HSK20] Mohsen Hoseini, Shahram Saeidi, and Do Sang Kim. On perturbed hybrid steepest descent method with minimization or superiorization for subdifferentiable functions. Numerical Algorithms, 85(1):353–374, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00818-
3.

Huertas:2019:NAP

[HSL19] Edmundo J. Huertas and Anier Soria-Lorente. New analytic properties of nonstandard Sobolev-type Charlier orthogonal polynomials. Numerical Algorithms, 82(1):41–68, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Humes:2004:SIH

[HSS04] Carlos Humes, Jr., Paulo J. S. Silva, and Benar F. Svaiter. Some inexact hybrid proximal augmented Lagrangian algorithms. Numerical Algorithms, 35(2–4):175–184, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/12/abstract.htm.

Huard:2013:RQD

[HSSB13] Mathieu Huard, Nathalie Sprynski, Nicolas Szafran, and Luc Biard. Reconstruction of quasi developable surfaces from ribbon curves. Numerical Algorithms, 63(3):483–506, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9633-3.

Hosseini:2015:MBM

[HST15] S. A. Hosseini, S. Shahmorad, and F. Talati. A matrix based method for two dimensional nonlinear Volterra–Fredholm integral equations. Numerical Algorithms, 68(3):511–529, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9858-4.

Hetmaniok:2014:UHA

[HSTW14] Edyta Hetmaniok, Damian Slota, Tomasz Trawiński, and Roman Witula. Usage of the homotopy analysis method for solving the nonlinear and linear integral equations of the second kind. Numerical Algorithms, 67(1):163–185, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-013-9781-0.pdf.

Huang:2023:ETK

[HSY23] Xing Huang, Yongqiang Suo, and Chenggui Yuan. Estimate of transition kernel for Euler–Maruyama scheme for SDEs driven by α-stable noise and applications. Numerical Algorithms, 94(3):1381–1402, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01539-4.

Hon:2003:AGA

[HSZ03] Y. C. Hon, R. Schaback, and X. Zhou. An adaptive greedy algorithm for solving large RBF collocation problems. Numerical Algorithms, 32(1):13–25, January 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
44/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/44/2/fulltext.pdf.

Hadjidimos:2016:SLC

[HT16] Apostolos Hadjidimos and Michael Tzoumas. The solution of the linear complementarity problem by the matrix analogue of the accelerated overrelaxation iterative method. Numerical Algorithms, 73(3):665–684, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0112-0.

He:2019:SPM

[HT19] Songnian He and Hanlin Tian. Selective projection methods for solving a class of variational inequalities. Numerical Algorithms, 80(2):617–634, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hadjidimos:2021:ETP

[HT21] Apostolos Hadjidimos and Michael Tzoumas. On equivalence of three-parameter iterative methods for singular symmetric saddle-point problem. Numerical Algorithms, 86(4):1391–1419, April 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00938-1.

Halder:2023:NSN

[HT23] Joydev Halder and Suman Kumar Tumuluri. Numerical solution to a nonlinear McKendrick–Von Foerster equation with diffusion. Numerical Algorithms, 92(2):1007–1039, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01328-5.

Huang:2013:TFD

[HTVY13] Jianfei Huang, Yifa Tang, Luis Vázquez, and Jiye Yang. Two finite difference schemes for time fractional diffusion-wave equation. Numerical Algorithms, 64(4):707–720, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9689-0.

Hu:2022:TGM

[Hu22] Hanzhang Hu. Two-grid method for compressible miscible displacement problem by mixed finite element methods and expanded mixed finite element method of characteristics. Numerical Algorithms, 89(2):611–636, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01127-4.

Huang:1994:SCG

[Hua94] Zhi Jian Huang. The semilocal convergence of a generalization of Brent’s and Brown’s methods. Numerical Algorithms, 6(1– 2):37–62, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:1996:HSP

[Hua96] Wanzhen Huang. Heuristic solutions to polynomial moment problems with some convex entropic objectives. Numerical Algorithms, 12(3–4):297–308, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2018:SGQ

[Hua18] Rong Huang. On the sensitivity of generators for the QR factorization of quasiseparable matrices with total nonpositivity. Numerical Algorithms, 77(3):905–924, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2020:VPU

[Hua20] Na Huang. Variable parameter Uzawa method for solving a class of block three-by-three saddle point problems. Numerical Algorithms, 85(4):1233–1254, December 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-019-00863-y.

Huang:2021:NSM

[Hua21] Baohua Huang. Numerical study on Moore–Penrose inverse of tensors via Einstein product. Numerical Algorithms, 87(4): 1767–1797, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01074-0.

Huckle:1992:NSP

[Huc92] Thomas Huckle. A note on skewcirculant preconditioners for elliptic problems. Numerical Algorithms, 2(3–4):279–286, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hunter:1995:ELF

[Hun95] D. B. Hunter. The evaluation of Legendre functions of the second kind. Numerical Algorithms, 10(1–2):41–49, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Huber:1998:PEG

[HV98] Birkett Huber and Jan Verschelde. Polyhedral end games for polynomial continuation. Numerical Algorithms, 18 (1):91–108, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
13/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/13/1/fulltext.pdf.

Humet:2012:WUT

[HV12] Matthias Humet and Marc Van Barel. When is the Uvarov transformation positive definite? Numerical Algorithms, 59(1):51–62, January 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=1&spage=51.

Haussermann:2015:ELE

[HV15] John Haussermann and Robert A. Van Gorder. Efficient lowerror analytical-numerical approximations for radial solutions of nonlinear Laplace equations. Numerical Algorithms, 70(2): 227–248, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9944-7.

Hendy:2022:STD

[HV22] A. S. Hendy and K. Van Bockstal. A solely time-dependent source reconstruction in a multiterm time-fractional order diffusion equation with non-smooth solutions. Numerical Algorithms, 90(2):809–832, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01210-
w.

Hogervorst:1993:DPP

[HvD93] Bas J. Hogervorst and Ruud van Damme. Degenerate polynomial patches of degree 11 for almost GC^{2} interpolation over triangles. Numerical Algorithms, 5(1–4):557–568, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Hernandez-Veron:2015:SCT

[HVM15] M. A. Hernández-Verón and Eulalia Martínez. On the semilocal convergence of a three steps Newton-type iterative process under mild convergence conditions. Numerical Algorithms, 70(2):377–392, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9952-7.

Hernandez-Veron:2024:SND

[HVMMV24] M. A. Hernández-Verón, Á. A. Magreñán, Eulalia Martínez, and Eva G. Villalba. Solving non-differentiable Hammerstein integral equations via first-order divided differences. Numerical Algorithms, 97(2):567–594, October 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01715-6.

Hernandez-Veron:2017:SCK

[HVMT17] M. A. Hernández-Verón, Eulalia Martínez, and Carles Teruel. Semilocal convergence of a k-step iterative process and its application for solving a special kind of conservative problems. Numerical Algorithms, 76(2):309–331, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Hernandez-Veron:2023:KTM

[HVYMS23] M. A. Hernández-Verón, Nisha Yadav, Eulalia Martínez, and Sukhjit Singh. Kurchatov-type methods for non-differentiable Hammerstein-type integral equations. Numerical Algorithms, 93(1):131–155, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01406-8.

Han:2000:NOA

[HW00] Lei Han and G. W. Wasilkowski. A new optimal algorithm for weighted approximation and integration over R. Numerical Algorithms, 23(4):393–406, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
26/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/26/2/fulltext.pdf.

Hon:2018:CPA

[HW18] Sean Hon and Andrew Wathen. Circulant preconditioners for analytic functions of Toeplitz matrices. Numerical Algorithms, 79(4):1211–1230, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2015:NNL

[HWC15] Shuai Huang, Zhong Wan, and Xiaohong Chen. A new nonmonotone line search technique for unconstrained optimization. Numerical Algorithms, 68(4):671–689, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9866-4.

Hu:2025:PEE

[HWC25] Xianfa Hu, Wansheng Wang, and Jiliang Cao. A posteriori error estimates for the exponential midpoint method for linear and semilinear parabolic equations. Numerical Algorithms, 99 (4):1967–1982, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01940-7.

He:2019:OPS

[HWCR19] Songnian He, Tao Wu, Yeol Je Cho, and Themistocles M. Rassias. Optimal parameter selections for a general Halpern iteration. Numerical Algorithms, 82(4):1171–1188, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

He:2025:IRK

[HWD25] Songnian He, Ziting Wang, and Qiao-Li Dong. Inertial randomized Kaczmarz algorithms for solving coherent linear systems. Numerical Algorithms, 99(1):133–163, May 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01872-2.

Huang:2017:GVD

[HWXC17] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. A generalized variant of the deteriorated PSS preconditioner for nonsymmetric saddle point problems. Numerical Algorithms, 75(4):1161–1191, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2018:MGS

[HWXC18] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. A modified generalized shift-splitting preconditioner for nonsymmetric saddle point problems. Numerical Algorithms, 78(1):297–331, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2019:MPI

[HWXC19] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Modified PHSS iterative methods for solving nonsingular and singular saddle point problems. Numerical Algorithms, 80 (2):485–519, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2024:ULT

[HWZ24] Yi Huang, Wansheng Wang, and Yanming Zhang. Unconditional long-time stability-preserving second-order BDF fully discrete method for fractional Ginzburg–Landau equation. Numerical Algorithms, 97(1):167–189, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01699-3.

Huang:2021:DDS

[HY21] Xing Huang and Fen-Fen Yang. Distribution-dependent SDEs with Hölder continuous drift and α-stable noise. Numerical Algorithms, 86(2):813–831, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00913-
w.

Huang:2020:EMN

[HYJ20] Jianfei Huang, Dandan Yang, and Laurent O. Jay. Efficient methods for nonlinear time fractional diffusion-wave equations and their fast implementations. Numerical Algorithms, 85(2): 375–397, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00817-4.

Hou:2026:MTR

[HYT26] Wen-Qiong Hou, Fan Yang, and Zhen-Ji Tian. The modified Tikhonov regularization method for backward heat conduction problem with a complete parabolic equation in R^{n}. Numerical Algorithms, 102(2):981–1010, June 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02164-z.

Hao:2020:HCM

[HYW20] Yue Hao, Ai-Li Yang, and Yu-Jiang Wu. How to compute the minimum norm least squares solution of singular linear system by using the preconditioned HSS method? Numerical Algorithms, 83(3):1205–1221, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Helfrich:1993:TRM

[HZ93] H.-P. Helfrich and D. Zwick. A trust region method for implicit orthogonal distance regression. Numerical Algorithms, 5(1–4): 535–545, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Helfrich:1995:TRA

[HZ95] H.-P. Helfrich and D. Zwick. Trust region algorithms for the nonlinear least distance problem. Numerical Algorithms, 9(1– 2):171–179, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Huang:2015:MCF

[HZ15] Zhengda Huang and Xiaoyan Zhou. On the minimum convergence factor of a class of GSOR-like methods for augmented systems. Numerical Algorithms, 70(1):113–132, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9937-6.

Hu:2020:IAT

[HZ20] Xindi Hu and Shengfeng Zhu. Isogeometric analysis for timefractional partial differential equations. Numerical Algorithms, 85(3):909–930, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00844-1.

Huang:2024:TRE

[HZ24] Guang-Xin Huang and Shuang-You Zhong. Tensor randomized extended Kaczmarz methods for large inconsistent tensor linear equations with t-product. Numerical Algorithms, 96(4): 1755–1778, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01684-w.

Hang:2023:CCD

[HZPW23] Tongtong Hang, Zhongguo Zhou, Hao Pan, and Yan Wang. The conservative characteristic difference method and analysis for solving two-sided space-fractional advection–diffusion equations. Numerical Algorithms, 92(3):1723–1755, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01363-2.

Huo:2020:MAB

[HZX20] Guanze Huo, Yongkui Zou, and Yingxiang Xu. B-method approach to blow-up solutions of fourth-order semilinear parabolic equations. Numerical Algorithms, 85(4):1365–1384, December 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00868-7.

Hou:2021:HES

[HZX21] Dianming Hou, Hongyi Zhu, and Chuanju Xu. Highly efficient schemes for time-fractional Allen–Cahn equation using extended SAV approach. Numerical Algorithms, 88(3):1077– 1108, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01068-y.

Ibrahim:2024:TCS

[IAAH24] Abdulkarim Hassan Ibrahim, Mohammed Alshahrani, and Suliman Al-Homidan. Two classes of spectral three-term derivative-free method for solving nonlinear equations with application. Numerical Algorithms, 96(4):1625–1645, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01679-7.

Iqbal:2020:EAF

[IAH20] Hira Iqbal, Mujahid Abbas, and S. M. Husnine. Existence and approximation of fixed points of multivalued generalized α-nonexpansive mappings in Banach spaces. Numerical Algorithms, 85(3):1029–1049, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00854-
z.

Ibragimov:2025:GAM

[Ibr25] Alimzhan Ibragimov. Generalized analysis of methods for solving systems of nonlinear equations with point and interval coefficients. Numerical Algorithms, 99(2):627–649, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01891-z.

Ilie:2006:NSI

[ICR06] Silvana Ilie, Robert M. Corless, and Greg Reid. Numerical solutions of index-1 differential algebraic equations can be computed in polynomial time. Numerical Algorithms, 41(2):161–171, February 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=2&spage=161.

Ishteva:2009:DGN

[IDAV09] Mariya Ishteva, Lieven De Lathauwer, P.-A. Absil, and Sabine Van Huffel. Differential-geometric Newton method for the best rank-(R_{1}, R_{2}, R_{3}) approximation of tensors. Numerical Algorithms, 51(2):179–194, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=2&spage=179.

Tributes to Gene H. Golub, Part II.

Imakura:2016:EBR

[IDS16] Akira Imakura, Lei Du, and Tetsuya Sakurai. Error bounds of Rayleigh–Ritz type contour integral-based eigensolver for solving generalized eigenvalue problems. Numerical Algorithms, 71(1):103–120, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-9987-4.

Ihsberner:2007:REA

[Ihs07] Katja Ihsberner. Roundoff error analysis of fast DCT algorithms in fixed point arithmetic. Numerical Algorithms, 46(1):1–22, September 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=1&spage=1.

Iiduka:2024:TAA

[Iid24] Hideaki Iiduka. Theoretical analysis of Adam using hyperparameters close to one without Lipschitz smoothness. Numerical Algorithms, 95(1):383–421, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01575-0.

Izzo:2019:TIE

[IJ19] G. Izzo and Z. Jackiewicz. Transformed implicit-explicit DIM-SIMs with strong stability preserving explicit part. Numerical Algorithms, 81(4):1343–1359, August 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ilie:2015:ATS

[IJE15] Silvana Ilie, Kenneth R. Jackson, and Wayne H. Enright. Adaptive time-stepping for the strong numerical solution of stochastic differential equations. Numerical Algorithms, 68(4): 791–812, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9872-6.

Iusem:2016:NSQ

[IJSS16] Alfredo N. Iusem, Joaquim J. Júdice, Valentina Sessa, and Hanif D. Sherali. On the numerical solution of the quadratic eigenvalue complementarity problem. Numerical Algorithms, 72(3):721–747, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0064-9.

Izgin:2024:NGL

[IKMS24] Thomas Izgin, Stefan Kopecz, Andreas Meister, and Amandine Schilling. On the non-global linear stability and spurious fixed points of MPRK schemes with negative RK parameters. Numerical Algorithms, 96(3):1221–1242, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01770-7.

Iske:2005:MSD

[IL05] Armin Iske and Jeremy Levesley. Multilevel scattered data approximation by adaptive domain decomposition. Numerical Algorithms, 39(1–3):187–198, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ilati:2020:AAI

[Ila20] Mohammad Ilati. Analysis and application of the interpolating element-free Galerkin method for extended Fisher– Kolmogorov equation which arises in brain tumor dynamics modeling. Numerical Algorithms, 85(2):485–502, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00823-6.

Iavernaro:2002:GBD

[IM02] Felice Iavernaro and Francesca Mazzia. Generalization of backward differentiation formulas for parallel computers. Numerical Algorithms, 31(1–4):139–155, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/24/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/43/
24/fulltext.pdf.

Iavernaro:2002:EQE

[IMT02] Felice Iavernaro, Francesca Mazzia, and Donato Trigiante. Eigenvalues and quasi-eigenvalues of banded Toeplitz matrices: Some properties and applications. Numerical Algorithms, 31(1–4):157–170, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
25/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/25/fulltext.pdf.

Imakura:2023:CMB

[IMT23] Akira Imakura, Keiichi Morikuni, and Akitoshi Takayasu. Complex moment-based methods for differential eigenvalue problems. Numerical Algorithms, 92(1):693–721, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01456-y.

Iserles:1995:ERB

[IN95] Arieh Iserles and Syvert P. Nørsett. Explicit representations of biorthogonal polynomials. Numerical Algorithms, 10(1–2):

51–67, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Ivan:2021:RIP

[IN21] Mircea Ivan and Vicuta Neagos. A representation of the interpolation polynomial. Numerical Algorithms, 88(3):1215– 1231, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01072-2.

Isoni:2001:SCN

[INR01] Tiziana Isoni, Pierpaolo Natalini, and Paolo E. Ricci. Symbolic computation of Newton sum rules for the zeros of polynomial eigenfunctions of linear differential operators. Numerical Algorithms, 28(1–4):215–227, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/37/13/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/37/
13/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/13/abstract.htm;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/13/
fulltext.pdf.

Ivanov:2016:HES

[IP16] Kamen G. Ivanov and Pencho P. Petrushev. Highly effective stable evaluation of bandlimited functions on the sphere. Numerical Algorithms, 71(3):585–611, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0011-9.

Inoan:2013:RAB

[IR13] Daniela Inoan and Ioan Rasa. A recursive algorithm for Bernstein operators of second kind. Numerical Algorithms, 64(4): 699–706, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9688-1.

Imakura:2017:BKT

[IS17] Akira Imakura and Tetsuya Sakurai. Block Krylov-type complex moment-based eigensolvers for solving generalized eigenvalue problems. Numerical Algorithms, 75(2):413–433, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Iiduka:2022:RSF

[IS22] Hideaki Iiduka and Hiroyuki Sakai. Riemannian stochastic fixed point optimization algorithm. Numerical Algorithms, 90 (4):1493–1517, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01238-y.

Iserles:1996:BPR

[Ise96] Arieh Iserles. Biorthogonal polynomials: Recent developments. Numerical Algorithms, 11(1–4):215–228, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Ishizaka:2026:NMU

[Ish26] Hiroki Ishizaka. Nitsche’s method under a semi-regular mesh condition. Numerical Algorithms, 102(1):291–334, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02131-8.

Istace:1993:CBC

[IT93] M.-P. Istace and J.-P. Thiran. On computing best Chebyshev complex rational approximants. Numerical Algorithms, 5(1–4): 299–308, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Idiz:2024:NMB

[ITA24] Fatih Idiz, Gamze Tanoglu, and Nasser Aghazadeh. A numerical method based on Legendre wavelet and quasilinearization technique for fractional Lane–Emden type equations. Numerical Algorithms, 95(1):181–206, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01568-z.

Izuchukwu:2019:PTA

[IUM^{+}19] C. Izuchukwu, G. C. Ugwunnadi, O. T. Mewomo, A. R. Khan, and M. Abbas. Proximal-type algorithms for split minimization problem in P -uniformly convex metric spaces. Numerical Algorithms, 82(3):909–935, November 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

intHout:2012:SCF

[iV12] Karel J. in ’t Hout and Kim Volders. Stability of central finite difference schemes for the Heston PDE. Numerical Algorithms, 60(1):115–133, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=1&spage=115.

Ivan:2015:NHI

[Iva15] Mircea Ivan. A note on the Hermite interpolation. Numerical Algorithms, 69(3):517–522, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9909-x.

Ivanov:2017:USC

[Iva17] Stoil I. Ivanov. A unified semilocal convergence analysis of a family of iterative algorithms for computing all zeros of a polynomial simultaneously. Numerical Algorithms, 75(4): 1193–1204, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ivanov:2024:FHO

[Iva24] Stoil I. Ivanov. Families of high-order simultaneous methods with several corrections. Numerical Algorithms, 97(2):945–958, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01734-3.

Idriss:2004:VII

[IW04] Ismail I. Idriss and Wolfgang V. Walter. Validated infeasible interior-point predictor–corrector methods for linear programming. Numerical Algorithms, 37(1–4):177–185, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/13/abstract.htm.

Ixaru:2019:ETF

[Ixa19] L. Gr. Ixaru. Exponential and trigonometrical fittings: userfriendly expressions for the coefficients. Numerical Algorithms, 82(3):1085–1096, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ixaru:2021:NCC

[Ixa21] L. Gr. Ixaru. Numerical computation of the coefficients in exponential fitting. Numerical Algorithms, 87(3):1097–1106, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01000-w.

Ibrahim:2015:HSC

[IY15] Iman H. Ibrahim and Fatma M. Yousry. Hybrid special class for solving differential-algebraic equations. Numerical Algorithms, 69(2):301–320, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9897-x.

Ikram:2024:CLI

[IZ24] Boukhelkhal Ikram and Rebiha Zeghdane. Correction to: Lagrange interpolation polynomials for solving nonlinear stochastic integral equations. Numerical Algorithms, 96(2):619, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01713-8. See [BZ24].

Jolaoso:2022:ESE

[JA22] Lateef Olakunle Jolaoso and Maggie Aphane. An explicit subgradient extragradient algorithm with self-adaptive stepsize for pseudomonotone equilibrium problems in Banach spaces. Numerical Algorithms, 89(2):583–610, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01126-5.

Jaiswal:2016:SCE

[Jai16] J. P. Jaiswal. Semilocal convergence of an eighth-order method in Banach spaces and its computational efficiency. Numerical Algorithms, 71(4):933–951, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0031-5.

See erratum [Jai17].

Jaiswal:2017:ESC

[Jai17] J. P. Jaiswal. Erratum to: Semilocal convergence of an eighthorder method in Banach spaces and its computational efficiency. Numerical Algorithms, 74(2):639–641, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0266-4.pdf. See [Jai16].

Janka:2003:FCM

[Jan03] Aleš Janka. A flux correction multigrid for compressible flow. Numerical Algorithms, 33(1–4):319–330, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/26/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
26/fulltext.pdf.

Jator:2015:ITD

[Jat15] S. N. Jator. Implicit third derivative Runge–Kutta–Nyström method with trigonometric coefficients. Numerical Algorithms, 70(1):133–150, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9938-5.

Jawecki:2022:SDB

[Jaw22] Tobias Jawecki. A study of defect-based error estimates for the Krylov approximation of ϕ-functions. Numerical Algorithms, 90(1):323–361, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01190-x.

Jay:2002:ISS

[Jay02] Laurent O. Jay. Iterative solution of SPARK methods applied to DAEs. Numerical Algorithms, 31(1–4):171–191, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/43/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
43/1/fulltext.pdf.

Jay:2021:SCS

[Jay21] Laurent O. Jay. Symplecticness conditions of some low order partitioned methods for non-autonomous Hamiltonian systems. Numerical Algorithms, 86(2):495–514, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00898-6.

Jarry-Bolduc:2022:ADH

[JB22] Gabriel Jarry-Bolduc. Approximating the diagonal of a Hessian: which sample set of points should be used. Numerical Algorithms, 91(3):1349–1361, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01304-z.

Jiang:2024:MMT

[JB24] Yingjun Jiang and Huiran Bai. Multigrid methods for time fractional conservation laws. Numerical Algorithms, 97(3): 1301–1322, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01750-x.

Jbilou:1993:GPA

[Jbi93] Khalide Jbilou. A general projection algorithm for solving systems of linear equations. Numerical Algorithms, 4(4):361–377, May 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jbilou:2003:BKS

[Jbi03] K. Jbilou. Block Krylov subspace methods for large algebraic Riccati equations. Numerical Algorithms, 34(2– 4):339–353, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/3/fulltext.pdf.

Jani:2017:BOM

[JBJB17] M. Jani, E. Babolian, S. Javadi, and D. Bhatta. Banded operational matrices for Bernstein polynomials and application to the fractional advection-dispersion equation. Numerical Algorithms, 75(4):1041–1063, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jezequel:2004:CII

[JC04] F. Jézéquel and J.-M. Chesneaux. Computation of an infinite integral using Romberg’s method. Numerical Algorithms, 36(3):265–283, July 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/57/A/5/
abstract.htm.

Janna:2015:EGP

[JCF15] Carlo Janna, Nicola Castelletto, and Massimiliano Ferronato. The effect of graph partitioning techniques on parallel Block FSAI preconditioning: a computational study. Numerical Algorithms, 68(4):813–836, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9873-5.

Jiang:2023:IAF

[JCH23] Fan Jiang, Xingju Cai, and Deren Han. Inexact asymmetric forward–backward-adjoint splitting algorithms for saddle point problems. Numerical Algorithms, 94(1):479–509, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01509-w.

Ji:2016:NAI

[JCL16] Haifeng Ji, Jinru Chen, and Zhilin Li. A new augmented immersed finite element method without using SVD interpolations. Numerical Algorithms, 71(2):395–416, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9999-0.

Jin:2025:DSG

[JFSD25] Zheng-Fen Jin, Yibao Fan, Youlin Shang, and Weiwei Ding. A dual symmetric Gauss–Seidel technique-based proximal ADMM for robust fused lasso estimation. Numerical Algorithms, 98(3):1337–1360, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01836-

Jang:2025:RFG

[JGMC25] Yongseok Jang, Laura Grigori, Emeric Martin, and Cédric Content. Randomized flexible GMRES with deflated restarting. Numerical Algorithms, 98(1):431–465, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01801-3.

Jo:2022:EMA

[JH22] Gwanghyun Jo and Youn Doh Ha. Effective multigrid algorithms for algebraic system arising from static peridynamic systems. Numerical Algorithms, 89(2):885–904, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01138-1.

Jiang:2026:GQM

[JH26] Yunjing Jiang and Guo He. Generalized quadrature method for Volterra integral equation with highly oscillatory Besseltrigonometric kernels. Numerical Algorithms, 101(2):741–760, February 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02020-0.

Jiang:2015:ADM

[JHLL15] Le Jiang, Jin Huang, Xiao-Guang Lv, and Jun Liu. Alternating direction method for the high-order total variation-based Poisson noise removal problem. Numerical Algorithms, 69(3): 495–516, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9908-y.

Hou:2006:IML

[jHyPlZ06] Jin jun Hou, Zhen yun Peng, and Xiang lin Zhang. An iterative method for the least squares symmetric solution of matrix equation AXB = CAXB = C. Numerical Algorithms, 42(2):181–192, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=2&spage=181.

Jia:2006:UCP

[Jia06] Zhongxiao Jia. Using cross-product matrices to compute the SVD. Numerical Algorithms, 42(1):31–61, May 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=1&spage=31.

Jia:2020:BFA

[Jia20a] Ji-Teng Jia. A breakdown-free algorithm for computing the determinants of periodic tridiagonal matrices. Numerical Algorithms, 83(1):149–163, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jia:2020:RPK

[Jia20b] Zhongxiao Jia. Regularization properties of Krylov iterative solvers CGME and LSMR for linear discrete illposed problems with an application to truncated randomized SVDs. Numerical Algorithms, 85(4):1281–1310, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00865-w.

Jia:2013:SAS

[JJ13] Ji-Teng Jia and Yao-Lin Jiang. Symbolic algorithm for solving cyclic penta-diagonal linear systems. Numerical Algorithms, 63 (2):357–367, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9626-2.

Jena:2024:IHT

[JJ24] Hrushikesh Jena and Mahendra Kumar Jena. An introduction to a hybrid trigonometric box spline surface producing subdivision scheme. Numerical Algorithms, 95(1):73–116, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01565-2.

Jia:2024:IUE

[JJC24] Junqing Jia, Xiaoyun Jiang, and Xiaoqing Chi. Improved uniform error bounds of Lawson-type exponential integrator method for long-time dynamics of the high-dimensional space fractional sine-Gordon equation. Numerical Algorithms, 97(3):

1179–1214, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01745-0.

Johnson:1997:TDI

[JJK97] Mark E. Johnson, Michael S. Jolly, and Ioannis G. Kevrekidis. Two-dimensional invariant manifolds and global bifurcations: some approximation and visualization studies. Numerical Algorithms, 14(1–3):125–140, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/7/fulltext.pdf. Dynamical

numerical analysis (Atlanta, GA, 1995).

Jaklic:2018:PPC

[JK18] Gasper Jaklic and Jernej Kozak. On parametric polynomial circle approximation. Numerical Algorithms, 77(2):433–450, February 2018. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-017-0322-0.

Jo:2019:GMA

[JK19] Gwanghyun Jo and Do Y. Kwak. Geometric multigrid algorithms for elliptic interface problems using structured grids. Numerical Algorithms, 81(1):211–235, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jaiswal:2026:ECW

[JKC26] Aishwarya Jaiswal, Sunil Kumar, and Carmelo Clavero. Efficient component-wise splitting approach to solve coupled singularly perturbed parabolic reaction-diffusion systems with interior layers. Numerical Algorithms, 102(1):335–363, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02133-6.

Jaklic:2007:TPL

[JKK^{+}07] Gasper Jaklic, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Zagar. Three-pencil lattices on triangulations. Numerical Algorithms, 45(1–4):49–60, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=49.

Jaklic:2008:BCL

[JKK^{+}08] Gasper Jaklic, Jernej Kozak, Marjeta Krajnc, Vito Vitrih, and Emil Zagar. Barycentric coordinates for Lagrange interpolation over lattices on a simplex. Numerical Algorithms, 48(1–3):93–104, July 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=93.

Jarlebring:2018:DNQ

[JKM18] E. Jarlebring, A. Koskela, and G. Mele. Disguised and new quasi-Newton methods for nonlinear eigenvalue problems. Numerical Algorithms, 79(1):311–335, September 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0438-2.pdf.

Jangveladze:2013:FEA

[JKNR13] Temur Jangveladze, Zurab Kiguradze, Beny Neta, and Simeon Reich. Finite element approximations of a nonlinear diffusion model with memory. Numerical Algorithms, 64(1):127– 155, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9658-7.

Jator:2012:ASO

[JL12] Samuel N. Jator and Jiang Li. An algorithm for second order initial and boundary value problems with an automatic error estimate based on a third derivative method. Numerical Algorithms, 59(3):333–346, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=333.

Jia:2015:PEE

[JL15] Zhongxiao Jia and Hui Lv. A posteriori error estimates of Krylov subspace approximations to matrix functions. Numerical Algorithms, 69(1):1–28, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9878-0.

Jia:2016:NAD

[JL16] Ji-Teng Jia and Su-Mei Li. Numerical algorithm for the determinant evaluation of cyclic pentadiagonal matrices with Toeplitz structure. Numerical Algorithms, 71(2):337–348, February 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9995-4.

Jia:2021:JBP

[JL21] Zhongxiao Jia and Haibo Li. The joint bidiagonalization process with partial reorthogonalization. Numerical Algorithms, 88(2):965–992, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01064-8.

Jacumin:2026:AFD

[JL26] Thomas Jacumin and Andreas Langer. An adaptive finite difference method for total variation minimization. Numerical Algorithms, 101(2):1337–1372, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02044-6.

Jia:2019:NFS

[JLFL19] Hongen Jia, Jichun Li, Zhiwei Fang, and Ming Li. A new FDTD scheme for Maxwell’s equations in Kerr-type nonlinear media. Numerical Algorithms, 82(1):223–243, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jiang:2022:NFH

[JLJ22] Xianzhen Jiang, Wei Liao, and Jinbao Jian. A new family of hybrid three-term conjugate gradient methods with applications in image restoration. Numerical Algorithms, 91(1): 161–191, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01258-2.

Jeannerod:2016:SEB

[JLMP16] Claude-Pierre Jeannerod, Nicolas Louvet, Jean-Michel Muller, and Antoine Plet. Sharp error bounds for complex floatingpoint inversion. Numerical Algorithms, 73(3):735–760, November 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0115-x.

Jentzen:2020:EMB

[JLP20] Arnulf Jentzen, Felix Lindner, and Primoz Pusnik. Exponential moment bounds and strong convergence rates for tamed-truncated numerical approximations of stochastic convolutions. Numerical Algorithms, 85(4):1447–1473, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00871-y.

Jiang:2025:UPP

[JLWZ25] Yingsong Jiang, Ruishu Liu, Xiaojie Wang, and Jinghua Zhuo. Unconditionally positivity-preserving approximations of the Aït–Sahalia type model: Explicit Milstein-type schemes. Numerical Algorithms, 98(4):1985–2003, April 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01861-5.

Jiang:2022:NAI

[JLX22] Xiaoying Jiang, Xiaowen Li, and Xiang Xu. Numerical algorithms for inverse Sturm–Liouville problems. Numerical Algorithms, 89(3):1287–1309, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01153-
2.

Liu:2017:SMU

[jLyLqW17] Fu jun Liu, Hui yuan Li, and Zhong qing Wang. Spectral methods using generalized Laguerre functions for second and fourth order problems. Numerical Algorithms, 75(4):1005–1040, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jia:2023:HMS

[JLZZ23] Zhigang Jia, Xuan Liu, Jingfei Zhu, and Meixiang Zhao. Harmonic multi-symplectic Lanczos algorithm for quaternion singular triplets. Numerical Algorithms, 93(3):1309–1335, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01469-7.

Joly:1993:CCG

[JM93] Pascal Joly and Gérard Meurant. Complex conjugate gradient methods. Numerical Algorithms, 4(4):379–406, May 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Joly:2000:WPS

[JM00] Pascal Joly and Roland Masson. Wavelet preconditioning of the Stokes problem in ψ-ω formulation. Numerical Algorithms, 24(4):357–369, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
29/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/29/4/fulltext.pdf. Sparse

matrices and differential equations (Rennes, 1999).

Jiang:2018:WRP

[JM18a] Yao-Lin Jiang and Zhen Miao. Waveform relaxation of partial differential equations. Numerical Algorithms, 79(4):1087–1106, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jouymandi:2018:RAS

[JM18b] Zeynab Jouymandi and Fridoun Moradlou. Retraction algorithms for solving variational inequalities, pseudomonotone equilibrium problems, and fixed-point problems in Banach spaces. Numerical Algorithms, 78(4):1153–1182, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jing:2025:EOS

[JML25] Xia Jing, Xiaohua Ma, and Xia Liu. An efficient outcome-space branch-and-bound algorithm for solving a class of large-scale linear multiplicative programs. Numerical Algorithms, 100(1): 277–309, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01961-2.

Jaros:2016:AGI

[JMS16] Patrycja Jaros, Łukasz Maślanka, and Filip Strobin. Algorithms generating images of attractors of generalized iterated function systems. Numerical Algorithms, 73(2):477–499, October 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0104-0.

Jourhmane:1999:AMI

[JN99] M. Jourhmane and A. Nachaoui. An alternating method for an inverse Cauchy problem. Numerical Algorithms, 21 (1–4):247–260, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
26/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/26/fulltext.pdf.

Jia:2019:LML

[JNS19] Zhigang Jia, Michael K. Ng, and Guang-Jing Song. Lanczos method for large-scale quaternion singular value decomposition. Numerical Algorithms, 82(2):699–717, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jones:1992:ASP

[JNW92] William B. Jones, Olav Njåstad, and Haakon Waadeland. Asymptotics for Szegő polynomial zeros. Numerical Algorithms, 3(1–4):255–264, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Johansson:2015:RHP

[Joh15] Fredrik Johansson. Rigorous high-precision computation of the Hurwitz zeta function and its derivatives. Numerical Algorithms, 69(2):253–270, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9893-1.

Johansson:2020:CLW

[Joh20] Fredrik Johansson. Computing the Lambert W function in arbitrary-precision complex interval arithmetic. Numerical Algorithms, 83(1):221–242, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jonasson:1993:PCG

[Jón93] Kristján Jónasson. A projected conjugate gradient method for sparse minimax problems. Numerical Algorithms, 5(1–4):309– 323, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Joshi:2022:MPV

[Jos22] Bhuwan Chandra Joshi. Mathematical programs with vanishing constraints involving strongly invex functions. Numerical Algorithms, 91(2):505–530, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01271-5.

Jebelean:2014:NSS

[JP14] Petru Jebelean and Constantin Popa. Numerical solutions to singular φ-Laplacian with Dirichlet boundary conditions. Numerical Algorithms, 67(2):305–318, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9792-x.

Jiranek:2010:AVS

[JR10] Pavel Jiránek and Miroslav Rozlozník. Adaptive version of Simpler GMRES. Numerical Algorithms, 53(1):93–112, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=53&
issue=1&spage=93.

Jerinkic:2020:PVS

[JR20] Natasa Krklec Jerinkić and Andrea Roznjik. Penalty variable sample size method for solving optimization problems with equality constraints in a form of mathematical expectation. Numerical Algorithms, 83(2):701–718, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jandron:2017:ADS

[JRB17] Michael A. Jandron, Anthony A. Ruffa, and James Baglama. An asynchronous direct solver for banded linear systems. Numerical Algorithms, 76(1):211–235, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Judice:2008:SSE

[JRRS08] Joaquim J. Júdice, Marcos Raydan, Silvério S. Rosa, and Sandra A. Santos. On the solution of the symmetric eigenvalue complementarity problem by the spectral projected gradient algorithm. Numerical Algorithms, 47(4):391–407, April 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=4&
spage=391.

Jbilou:2009:VEE

[JRS09] K. Jbilou, L. Reichel, and H. Sadok. Vector extrapolation enhanced TSVD for linear discrete ill-posed problems. Numerical Algorithms, 51(2):195–208, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=2&spage=195.

Tributes to Gene H. Golub, Part II.

Jbilou:2015:MPE

[JS15] K. Jbilou and H. Sadok. Matrix polynomial and epsilontype extrapolation methods with applications. Numerical Algorithms, 68(1):107–119, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9879-z.

Jesus:2021:NSA

[JS21] Carla Jesus and Ercília Sousa. Numerical solutions for asymmetric Lévy flights. Numerical Algorithms, 87(3):967–999, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00995-6.

Jones:2023:IAK

[JS23] Felix G. Jones and Gideon Simpson. Iterate averaging, the Kalman filter, and 3DVAR for linear inverse problems. Numerical Algorithms, 92(2):1105–1125, February 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01332-9.

Jator:2013:TFB

[JSF13] Samuel N. Jator, S. Swindell, and R. French. Trigonometrically fitted block Numerov type method for y′′ = f(x, y, y′). Numerical Algorithms, 62(1):13–26, January 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9562-1/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=13-26.

Jiang:2022:CAS

[JSZ22] Jie Jiang, Hailin Sun, and Bin Zhou. Convergence analysis of sample average approximation for a class of stochastic nonlinear complementarity problems: from two-stage to multistage. Numerical Algorithms, 89(1):167–194, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01110-z;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01110-z.pdf.

Jackiewicz:1996:VSC

[JT96] Z. Jackiewicz and S. Tracogna. Variable stepsize continuous two-step Runge–Kutta methods for ordinary differential equations. Numerical Algorithms, 12(3–4):347–368, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jiang:2025:CEA

[JTW25a] Nan Jiang, Aziz Takhirov, and Jiajia Waters. Correction to: Ensemble algorithm for parameterized convection problems with energy stable open boundary conditions. Numerical Algorithms, 100(3):1387–1388, November 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-02000-w. See [JTW25b].

Jiang:2025:EAP

[JTW25b] Nan Jiang, Aziz Takhirov, and Jiajia Waters. Ensemble algorithm for parameterized convection problems with energy stable open boundary conditions. Numerical Algorithms, 100(3): 1365–1385, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01991-w. See

correction [JTW25a].

Jarlebring:2022:IAE

[JU22] Elias Jarlebring and Parikshit Upadhyaya. Implicit algorithms for eigenvector nonlinearities. Numerical Algorithms, 90(1):301–321, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01189-4.

Jeannerod:1998:GEE

[JV98] C. P. Jeannerod and J. Visconti. Global error estimation for index-1 and -2 DAEs. Numerical Algorithms, 19 (1–4):111–125, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/11/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Jafari-Varzaneh:2015:NMC

[JVH15] H. A. Jafari-Varzaneh and S. M. Hosseini. A new map for the Chebyshev pseudospectral solution of differential equations with large gradients. Numerical Algorithms, 69(1):95–108, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9883-3.

Jia:2025:NAF

[JW25] Ji-Teng Jia and Yi-Fan Wang. Numerical algorithms for the fast and reliable solution of periodic tridiagonal Toeplitz linear systems. Numerical Algorithms, 98(1):307–324, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01795-y.

Jiang:2021:FOI

[JWCZ21] Fan Jiang, Zhongming Wu, Xingju Cai, and Hongchao Zhang. A first-order inexact primal–dual algorithm for a class of convex-concave saddle point problems. Numerical Algorithms, 88(3):1109–1136, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01069-x.

Jiang:2020:SDS

[JWL20] Yanan Jiang, Lihui Weng, and Wei Liu. Stationary distribution of the stochastic theta method for nonlinear stochastic differential equations. Numerical Algorithms, 83(4):1531–1553, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Jiang:2021:FFG

[JWY21] Ying Jiang, Bo Wang, and Dandan Yu. A fast Fourier–Galerkin method solving boundary integral equations for the Helmholtz equation with exponential convergence. Numerical Algorithms, 88(3):1457–1491, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01082-0.

Jia:2023:FAT

[JWZ23] Jinhong Jia, Hong Wang, and Xiangcheng Zheng. A fast algorithm for time-fractional diffusion equation with spacetime-dependent variable order. Numerical Algorithms, 94(4): 1705–1730, December 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01552-7.

Jia:2024:EBF

[JXNX24] Ji-Teng Jia, Rong Xie, Shuo Ni, and Xiao-Yan Xu. An efficient breakdown-free algorithm for numerically evaluating the determinants of (p, q)-pentadiagonal matrices. Numerical Algorithms, 97(4):2031–2049, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01777-0.

Jia:2023:BBN

[JXX^{+}23] Ji-Teng Jia, Rong Xie, Xiao-Yan Xu, Shuo Ni, and Jie Wang. A bidiagonalization-based numerical algorithm for computing the inverses of (p, q)-tridiagonal matrices. Numerical Algorithms, 93(2):899–917, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01446-0.

Jiang:2026:NKB

[JXZ26] Xingzhou Jiang, Yang Xu, and Jingjun Zhao. A new kind of block fractional generalized Adams method for solving fractional differential equations. Numerical Algorithms, 101(3): 1885–1914, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02067-z.

Jiang:2023:ACS

[JY23] Nan Jiang and Huanhuan Yang. Artificial compressibility SAV ensemble algorithms for the incompressible Navier– Stokes equations. Numerical Algorithms, 92(4):2161–2188, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01382-z.

Jiang:2021:FAF

[JYLC21] Xiaowei Jiang, Yueting Yang, Yunlong Lu, and Mingyuan Cao. Flattened aggregate function method for nonlinear programming with many complicated constraints. Numerical Algorithms, 86(1):103–122, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00881-
1.

Jin:2016:CDZ

[JZ16] Long Jin and Yunong Zhang. Continuous and discrete Zhang dynamics for real-time varying nonlinear optimization. Numerical Algorithms, 73(1):115–140, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0088-1.

Ji:2025:NCI

[JZ25] Bingquan Ji and Xuan Zhao. L^{2} norm convergence of IMEX BDF2 scheme with variable-step for the incompressible Navier– Stokes equations. Numerical Algorithms, 98(4):1917–1942, April 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01858-0.

Jia:2020:FMV

[JZF^{+}20] Jinhong Jia, Xiangcheng Zheng, Hongfei Fu, Pingfei Dai, and Hong Wang. A fast method for variable-order space-fractional diffusion equations. Numerical Algorithms, 85(4):1519–1540, December 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00875-z.

Ji:2023:NEM

[JZYY23] Bingjie Ji, Jiansong Zhang, Yue Yu, and Yun Yu. A new expanded mixed finite element method for Kirchhoff type parabolic equation. Numerical Algorithms, 92(4):2405–2432, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01396-7.

Zokovic:2015:SNP

[– DK15] Dragomir Ž. – Doković and Ilias S. Kotsireas. Some new periodic Golay pairs. Numerical Algorithms, 69(3):523–530, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9910-4.

Kumar:2026:FDC

[KA26] Manoj Kumar and Syed Abbas. Fractal dimension and control of Julia set generated by the discrete competitive model. Numerical Algorithms, 101(1):643–678, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02018-8.

Kacewicz:2018:AMP

[Kac18] Bolesław Kacewicz. Adaptive mesh point selection for the efficient solution of scalar IVPs. Numerical Algorithms, 77(1): 57–75, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Khosravian-Arab:2018:NAI

[KADE18] Hassan Khosravian-Arab, Mehdi Dehghan, and M. R. Eslahchi. A new approach to improve the order of approximation of the Bernstein operators: theory and applications. Numerical Algorithms, 77(1):111–150, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kazmi:2018:HIM

[KAF18a] K. R. Kazmi, Rehan Ali, and Mohd Furkan. Hybrid iterative method for split monotone variational inclusion problem and hierarchical fixed point problem for a finite family of nonexpansive mappings. Numerical Algorithms, 79(2):499–527, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kazmi:2018:KMT

[KAF18b] K. R. Kazmi, Rehan Ali, and Mohd Furkan. Krasnoselski– Mann type iterative method for hierarchical fixed point problem and split mixed equilibrium problem. Numerical Algorithms, 77(1):289–308, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kalantari:2000:NFA

[Kal00] Bahman Kalantari. New formulas for approximation of π and other transcendental numbers. Numerical Algorithms, 24(1–2):59–81, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
27/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/27/5/fulltext.pdf. Computa-

tional methods from rational approximation theory (Wilrijk, 1999).

Kalantari:2022:CLO

[KAL22] Bahman Kalantari, Fedor Andreev, and Chun Lau. Characterization of local optima of polynomial modulus over a disc. Numerical Algorithms, 90(2):773–787, June 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01208-4.

Kamrani:2015:NSS

[Kam15] Minoo Kamrani. Numerical solution of stochastic fractional differential equations. Numerical Algorithms, 68(1):81–93, January 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9839-7.

Kaouane:2020:TMB

[Kao20] Y. Kaouane. A tangential method for the balanced truncation in model reduction. Numerical Algorithms, 83(2):629–652, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Karatsuba:2000:CEC

[Kar00] Ekatharine A. Karatsuba. On the computation of the Euler constant γ. Numerical Algorithms, 24(1–2):83–97, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/6/fulltext.pdf. Computational methods from rational approximation theory (Wilrijk, 1999).

Karatsuba:2007:ASJ

[Kar07] Ekatherina A. Karatsuba. On an approach to the study of the Jaynes–Cummings sum in quantum optics. Numerical Algorithms, 45(1–4):127–137, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=127.

Karageorghis:2009:EMA

[Kar09] Andreas Karageorghis. Efficient MFS algorithms in regular polygonal domains. Numerical Algorithms, 50(2):215–240, February 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
50&issue=2&spage=215.

Karageorghis:2010:EKT

[Kar10] Andreas Karageorghis. Efficient Kansa-type MFS algorithm for elliptic problems. Numerical Algorithms, 54(2):261–278, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=54&
issue=2&spage=261.

Karageorghis:2013:ETC

[Kar13] A. Karageorghis. Efficient Trefftz collocation algorithms for elliptic problems in circular domains. Numerical Algorithms, 64(3):427–453, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9673-8.

Karageorghis:2015:MFS

[Kar15] Andreas Karageorghis. The method of fundamental solutions for elliptic problems in circular domains with mixed boundary conditions. Numerical Algorithms, 68(1):185–211, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9900-6.

Kazmi:2024:FHO

[Kaz24] Kamran Kazmi. A fast and high-order IMEX method for non-linear time–space-fractional reaction–diffusion equations. Numerical Algorithms, 95(1):243–266, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01570-5.

Kerr:2002:NMM

[KB02] Minnie Kerr and Kevin Burrage. New multivalue methods for differential algebraic equations. Numerical Algorithms, 31(1–4):193–213, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
43/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/43/2/fulltext.pdf.

Kumar:2020:LMM

[KB20] Alpesh Kumar and Akanksha Bhardwaj. A local meshless method for time fractional nonlinear diffusion wave equation. Numerical Algorithms, 85(4):1311–1334, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00866-9.

Khoshsimaye-Bargard:2023:DSD

[KBA23] Maryam Khoshsimaye-Bargard and Ali Ashrafi. A descent spectral Dai–Liao method based on the quasi-Newton aspects. Numerical Algorithms, 94(1):397–411, September 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01506-z.

Khoshsimaye-Bargard:2026:FDS

[KBA26] Maryam Khoshsimaye-Bargard and Ali Ashrafi. A family of descent spectral three-term conjugate gradient methods based on the quasi-Newton aspects with applications to nonnegative matrix factorization and image restoration. Numerical Algorithms, 101(4):2381–2412, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02091-z.

Kalantzis:2013:ADU

[KBCG13] V. Kalantzis, C. Bekas, A. Curioni, and E. Gallopoulos. Accelerating data uncertainty quantification by solving linear systems with multiple right-hand sides. Numerical Algorithms, 62 (4):637–653, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9687-2.

Kundu:2017:ABF

[KBP17] Sudeep Kundu, Saumya Bajpai, and Amiya K. Pani. Asymptotic behavior and finite element error estimates of Kelvin– Voigt viscoelastic fluid flow model. Numerical Algorithms, 75 (3):619–653, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Knaepkens:2023:REB

[KC23] Ferre Knaepkens and Annie Cuyt. On the robustness of exponential base terms and the Padé denominator in some least squares sense. Numerical Algorithms, 92(1):747–766, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01455-z.

Kong:2024:ROM

[KC24] Ling Chang Kong and Xiao Shan Chen. Riemannian optimization methods for the truncated Takagi factorization. Numerical Algorithms, 97(1):223–242, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01701-y.

Kansal:2021:NFS

[KCBT21] Munish Kansal, Alicia Cordero, Sonia Bhalla, and Juan R. Torregrosa. New fourth- and sixth-order classes of iterative methods for solving systems of nonlinear equations and their stability analysis. Numerical Algorithms, 87(3):1017–1060, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00997-4.

Kazem:2016:UGF

[KCHD16] Saeed Kazem, Edmund Chadwick, Ali Hatam, and Mehdi Dehghan. Using generating functions to convert an implicit (3, 3) finite difference method to an explicit form on diffusion equation with different boundary conditions. Numerical Algorithms, 71(4):827–854, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0026-2.

Kchouk:2014:PIC

[KD14] Bilel Kchouk and Jean-Pierre Dussault. On per-iteration complexity of high order Chebyshev methods for sparse functions with banded Hessians. Numerical Algorithms, 66(4):885–909, August 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9767-y.

Kim:2018:PEA

[KD18] Do Sang Kim and Bui Van Dinh. Parallel extragradient algorithms for multiple set split equilibrium problems in Hilbert spaces. Numerical Algorithms, 77(3):741–761, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kimiaei:2016:TRA

[KE16] Morteza Kimiaei and Hamid Esmaeili. A trust-region approach with novel filter adaptive radius for system of nonlinear equations. Numerical Algorithms, 73(4):999–1016, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0126-7.

Ke:2021:CAM

[Ke21] Yi-Fen Ke. Convergence analysis on matrix splitting iteration algorithm for semidefinite linear complementarity problems. Numerical Algorithms, 86(1):257–279, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00888-8.

Keenan:1994:MMQ

[Kee94] Philip T. Keenan. Mixed methods on quadrilaterals and hexahedra. Numerical Algorithms, 7(2–4):269–293, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Keller:2007:MII

[Kel07] Pawel Keller. A method for indefinite integration of oscillatory and singular functions. Numerical Algorithms, 46(3):219–251, November 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
46&issue=3&spage=219.

Kovalnogov:2024:RKP

[KFK^{+}24] Vladislav N. Kovalnogov, Ruslan V. Fedorov, Tamara V. Karpukhina, Theodore E. Simos, and Charalampos Tsitouras. Runge–Kutta pairs of orders 9(8) for use in quadruple precision computations. Numerical Algorithms, 95(4):1905–1919, April 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01632-8.

Kar:2023:IBZ

[KG23] Pinaki Prasad Kar and Priyabrat Gochhayat. Interlacing and bounds of zeros of quasi-orthogonal little q-Jacobi polynomials. Numerical Algorithms, 93(3):1157–1170, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01460-2.

Khachan:2003:CTN

[KGD03] Mohammed Khachan, François Guibault, and Hafsa Deddi. Characterization of trimmed NURBS surface boundaries using topological criteria. Numerical Algorithms, 34(2– 4):355–366, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/4/fulltext.pdf.

Kokosinski:2025:FAA

[KGG25] Zbigniew Kokosiński, Paweł Gepner, and Nataliia Gavkalova. Fast and accurate approximation algorithms for computing floating point square root. Numerical Algorithms, 99(4): 1791–1804, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01932-7.

Khasi:2014:NAH

[KGH14] M. Khasi, F. Ghoreishi, and M. Hadizadeh. Numerical analysis of a high order method for state-dependent delay integral equations. Numerical Algorithms, 66(1):177–201, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9729-4.

Kumar:2021:CDI

[KGMH21] Abhimanyua Kumar, D. K. Gupta, Eulalia Martínez, and José L. Hueso. Convergence and dynamics of improved Chebyshev–secant-type methods for non differentiable operators. Numerical Algorithms, 86(3):1051–1070, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00922-9.

Kumari:2024:VAT

[KGN^{+}24] Sudesh Kumari, Krzysztof Gdawiec, Ashish Nandal, Naresh Kumar, and Renu Chugh. On the viscosity approximation type iterative method and its non-linear behaviour in the generation of Mandelbrot and Julia sets. Numerical Algorithms, 96(1):211–236, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01644-4.

Khan:2011:ALD

[KH11] Majid Khan and Mazhar Hussain. Application of Laplace decomposition method on semi-infinite domain. Numerical Algorithms, 56(2):211–218, February 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=2&spage=211.

Kovac:2018:JMS

[KH18] Erna Begović Kovac and Vjeran Hari. Jacobi method for symmetric 4 × 4 matrices converges for every cyclic pivot strategy. Numerical Algorithms, 78(3):701–720, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kheirfam:2020:FNS

[KH20] Behrouz Kheirfam and Masoumeh Haghighi. A full-Newton step infeasible interior-point method based on a trigonometric kernel function without centering steps. Numerical Algorithms, 85(1):59–75, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00802-x.

Khashin:2013:BAB

[Kha13] Sergey Khashin. Butcher algebras for Butcher systems. Numerical Algorithms, 63(4):679–689, August 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9647-x.

Khashin:2014:SER

[Kha14] Sergey Khashin. Symmetries of explicit Runge–Kutta methods. Numerical Algorithms, 65(3):597–609, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9829-9.

Khasi:2025:CAB

[Kha25] Manoochehr Khasi. A computational approach based on the Legendre–Galerkin method for solving a distributed optimal control problem constrained by the biharmonic equation. Numerical Algorithms, 98(3):1231–1259, March 2025.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01832-w.

Kheirfam:2012:PDI

[Khe12a] Behrouz Kheirfam. Primal-dual interior-point algorithm for semidefinite optimization based on a new kernel function with trigonometric barrier term. Numerical Algorithms, 61(4): 659–680, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9557-y/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=61&issue=4&spage=659-680.

Kheirfam:2012:SII

[Khe12b] Behrouz Kheirfam. Simplified infeasible interior-point algorithm for SDO using full Nesterov–Todd step. Numerical Algorithms, 59(4):589–606, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=589.

Kheirfam:2014:PCI

[Khe14] Behrouz Kheirfam. A predictor–corrector interior-point algorithm for P_{∗}(κ)-horizontal linear complementarity problem. Numerical Algorithms, 66(2):349–361, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9738-3.

Kheirfam:2016:IFN

[Khe16] Behrouz Kheirfam. An improved full-Newton step O(n) infeasible interior-point method for horizontal linear complementarity problem. Numerical Algorithms, 71(3):491–503, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0005-7.

Kheirfam:2017:IFN

[Khe17] Behrouz Kheirfam. An infeasible full-NT step interior point algorithm for CQSCO. Numerical Algorithms, 74(1):93–109, January 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0140-9.

Kurtoglu:2020:MEC

[KHM20] Dilan Kiliç Kurtoğlu, A. Ihsan Hasçelik, and Gradimir V. Milovanović. A method for efficient computation of integrals with oscillatory and singular integrand. Numerical Algorithms, 85 (4):1155–1173, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00859-8.

Kieffer:2023:CNS

[Kie23] Jean Kieffer. Certified Newton schemes for the evaluation of low-genus theta functions. Numerical Algorithms, 93(2):833– 862, June 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01443-3.

Kim:2021:FSE

[Kim21] Minho Kim. Fast and stable evaluation of splines and their derivatives generated by the seven-direction quartic box-spline. Numerical Algorithms, 86(2):887–909, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00916-7.

Kimiaei:2026:IRA

[Kim26] Morteza Kimiaei. An improved randomized algorithm with noise level tuning for large-scale noisy unconstrained DFO problems. Numerical Algorithms, 101(1):565–603, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02016-w.

Krejic:2015:NLS

[KJ15] Natasa Krejić and Natasa Krklec Jerinkić. Nonmonotone line search methods with variable sample size. Numerical Algorithms, 68(4):711–739, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9869-1.

Kamrani:2018:IMM

[KJ18] Minoo Kamrani and Nahid Jamshidi. Implicit Milstein method for stochastic differential equations via the Wong–Zakai approximation. Numerical Algorithms, 79(2):357–374, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Karageorghis:2018:KRA

[KJC18] Andreas Karageorghis, Malgorzata A. Jankowska, and C. S. Chen. Kansa–RBF algorithms for elliptic problems in regular polygonal domains. Numerical Algorithms, 79(2):399–421, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kar:2023:SIZ

[KJG23] Pinaki Prasad Kar, Kerstin Jordaan, and Priyabrat Gochhayat. Stieltjes interlacing of zeros of little q-Jacobi and q-Laguerre polynomials from different sequences. Numerical Algorithms, 92(1):723–746, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01387-8.

Krejic:2023:SPS

[KJO23] Natasa Krejić, Natasa Krklec Jerinkić, and Tijana Ostojić. Spectral projected subgradient method for nonsmooth convex optimization problems. Numerical Algorithms, 93(1):347– 365, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01419-3.

Karavelas:2000:SSP

[KK00] M. I. Karavelas and P. D. Kaklis. Spatial shape-preserving interpolation using ν-splines. Numerical Algorithms, 23 (2–3):217–250, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
25/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/25/4/fulltext.pdf.

Kumar:2012:PRN

[KK12] Sunil Kumar and Mukesh Kumar. Parameter-robust numerical method for a system of singularly perturbed initial value problems. Numerical Algorithms, 59(2):185–195, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=2&
spage=185.

Kumar:2016:ASN

[KK16] Sunil Kumar and Mukesh Kumar. Analysis of some numerical methods on layer adapted meshes for singularly perturbed quasilinear systems. Numerical Algorithms, 71(1):139–150, January 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9989-2.

Kumar:2017:SOU

[KK17] Sunil Kumar and Mukesh Kumar. A second order uniformly convergent numerical scheme for parameterized singularly perturbed delay differential problems. Numerical Algorithms, 76 (2):349–360, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Khieu:2022:FFM

[KK22a] Tran Thi Khieu and Tra Quoc Khanh. Fractional filter method for recovering the historical distribution for diffusion equations with coupling operator of local and nonlocal type. Numerical Algorithms, 89(4):1743–1767, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01171-0.

Kumari:2022:EBS

[KK22b] Archna Kumari and V. K. Kukreja. Error bounds for septic Hermite interpolation and its implementation to study modified Burgers’ equation. Numerical Algorithms, 89(4):1799– 1821, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01173-y.

Kim:2023:AEF

[KK23] Hyunjun Kim and Minho Kim. Analytic evaluation of the FCC Voronoi-splines. Numerical Algorithms, 94(4):2005–2030, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01562-5.

Kumar:2017:NAK

[KKA17] Sunil Kumar, Amit Kumar, and Ioannis K. Argyros. A new analysis for the Keller–Segel model of fractional order. Numerical Algorithms, 75(1):213–228, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kansal:2016:EDF

[KKB16] Munish Kansal, V. Kanwar, and Saurabh Bhatia. Efficient derivative-free variants of Hansen–Patrick’s family with memory for solving nonlinear equations. Numerical Algorithms, 73(4):1017–1036, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0127-6.

Kumar:2025:SOP

[KKD25] Shashikant Kumar, Sunil Kumar, and Pratibhamoy Das. Second-order a priori and a posteriori error estimations for integral boundary value problems of nonlinear singularly perturbed parameterized form. Numerical Algorithms, 99(3): 1343–1363, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01918-5.

Kubelik:2022:CFM

[KKK22] P. Kubelík, V. G. Kurbatov, and I. V. Kurbatova. Calculating a function of a matrix with a real spectrum. Numerical Algorithms, 90(3):905–930, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01214-
6.

Kressner:2020:LRU

[KKM20] Daniel Kressner, Patrick Kürschner, and Stefano Massei. Lowrank updates and divide-and-conquer methods for quadratic matrix equations. Numerical Algorithms, 84(2):717–741, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00776-w.

Kobayashi:2017:MVA

[KKO17] Ryo Kobayashi, Takuma Kimura, and Shin’ichi Oishi. A method for verifying the accuracy of numerical solutions of symmetric saddle point linear systems. Numerical Algorithms, 76(1):33–51, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kachakhidze:2016:CMO

[KKPT16] N. Kachakhidze, N. Khomeriki, J. Peradze, and Z. Tsiklauri. Chipot’s method for a one-dimensional Kirchhoff static equation. Numerical Algorithms, 73(4):1091–1106, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0131-x.

Kalyani:2025:BIM

[KKR25] Mehdi Najafi Kalyani, Malihe Nobakht Kooshkghazi, and Lothar Reichel. Block iterative methods for solving multilinear systems. Numerical Algorithms, 100(2):803–829, October 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02077-x.

Kumar:2022:PEE

[KKS22] Sunil Kumar, Shashikant Kumar, and Sumit. A posteriori error estimation for quasilinear singularly perturbed problems with integral boundary condition. Numerical Algorithms, 89(2): 791–809, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01134-5.

Kumar:2024:PPE

[KKS24] Sunil Kumar, Shashikant Kumar, and Sumit. A priori and a posteriori error estimation for singularly perturbed delay integro-differential equations. Numerical Algorithms, 95(4): 1561–1582, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01620-y.

Kammerer:2022:SES

[KKV22] Lutz Kämmerer, Felix Krahmer, and Toni Volkmer. A sample efficient sparse FFT for arbitrary frequency candidate sets in high dimensions. Numerical Algorithms, 89(4):1479–1520, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01162-1.

Kulkarni:1991:S

[KL91] Rekha Kulkarni and Pierre-Jean Laurent. Q-splines. Numerical Algorithms, 1(1):45–73, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kenney:1994:HTI

[KL94] C. S. Kenney and A. J. Lamb. A hyperbolic tangent identity and the geometry of Padé sign function iterations. Numerical Algorithms, 7(2–4):111–128, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Krivsky:2004:UIA

[KL04] Stefanie Krivsky and Bruno Lang. Using interval arithmetic for determining the structure of convex hulls. Numerical Algorithms, 37(1–4):233–240, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/18/abstract.htm.

Kou:2006:VCM

[KL06] Jisheng Kou and Yitian Li. A variant of Chebyshev’s method with sixth-order convergence. Numerical Algorithms, 43(3):273–278, November 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=3&spage=273.

Kehlet:2017:PEA

[KL17] Benjamin Kehlet and Anders Logg. A posteriori error analysis of round-off errors in the numerical solution of ordinary differential equations. Numerical Algorithms, 76(1):191–210, September 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-016-0250-4.pdf.

Kelly:2022:AEM

[KL22] Cónall Kelly and Gabriel J. Lord. Adaptive Euler methods for stochastic systems with non-globally Lipschitz coefficients. Numerical Algorithms, 89(2):721–747, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01131-8;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01131-8.pdf.

Karageorghis:2025:SCP

[KL25] Andreas Karageorghis and Daniel Lesnic. Solution of the Cauchy problem for the Brinkman equations using an alternating method of fundamental solutions. Numerical Algorithms, 98(3):1361–1393, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01837-5.

Karaman:2024:SSO

[KLA24] Rabia Aktas Karaman, Esra Güldogan Lekesiz, and Yelda Aygar. A study on Sobolev orthogonal polynomials on a triangle. Numerical Algorithms, 97(2):915–944, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01732-5.

Kolb:2010:IBS

[KLB10] Oliver Kolb, Jens Lang, and Pia Bales. An implicit box scheme for subsonic compressible flow with dissipative source term. Numerical Algorithms, 53(2–3):293–307, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=293.

Khaliq:2017:FOI

[KLF17] A. Q. M. Khaliq, X. Liang, and K. M. Furati. A fourthorder implicit-explicit scheme for the space fractional nonlinear Schrödinger equations. Numerical Algorithms, 75(1):147–172, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kovacs:2010:SCF

[KLL10] Mihály Kovács, Stig Larsson, and Fredrik Lindgren. Strong convergence of the finite element method with truncated noise for semilinear parabolic stochastic equations with additive noise. Numerical Algorithms, 53(2–3):309–320, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=309.

Koch:2025:SSP

[KLL25] Michael Koch, Sabine Le Borne, and Willi Leinen. Smaller stencil preconditioners for linear systems in RBF–FD discretizations. Numerical Algorithms, 98(3):1313–1335, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01835-7.

Kupper:2007:SSC

[KLR07] Dominique Küpper, Jürgen Lehn, and Andreas Rößler. A step size control algorithm for the weak approximation of stochastic differential equations. Numerical Algorithms, 44(4):335–346, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=44&
issue=4&spage=335.

Kresoja:2017:ASA

[KLS17] Milena Kresoja, Zorana Luzanin, and Irena Stojkovska. Adaptive stochastic approximation algorithm. Numerical Algorithms, 76(4):917–937, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kang:2014:SMC

[KLSS14] Kingston Kang, William Lothian, Jessica Sears, and Brian D. Sutton. Simultaneous multidiagonalization for the CS decomposition. Numerical Algorithms, 66(3):479–493, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9744-5.

Kaufman:1995:OBA

[KLT95] E. H. Kaufman, Jr., D. J. Leeming, and G. D. Taylor. An ODEbased approach to nonlinearly constrained minimax problems. Numerical Algorithms, 9(1–2):25–37, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Kaufman:2003:AMR

[KLT03] E. H. Kaufman, Jr., D. J. Leeming, and G. D. Taylor. Adaptive monotone rational approximation on finite sets. Numerical Algorithms, 32(1):1–12, January 2003. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
44/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/44/1/fulltext.pdf.

Kum:2023:RLP

[KLW^{+}23] Sangho Kum, Chong Li, Jinhua Wang, Jen-Chih Yao, and Linglingzhi Zhu. Riemannian linearized proximal algorithms for nonnegative inverse eigenvalue problem. Numerical Algorithms, 94(4):1819–1848, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01556-3.

Krajewski:1995:PSR

[KLZV95] W. Krajewski, A. Lepschy, M. Redivo Zaglia, and U. Viaro. A program for solving the L_{2} reduced-order model problem with fixed denominator degree. Numerical Algorithms, 9(3–4):355– 377, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kalovics:2004:BVF

[KM04] Ferenc Kálovics and Gabriella Mészáros. Box valued functions in solving systems of equations and inequalities. Numerical Algorithms, 36(1):1–12, May 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/52/A/2/abstract.htm.

Kravvaritis:2009:CMP

[KM09] Christos Kravvaritis and Marilena Mitrouli. Compound matrices: properties, numerical issues and analytical computations. Numerical Algorithms, 50(2):155–177, February 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=155.

Kravvaritis:2013:CPC

[KM13] Christos Kravvaritis and Marilena Mitrouli. On the complete pivoting conjecture for Hadamard matrices: further progress and a good pivots property. Numerical Algorithms, 62(4): 571–582, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9643-1.

Ke:2017:IMR

[KM17] Yi-Fen Ke and Chang-Feng Ma. An inexact modified relaxed splitting preconditioner for the generalized saddle point problems from the incompressible Navier–Stokes equations. Numerical Algorithms, 75(4):1103–1121, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kopecz:2019:ETS

[KM19] Stefan Kopecz and Andreas Meister. On the existence of threestage third-order modified Patankar–Runge–Kutta schemes. Numerical Algorithms, 81(4):1473–1484, August 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kaur:2024:HOH

[KM24a] Deepti Kaur and R. K. Mohanty. High-order half-step compact numerical approximation for fourth-order parabolic PDEs. Numerical Algorithms, 95(3):1127–1153, March 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01602-0.

Kayal:2024:NAS

[KM24b] Arnab Kayal and Moumita Mandal. A new approach of shifted Jacobi spectral Galerkin methods (SJSGM) for weakly singular Fredholm integral equation with non-smooth solution. Numerical Algorithms, 96(4):1553–1582, August 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01677-9.

Khorramizadeh:2013:EAS

[KMA13] M. Khorramizadeh and N. Mahdavi-Amiri. An efficient algorithm for sparse null space basis problem using ABS methods. Numerical Algorithms, 62(3):469–485, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9599-1/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=469-485.

Kaafi:2024:OJG

[KMH24] R. Kaafi, P. Mokhtary, and E. Hesameddini. Operational Jacobi Galerkin method for a class of cordial Volterra integral equations. Numerical Algorithms, 96(2):827–843, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01667-x.

Khakzad:2025:SSP

[KMH25] Pari Khakzad, Afsaneh Moradi, and Gholamreza Hojjati. Strong stability preserving second derivative multistep methods. Numerical Algorithms, 100(1):85–108, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01956-z.

Khiar:2025:ACN

[KMRAR25] Y. Khiar, E. Mainar, E. Royo-Amondarain, and B. Rubio. On the accurate computation of the Newton form of the Lagrange interpolant. Numerical Algorithms, 98(3):1553–1573, March 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01843-7.

Kopteva:2005:GER

[KMS05] Natalia Kopteva, Niall Madden, and Martin Stynes. Grid equidistribution for reaction–diffusion problems in one dimension. Numerical Algorithms, 40(3):305–322, November 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=3&
spage=305.

Kressner:2023:MPL

[KMS23] Daniel Kressner, Yuxin Ma, and Meiyue Shao. A mixed precision LOBPCG algorithm. Numerical Algorithms, 94(4): 1653–1671, December 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01550-9.

Koleva:2017:FOC

[KMV17] Miglena N. Koleva, Walter Mudzimbabwe, and Lubin G. Vulkov. Fourth-order compact schemes for a parabolicordinary system of European option pricing liquidity shocks model. Numerical Algorithms, 74(1):59–75, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0138-3.

Ke:2018:MBM

[KMZ18] Yi-Fen Ke, Chang-Feng Ma, and Huai Zhang. The modulusbased matrix splitting iteration methods for second-order cone linear complementarity problems. Numerical Algorithms, 79 (4):1283–1303, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kubinova:2018:RRM

[KN18] Marie Kubínová and James G. Nagy. Robust regression for mixed Poisson–Gaussian model. Numerical Algorithms, 79(3): 825–851, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kunis:2021:CCN

[KN21a] Stefan Kunis and Dominik Nagel. Correction to: On the condition number of Vandermonde matrices with pairs of nearly-colliding nodes. Numerical Algorithms, 87(1):497, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01015-3;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01015-3.pdf. See

[KN21b].

Kunis:2021:CNV

[KN21b] Stefan Kunis and Dominik Nagel. On the condition number of Vandermonde matrices with pairs of nearly-colliding nodes. Numerical Algorithms, 87(1):473–496, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00974-x;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00974-x.pdf. See

correction [KN21a].

Kaur:2023:NNS

[KN23] Jaspreet Kaur and Srinivasan Natesan. A novel numerical scheme for time-fractional Black–Scholes PDE governing European options in mathematical finance. Numerical Algorithms, 94(4):1519–1549, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01545-6.

Karouma:2018:NCE

[KNBGV18] Abdulrahman Karouma, Truong Nguyen-Ba, Thierry Giordano, and Rémi Vaillancourt. A new class of efficient one-step contractivity preserving high-order time discretization methods of order 5 to 14. Numerical Algorithms, 79(1):251–280, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Knobloch:2023:ASM

[Kno23] Petr Knobloch. An algebraically stabilized method for convection–diffusion-reaction problems with optimal experimental convergence rates on general meshes. Numerical Algorithms, 94(2):547–580, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01511-
2.

Kobbelt:1997:SEB

[Kob97] Leif Kobbelt. Stable evaluation of box-splines. Numerical Algorithms, 14(4):377–382, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
6/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/6/6/fulltext.pdf.

Kheirfam:2021:ASI

[KOK21] Behrouz Kheirfam, Naser Osmanpour, and Mohammad Keyanpour. An arc-search infeasible interior-point method for semidefinite optimization with the negative infinity neighborhood. Numerical Algorithms, 88(1):143–163, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01033-1.

Kolev:2004:IIL

[Kol04a] Lubomir V. Kolev. An improved interval linearization for solving nonlinear problems. Numerical Algorithms, 37(1–4):213–224, December 2004. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp009.kluweronline.com/IPS/content/ext/
x/J/5058/I/58/A/16/abstract.htm.

Kolev:2004:SLS

[Kol04b] Lubomir V. Kolev. Solving linear systems whose elements are nonlinear functions of intervals. Numerical Algorithms, 37(1–4):199–212, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/15/abstract.htm.

Kolotilina:2006:BPR

[Kol06] Lilia Yu. Kolotilina. Bounds for the Perron root, singularity/nonsingularity conditions, and eigenvalue inclusion sets. Numerical Algorithms, 42(3–4):247–280, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=3&spage=247.

Koumandos:2007:CIG

[Kou07] Stamatis Koumandos. On a conjectured inequality of Gautschi and Leopardi for Jacobi polynomials. Numerical Algorithms, 44(3):249–253, March 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=3&spage=249.

Kowalski:2000:SBR

[Kow00] Marek A. Kowalski. Sensitivity of best recovery in the Sobolev spaces W_{∞}^{r,d}, W̃_{∞}^{r,d} for perturbed sampling. Numerical Algorithms, 23(2–3):251–261, June 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
25/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/25/2/fulltext.pdf.

Kaagstrom:1996:CES

[KP96a] Bo Kågström and Peter Poromaa. Computing eigenspaces with specified eigenvalues of a regular matrix pair (A, B) and condition estimation: theory, algorithms and software. Numerical Algorithms, 12(3–4):369–407, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Karageorghis:1996:EDM

[KP96b] Andreas Karageorghis and Marcin Paprzycki. An efficient direct method for fully conforming spectral collocation schemes. Numerical Algorithms, 12(3–4):309–319, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kzaz:2003:CAG

[KP03] M. Kzaz and M. Prévost. Convergence acceleration of Gauss–Chebyshev quadrature formulae. Numerical Algorithms, 34(2–4):379–391, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/6/fulltext.pdf.

Kim:2009:FSE

[KP09] Minho Kim and Jörg Peters. Fast and stable evaluation of boxsplines via the BB-form. Numerical Algorithms, 50(4):381–399, April 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=50&
issue=4&spage=381.

Kacewicz:2016:ORS

[KP16] Bolesław Kacewicz and Pawel Przybylowicz. On the optimal robust solution of IVPs with noisy information. Numerical Algorithms, 71(3):505–518, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0006-6.

Kwak:2022:FCD

[KP22] Do Y. Kwak and Hyeokjoo Park. A formal construction of a divergence-free basis in the nonconforming virtual element method for the Stokes problem. Numerical Algorithms, 91(1): 449–471, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01269-z.

Kumari:2025:ETA

[KP25] Sarita Kumari and Rajesh K. Pandey. An efficient temporal approximation for weakly singular time-fractional semilinear diffusion-wave equation with variable coefficients. Numerical Algorithms, 100(1):189–225, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01959-w.

Kileel:2026:SPM

[KP26] Joe Kileel and João M. Pereira. Subspace power method for symmetric tensor decomposition. Numerical Algorithms, 102 (2):1011–1048, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02165-y.

Kouibia:2020:VMS

[KPA20] A. Kouibia, M. Pasadas, and R. Akhrif. A variational method for solving two-dimensional Bratu’s problem. Numerical Algorithms, 84(4):1589–1599, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00957-
y.

Kesornprom:2020:CAG

[KPC20] Suparat Kesornprom, Nattawut Pholasa, and Prasit Cholamjiak. On the convergence analysis of the gradient-CQ algorithms for the split feasibility problem. Numerical Algorithms, 84 (3):997–1017, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00790-y.

Kamil:2025:ECA

[KPD25] Safa Hussein Kamil, Saeed Pishbin, and Parviz Darania. An effective computational approach based on new multistep collocation method for solving integral–algebraic equations. Numerical Algorithms, 100(3):1333–1363, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01990-x.

Khebchareon:2024:ADI

[KPFF24] Morrakot Khebchareon, Amiya K. Pani, Graeme Fairweather, and Ryan I. Fernandes. An alternating direction implicit finite element Galerkin method for the linear Schrödinger equation. Numerical Algorithms, 97(3):1039–1066, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01740-5.

Kreinovich:2004:FAL

[KPFG04] Vladik Kreinovich, Eric J. Pauwels, Scott A. Ferson, and Lev Ginzburg. A feasible algorithm for locating concave and convex zones of interval data and its use in statistics-based clustering. Numerical Algorithms, 37(1–4):225–232, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/17/abstract.htm.

Kouibia:2003:COC

[KPR03] A. Kouibia, M. Pasadas, and M. L. Rodríguez. Construction of ODE curves. Numerical Algorithms, 34(2–4):367–377, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/48/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
48/5/fulltext.pdf.

Kressner:2014:IVL

[KPS14] Daniel Kressner, Marija Miloloza Pandur, and Meiyue Shao. An indefinite variant of LOBPCG for definite matrix pencils. Numerical Algorithms, 66(4):681–703, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9754-3.

Kim:2022:EAP

[KPS22] Dongho Kim, Eun-Jae Park, and Boyoon Seo. Error analysis for the pseudostress formulation of unsteady Stokes problem. Numerical Algorithms, 91(2):959–996, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01288-w.

Kouibia:2003:CSP

[KPT03] A. Kouibia, M. Pasadas, and J. J. Torrens. Construction of surfaces with parallelism conditions. Numerical Algorithms, 33(1–4):331–342, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/

27/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/27/fulltext.pdf.

Kircheis:2023:NFF

[KPT23] Melanie Kircheis, Daniel Potts, and Manfred Tasche. Nonuniform fast Fourier transforms with nonequispaced spatial and frequency data and fast sinc transforms. Numerical Algorithms, 92(4):2307–2339, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01389-
6.

Kounchev:2007:CPS

[KR07] O. Kounchev and H. Render. Convergence of polyharmonic splines on semi-regular grids Z × aZ^{n} for a ⊗ 0. Numerical Algorithms, 44(3):255–272, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=3&spage=255.

Kemppainen:2011:SCM

[KR11] Jukka T. Kemppainen and Keijo Matti Ruotsalainen. On the spline collocation method for the single layer equation related to time-fractional diffusion. Numerical Algorithms, 57(3):313– 327, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
57&issue=3&spage=313.

Kulkarni:2020:RED

[KR20] Rekha P. Kulkarni and Gobinda Rakshit. Richardson extrapolation for the discrete iterated modified projection solution. Numerical Algorithms, 85(1):171–189, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00808-5.

Kastner:2023:AAA

[KR23] Felix Kastner and Andreas Rößler. An analysis of approximation algorithms for iterated stochastic integrals and a Julia and Matlab simulation toolbox. Numerical Algorithms, 93(1):27– 66, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01401-z.

Krivovichev:2026:AGJ

[Kri26] Gerasim Krivovichev. The application of Goeken–Johnson’s Runge–Kutta methods in unconstrained convex optimization. Numerical Algorithms, 101(1):149–197, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-02001-9.

Kuian:2019:FFR

[KRS19] Mykhailo Kuian, Lothar Reichel, and Sergij V. Shiyanovskii. Fast factorization of rectangular Vandermonde matrices with Chebyshev nodes. Numerical Algorithms, 81(2):547–559, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kloeden:1997:NSC

[KS97] Peter E. Kloeden and Björn Schmalfuß. Nonautonomous systems, cocycle attractors and variable time-step discretization. Numerical Algorithms, 14(1–3):141–152, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/5/8/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/5/

8/fulltext.pdf. Dynamical numerical analysis (Atlanta, GA, 1995).

Kulikov:2006:OLV

[KS06] Gennady Yu. Kulikov and Sergey K. Shindin. One-leg variable-coefficient formulas for ordinary differential equations and local-global step size control. Numerical Algorithms, 43(1):99–121, September 2006. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=1&spage=99.

Khan:2012:NPQ

[KS12] Arshad Khan and Talat Sultana. Non-polynomial quintic spline solution for the system of third order boundaryvalue problems. Numerical Algorithms, 59(4):541–559, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=4&
spage=541.

Khattri:2014:AFD

[KS14] Sanjay K. Khattri and Trond Steihaug. Algorithm for forming derivative-free optimal methods. Numerical Algorithms, 65(4): 809–824, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9715-x.

Kanth:2018:ANS

[KS18a] A. S. V. Ravi Kanth and Deepika Sirswal. Analysis and numerical simulation for a class of time fractional diffusion equation via tension spline. Numerical Algorithms, 79(2):479–497, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kolev:2018:ESP

[KS18b] Lubomir Kolev and Iwona Skalna. Exact solution to a parametric linear programming problem. Numerical Algorithms, 78 (4):1183–1194, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Khalsaraei:2020:ESS

[KS20] Mohammad Mehdizadeh Khalsaraei and Ali Shokri. An explicit six-step singularly P -stable Obrechkoff method for the numerical solution of second-order oscillatory initial value problems. Numerical Algorithms, 84(3):871–886, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00784-w.

Kostic:2023:MRB

[KS23] Vladimir R. Kostić and Saverio Salzo. The method of randomized Bregman projections for stochastic feasibility problems. Numerical Algorithms, 93(3):1269–1307, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01468-8.

Kamdem:2025:SCN

[KS25] Jules Djoko Kamdem and Toni Sayah. Stopping criteria for nonlinear variational problems: iterative approach. Numerical Algorithms, 100(1):487–527, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01969-8.

Kim:2026:EIS

[KS26] Mi-Young Kim and Dongwook Shin. An edgewise iterative scheme for the discontinuous Galerkin method with Lagrange multiplier for Poisson’s equation. Numerical Algorithms, 101 (1):605–642, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02017-9.

Kanwar:2008:SGC

[KSB08] V. Kanwar, Sukhjit Singh, and S. Bakshi. Simple geometric constructions of quadratically and cubically convergent iterative functions to solve nonlinear equations. Numerical Algorithms, 47(1):95–107, January 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=1&spage=95.

Khater:2007:NSI

[KSCS07] A. H. Khater, A. B. Shamardan, D. K. Callebaut, and M. R. A. Sakran. Numerical solutions of integral and integrodifferential equations using Legendre polynomials. Numerical Algorithms, 46(3):195–218, November 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=3&spage=195.

Khater:2008:SIE

[KSCS08] A. H. Khater, A. B. Shamardan, D. K. Callebaut, and M. R. A. Sakran. Solving integral equations with logarithmic kernels by Chebyshev polynomials. Numerical Algorithms, 47(1):81–93, January 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=1&spage=81.

Kumar:2026:ATO

[KSJ26] Naresh Kumar, Jasbir Singh, and Ram Jiwari. An ADI type operator splitting WG-FEM for 2D nonlinear unsteady singularly perturbed problem. Numerical Algorithms, 101(3):

1701–1738, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02061-5.

Karageorghis:2006:MDM

[KST06] A. Karageorghis, Y.-S. Smyrlis, and T. Tsangaris. A matrix decomposition MFS algorithm for certain linear elasticity problems. Numerical Algorithms, 43(2):123–149, October 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=43&issue=2&
spage=123.

Karvonen:2021:CKB

[KST21a] Toni Karvonen, Simo Särkkä, and Ken’ichiro Tanaka. Correction to: Kernel-based interpolation at approximate Fekete points. Numerical Algorithms, 87(1):469–471, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01034-0;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01034-0.pdf. See

[KST21b].

Karvonen:2021:KBI

[KST21b] Toni Karvonen, Simo Särkkä, and Ken’ichiro Tanaka. Kernelbased interpolation at approximate Fekete points. Numerical Algorithms, 87(1):445–468, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00973-

y. See correction [KST21a].

K:2023:TOI

[KSV23] Dhivya Prabhu K, Sanjeev Singh, and Antony Vijesh V. A third-order iterative algorithm for inversion of cumulative central beta distribution. Numerical Algorithms, 94(3):1331–1353, November 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01537-6.

Kuo:2007:PSM

[KSW07] Frances Y. Kuo, Ian H. Sloan, and Henryk Woźniakowski. Periodization strategy may fail in high dimensions. Numerical Algorithms, 46(4):369–391, December 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=4&spage=369.

Kressner:2009:IQA

[KSW09] Daniel Kressner, Christian Schröder, and David S. Watkins. Implicit QR algorithms for palindromic and even eigenvalue problems. Numerical Algorithms, 51(2):209–238, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=2&

spage=209. Tributes to Gene H. Golub, Part II.

Kanwar:2007:MFM

[KT07] V. Kanwar and S. K. Tomar. Modified families of multipoint iterative methods for solving nonlinear equations. Numerical Algorithms, 44(4):381–389, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=4&spage=381.

Kubica:2015:PHT

[Kub15] Bartlomiej Jacek Kubica. Presentation of a highly tuned multithreaded interval solver for underdetermined and welldetermined nonlinear systems. Numerical Algorithms, 70 (4):929–963, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/link.springer.com/article/10.1007/s11075-015-9980-
y;   http://link.springer.com/content/pdf/10.1007/
s11075-015-9980-y.pdf.

Kubica:2023:ADH

[Kub23] Bartłomiej Jacek Kubica. Algorithmic differentiation and hullconsistency enforcing using C++ template meta-programming. Numerical Algorithms, 94(4):1673–1704, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01551-8.

Kuhlman:2013:RIL

[Kuh13] Kristopher L. Kuhlman. Review of inverse Laplace transform algorithms for Laplace-space numerical approaches. Numerical Algorithms, 63(2):339–355, June 2013. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9625-3.

Kulikov:2010:LTE

[Kul10] Gennady Yu. Kulikov. Local theory of extrapolation methods. Numerical Algorithms, 53(2–3):321–342, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=321.

Kunoth:2001:WTF

[Kun01] Angela Kunoth. Wavelet techniques for the fictitiousdomain-Lagrange-multiplier-approach. Numerical Algorithms, 27(3):291–316, July 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.lwwonline.com/content/getfile/5058/34/3/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/34/3/fulltext.pdf.

Kunoth:2005:AWS

[Kun05] Angela Kunoth. Adaptive wavelet schemes for an elliptic control problem with Dirichlet boundary control. Numerical Algorithms, 39(1–3):199–220, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kravanja:2000:CVM

[KV00] Peter Kravanja and Marc Van Barel. Coupled Vandermonde matrices and the superfast computation of Toeplitz determinants. Numerical Algorithms, 24(1–2):99–116, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/7/fulltext.pdf. Computational methods from rational approximation theory (Wilrijk, 1999).

Kandilarov:2004:IIM

[KV04] Juri D. Kandilarov and Lubin G. Vulkov. The immersed interface method for a nonlinear chemical diffusion equation with local sites of reactions. Numerical Algorithms, 36(4):285–307, August 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Krajewski:2007:MMR

[KV07] Wiesław Krajewski and Umberto Viaro. On MIMO model reduction by the weighted equation-error approach. Numerical Algorithms, 44(1):83–98, January 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=1&spage=83.

Kvaerno:2012:SET

[KV12] Anne Kværnø and J. H. Verner. Subquadrature expansions for TSRK methods. Numerical Algorithms, 59(3):487–504, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=3&
spage=487.

Kvasov:2001:ADG

[Kva01] Boris I. Kvasov. Approximation by discrete GB-splines. Numerical Algorithms, 27(2):169–188, June 2001. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp007.lwwonline.com/content/
getfile/5058/35/5/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/35/5/fulltext.pdf.

Kvasov:2014:DWC

[Kva14] Boris I. Kvasov. Discrete weighted cubic splines. Numerical Algorithms, 67(4):863–888, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9830-3.

Kiesel:1996:NKP

[KW96] Harry Kiesel and Jet Wimp. A note on Koornwinder’s polynomials with weight function (1 − x)^{α}(1 + x)^{β} + Mδ(x + 1) + Nδ(x − 1). Numerical Algorithms, 11(1–4):229–241, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Kaagstrom:2000:EPC

[KW00] Bo Kågström and Petter Wiberg. Extracting partial canonical structure for large scale eigenvalue problems. Numerical Algorithms, 24(3):195–237, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Kieffer:2004:GNS

[KW04] Michel Kieffer and Eric Walter. Guaranteed nonlinear state estimator for cooperative systems. Numerical Algorithms, 37(1–4):187–198, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/14/abstract.htm.

Kosari:2025:EHN

[KXD25] Saeed Kosari, Peng Xu, and MohammadHossein Derakhshan. An efficient hybrid numerical approach for solving twodimensional fractional cable model involving time-fractional operator of distributed order with error analysis. Numerical Algorithms, 99(3):1237–1267, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01913-
w.

Kang:2021:EQR

[KXXW21] Hongchao Kang, Chunzhi Xiang, Zhenhua Xu, and Hong Wang. Efficient quadrature rules for the singularly oscillatory Bessel transforms and their error analysis. Numerical Algorithms, 88(3):1493–1521, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01083-
z.

Kielbasinski:2003:NBH

[KZ03] Andrzej Kiełbasiński and Krystyna Zietak. Numerical behaviour of Higham’s scaled method for polar decomposition. Numerical Algorithms, 32(2–4):105–140, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
45/1/fulltext.pdf.

Kuzelewski:2021:FAP

[KZ21] Andrzej Kuzelewski and Eugeniusz Zieniuk. The FMM accelerated PIES with the modified binary tree in solving potential problems for the domains with curvilinear boundaries. Numerical Algorithms, 88(3):1025–1050, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01066-6;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01066-6.pdf.

Kzaz:1997:GQA

[Kza97] M. Kzaz. Gaussian quadrature and acceleration of convergence. Numerical Algorithms, 15(1):75–89, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/30/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
30/5/fulltext.pdf.

Karimi:2021:WRS

[KZS21] Milad Karimi, Fatemeh Zallani, and Khosro Sayevand. Wavelet regularization strategy for the fractional inverse diffusion problem. Numerical Algorithms, 87(4):1679–1705, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01025-1.

LaCruz:2017:SAL

[La 17] William La Cruz. A spectral algorithm for large-scale systems of nonlinear monotone equations. Numerical Algorithms, 76(4): 1109–1130, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liang:2022:EIS

[LA22a] Zhao-Zheng Liang and Owe Axelsson. Exact inverse solution techniques for a class of complex valued block two-by-two linear systems. Numerical Algorithms, 90(1):79–98, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01180-z.

Long:2022:ERN

[LA22b] Dang Quang Long and Dang Quang A. Existence results and numerical method for solving a fourth-order nonlinear integrodifferential equation. Numerical Algorithms, 90(2):563–576, June 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01198-3.

LeGuyader:2005:ULS

[LAG05] Carole Le Guyader, Dominique Apprato, and Christian Gout. Using a level set approach for image segmentation under interpolation conditions. Numerical Algorithms, 39(1–3):221–235, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lobos:2022:MGQ

[LAH22] Claudio Lobos, Cristopher Arenas, and Nancy Hitschfeld. Measuring geometrical quality of different 3D linear element types. Numerical Algorithms, 90(1):423–446, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01193-8.

Lai:1992:FSN

[Lai92] Ming Jun Lai. Fortran subroutines for B-nets of box splines on three- and four-directional meshes. Numerical Algorithms, 2 (1):33–38, February 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lambers:2009:KSS

[Lam09] James V. Lambers. Krylov subspace spectral methods for the time-dependent Schrödinger equation with non-smooth potentials. Numerical Algorithms, 51(2):239–280, June 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=2&

spage=239. Tributes to Gene H. Golub, Part II.

Liang:2018:RSP

[LAN18] Zhao-Zheng Liang, Owe Axelsson, and Maya Neytcheva. A robust structured preconditioner for time-harmonic parabolic optimal control problems. Numerical Algorithms, 79(2):575– 596, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-017-0451-5.pdf.

Laurie:2007:CRK

[Lau07] Dirk Laurie. Calculation of Radau–Kronrod and Lobatto– Kronrod quadrature formulas. Numerical Algorithms, 45(1– 4):139–152, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-007-9105-3.

Lazzaro:1999:BBF

[Laz99] Damiana Lazzaro. Biorthogonal M-band filter construction using the lifting scheme. Numerical Algorithms, 22 (1):53–72, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
21/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/21/5/fulltext.pdf.

Levrie:1993:NTF

[LB93] Paul Levrie and Adhemar Bultheel. A note on Thiele nfractions. Numerical Algorithms, 4(3):225–239, March 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2025:SRB

[LB25] Xue Li and Wei Bian. Smoothing randomized block-coordinate proximal gradient algorithms for nonsmooth nonconvex composite optimization. Numerical Algorithms, 100(1):369–393, September 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01965-y.

Li:2024:SCN

[LBF24] Yang Li, Yanhong Bai, and Minfu Feng. A stabilized Crank–Nicolson virtual element method for the unsteady Navier–Stokes problems with high Reynolds number. Numerical Algorithms, 96(4):1779–1817, August 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01685-9.

Lawrence:2014:SRB

[LC14] Piers W. Lawrence and Robert M. Corless. Stability of rootfinding for barycentric Lagrange interpolants. Numerical Algorithms, 65(3):447–464, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9770-3.

Li:2019:KPM

[LC19] Lu Li and Elena Celledoni. Krylov projection methods for linear Hamiltonian systems. Numerical Algorithms, 81(4):1361– 1378, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2021:CLC

[LC21] Yingzhi Liu and Xiao-Chuan Cai. A central-line coarse preconditioner for Stokes flows in artery-like domains. Numerical Algorithms, 87(1):137–160, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00961-2.

Lyu:2025:RSD

[LC25] Kaiyu Lyu and Hao Cheng. Recovering a space-dependent source term for distributed order time-space fractional diffusion equation. Numerical Algorithms, 99(3):1289–1315, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01916-7.

Liu:2023:VEM

[LCGH23] Wanxiang Liu, Yanping Chen, Qiling Gu, and Yunqing Huang. Virtual element method for nonlinear Sobolev equation on polygonal meshes. Numerical Algorithms, 94(4):1731–1761, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01553-6.

Lin:2020:PEE

[LCH20] Xiuxiu Lin, Yanping Chen, and Yunqing Huang. A posteriori error estimates of hp spectral element methods for optimal control problems with L^{2}-norm state constraint. Numerical Algorithms, 83(3):1145–1169, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2021:GST

[LCHH21] Ya-Jie Liu, Xiang-Ke Chang, Yi He, and Xing-Biao Hu. Generalizations of Shanks transformation and corresponding convergence acceleration algorithms via pfaffians. Numerical Algorithms, 88(4):1733–1756, December 2021. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01092-
y.

Liu:2021:SCF

[LCL21] Xindong Liu, Zili Chen, and Jinxing Liu. On the split common fixed point problem for strict quasi-φ-pseudocontractive mappings in Banach spaces. Numerical Algorithms, 87(3): 1129–1145, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01002-8.

Li:2018:SRB

[LCVL18] Wen Li, Yanmei Chen, Seakweng Vong, and Qilun Luo. Some refined bounds for the perturbation of the orthogonal projection and the generalized inverse. Numerical Algorithms, 79(2): 657–677, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2020:FDA

[LCW20] Wei Liu, Jintao Cui, and Zhifeng Wang. A finite difference approximation of reduced coupled model for slightly compressible Forchheimer fractures in karst aquifer system. Numerical Algorithms, 84(1):133–163, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Luo:2021:NBC

[LCW21] Hezhi Luo, Sikai Chen, and Huixian Wu. A new branchand-cut algorithm for non-convex quadratic programming via alternative direction method and semidefinite relaxation. Numerical Algorithms, 88(2):993–1024, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01065-7.

Li:2023:POD

[LCW23] Qing Li, Huanzhen Chen, and Hong Wang. A proper orthogonal decomposition-compact difference algorithm for plate vibration models. Numerical Algorithms, 94(3):1489–1518, November 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01544-7.

Liu:2025:UPP

[LCW25] Ruishu Liu, Yulin Cao, and Xiaojie Wang. Unconditionally positivity-preserving explicit Euler-type schemes for a generalized Aït-Sahalia model. Numerical Algorithms, 98(2):651–675, February 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01810-2.

Li:2026:NMA

[LCZ26] Lin Li, Zhong Chen, and Biao Zhang. A novel meshless approach based on new fractional spaces for solving nonlinear Riesz space distributed order reaction-diffusion equations with non-smooth solutions and stability analysis. Numerical Algorithms, 101(2):883–914, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02026-8.

Liu:2023:UOE

[LCZZ23] Nan Liu, Yanping Chen, Jiwei Zhang, and Yanmin Zhao. Unconditionally optimal H^{1}-error estimate of a fast nonuniform L2-1_{σ} scheme for nonlinear subdiffusion equations. Numerical Algorithms, 92(3):1655–1677, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01359-y.

Liu:2020:NMT

[LD20] Xing Liu and Weihua Deng. Numerical methods for the two-dimensional Fokker–Planck equation governing the probability density function of the tempered fractional Brownian motion. Numerical Algorithms, 85(1):23–38, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00800-z.

Li:2021:LPF

[LD21] Qingtao Li and Guangzhi Du. Local and parallel finite element methods based on two-grid discretizations for the nonstationary Navier–Stokes equations. Numerical Algorithms, 88 (4):1915–1936, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01100-1.

Lin:2010:SDP

[LDC10] Matthew M. Lin, Bo Dong, and Moody T. Chu. Semi-definite programming techniques for structured quadratic inverse eigenvalue problems. Numerical Algorithms, 53(4):419–437, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=53&
issue=4&spage=419.

Li:2023:FEM

[LDH23] Min Li, Xinjie Dai, and Chengming Huang. Fast Euler– Maruyama method for weakly singular stochastic Volterra integral equations with variable exponent. Numerical Algorithms, 92(4):2433–2455, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01397-6.

Li:2017:NEB

[LDL17] Chaoqian Li, Pingfan Dai, and Yaotang Li. New error bounds for linear complementarity problems of Nekrasov matrices and B-nekrasov matrices. Numerical Algorithms, 74(4):997–1009, April 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0181-0.

Liu:2019:SSO

[LDL^{+}19] Yang Liu, Yanwei Du, Hong Li, Fawang Liu, and Yajun Wang. Some second-order θ schemes combined with finite element method for nonlinear fractional cable equation. Numerical Algorithms, 80(2):533–555, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lee:2016:EGL

[LDN16] Jaechoul Lee, Anthony Dini, and William Negri. An efficient generalized least squares algorithm for periodic trended regression with autoregressive errors. Numerical Algorithms, 71(1): 59–75, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9984-7.

Li:2018:EEP

[LDW18] Jiyong Li, Shuo Deng, and Xianfen Wang. Extended explicit pseudo two-step RKN methods for oscillatory systems y′′ + My = f(y). Numerical Algorithms, 78(3):673–700, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2023:MFM

[LDX23] Xindong Li, Mingyang Du, and Wenwen Xu. A multipoint flux mixed finite element method with mass-conservative characteristic finite element method for incompressible miscible displacement problem. Numerical Algorithms, 93(4):1795–1810, August 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01489-3.

Liang:2024:IMS

[LDZ24] Maolin Liang, Lifang Dai, and Ruijuan Zhao. Iterative methods for solving tensor equations based on exponential acceleration. Numerical Algorithms, 97(1):29–49, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01692-w.

Du:2021:TSO

[lDzS21] Rui lian Du and Zhi zhong Sun. Temporal second-order difference methods for solving multi-term time fractional mixed diffusion and wave equations. Numerical Algorithms, 88(1): 191–226, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01037-x.

leMehaute:1991:TFS

[le 91] A. le Méhauté. Taylorian fields and subdivision algorithms. Numerical Algorithms, 1(3):225–235, September 1991. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

LeFerrand:1992:QCT

[Le 92] Hervé Le Ferrand. The quadratic convergence of the topological ϵ-algorithm for systems of nonlinear equations. Numerical Algorithms, 3(1–4):273–283, December 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

LeVey:1998:SRS

[Le 98] G. Le Vey. Some remarks on solvability and various indices for implicit differential equations. Numerical Algorithms, 19 (1–4):127–145, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
13/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/13/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Ferrand:2019:RIM

[Le 19] Hervé Le Ferrand. The rational iteration method by Georges Lemaître. Numerical Algorithms, 80(1):235–251, January 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lee:1994:FPS

[Lee94] Daeshik Lee. Fast parallel solution of the Poisson equation on irregular domains. Numerical Algorithms, 8(2–4):347–362, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lei:2015:IFM

[Lei15] Yuan Lei. The inexact fixed matrix iteration for solving large linear inequalities in a least squares sense. Numerical Algorithms, 69(1):227–251, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9892-2.

Leong:2021:DQN

[LEK21] Wah June Leong, Sharareh Enshaei, and Sie Long Kek. Diagonal quasi-Newton methods via least change updating principle with weighted Frobenius norm. Numerical Algorithms, 86(3): 1225–1241, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00930-9.

Lenarduzzi:1993:PSN

[Len93] L. Lenarduzzi. Practical selection of neighbourhoods for local regression in the bivariate case. Numerical Algorithms, 5(1–4): 205–213, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Leopold:2003:PRR

[Leo03] Elie Leopold. Perturbed recurrence relations. Numerical Algorithms, 33(1–4):357–366, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
29/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/29/fulltext.pdf.

Leopardi:2007:PWQ

[Leo07a] Paul C. Leopardi. Positive weight quadrature on the sphere and monotonicities of Jacobi polynomials. Numerical Algorithms, 45(1–4):75–87, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=75.

Leopold:2007:PRR

[Leo07b] Elie Leopold. Perturbed recurrence relations. II. The general case. Numerical Algorithms, 44(4):347–366, April 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=44&issue=4&
spage=347.

Leopold:2008:PRR

[Leo08] Elie Leopold. Perturbed recurrence relations. III. The general case—some new applications. Numerical Algorithms, 48(4):383–402, August 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=383.

Luzyanina:2003:CSD

[LER03] T. Luzyanina, K. Engelborghs, and D. Roose. Computing stability of differential equations with bounded distributed delays. Numerical Algorithms, 34(1):41–66, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.kluweronline.
com/content/getfile/5058/47/4/abstract.htm;   http:
//ipsapp008.kluweronline.com/content/getfile/5058/
47/4/fulltext.pdf.

Levesley:1995:CKB

[Lev95] J. Levesley. Convolution kernels based on thin-plate splines. Numerical Algorithms, 10(3–4):401–419, October 1995. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Levesley:2005:WTW

[Lev05] Jeremy Levesley. Where there’s a Will there’s a way – the research of Will Light. Numerical Algorithms, 39(1–3):7–34, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2019:DFI

[LF19] Jinkui Liu and Yuming Feng. A derivative-free iterative method for nonlinear monotone equations with convex constraints. Numerical Algorithms, 82(1):245–262, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See comment [AAFL23].

Lu:2025:FAM

[LF25] Zhenhao Lu and Wenping Fan. A fast algorithm for multiterm time-space fractional diffusion equation with fractional boundary condition. Numerical Algorithms, 98(3):1171–1194, March 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01830-y.

Lui:1995:HMN

[LG95] S. H. Lui and G. H. Golub. Homotopy method for the numerical solution of the eigenvalue problem of self-adjoint partial differential operators. Numerical Algorithms, 10(3–4):363–378, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Livne:2004:SB

[LG04] Oren E. Livne and Gene H. Golub. Scaling by binormalization. Numerical Algorithms, 35(1):97–120, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/49/A/7/abstract.htm.

LeGuyader:2008:GAC

[LG08] Carole Le Guyader and Christian Gout. Geodesic active contour under geometrical conditions: theory and 3D applications. Numerical Algorithms, 48(1–3):105–133, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=105.

Li:2017:MSM

[LG17] Yang Li and Xue-Ping Guo. Multi-step modified Newton–HSS methods for systems of nonlinear equations with positive definite Jacobian matrices. Numerical Algorithms, 75(1):55–80, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2018:AMN

[LG18] Ya-Min Li and Xue-Ping Guo. On the accelerated modified Newton–HSS method for systems of nonlinear equations. Numerical Algorithms, 79(4):1049–1073, December 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2019:EPT

[LG19] Jiyong Li and Yachao Gao. Energy-preserving trigonometrically fitted continuous stage Runge–Kutta–Nyström methods for oscillatory Hamiltonian systems. Numerical Algorithms, 81 (4):1379–1401, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lewanowicz:2000:RRC

[LGA^{+}00] Stanisław Lewanowicz, Eduardo Godoy, Iván Area, André Ronveaux, and Alejandro Zarzo. Recurrence relations for the coefficients of the Fourier series expansions with respect to q-classical orthogonal polynomials. Numerical Algorithms, 23(1):31–50, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
24/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/24/2/fulltext.pdf.

Li:2024:ADG

[LGC24] Rui Li, Yali Gao, and Zhangxin Chen. Adaptive discontinuous Galerkin finite element methods for the Allen–Cahn equation on polygonal meshes. Numerical Algorithms, 95(4):1981– 2014, April 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01635-5.

Li:2025:EMM

[LGHS25] Min Li, Shangjiang Guo, Peng Hu, and Haiyan Song. An Euler–Maruyama method for Caputo–Hadamard fractional stochastic differential equations on exponential meshes and its fast approximation. Numerical Algorithms, 99(2):573–601, June 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01889-7.

Lei:2023:FOS

[LGL23] Ziyi Lei, Siqing Gan, and Jing Liu. First order strong approximation of Ait–Sahalia-type interest rate model with Poisson jumps. Numerical Algorithms, 94(1):93–130, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01494-6.

Long:2025:FTR

[LGM25] Le Dinh Long, Yusuf Gürefe, and B. Parsa Moghaddam. Fractional Tikhonov regularization and error estimation in inverse source problems for biharmonic equations: a priori and a posteriori analysis under deterministic and random perturbations. Numerical Algorithms, 100(1):63–84, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01955-0.

Liu:2011:EAJ

[LGP11] Da-Yan Liu, Olivier Gibaru, and Wilfrid Perruquetti. Error analysis of Jacobi derivative estimators for noisy signals. Numerical Algorithms, 58(1):53–83, September 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=1&spage=53.

Lin:2014:CDF

[LGW14] Hongwei Lin, Yuelin Gao, and Yuping Wang. A continuously differentiable filled function method for global optimization. Numerical Algorithms, 66(3):511–523, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9746-3.

Liu:2025:CAA

[LGZ25] Benqi Liu, Kai Gong, and Liwei Zhang. Convergence analysis of the augmented Lagrangian method for ℓ_{p}-norm cone optimization problems with p ≥ 2. Numerical Algorithms, 99 (3):1203–1236, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01912-x.

Liu:2023:MDM

[LH23] Chengyu Liu and Guanghui Hu. An MP–DWR method for hadaptive finite element methods. Numerical Algorithms, 94(3): 1309–1329, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01536-7.

Liu:2011:PPN

[LHL11] Hongwei Liu, Yakui Huang, and Xiangli Li. Partial projected Newton method for a class of stochastic linear complementarity problems. Numerical Algorithms, 58(4):593–618, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=58&issue=4&
spage=593.

Liu:2024:MPS

[LHlLG24] Xu Liu, Qi Hong, Hong lin Liao, and Yuezheng Gong. A multi-physical structure-preserving method and its analysis for the conservative Allen–Cahn equation with nonlocal constraint. Numerical Algorithms, 97(3):1431–1451, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01757-4.

Li:2020:RTG

[LHM20] Meng Li, Chengming Huang, and Wanyuan Ming. A relaxation-type Galerkin FEM for nonlinear fractional Schrödinger equations. Numerical Algorithms, 83(1):99–124, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2020:SOA

[LHR20] Yi Li, Yanren Hou, and Yao Rong. A second-order artificial compression method for the evolutionary Stokes–Darcy system. Numerical Algorithms, 84(3):1019–1048, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00791-x.

Luo:2013:SEA

[LHW13] Xin Luo, Jin Huang, and Chuan-Long Wang. Splitting extrapolation algorithms for solving the boundary integral equations of anisotropic Darcy’s equation on polygons by mechanical quadrature methods. Numerical Algorithms, 62 (1):27–43, January 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9563-0/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=1&spage=27-43.

Li:2017:GFE

[LHW17] Meng Li, Chengming Huang, and Pengde Wang. Galerkin finite element method for nonlinear fractional Schrödinger equations. Numerical Algorithms, 74(2):499–525, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0160-5.

Li:2010:NPS

[LHZ10] Jiao-Fen Li, Xi-Yan Hu, and Lei Zhang. The nearness problems for symmetric centrosymmetric with a special submatrix constraint. Numerical Algorithms, 55(1):39–57, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=1&
spage=39.

Li:2020:FCN

[LHZ20a] Meng Li, Chengming Huang, and Yongliang Zhao. Fast conservative numerical algorithm for the coupled fractional Klein– Gordon–Schrödinger equation. Numerical Algorithms, 84(3): 1081–1119, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00793-9.

Lin:2020:EKA

[LHZ20b] Chuan Lin, Gabor T. Herman, and Marcelo V. W. Zibetti. Enhancement of the Kaczmarz algorithm with projection adjustment. Numerical Algorithms, 85(2):713–736, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00834-3.

Li:2021:ATV

[LHZ^{+}21] Xin Li, Ting-Zhu Huang, Xi-Le Zhao, Teng-Yu Ji, Yu-Bang Zheng, and Liang-Jian Deng. Adaptive total variation and second-order total variation-based model for low-rank tensor completion. Numerical Algorithms, 86(1):1–24, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00876-y.

Li:1995:SIM

[Li95] Zi Cai Li. Splitting-integrating method for inverse transformation of n-dimensional digital images and patterns. Numerical Algorithms, 9(3–4):181–198, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:1996:ADT

[Li96] Zi-Cai Li. Analysis of discrete techniques for image transformation. Numerical Algorithms, 13(3–4):225–263, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:1997:IIM

[Li97] Zhilin Li. Immersed interface methods for moving interface problems. Numerical Algorithms, 14(4):269–293, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.

com/content/getfile/5058/6/1/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/6/
1/fulltext.pdf.

Livermore:2010:QNO

[LI10] Philip W. Livermore and Glenn R. Ierley. Quasi-L^{p} norm orthogonal Galerkin expansions in sums of Jacobi polynomials. Numerical Algorithms, 54(4):533–569, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=4&spage=533.

Li:2017:TFM

[Li17] Jiyong Li. Trigonometrically fitted multi-step Runge–Kutta methods for solving oscillatory initial value problems. Numerical Algorithms, 76(1):237–258, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lie:2000:FTO

[Lie00] Knut-Andreas Lie. Front tracking for one-dimensional quasilinear hyperbolic equations with variable coefficients. Numerical Algorithms, 24(3):275–298, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Light:1993:TGA

[Lig93] W. A. Light. Techniques for generating approximations via convolution kernels. Numerical Algorithms, 5(1–4):247–261, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Linh:1998:SQA

[Lin98] Vu Hoang Linh. On some questions arising in numerical realization of amplitude-phase methods. Numerical Algorithms, 17(1–2):171–191, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
31/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/31/8/fulltext.pdf.

Lin:2001:PBT

[Lin01] Fu-Rong Lin. Preconditioners for block Toeplitz systems based on circulant preconditioners. Numerical Algorithms, 26(4):365–379, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
32/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/32/5/fulltext.pdf;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/32/
5/abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/32/5/fulltext.pdf.

Linss:2005:SCU

[Lin05] Torsten Linß. Sufficient conditions for uniform convergence on layer-adapted meshes for one-dimensional reaction–diffusion problems. Numerical Algorithms, 40(1):23–32, September 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=40&issue=1&
spage=23.

Linss:2009:AFC

[Lin09] Torsten Linß. Analysis of a FEM for a coupled system of singularly perturbed reaction-diffusion equations. Numerical Algorithms, 50(3):283–291, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=3&spage=283.

Ling:2016:INI

[Lin16] Aifan Ling. An inexact non-interior continuation method for semidefinite programming: convergence analysis and numerical results. Numerical Algorithms, 73(1):219–244, September 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0093-4.

Liseikin:2025:TNA

[Lis25] V. D. Liseikin. Theoretical and numerical analysis of problems with a boundary turning point and a variable diffusion coefficient. Numerical Algorithms, 100(1):339–351, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01963-0.

Liu:2011:TMP

[Liu11] Qunfeng Liu. Two minimal positive bases based direct search conjugate gradient methods for computationally expensive functions. Numerical Algorithms, 58(4):461–474, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=58&issue=4&
spage=461.

Liu:2014:RLM

[Liu14] Tao-Wen Liu. A regularized limited memory BFGS method for nonconvex unconstrained minimization. Numerical Algorithms, 65(2):305–323, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9706-y.

Liu:2021:NES

[Liu21] Zhengguang Liu. Novel energy stable schemes for Swift– Hohenberg model with quadratic-cubic nonlinearity based on the H^{−1}-gradient flow approach. Numerical Algorithms, 87 (2):633–650, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00981-y.

Liang:2011:AAS

[LJ11] Songxin Liang and David J. Jeffrey. An analytical approach for solving nonlinear boundary value problems in finite domains. Numerical Algorithms, 56(1):93–106, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=93.

Liu:2022:CNT

[LJ22] Qiaohua Liu and Zhigang Jia. On condition numbers of the total least squares problem with linear equality constraint. Numerical Algorithms, 90(1):363–385, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01191-w.

Long:2021:MAM

[LJbL21] Guangqing Long, Rong Jie, and Li bin Liu. Multilevel augmentation methods for eigen-problems of compact integral operators. Numerical Algorithms, 88(3):1523–1540, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01084-y.

Liu:2017:FSA

[LJW17] Jun Liu, Yao-Lin Jiang, and Yan Wang. A fast and stable algorithm for linear parabolic partial differential equations. Numerical Algorithms, 75(3):699–729, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2021:WRF

[LJWW21] Jun Liu, Yao-Lin Jiang, Xiao-Long Wang, and Yan Wang. Waveform relaxation for fractional sub-diffusion equations. Numerical Algorithms, 87(4):1445–1478, August 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-01014-4.

Li:2025:NAR

[LJX25] Xiaowen Li, Xiaoying Jiang, and Xiang Xu. Numerical algorithms for recovering a fractional Sturm–Liouville operator based on trace formulae. Numerical Algorithms, 99(3):1601– 1622, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01926-5.

Liu:2024:ISC

[LJY24] Wenli Liu, Jinbao Jian, and Jianghua Yin. An inertial spectral conjugate gradient projection method for constrained nonlinear pseudo-monotone equations. Numerical Algorithms, 97(3): 985–1015, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01736-1.

Lee:2020:DAU

[LK20] Min-Young Lee and Young Ik Kim. The dynamical analysis of a uniparametric family of three-point optimal eighthorder multiple-root finders under the Möbius conjugacy map on the Riemann sphere. Numerical Algorithms, 83(3):1063–1090, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liang:2017:LEE

[LKBF17] X. Liang, A. Q. M. Khaliq, H. Bhatt, and K. M. Furati. The locally extrapolated exponential splitting scheme for multidimensional nonlinear space-fractional Schrödinger equations. Numerical Algorithms, 76(4):939–958, December 2017. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Luan:2021:FMP

[LKK21] Tran Nhat Luan, Tran Thi Khieu, and Tra Quoc Khanh. A filter method with a priori and a posteriori parameter choice for the regularization of Cauchy problems for biharmonic equations. Numerical Algorithms, 86(4):1721–1746, April 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00951-4.

Lukas:2015:PFB

[LKKM15] Dalibor Lukás, Petr Kovár, Tereza Kovárová, and Michal Merta. A parallel fast boundary element method using cyclic graph decompositions. Numerical Algorithms, 70(4):807–824, December 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9974-9.

Li:2023:ICN

[LKQ23] Shuguang Li, Oleg V. Kravchenko, and Kai Qu. On the L^{∞} convergence of a novel fourth-order compact and conservative difference scheme for the generalized Rosenau–KdV– RLW equation. Numerical Algorithms, 94(2):789–816, October 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01520-1.

Lewanowicz:2017:CAR

[LKW17] Stanisław Lewanowicz, Paweł Keller, and Paweł Woźny. Constrained approximation of rational triangular Bézier surfaces by polynomial triangular Bézier surfaces. Numerical Algorithms, 75(1):93–111, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:1993:HAS

[LL93] Kuiyuan Li and Tien-Yien Li. A homotopy algorithm for a symmetric generalized eigenproblem. Numerical Algorithms, 4 (1–2):167–195, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Leriche:2005:FSE

[LL05] E. Leriche and G. Labrosse. Fundamental Stokes eigenmodes in the square: Which expansion is more accurate, Chebyshev or Reid–Harris? Numerical Algorithms, 38(1–3):111–131, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2007:CBU

[LL07] Tao-Wen Liu and Dong-Hui Li. A cautious BFGS update for reduced Hessian SQP. Numerical Algorithms, 44(1):11–28, January 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
44&issue=1&spage=11.

Li:2014:HVD

[LL14] Zhe Li and Qi Liu. A heuristic verification of the degree of the approximate GCD of two univariate polynomials. Numerical Algorithms, 67(2):319–334, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9793-9.

Li:2016:WCD

[LL16] Chaoqian Li and Yaotang Li. Weakly chained diagonally dominant B-matrices and error bounds for linear complementarity problems. Numerical Algorithms, 73(4):985–998, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0125-8.

Liu:2018:EGM

[LL18] Zexian Liu and Hongwei Liu. An efficient gradient method with approximate optimal stepsize for large-scale unconstrained optimization. Numerical Algorithms, 78(1):21–39, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2020:EBL

[LL20a] Jicheng Li and Ge Li. Error bounds for linear complementarity problems of S-QN matrices. Numerical Algorithms, 83(3):935– 955, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2020:RAI

[LL20b] Ji-Chuan Liu and Xiao-Chen Li. Reconstruction algorithms of an inverse source problem for the Helmholtz equation. Numerical Algorithms, 84(3):909–933, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00786-8.

Liu:2020:EMS

[LL20c] Zhengguang Liu and Xiaoli Li. Efficient modified stabilized invariant energy quadratization approaches for phase-field crystal equation. Numerical Algorithms, 85(1):107–132, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00804-9.

Li:2022:HAA

[LL22a] Bin Li and Yuan Lei. Hybrid algorithms with active set prediction for solving linear inequalities in a least squares sense. Numerical Algorithms, 90(3):1327–1356, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01232-4.

Li:2022:EHS

[LL22b] Meng Li and Xianbing Luo. An EMC-HDG scheme for the convection-diffusion equation with random diffusivity. Numerical Algorithms, 90(4):1755–1776, August 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01250-2.

Liu:2022:SSS

[LL22c] Zhengguang Liu and Xiaoli Li. Step-by-step solving schemes based on scalar auxiliary variable and invariant energy quadratization approaches for gradient flows. Numerical Algorithms, 89(1):65–86, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01106-9.

Li:2026:GBK

[LL26a] Ran-Ran Li and Hao Liu. The global block Kaczmarz method using double greedy strategy. Numerical Algorithms, 101(3): 1949–1963, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02069-x.

Liang:2026:CTS

[LL26b] Juan Liang and Yonghui Ling. On the convergence of twostep modified Newton method for nonsymmetric algebraic Riccati equations from transport theory. Numerical Algorithms, 102(2):809–845, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02154-1.

Lai:2021:STF

[LLAL21] Junjiang Lai, Fawang Liu, Vo V. Anh, and Qingxia Liu. A space-time finite element method for solving linear Riesz space fractional partial differential equations. Numerical Algorithms, 88(1):499–520, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01047-9.

Liu:2020:RAG

[LLC20] Li-Bin Liu, Guangqing Long, and Zhongdi Cen. A robust adaptive grid method for a nonlinear singularly perturbed differential equation with integral boundary condition. Numerical Algorithms, 83(2):719–739, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2023:SSF

[LLD23] Xu Li, Yi-Xin Li, and Yan Dou. Shift-splitting fixed point iteration method for solving generalized absolute value equations. Numerical Algorithms, 93(2):695–710, June 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01435-3.

Lin:2023:PNS

[lLHNS23] Xue lei Lin, Xin Huang, Michael K. Ng, and Hai-Wei Sun. A τ-preconditioner for a non-symmetric linear system arising from multi-dimensional Riemann–Liouville fractional diffusion equation. Numerical Algorithms, 92(1):795–813, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01342-7.

Liang:2007:EAG

[lLhYfD07] Mao lin Liang, Chuan hua You, and Li fang Dai. An efficient algorithm for the generalized centro-symmetric solution of matrix equation AXB = C. Numerical Algorithms, 44(2):173–184, February 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=2&spage=173.

Li:2018:NSM

[LLL18] Ming Li, Hongwei Liu, and Zexian Liu. A new subspace minimization conjugate gradient method with nonmonotone line search for unconstrained optimization. Numerical Algorithms, 79(1):195–219, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ling:2022:SCA

[LLL22] Yonghui Ling, Juan Liang, and Weihua Lin. On semilocal convergence analysis for two-step Newton method under generalized Lipschitz conditions in Banach spaces. Numerical Algorithms, 90(2):577–606, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01199-
2.

Li:2026:HIA

[LLL26] Linhao Li, Pengjie Liu, and Bohan Li. A hybrid-inertial accelerated spectral CG projection method with restart mechanism and its application to compressed sensing. Numerical Algorithms, 101(4):2181–2214, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02082-0.

Liu:2017:MGA

[LLLD17] Zisheng Liu, Jicheng Li, Wenbo Li, and Pingfan Dai. A modified greedy analysis pursuit algorithm for the cosparse analysis model. Numerical Algorithms, 74(3):867–887, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0174-z.

Liao:2017:SFD

[lLLVZ17] Hong lin Liao, Pin Lyu, Seakweng Vong, and Ying Zhao. Stability of fully discrete schemes with interpolation-type fractional formulas for distributed-order subdiffusion equations. Numerical Algorithms, 75(4):845–878, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2021:SOF

[LlLZ21] Xin Li, Hong lin Liao, and Luming Zhang. A second-order fast compact scheme with unequal time-steps for subdiffusion problems. Numerical Algorithms, 86(3):1011–1039, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00920-x.

Li:2025:BPS

[LLM25] Shu-Jiao Li, Kang-Ya Lu, and Cun-Qiang Miao. Banded preconditioning with shift compensation for solving discrete Riesz space-fractional diffusion equations. Numerical Algorithms, 98 (2):839–876, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01817-9.

Li:2017:PCS

[LLQ17] Yaotang Li, Qilong Liu, and Liqun Qi. Programmable criteria for strong H-tensors. Numerical Algorithms, 74(1):199–221, January 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0145-4.

Li:2011:FSM

[LLS11] Xiangli Li, Hongwei Liu, and Xiaojun Sun. Feasible smooth method based on Barzilai–Borwein method for stochastic linear complementarity problem. Numerical Algorithms, 57(2):207–215, June 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=2&spage=207.

Li:2024:RBK

[LLWC24] Chaoqian Li, Yonghe Liu, Fengsheng Wu, and Maolin Che. Randomized block Krylov subspace algorithms for low-rank quaternion matrix approximations. Numerical Algorithms, 96 (2):687–717, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01662-2.

Li:2020:STF

[LLX20] Binjie Li, Hao Luo, and Xiaoping Xie. A space–time finite element method for fractional wave problems. Numerical Algorithms, 85(3):1095–1121, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00857-
w.

Luo:2022:CNM

[lLXhL22] Xin long Luo, Hang Xiao, and Jia hui Lv. Continuation Newton methods with the residual trust-region time-stepping scheme for nonlinear equations. Numerical Algorithms, 89(1): 223–247, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01112-x.

Luo:2024:CNM

[lLXZ24] Xin long Luo, Hang Xiao, and Sen Zhang. Continuation Newton methods with deflation techniques for global optimization problems. Numerical Algorithms, 97(4):1715–1790, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01768-1.

Li:2022:GMB

[LLY22] Rui Li, Zhi-Lin Li, and Jun-Feng Yin. A generalized modulusbased Newton method for solving a class of non-linear complementarity problems with P -matrices. Numerical Algorithms, 89(2):839–853, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01136-3.

Li:2025:CGM

[LLY25] Yuping Li, Hui Liang, and Huifang Yuan. On the convergence of Galerkin methods for auto-convolution Volterra integrodifferential equations. Numerical Algorithms, 99(1):183–205, May 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01874-0.

Lang:2026:TSA

[LLY26] Yanhua Lang, Xing Liu, and Huizi Yang. Threshold stability analysis of an unconditionally positivity-preserving method for a nonlinear reaction–diffusion–advection algae– mussel model under Danckwerts boundary conditions. Numerical Algorithms, 101(4):2737–2765, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02108-7.

Li:1994:AGS

[LLZ94] Kuiyuan Li, Tien-Yien Li, and Zhong Gang Zeng. An algorithm for the generalized symmetric tridiagonal eigenvalue problem. Numerical Algorithms, 8(2–4):269–291, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lee:2018:RAP

[LLZ18] Tsung-Lin Lee, Tien-Yien Li, and Zhonggang Zeng. RankRev: a Matlab package for computing the numerical rank and updating/downdating. Numerical Algorithms, 77(2):559–576, February 2018. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-017-0328-7.

Lawton:1997:CCQ

[LM97] Wayne Lawton and Charles A. Micchelli. Construction of conjugate quadrature filters with specified zeros. Numerical Algorithms, 14(4):383–399, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
6/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/6/8/fulltext.pdf.

LeCalvez:1999:IRD

[LM99] C. Le Calvez and B. Molina. Implicitly restarted and deflated GMRES. Numerical Algorithms, 21(1–4):261–285, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/16/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
16/fulltext.pdf.

Li:2000:ARB

[LM00] Xin Li and Charles A. Micchelli. Approximation by radial bases and neural networks. Numerical Algorithms, 25 (1–4):241–262, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/18/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/18/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Lamberti:2001:SPF

[LM01] P. Lamberti and C. Manni. Shape-preserving C^{2} functional interpolation via parametric cubics. Numerical Algorithms, 28(1–4):229–254, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.lwwonline.com/content/getfile/5058/37/
14/abstract.htm;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/14/fulltext.pdf.

Lavor:2004:FTM

[LM04] Carlile Lavor and Nelson Maculan. A function to test methods applied to global minimization of potential energy of molecules. Numerical Algorithms, 35(2–4):287–300, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/6/abstract.htm.

Lorch:2008:MSR

[LM08] Lee Lorch and Martin E. Muldoon. Monotonic sequences related to zeros of Bessel functions. Numerical Algorithms, 49(1–4):221–233, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=221.

Lamour:2011:NAI

[LM11] René Lamour and Dagmar Monett. A new algorithm for index determination in DAEs using algorithmic differentiation. Numerical Algorithms, 58(2):261–292, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=2&spage=261.

Loureiro:2012:PSA

[LM12] Ana Filipa Loureiro and Pascal Maroni. Polynomial sequences associated with the classical linear functionals. Numerical Algorithms, 60(2):297–314, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=2&spage=297.

Liu:2013:NTR

[LM13] Jinghui Liu and Changfeng Ma. A nonmonotone trust region method with new inexact line search for unconstrained optimization. Numerical Algorithms, 64(1):1–20, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9652-0.

Liang:2014:LTS

[LM14] Songxin Liang and Junchi Ma. Laplace transform for the solution of higher order deformation equations arising in the homotopy analysis method. Numerical Algorithms, 67(1):49– 57, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9773-0.

Lu:2015:SNQ

[LM15] Dawei Lu and Congxu Ma. Some new quicker continued fraction approximations for the gamma function related to the Nemes’ formula. Numerical Algorithms, 70(4):825–833, December 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9975-8.

Li:2017:PUL

[LM17a] Jing-Tao Li and Chang-Feng Ma. The parameterized upper and lower triangular splitting methods for saddle point problems. Numerical Algorithms, 76(2):413–425, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2017:ASS

[LM17b] Wei Liu and Xuerong Mao. Almost sure stability of the Euler– Maruyama method with random variable stepsize for stochastic differential equations. Numerical Algorithms, 74(2):573– 592, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0162-3.

Li:2019:EPR

[LM19] Cheng-Liang Li and Chang-Feng Ma. Efficient parameterized rotated shift-splitting preconditioner for a class of complex symmetric linear systems. Numerical Algorithms, 80(2): 337–354, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2021:BCB

[LM21] Ling Liu and Junjie Ma. Block collocation boundary value solutions of the first-kind Volterra integral equations. Numerical Algorithms, 86(2):911–932, February 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-00917-6.

Lv:2025:PIF

[LM25] Xin-Mei Lv and Shu-Xin Miao. Preconditioned inexact fixed point iteration method for solving tensor absolute value equation. Numerical Algorithms, 100(3):1069–1084, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01980-z.

Laudadio:2025:CAS

[LMD25] T. Laudadio, N. Mastronardi, and P. Van Dooren. Computational aspects of simultaneous Gaussian quadrature. Numerical Algorithms, 100(2):621–643, October 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01785-0.

Laurent:1997:CSS

[LMM97] Pierre-Jean Laurent, Marie-Laurence Mazure, and Voichiţa Teodora Maxim. Chebyshev splines and shape parameters. Numerical Algorithms, 15(3–4):373–383, January 1997. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lopez-Moreno:2005:AEM

[LMMD05] Antonio-Jesús López-Moreno and Francisco-Javier Muñoz-Delgado. Asymptotic expansion of multivariate Kantorovich type operators. Numerical Algorithms, 39(1–3):237–252, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2011:FOI

[LMMH11] Xiaowu Li, Chunlai Mu, Jinwen Ma, and Linke Hou. Fifthorder iterative method for finding multiple roots of nonlinear equations. Numerical Algorithms, 57(3):389–398, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=3&
spage=389.

Li:2019:IFD

[LMUZ19] Jian Li, Mingzhi Mao, Frank Uhlig, and Yunong Zhang. A 5instant finite difference formula to find discrete time-varying generalized matrix inverses, matrix inverses, and scalar reciprocals. Numerical Algorithms, 81(2):609–629, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lemmerling:2000:FAS

[LMV00] Philippe Lemmerling, Nicola Mastronardi, and Sabine Van Huffel. Fast algorithm for solving the Hankel/Toeplitz Structured Total Least Squares problem. Numerical Algorithms, 23(4):371–392, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
26/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/26/3/fulltext.pdf.

Laudadio:2023:CGQ

[LMV23] Teresa Laudadio, Nicola Mastronardi, and Paul Van Dooren. Computing Gaussian quadrature rules with high relative accuracy. Numerical Algorithms, 92(1):767–793, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01297-9.

Laudadio:2024:CMM

[LMV24] Teresa Laudadio, Nicola Mastronardi, and Paul Van Dooren. On computing modified moments for half-range Hermite weights. Numerical Algorithms, 95(3):1435–1457, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01615-9.

Laudadio:2025:CZL

[LMV25] T. Laudadio, N. Mastronardi, and P. Van Dooren. On computing the zeros of Laguerre–Sobolev polynomials. Numerical Algorithms, 100(4):1507–1526, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02021-z.

Levin:1995:CPI

[LN95] David Levin and Edmond Nadler. Convexity preserving interpolation by algebraic curves and surfaces. Numerical Algorithms, 9(1–2):113–139, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Lin:2010:IPT

[LN10] Fu-Rong Lin and Michael K. Ng. Inverse product Toeplitz preconditioners for non-Hermitian Toeplitz systems. Numerical Algorithms, 54(2):279–295, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=2&spage=279.

Liu:2022:PIV

[LN22] Xin Liu and Yufeng Nie. Pressure-independent velocity error estimates for (Navier–)Stokes nonconforming virtual element discretization with divergence free. Numerical Algorithms, 90(2):477–506, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01195-6.

Liesen:2023:CLC

[LNS23] Jörg Liesen, Mohamed M. S. Nasser, and Olivier Sète. Computing the logarithmic capacity of compact sets having (infinitely) many components with the charge simulation method. Numerical Algorithms, 93(2):581–614, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01428-2.

Lo:1997:LPT

[Lo97] Martin W. Lo. Libration point trajectory design. Numerical Algorithms, 14(1–3):153–164, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/9/fulltext.pdf. Dynamical

numerical analysis (Atlanta, GA, 1995).

Loczi:2018:EOV

[Lóc18] Lajos Lóczi. Exact optimal values of step-size coefficients for boundedness of linear multistep methods. Numerical Algorithms, 77(4):1093–1116, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Loczi:2020:OSS

[Lóc20] Lajos Lóczi. Optimal subsets in the stability regions of multistep methods. Numerical Algorithms, 84(2):679–715, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-019-00775-x;   http://link.springer.
com/content/pdf/10.1007/s11075-019-00775-x.pdf.

Lohi:2022:SIH

[Loh22] Jonni Lohi. Systematic implementation of higher order Whitney forms in methods based on discrete exterior calculus. Numerical Algorithms, 91(3):1261–1285, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01301-2.

Lorentzen:1995:CLP

[Lor95] Lisa Lorentzen. Computation of limit periodic continued fractions. A survey. Numerical Algorithms, 10(1–2):69–111, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Lorin:2019:DDM

[Lor19] Emmanuel Lorin. Domain decomposition method for the Nbody time-independent and time-dependent Schrödinger equations. Numerical Algorithms, 81(2):655–694, June 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lowitzsch:2005:DTM

[Low05] Svenja Lowitzsch. A density theorem for matrix-valued radial basis functions. Numerical Algorithms, 39(1–3):253–256, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Landi:2008:PNC

[LP08] G. Landi and E. Loli Piccolomini. A projected Newton–CG method for nonnegative astronomical image deblurring. Numerical Algorithms, 48(4):279–300, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=279.

Landi:2012:INP

[LP12] Germana Landi and Elena Loli Piccolomini. An improved Newton projection method for nonnegative deblurring of Poisson-corrupted images with Tikhonov regularization. Numerical Algorithms, 60(1):169–188, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=1&spage=169.

Landi:2013:NMS

[LP13] Germana Landi and Elena Loli Piccolomini. NPTool: a Matlab software for nonnegative image restoration with Newton projection methods. Numerical Algorithms, 62(3):487–504, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9602-x/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=487-504.

Lerkchaiyaphum:2018:IAS

[LP18] Kritsada Lerkchaiyaphum and Withun Phuengrattana. Iterative approaches to solving convex minimization problems and fixed point problems in complete CAT(0) spaces. Numerical Algorithms, 77(3):727–740, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2020:PRP

[LP20] Jun Liu and John W. Pearson. Parameter-robust preconditioning for the optimal control of the wave equation. Numerical Algorithms, 83(3):1171–1203, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2016:IIN

[LPGL16] Chaoqian Li, Hui Pei, Aning Gao, and Yaotang Li. Improvements on the infinity norm bound for the inverse of Nekrasov matrices. Numerical Algorithms, 71(3):613–630, March 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0012-8.

Lizarte:2021:RPC

[LPP21] Fátima Lizarte, Teresa E. Pérez, and Miguel A. Piñar. The radial part of a class of Sobolev polynomials on the unit ball. Numerical Algorithms, 87(4):1369–1389, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01011-7.

Li:2024:SHO

[LPRÁ24] Zhilin Li, Kejia Pan, and Juan Ruiz-Álvarez. Stable high order FD methods for interface and internal layer problems based on non-matching grids. Numerical Algorithms, 96(4):1647–1674, August 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01680-0.

Laskari:2003:EOG

[LPV03] E. C. Laskari, K. E. Parsopoulos, and M. N. Vrahatis. Evolutionary operators in global optimization with dynamic search trajectories. Numerical Algorithms, 34(2– 4):393–403, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/48/
28/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/28/fulltext.pdf.

Lv:2019:IBM

[LPXX19] Jian Lv, Li-Ping Pang, Na Xu, and Ze-Hao Xiao. An infeasible bundle method for nonconvex constrained optimization with application to semi-infinite programming problems. Numerical Algorithms, 80(2):397–427, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lin:2016:STG

[LQ16] Qiu Lin and Ruisheng Qi. Stancu type generalization of the q-Phillips operators. Numerical Algorithms, 72(1):181–193, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0040-4.

Liu:2020:SCE

[LQ20] Liya Liu and Xiaolong Qin. Strong convergence of an extragradient-like algorithm involving pseudo-monotone mappings. Numerical Algorithms, 83(4):1577–1590, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2026:RLR

[LQ26] Wei Li and Gangrong Qu. Robust low-rank matrix completion based on regionally weighted truncated nuclear norm regularization. Numerical Algorithms, 102(1):1–23, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02119-4.

Lun:2014:IZJ

[LR14] Yen Chi Lun and Fernando Rodrigo Rafaeli. Inequalities for zeros of Jacobi polynomials via Sturm’s theorem: Gautschi’s conjectures. Numerical Algorithms, 67(3):549–563, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9807-7.

Li:2018:TTS

[LR18] Xiaoli Li and Hongxing Rui. Two temporal second-order H^{1}- galerkin mixed finite element schemes for distributed-order fractional sub-diffusion equations. Numerical Algorithms, 79 (4):1107–1130, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2019:FCB

[LRC19] Xiaoli Li, Hongxing Rui, and Shuangshuang Chen. A fully conservative block-centered finite difference method for simulating Darcy–Forchheimer compressible wormhole propagation. Numerical Algorithms, 82(2):451–478, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Luengo:2002:PSG

[LRGH02] F. Luengo, M. Raydan, W. Glunt, and T. L. Hayden. Preconditioned spectral gradient method. Numerical Algorithms, 30(3–4):241–258, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
42/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/42/2/fulltext.pdf.

Li:2019:TAD

[LRL19] Xiaoli Li, Hongxing Rui, and Zhengguang Liu. Two alternating direction implicit spectral methods for two-dimensional distributed-order differential equation. Numerical Algorithms, 82(1):321–347, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Linh:2022:ATV

[LRL22] Ha Manh Linh, Simeon Reich, and Nguyen Phuong Lan. Analysis of two variants of an inertial projection algorithm for finding the minimum-norm solutions of variational inequality and fixed point problems. Numerical Algorithms, 89(4):1695– 1721, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01169-8.

Lotfi:2016:PAE

[LRM16] Taher Lotfi, Mehdi Rashidi, and Katauoun Mahdiani. A posteriori analysis: error estimation for the eighth order boundary value problems in reproducing kernel space. Numerical Algorithms, 73(2):391–406, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0100-4.

Lebtahi:2019:FRG

[LRT19] Leila Lebtahi, Óscar Romero, and Néstor Thome. Further results on generalized centro-invertible matrices. Numerical Algorithms, 80(4):1309–1328, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2025:INB

[LRW25] Qin Li, Wenwen Ran, and Feng Wang. Infinity norm bounds for the inverse of quasi-SDD_{k} matrices with applications. Numerical Algorithms, 99(4):2173–2197, August 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01949-y.

Liu:2018:DEA

[LRY18] Yanzhi Liu, Jason Roberts, and Yubin Yan. Detailed error analysis for a fractional Adams method with graded meshes. Numerical Algorithms, 78(4):1195–1216, August 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0419-5.pdf.

Levin:2026:WEN

[LRY26] David Levin, José M. Ramón, and Dionisio F. Yáñez. Weighted essentially non-oscillatory Shepard method. Numerical Algorithms, 102(1):593–609, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02141-
6.

Linss:2012:ASP

[LRZ12] Torsten Linß, Goran Radojev, and Helena Zarin. Approximation of singularly perturbed reaction-diffusion problems by quadratic C^{1}-splines. Numerical Algorithms, 61(1):35–55, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=1&spage=35.

Lavallee:2003:CPS

[LS03a] P.-F. Lavallée and M. Sadkane. Computation of pseudospectra by spectral dichotomy methods in a parallel environment. Numerical Algorithms, 33(1–4):343–355, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/28/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
28/fulltext.pdf.

Little:2003:BPS

[LS03b] Leigh Little and Yousef Saad. Block preconditioners for saddle point problems. Numerical Algorithms, 33(1– 4):367–379, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
30/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/30/fulltext.pdf.

Liu:2007:IIP

[LS07] Zhongyi Liu and Wenyu Sun. An infeasible interior-point algorithm with full-Newton step for linear optimization. Numerical Algorithms, 46(2):173–188, October 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=2&spage=173.

Luo:2011:STL

[LS11] Xiaolin Luo and Pavel V. Shevchenko. A short tale of long tail integration. Numerical Algorithms, 56(4):577–590, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=577.

Laayouni:2014:PAO

[LS14] Lahcen Laayouni and Daniel B. Szyld. On the performance of the algebraic optimized Schwarz methods with applications. Numerical Algorithms, 67(4):889–916, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9831-2.

Li:2015:VEB

[LS15a] Zhe Li and Haifeng Sang. Verified error bounds for singular solutions of nonlinear systems. Numerical Algorithms, 70(2): 309–331, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9948-3.

Lin:2015:BFE

[LS15b] Runchang Lin and Martin Stynes. A balanced finite element method for a system of singularly perturbed reaction-diffusion two-point boundary value problems. Numerical Algorithms, 70(4):691–707, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-9969-6.

Liu:2020:SAT

[LS20] Qian Liu and Dongyang Shi. Superconvergence analysis of a two-grid method for an energy-stable Ciarlet–Raviart type scheme of Cahn–Hilliard equation. Numerical Algorithms, 85(2):607–622, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00829-0.

Labovsky:2025:NTF

[LS25a] Alexander E. Labovsky and Kyle J. Schwiebert. Note on two formulations of Crank–Nicolson method for Navier–Stokes equations. Numerical Algorithms, 100(3):1145–1156, November 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01983-w.

Luo:2025:ULA

[LS25b] Yusong Luo and Weiping Shen. An Ulm-like algorithm for generalized inverse eigenvalue problems. Numerical Algorithms, 98 (3):1611–1641, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01845-5.

Lotfi:2015:CST

[LSG15] Taher Lotfi, Fazlollah Soleymani, and Mohammad Ghorbanzadeh. On the construction of some tri-parametric iterative methods with memory. Numerical Algorithms, 70(4):835–845, December 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9976-7.

Li:2025:QAS

[LSL25] Xinying Li, Caiqin Song, and Hongjun Liu. QMGI algorithm for solving quaternion equation and its application in color image encryption. Numerical Algorithms, 99(3):1393– 1428, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01920-x.

Lu:2016:QCF

[LSM16] Dawei Lu, Lixin Song, and Congxu Ma. A quicker continued fraction approximation of the gamma function related to the Windschitl’s formula. Numerical Algorithms, 72(4):865–874, August 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0070-y.

Lotfi:2015:NCT

[LSSS15] Taher Lotfi, Somayeh Sharifi, Mehdi Salimi, and Stefan Siegmund. A new class of three-point methods with optimal convergence order eight and its dynamics. Numerical Algorithms, 68(2):261–288, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9843-y.

Liu:2016:GCS

[LSW16] Wenjie Liu, Jiebao Sun, and Boying Wu. Galerkin–Chebyshev spectral method and block boundary value methods for twodimensional semilinear parabolic equations. Numerical Algorithms, 71(2):437–455, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0002-x.

Li:2010:DFA

[LSX10] Huiyuan Li, Jiachang Sun, and Yuan Xu. Discrete Fourier analysis with lattices on planar domains. Numerical Algorithms, 55(2–3):279–300, November 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=279.

Liu:2023:FTT

[LSY^{+}23] Pengjie Liu, Hu Shao, Zihang Yuan, Xiaoyu Wu, and Tianlei Zheng. A family of three-term conjugate gradient projection methods with a restart procedure and their relaxed-inertial extensions for the constrained nonlinear pseudo-monotone equations with applications. Numerical Algorithms, 94(3):1055– 1083, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01527-8.

Li:2019:ITD

[LSZW19] Y. S. Li, L. L. Sun, Z. Q. Zhang, and T. Wei. Identification of the time-dependent source term in a multi-term time-fractional diffusion equation. Numerical Algorithms, 82(4):1279–1301, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Linh:2020:CCH

[LT20] Vu Hoang Linh and Nguyen Duy Truong. On convergence of continuous half-explicit Runge–Kutta methods for a class of delay differential-algebraic equations. Numerical Algorithms, 85(1):277–303, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00813-8.

Liu:2024:SSG

[LT24] Hailiang Liu and Xuping Tian. SGEM: stochastic gradient with energy and momentum. Numerical Algorithms, 95 (4):1583–1610, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01621-x.

Lima:2010:FES

[LTFL10] Pedro Miguel Lima, M. Filomena Teodoro, Neville J. Ford, and Patricia M. Lumb. Finite element solution of a linear mixed-type functional differential equation. Numerical Algorithms, 55(2–3):301–320, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=301.

Livieris:2018:DHC

[LTP18] Ioannis E. Livieris, Vassilis Tampakas, and Panagiotis Pintelas. A descent hybrid conjugate gradient method based on the memoryless BFGS update. Numerical Algorithms, 79(4): 1169–1185, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lu:2015:SRC

[Lu15] Chengbo Lu. Some results on certain generalized circulant matrices. Numerical Algorithms, 68(3):467–479, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9855-7.

Lucet:1997:FTF

[Luc97] Yves Lucet. Faster than the fast Legendre transform, the linear-time Legendre transform. Numerical Algorithms, 16 (2):171–185, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/6/fulltext.pdf.

Lucet:2006:FME

[Luc06] Yves Lucet. Fast Moreau envelope computation I: numerical algorithms. Numerical Algorithms, 43(3):235–249, November 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=43&issue=3&
spage=235.

Lui:2002:SAM

[Lui02] S. H. Lui. On Schwarz alternating methods for the subsonic full potential equation. Numerical Algorithms, 30(1):59–69, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/40/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
40/5/fulltext.pdf.

Lund:2023:ARB

[Lun23] Kathryn Lund. Adaptively restarted block Krylov subspace methods with low-synchronization skeletons. Numerical Algorithms, 93(2):731–764, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01437-
1.

Lemmerling:2001:AST

[LV01] Philippe Lemmerling and Sabine Van Huffel. Analysis of the structured total least squares problem for Hankel/ Toeplitz matrices. Numerical Algorithms, 27(1):89–114, May 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/33/5/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/33/
5/fulltext.pdf.

Le:2015:ADN

[LV15] Thanh Hieu Le and Marc Van Barel. An algorithm for decomposing a non-negative polynomial as a sum of squares of rational functions. Numerical Algorithms, 69(2):397–413, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9903-3.

Lyu:2018:LSO

[LV18] Pin Lyu and Seakweng Vong. A linearized second-order scheme for nonlinear time fractional Klein–Gordon type equations. Numerical Algorithms, 78(2):485–511, June 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lyu:2021:FLN

[LV21] Pin Lyu and Seakweng Vong. A fast linearized numerical method for nonlinear time-fractional diffusion equations. Numerical Algorithms, 87(1):381–408, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00971-0.

Lee:1995:TNO

[LW95] David Lee and Henryk Woźniakowski. Testing nonlinear operators. Numerical Algorithms, 9(3–4):319–342, ???? 1995.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lewanowicz:2004:RRC

[LW04] Stanisław Lewanowicz and Paweł Woźny. Recurrence relations for the coefficients in series expansions with respect to semi-classical orthogonal polynomials. Numerical Algorithms, 35(1):61–79, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/49/A/4/
abstract.htm.

Lin:2012:BPB

[LW12] Fu-Rong Lin and Chi-Xi Wang. BTTB preconditioners for BTTB systems. Numerical Algorithms, 60(1):153–167, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=1&
spage=153.

Li:2013:AFT

[LW13] Jiyong Li and Xinyuan Wu. Adapted Falkner-type methods solving oscillatory second-order differential equations. Numerical Algorithms, 62(3):355–381, March 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9583-9/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=355-381.

Li:2014:EAE

[LW14] Jiyong Li and Xinyuan Wu. Error analysis of explicit TSERKN methods for highly oscillatory systems. Numerical Algorithms, 65(3):465–483, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9761-4.

Li:2016:MSH

[LW16] Jiyong Li and Xianfen Wang. Multi-step hybrid methods for special second-order differential equations y′′(t) = f(t, y(t)). Numerical Algorithms, 73(3):711–733, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0114-y.

Liu:2017:TGD

[LW17] Huipo Liu and Shuanghu Wang. A two-grid discretization scheme for optimal control problems of elliptic equations. Numerical Algorithms, 74(3):699–716, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0168-x.

LeBorne:2020:MIS

[LW20] Sabine Le Borne and Michael Wende. Multilevel interpolation of scattered data using H-matrices. Numerical Algorithms, 85(4):1175–1193, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00860-1.

Liu:2022:USO

[LW22] Wenjie Liu and Boying Wu. Unconditional stability and optimal error estimates of a Crank–Nicolson Legendre–Galerkin method for the two-dimensional second-order wave equation. Numerical Algorithms, 90(1):137–158, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01182-x.

Lu:2025:HMM

[LW25] Peipei Lu and Juan Wang. Homogeneous multigrid method for hybridizable interior penalty method. Numerical Algorithms, 99(4):2011–2038, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01942-
5.

Lewanowicz:2008:MGB

[LWAG08] Stanisław Lewanowicz, Pawel Woźny, Iván Area, and Eduardo Godoy. Multivariate generalized Bernstein polynomials: identities for orthogonal polynomials of two variables. Numerical Algorithms, 49(1–4):199–220, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=199.

Lux:2021:ISH

[LWC^{+}21] Thomas C. H. Lux, Layne T. Watson, Tyler H. Chang, Yili Hong, and Kirk Cameron. Interpolation of sparse high-dimensional data. Numerical Algorithms, 88(1):281–313, September 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01040-2.

Lei:2023:FDS

[LWD23] Su Lei, Yanyan Wang, and Rui Du. A finite difference scheme for the two-dimensional Gray–Scott equation with fractional Laplacian. Numerical Algorithms, 94(3):1185–1215, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01532-x.

Li:2018:VEB

[LWG18] Zhe Li, Baocheng Wan, and Ruimei Gao. Verified error bounds for eigenvalues of geometric multiplicity q and corresponding invariant subspaces. Numerical Algorithms, 77(2):631–652, February 2018. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-017-0332-y.

Lin:2021:CNW

[LWJ21] Fu-Rong Lin, Qiu-Ya Wang, and Xiao-Qing Jin. Crank– Nicolson-weighted-shifted–Grünwald-difference schemes for space Riesz variable-order fractional diffusion equations. Numerical Algorithms, 87(2):601–631, June 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-00980-z.

Lewanowicz:2012:PAR

[LWK12] Stanisław Lewanowicz, Pawel Woźny, and Pawel Keller. Polynomial approximation of rational Bézier curves with constraints. Numerical Algorithms, 59(4):607–622, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=4&
spage=607.

Li:2019:NPE

[LWLT19] T. Li, Y. Wang, F. Liu, and I. Turner. Novel parameter estimation techniques for a multi-term fractional dynamical epidemic model of dengue fever. Numerical Algorithms, 82(4): 1467–1495, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2024:CAH

[LWLW24] Dandan Li, Songhua Wang, Yong Li, and Jiaqi Wu. A convergence analysis of hybrid gradient projection algorithm for constrained nonlinear equations with applications in compressed sensing. Numerical Algorithms, 95(3):1325–1345, March 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01610-0.

Li:2010:NAG

[LWM10] Jianling Li, Peng Wang, and Lin Ma. A new algorithm for the general quadratic programming problems with box constraints. Numerical Algorithms, 55(1):79–85, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=1&spage=79.

Long:2013:FMK

[LWN13] Guangqing Long, Weifen Wu, and Gnaneshwar Nelakanti. A fast multiscale Kantorovich method for weakly singular integral equations. Numerical Algorithms, 63(1):49–63, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9610-x.

Liu:2015:CGM

[LWQR15] Hao Liu, Haijun Wang, Xiaoyan Qian, and Feng Rao. A conjugate gradient method with sufficient descent property. Numerical Algorithms, 70(2):269–286, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9946-5.

Liu:2018:EPP

[LWS18] Kai Liu, Xinyuan Wu, and Wei Shi. Extended phase properties and stability analysis of RKN-type integrators for solving general oscillatory second-order initial value problems. Numerical Algorithms, 77(1):37–56, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2013:LSS

[LWwCL13] Tiexiang Li, Peter Chang-Yi Weng, Eric King wah Chu, and Wen-Wei Lin. Large-scale Stein and Lyapunov equations, Smith method, and applications. Numerical Algorithms, 63(4): 727–752, August 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9650-2.

Li:2026:AFE

[LWY26] Hengguang Li, Charuka D. Wickramasinghe, and Peimeng Yin. Analysis of a C^{0} finite element method for the biharmonic problem with Dirichlet boundary conditions. Numerical Algorithms, 101(3):1739–1784, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02062-4.

Li:2018:EAD

[LWZ18] Jingshi Li, Xiaoshen Wang, and Kai Zhang. An efficient alternating direction method of multipliers for optimal control problems constrained by random Helmholtz equations. Numerical Algorithms, 78(1):161–191, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Luo:2021:CDA

[LWZ21] Jianfeng Luo, Xiaozhou Wang, and Yi Zhao. Convergence of discrete approximation for differential linear stochastic complementarity systems. Numerical Algorithms, 87(1):223–262, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00965-y.

Liu:2023:SOF

[LWZ23] Jianfeng Liu, Tingchun Wang, and Teng Zhang. A secondorder finite difference scheme for the multi-dimensional nonlinear time-fractional Schrödinger equation. Numerical Algorithms, 92(2):1153–1182, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01335-
6.

Li:2026:EAO

[LWZ26] Dongping Li, Xue Wang, and Xiuying Zhang. Efficient algorithm for the oscillatory matrix functions. Numerical Algorithms, 102(2):1087–1110, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02171-
0.

Liu:2017:FMM

[LX17] Guidong Liu and Shuhuang Xiang. Fast multipole methods for approximating a function from sampling values. Numerical Algorithms, 76(3):727–743, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liang:2023:KIU

[LX23] Jiangli Liang and Shuhuang Xiang. A kernel-independent uniform fast multipole method based on barycentric rational interpolation. Numerical Algorithms, 93(4):1595–1611, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01481-x.

Liu:2024:PVS

[LX24] Yuncheng Liu and Fuquan Xia. Proximal variable smoothing method for three-composite nonconvex nonsmooth minimization with a linear operator. Numerical Algorithms, 96(1):237– 266, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01645-3.

Lv:2020:IBM

[LXP20] Jian Lv, Ze-Hao Xiao, and Li-Ping Pang. An incremental bundle method for portfolio selection problem under second-order stochastic dominance. Numerical Algorithms, 85(2):653–681, October 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00831-6.

Lu:2015:FSM

[LXQ15] Ling Lu, Wei Xu, and Sanzheng Qiao. A fast SVD for multilevel block Hankel matrices with minimal memory storage. Numerical Algorithms, 69(4):875–891, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9930-0.

Li:2023:SPS

[LXX23] Shiru Li, Yong Xia, and Zi Xu. Simultaneous perturbation stochastic approximation: towards one-measurement per iteration. Numerical Algorithms, 94(3):1085–1101, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01528-7.

Liang:2021:PTA

[LXZZ21] Ying Liang, Hua Xiang, Shiyang Zhang, and Jun Zou. Preconditioners and their analyses for edge element saddlepoint systems arising from time-harmonic Maxwell’s equations. Numerical Algorithms, 86(1):281–302, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00889-7.

Li:2017:AMB

[LY17] Rui Li and Jun-Feng Yin. Accelerated modulus-based matrix splitting iteration methods for a restricted class of nonlinear complementarity problems. Numerical Algorithms, 75(2):339– 358, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2018:IAA

[LY18] Hao Li and Yidu Yang. C^{0} IPG adaptive algorithms for the biharmonic eigenvalue problem. Numerical Algorithms, 78(2): 553–567, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lyashenko:1997:SSG

[Lya97] Andrei A. Lyashenko. On the stability of the spectral Galerkin approximation. Numerical Algorithms, 14 (1–3):165–178, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/10/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/10/fulltext.pdf. Dynami-

cal numerical analysis (Atlanta, GA, 1995).

Li:2020:NEB

[LYH^{+}20] Chaoqian Li, Shaorong Yang, Hui Huang, Yaotang Li, and Yimin Wei. Note on error bounds for linear complementarity problems of Nekrasov matrices. Numerical Algorithms, 83(1): 355–372, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2015:TGA

[LYL15] Wei Liu, Zhe Yin, and Jin Li. A two-grid algorithm based on expanded mixed element discretizations for strongly nonlinear elliptic equations. Numerical Algorithms, 70(1):93–111, September 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9936-7.

Liang:2025:HRM

[LYL25] Ying-Qing Liang, Fan Yang, and Xiao-Xiao Li. A hybrid regularization method for identifying the source term and the initial value simultaneously for fractional pseudo-parabolic equation with involution. Numerical Algorithms, 99(4):2039–2061, August 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01944-3.

Lyness:2008:NEF

[Lyn08] J. N. Lyness. Numerical evaluation of a fixed-amplitude variable-phase integral. Numerical Algorithms, 49(1–4):235– 249, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=49&issue=1&spage=235.

Li:2014:LPI

[LYW14] Xu Li, Ai-Li Yang, and Yu-Jiang Wu. Lopsided PMHSS iteration method for a class of complex symmetric linear systems. Numerical Algorithms, 66(3):555–568, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9748-1.

Lipu:2012:FNS

[LYY12] Zhang Lipu, Bai Yanqin, and Xu Yinghong. A full-Newton step infeasible interior-point algorithm for monotone LCP based on a locally-kernel function. Numerical Algorithms, 61(1):57–81, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=1&spage=57.

Lin:2009:SNP

[LZ09] Yingzhen Lin and Yongfang Zhou. Solving nonlinear pseudoparabolic equations with nonlocal boundary conditions in reproducing kernel space. Numerical Algorithms, 52(2):173–186, October 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=2&spage=173.

Liu:2012:NCC

[LZ12] Yuanyuan Liu and Yongkui Zou. Numerical computation of connecting orbits on a manifold. Numerical Algorithms, 61(3):429–464, November 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=429.

Li:2014:SNE

[LZ14] Wen Li and Hua Zheng. Some new error bounds for linear complementarity problems of H-matrices. Numerical Algorithms, 67(2):257–269, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9786-8.

Liang:2015:SCA

[LZ15] Zhao-Zheng Liang and Guo-Feng Zhang. Semi-convergence analysis of the GPIU method for singular nonsymmetric saddle-point problems. Numerical Algorithms, 70(1):151–169, September 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9939-4.

Liang:2016:SIM

[LZ16a] Zhao-Zheng Liang and Guo-Feng Zhang. On SSOR iteration method for a class of block two-by-two linear systems. Numerical Algorithms, 71(3):655–671, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0015-5.

Liao:2016:MDD

[LZ16b] Shijun Liao and Yinlong Zhao. On the method of directly defining inverse mapping for nonlinear differential equations. Numerical Algorithms, 72(4):989–1020, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0077-4.

Li:2018:ASI

[LZ18a] Dan Li and Detong Zhu. An affine scaling interior trust-region method combining with line search filter technique for optimization subject to bounds on variables. Numerical Algorithms, 77(4):1159–1182, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liang:2018:SCA

[LZ18b] Zhao-Zheng Liang and Guo-Feng Zhang. Semi-convergence analysis of preconditioned deteriorated PSS iteration method for singular saddle point problems. Numerical Algorithms, 78 (2):379–404, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2018:PES

[LZ18c] Xiaowei Liu and Jin Zhang. Pointwise estimates of SDFEM on Shishkin triangular meshes for problems with characteristic layers. Numerical Algorithms, 78(2):465–483, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liang:2019:RAB

[LZ19a] Zhao-Zheng Liang and Guo-Feng Zhang. Robust additive block triangular preconditioners for block two-by-two linear systems. Numerical Algorithms, 82(2):503–537, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liao:2019:NBD

[LZ19b] Li-Dan Liao and Guo-Feng Zhang. A note on block diagonal and block triangular preconditioners for complex symmetric linear systems. Numerical Algorithms, 80(4):1143–1154, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2021:STT

[LZ21] Qun Li and Bing Zheng. Scaled three-term derivative-free methods for solving large-scale nonlinear monotone equations. Numerical Algorithms, 87(3):1343–1367, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01010-8.

Li:2022:ITS

[LZ22a] Nan Li and Lihong Zhi. Improved two-step Newton’s method for computing simple multiple zeros of polynomial systems. Numerical Algorithms, 91(1):19–50, September 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01253-7.

Li:2022:RTV

[LZ22b] Rong Li and Bing Zheng. The l_{2,p} regularized total variation with overlapping group sparsity prior for image restoration with impulse noise. Numerical Algorithms, 91(4):1779–1814, December 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01322-x.

Li:2022:MCG

[LZ22c] Xiangli Li and Wenjuan Zhao. A modified conjugate gradient method based on the self-scaling memoryless BFGS update. Numerical Algorithms, 90(3):1017–1042, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01220-8.

Lv:2022:SBE

[LZ22d] Peng Lv and Bing Zheng. Structured backward error analysis for a class of block three-by-three saddle point problems. Numerical Algorithms, 90(1):59–78, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01179-6.

Li:2023:UAE

[LZ23a] Jiyong Li and Liqing Zhu. A uniformly accurate exponential wave integrator Fourier pseudo-spectral method with structure-preservation for long-time dynamics of the Dirac equation with small potentials. Numerical Algorithms, 92(2): 1367–1401, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01345-4.

Liu:2023:SWG

[LZ23b] Xiaowei Liu and Jin Zhang. Supercloseness of weak Galerkin method on bakhvalov-type mesh for a singularly perturbed problem in 1D. Numerical Algorithms, 93(1):367–395, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01420-w.

Liu:2023:UCO

[LZ23c] Xiaowei Liu and Jin Zhang. Uniform convergence of optimal order for a finite element method on a Bakhvalov-type mesh for a singularly perturbed convection-diffusion equation with parabolic layers. Numerical Algorithms, 94(1):459–478, September 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01508-x.

Li:2024:ABM

[LZ24] Hongwei Li and Xiangyu Zhang. Artificial boundary method for the Zakharov–rubenchik equations. Numerical Algorithms, 97(3):1017–1037, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01739-
y.

Li:2025:SAP

[LZ25] Rong Li and Bing Zheng. Spatially adapted parameters selection based on the local constraints for Gaussian plus impulse image deblurring. Numerical Algorithms, 99(3):1531– 1549, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01924-7.

Liu:2024:OCA

[LZCW24] Xuan Liu, Yongkui Zou, Shimin Chai, and Huimin Wang. Optimal convergence analysis of weak Galerkin finite element methods for parabolic equations with lower regularity. Numerical Algorithms, 97(3):1323–1339, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01751-w.

Zhou:2026:EAF

[lZfLqL26] Xue lin Zhou, Jiao fen Li, and Chao qian Li. An efficient algorithm for fitting the three-way GIPSCAL problem with missing values from asymmetric multidimensional scaling. Numerical Algorithms, 102(1):425–466, May 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02136-3.

Li:2020:FSO

[LZL20] Rui-Xia Li, Guo-Feng Zhang, and Zhao-Zheng Liang. Fast solver of optimal control problems constrained by Ohta– Kawasaki equations. Numerical Algorithms, 85(3):787–809, November 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00837-0.

Liao:2022:GVT

[LZL22] Si-Wei Liao, Guo-Feng Zhang, and Zhao-Zheng Liang. A generalized variant of two-sweep modulus-based matrix splitting iteration method for solving horizontal linear complementarity problems. Numerical Algorithms, 90(3):1279–1303, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01230-6.

Liao:2023:PGM

[LZL23] Si-Wei Liao, Guo-Feng Zhang, and Zhao-Zheng Liang. A preconditioned general modulus-based matrix splitting iteration method for solving horizontal linear complementarity problems. Numerical Algorithms, 93(2):919–947, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01447-z.

Li:2020:SNE

[LZlL20] Xin Li, Luming Zhang, and Hong lin Liao. Sharp H^{1}-norm error estimate of a cosine pseudo-spectral scheme for 2D reactionsubdiffusion equations. Numerical Algorithms, 83(3):1223– 1248, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Li:2023:TDS

[LZM23] Qi Li, Supei Zheng, and Liquan Mei. Three decoupled, second-order accurate, and energy stable schemes for the conserved Allen–Cahn-type block copolymer (BCP) model. Numerical Algorithms, 92(2):1233–1259, February 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01338-3.

Li:2022:AAP

[LZOY22] Zhizhi Li, Huai Zhang, and Le Ou-Yang. Anderson accelerating the preconditioned modulus approach for linear complementarity problems on second-order cones. Numerical Algorithms, 91 (2):803–839, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01283-1.

Liu:2026:HOC

[LZQ26] Mengmeng Liu, Xiangcheng Zheng, and Wenlin Qiu. Highorder compact difference method for convection–reaction– subdiffusion equation with variable exponent and coefficients. Numerical Algorithms, 101(3):1677–1699, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02060-6.

Zhou:2025:TRF

[lZqL25] Xue lin Zhou and Chao qian Li. A trust-region framework for iteration solution of the direct INDSCAL problem in metric multidimensional scaling. Numerical Algorithms, 99(3):1429– 1462, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01921-w.

Law:1992:ERF

[LZRJ92] A. G. Law, C. N. Zhang, A. Rezazadeh, and L. Jódar. Evaluation of a rational function. Numerical Algorithms, 3(1–4): 265–272, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Liu:2025:IPM

[LZT25] Jinkui Liu, Ning Zhang, and Bing Tang. An inertia projection method for nonlinear pseudo-monotone equations with convex constraints. Numerical Algorithms, 99(4):1805–1828, August 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01934-5.

Liao:2020:TEC

[LZW20] Feng Liao, Luming Zhang, and Tingchun Wang. Two energy-conserving and compact finite difference schemes for two-dimensional Schrödinger–Boussinesq equations. Numerical Algorithms, 85(4):1335–1363, December 2020. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-019-00867-8.

Liu:2022:UNB

[LZX22] Jianzhou Liu, Qi Zhou, and Yebo Xiong. Upper norm bounds for the inverse of locally doubly strictly diagonally dominant matrices with its applications in linear complementarity problems. Numerical Algorithms, 90(4):1465–1491, August 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01237-z.

Liu:2023:ASQ

[LZX23] Qiaohua Liu, Qian Zhang, and Xiangjian Xu. Accuracy and stability of quaternion Gaussian elimination. Numerical Algorithms, 94(3):1159–1183, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01531-y.

Liang:2019:AIM

[LZZ19] Maolin Liang, Bing Zheng, and Ruijuan Zhao. Alternating iterative methods for solving tensor equations with applications. Numerical Algorithms, 80(4):1437–1465, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Liu:2023:HEA

[LZZ23] Zhengguang Liu, Nan Zheng, and Zhaojie Zhou. A highly efficient and accurate new SAV approach for the modified phase field crystal model. Numerical Algorithms, 93(2):543– 562, June 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01426-4.

Liu:2024:HOH

[LZZ24] Wenju Liu, Tengjin Zhao, and Zhiyue Zhang. High order hybrid asymptotic augmented finite volume methods for nonlinear degenerate wave equations. Numerical Algorithms, 96 (1):135–184, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01642-6.

Montaner:1995:FDT

[MA95] J. M. Montaner and M. Alfaro. On five-diagonal Toeplitz matrices and orthogonal polynomials on the unit circle. Numerical Algorithms, 10(1–2):137–153, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Mustapha:2012:FDM

[MA12] Kassem Mustapha and Jaafar AlMutawa. A finite difference method for an anomalous sub-diffusion equation, theory and applications. Numerical Algorithms, 61(4):525–543, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9547-0/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=4&spage=525-543.

Mohebbi:2013:CFD

[MA13] Akbar Mohebbi and Mostafa Abbaszadeh. Compact finite difference scheme for the solution of time fractional advectiondispersion equation. Numerical Algorithms, 63(3):431–452, July 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9631-5.

Magrenan:2015:ETW

[MA15] Ángel Alberto Magreñán and Ioannis K. Argyros. An extension of a theorem by Wang for Smale’s α-theory and applications. Numerical Algorithms, 68(1):47–60, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9837-9.

Magrenan:2016:ICA

[MA16] Ángel Alberto Magreñán and Ioannis K. Argyros. Improved convergence analysis for Newton-like methods. Numerical Algorithms, 71(4):811–826, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0025-3.

Ma:2020:CSP

[Ma20] Feng Ma. Convergence study on the proximal alternating direction method with larger step size. Numerical Algorithms, 85(2): 399–425, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00819-2.

Mohammad:2022:GCD

[MA22] Hassan Mohammad and Aliyu Muhammed Awwal. Globally convergent diagonal Polak–Ribière–Polyak like algorithm for nonlinear equations. Numerical Algorithms, 91(4):1441–1460, December 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01309-8.

MacLeod:1996:RAS

[Mac96] Allan J. MacLeod. Rational approximations, software and test methods for sine and cosine integrals. Numerical Algorithms, 12(3–4):259–272, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mobeen:2025:NBC

[MAF25] Ashia Mobeen, Shahbaz Ahmad, and Faisal Fairag. Nonblind constraint image deblurring problem with mean curvature functional. Numerical Algorithms, 98(4):1703–1723, April 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01848-2.

Mustapha:2016:DGM

[MAFN16] K. Mustapha, B. Abdallah, K. M. Furati, and M. Nour. A discontinuous Galerkin method for time fractional diffusion equations with variable coefficients. Numerical Algorithms, 73(2): 517–534, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0106-y.

Mahmudov:2010:MQB

[Mah10] Nazim Mahmudov. The moments for q-Bernstein operators in the case 0 < q < 1. Numerical Algorithms, 53(4):439–450, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=53&
issue=4&spage=439.

Majidi:2022:GSD

[MAH22] Tahereh Majidi, Ali Abdi, and Gholamreza Hojjati. Generalized second derivative linear multistep methods for ordinary differential equations. Numerical Algorithms, 91(1):227–250, September 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01260-8.

Maignan:2001:RSE

[Mai01] Aude Maignan. Real solving of elementary-algebraic systems. Numerical Algorithms, 27(2):153–167, June 2001. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp007.lwwonline.com/content/
getfile/5058/35/4/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/35/4/fulltext.pdf.

Majidian:2013:NAH

[Maj13] Hassan Majidian. Numerical approximation of highly oscillatory integrals on semi-finite intervals by steepest descent method. Numerical Algorithms, 63(3):537–548, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9639-x.

Majidian:2014:MEM

[Maj14] Hassan Majidian. Modified Euler’s method with a graded mesh for a class of Volterra integral equations with weakly singular kernel. Numerical Algorithms, 67(2):405–422, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9799-3.

Majidian:2018:ACO

[Maj18] Hassan Majidian. Automatic computing of oscillatory integrals. Numerical Algorithms, 77(3):867–884, March 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mursaleen:2020:APE

[MAK20] Mohammad Mursaleen, Khursheed J. Ansari, and Asif Khan. Approximation properties and error estimation of q-Bernstein shifted operators. Numerical Algorithms, 84(1):207–227, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Markov:2004:SAS

[MAL04] Svetoslav Markov, Rene Alt, and Jean-Luc Lamotte. Stochastic arithmetic: s-spaces and some applications. Numerical Algorithms, 37(1–4):275–284, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/22/abstract.htm.

Malik:2021:ASS

[Mal21] Avinash Malik. Adaptive step size numerical integration for stochastic differential equations with discontinuous drift and diffusion. Numerical Algorithms, 87(2):849–872, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00990-x.

Mantica:2007:FMP

[Man07] Giorgio Mantica. Fractal measures and polynomial sampling: I.F.S.–Gaussian integration. Numerical Algorithms, 45(1–4):

269–281, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-007-9111-5.

Mantica:2010:DSN

[Man10] Giorgio Mantica. Dynamical systems and numerical analysis: the study of measures generated by uncountable I.F.S. Numerical Algorithms, 55(2–3):321–335, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=321.

Manh:2017:HIS

[Man17] Phung Van Manh. Hermite interpolation with symmetric polynomials. Numerical Algorithms, 76(3):709–725, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mannel:2021:CPB

[Man21] Florian Mannel. Convergence properties of the Broydenlike method for mixed linear-nonlinear systems of equations. Numerical Algorithms, 88(2):853–881, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01060-y;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01060-y.pdf.

Maroni:1992:TDO

[Mar92] P. Maroni. Two-dimensional orthogonal polynomials, their associated sets and the co-recursive sets. Numerical Algorithms, 3(1–4):299–311, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Marletta:1993:ASR

[Mar93] Marco Marletta. Automatic solution of regular and singular vector Sturm–Liouville problems. Numerical Algorithms, 4(1– 2):65–99, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Maroni:1996:RFD

[Mar96] P. Maroni. On a regular form defined by a pseudo-function. Numerical Algorithms, 11(1–4):243–254, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Marciniak:2004:IIM

[Mar04a] Andrzej Marciniak. Implicit interval methods for solving the initial value problem. Numerical Algorithms, 37(1–4):241–251, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp009.kluweronline.com/IPS/content/ext/
x/J/5058/I/58/A/19/abstract.htm.

Markot:2004:OPE

[Mar04b] Mihály Csaba Markót. Optimal packing of 28 equal circles in a unit square — the first reliable solution. Numerical Algorithms, 37(1–4):253–261, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/20/abstract.htm.

Markov:2004:QSC

[Mar04c] Svetoslav Markov. On quasivector spaces of convex bodies and zonotopes. Numerical Algorithms, 37(1–4):263–274, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/21/abstract.htm.

Mastroianni:1995:GCF

[Mas95] G. Mastroianni. Generalized Christoffel functions and error of positive quadrature. Numerical Algorithms, 10(1–2):113–126, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Mason:2005:MPC

[Mas05] J. C. Mason. The minimality properties of Chebyshev polynomials and their lacunary series. Numerical Algorithms, 38 (1–3):61–78, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mahdavi-Amiri:2017:ACP

[MAS17] N. Mahdavi-Amiri and M. Shaeiri. An adaptive competitive penalty method for nonsmooth constrained optimization. Numerical Algorithms, 75(1):305–336, May 2017. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Matos:1991:ARV

[Mat91] Ana C. Matos. Acceleration results for the vector E-algorithm. Numerical Algorithms, 1(3):237–260, September 1991. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Matos:1992:CAP

[Mat92] Ana C. Matos. Convergence and acceleration properties for the vector ϵ-algorithm. Numerical Algorithms, 3(1–4):313–319, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Matos:1996:SCR

[Mat96] Ana C. Matos. Some convergence results for the generalized Padé-type approximants. Numerical Algorithms, 11(1–4):255– 269, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Matthies:2001:MFE

[Mat01] Gunar Matthies. Mapped finite elements on hexahedra. necessary and sufficient conditions for optimal interpolation errors. Numerical Algorithms, 27(4):317–327, August 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/1/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
1/fulltext.pdf.

Mathar:2004:NRI

[Mat04] Richard J. Mathar. Numerical representations of the incomplete gamma function of complex-valued argument. Numerical Algorithms, 36(3):247–264, July 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/57/A/4/abstract.htm.

Matejas:2015:AOS

[Mat15] Josip Matejas. Accuracy of one step of the Falk–Langemeyer method. Numerical Algorithms, 68(4):645–670, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9865-5.

Mazure:1999:BBM

[Maz99] Marie-Laurence Mazure. Bernstein bases in Müntz spaces. Numerical Algorithms, 22(3–4):285–304, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/23/7/abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/5058/23/7/fulltext.
pdf.

Mazure:2002:BPP

[Maz02] Marie-Laurence Mazure. Blossoming: from polynomials to power series. Numerical Algorithms, 30(2):141–155, June 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/41/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
41/2/fulltext.pdf.

Mazure:2005:BS

[Maz05a] Marie-Laurence Mazure. Blossoming stories. Numerical Algorithms, 39(1–3):257–288, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mazure:2005:TEP

[Maz05b] Marie-Laurence Mazure. Towards existence of piecewise Chebyshevian B-spline bases. Numerical Algorithms, 39(4): 399–414, August 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mazure:2009:BTO

[Maz09a] Marie-Laurence Mazure. Bernstein-type operators in Chebyshev spaces. Numerical Algorithms, 52(1):93–128, September 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=1&
spage=93.

Mazure:2009:FRR

[Maz09b] Marie-Laurence Mazure. A few remarks on recurrence relations for geometrically continuous piecewise Chebyshevian B-splines.

Numerical Algorithms, 51(4):401–428, August 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=4&spage=401.

Mazure:2011:DEC

[Maz11a] Marie-Laurence Mazure. Dimension elevation for Chebyshevian splines. Numerical Algorithms, 56(1):1–16, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=1&
spage=1.

Mazure:2011:GNC

[Maz11b] Marie-Laurence Mazure. On a general new class of quasi Chebyshevian splines. Numerical Algorithms, 58(3):399–438, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
58&issue=3&spage=399.

Mazure:2012:PSE

[Maz12] Marie-Laurence Mazure. Polynomial splines as examples of Chebyshevian splines. Numerical Algorithms, 60(2):241–262, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=60&
issue=2&spage=241.

Mazure:2018:PCS

[Maz18] Marie-Laurence Mazure. Piecewise Chebyshevian splines: interpolation versus design. Numerical Algorithms, 77(4):1213– 1247, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mehne:2006:NMS

[MB06] H. Hashemi Mehne and A. Hashemi Borzabadi. A numerical method for solving optimal control problems using state parametrization. Numerical Algorithms, 42(2):165–169, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=42&issue=2&
spage=165.

MacNamara:2009:KSS

[MB09] Shev MacNamara and Kevin Burrage. Krylov and steadystate techniques for the solution of the chemical master equation for the mitogen-activated protein kinase cascade. Numerical Algorithms, 51(3):281–307, July 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=3&spage=281.

Tributes to Gene H. Golub, Part III.

Ma:2019:PCB

[MBG19] Feng Ma, Yiming Bi, and Bin Gao. A prediction-correction-based primal-dual hybrid gradient method for linearly constrained convex minimization. Numerical Algorithms, 82(2): 641–662, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Madhu:2017:IDS

[MBJ17] Kalyanasundaram Madhu, D. K. R. Babajee, and Jayakumar Jayaraman. An improvement to double-step Newton method and its multi-step version for solving system of nonlinear equations and its applications. Numerical Algorithms, 74(2):593– 607, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0163-2.

Millwater:2025:CMP

[MBM25] H. Millwater, M. Balcer, and N. Millwater. Calculation of machine precision second order derivatives using dual-complex numbers. Numerical Algorithms, 99(4):1871–1924, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01937-2.

Maclean:2021:TEF

[MBR21] John Maclean, J. E. Bunder, and A. J. Roberts. A toolbox of equation-free functions in Matlab /Octave for efficient system level simulation. Numerical Algorithms, 87(4):1729–1748, August 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01027-z.

Marcuzzi:2005:PSA

[MC05a] F. Marcuzzi and M. Morandi Cecchi. A parallel solver for adaptive finite element discretizations. Numerical Algorithms, 40(3):217–231, November 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=3&spage=217.

Mason:2005:LAAa

[MC05b] J. C. Mason and A. Crampton. Laurent–Padé approximants to four kinds of Chebyshev polynomial expansions. Part I. Maehly type approximants. Numerical Algorithms, 38(1–3):3– 18, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mason:2005:LAAb

[MC05c] J. C. Mason and A. Crampton. Laurent–Padé approximants to four kinds of Chebyshev polynomial expansions. Part II: Clenshaw–Lord type approximants. Numerical Algorithms, 38 (1–3):19–29, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mahmoud:2008:VII

[MC08] Sayed Mahmoud and Xiaojun Chen. A verified inexact implicit Runge–Kutta method for nonsmooth ODEs. Numerical Algorithms, 47(3):275–290, March 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=3&spage=275.

Mandal:2026:DSF

[MCC26] Nitai Mandal, Subham Chatterjee, and Gorachand Chakraborty. On the dynamics of Steffensen’s family of root-finding methods for entire function. Numerical Algorithms, 102(1):409– 424, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02135-4.

Moradi:2024:JFE

[MCDS24] Afsaneh Moradi, Jeremy Chouchoulis, Raffaele D’Ambrosio, and Jochen Schütz. Jacobian-free explicit multiderivative general linear methods for hyperbolic conservation laws. Numerical Algorithms, 97(4):1823–1858, December 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01771-6.

Martinez:2004:EUG

[MCG^{+}04] J. A. Martínez, L. G. Casado, I. García, Ya. D. Sergeyev, and B. Tóth. On an efficient use of gradient information for accelerating interval global optimization algorithms. Numerical Algorithms, 37(1–4):61–69, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/4/abstract.htm.

McLachlan:2002:FHO

[McL02] Robert I. McLachlan. Families of high-order composition methods. Numerical Algorithms, 31(1–4):233–246, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/43/4/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
43/4/fulltext.pdf.

McLachlan:2025:RKM

[McL25] Robert I. McLachlan. Runge–Kutta methods determined from extended phase space methods for Hamiltonian systems. Numerical Algorithms, 100(3):983–993, November 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01976-9.

Ming:2016:CMM

[MClXzJe16] Li Ming, Li Chen-liang, Cui Xiang-zhao, and Zhao Jin-e. Cascadic multigrid methods combined with sixth order compact scheme for Poisson equation. Numerical Algorithms, 71(4): 715–727, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0018-2.

Moir:2020:SNI

[MCMX20] Robert H. C. Moir, Robert M. Corless, Marc Moreno Maza, and Ning Xie. Symbolic–numeric integration of rational functions. Numerical Algorithms, 83(4):1295–1320, April 2020.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mao:2022:PEE

[MCW22] Wenting Mao, Yanping Chen, and Huasheng Wang. A posteriori error estimations of the Petrov–Galerkin methods for fractional Helmholtz equations. Numerical Algorithms, 89(3): 1095–1127, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01147-0.

Martinez:2012:HIL

[Md12] Jose Mario Martínez and Leandro da Fonseca Prudente. Handling infeasibility in a large-scale nonlinear optimization algorithm. Numerical Algorithms, 60(2):263–277, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=2&
spage=263.

Mina-Diaz:2015:LCP

[MD15] Erwin Miña-Díaz. On the leading coefficient of polynomials orthogonal over domains with corners. Numerical Algorithms, 70 (1):1–8, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9932-y.

Macias-Diaz:2021:NEV

[MD21a] Jorge E. Macías-Díaz. A numerically efficient variational algorithm to solve a fractional nonlinear elastic string equation. Numerical Algorithms, 86(1):75–102, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00880-2.

Maleki:2021:AAC

[MD21b] Mohammad Maleki and Ali Davari. Analysis of an adaptive collocation solution for retarded and neutral delay systems. Numerical Algorithms, 88(1):67–91, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01030-4.

Moghaderi:2016:FET

[MDH16] Hamid Moghaderi, Mehdi Dehghan, and Masoud Hajarian. A fast and efficient two-grid method for solving d-dimensional Poisson equations. Numerical Algorithms, 72(3):483–537, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0057-8.

Mohanty:2015:OCM

[MDL15] R. K. Mohanty, Weizhong Dai, and Don Liu. Operator compact method of accuracy two in time and four in space for the solution of time dependent Burgers–Huxley equation. Numerical Algorithms, 70(3):591–605, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9963-z.

Maroni:2008:CCB

[MdR08] P. Maroni and Z. da Rocha. Connection coefficients between orthogonal polynomials and the canonical sequence: an approach based on symbolic computation. Numerical Algorithms, 47(3):291–314, March 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=3&spage=291.

Maroni:2013:CCO

[MdR13] Pascal Maroni and Zélia da Rocha. Connection coefficients for orthogonal polynomials: symbolic computations, verifications and demonstrations in the Mathematica language. Numerical Algorithms, 63(3):507–520, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9634-2.

Mi:2023:LPF

[MDR23] Shilin Mi, Guangzhi Du, and Yao Rong. Local and parallel finite element algorithms for magnetohydrodynamic flows with low electromagnetic Reynolds number. Numerical Algorithms, 93(4):1661–1683, August 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01484-8.

MacDonald:1992:IOR

[ME92] Craig MacDonald and Wayne H. Enright. Implications of order reduction for implicit Runge–Kutta methods. Numerical Algorithms, 2(3–4):351–369, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mason:1995:CNM

[ME95] J. C. Mason and G. H. Elliott. Constrained near-minimax approximation by weighted expansion and interpolation using Chebyshev polynomials of the second, third, and fourth kinds. Numerical Algorithms, 9(1–2):39–54, May 1995. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Medkova:2010:IRS

[Med10] Dagmar Medková. Integral representation of a solution of the Neumann problem for the Stokes system. Numerical Algorithms, 54(4):459–484, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=4&spage=459.

Mehrkanoon:2011:DVS

[Meh11] Siamak Mehrkanoon. A direct variable step block multistep method for solving general third-order ODEs. Numerical Algorithms, 57(1):53–66, May 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=1&spage=53.

Messaoudi:2019:GNA

[MEJS19] Abderrahim Messaoudi, Mohammed Errachid, Khalide Jbilou, and Hassane Sadok. GRPIA: a new algorithm for computing interpolation polynomials. Numerical Algorithms, 80(1): 253–278, January 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Melkonian:2010:SBF

[Mel10]  Dmitriy Melkonian.  Similar basis function algorithm

for numerical estimation of Fourier integrals. Numerical Algorithms, 54(1):73–100, May 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=73.

Melman:2014:IPT

[Mel14] A. Melman. Implementation of Pellet’s theorem. Numerical Algorithms, 65(2):293–304, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9705-z.

Melman:2024:EAC

[Mel24] A. Melman. An efficient approximation to the Cauchy radius. Numerical Algorithms, 96(1):1–11, May 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01636-4.

Meng:2026:CIM

[Men26] Jie Meng. On a class of infinite M-matrices with an extended quasi-Toeplitz structure. Numerical Algorithms, 101 (3):1965–1988, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02071-3.

Merrien:1992:FHI

[Mer92] J. L. Merrien. A family of Hermite interpolants by bisection algorithms. Numerical Algorithms, 2(2):187–200, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Merrien:1994:DHI

[Mer94] Jean-Louis Merrien. Dyadic Hermite interpolants on a triangulation. Numerical Algorithms, 7(2–4):391–410, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Metzger:2019:SDR

[Met19] Stefan Metzger. On stable, dissipation reducing splitting schemes for two-phase flow of electrolyte solutions. Numerical Algorithms, 80(4):1361–1390, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Meurant:1997:CBN

[Meu97] Gérard Meurant. The computation of bounds for the norm of the error in the conjugate gradient algorithm. Numerical Algorithms, 16(1):77–87, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
9/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/9/6/fulltext.pdf. Sparse ma-

trices in industry (Lille, 1997).

Meurant:1999:NEC

[Meu99] Gérard Meurant. Numerical experiments in computing bounds for the norm of the error in the preconditioned conjugate gradient algorithm. Numerical Algorithms, 22 (3–4):353–365, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/6/fulltext.pdf.

Meurant:2002:MAP

[Meu02] Gérard Meurant. A multilevel AINV preconditioner. Numerical Algorithms, 29(1–3):107–129, March 2002. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp008.lwwonline.com/content/
getfile/5058/38/8/abstract.htm;   http://ipsapp008.
lwwonline.com/content/getfile/5058/38/8/fulltext.pdf.

Meurant:2005:ENE

[Meu05] Gérard Meurant. Estimates of the l_{2} norm of the error in the conjugate gradient algorithm. Numerical Algorithms, 40(2):157–169, October 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=157.

Meurant:2009:ETI

[Meu09a] Gérard Meurant. Estimates of the trace of the inverse of a symmetric matrix using the modified Chebyshev algorithm. Numerical Algorithms, 51(3):309–318, July 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=3&spage=309.

Tributes to Gene H. Golub, Part III.

Meurant:2009:GHN

[Meu09b] Gérard Meurant. Gene H. Golub 1932–2007. Numerical Algorithms, 51(1):1–4, May 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=1&spage=1. Trib-

utes to Gene H. Golub, Part I.

Meurant:2012:CIV

[Meu12] Gérard Meurant. The computation of isotropic vectors. Numerical Algorithms, 60(2):193–204, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=2&spage=193.

Meurant:2019:Pb

[Meu19a] G. Meurant. Postface. Numerical Algorithms, 80(1):303, January 2019. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-018-0582-3.pdf.

Meurant:2019:Pa

[Meu19b] G. Meurant. Preface. Numerical Algorithms, 80(1):1–3, January 2019. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-018-0545-8.pdf.

Meurant:2020:PCB

[Meu20] Gérard Meurant. On prescribing the convergence behavior of the conjugate gradient algorithm. Numerical Algorithms, 84(4):1353–1380, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00851-2.

Meurant:2023:DCS

[Meu23] Gérard Meurant. Detection and correction of silent errors in the conjugate gradient algorithm. Numerical Algorithms, 92 (1):869–891, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01380-1.

Meurant:2025:NGM

[Meu25] Gérard Meurant. A note on the Grcar matrix. Numerical Algorithms, 100(4):1459–1476, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01999-2. See correction [Meu26].

Meurant:2026:CNG

[Meu26] Gérard Meurant. Correction to: A note on the Grcar matrix. Numerical Algorithms, 101(3):2153, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02076-y. See [Meu25].

Mezzadri:2022:MBF

[Mez22] Francesco Mezzadri. A modulus-based formulation for the vertical linear complementarity problem. Numerical Algorithms, 90(4):1547–1568, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01240-
4.

Maxim:2023:TRP

[MFBB23] Voichiţa Maxim, Yuemeng Feng, Hussein Banjak, and Elie Bretin. Tomographic reconstruction from Poisson distributed data: a fast and convergent EM–TV dual approach. Numerical Algorithms, 94(2):701–731, October 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01517-w.

Mayrhofer:2015:USA

[MFK^{+}15] Arno Mayrhofer, Martin Ferrand, Christophe Kassiotis, Damien Violeau, and François-Xavier Morel. Unified semianalytical wall boundary conditions in SPH: analytical extension to 3-D. Numerical Algorithms, 68(1):15–34, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9835-y.

Mazzia:2007:CNI

[MFPG07] Annamaria Mazzia, Massimiliano Ferronato, Giorgio Pini, and Giuseppe Gambolati. A comparison of numerical integration rules for the meshless local Petrov–Galerkin method. Numerical Algorithms, 45(1–4):61–74, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=61.

Muhlbach:1991:MPI

[MG91] G. Mühlbach and M. Gasca. Multivariate polynomial interpolation under projectivities. I. Lagrange and Newton interpolation formulas. Numerical Algorithms, 1(4):375–399, November 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Muhlbach:1994:MPI

[MG94] G. Mühlbach and M. Gasca. Multivariate polynomial interpolation under projectivities. III. Remainder formulas. Numerical Algorithms, 8(1):103–109, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Misirli:2011:MAB

[MG11a] Emine Misirli and Yusuf Gurefe. Multiplicative Adams Bashforth–Moulton methods. Numerical Algorithms, 57(4): 425–439, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=57&issue=4&spage=425.

Mokhtary:2011:CLS

[MG11b] Payam Mokhtary and Farideh Ghoreishi. The L^{2}-convergence of the Legendre spectral Tau matrix formulation for nonlinear fractional integro differential equations. Numerical Algorithms, 58(4):475–496, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=475.

Moudafi:2018:RSF

[MG18] Abdellatif Moudafi and Aviv Gibali. l_{1}–l_{2} regularization of split feasibility problems. Numerical Algorithms, 78(3):739– 757, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mezzadri:2020:MBM

[MG20] Francesco Mezzadri and Emanuele Galligani. Modulus-based matrix splitting methods for horizontal linear complementarity problems. Numerical Algorithms, 83(1):201–219, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mezzadri:2021:MBM

[MG21] Francesco Mezzadri and Emanuele Galligani. Modulus-based matrix splitting methods for a class of horizontal nonlinear complementarity problems. Numerical Algorithms, 87(2):667– 687, June 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00983-w.

Morin:2022:CSB

[MG22] Martin Morin and Pontus Giselsson. Cocoercivity, smoothness and bias in variance-reduced stochastic gradient methods. Numerical Algorithms, 91(2):749–772, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01280-4.

Muniyasamy:2025:UCA

[MGG25] Muniyasamy M, Santhosh George, and Chandhini G. Unified convergence analysis of a class of iterative methods. Numerical Algorithms, 99(2):683–715, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01893-x.

Ma:2020:EMC

[MGL20] Jie Ma, Feng Gao, and Yongshu Li. An efficient method for computing the outer inverse A_{T,S}^{(2)} through Gauss–Jordan elimination. Numerical Algorithms, 85(1):77–106, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00803-w.

Mokhtary:2008:SMI

[MH08] P. Mokhtary and S. M. Hosseini. “rescale and modify” implementation of IRKS methods. Numerical Algorithms, 47(4):315–325, April 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=4&spage=315.

Ma:2021:ALC

[MH21] Xiaohua Ma and Chengming Huang. An accurate Legendre collocation method for third-kind Volterra integro-differential equations with non-smooth solutions. Numerical Algorithms, 88(4):1571–1593, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01086-w.

Muoi:2022:PAM

[MH22] Pham Quy Muoi and Duong Xuan Hiep. Proximal algorithm for minimization problems in l_{0}-regularization for nonlinear inverse problems. Numerical Algorithms, 91(1):367–388, September 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01266-2.

Ma:2023:EHV

[MH23] Zheng Ma and Chengming Huang. An hp-version fractional collocation method for Volterra integro-differential equations with weakly singular kernels. Numerical Algorithms, 92(4): 2377–2404, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01394-9.

Movahedinejad:2016:SDG

[MHA16] A. Movahedinejad, G. Hojjati, and A. Abdi. Second derivative general linear methods with inherent Runge–Kutta stability. Numerical Algorithms, 73(2):371–389, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0099-6.

May:2023:CPM

[MHR23] Ian C. T. May, Ronald D. Haynes, and Steven J. Ruuth. A closest point method library for PDEs on surfaces with parallel domain decomposition solvers and preconditioners. Numerical Algorithms, 93(2):615–637, June 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01429-1.

Mathelin:2005:SAU

[MHZ05] Lionel Mathelin, M. Yousuff Hussaini, and Thomas A. Zang. Stochastic approaches to uncertainty quantification in CFD simulations. Numerical Algorithms, 38(1–3):209–236, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Miao:2019:FKL

[Mia19] Cun-Qiang Miao. Filtered Krylov-like sequence method for symmetric eigenvalue problems. Numerical Algorithms, 82(3): 791–807, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Micchelli:1991:UREa

[Mic91] Charles A. Micchelli. Using the refinement equation for the construction of pre-wavelets. Numerical Algorithms, 1(1):75– 116, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Micchelli:1993:OEL

[Mic93] Charles A. Micchelli. Optimal estimation of linear operators from inaccurate data: a second look. Numerical Algorithms, 5(1–4):375–390, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Micula:2023:NST

[Mic23] Sanda Micula. Numerical solution of two-dimensional Hammerstein integral equations via quadratic spline collocation. Numerical Algorithms, 93(3):1225–1241, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01465-x.

Milovanovic:1995:SSG

[Mil95] Gradimir V. Milovanović. Summation of series and Gaussian quadratures. II. Numerical Algorithms, 10(1–2):127–136, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Milewski:2013:SCA

[Mil13] SłAwomir Milewski. Selected computational aspects of the meshless finite difference method. Numerical Algorithms, 63(1):107–126, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/link.springer.com/article/10.1007/s11075-012-9614-
6;   http://link.springer.com/content/pdf/10.1007/
s11075-012-9614-6.pdf.

Milovanovic:2017:SNC

[Mil17] Gradimir V. Milovanović. Symbolic–numeric computation of orthogonal polynomials and Gaussian quadratures with respect to the cardinal B-spline. Numerical Algorithms, 76(2):333–347, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Millan:2018:TAS

[Mil18] R. Díaz Millán. Two algorithms for solving systems of inclusion problems. Numerical Algorithms, 78(4):1111–1127, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Milosevic:2019:DBE

[Mil19] Marija Milosević. Divergence of the backward Euler method for ordinary stochastic differential equations. Numerical Algorithms, 82(4):1395–1407, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Milewski:2020:MSA

[Mil20] Sławomir Milewski. A Matlab software for approximate solution of 2D elliptic problems by means of the meshless Monte Carlo random walk method. Numerical Algorithms, 83(2): 565–591, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-019-00694-x.pdf.

Mitrouli:2011:STD

[Mit11] Marilena Mitrouli. A sign test for detecting the equivalence of Sylvester Hadamard matrices. Numerical Algorithms, 57(2):169–186, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=57&issue=2&spage=169.

Miyatake:2024:NFF

[Miy24] Yuto Miyatake. A new family of fourth-order energy-preserving integrators. Numerical Algorithms, 96(3):1269–1293, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01824-w.

Marciniak:2018:IVM

[MJ18] Andrzej Marciniak and Malgorzata A. Jankowska. Interval versions of Milne’s multistep methods. Numerical Algorithms, 79(1):87–105, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-017-0429-3.
pdf.

Marciniak:2020:IMA

[MJ20] Andrzej Marciniak and Malgorzata A. Jankowska. Interval methods of Adams–Bashforth type with variable step sizes. Numerical Algorithms, 84(2):651–678, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-019-00774-y;   http://link.springer.
com/content/pdf/10.1007/s11075-019-00774-y.pdf.

Mboup:2009:NDA

[MJF09] Mamadou Mboup, Cédric Join, and Michel Fliess. Numerical differentiation with annihilators in noisy environment. Numerical Algorithms, 50(4):439–467, April 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=4&spage=439.

Marciniak:2017:IPC

[MJH17] Andrzej Marciniak, Małgorzata A. Jankowska, and Tomasz Hoffmann. On interval predictor–corrector methods. Numerical Algorithms, 75(3):777–808, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-016-
0220-x.pdf.

Ma:2023:MIT

[MJJ^{+}23] Guodong Ma, Jiachen Jin, Jinbao Jian, Jianghua Yin, and Daolan Han. A modified inertial three-term conjugate gradient projection method for constrained nonlinear equations with applications in compressed sensing. Numerical Algorithms, 92 (3):1621–1653, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01356-1.

Mitrouli:1994:CMA

[MK94] M. Mitrouli and G. Kalogeropoulos. A compound matrix algorithm for the computation of the Smith form of a polynomial matrix. Numerical Algorithms, 7(2–4):145–159, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mitrouli:1997:CSN

[MK97] M. Mitrouli and C. Koukouvinos. On the computation of the Smith normal form of compound matrices. Numerical Algorithms, 16(2):95–105, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/2/fulltext.pdf.

Mansouri:1998:EPN

[MK98] N. Mansouri and J. P. Kernévez. Estimation of parameters in nonlinear problems. Numerical Algorithms, 17(3– 4):333–343, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
12/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/12/7/fulltext.pdf.

Mohanty:2017:NTV

[MK17] R. K. Mohanty and Deepti Kaur. Numerov type variable mesh approximations for 1D unsteady quasi-linear biharmonic problem: application to Kuramoto–Sivashinsky equation. Numerical Algorithms, 74(2):427–459, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0154-3.

Maleknejad:2014:SID

[MKA14] Khosrow Maleknejad, Iraj Najafi Khalilsaraye, and Mahdieh Alizadeh. On the solution of the integro-differential equation with an integral boundary condition. Numerical Algorithms, 65(2):355–374, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9709-8.

Mustafa:2026:BDB

[MKA26] Rakib Mustafa, Faheem Khan, and Hijaz Ahmad. Bounding the distance between the k-th control mesh and the limit surface of Li’s subdivision. Numerical Algorithms, 101(3): 1785–1807, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02063-3.

Malzoumati-Khiaban:2019:LTA

[MKBY19] M. Malzoumati-Khiaban, A. Foroush Bastani, and M. R. Yaghouti. Long-term adaptive symplectic numerical integration of linear stochastic oscillators driven by additive white noise. Numerical Algorithms, 80(3):1059–1095, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

M:2024:NMA

[MKG24] Sivalingam S. M, Pushpendra Kumar, and V. Govindaraj. A novel method to approximate fractional differential equations based on the theory of functional connections. Numerical Algorithms, 95(1):527–549, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01580-
3.

Mulupuru:2004:IAP

[MKO04] Bharat C. Mulupuru, Vladik Kreinovich, and Roberto Osegueda. Interval approach to phase measurements can lead to arbitrarily complex sets — a theorem and ways around it. Numerical Algorithms, 37(1–4):285–299, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/23/abstract.htm.

Monovasilis:2018:TFC

[MKS18] Th. Monovasilis, Z. Kalogiratou, and T. E. Simos. Trigonometrical fitting conditions for two derivative Runge–Kutta methods. Numerical Algorithms, 79(3):787–800, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Morgado:2010:FDS

[ML10] M. Luísa Morgado and Pedro Miguel Lima. Finite difference solution of a singular boundary value problem for the p-Laplace operator. Numerical Algorithms, 55(2–3):337–348, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=2&
spage=337.

Ma:2020:FCB

[ML20] Junjie Ma and Huilan Liu. Fractional collocation boundary value methods for the second kind Volterra equations with weakly singular kernels. Numerical Algorithms, 84(2):743– 760, June 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00777-9.

Mainge:2022:AMF

[ML22] Paul-Emile Maingé and Florian Labarre. Accelerated methods with fastly vanishing subgradients for structured nonsmooth minimization. Numerical Algorithms, 90(1):99–136, May 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01181-y.

Moreno:2019:NAS

[MLM19] J. Moreno, Miguel A. López, and R. Martínez. A new algorithm for solving all the real roots of a nonlinear system of equations in a given feasible region. Numerical Algorithms, 82 (1):123–154, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ma:2025:PGR

[MLM25] Shan Ma, Si Li, and Feng Ma. Preconditioned golden ratio primal–dual algorithm with linesearch. Numerical Algorithms, 98(3):1281–1311, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01834-8.

Maple:1999:SHS

[MM99] Carsten R. Maple and Marco Marletta. Solving Hamiltonian systems arising from ODE eigenproblems. Numerical Algorithms, 22(3–4):263–284, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/2/fulltext.pdf.

Magnus:2000:EFI

[MM00] Alphonse P. Magnus and Jean Meinguet. The elliptic functions and integrals of the “1/9” problem. Numerical Algorithms, 24(1–2):117–139, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
27/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/27/8/fulltext.pdf. Computa-

tional methods from rational approximation theory (Wilrijk, 1999).

Maxim:2004:SSI

[MM04] Voichiţa Maxim and Marie-Laurence Mazure. Subdivision schemes and irregular grids. Numerical Algorithms, 35(1): 1–28, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/49/A/2/
abstract.htm.

Maroni:2008:SSC

[MM08a] P. Maroni and M. Mejri. The symmetric D_{ω}-semi-classical orthogonal polynomials of class one. Numerical Algorithms, 49(1–4):251–282, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=251.

Mastroianni:2008:SNA

[MM08b] G. Mastroianni and G. Monegato. Some new applications of truncated Gauss–Laguerre quadrature formulas. Numerical Algorithms, 49(1–4):283–297, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=283.

Mastroianni:2009:SNA

[MM09a] G. Mastroianni and G. Monegato. Some new applications of truncated Gauss–Laguerre quadrature formulas. Numerical Algorithms, 52(3):507, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=507.

McLean:2009:CAD

[MM09b] William McLean and Kassem Mustapha. Convergence analysis of a discontinuous Galerkin method for a sub-diffusion equation. Numerical Algorithms, 52(1):69–88, September 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=1&
spage=69.

Mustapha:2011:PLD

[MM11] Kassem Mustapha and William McLean. Piecewise-linear, discontinuous Galerkin method for a fractional diffusion equation. Numerical Algorithms, 56(2):159–184, February 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=2&spage=159.

Matei:2012:ACN

[MM12] Basarab Mateï and Sylvain Meignen. Analysis of a class of nonlinear and non-separable multiscale representations. Numerical Algorithms, 60(3):391–418, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=3&spage=391.

Maruster:2017:LCG

[MM17] St. Maruster and L. Maruster. Local convergence of generalized Mann iteration. Numerical Algorithms, 76(4):905–916, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Maruster:2019:PSP

[MM19] Stefan Maruster and Laura Maruster. On the problem of starting points for iterative methods. Numerical Algorithms, 81(3): 1149–1155, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-018-0589-9.pdf.

McLean:2022:USS

[MM22] William McLean and Kassem Mustapha. Uniform stability for a spatially discrete, subdiffusive Fokker–Planck equation. Numerical Algorithms, 89(4):1441–1463, April 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01160-3.

McLean:2023:EPD

[MM23] William McLean and Kassem Mustapha. Error profile for discontinuous Galerkin time stepping of parabolic PDEs. Numerical Algorithms, 93(1):157–177, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01410-
y.

Marin:2017:UPM

[MMGH17] J. Marín, J. Mas, D. Guerrero, and K. Hayami. Updating preconditioners for modified least squares problems. Numerical Algorithms, 75(2):491–508, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mokhtary:2020:CAN

[MMLM20] P. Mokhtary, B. P. Moghaddam, A. M. Lopes, and J. A. Tenreiro Machado. A computational approach for the non-smooth solution of non-linear weakly singular Volterra integral equation with proportional delay. Numerical Algorithms, 83(3): 987–1006, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Miculescu:2020:NAG

[MMU20] Radu Miculescu, Alexandru Mihail, and Silviu-Aurelian Urziceanu. A new algorithm that generates the image of the attractor of a generalized iterated function system. Numerical Algorithms, 83(4):1399–1413, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Marco:2017:API

[MMV17] Ana Marco, José-Javier Martínez, and Raquel Viana. Accurate polynomial interpolation by using the Bernstein basis. Numerical Algorithms, 75(3):655–674, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Marco:2019:LSP

[MMV19] Ana Marco, José-Javier Martínez, and Raquel Viaña. Least squares problems involving generalized Kronecker products and application to bivariate polynomial regression. Numerical Algorithms, 82(1):21–39, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Marsland:2020:PIB

[MMW20] Stephen Marsland, Robert I. McLachlan, and Matthew C. Wilkins. Parallelization, initialization, and boundary treatments for the diamond scheme. Numerical Algorithms, 84(2): 761–779, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00778-8.

Murofushi:1992:ISE

[MN92] Makoto Murofushi and Hideko Nagasaka. On the internal stepsize of an extrapolation algorithm for IVP in ODE. Numerical Algorithms, 3(1–4):321–334, December 1992. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Minesaki:2001:DRT

[MN01] Yukitaka Minesaki and Yoshimasa Nakamura. The discrete relativistic Toda molecule equation and a Padé approximation algorithm. Numerical Algorithms, 27(3):219– 235, July 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/34/1/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/34/
1/fulltext.pdf.

Mittal:2008:SCS

[MN08] Ramesh Chand Mittal and Ruchi Nigam. Solution of a class of singular boundary value problems. Numerical Algorithms, 47(2):169–179, February 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=169.

Mukherjee:2011:UEE

[MN11] Kaushik Mukherjee and Srinivasan Natesan. ϵ-uniform error estimate of hybrid numerical scheme for singularly perturbed parabolic problems with interior layers. Numerical Algorithms, 58(1):103–141, September 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=1&spage=103.

Majee:2017:HVI

[MN17] P. Majee and C. Nahak. A hybrid viscosity iterative method with averaged mappings for split equilibrium problems and fixed point problems. Numerical Algorithms, 74(2):609–635, February 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0164-1.

Monjezi:2022:PBB

[MN22] N. Hoseini Monjezi and S. Nobakhtian. A proximal bundlebased algorithm for nonsmooth constrained multiobjective optimization problems with inexact data. Numerical Algorithms, 89(2):637–674, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01128-3.

Maji:2023:ANS

[MN23a] Sandip Maji and Srinivasan Natesan. Analytical and numerical solution techniques for a class of time-fractional integro-partial differential equations. Numerical Algorithms, 94(1):229–256, September 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01498-w.

Meng:2023:NMC

[MN23b] Chang Meng and James Nagy. Numerical methods for CT reconstruction with unknown geometry parameters. Numerical Algorithms, 92(1):831–847, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01451-3.

Monjezi:2023:NPB

[MN23c] N. Hoseini Monjezi and S. Nobakhtian. New proximal bundle algorithm based on the gradient sampling method for nonsmooth nonconvex optimization with exact and inexact information. Numerical Algorithms, 94(2):765–787, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01519-8.

Mirzaee:2023:ASN

[MNS23] Farshid Mirzaee, Shiva Naserifar, and Erfan Solhi. Accurate and stable numerical method based on the Floater– Hormann interpolation for stochastic Itô–Volterra integral equations. Numerical Algorithms, 94(1):275–292, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01500-5.

Murashige:2004:NVS

[MO04] Sunao Murashige and Shin’ichi Oishi. Numerical verification of solutions of periodic integral equations with a singular kernel. Numerical Algorithms, 37(1–4):301–310, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/24/abstract.htm.

McLachlan:2010:PDP

[MO10] Robert I. McLachlan and Dion R. J. O’Neale. Preservation and destruction of periodic orbits by symplectic integrators. Numerical Algorithms, 53(2–3):343–362, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=343.

McLachlan:2019:SIB

[MO19] Robert I. McLachlan and Christian Offen. Symplectic integration of boundary value problems. Numerical Algorithms, 81(4): 1219–1233, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mohanty:2010:UAA

[Moh10] Ranjan Kumar Mohanty. On the use of AGE algorithm with a high accuracy Numerov type variable mesh discretization for 1D non-linear parabolic equations. Numerical Algorithms, 54(3):379–393, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=379.

Mokhtary:2016:NTW

[Mok16] P. Mokhtary. Numerical treatment of a well-posed Chebyshev Tau method for Bagley–Torvik equation with high-order of accuracy. Numerical Algorithms, 72(4):875–891, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0072-9.

Monegato:1996:SNI

[Mon96] Giovanni Monegato. A stable Nyström interpolant for some Mellin convolution equations. Numerical Algorithms, 11(1–4): 271–283, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Monegato:2001:OCA

[Mon01] Giovanni Monegato. An overview of the computational aspects of Kronrod quadrature rules. Numerical Algorithms, 26(2):173–196, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
1/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/1/7/fulltext.pdf.

Moore:2007:IPS

[Moo07] Peter K. Moore. The impact of parameter selection on the performance of an automatic adaptive code for solving reaction-diffusion equations in three dimensions. Numerical Algorithms, 46(2):121–139, October 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=2&spage=121.

Moore:2020:DGI

[Moo20] Stephen Edward Moore. Discontinuous Galerkin isogeometric analysis for elliptic problems with discontinuous diffusion coefficients on surfaces. Numerical Algorithms, 85(3):1075– 1094, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00856-x.

Mortici:2011:NSS

[Mor11] Cristinel Mortici. A new Stirling series as continued fraction. Numerical Algorithms, 56(1):17–26, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=17.

Morikuni:2017:MMS

[Mor17] Keiichi Morikuni. Multistep matrix splitting iteration preconditioning for singular linear systems. Numerical Algorithms, 75 (2):457–475, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Moratis:2025:ABA

[Mor25] Kimon Moratis. Analysis of the basins of attraction generated by a pseudo Newton–Raphson method iterating over complex order derivatives. Numerical Algorithms, 100(3):1157–1178, November 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01984-9.

Motsa:2014:PUS

[Mot14] Sandile S. Motsa. On the practical use of the spectral homotopy analysis method and local linearisation method for unsteady boundary-layer flows caused by an impulsively stretching plate. Numerical Algorithms, 66(4):865–883, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9766-z.

Matos:1992:APC

[MP92] Ana C. Matos and Marc Prévost. Acceleration property for the columns of the E-algorithm. Numerical Algorithms, 2(3– 4):393–408, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mainar:1999:EAC

[MP99] E. Mainar and J. M. Peña. Error analysis of corner cutting algorithms. Numerical Algorithms, 22(1):41–52, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/21/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
21/1/fulltext.pdf.

MagolumongaMade:2000:ECT

[MP00] Mardochée Magolu monga Made and Ben Polman. Experimental comparison of three-dimensional point and line modified incomplete factorizations. Numerical Algorithms, 23(1):51–70, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
24/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/24/3/fulltext.pdf.

Mezher:2002:PRP

[MP02] Dany Mezher and Bernard Philippe. PAT — a reliable pathfollowing algorithm. Numerical Algorithms, 29(1–3):131–152, March 2002. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp008.lwwonline.
com/content/getfile/5058/38/9/abstract.htm;   http:
//ipsapp008.lwwonline.com/content/getfile/5058/38/
9/fulltext.pdf.

Mallat:2007:RBM

[MP07] Stéphane Mallat and Gabriel Peyré. A review of Bandlet methods for geometrical image representation. Numerical Algorithms, 44(3):205–234, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=3&spage=205.

Mandloi:2008:PTI

[MP08] A. Mandloi and A. K. Pathak. (0, 2) Pál-type interpolation on a circle in the complex plane involving Möbius transforms. Numerical Algorithms, 47(2):181–190, February 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=181.

Mollet:2013:EAN

[MP13] Christian Mollet and Roland Pabel. Efficient application of nonlinear stationary operators in adaptive wavelet methods — the isotropic case. Numerical Algorithms, 63(4):615–643, August 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9645-z.

Mishra:2014:DTM

[MP14] Vishnu Narayan Mishra and Prashantkumar Patel. The Durrmeyer type modification of the q-baskakov type operators with two parameter α and β. Numerical Algorithms, 67(4):753– 769, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9821-9.

Mishra:2022:CDS

[MP22] Soumyarani Mishra and Ambit K. Pany. Completely discrete schemes for 2D Sobolev equations with Burgers’ type nonlinearity. Numerical Algorithms, 90(3):963–987, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01218-2.

Marin:2025:BRT

[MP25] Liviu Marin and Andrei Tiberiu Pantea. Boundary reconstruction in two-dimensional steady-state anisotropic heat conduction. Numerical Algorithms, 98(3):1195–1230, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01831-x.

Morovati:2016:BBM

[MPB16] Vahid Morovati, Latif Pourkarimi, and Hadi Basirzadeh. Barzilai and Borwein’s method for multiobjective optimization problems. Numerical Algorithms, 72(3):539–604, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0058-7.

Marcellan:2012:QCP

[MPC12] Francisco Marcellán and Natalia C. Pinzón-Cortés. (1, 1) − qcoherent pairs. Numerical Algorithms, 60(2):223–239, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=2&
spage=223.

Mainar:2022:TPA

[MPR22] E. Mainar, J. M. Peña, and B. Rubio. Total positivity and accurate computations with Gram matrices of Bernstein bases. Numerical Algorithms, 91(2):841–859, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01284-0.

Mainar:2024:HRA

[MPR24] E. Mainar, J. M. Peña, and B. Rubio. High relative accuracy through Newton bases. Numerical Algorithms, 95(2):747– 772, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01588-9.

Mena:2020:GAS

[MPS20] Hermann Mena, Lena-Maria Pfurtscheller, and Tony Stillfjord. GPU acceleration of splitting schemes applied to differential matrix equations. Numerical Algorithms, 83(1):395–419, January 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-019-00687-w.pdf.

Meurant:2021:AEE

[MPT21] Gérard Meurant, Jan Papez, and Petr Tichý. Accurate error estimation in CG. Numerical Algorithms, 88(3):1337–1359, November 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01078-w.

Molina:1996:PBB

[MR96] Brigida Molina and Marcos Raydan. Preconditioned Barzilai– Borwein method for the numerical solution of partial differential equations. Numerical Algorithms, 13(1–2):45–60, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Molina:2002:SVK

[MR02] Brígida Molina and Marcos Raydan. Spectral variants of Krylov subspace methods. Numerical Algorithms, 29 (1–3):197–208, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.lwwonline.com/content/getfile/5058/38/
11/abstract.htm;   http://ipsapp008.lwwonline.com/
content/getfile/5058/38/11/fulltext.pdf.

Mansouri:2009:NFN

[MR09] H. Mansouri and C. Roos. A new full-Newton step O(n) infeasible interior-point algorithm for semidefinite optimization. Numerical Algorithms, 52(2):225–255, October 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=2&spage=225.

Mazzieri:2012:DAT

[MR12] Ilario Mazzieri and Francesca Rapetti. Dispersion analysis of triangle-based spectral element methods for elastic wave propagation. Numerical Algorithms, 60(4):631–650, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=4&
spage=631.

Mason:1993:OSB

[MRS93] J. C. Mason, G. Rodriguez, and S. Seatzu. Orthogonal splines based on B-splines—with applications to least squares, smoothing and regularisation problems. Numerical Algorithms, 5(1–4):25–40, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Morigi:2006:TPS

[MRS06] Serena Morigi, Lothar Reichel, and Fiorella Sgallari. A truncated projected SVD method for linear discrete ill-posed problems. Numerical Algorithms, 43(3):197–213, November 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=43&issue=3&
spage=197.

Morigi:2007:OPR

[MRS07] S. Morigi, L. Reichel, and F. Sgallari. Orthogonal projection regularization operators. Numerical Algorithms, 44(2):99–114, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
44&issue=2&spage=99.

Morigi:2010:NRC

[MRS10] S. Morigi, L. Reichel, and F. Sgallari. Noise-reducing cascadic multilevel methods for linear discrete ill-posed problems. Numerical Algorithms, 53(1):1–22, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=1&spage=1.

Mouhssine:2025:VEM

[MRS25] Abdellatif Mouhssine, Ahmed Ratnani, and Hassane Sadok. Vector extrapolation methods applied to geometric multigrid solvers for isogeometric analysis. Numerical Algorithms, 100 (4):1813–1836, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02132-
7.

Micchelli:1991:UREb

[MRU91] Charles A. Micchelli, Christophe Rabut, and Florencio I. Utreras. Using the refinement equation for the construction of pre-wavelets. III. Elliptic splines. Numerical Algorithms, 1(4): 331–351, November 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mach:2023:ACA

[MRV23] T. Mach, L. Reichel, and Marc Van Barel. Adaptive cross approximation for Tikhonov regularization in general form. Numerical Algorithms, 92(1):815–830, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01395-8.

Marcellan:1992:QDO

[MS92] Francisco Marcellán and Gabriela Sansigre. Quadratic decomposition of orthogonal polynomials: a matrix approach. Numerical Algorithms, 3(1–4):285–298, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Mulansky:1996:CMC

[MS96] Bernd Mulansky and Jochen W. Schmidt. Constructive methods in convex C^{2} interpolation using quartic splines. Numerical Algorithms, 12(1–2):111–124, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mazroui:2001:ECS

[MS01a] A. Mazroui and D. Sbibih. Existence and construction of Δ_{1}- splines of class C^{k} on a three-directional mesh. Numerical Algorithms, 26(1):21–48, January 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
4/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/4/3/fulltext.pdf.

Molinaro:2001:FMR

[MS01b] Anna Molinaro and Yaroslav D. Sergeyev. Finding the minimal root of an equation with the multiextremal and nondifferentiable left-hand part. Numerical Algorithms, 28(1– 4):255–272, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/37/
15/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/37/15/fulltext.pdf.

Miellou:2002:ORP

[MS02] J. C. Miellou and P. Spiteri. Optimization of the relaxation parameter for S.S.O.R. and A.D.I. preconditioning. Numerical Algorithms, 29(1–3):153–195, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.lwwonline.com/content/getfile/5058/38/
10/abstract.htm;   http://ipsapp008.lwwonline.com/
content/getfile/5058/38/10/fulltext.pdf.

Mehrmann:2006:THO

[MS06] Volker Mehrmann and Chunchao Shi. Transformation of high order linear differential-algebraic systems to first order. Numerical Algorithms, 42(3–4):281–307, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=3&spage=281.

Moreno:2011:IFP

[MS11] Joaquin Moreno and A. Saiz. Inverse functions of polynomials and its applications to initialize the search of solutions of polynomials and polynomial systems. Numerical Algorithms, 58(2):203–233, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=2&spage=203.

Motsa:2013:SMQ

[MS13] Sandile S. Motsa and Precious Sibanda. Some modifications of the quasilinearization method with higher-order convergence for solving nonlinear BVPs. Numerical Algorithms, 63(3): 399–417, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9629-z.

Meurant:2014:FVG

[MS14a] Gérard Meurant and Alvise Sommariva. Fast variants of the Golub and Welsch algorithm for symmetric weight functions in Matlab. Numerical Algorithms, 67(3):491–506, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9804-x.

Mirzaei:2014:SHC

[MS14b] Davoud Mirzaei and Robert Schaback. Solving heat conduction problems by the Direct Meshless Local Petrov–Galerkin (DMLPG) method. Numerical Algorithms, 65(2):275–291, February 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9711-1.

Mortici:2015:PAG

[MS15] Cristinel Mortici and Hari M. Srivastava. A product approximation of the gamma function. Numerical Algorithms, 69 (3):595–610, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9915-z.

Messaoudi:2017:RPI

[MS17] Abderrahim Messaoudi and Hassane Sadok. Recursive polynomial interpolation algorithm (RPIA). Numerical Algorithms, 76(3):675–694, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ma:2020:DDG

[MS20a] Guanglong Ma and Martin Stynes. A direct discontinuous Galerkin finite element method for convection-dominated twopoint boundary value problems. Numerical Algorithms, 83(2): 741–765, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mirchi:2020:NPE

[MS20b] Hamid Mirchi and Davod Khojasteh Salkuyeh. A new preconditioner for elliptic PDE-constrained optimization problems. Numerical Algorithms, 83(2):653–668, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Martinez:2022:TCA

[MS22] José Mario Martínez and Lúcio Tunes Santos. On two conjectures about Dennis–Moré conditions. Numerical Algorithms, 91(3):1407–1425, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01307-
w.

Maurya:2023:HOA

[MS23a] Rahul Kumar Maurya and Vineet Kumar Singh. A highorder adaptive numerical algorithm for fractional diffusion wave equation on non-uniform meshes. Numerical Algorithms, 92(3):1905–1950, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01372-1.

Messaoudi:2023:RNA

[MS23b] Abderrahim Messaoudi and Hassane Sadok. RMPIA: a new algorithm for computing the Lagrange matrix interpolation polynomials. Numerical Algorithms, 92(1):849–867, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01357-0.

Mirzaei:2023:FDM

[MS23c] Davoud Mirzaei and Navid Soodbakhsh. A fault detection method based on partition of unity and kernel approximation. Numerical Algorithms, 93(4):1759–1794, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01488-4.

Meng:2024:BEN

[MS24a] Xiangyun Meng and Martin Stynes. Balanced and energy norm error bounds for a spatial FEM with Crank–Nicolson and BDF2 time discretisation applied to a singularly perturbed reaction-diffusion problem. Numerical Algorithms, 95 (3):1155–1176, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01603-z.

Mustafa:2024:TDR

[MS24b] Ashif Mustafa and Manideepa Saha. A two-dimensional randomized extended Gauss–Seidel algorithm for solving least squares problems. Numerical Algorithms, 96(2):665–686, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01661-3.

Ma:2026:GBR

[MS26] Junqiao Ma and YouLin Shang. Global branch-reduction-bound algorithm to tackle the sum of the general affine ratios programming. Numerical Algorithms, 101(4):2687–2708, April 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02104-x.

Ma:2020:TSI

[MsC20] Wei Ma and Xiao shan Chen. Two-step inexact Newton-type method for inverse singular value problems. Numerical Algorithms, 84(3):847–870, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00783-x.

MacCarthy:1993:SMI

[MSCB93] B. L. MacCarthy, C. S. Syan, and M. Caulfield-Browne. Splines in motion — an introduction to MODUS and some unresolved approximation problems. Numerical Algorithms, 5(1–4):41–49, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Motsa:2012:NAM

[MSM12] Sandile S. Motsa, Precious Sibanda, and Gerald T. Marewo. On a new analytical method for flow between two inclined walls. Numerical Algorithms, 61(3):499–514, November 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=3&
spage=499.

Milosavljevic:2025:AGQ

[MSM25] Aleksandra S. Milosavljević, Marija P. Stanić, and Tatjana V. Tomović Mladenović. Anti-Gaussian quadrature rules related to orthogonality on the semicircle. Numerical Algorithms, 99(4):2137–2172, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01948-
z.

Motsa:2012:ISH

[MSMS12] Sandile Sydney Motsa, Stanford Shateyi, Gerald T. Marewo, and Precious Sibanda. An improved spectral homotopy analysis method for MHD flow in a semi-porous channel. Numerical Algorithms, 60(3):463–481, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=3&spage=463.

Mishra:2025:ODR

[MSP25] S. K. Mishra, D. Singh, and Pankaj. Optimality and duality results for fractional programming problems under Eunivexity. Numerical Algorithms, 98(3):1467–1480, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01840-w.

Monegato:2011:LCQ

[MSS11] Giovanni Monegato, Letizia Scuderi, and Marija P. Stanić. Lubich convolution quadratures and their application to problems described by space–time BIEs. Numerical Algorithms, 56(3):405–436, March 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=3&spage=405.

Messaoudi:2018:NAC

[MSS18] A. Messaoudi, R. Sadaka, and H. Sadok. New algorithm for computing the Hermite interpolation polynomial. Numerical Algorithms, 77(4):1069–1092, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mazroui:2003:BSS

[MST03] A. Mazroui, D. Sbibih, and A. Tijini. C^{k}–B-splines with square support on a three-direction mesh of the plane. Numerical Algorithms, 34(1):67–84, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.kluweronline.com/content/getfile/5058/
47/5/abstract.htm;   http://ipsapp008.kluweronline.
com/content/getfile/5058/47/5/fulltext.pdf.

Miyatake:2020:ASM

[MSZ20] Yuto Miyatake, Tomohiro Sogabe, and Shao-Liang Zhang. Adaptive SOR methods based on the Wolfe conditions. Numerical Algorithms, 84(1):117–132, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Model:1993:IPM

[MT93] R. Model and L. Trahms. An inverse problem of magnetic source localization. Numerical Algorithms, 5(1–4):603–610, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Macutan:1998:TFI

[MT98] Y. O. Macutan and G. Thomas. Theory of formal integrability and DAEs: Effective computations. Numerical Algorithms, 19(1–4):147–157, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
12/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/12/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Mazzia:2004:HMS

[MT04] Francesca Mazzia and Donato Trigiante. A hybrid mesh selection strategy based on conditioning for boundary value ODE problems. Numerical Algorithms, 36(2):169–187, June 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/54/A/5/abstract.htm.

Muhlbach:2006:CEB

[MT06] Günter W. Mühlbach and Yuehong Tang. Computing ECT-B-splines recursively. Numerical Algorithms, 41(1):35–78, January 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=41&
issue=1&spage=35.

Matthies:2012:TLL

[MT12] Gunar Matthies and Lutz Tobiska. A two-level local projection stabilisation on uniformly refined triangular meshes. Numerical Algorithms, 61(3):465–478, November 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=465.

Meurant:2013:CQB

[MT13] Gérard Meurant and Petr Tichý. On computing quadraturebased bounds for the A-norm of the error in conjugate gradients. Numerical Algorithms, 62(2):163–191, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9591-9/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=163-191. See erratum [MT14].

Meurant:2014:ECQ

[MT14] Gérard Meurant and Petr Tichý. Erratum to: On computing quadrature-based bounds for the A-norm of the error in conjugate gradients. Numerical Algorithms, 66(3):679–680, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11075-014-9868-2.pdf. See [MT13].

Meurant:2015:REP

[MT15] Gérard Meurant and Jurjen Duintjer Tebbens. The role eigenvalues play in forming GMRES residual norms with nonnormal matrices. Numerical Algorithms, 68(1):143–165, January 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9891-3.

McCoid:2018:PSM

[MT18] Conor McCoid and Manfred R. Trummer. Preconditioning of spectral methods via Birkhoff interpolation. Numerical Algorithms, 79(2):555–573, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Meurant:2019:AER

[MT19] Gérard Meurant and Petr Tichý. Approximating the extreme Ritz values and upper bounds for the A-norm of the error in CG. Numerical Algorithms, 82(3):937–968, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Meurant:2023:BGR

[MT23] Gérard Meurant and Petr Tichý. The behavior of the Gauss–Radau upper bound of the error norm in CG. Numerical Algorithms, 94(2):847–876, October 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01522-z.

Tang:2014:FSG

[mTLbJlL14] Chun ming Tang, Shuai Liu, Jin bao Jian, and Jian ling Li. A feasible SQP–GS algorithm for nonconvex, nonsmooth constrained optimization. Numerical Algorithms, 65(1):1–22, January 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9692-5.

Manh:2022:TTH

[MTTC22] Phung Van Manh, Nguyen Van Trao, Phan Thanh Tung, and Le Ngoc Cuong. Taylor type and Hermite type interpolants in R^{n}. Numerical Algorithms, 89(1):145–166, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01109-6.

Muhlbach:1999:RRG

[Müh99] G. Mühlbach. A recurrence relation for generalized divided differences with respect to ECT-systems. Numerical Algorithms, 22(3–4):317–326, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
23/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/23/4/fulltext.pdf.

Muller:2000:SEC

[Mül00] Jürgen Müller. Series expansions for computing Bessel functions of variable order on bounded intervals. Numerical Algorithms, 24(3):299–308, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Macdougall:2002:GEE

[MV02] T. Macdougall and J. H. Verner. Global error estimators for order 7, 8 Runge–Kutta pairs. Numerical Algorithms, 31(1–4):215–231, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
43/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/43/3/fulltext.pdf.

Mallory:2013:CEH

[MV13] Kristina Mallory and Robert A. Van Gorder. Control of error in the homotopy analysis of solutions to the Zakharov system with dissipation. Numerical Algorithms, 64(4):633– 657, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9683-6.

Mallory:2014:OHA

[MV14] Kristina Mallory and Robert A. Van Gorder. Optimal homotopy analysis and control of error for solutions to the nonlocal Whitham equation. Numerical Algorithms, 66(4):843– 863, August 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9765-0.

Messina:2017:SCS

[MV17] Eleonora Messina and Antonia Vecchio. A sufficient condition for the stability of direct quadrature methods for Volterra integral equations. Numerical Algorithms, 74(4):1223–1236, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0193-9.

Moutier:2021:DSN

[MVG21] François Moutier, Arnaud Vandaele, and Nicolas Gillis. Offdiagonal symmetric nonnegative matrix factorization. Numerical Algorithms, 88(2):939–963, October 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-01063-9.

Mullen:2009:VPA

[MvS09] Katharine M. Mullen and Ivo H. M. van Stokkum. The variable projection algorithm in time-resolved spectroscopy, microscopy and mass spectrometry applications. Numerical Algorithms, 51(3):319–340, July 2009. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=51&issue=3&spage=319.

Tributes to Gene H. Golub, Part III.

Mastronardi:2005:DCA

[MVV05a] N. Mastronardi, E. Van Camp, and M. Van Barel. Divide and conquer algorithms for computing the eigendecomposition of symmetric diagonal-plus-semiseparable matrices. Numerical Algorithms, 39(4):379–398, August 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mastronardi:2005:NRC

[MVV05b] Nicola Mastronardi, Marc Van Barel, and Raf Vandebril. A note on the recursive calculation of dominant singular subspaces. Numerical Algorithms, 38(4):237–242, April 2005. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mastronardi:2025:CZC

[MVV25] N. Mastronardi, M. Van Barel, and P. Van Dooren. On computing the zeros of a class of Sobolev orthogonal polynomials. Numerical Algorithms, 100(4):1765–1781, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02106-9.

Martin-Vaquero:2008:EFG

[MVVA08] J. Martín-Vaquero and J. Vigo-Aguiar. Exponential fitted Gauss, Radau and Lobatto methods of low order. Numerical Algorithms, 48(4):327–346, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=327.

Mastronardi:2024:RSP

[MVVV24] Nicola Mastronardi, Marc Van Barel, Raf Vandebril, and Paul Van Dooren. Rational QZ steps with perfect shifts. Numerical Algorithms, 95(3):1079–1102, March 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01600-2.

Mattheij:1998:SSL

[MW98] R. M. M. Mattheij and P. M. E. J. Wijckmans. Sensitivity of solutions of linear DAE to perturbations of the system matrices. Numerical Algorithms, 19(1–4):159–171, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/16/10/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/16/

10/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Meng:2016:BTE

[MW16] Wenhui Meng and Liantang Wang. Bounds for truncation errors of Graf’s and Neumann’s addition theorems. Numerical Algorithms, 72(1):91–106, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0035-1.

Marples:2024:PAU

[MW24] Callum Robert Marples and Philip Michael Williams. Patch area and uniform sampling on the surface of any ellipsoid.

Numerical Algorithms, 95(4):1801–1827, April 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01628-4. See correction [MW26].

Marples:2026:CPA

[MW26] Callum Robert Marples and Philip Michael Williams. Correction to: Patch area and uniform sampling on the surface of any ellipsoid. Numerical Algorithms, 102(1):749–750, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02150-5. See [MW24].

Manebo:2024:CAS

[MWDD24] Wubeshet Seyoum Manebo, Mesfin Mekuria Woldaregay, Tekle Gemechu Dinka, and Gemechis File Duressa. A computational approach to solving a second-order singularly perturbed Fredholm integro-differential equation with discontinuous source term. Numerical Algorithms, 97(3):1415–1430, November 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01756-5.

Ma:2019:SFP

[MWsC19] Zhaoli Ma, Lin Wang, and Shih sen Chang. On the split feasibility problem and fixed point problem of quasi-φ-nonexpansive mapping in Banach spaces. Numerical Algorithms, 80(4):1203–1218, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Mu:2013:CSW

[MWWY13] Lin Mu, Junping Wang, Yanqiu Wang, and Xiu Ye. A computational study of the weak Galerkin method for second-order elliptic equations. Numerical Algorithms, 63(4):753–777, August 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9651-1.

Miao:2023:SOB

[MWZL23] Ni Miao, Danxia Wang, Haifeng Zhang, and Jing Liu. A second-order BDF convex splitting numerical scheme for the Ericksen–Leslie equation. Numerical Algorithms, 94(1):293– 314, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01501-4.

Meng:2025:VPU

[MXL25] Lingsheng Meng, Kailiang Xin, and Jun Li. Variable parameter Uzawa method for solving the indefinite least squares problem. Numerical Algorithms, 99(2):1043–1061, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01905-w.

Martinez:2004:P

[MY04] José Mario Martínez and Jin Yun Yuan. Preface. Numerical Algorithms, 35(2–4):137, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/50/A/1/abstract.htm.

Ma:2022:EEF

[MY22] Ying Ma and Jia Yin. Error estimates of finite difference methods for the Dirac equation in the massless and nonrelativistic regime. Numerical Algorithms, 89(4):1415–1440, April 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01159-w.

Mikram:1999:CNF

[MZ99] Jilali Mikram and Fouad Zinoun. Computation of normal forms of Hamiltonian systems in the presence of Poisson commuting integrals. Numerical Algorithms, 21(1– 4):287–310, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
23/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/23/fulltext.pdf.

Migorski:2019:RMN

[MZ19] Stanisław Migórski and Shengda Zeng. Rothe method and numerical analysis for history-dependent hemivariational inequalities with applications to contact mechanics. Numerical Algorithms, 82(2):423–450, October 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-019-00667-0.pdf.

Ma:2026:OSB

[MZG26] Suxia Ma, Bo Zhang, and Yuelin Gao. Output-space branchand-bound algorithm adopting an adaptive branching rule for solving general linear fractional-multiplicative programs. Numerical Algorithms, 102(1):467–504, May 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02137-2.

Meng:2020:CNM

[MZW20] Lingsheng Meng, Bing Zheng, and Yimin Wei. Condition numbers of the multidimensional total least squares problems having more than one solution. Numerical Algorithms, 84(3): 887–908, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00785-9.

Meng:2025:ERS

[MZZ25] Zixiang Meng, Zhihan Zhou, and Fengxia Zhang. An effective real structure-preserving algorithm for the quaternion indefinite least squares problem. Numerical Algorithms, 99(4): 1697–1742, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01929-2.

Nosrati:2024:NSS

[NA24] Mahsa Nosrati and Keyvan Amini. A new structured spectral conjugate gradient method for nonlinear least squares problems. Numerical Algorithms, 97(2):897–914, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01729-0.

Nwaigwe:2026:FOP

[NA26] Chinedu Nwaigwe and Abdon Atangana. Fourth-order predictor-corrector method for initial value ordinary differential equation problems. Numerical Algorithms, 101(2): 1303–1336, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02043-7.

Noeiaghdam:2019:FOC

[NAA19] Samad Noeiaghdam, Mohammad Ali Fariborzi Araghi, and Saeid Abbasbandy. Finding optimal convergence control parameter in the homotopy analysis method to solve integral equations based on the stochastic arithmetic. Numerical Algorithms, 81(1):237–267, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Nachaoui:1999:ISD

[Nac99] Abdeljalil Nachaoui. Iterative solution of the drift–diffusion equations. Numerical Algorithms, 21(1–4):323–341, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/20/17/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/20/
17/fulltext.pdf.

Nachaoui:2003:III

[Nac03] Abdeljalil Nachaoui. An improved implementation of an iterative method in boundary identification problems. Numerical Algorithms, 33(1–4):381–398, August 2003. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/46/31/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/31/fulltext.pdf.

Nikazad:2022:NSS

[NAE22] T. Nikazad, M. Abbasi, and T. Elfving. A new step size rule for the superiorization method and its application in computerized tomography. Numerical Algorithms, 90(3):1253–1277, July 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01229-z.

Nasu:2021:TGF

[NAHZ21] Nasrin Jahan Nasu, Md. Abdullah Al Mahbub, Shahid Hussain, and Haibiao Zheng. Two-grid finite element method for the dual-permeability-Stokes fluid flow model. Numerical Algorithms, 88(4):1703–1731, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01091-
z.

Nachaoui:2005:PSD

[NAR05a] A. Nachaoui, J. Abouchabaka, and N. Rafalia. Parallel solvers for the depletion region identification in metal semiconductor field effect transistors. Numerical Algorithms, 40(2):187–199, October 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=187.

Narcowich:2005:RDE

[Nar05b] Francis J. Narcowich. Recent developments in error estimates for scattered-data interpolation via radial basis functions. Numerical Algorithms, 39(1–3):307–315, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Natsis:2007:SWS

[Nat07] D. G. Natsis. Solitary wave solutions of the one-dimensional Boussinesq equations. Numerical Algorithms, 44(3):281–289, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
44&issue=3&spage=281.

Nguyen-Ba:2016:VSH

[NB16] Truong Nguyen-Ba. On variable step Hermite–Birkhoff solvers combining multistep and 4-stage DIRK methods for stiff ODEs. Numerical Algorithms, 71(4):855–888, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0027-1.

Ndukum:2017:FTF

[NBJA17] P. L. Ndukum, T. A. Biala, S. N. Jator, and R. B. Adeniyi. On a family of trigonometrically fitted extended backward differentiation formulas for stiff and oscillatory initial value problems. Numerical Algorithms, 74(1):267–287, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0148-1.

Narang:2017:NED

[NBK17] Mona Narang, Saurabh Bhatia, and Vinay Kanwar. New efficient derivative free family of seventh-order methods for solving systems of nonlinear equations. Numerical Algorithms, 76(1): 283–307, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ng:1994:FIM

[NC94] Michael K. Ng and Raymond H. Chan. Fast iterative methods for least squares estimations. Numerical Algorithms, 6(3–4): 353–378, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Nguyen:2011:CCR

[NCC11] Hoa Thang Nguyen, Annie Cuyt, and Oliver Salazar Celis. Comonotone and coconvex rational interpolation and approximation. Numerical Algorithms, 58(1):1–21, September 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=58&issue=1&
spage=1.

Nedaiasl:2021:GFE

[ND21] Khadijeh Nedaiasl and Raziyeh Dehbozorgi. Galerkin finite element method for nonlinear fractional differential equations. Numerical Algorithms, 88(1):113–141, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01032-2.

Neher:2004:CMV

[Neh04] M. Neher. A complex mean value form for curves. Numerical Algorithms, 37(1–4):337–343, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/27/abstract.htm.

Nik:2014:SHA

[NEMS14] H. Saberi Nik, S. Effati, S. S. Motsa, and M. Shirazian. Spectral homotopy analysis method and its convergence for solving a class of nonlinear optimal control problems. Numerical Algorithms, 65(1):171–194, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9700-4.

Nwawuru:2026:SAS

[NEY26] Francis O Nwawuru, Jeremiah N. Ezeora, and Jen-Chih Yao. Self-adaptive subgradient extragradient algorithm for solving equilibrium and fixed point problems in Hilbert spaces. Numerical Algorithms, 101(3):1567–1605, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02055-3.

Nguyen:2023:IPT

[NG23] Van-Thanh Nguyen and Laura Grigori. Interpretation of parareal as a two-level additive Schwarz in time preconditioner and its acceleration with GMRES. Numerical Algorithms, 94 (1):29–72, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01492-8.

Ngondiep:2023:HON

[Ngo23] Eric Ngondiep. A high-order numerical scheme for multidimensional convection–diffusion-reaction equation with timefractional derivative. Numerical Algorithms, 94(2):681–700, October 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01516-x.

Nguyen:2016:VQB

[Ngu16] Quang Van Nguyen. Variable quasi-Bregman monotone sequences. Numerical Algorithms, 73(4):1107–1130, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0132-9.

Najafi:2024:IRC

[NH24] Shahabeddin Najafi and Masoud Hajarian. An improved Riemannian conjugate gradient method and its application to robust matrix completion. Numerical Algorithms, 96(4): 1887–1900, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01688-6.

Nedzhibov:2006:SFM

[NHP06] Gyurhan H. Nedzhibov, Vejdi I. Hasanov, and Milko G. Petkov. On some families of multi-point iterative methods for solving nonlinear equations. Numerical Algorithms, 42(2):127– 136, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
42&issue=2&spage=127.

Niethammer:1993:NIS

[Nie93] Wilhelm Niethammer. A note on the implementation of the successive overrelaxation method for linear complementarity problems. Numerical Algorithms, 4(1–2):197–200, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Niethammer:2000:ONS

[Nie00] Wilhelm Niethammer. Old and new from SOR. Numerical Algorithms, 25(1–4):263–277, September 2000. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/3/17/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/17/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Nagata:2012:EAM

[NIN12] Munehiro Nagata, Masashi Iwasaki, and Yoshimasa Nakamura. Error analysis of the mdLVs algorithm for computing bidiagonal singular values. Numerical Algorithms, 61(2):261–274, October 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=61&
issue=2&spage=261.

Ngwane:2013:BHM

[NJ13] F. F. Ngwane and S. N. Jator. Block hybrid method using trigonometric basis for initial value problems with oscillating solutions. Numerical Algorithms, 63(4):713–725, August 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9649-8.

Nikazad:2016:CNE

[NK16] T. Nikazad and M. Karimpour. Controlling noise error in block iterative methods. Numerical Algorithms, 73(4):907– 925, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0122-y.

Nikazad:2021:COA

[NK21] T. Nikazad and M. Karimpour. Column-oriented algebraic iterative methods for nonnegative constrained least squares problems. Numerical Algorithms, 86(3):1265–1284, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00932-7.

Neumaier:2024:GLC

[NKA24] Arnold Neumaier, Morteza Kimiaei, and Behzad Azmi. Globally linearly convergent nonlinear conjugate gradients without Wolfe line search. Numerical Algorithms, 97(4):1607–1633, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01764-5.

Ngoc:2025:MPI

[NKM25] Nguyen Anh Ngoc, Nguyen Van Khiem, and Phung Van Manh. Multivariate polynomial interpolation based on Radon projections. Numerical Algorithms, 99(4):1925–1941, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01938-1.

Nedialkov:2004:IAA

[NKS04] Nedialko S. Nedialkov, Vladik Kreinovich, and Scott A. Starks. Interval arithmetic, affine arithmetic, Taylor series methods: Why, what next? Numerical Algorithms, 37(1–4):325–336, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/26/abstract.htm.

Nigro:1997:PSI

[NL97] Abdelmalek Nigro and Pierre-Jean Laurent. Progressive stable interpolation. Numerical Algorithms, 14(4):343–359, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/6/4/abstract.htm;   http:/

/ipsapp007.kluweronline.com/content/getfile/5058/6/
4/fulltext.pdf.

Nemati:2021:NSC

[NLT21] S. Nemati, Pedro M. Lima, and Delfim F. M. Torres. Numerical solution of a class of third-kind Volterra integral equations using Jacobi wavelets. Numerical Algorithms, 86(2):675– 691, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00906-9.

Naseri:2014:OTS

[NM14] R. Naseri and A. Malek. Optimization technique in solving laminar boundary layer problems with temperature dependent viscosity. Numerical Algorithms, 66(3):663–678, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9753-4.

Naserifar:2026:ICM

[NM26] Shiva Naserifar and Farshid Mirzaee. Innovative computational model for variable order fractional Brownian motion and solving stochastic differential equations. Numerical Algorithms, 101(4):2215–2244, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02083-z.

Nhan:2018:BLP

[NMM18] Thái Anh Nhan, Scott MacLachlan, and Niall Madden. Boundary layer preconditioners for finite-element discretizations of singularly perturbed reaction–diffusion problems. Numerical Algorithms, 79(1):281–310, September 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Nayak:2025:RSS

[NMM25] Snigdhashree Nayak, Debasisha Mishra, and Nachiketa Mishra. A rotated shift-splitting method for complex symmetric linear systems. Numerical Algorithms, 98(1):85–115, January 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01786-z.

Noferini:2026:FNS

[NN26] Vanni Noferini and Lauri Nyman. Finding the nearest Ω-stable pencil with Riemannian optimization. Numerical Algorithms, 102(1):569–592, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02140-7.

Nigam:2023:SRN

[NNCN23] Ritu Nigam, Nilofar Nahid, Samiran Chakraborty, and Gnaneshwar Nelakanti. Superconvergence results for nonlinear Hammerstein integral equations on unbounded domain. Numerical Algorithms, 94(3):1243–1279, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01534-9.

Norsett:1996:DOP

[NØ96] Syvert P. Nørsett and Bente Østigård. Dual-orthogonal polynomials. Numerical Algorithms, 11(1–4):311–326, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Norfolk:2000:APS

[Nor00] Timothy S. Norfolk. Asymptotics of the partial sums of a set of integral transforms. Numerical Algorithms, 25(1– 4):279–291, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/16/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/16/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Notaris:1995:SPG

[Not95] Sotirios E. Notaris. Stieltjes polynomials and Gauss–Kronrod quadrature formulae for measures induced by Chebyshev polynomials. Numerical Algorithms, 10(1–2):167–186, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Notaris:2008:PWI

[Not08] Sotirios E. Notaris. Positivity of the weights of interpolatory quadrature formulae with Bernstein–Szegő abscissae. Numerical Algorithms, 49(1–4):315–329, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=315.

Notaris:2012:ENQ

[Not12] Sotirios E. Notaris. The error norm of quadrature formulae. Numerical Algorithms, 60(4):555–578, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=555.

Notay:2022:RCP

[Not22] Yvan Notay. Rigorous convergence proof of space-time multigrid with coarsening in space. Numerical Algorithms, 89(2): 675–699, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01129-2.

Novakovic:2026:AEK

[Nov26] Vedran Novaković. Arithmetical enhancements of the Kogbetliantz method for the SVD of order two. Numerical Algorithms, 101(2):1131–1156, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02035-7.

Nowak:2006:MCA

[Now06] Rafal Nowak. A method of convergence acceleration of some continued fractions. Numerical Algorithms, 41(3):297–317, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
41&issue=3&spage=297.

Nowak:2013:MCA

[Now13] Rafal Nowak. A method of convergence acceleration of some continued fractions II. Numerical Algorithms, 63(4):573–600, August 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-012-9642-2.pdf.

Nowak:2019:CAA

[Now19] Rafal Nowak. Convergence acceleration of alternating series. Numerical Algorithms, 81(2):591–608, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Navas-Palencia:2018:FAA

[NP18] Guillermo Navas-Palencia. Fast and accurate algorithm for the generalized exponential integral E_{ν}(x) for positive real order. Numerical Algorithms, 77(2):603–630, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-017-0331-z.

Nigam:2022:SEM

[NP22] Nilima Nigam and Sara Pollock. A simple extrapolation method for clustered eigenvalues. Numerical Algorithms, 89(1): 115–143, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01108-7.

Nagy:2004:IMI

[NPP04] James G. Nagy, Katrina Palmer, and Lisa Perrone. Iterative methods for image deblurring: a Matlab object-oriented approach. Numerical Algorithms, 36(1):73–93, May 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/52/A/5/abstract.htm.

Natalini:2008:DPL

[NPR08] Pierpaolo Natalini, Roberto Patrizi, and Paolo E. Ricci. The Dirichlet problem for the Laplace equation in a starlike domain of a Riemann surface. Numerical Algorithms, 49(1–4):299–313, December 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=299.

Neitzel:2009:STD

[NPS09] Ira Neitzel, Uwe Prüfert, and Thomas Slawig. Strategies for time-dependent PDE control with inequality constraints using an integrated modeling and simulation environment. Numerical Algorithms, 50(3):241–269, March 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=3&spage=241.

Noschese:2012:IPR

[NR12] Silvia Noschese and Lothar Reichel. Inverse problems for regularization matrices. Numerical Algorithms, 60(4):531–544, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=60&
issue=4&spage=531.

Novati:2014:AAT

[NR14] Paolo Novati and Maria Rosaria Russo. Adaptive Arnoldi– Tikhonov regularization for image restoration. Numerical Algorithms, 65(4):745–757, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9712-0.

Noschese:2024:EMG

[NR24a] Silvia Noschese and Lothar Reichel. Enhancing multiplex global efficiency. Numerical Algorithms, 96(1):397–416, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01651-5.

Noschese:2024:NAA

[NR24b] Silvia Noschese and Lothar Reichel. Network analysis with the aid of the path length matrix. Numerical Algorithms, 95(1):451–470, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01577-y.

Noschese:2025:CMT

[NR25a] Silvia Noschese and Lothar Reichel. Communication in multiplex transportation networks. Numerical Algorithms, 100(4): 1439–1458, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01943-4.

Noschese:2025:EIC

[NR25b] Silvia Noschese and Lothar Reichel. Edge importance in complex networks. Numerical Algorithms, 99(1):377–410, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01881-1.

Neuman:2012:ARR

[NRS12] Arthur Neuman, Lothar Reichel, and Hassane Sadok. Algorithms for range restricted iterative methods for linear discrete ill-posed problems. Numerical Algorithms, 59(2):325–331, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
59&issue=2&spage=325.

Ngo:2023:FOC

[NRV23] Hoa T. B. Ngo, Mohsen Razzaghi, and Thieu N. Vo. Fractionalorder Chelyshkov wavelet method for solving variable-order fractional differential equations and an application in variableorder fractional relaxation system. Numerical Algorithms, 92 (3):1571–1588, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01354-3.

Nouisser:2001:ECS

[NS01] O. Nouisser and D. Sbibih. Existence and construction of simple B-splines of class C^{k} on a four-directional mesh of the plane. Numerical Algorithms, 27(4):329–358, August 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/36/2/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/36/
2/fulltext.pdf.

Novakovic:2022:KTA

[NS22] Vedran Novaković and Sanja Singer. A Kogbetliantz-type algorithm for the hyperbolic SVD. Numerical Algorithms, 90 (2):523–561, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01197-4.

Naghdi:2024:TNA

[NS24] Niloofar Naghdi and Sedaghat Shahmorad. Theoretical and numerical analysis of a first-kind linear Volterra functional integral equation with weakly singular kernel and vanishing delay. Numerical Algorithms, 97(1):453–473, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01710-x.

Nandi:2020:TSA

[NSM20] Ashish Kumar Nandi, Jajati Keshari Sahoo, and Debasisha Mishra. Three-step alternating iterations for index 1 and nonsingular matrices. Numerical Algorithms, 84(2):457–483, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00764-0.

Nouisser:2003:FSQ

[NSS03] O. Nouisser, D. Sbibih, and Paul Sablonnière. A family of spline quasi-interpolants on the sphere. Numerical Algorithms, 33(1–4):399–413, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
32/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/32/fulltext.pdf.

Noupelah:2021:OSC

[NT21] Aurelien Junior Noupelah and Antoine Tambue. Optimal strong convergence rates of some Euler-type timestepping schemes for the finite element discretization SPDEs driven by additive fractional Brownian motion and Poisson random measure. Numerical Algorithms, 88(1):315–363, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-01041-1;   http://link.springer.
com/content/pdf/10.1007/s11075-020-01041-1.pdf.

Nawaz:2025:CAP

[NUG25] Bashir Nawaz, Kifayat Ullah, and Krzysztof Gdawiec. Convergence analysis of Picard–SP iteration process for generalized α-nonexpansive mappings. Numerical Algorithms, 98(4):1943– 1964, April 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01859-z.

Nhan:2021:BMC

[NV21] Thái Anh Nhan and Relja Vulanović. The Bakhvalov mesh: a complete finite-difference analysis of two-dimensional singularly perturbed convection-diffusion problems. Numerical Algorithms, 87(1):203–221, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00964-z.

Nakao:2004:EAN

[NW04] Mitsuhiro T. Nakao and Yoshitaka Watanabe. An efficient approach to the numerical verification for solutions of elliptic differential equations. Numerical Algorithms, 37(1–4):311–323, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/25/abstract.htm.

Nowak:2017:NPC

[NW17] Rafał Nowak and Paweł Woźny. New properties of a certain method of summation of generalized hypergeometric series. Numerical Algorithms, 76(2):377–391, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Naumann:2019:MFS

[NW19] Andreas Naumann and Jörg Wensch. Multirate finite step methods. Numerical Algorithms, 81(4):1547–1571, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-019-00763-1.

Niu:2026:SKM

[NYL26] Yu-Qi Niu, Susu Yin, and Peng Lv. A sampling Kaczmarz– Motzkin algorithm for tensor linear systems. Numerical Algorithms, 102(2):1185–1206, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02174-
x.

Ng:2019:SMC

[NZ19] Michael K. Ng and Zhaochen Zhu. Sparse matrix computation for air quality forecast data assimilation. Numerical Algorithms, 80(3):687–707, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Nadeem:2011:ETR

[NZF11] Sohail Nadeem, Shehla Zaheer, and Tiegang Fang. Effects of thermal radiation on the boundary layer flow of a Jeffrey fluid over an exponentially stretching surface. Numerical Algorithms, 57(2):187–205, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=2&spage=187.

Oyewole:2022:TEM

[OAMA22] O. K. Oyewole, H. A. Abass, A. A. Mebawondu, and K. O. Aremu. A Tseng extragradient method for solving variational inequality problems in Banach spaces. Numerical Algorithms, 89(2):769–789, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01133-6.

Oara:1994:PDS

[Oar94] Cristian Oara. Proper deflating subspaces: properties, algorithms and applications. Numerical Algorithms, 7(2–4):355– 373, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Oviedo:2022:FOW

[OAR22] Harry Oviedo, Roberto Andreani, and Marcos Raydan. A family of optimal weighted conjugate-gradient-type methods for strictly convex quadratic minimization. Numerical Algorithms, 90(3):1225–1252, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01228-0.

Ozyapici:2016:FPR

[OB16] Ali Ozyapici and Bülent Bilgehan. Finite product representation via multiplicative calculus and its applications to exponential signal processing. Numerical Algorithms, 71(2):475– 489, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0004-8.

Ortiz-Bernardin:2019:VEO

[OBAHK^{+}19] A. Ortiz-Bernardin, C. Alvarez, N. Hitschfeld-Kahler, A. Russo, R. Silva-Valenzuela, and E. Olate-Sanzana. Veamy: an extensible object-oriented C++ library for the virtual element method. Numerical Algorithms, 82(4):1189–1220, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Oktay:2024:MPR

[OC24] Eda Oktay and Erin Carson. Mixed precision Rayleigh quotient iteration for total least squares problems. Numerical Algorithms, 96(2):777–798, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01665-
z.

Ou:2024:CFS

[OCWV24a] Caixia Ou, Dakang Cen, Zhibo Wang, and Seakweng Vong. Correction to: Fitted schemes for Caputo–Hadamard fractional differential equations. Numerical Algorithms, 97(1):165, September 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01748-x. See [OCWV24b].

Ou:2024:FSC

[OCWV24b] Caixia Ou, Dakang Cen, Zhibo Wang, and Seakweng Vong. Fitted schemes for Caputo–Hadamard fractional differential equations. Numerical Algorithms, 97(1):135–164, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01696-6. See correction [OCWV24a].

Odlyzko:2000:IBB

[Odl00] A. M. Odlyzko. An improved bound for the de Bruijn–Newman constant. Numerical Algorithms, 25(1–4):293–303, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/15/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/15/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Oviedo:2021:TAS

[ODL21] Harry Oviedo, Oscar Dalmau, and Hugo Lara. Two adaptive scaled gradient projection methods for Stiefel manifold constrained optimization. Numerical Algorithms, 87(3):1107– 1127, July 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01001-9.

Oliveira:2013:NAT

[OdZdRV13] S. P. Oliveira, A. Ruiz de Zárate, A. C. da Rocha, and D. G. Alfaro Vigo. A note on the alternate trapezoidal quadrature method for Fredholm integral eigenvalue problems. Numerical Algorithms, 62(4):601–614, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9681-8.

Okuonghae:2014:SDG

[OI14] R. I. Okuonghae and M. N. O. Ikhile. Second derivative general linear methods. Numerical Algorithms, 67(3):637–654, November 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9814-8.

Ogwo:2021:IMF

[OIM21] G. N. Ogwo, C. Izuchukwu, and O. T. Mewomo. Inertial methods for finding minimum-norm solutions of the split variational inequality problem beyond monotonicity. Numerical Algorithms, 88(3):1419–1456, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01081-
1.

Ouermi:2023:EBH

[OKB23] T. A. J. Ouermi, Robert M. Kirby, and Martin Berzins. ENObased high-order data-bounded and constrained positivitypreserving interpolation. Numerical Algorithms, 92(3):1517– 1551, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01352-5.

Omelyan:2021:ANS

[OKP21] Igor Omelyan, Yuri Kozitsky, and Krzysztof Pilorz. Algorithm for numerical solutions to the kinetic equation of a spatial population dynamics model with coalescence and repulsive jumps. Numerical Algorithms, 87(2):895–919, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-020-00992-9;   http://link.springer.
com/content/pdf/10.1007/s11075-020-00992-9.pdf.

Ou:2021:NAZ

[OL21] Yigui Ou and Haichan Lin. A neurodynamic approach to zero-one quadratic programming. Numerical Algorithms, 88 (3):1251–1274, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01075-z.

Ou:2023:UCA

[OL23] Yigui Ou and Lin Li. A unified convergence analysis of the derivative-free projection-based method for constrained nonlinear monotone equations. Numerical Algorithms, 93(4): 1639–1660, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01483-9.

Olszanskyj:1994:RMM

[OLB94] Serge J. Olszanskyj, James M. Lebak, and Adam W. Bojańczyk. Rank-k modification methods for recursive least squares problems. Numerical Algorithms, 7(2–4):325–354, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ong:2018:PIN

[OM18] Benjamin W. Ong and Bankim C. Mandal. Pipeline implementations of Neumann–Neumann and Dirichlet–Neumann waveform relaxation methods. Numerical Algorithms, 78(1):1–20, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ouyang:2026:EBC

[OM26] Chunping Ouyang and Junjie Ma. Efficient block collocation methods for highly oscillatory Volterra integral equations with Bessel kernels. Numerical Algorithms, 102(2):1111–1133, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02172-z.

Ofem:2024:RTS

[OMU^{+}24] Austine Efut Ofem, Akindele Adebayo Mebawondu, Godwin Chidi Ugwunnadi, Prasit Cholamjiak, and Ojen Kumar Narain. Relaxed Tseng splitting method with double inertial steps for solving monotone inclusions and fixed point problems. Numerical Algorithms, 96(4):1465–1498, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01674-y.

Oh:2021:FEM

[OMW21] Minah Oh, Lina Ma, and Kening Wang. P_{1} finite element methods for a weighted elliptic state-constrained optimal control problem. Numerical Algorithms, 87(1):1–17, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00955-0.

Ojha:2026:FOC

[ON26] Ramesh Ojha and Srinivasan Natesan. A fourth-order collocation scheme for generalized Black–Scholes model in financial decision-making. Numerical Algorithms, 102(1):221– 245, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02128-3.

Ozaki:2022:GTM

[OO22] Katsuhisa Ozaki and Takeshi Ogita. Generation of test matrices with specified eigenvalues using floating-point arithmetic. Numerical Algorithms, 90(1):241–262, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01186-7.

Ozaki:2011:AAS

[OOO11] Katsuhisa Ozaki, Takeshi Ogita, and Shiníchi Oishi. An algorithm for automatically selecting a suitable verification method for linear systems. Numerical Algorithms, 56(3):363– 382, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
56&issue=3&spage=363.

Ozaki:2012:EFT

[OOOR12] Katsuhisa Ozaki, Takeshi Ogita, Shin’ichi Oishi, and Siegfried M. Rump. Error-free transformations of matrix multiplication by using fast routines of matrix multiplication and its applications. Numerical Algorithms, 59(1):95–118, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=1&
spage=95.

Orive:2022:GKQ

[OPSM22] Ramón Orive, Aleksandar V. Pejčev, Miodrag M. Spalević, and Ljubica Mihić. On the Gauss–Kronrod quadrature formula for a modified weight function of Chebyshev type. Numerical Algorithms, 91(4):1855–1877, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01325-8.

ORiordan:2012:PUN

[OQ12] Eugene O’Riordan and Jason Quinn. Parameter-uniform numerical methods for some singularly perturbed nonlinear initial value problems. Numerical Algorithms, 61(4): 579–611, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9552-3/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=61&issue=4&spage=579-611.

Onunwor:2017:CTS

[OR17] Enyinda Onunwor and Lothar Reichel. On the computation of a truncated SVD of a large linear discrete ill-posed problem. Numerical Algorithms, 75(2):359–380, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Oyewole:2024:TSE

[OR24] Olawale K. Oyewole and Simeon Reich. Two subgradient extragradient methods based on the golden ratio technique for solving variational inequality problems. Numerical Algorithms, 97(3):1215–1236, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01746-z.

Orban:2015:LML

[Orb15] Dominique Orban. Limited-memory LDL^{T} factorization of symmetric quasi-definite matrices with application to constrained optimization. Numerical Algorithms, 70(1):9–41, September 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9933-x.

Ozyapici:2014:MVM

[ÖRBB14] Ali Özyapici, Mustafa Riza, Bülent Bilgehan, and Agamirza E. Bashirov. On multiplicative and Volterra minimization methods. Numerical Algorithms, 67(3):623–636, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9813-9.

Ouahidi:2025:TI

[OS25] Salma Ouahidi and Rachid Sadaka. Tensor interpolation. Numerical Algorithms, 100(4):1599–1615, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02070-4.

Osada:1992:ELT

[Osa92a] Naoki Osada. Extensions of Levin’s transformations to vector sequences. Numerical Algorithms, 2(2):121–131, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Osada:1992:EMS

[Osa92b] Naoki Osada. Extrapolation methods for some singular fixed point sequences. Numerical Algorithms, 3(1–4):335–344, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Osada:2012:EHC

[Osa12] Naoki Osada. The early history of convergence acceleration methods. Numerical Algorithms, 60(2):205–221, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=2&
spage=205.

Ostrovska:2007:ALF

[Ost07] Sofiya Ostrovska. The approximation of logarithmic function by q-Bernstein polynomials in the case q > 1. Numerical Algorithms, 44(1):69–82, January 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=1&spage=69.

Oswald:2001:RMB

[Osw01] P. Oswald. Remarks on multilevel bases for divergencefree finite elements. Numerical Algorithms, 27(2):131–152, June 2001. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/35/2/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/35/
2/fulltext.pdf.

Ouazzi:1999:MFS

[Oua99] Abderrahim Ouazzi. A mixed formulation of the Stokes equation in terms of (w, p, u). Numerical Algorithms, 21 (1–4):343–352, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
19/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/19/fulltext.pdf.

Oviedo:2022:GCR

[Ovi22] Harry Oviedo. Global convergence of Riemannian line search methods with a Zhang–Hager-type condition. Numerical Algorithms, 91(3):1183–1203, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01298-8.

Pan:1996:MFF

[Pan96] K. Pan. On the method of finding frequencies with large amplitudes. Numerical Algorithms, 12(1–2):89–96, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pany:2018:FDS

[Pan18] Ambit Kumar Pany. Fully discrete second-order backward difference method for Kelvin–Voigt fluid flow model. Numerical Algorithms, 78(4):1061–1086, August 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pandur:2020:DHQ

[Pan20] Marija Miloloza Pandur. Detecting a hyperbolic quadratic eigenvalue problem by using a subspace algorithm. Numerical Algorithms, 83(2):767–787, February 2020. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Panic:2026:LPB

[Pan26] Stefan Panic. Legendre polynomial based approximation of the Hessian for quasi-Newton optimization. Numerical Algorithms, 102(2):1433–1463, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02182-x.

Paricio:2016:BNR

[Par16] Luis Javier Hernández Paricio. Bivariate Newton–Raphson method and toroidal attraction basins. Numerical Algorithms, 71(2):349–381, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-9996-3.

Paszkowski:1992:CAC

[Pas92] Stefan Paszkowski. Convergence acceleration of continued fractions of Poincaré’s type 1. Numerical Algorithms, 2(2):155–170, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Paszkowski:1995:QHS

[Pas95] S. Paszkowski. Quasipower and hypergeometric series— construction and evaluation. Numerical Algorithms, 10(3–4): 337–361, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Paszkowski:1999:BGI

[Pas99] Stefan Paszkowski. The Bloch–Grüneisen integrals. some analytic expressions. Numerical Algorithms, 20(4):369–378, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/19/8/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/19/8/fulltext.pdf.

Paszkowski:2003:CAS

[Pas03] Stefan Paszkowski. Convergence acceleration of some continued fractions. Numerical Algorithms, 32(2–4):193–247, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
45/5/fulltext.pdf.

Paszkowski:2006:EHI

[Pas06] Stefan Paszkowski. Evaluation of a Harrison integral. Numerical Algorithms, 41(4):353–358, April 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=4&spage=353.

Paszkowski:2008:CAO

[Pas08] Stefan Paszkowski. Convergence acceleration of orthogonal series. Numerical Algorithms, 47(1):35–62, January 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=1&
spage=35.

Paszkowski:2011:UMC

[Pas11] Stefan Paszkowski. Untypical methods of convergence acceleration. Numerical Algorithms, 56(2):185–209, February 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=2&
spage=185.

Patra:2025:NAI

[Pat25] Suchismita Patra. A numerical approach for investigating multiple eigenpairs of a quasilinear elliptic system. Numerical Algorithms, 98(2):987–1012, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01822-y.

Pavaloiu:2013:ANT

[PC13] Ion Pavaloiu and Emil Catinas. On an Aitken–Newton type method. Numerical Algorithms, 62(2):253–260, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9577-7/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=2&spage=253-260.

Pinto:2020:PAS

[PCDH20] M. Campos Pinto, F. Charles, B. Després, and M. Herda. A projection algorithm on the set of polynomials with two bounds. Numerical Algorithms, 85(4):1475–1498, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00872-x.

Piri:2019:AFP

[PDRG19] H. Piri, B. Daraby, S. Rahrovi, and M. Ghasemi. Approximating fixed points of generalized α-nonexpansive mappings in Banach spaces by new faster iteration process. Numerical Algorithms, 81(3):1129–1148, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pham:2023:PSA

[PDS^{+}23] Tan Nhat Pham, Minh N. Dao, Rakibuzzaman Shah, Nargiz Sultanova, Guoyin Li, and Syed Islam. A proximal subgradient algorithm with extrapolation for structured nonconvex nonsmooth problems. Numerical Algorithms, 94(4):1763–1795, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01554-5.

Pearson:2013:RBF

[Pea13] John W. Pearson. A radial basis function method for solving PDE-constrained optimization problems. Numerical Algorithms, 64(3):481–506, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9675-6.

Parvizi:2015:NSF

[PED15] M. Parvizi, M. R. Eslahchi, and Mehdi Dehghan. Numerical solution of fractional advection–diffusion equation with a nonlinear source term. Numerical Algorithms, 68(3):601–629, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9863-7.

Pei:2024:SID

[Pei24] Wenlong Pei. The semi-implicit DLN algorithm for the Navier– Stokes equations. Numerical Algorithms, 97(4):1673–1713, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01766-3.

Pejcev:2014:EBG

[Pej14] Aleksandar Pejcev. Error bounds for Gauss-type quadratures with Bernstein–Szegő weights. Numerical Algorithms, 66(3): 569–590, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9749-0.

Penski:1998:NSH

[Pen98] Christian Penski. A numerical scheme for highly oscillatory DAEs and its application to circuit simulation. Numerical Algorithms, 19(1–4):173–181, September 1998. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/16/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/3/fulltext.pdf. Differen-

tial algebraic equations (Grenoble, 1997).

Peng:2013:RLS

[Pen13] Zhuohua Peng. The reflexive least squares solutions of the matrix equation A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{l}X_{l}B_{l} = C with a submatrix constraint. Numerical Algorithms, 64(3):455– 480, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9674-7.

Pepino:2023:AST

[Pep23] Rafael Tristão Pepino. Acceleration of sequences with transformations involving hypergeometric functions. Numerical Algorithms, 92(1):893–915, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01334-7.

Petras:1995:GIC

[Pet95] Knut Petras. Gaussian integration of Chebyshev polynomials and analytic functions. Numerical Algorithms, 10(1–2):187– 202, July 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Petras:2001:FCC

[Pet01] Knut Petras. Fast calculation of coefficients in the Smolyak algorithm. Numerical Algorithms, 26(2):93–109, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/1/1/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/1/
1/fulltext.pdf.

Papageorgiou:1998:SSD

[PFT98] G. Papageorgiou, I. Th. Famelis, and Ch. Tsitouras. A P -stable singly diagonally implicit Runge–Kutta–Nyström method. Numerical Algorithms, 17(3–4):345–353, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/12/8/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
12/8/fulltext.pdf.

Proot:2005:ALS

[PG05] Michael M. J. Proot and Marc I. Gerritsma. Application of the least-squares spectral element method using Chebyshev polynomials to solve the incompressible Navier–Stokes equations. Numerical Algorithms, 38(1–3):155–172, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Parhi:2012:CTO

[PG12] S. K. Parhi and D. K. Gupta. Convergence of a third order method for fixed points in Banach spaces. Numerical Algorithms, 60(3):419–434, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=3&spage=419.

Pour:2015:NHS

[PG15] Hossein Noormohammadi Pour and Hossein Sadeghi Goughery. New Hermitian and skew-Hermitian splitting methods for non-Hermitian positive-definite linear systems. Numerical Algorithms, 69(1):207–225, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9890-4.

Priyanka:2024:FFA

[PGC24] T. M. C. Priyanka, A. Gowrisankar, and Jinde Cao. Fractal functions associated with Reich contractions: an approximation of chaotic attractors. Numerical Algorithms, 96(4): 1869–1886, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01687-7. See

comments [PK25a] and correction [PK25b].

Pestano-Gabino:1997:MPA

[PGGC97] Celina Pestano-Gabino and Concepción González-Concepción. Matrix Padé approximation of rational functions. Numerical Algorithms, 15(2):167–192, December 1997. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/7/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/7/3/fulltext.pdf.

Pestano-Gabino:2011:NII

[PGGCGF11] Celina Pestano-Gabino, Concepción González-Concepción, and María Candelaria Gil-Fariña. A note on the initial identification of scalar component models. Numerical Algorithms, 58(3):439–449, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=3&spage=439.

Priyanka:2025:CFD

[PGP^{+}25] T. M. C. Priyanka, A. Gowrisankar, M. Guru Prem Prasad, Yongshun Liang, and Jinde Cao. Caputo fractional derivative of α-fractal spline. Numerical Algorithms, 99(1):207– 228, May 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01875-z.

Pasha:2025:TPT

[PGSU25] Mirjeta Pasha, Silvia Gazzola, Connor Sanderford, and Ugochukwu O. Ugwu. TRIPs-Py: Techniques for regularization of inverse problems in Python. Numerical Algorithms, 99(1):285–322, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01878-w.

Peyghami:2014:CAI

[PH14] M. Reza Peyghami and S. Fathi Hafshejani. Complexity analysis of an interior-point algorithm for linear optimization based on a new proximity function. Numerical Algorithms, 67(1): 33–48, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9772-1.

Petkovic:2020:CMR

[PH20] Ivan Petković and – Dorde Herceg. Computers in mathematical research: the study of three-point root-finding methods. Numerical Algorithms, 84(3):1179–1198, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00796-6.

Petkovic:1998:SCS

[PHI98] Miodrag S. Petković, – Dor– de Herceg, and Snežana Ilić. Safe convergence of simultaneous methods for polynomial zeros. Numerical Algorithms, 17(3–4):313–331, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/12/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
12/5/fulltext.pdf.

Pietrus:1996:GCM

[Pie96] Alain Pietrus. A globally convergent method for solving nonlinear equations without the differentiability condition. Numerical Algorithms, 13(1–2):61–76, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pietrus:1999:STA

[Pié99] Alain Piétrus. Solving triangular algebraic systems by means of simultaneous iterations. Numerical Algorithms, 20 (4):353–368, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/

19/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/19/6/fulltext.pdf.

Pishbin:2016:NSS

[Pis16] S. Pishbin. Numerical solution and structural analysis of two-dimensional integral-algebraic equations. Numerical Algorithms, 73(2):305–322, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0096-9.

Pang:2022:FSA

[PJ22] Gang Pang and Songsong Ji. A fast stable accurate artificial boundary condition for the linearized Green–Naghdi system. Numerical Algorithms, 90(4):1437–1463, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01236-0.

Palitta:2021:CKM

[PK21] Davide Palitta and Patrick Kürschner. On the convergence of Krylov methods with low-rank truncations. Numerical Algorithms, 88(3):1383–1417, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01080-
2;   http://link.springer.com/content/pdf/10.1007/
s11075-021-01080-2.pdf.

Pischke:2022:CQA

[PK22a] Nicholas Pischke and Ulrich Kohlenbach. Correction to: Quantitative analysis of a subgradient-type method for equilibrium problems. Numerical Algorithms, 90(1):221, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01202-w. See [PK22b].

Pischke:2022:QAS

[PK22b] Nicholas Pischke and Ulrich Kohlenbach. Quantitative analysis of a subgradient-type method for equilibrium problems. Numerical Algorithms, 90(1):197–219, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01184-9. See correction [PK22a].

Prithvi:2025:CFF

[PK25a] B. V. Prithvi and S. K. Katiyar. Comments on “Fractal functions associated with Reich contractions: an approximation of chaotic attractors”. Numerical Algorithms, 98(1):35–43, January 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01780-5. See [PGC24] and

correction [PK25b].

Prithvi:2025:CCF

[PK25b] B. V. Prithvi and S. K. Katiyar. Correction to: Comments on “Fractal functions associated with Reich contractions: an approximation of chaotic attractors”. Numerical Algorithms, 98(1):45, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01798-9. See

[PK25a, PGC24].

Pakkaranang:2018:PPA

[PKC18] Nuttapol Pakkaranang, Poom Kumam, and Yeol Je Cho. Proximal point algorithms for solving convex minimization problem and common fixed points problem of asymptotically quasinonexpansive mappings in CAT(0) spaces with convergence analysis. Numerical Algorithms, 78(3):827–845, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Padcharoen:2019:SCF

[PKC19] Anantachai Padcharoen, Poom Kumam, and Yeol Je Cho. Split common fixed point problems for demicontractive operators. Numerical Algorithms, 82(1):297–320, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pak:2020:ETO

[PKR20] Chol-Kyu Pak, Mun-Chol Kim, and Chang-Ho Rim. An efficient third-order scheme for BSDEs based on nonequidistant difference scheme. Numerical Algorithms, 85(2):467–483, October 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00822-7.

Pan:1999:BAS

[PL99] Victor Y. Pan and Elliot Linzer. Bisection acceleration for the symmetric tridiagonal eigenvalue problem. Numerical Algorithms, 22(1):13–39, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
21/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/21/4/fulltext.pdf.

Perez:2004:RAN

[PL04] Rosana Pérez and Véra Lucia Rocha Lopes. Recent applications and numerical implementation of quasi-Newton methods for solving nonlinear systems of equations. Numerical Algorithms, 35(2–4):261–285, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/50/A/5/abstract.htm.

Pang:2025:DLB

[PL25] Liping Pang and Shuangyu Li. Discrete level-bundle method for mean-CVaR portfolio optimization with cardinality constraint. Numerical Algorithms, 99(4):2097–2135, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01947-0.

Plato:1999:CGM

[Pla99] Robert Plato. The conjugate gradient method for linear ill-posed problems with operator perturbations. Numerical Algorithms, 20(1):1–22, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
17/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/17/4/fulltext.pdf.

Plesniak:2003:VPL

[Ple03] W. Pleśniak. Volume of polynomial lemniscates in C^{n}. Numerical Algorithms, 33(1–4):415–420, August 2003. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/46/33/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/46/33/fulltext.pdf.

Plesniak:2012:NOM

[Ple12] Wiesław Pleśniak. Nearly optimal meshes in subanalytic sets. Numerical Algorithms, 60(4):545–553, August 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=545.

Pham:2020:CPA

[PLH20] Quy Muoi Pham, Delf Lachmund, and Dinh Nho Hào. Convergence of proximal algorithms with stepsize controls for nonlinear inverse problems and application to sparse non-negative matrix factorization. Numerical Algorithms, 85(4):1255–1279, December 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00864-x.

Plonka:1993:EAP

[Plo93] Gerlind Plonka. An efficient algorithm for periodic Hermite spline interpolation with shifted nodes. Numerical Algorithms, 5(1–4):51–62, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Plonka:1994:OSP

[Plo94] Gerlind Plonka. Optimal shift parameters for periodic spline interpolation. Numerical Algorithms, 6(3–4):297–316, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Potschka:2011:TPF

[PLVB11] Andreas Potschka, Filip Logist, Jan F. Van Impe, and Hans Georg Bock. Tracing the Pareto frontier in biobjective optimization problems by ODE techniques. Numerical Algorithms, 57(2):217–233, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=2&spage=217.

Peng:2024:UCS

[PLZ^{+}24] Shanshan Peng, Meng Li, Yanmin Zhao, Fawang Liu, and Fangfang Cao. Unconditionally convergent and superconvergent finite element method for nonlinear time-fractional parabolic equations with distributed delay. Numerical Algorithms, 95(4):1643–1714, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01624-
8.

Petkovic:2005:HOF

[PM05] Miodrag S. Petković and Dušan M. Milošević. A higher order family for the simultaneous inclusion of multiple zeros of polynomials. Numerical Algorithms, 39(4):415–435, August 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Padway:2022:ALF

[PM22] Emmett Padway and Dimitri Mavriplis. Approximate linearization of fixed-point iterations. Numerical Algorithms, 91 (2):583–614, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01274-2.

Priyadarshana:2024:EFS

[PM24] S. Priyadarshana and J. Mohapatra. An efficient fractional step numerical algorithm for time-delayed singularly perturbed 2D convection-diffusion-reaction problem with two small parameters. Numerical Algorithms, 97(2):687–726, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01720-9.

Perczynski:2025:NIH

[PM25] Rafał Perczyński and Grzegorz Madejski. Numerical integrator for highly oscillatory differential equations based on the Neumann series. Numerical Algorithms, 98(3):1481–1505, March 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01841-9.

Petrovic:2026:DEM

[PM26] Aleksandra Petrović and Marija Milosević. Divergence of the Euler–Maruyama method for neutral stochastic differential equations with unbounded delay and Markovian switching. Numerical Algorithms, 101(4):2767–2790, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02110-z.

Petkovic:2011:NHO

[PMM11] Miodrag S. Petković, Mimica R. Milosević, and Dusan M. Milosević. New higher-order methods for the simultaneous inclusion of polynomial zeros. Numerical Algorithms, 58(2):179–201, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=58&
issue=2&spage=179.

Paluszny:2005:LCP

[PMO05] Marco Paluszny, Guillermo Montilla, and José Rafael Ortega. Lemniscates 3D: a CAGD primitive? Numerical Algorithms, 39(1–3):317–327, July 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pfluger:1993:DSP

[PN93] P. R. Pfluger and M. Neamtu. On degenerate surface patches. Numerical Algorithms, 5(1–4):569–575, November 1993. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Prangprakhon:2021:ESC

[PN21] Mootta Prangprakhon and Nimit Nimana. Extrapolated sequential constraint method for variational inequality over the intersection of fixed-point sets. Numerical Algorithms, 88 (3):1051–1075, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01067-z.

Pan:2017:FPI

[PNW17] Jianyu Pan, Michael Ng, and Hong Wang. Fast preconditioned iterative methods for finite volume discretization of steady-state space-fractional diffusion equations. Numerical Algorithms, 74(1):153–173, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0143-6.

Pockaj:2014:HRS

[Poc14] Karla Pockaj. Hermite G^{1} rational spline motion of degree six. Numerical Algorithms, 66(4):721–739, August 2014.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9756-1.

Pogorelov:1998:DAE

[Pog98] D. Pogorelov. Differential-algebraic equations in multibody system modeling. Numerical Algorithms, 19(1–4): 183–194, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
21/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/21/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Poloni:2010:NSI

[Pol10] Federico Poloni. A note on the O(n)-storage implementation of the GKO algorithm and its adaptation to Trummer-like matrices. Numerical Algorithms, 55(1):115–139, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=1&
spage=115.

Popova:2004:PIL

[Pop04] Evgenija D. Popova. Parametric interval linear solver. Numerical Algorithms, 37(1–4):345–356, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/28/abstract.htm.

Popova:2015:UPT

[Pop15] Evgenija D. Popova. On the unbounded parametric tolerable solution set. Numerical Algorithms, 69(1):169–182, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9888-y.

Pearson:2017:NMC

[POP17] John W. Pearson, Sheehan Olver, and Mason A. Porter. Numerical methods for the computation of the confluent and Gauss hypergeometric functions. Numerical Algorithms, 74 (3):821–866, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0173-0.

Popa:2018:CRK

[Pop18a] Constantin Popa. Convergence rates for Kaczmarz-type algorithms. Numerical Algorithms, 79(1):1–17, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). See correction [Pop19].

Popova:2018:ESS

[Pop18b] Evgenija D. Popova. Enclosing the solution set of parametric interval matrix equation A(p)X = B(p). Numerical Algorithms, 78(2):423–447, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Popa:2019:CCR

[Pop19] Constantin Popa. Correction to: Convergence rates for Kaczmarz-type algorithms. Numerical Algorithms, 82(3):1117– 1120, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-019-00747-1.pdf. See

[Pop18a].

Popova:2021:PED

[Pop21] Evgenija D. Popova. Proving endpoint dependence in solving interval parametric linear systems. Numerical Algorithms, 86(3):1339–1358, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00936-3.

Potschka:2019:BSC

[Pot19] Andreas Potschka. Backward step control for Hilbert space problems. Numerical Algorithms, 81(1):151–180, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Powell:1993:TLE

[Pow93] M. J. D. Powell. Truncated Laurent expansions for the fast evaluation of thin plate splines. Numerical Algorithms, 5(1–4): 99–120, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Pinar:1992:HOP

[PP92] Miguel A. Piñar and Teresa E. Pérez. On higher order Padétype approximants with some prescribed coefficients in the numerator. Numerical Algorithms, 3(1–4):345–352, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Palmer:2005:NAS

[PP05] A. S. Palmer and T. N. Phillips. Numerical approximation of the spectra of Phan-Thien Tanner liquids. Numerical Algorithms, 38(1–3):133–153, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Petkovic:2006:CSG

[PP06] Miodrag S. Petković and Ljiljana D. Petković. On the convergence of the sequences of Gerschgorin-like disks. Numerical Algorithms, 42(3–4):363–377, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=3&spage=363.

Petra:2016:SPK

[PP16] Stefania Petra and Constantin Popa. Single projection Kaczmarz extended algorithms. Numerical Algorithms, 73(3): 791–806, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0118-7. See

erratum [PP17].

Petra:2017:ESP

[PP17] Stefania Petra and Constantin Popa. Erratum to: Single projection Kaczmarz extended algorithms. Numerical Algorithms, 76(2):599–604, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-017-

0390-1.pdf. See [PP16].

Petkovic:2018:DSS

[PP18] Miodrag S. Petković and Ljiljana D. Petković. Dynamic study of Schröder’s families of first and second kind. Numerical Algorithms, 78(3):847–865, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pozza:2021:GQG

[PP21] Stefano Pozza and Miroslav Pranić. The Gauss quadrature for general linear functionals, Lanczos algorithm, and minimal partial realization. Numerical Algorithms, 88(2):647–678, October 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01052-y.

Patel:2024:JSP

[PP24] Subhashree Patel and Bijaya Laxmi Panigrahi. Jacobi spectral projection methods for Fredholm integral equations of the first kind. Numerical Algorithms, 96(1):33–57, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01638-2.

Petrot:2023:DDC

[PPPN23] Narin Petrot, Mootta Prangprakhon, Porntip Promsinchai, and Nimit Nimana. A dynamic distributed conjugate gradient method for variational inequality problem over the common fixed-point constraints. Numerical Algorithms, 93(2):639– 668, June 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01430-8.

Popa:2015:RSL

[PPR15] Constantin Popa, Tobias Preclik, and Ulrich Rüde. Regularized solution of LCP problems with application to rigid body dynamics. Numerical Algorithms, 69(1):145–156, May 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9886-0.

Peng:2025:NAS

[PPS25] Xin Peng, Paolo Piersanti, and Xiaoqin Shen. Numerical approximation of the solution of Koiter’s model for an elliptic membrane shell in absence of friction subjected to an obstacle via the penalty method. Numerical Algorithms, 100(1): 109–132, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01957-y.

Potts:2009:FAN

[PPV09] Daniel Potts, Jürgen Prestin, and Antje Vollrath. A fast algorithm for nonequispaced Fourier transforms on the rotation group. Numerical Algorithms, 52(3):355–384, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=3&
spage=355.

Prestin:1995:TIW

[PQ95] J. Prestin and E. Quak. Trigonometric interpolation and wavelet decompositions. Numerical Algorithms, 9(3–4):293– 317, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pang:2022:AOM

[PQS22] Hong-Kui Pang, Hai-Hua Qin, and Hai-Wei Sun. All-at-once method for variable-order time fractional diffusion equations. Numerical Algorithms, 90(1):31–57, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01178-7.

Potchinkov:1993:GMV

[PR93] A. Potchinkov and R. Reemtsen. A globally most violated cutting plane method for complex minimax problems with application to digital filter design. Numerical Algorithms, 5(1–4): 611–620, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Palomares:2003:TDM

[PR03] A. Palomares and V. Ramírez. Thresholds of the divisor methods. Numerical Algorithms, 34(2–4):405–415, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/48/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
48/7/fulltext.pdf.

Pasquetti:2010:SEM

[PR10] Richard Pasquetti and Francesca Rapetti. Spectral element methods on unstructured meshes: which interpolation points? Numerical Algorithms, 55(2–3):349–366, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=2&
spage=349.

Petkovic:2014:GCN

[PR14] Miodrag S. Petković and Lidija Z. Rancić. On the guaranteed convergence of new two-point root-finding methods for polynomial zeros. Numerical Algorithms, 67(1):187–222, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9782-z.

Potzsche:2025:NDI

[PR25] Christian Pötzsche and David Rackl. Nyström discretization of integrodifference equations: numerical continuation of periodic solutions and Floquet multipliers. Numerical Algorithms, 98(3):1429–1465, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01839-3.

Prestin:1993:LIE

[Pre93] Jürgen Prestin. Lagrange interpolation on extended generalized Jacobi nodes. Numerical Algorithms, 5(1–4):179–190, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Petrovic:2018:HAG

[PRK^{+}18] Milena Petrović, Vladimir Rakocević, Natasa Kontrec, Stefan Panić, and Dejan Ilić. Hybridization of accelerated gradient descent method. Numerical Algorithms, 79(3):769–786, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Petrovic:2020:NHP

[PRVI20] Milena J. Petrović, Vladimir Rakocević, Dragana Valjarević, and Dejan Ilić. A note on hybridization process applied on transformed double step size model. Numerical Algorithms, 85(2):449–465, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00821-8.

Pryce:1998:SHI

[Pry98] J. D. Pryce. Solving high-index DAEs by Taylor series. Numerical Algorithms, 19(1–4):195–211, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/16/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

16/6/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Przybylowicz:2009:LIA

[Prz09] Paweł Przybyłowicz. Linear information for approximation of the Itô integrals. Numerical Algorithms, 52(4):677–699, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
52&issue=4&spage=663.

Przybylowicz:2016:OGA

[Prz16] Pawel Przybylowicz. Optimal global approximation of stochastic differential equations with additive Poisson noise. Numerical Algorithms, 73(2):323–348, October 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-016-0097-8.

Prokhorov:2000:MA

[PS00] V. A. Prokhorov and E. B. Saff. On meromorphic approximation. Numerical Algorithms, 25(1–4):305–321, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/14/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/14/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Pittaluga:2001:NAC

[PS01] Giovanna Pittaluga and Laura Sacripante. A numerical algorithm for cubature by bivariate splines on nonuniform partitions. Numerical Algorithms, 28(1–4):273–284, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/37/16/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/37/
16/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/37/16/abstract.htm;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/16/
fulltext.pdf.

Petkovic:2006:IAL

[PS06] Marko D. Petković and Predrag S. Stanimirović. Interpolation algorithm of Leverrier–Faddeev type for polynomial matrices. Numerical Algorithms, 42(3–4):345–361, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=42&issue=3&spage=345.

Pittaluga:2009:ASF

[PS09] Giovanna Pittaluga and Laura Sacripante. An algorithm for solving Fredholm integro-differential equations. Numerical Algorithms, 50(2):115–126, February 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=2&spage=115.

Petkovic:2016:SED

[PS16] Miodrag S. Petković and Janak Raj Sharma. On some efficient derivative-free iterative methods with memory for solving systems of nonlinear equations. Numerical Algorithms, 71(2): 457–474, February 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-0003-9.

Plociniczak:2017:NSI

[PS17] Lukasz Plociniczak and Szymon Sobieszek. Numerical schemes for integro-differential equations with Erdélyi–Kober fractional operator. Numerical Algorithms, 76(1):125–150, September 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-016-0247-z.pdf.

Piersanti:2020:NMS

[PS20] Paolo Piersanti and Xiaoqin Shen. Numerical methods for static shallow shells lying over an obstacle. Numerical Algorithms, 85(2):623–652, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00830-
7;   http://link.springer.com/content/pdf/10.1007/
s11075-019-00830-7.pdf.

Plaskota:2022:AAU

[PS22] Leszek Plaskota and Paweł Samoraj. Automatic approximation using asymptotically optimal adaptive interpolation. Numerical Algorithms, 89(1):277–302, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01114-9;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01114-9.pdf.

Piccinini:2025:TLT

[PS25] Lorenzo Piccinini and Valeria Simoncini. TT-LSQR for tensor least squares problems and application to data mining. Numerical Algorithms, 100(4):2069–2093, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02204-8.

Patel:2026:EHN

[PSB26] Subham Patel, Sudarshan Santra, and Ratikanta Behera. An efficient hybrid numerical approach for time-fractional subdiffusion equations with multi-singularities. Numerical Algorithms, 102(1):85–122, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02122-
9.

Pandey:2010:SAN

[PSS10] Rajesh K. Pandey, Om Prakash Singh, and Vineet K. Singh. A stable algorithm for numerical evaluation of Hankel transforms using Haar wavelets. Numerical Algorithms, 53(4):451–466, April 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=53&
issue=4&spage=451.

Prado:2022:CAP

[PSS22] Renan W. Prado, Sandra A. Santos, and Lucas E. A. Simões. On the convergence analysis of a penalty algorithm for nonsmooth optimization and its performance for solving hardsphere problems. Numerical Algorithms, 91(2):933–957, October 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01287-x.

Pietrus:2025:IQN

[PSS25] A. Piétrus, P. S. M. Santos, and G. N. Silva. Inexact quasi-Newton methods under a relaxed metric regularity assumption. Numerical Algorithms, 100(3):1045–1067, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01979-6.

Pulverer:2011:AGC

[PSW11] Gernot Pulverer, Gustaf Söderlind, and Ewa Weinmüller. Automatic grid control in adaptive BVP solvers. Numerical Algorithms, 56(1):61–92, January 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=61.

Pandiya:2023:FGM

[PSWE23] Ridwan Pandiya, Salmah Salmah, Widodo Widodo, and Irwan Endrayanto. Finding global minima with an inflection pointbased filled function algorithm. Numerical Algorithms, 92(2): 1403–1424, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01346-3.

Pei:2023:FSA

[PSZ23] Yonggang Pei, Shaofang Song, and Detong Zhu. A filter sequential adaptive cubic regularization algorithm for nonlinear constrained optimization. Numerical Algorithms, 93(4): 1481–1507, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01475-9.

Pham:2017:SEL

[PT17] Ky Anh Pham and Ngoc Hai Trinh. Splitting extragradientlike algorithms for strongly pseudomonotone equilibrium problems. Numerical Algorithms, 76(1):67–91, September 2017.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Petrov:2018:SSP

[PT18] Miroslav S. Petrov and Todor D. Todorov. Stable subdivision of 4D polytopes. Numerical Algorithms, 79(2):633–656, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pade:2019:WRC

[PT19] Jonas Pade and Caren Tischendorf. Waveform relaxation: a convergence criterion for differential–algebraic equations. Numerical Algorithms, 81(4):1327–1342, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Papez:2024:EEN

[PT24] Jan Papez and Petr Tichý. Estimating error norms in CGlike algorithms for least-squares and least-norm problems. Numerical Algorithms, 97(1):1–28, September 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01691-x.

Pham-Trong:2001:PGT

[PTSB01] Valérie Pham-Trong, Nicolas Szafran, and Luc Biard. Pseudo-geodesics on three-dimensional surfaces and pseudogeodesic meshes. Numerical Algorithms, 26(4):305–315, April 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/32/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
32/1/fulltext.pdf;   http://ipsapp007.lwwonline.com/
content/getfile/5058/32/1/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/32/1/fulltext.pdf.

Petrov:2022:EFD

[PTW22] Miroslav S. Petrov, Todor D. Todorov, and Freddie D. Witherden. Enabling four-dimensional conformal hybrid meshing with cubic pyramids. Numerical Algorithms, 91(2):671–709, October 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01278-y.

Piquet:1998:MPK

[PV98] Jean Piquet and Xavier Vasseur. Multigrid preconditioned Krylov subspace methods for three-dimensional numerical solutions of the incompressible Navier–Stokes equations. Numerical Algorithms, 17(1–2):1–32, March 1998. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL

http://ipsapp007.kluweronline.com/content/getfile/
5058/31/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/31/6/fulltext.pdf.

Prevost:1999:PPP

[PV99] M. Prévost and D. Vekemans. Partial Padé prediction. Numerical Algorithms, 20(1):23–50, March 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
17/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/17/3/fulltext.pdf.

Piquet:2000:NMM

[PV00] Jean Piquet and Xavier Vasseur. A nonstandard multigrid method with flexible multiple semicoarsening for the numerical solution of the pressure equation in a Navier– Stokes solver. Numerical Algorithms, 24(4):333–355, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/29/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

29/3/fulltext.pdf. Sparse matrices and differential equations (Rennes, 1999).

Prevost:2003:APN

[PV03] Marc Prévost and Denis Vekemans. Approximation and prediction of the numerical solution of some Burgers problems. Numerical Algorithms, 33(1–4):421–431, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/34/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/46/
34/fulltext.pdf.

Pandey:2022:RSR

[PV22a] K. K. Pandey and P. Viswanathan. In reference to a selfreferential approach towards smooth multivariate approximation. Numerical Algorithms, 91(1):251–281, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01261-7.

Papez:2022:IGE

[PV22b] Jan Papez and Martin Vohralík. Inexpensive guaranteed and efficient upper bounds on the algebraic error in finite element discretizations. Numerical Algorithms, 89(1):371–407, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01118-5.

Papadimitriou:2023:PNA

[PV23] Dimitri Papadimitriou and B´ ăng Công Vũ. A penalized nonlinear ADMM algorithm applied to the multi-constrained traffic assignment problem. Numerical Algorithms, 92(4):2219– 2242, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01384-x.

Plaskota:2004:SAI

[PW04] Leszek Plaskota and Grzegorz W. Wasilkowski. Smolyak’s algorithm for integration and L_{1}-approximation of multivariate functions with bounded mixed derivatives of second order. Numerical Algorithms, 36(3):229–246, July 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/57/A/3/abstract.htm.

Plaskota:2014:EAM

[PW14] L. Plaskota and G. W. Wasilkowski. Efficient algorithms for multivariate and ∞-variate integration with exponential weight. Numerical Algorithms, 67(2):385–403, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-013-9798-4.pdf.

Plonka:2016:DSF

[PW16] Gerlind Plonka and Katrin Wannenwetsch. A deterministic sparse FFT algorithm for vectors with small support. Numerical Algorithms, 71(4):889–905, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0028-0.

Pang:2022:FPB

[PW22] Li-Ping Pang and Qi Wu. A feasible proximal bundle algorithm with convexification for nonsmooth, nonconvex semiinfinite programming. Numerical Algorithms, 90(1):387–422, May 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01192-9.

Plonka:2018:DSF

[PWCsL18] Gerlind Plonka, Katrin Wannenwetsch, Annie Cuyt, and Wen shin Lee. Deterministic sparse FFT for M-sparse vectors. Numerical Algorithms, 78(1):133–159, May 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Pan:2025:EAT

[PXH25] Kejia Pan, Jiajia Xie, and Hongling Hu. An efficient acceleration technique of implicit schemes for quasi-linear parabolic problems. Numerical Algorithms, 100(3):921–935, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01973-y.

Peiping:2023:SBB

[PYD23] Shen Peiping, Wang Yafei, and Wu Dianxiao. A spatial branch and bound algorithm for solving the sum of linear ratios optimization problem. Numerical Algorithms, 93(3):1373–1400, July 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01471-z.

Pollock:2020:MAA

[PZ20] Sara Pollock and Yunrong Zhu. A matrix analysis approach to discrete comparison principles for nonmonotone PDE. Numerical Algorithms, 83(3):1007–1027, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Percan:2015:AEC

[PZL15] Yvonne Percan, Stefan Zellmann, and Ulrich Lang. Asymptotic error of cubic B-spline interpolation using prefiltering. Numerical Algorithms, 70(1):191–203, September 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9941-x;   http://link.springer.com/
content/pdf/10.1007/s11075-014-9941-x.pdf.

Qureshi:2024:NOR

[QAS^{+}24] Sania Qureshi, Ioannis K. Argyros, Amanullah Soomro, Krzysztof Gdawiec, Asif Ali Shaikh, and Evren Hincal. A new optimal root-finding iterative algorithm: local and semilocal analysis with polynomiography. Numerical Algorithms, 95 (4):1715–1745, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01625-7.

Qiang:2012:MMM

[QL12] Hua Qiang and Kewei Liang. Moving mesh method for problems with blow-up on unbounded domains. Numerical Algorithms, 59(1):63–77, January 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=1&spage=63.

Qian:2011:SAA

[QLZX11] Y. H. Qian, S. K. Lai, W. Zhang, and Y. Xiang. Study on asymptotic analytical solutions using HAM for strongly nonlinear vibrations of a restrained cantilever beam with an intermediate lumped mass. Numerical Algorithms, 58(3):293–314, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
58&issue=3&spage=293.

Qiao:2023:CNA

[QQX23] Leijie Qiao, Wenlin Qiu, and Da Xu. Crank–Nicolson ADI finite difference/compact difference schemes for the 3D tempered integrodifferential equation associated with Brownian motion. Numerical Algorithms, 93(3):1083–1104, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01454-0.

Qiu:2008:GAE

[QW08] W.-Y. Qiu and R. Wong. Global asymptotic expansions of the Laguerre polynomials — a Riemann–Hilbert approach. Numerical Algorithms, 49(1–4):331–372, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=49&issue=1&
spage=331.

Qiu:2020:TTG

[QXGZ20] Wenlin Qiu, Da Xu, Jing Guo, and Jun Zhou. A time twogrid algorithm based on finite difference method for the twodimensional nonlinear time-fractional mobile/immobile transport model. Numerical Algorithms, 85(1):39–58, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00801-y.

Qiu:2019:NFS

[QZG^{+}19] Binbin Qiu, Yunong Zhang, Jinjin Guo, Zhi Yang, and Xiaodong Li. New five-step DTZD algorithm for future nonlinear minimization with quartic steady-state error pattern. Numerical Algorithms, 81(3):1043–1065, July 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ren:2012:ILA

[RA12] Hongmin Ren and Ioannis K. Argyros. Improved local analysis for a certain class of iterative methods with cubic convergence. Numerical Algorithms, 59(4):505–521, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=505.

Rabinowitz:1992:EMN

[Rab92a] Philip Rabinowitz. Extrapolation methods in numerical integration. Numerical Algorithms, 3(1–4):17–28, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Rabut:1992:EHC

[Rab92b] Christophe Rabut. Elementary m-harmonic cardinal Bsplines. Numerical Algorithms, 2(1):39–61 (or 39–62??), February 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rabut:1992:HLH

[Rab92c] Christophe Rabut. High level m-harmonic cardinal B-splines. Numerical Algorithms, 2(1):63–84, February 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rabut:2005:LTP

[Rab05] Christophe Rabut. Locally tensor product functions. Numerical Algorithms, 39(1–3):329–348, July 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rabago:2023:NSE

[Rab23] Julius Fergy T. Rabago. Numerical solution to the exterior Bernoulli problem using the Dirichlet–Robin energy gap cost functional approach in two and three dimensions. Numerical Algorithms, 94(1):175–227, September 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01497-x.

Radunovic:2008:MEB

[Rad08] Desanka Radunović. Multiresolution exponential B-splines and singularly perturbed boundary problem. Numerical Algorithms, 47(2):191–210, February 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=191.

Rahman:2011:SPC

[Rah11a] Talal Rahman. Schwarz preconditioned CG algorithm for the mortar finite element. Numerical Algorithms, 58(2):235–260, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
58&issue=2&spage=235.

Ren:2011:DFI

[RAH11b] Hongmin Ren, Ioannis K. Argyros, and Saïd Hilout. A derivative free iterative method for solving least squares problems. Numerical Algorithms, 58(4):555–571, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=58&issue=4&
spage=555.

Rui:2026:MRC

[RALJ26] Liu Rui, Yang Ai-Li, and Ma Jian. Maximal residual coordinate descent method with k-means clustering for solving large linear least-squares problems. Numerical Algorithms, 101(2): 1157–1173, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02036-6.

Roul:2017:NNA

[RB17] Pradip Roul and Dipak Biswal. A new numerical approach for solving a class of singular two-point boundary value problems. Numerical Algorithms, 75(3):531–552, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Raffo:2021:WQI

[RB21] Andrea Raffo and Silvia Biasotti. Weighted quasi-interpolant spline approximations: Properties and applications. Numerical Algorithms, 87(2):819–847, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00989-
4.

Rashidinia:2014:ASG

[RBN14] J. Rashidinia, A. Barati, and M. Nabati. Application of Sinc– Galerkin method to singularly perturbed parabolic convectiondiffusion problems. Numerical Algorithms, 66(3):643–662, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9752-5.

Rada:2014:VMN

[RC14] Lavdie Rada and Ke Chen. A variational model and its numerical solution for local, selective and automatic segmentation. Numerical Algorithms, 66(2):399–430, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9741-8.

Ren:2022:ISP

[RCW22] Bi-Cong Ren, Fang Chen, and Xiao-Liang Wang. Improved splitting preconditioner for double saddle point problems arising from liquid crystal director modeling. Numerical Algorithms, 91(3):1363–1379, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01305-y.

Reiser:2004:IGM

[RDdRC04] Renata Hax Sander Reiser, Graçaliz Pereira Dimuro, and Antônio Carlos da Rocha Costa. The interval geometric machine model. Numerical Algorithms, 37(1–4):357–366, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp009.kluweronline.
com/IPS/content/ext/x/J/5058/I/58/A/29/abstract.htm.

Rebillard:1997:RAC

[Reb97] Luc Rebillard. Rational approximation in the complex plane using a τ-method and computer algebra. Numerical Algorithms, 16(2):187–208, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/7/fulltext.pdf.

Recchioni:2001:QCM

[Rec01] Maria Cristina Recchioni. Quadratically convergent method for simultaneously approaching the roots of polynomial solutions of a class of differential equations: Application to orthogonal polynomials. Numerical Algorithms, 28(1– 4):285–308, December 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/37/
17/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/37/17/fulltext.pdf;   http://
ipsapp007.lwwonline.com/content/getfile/5058/37/17/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/37/17/fulltext.pdf.

Reemtsen:1992:CPM

[Ree92] Rembert Reemtsen. A cutting plane method for solving minimax problems in the complex plane. Numerical Algorithms, 2(3–4):409–436, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rhouni:2026:RNA

[REE26] Omar Rhouni, Mohammed Errachid, and Mustapha Esghir. RHBPIA: a new algorithm for computing Hermite–Birkhoff interpolation polynomials. Numerical Algorithms, 101(4):2791– 2818, April 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02114-9.

Reif:1997:UBS

[Rei97] Ulrich Reif. Uniform B-spline approximation in Sobolev spaces. Numerical Algorithms, 15(1):1–14, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/30/1/abstract.htm;   http://ipsapp007.
kluweronline.com/content/getfile/5058/30/1/fulltext.
pdf.

Reich:1998:MPE

[Rei98] Sebastian Reich. Modified potential energy functions for constrained molecular dynamics. Numerical Algorithms, 19 (1–4):213–221, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
16/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/16/5/fulltext.pdf. Differen-

tial algebraic equations (Grenoble, 1997).

Rashidinia:2021:NOV

[REM21] Jalil Rashidinia, Tahereh Eftekhari, and Khosrow Maleknejad. A novel operational vector for solving the general form of distributed order fractional differential equations in the time domain based on the second kind Chebyshev wavelets. Numerical Algorithms, 88(4):1617–1639, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01088-8.

Revol:2003:INI

[Rev03] N. Revol. Interval Newton iteration in multiple precision for the univariate case. Numerical Algorithms, 34(2– 4):417–426, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/8/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/8/fulltext.pdf.

Ran:2023:SSO

[RF23] Maohua Ran and Zhouping Feng. A stable second-order difference scheme for the generalized time-fractional non-Fickian delay reaction-diffusion equations. Numerical Algorithms, 93 (3):993–1012, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01450-4.

R:2023:NSA

[RFS23] Sahu D. R., Babu Feeroz, and Sharma Shikher. A new self-adaptive iterative method for variational inclusion problems on Hadamard manifolds with applications. Numerical Algorithms, 94(3):1435–1460, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01542-9.

Rump:2010:VEB

[RG10] Siegfried M. Rump and Stef Graillat. Verified error bounds for multiple roots of systems of nonlinear equations. Numerical Algorithms, 54(3):359–377, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=359.

Rashidinia:2010:HNS

[RGJ10] Jalil Rashidinia, M. Ghasemi, and R. Jalilian. An O(h^{6}) numerical solution of general nonlinear fifth-order two point boundary value problems. Numerical Algorithms, 55(4): 403–428, December 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=55&issue=4&spage=403.

Rhanizar:2022:MVH

[Rha22] Bouchta Rhanizar. On a modified version of the Henrici’s transformation into optimization. Numerical Algorithms, 90 (2):507–522, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01196-5.

Ren:2015:ESN

[RhG15] Jincheng Ren and Guang hua Gao. Efficient and stable numerical methods for the two-dimensional fractional Cattaneo equation. Numerical Algorithms, 69(4):795–818, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9926-9.

Riahi:2016:NAI

[Ria16] Mohamed Kamel Riahi. A new approach to improve illconditioned parabolic optimal control problem via time domain decomposition. Numerical Algorithms, 72(3):635–666, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0060-0.

Rahaman:2019:YAI

[RIAA19] Mijanur Rahaman, Mohd. Ishtyak, Rais Ahmad, and Imran Ali. The Yosida approximation iterative technique for split monotone Yosida variational inclusions. Numerical Algorithms, 82(1):349–369, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rim:2025:AAI

[Rim25] Kyung Soo Rim. An algorithm for approximating implicit functions by polynomials without using higher order differentiability information. Numerical Algorithms, 98(3):1261–1279, March 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01833-9.

Rippa:1993:SDI

[Rip93] Shmuel Rippa. Scattered data interpolation using minimum energy Powell–Sabin elements and data dependent triangulations. Numerical Algorithms, 5(1–4):577–587, ???? 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rizzardi:2018:DSC

[Riz18] Mariarosaria Rizzardi. Detection of the singularities of a complex function by numerical approximations of its Laurent coefficients. Numerical Algorithms, 77(4):955–982, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rostamy:2011:HMH

[RK11] Davood Rostamy and K. Karimi. Hypercomplex mathematics and HPM for the time-delayed Burgers equation with convergence analysis. Numerical Algorithms, 58(1):85–101, September 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=58&
issue=1&spage=85.

Raza:2025:AOE

[RKM25] Nusrat Raza, Manoj Kumar, and M. Mursaleen. Approximation operators explored through the lens of degenerate Hermite polynomials. Numerical Algorithms, 100(3):1179–1198, November 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01985-8.

Ramos:2016:OTS

[RKMS16] Higinio Ramos, Z. Kalogiratou, Th. Monovasilis, and T. E. Simos. An optimized two-step hybrid block method for solving general second order initial-value problems. Numerical Algorithms, 72(4):1089–1102, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0081-8.

Rashidinia:2011:TSS

[RM11] Jalil Rashidinia and Reza Mohammadi. Tension spline solution of nonlinear sine-Gordon equation. Numerical Algorithms, 56(1):129–142, January 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=1&spage=129.

Rodrigues:2013:TMN

[RM13] Maria João Rodrigues and José Matos. A tau method for nonlinear dynamical systems. Numerical Algorithms, 62(4): 583–600, April 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9663-x.

Ruby:2026:SJS

[RM26] Ruby and Moumita Mandal. Superconvergent Jacobi spectral Galerkin method for nonlinear fractional Volterra–Fredholm integro-differential equations. Numerical Algorithms, 102 (1):531–568, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02139-0.

Rashidinia:2013:SGM

[RMT13] Jalil Rashidinia, Khosro Maleknejad, and Narges Taheri. Sinc–Galerkin method for numerical solution of the Bratu’s problems. Numerical Algorithms, 62(1):1–11, January 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9560-3/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=1-11.

Rebrova:2021:BGS

[RN21] Elizaveta Rebrova and Deanna Needell. On block Gaussian sketching for the Kaczmarz method. Numerical Algorithms, 86(1):443–473, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00895-9.

Raina:2026:EOS

[RNT26] Aayushman Raina, Srinivasan Natesan, and Suayip Toprakseven. An efficient operator splitting weak Galerkin method for singularly perturbed 2D parabolic PDEs. Numerical Algorithms, 102(2):751–777, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02152-
3.

Roberts:1992:CAV

[Rob92] D. E. Roberts. Clifford algebras and vector-valued rational forms. II. Numerical Algorithms, 3(1–4):371–381, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Robinson:1997:CFI

[Rob97] James C. Robinson. Convergent families of inertial manifolds for convergent approximations. Numerical Algorithms, 14(1–3):179–188, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/11/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/11/fulltext.pdf. Dynami-

cal numerical analysis (Atlanta, GA, 1995).

Roberts:1998:VCA

[Rob98] D. E. Roberts. A vector Chebyshev algorithm. Numerical Algorithms, 17(1–2):33–50, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/31/
11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/31/11/fulltext.pdf.

Roberts:2002:VEA

[Rob02] D. E. Roberts. The vector epsilon algorithm — a residual approach. Numerical Algorithms, 29(1–3):209–227, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.lwwonline.com/
content/getfile/5058/38/12/abstract.htm;   http://
ipsapp008.lwwonline.com/content/getfile/5058/38/12/
fulltext.pdf.

Rahimkhani:2017:NOM

[ROB17] P. Rahimkhani, Y. Ordokhani, and E. Babolian. A new operational matrix based on Bernoulli wavelets for solving fractional delay differential equations. Numerical Algorithms, 74(1):223– 245, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0146-3.

Rahimkhani:2018:MLW

[ROB18] P. Rahimkhani, Y. Ordokhani, and E. Babolian. Müntz– Legendre wavelet operational matrix of fractional-order integration and its applications for solving the fractional pantograph differential equations. Numerical Algorithms, 77(4): 1283–1305, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rogers:1995:PTM

[Rog95] Jack W. Rogers, Jr. Potential transformation methods for large-scale constrained global optimization. Numerical Algorithms, 9(1–2):13–24, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Rohwer:2007:EMU

[Roh07] Carl Heinrich Rohwer. The estimation of moments of an unknown error distribution in the discrete pulse transform. Numerical Algorithms, 45(1–4):239–251, August 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-007-9116-0.

Ronning:1992:PFS

[Røn92a] Frode Rønning. PC-fractions and Szegő polynomials associated with starlike univalent functions. Numerical Algorithms, 3(1–4):383–391, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Ronveaux:1992:SIP

[Ron92b] A. Ronveaux. Sobolev inner products and orthogonal polynomials of Sobolev type. Numerical Algorithms, 3(1–4):393–399, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Ronveaux:2008:OPC

[Ron08] André Ronveaux. Orthogonal polynomials — centroid of their zeroes. Numerical Algorithms, 49(1–4):373–385, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=49&issue=1&
spage=373.

Rossini:1997:IDN

[Ros97] Milvia Rossini. Irregularity detection from noisy data in one and two dimensions. Numerical Algorithms, 16(3– 4):283–301, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
11/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/11/3/fulltext.pdf.

Roul:2025:HON

[Rou25] Pradip Roul. A high-order numerical scheme and its analysis for Caputo temporal-fractional Black–Scholes model: European double barrier knock-out option. Numerical Algorithms, 98(1):467–502, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01802-2.

Reddy:2026:CPC

[RP26] G. D. Reddy and D. Pradeep. A class of parameter choice strategies for the finite dimensional iterated weighted Tikhonov regularization scheme. Numerical Algorithms, 102(1):169–197, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02126-5.

Rey:1998:RRP

[RR98] Christian Rey and Franck Risler. A Rayleigh–Ritz preconditioner for the iterative solution to large scale nonlinear problems. Numerical Algorithms, 17(3–4):279–311, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/12/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
12/1/fulltext.pdf.

Risler:2000:IAA

[RR00] Franck Risler and Christian Rey. Iterative accelerating algorithms with Krylov subspaces for the solution to large-scale nonlinear problems. Numerical Algorithms, 23(1):1–30, March 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/24/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
24/1/fulltext.pdf.

Rabut:2008:PMA

[RR08] Christophe Rabut and Milvia Rossini. Polyharmonic multiresolution analysis: an overview and some new results. Numerical Algorithms, 48(1–3):135–160, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=135.

Reichel:2013:ONP

[RR13] Lothar Reichel and Giuseppe Rodriguez. Old and new parameter choice rules for discrete ill-posed problems. Numerical Algorithms, 63(1):65–87, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9612-8.

Ramos:2020:NSB

[RR20] Higinio Ramos and M. A. Rufai. Numerical solution of boundary value problems by using an optimized two-step block method. Numerical Algorithms, 84(1):229–251, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Roul:2022:NHO

[RR22] Pradip Roul and Vikas Rohil. A novel high-order numerical scheme and its analysis for the two-dimensional time-fractional reaction–subdiffusion equation. Numerical Algorithms, 90(4): 1357–1387, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01233-3.

Reinsch:2023:SVD

[RR23] Christian Reinsch and Mathias Richter. Singular value decomposition in extended double precision arithmetic. Numerical Algorithms, 93(3):1137–1155, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01459-
9.

Reichel:2009:EEL

[RRS09] L. Reichel, G. Rodriguez, and S. Seatzu. Error estimates for large-scale ill-posed problems. Numerical Algorithms, 51 (3):341–361, July 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=51&issue=3&spage=341. Tributes to Gene H.

Golub, Part III.

Rodriguez:2021:NDF

[RRZ21] Ana Alonso Rodríguez, Francesca Rapetti, and Elena Zappon. New degrees of freedom for high-order Whitney approximations of Darcy’s flows. Numerical Algorithms, 87(4): 1613–1634, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01022-4.

Rodriguez:1993:AMF

[RS93] Giuseppe Rodriguez and Sebastiano Seatzu. Approximation methods for the finite moment problem. Numerical Algorithms, 5(1–4):391–405, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Renaut:1997:ECP

[RS97] Rosemary Renaut and Yi Su. Evaluation of Chebyshev pseudospectral methods for third order differential equations. Numerical Algorithms, 16(3–4):255–281, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/11/2/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
11/2/fulltext.pdf.

Robbe:2002:CAG

[RS02] Mickaël Robbé and Miloud Sadkane. A convergence analysis of GMRES and FOM methods for Sylvester equations. Numerical Algorithms, 30(1):71–89, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
40/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/40/6/fulltext.pdf.

Rapun:2006:BIA

[RS06] María-Luisa Rapún and Francisco-Javier Sayas. Boundary integral approximation of a heat-diffusion problem in time-harmonic regime. Numerical Algorithms, 41(2):127–160, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
41&issue=2&spage=127.

Rai:2020:NAC

[RS20] Pratima Rai and Kapil K. Sharma. Numerical approximation for a class of singularly perturbed delay differential equations with boundary and interior layer(s). Numerical Algorithms, 85(1):305–328, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00815-6.

Rauhut:2021:TTN

[RS21] Holger Rauhut and Zeljka Stojanac. Tensor theta norms and low rank recovery. Numerical Algorithms, 88(1):25–66, September 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01029-x;   http://link.
springer.com/content/pdf/10.1007/s11075-020-01029-
x.pdf.

Roul:2025:NNM

[RS25] Pradip Roul and S. Sundar. Novel numerical methods based on graded, adaptive and uniform meshes for a time-fractional advection–diffusion equation subjected to weakly singular solution. Numerical Algorithms, 98(2):531–561, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01804-0.

Roul:2026:TEN

[RS26] Pradip Roul and S. Sundar. Two efficient non-uniform mesh numerical techniques for a multi-term time-fractional advection-diffusion equation: Moving mesh method. Numerical Algorithms, 101(2):1273–1302, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02042-8.

Rodriguez-Sanchez:2019:LAT

[RSCH^{+}19] Rafael Rodríguez-Sánchez, Sandra Catalán, José R. Herrero, Enrique S. Quintana-Ortí, and Andrés E. Tomás. Look-ahead in the two-sided reduction to compact band forms for symmetric eigenvalue problems and the SVD. Numerical Algorithms, 80(2):635–660, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ramos:2017:EPN

[RSKB17] Higinio Ramos, Gurjinder Singh, V. Kanwar, and Saurabh Bhatia. An embedded 3(2) pair of nonlinear methods for solving first order initial-value ordinary differential systems. Numerical Algorithms, 75(3):509–529, July 2017. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Rodriguez:2003:NTI

[RST03] G. Rodriguez, S. Seatzu, and D. Theis. A new technique for ill-conditioned linear systems. Numerical Algorithms, 33 (1–4):433–442, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
35/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/35/fulltext.pdf.

Reichel:2020:SSC

[RSZ20] Lothar Reichel, Hassane Sadok, and Wei-Hong Zhang. Simple stopping criteria for the LSQR method applied to discrete ill-posed problems. Numerical Algorithms, 84(4):1381–1395, August 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00852-1.

Rashidinia:2012:SNV

[RT12] Jalil Rashidinia and Ali Tahmasebi. Systems of nonlinear Volterra integro-differential equations. Numerical Algorithms, 59(2):197–212, February 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=2&spage=197.

Ravnik:2019:FBD

[RT19] Jure Ravnik and Jan Tibuat. Fast boundary-domain integral method for unsteady convection-diffusion equation with variable diffusivity using the modified Helmholtz fundamental solution. Numerical Algorithms, 82(4):1441–1466, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Reich:2020:NAS

[RT20] Simeon Reich and Truong Minh Tuyen. A new algorithm for solving the split common null point problem in Hilbert spaces. Numerical Algorithms, 83(2):789–805, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Reich:2022:GCI

[RT22] Simeon Reich and Truong Minh Tuyen. A generalized cyclic iterative method for solving variational inequalities over the solution set of a split common fixed point problem. Numerical Algorithms, 91(1):1–17, September 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01252-0.

Reich:2024:TNS

[RT24] Simeon Reich and Truong Minh Tuyen. Two new selfadaptive algorithms for solving the split feasibility problem in Hilbert space. Numerical Algorithms, 95(2):1011–1032, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01597-8.

Reich:2021:IPT

[RTCL21] Simeon Reich, Duong Viet Thong, Prasit Cholamjiak, and Luong Van Long. Inertial projection-type methods for solving pseudomonotone variational inequality problems in Hilbert space. Numerical Algorithms, 88(2):813–835, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01058-6.

Reich:2021:NAC

[RTD^{+}21] Simeon Reich, Duong Viet Thong, Qiao-Li Dong, Xiao-Huan Li, and Vu Tien Dung. New algorithms and convergence theorems for solving variational inequalities with non-Lipschitz mappings. Numerical Algorithms, 87(2):527–549, June 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00977-8.

Reich:2022:NSA

[RTTH22] Simeon Reich, Truong Minh Tuyen, Nguyen Thi Thu Thuy, and Mai Thi Ngoc Ha. A new self-adaptive algorithm for solving the split common fixed point problem with multiple output sets in Hilbert spaces. Numerical Algorithms, 89(3): 1031–1047, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01144-3.

Rump:2012:FIM

[Rum12] Siegfried M. Rump. Fast interval matrix multiplication. Numerical Algorithms, 61(1):1–34, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=1&spage=1.

Rump:2014:ICV

[Rum14] Siegfried M. Rump. Improved componentwise verified error bounds for least squares problems and underdetermined linear systems. Numerical Algorithms, 66(2):309–322, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9735-6.

Roland:2007:SPE

[RVF07] Ch. Roland, R. Varadhan, and C. E. Frangakis. Squared polynomial extrapolation methods with cycling: an application to the positron emission tomography problem. Numerical Algorithms, 44(2):159–172, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=44&issue=2&spage=159.

Rutka:2006:EJI

[RW06] Vita Rutka and Andreas Wiegmann. Explicit jump immersed interface method for virtual material design of the effective elastic moduli of composite materials. Numerical Algorithms, 43(4):309–330, December 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=4&spage=309.

Ruffer:2011:SVM

[RW11] Björn S. Rüffer and Fabian R. Wirth. Stability verification for monotone systems using homotopy algorithms. Numerical Algorithms, 58(4):529–543, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=529.

Ran:2024:BCC

[RW24] Yu-Hong Ran and Qian-Qian Wu. Block-circulant with circulant-block preconditioners for two-dimensional spatial fractional diffusion equations. Numerical Algorithms, 97(4): 1935–1953, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01774-3.

Ren:2009:NVJ

[RWB09] Hongmin Ren, Qingbiao Wu, and Weihong Bi. New variants of Jarratt’s method with sixth-order convergence. Numerical Algorithms, 52(4):585–603, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=4&spage=585.

Ren:2021:DSG

[RWTM21] Longfei Ren, Chengjing Wang, Peipei Tang, and Zheng Ma. A dual symmetric Gauss–Seidel alternating direction method of multipliers for hyperspectral sparse unmixing. Numerical Algorithms, 87(2):719–754, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00985-8.

Ren:2019:PGT

[RWTW19] Huan Ren, Xiang Wang, Xiao-Bin Tang, and Teng Wang. A preconditioned general two-step modulus-based matrix splitting iteration method for linear complementarity problems of H_{+}-matrices. Numerical Algorithms, 82(3):969–986, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Redivo-Zaglia:1999:P

[RZ99a] M. Redivo-Zaglia. Preface. Numerical Algorithms, 20(2–3):0, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/1/fulltext.pdf.

Redivo-Zaglia:1999:SLN

[RZ99b] M. Redivo-Zaglia. The software library numeralgo. Numerical Algorithms, 20(2–3):107–116, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
18/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/18/2/fulltext.pdf.

Reich:2016:MSA

[RZ16] Simeon Reich and Rafal Zalas. A modular string averaging procedure for solving the common fixed point problem for quasi-nonexpansive mappings in Hilbert space. Numerical Algorithms, 72(2):297–323, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0045-z.

Redivo-Zaglia:2023:P

[RZ23a] Michela Redivo-Zaglia. Preface. Numerical Algorithms, 92 (1):1–3, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01362-3.

Reich:2023:PEM

[RZ23b] Simeon Reich and Rafał Zalas. Polynomial estimates for the method of cyclic projections in Hilbert spaces. Numerical Algorithms, 94(3):1217–1242, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01533-w.

Redivo-Zaglia:2012:SMT

[RZR12] Michela Redivo-Zaglia and Giuseppe Rodriguez. smt: a Matlab toolbox for structured matrices. Numerical Algorithms, 59(4):639–659, April 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=639.

Ren:2025:SOM

[RZZ25] Huimin Ren, Qi Zhang, and Zhongguo Zhou. The secondorder modified upwind PPM characteristic difference method and analysis for solving convection-diffusion equations. Numerical Algorithms, 98(2):503–530, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01803-1.

Sharma:2014:NDF

[SA14] Janak Raj Sharma and Himani Arora. A novel derivative free algorithm with seventh order convergence for solving systems of nonlinear equations. Numerical Algorithms, 67(4):917– 933, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9832-1.

Souleymane:2023:BVP

[SA23] Kadri Harouna Souleymane and Kaïs Ammari. Boundary value problems initial condition identification by a waveletbased Galerkin method. Numerical Algorithms, 93(1):397– 414, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01421-9.

Spyropoulos:2026:NSM

[SA26] Antony Spyropoulos and Christos Antonopoulos. Numerical study of mixed precision GMRES(m) preconditioned by deflation. Numerical Algorithms, 101(4):2631–2657, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02102-z.

Saad:2023:RBJ

[Saa23] Yousef Saad. Revisiting the (block) Jacobi subspace rotation method for the symmetric eigenvalue problem. Numerical Algorithms, 92(1):917–944, January 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01377-w.

Sablonniere:1991:CFA

[Sab91] Paul Sablonnière. Comparison of four algorithms accelerating the convergence of a subset of logarithmic fixed point sequences. Numerical Algorithms, 1(2):177–197, ???? 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sablonniere:1992:ABI

[Sab92a] Paul Sablonnière. Asymptotic behaviour of iterated modified Δ^{2} and θ_{2} transforms on some slowly convergent sequences. Numerical Algorithms, 3(1–4):401–409, December 1992. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Sablonniere:1992:SNG

[Sab92b] Paul Sablonnière. A short note on the generalized Euler transform for the summation of power series. Numerical Algorithms, 2(3–4):241–254, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sablonniere:2003:ACH

[Sab03] Paul Sablonnière. An algorithm for the computation of Hermite–Padé approximations to the exponential function: Divided differences and Hermite–Padé; forms. Numerical Algorithms, 33(1–4):443–452, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
36/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/36/fulltext.pdf.

Sablonniere:2014:PTA

[Sab14] P. Sablonnière. Padé-type approximants for generalized Euler transforms. Numerical Algorithms, 66(2):339–347, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9737-4.

Sadri:2018:NNM

[SAC18] K. Sadri, A. Amini, and C. Cheng. A new numerical method for delay and advanced integro-differential equations.

Numerical Algorithms, 77(2):381–412, February 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-017-0320-2.

Sadok:1999:CNM

[Sad99] H. Sadok. CMRH: a new method for solving nonsymmetric linear systems based on the Hessenberg reduction algorithm. Numerical Algorithms, 20(4):303–321, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/19/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
19/1/fulltext.pdf.

Sadok:2005:ACM

[Sad05] H. Sadok. Analysis of the convergence of the minimal and the orthogonal residual methods. Numerical Algorithms, 40(2):201–216, October 2005. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=2&spage=201.

Schaumburg:2019:PIB

[SAE19] Herman D. Schaumburg, Afnan Al Marzouk, and Bela Erdelyi. Picard iteration-based variable-order integrator with dense output employing algorithmic differentiation. Numerical Algorithms, 80(2):377–396, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Safouhi:2010:BSC

[Saf10] Hassan Safouhi. Bessel, sine and cosine functions and extrapolation methods for computing molecular multi-center integrals. Numerical Algorithms, 54(1):141–167, May 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=141.

Sharifi:2025:IET

[SAHM25] Mohammad Sharifi, Ali Abdi, Gholamreza Hojjati, and Aida Mousavi. Implicit–explicit two-step peer methods with RK stability for implicit part. Numerical Algorithms, 98(4):2145– 2170, April 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01867-z.

Salam:1994:NCE

[Sal94] A. Salam. Non-commutative extrapolation algorithms. Numerical Algorithms, 7(2–4):225–251, July 1994. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Salam:1996:AAV

[Sal96] A. Salam. An algebraic approach to the vector ϵ-algorithm. Numerical Algorithms, 11(1–4):327–337, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1994).

Salam:2005:TNA

[Sal05] Ahmed Salam. On theoretical and numerical aspects of symplectic Gram–Schmidt-like algorithms. Numerical Algorithms, 39(4):437–462, August 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Salehi:2017:MPC

[Sal17] Rezvan Salehi. A meshless point collocation method for 2-D multi-term time fractional diffusion-wave equation. Numerical Algorithms, 74(4):1145–1168, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0190-z.

Salam:2025:EFV

[Sal25] Ahmed Salam. The explicit form of the vector ε-algorithm’s kernel: a conjecture and its solution in R^{4}. Numerical Algorithms, 100(4):1639–1655, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02074-0.

Sandu:2014:NLC

[San14] Adrian Sandu. A new look at the chemical master equation. Numerical Algorithms, 65(3):485–498, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9758-z.

Sang:2019:NBT

[San19] Caili Sang. A new Brauer-type Z-eigenvalue inclusion set for tensors. Numerical Algorithms, 80(3):781–794, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Santra:2025:AHO

[San25] Sudarshan Santra. Analysis of a higher-order scheme for multiterm time-fractional integro-partial differential equations with multi-term weakly singular kernels. Numerical Algorithms, 99(4):??, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01927-4.

Sarra:2006:DTV

[Sar06] Scott A. Sarra. Digital total variation filtering as postprocessing for Chebyshev pseudospectral methods for conservation laws. Numerical Algorithms, 41(1):17–33, January 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=41&issue=1&
spage=17.

Sastry:1993:ACH

[Sas93] Venkat V. S. S. Sastry. Algorithms for the computation of Hankel functions of complex order. Numerical Algorithms, 5(1–4): 621–628, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Sato:2024:CHO

[Sat24a] Shun Sato. Correction to: High-order linearly implicit exponential integrators conserving quadratic invariants with application to scalar auxiliary variable approach. Numerical Algorithms, 96(4):1945–1946, August 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01807-x. See [Sat24b].

Sato:2024:HOL

[Sat24b] Shun Sato. High-order linearly implicit exponential integrators conserving quadratic invariants with application to scalar auxiliary variable approach. Numerical Algorithms, 96 (3):1295–1329, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01781-4. See

correction [Sat24a].

Sauer:2007:AVA

[Sau07] Tomas Sauer. Approximate varieties, approximate ideals and dimension reduction. Numerical Algorithms, 45(1–4):295–313, August 2007. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-007-9112-4.

Shanazari:2021:MMN

[SB21] Kamal Shanazari and Siamak Banei. A meshfree method with a non-overlapping domain decomposition method based on TPS for solving the forward-backward heat equation in two-dimension. Numerical Algorithms, 86(4):1747–1767, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00952-3.

Smotzer:2024:AHI

[SB24] Thomas Smotzer and John Buoni. Another HSS iteration method. Numerical Algorithms, 97(2):889–896, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01728-1.

Sabiu:2026:HSP

[SBA26] Jamilu Sabi’u, Ado Balili, and Mohammed Alsubhi. A hybrid spectral projection-based method for solving monotone nonlinear systems and signal recovery. Numerical Algorithms, 101 (1):119–147, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01998-3.

Sadeghi:2024:IHD

[SBG24] Hamed Sadeghi, Sebastian Banert, and Pontus Giselsson. Incorporating history and deviations in forward-backward splitting. Numerical Algorithms, 96(4):1819–1867, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01686-8.

Sun:2019:CRA

[SBJC19] Tao Sun, Roberto Barrio, Hao Jiang, and Lizhi Cheng. Convergence rates of accelerated proximal gradient algorithms under independent noise. Numerical Algorithms, 81(2):631–654, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

SabateLandman:2025:RIP

[SBN25] Malena Sabaté Landman, Ariana N. Brown, and James G. Nagy. Randomized and inner-product free Krylov methods for large-scale inverse problems. Numerical Algorithms, 100(4): 1965–1985, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02167-w.

Selvan:2026:ASI

[SBR26] A. Antony Selvan, Ayush Bhandari, and R. Radha. Approximation from shift-invariant spaces with smooth generators. Numerical Algorithms, 101(3):1833–1864, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02065-1.

Schmidt:1998:ACR

[SBW98] J. W. Schmidt and M. Bastian-Walther. Algorithm for constructing range restricted histosplines. Numerical Algorithms, 17(3–4):241–260, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
12/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/12/3/fulltext.pdf.

Serra-Capizzano:2003:PBT

[SC03a] S. Serra-Capizzano. Practical band Toeplitz preconditioning and boundary layer effects. Numerical Algorithms, 34 (2–4):427–440, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
48/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/48/9/fulltext.pdf.

Shanazari:2003:MDC

[SC03b] Kamal Shanazari and Ke Chen. A minimal distance constrained adaptive mesh algorithm with application to the dual reciprocity method. Numerical Algorithms, 32(2–4):275–286, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/8/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
45/8/fulltext.pdf.

Sunthrayuth:2018:IMS

[SC18] Pongsakorn Sunthrayuth and Prasit Cholamjiak. Iterative methods for solving quasi-variational inclusion and fixed point problem in q-uniformly smooth Banach spaces. Numerical Algorithms, 78(4):1019–1044, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sahu:2021:IRC

[SCD^{+}21] D. R. Sahu, Y. J. Cho, Q. L. Dong, M. R. Kashyap, and X. H. Li. Inertial relaxed CQ algorithms for solving a split feasibility problem in Hilbert spaces. Numerical Algorithms, 87(3): 1075–1095, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00999-2.

Shi:2020:NSC

[SCDM20] Lei Shi, Zhong Chen, Xiaohua Ding, and Qiang Ma. A new stable collocation method for solving a class of nonlinear fractional delay differential equations. Numerical Algorithms, 85(4):1123–1153, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00858-
9;   http://link.springer.com/content/pdf/10.1007/
s11075-019-00858-9.pdf.

Shi:2023:IGQ

[SCF23] Chengxin Shi, Hao Cheng, and Wenping Fan. An iterative generalized quasi-boundary value regularization method for the backward problem of time fractional diffusion-wave equation in a cylinder. Numerical Algorithms, 94(4):1619–1651, December 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01549-2.

Schatzman:2002:NIR

[Sch02] Michelle Schatzman. Numerical integration of reaction– diffusion systems. Numerical Algorithms, 31(1–4):247–269, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/43/5/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/43/5/fulltext.pdf.

Schumaker:2008:CBS

[Sch08] Larry L. Schumaker. Computing bivariate splines in scattered data fitting and the finite-element method. Numerical Algorithms, 48(1–3):237–260, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=1&spage=237.

Schwarz:2009:CID

[Sch09] Diana Estévez Schwarz. Consistent initialization for DAEs in Hessenberg form. Numerical Algorithms, 52(4):629–648, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
52&issue=4&spage=629.

Schaback:2014:GSL

[Sch14] Robert Schaback. Greedy sparse linear approximations of functionals from nodal data. Numerical Algorithms, 67(3):531– 547, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9806-8.

Schweitzer:2017:TSS

[Sch17] Marcel Schweitzer. A two-sided short-recurrence extended Krylov subspace method for nonsymmetric matrices and its relation to rational moment matching. Numerical Algorithms, 76(1):1–31, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Schmidt:2025:GCN

[Sch25] Niko Schmidt. Global convergence of Newton’s method for the regularized p-Stokes equations. Numerical Algorithms, 99 (4):1983–2010, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01941-6.

Schaback:2026:GAL

[Sch26] Robert Schaback. Greedy adaptive local recovery of functions in Sobolev spaces. Numerical Algorithms, 102(1):611– 635, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02142-5.

Suparatulatorn:2018:MIP

[SCS18] Raweerote Suparatulatorn, Watcharaporn Cholamjiak, and Suthep Suantai. A modified S-iteration process for Gnonexpansive mappings in Banach spaces with graphs. Numerical Algorithms, 77(2):479–490, February 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-017-0324-y.

Serra-Capizzano:2000:CTA

[SCTP00] Stefano Serra-Capizzano and Cristina Tablino-Possio. Constructive techniques for approximating collocation linear systems. Numerical Algorithms, 25(1–4):323–339, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/13/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/13/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Shen:2017:TID

[SCW17] Qin-Qin Shen, Yang Cao, and Li Wang. Two improvements of the deteriorated PSS preconditioner for generalized saddle point problems. Numerical Algorithms, 75(1):33–54, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shahraki:2020:ASP

[SD20] M. Sayadi Shahraki and A. Delavarkhalafi. An arcsearch predictor–corrector infeasible-interior-point algorithm for P_{∗}(κ)-SCLCPs. Numerical Algorithms, 83(4):1555–1575, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shirilord:2025:SSC

[SD25] Akbar Shirilord and Mehdi Dehghan. Solving a system of complex matrix equations using a gradient-based method and its application in image restoration. Numerical Algorithms, 98 (4):1849–1896, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01856-2.

Shirilord:2026:SES

[SD26] Akbar Shirilord and Mehdi Dehghan. Solvability, exact solution and momentum acceleration gradient-based iterative algorithm for general matrix equation AXB + C\overline{X}D + EX^{T} F + GX^{∗}H = J and studying its applications in discrete-time antilinear systems and color image restoration. Numerical Algorithms, 101(4):2505–2566, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02097-7.

Shi:2023:DGS

[SDL^{+}23] Yang Shi, Chenling Ding, Shuai Li, Bin Li, and Xiaobing Sun. Discrete generalized-Sylvester matrix equation solved by RNN with a novel direct discretization numerical method. Numerical Algorithms, 93(3):971–992, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01449-x.

Sitthithakerngkiet:2018:CAG

[SDMMK18] Kanokwan Sitthithakerngkiet, Jitsupa Deepho, Juan Martínez-Moreno, and Poom Kumam. Convergence analysis of a general iterative algorithm for finding a common solution of split variational inclusion and optimization problems. Numerical Algorithms, 79(3):801–824, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shen:2024:CSA

[SDW24] Peiping Shen, Yaping Deng, and Dianxiao Wu. A criterion space algorithm for solving linear multiplicative programming problems. Numerical Algorithms, 96(4):1901–1923, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01689-5.

Segura:1998:GNM

[Seg98] Javier Segura. A global Newton method for the zeros of cylinder functions. Numerical Algorithms, 18(3– 4):259–276, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/2/fulltext.pdf.

Segura:2008:IZC

[Seg08] Javier Segura. Interlacing of the zeros of contiguous hypergeometric functions. Numerical Algorithms, 49(1–4):387–407, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
49&issue=1&spage=387.

Scolnik:2014:IOP

[SEG14] H. Scolnik, N. Echebest, and M. T. Guardarucci. On the incomplete oblique projections method for solving box constrained least squares problems. Numerical Algorithms, 66 (1):17–32, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9721-z.

Seiler:1998:PVM

[Sei98] Werner M. Seiler. Position versus momentum projections for constrained Hamiltonian systems. Numerical Algorithms, 19 (1–4):223–234, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
16/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/16/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Samaey:2002:NCC

[SER02] Giovanni Samaey, Koen Engelborghs, and Dirk Roose. Numerical computation of connecting orbits in delay differential equations. Numerical Algorithms, 30(3–4):335–352, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/42/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
42/6/fulltext.pdf.

Sarumi:2023:EHO

[SFMK23] Ibrahim O. Sarumi, Khaled M. Furati, Kassem Mustapha, and Abdul Q. M. Khaliq. Efficient high-order exponential time differencing methods for nonlinear fractional differential models. Numerical Algorithms, 92(2):1261–1288, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01339-2.

Shi:2023:OEE

[SFS23] Kaiwen Shi, Xinlong Feng, and Haiyan Su. Optimal error estimate of the penalty method for the 2D/ 3D time-dependent MHD equations. Numerical Algorithms, 93(3):1337–1371, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01470-0.

Simos:2003:ZDE

[SFT03] T. E. Simos, I. T. Famelis, and C. Tsitouras. Zero dissipative, explicit Numerov-type methods for second order IVPs with oscillating solutions. Numerical Algorithms, 34 (1):27–40, September 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.kluweronline.com/content/getfile/5058/
47/3/abstract.htm;   http://ipsapp008.kluweronline.
com/content/getfile/5058/47/3/fulltext.pdf.

Slimani:2022:MFS

[SFZ22] S. Slimani, M. Farhloul, and A. Zine. Mixed formulation of a stationary seawater intrusion problem in confined aquifers. Numerical Algorithms, 91(2):651–669, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01277-z.

Sorensen:2010:QCM

[SG10] Danny C. Sorensen and Roland Glowinski. A quadratically constrained minimization problem arising from PDE of Monge–Ampère type. Numerical Algorithms, 53(1):53–66, January 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=53&
issue=1&spage=53.

Song:2017:NLM

[SG17] Linsen Song and Yan Gao. A nonsmooth Levenberg– Marquardt method for vertical complementarity problems. Numerical Algorithms, 76(2):473–485, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Song:2018:SLM

[SG18] Linsen Song and Yan Gao. A smoothing Levenberg–Marquardt method for nonlinear complementarity problems. Numerical Algorithms, 79(4):1305–1321, December 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sun:2023:HOM

[SG23] Zhengjie Sun and Yuyan Gao. High order multiquadric trigonometric quasi-interpolation method for solving time-dependent partial differential equations. Numerical Algorithms, 93(4): 1719–1739, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01486-6.

Shubha:2015:DFI

[SGJ15] Vorkady S. Shubha, Santhosh George, and P. Jidesh. A derivative free iterative method for the implementation of Lavrentiev regularization method for ill-posed equations. Numerical Algorithms, 68(2):289–304, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9844-x.

Stenger:1999:OIP

[SGK^{+}99] F. Stenger, S.-Å. Gustafson, B. Keyes, M. O’Reilly, and K. Parker. ODE-IVP-PACK via sinc indefinite integration and Newton’s method. Numerical Algorithms, 20(2–3):241–268, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/7/fulltext.pdf.

Salam:2002:VEA

[SGM02] A. Salam and P. R. Graves-Morris. On the vector ε-algorithm for solving linear systems of equations. Numerical Algorithms, 29(1–3):229–247, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp008.lwwonline.com/content/getfile/5058/38/
13/abstract.htm;   http://ipsapp008.lwwonline.com/
content/getfile/5058/38/13/fulltext.pdf.

Sequeira:2022:SAC

[SGO22] Filánder A. Sequeira and Helen Guillén-Oviedo. Some aspects on the computational implementation of diverse terms arising in mixed virtual element formulations. Numerical Algorithms, 89(2):487–528, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01123-8.

Sharma:2013:EFO

[SGS13] Janak Raj Sharma, Rangan Kumar Guha, and Rajni Sharma. An efficient fourth order weighted-Newton method for systems of nonlinear equations. Numerical Algorithms, 62(2): 307–323, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9585-7/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=2&spage=307-323.

Saupe:2025:ATS

[SGS25] Thomas Saupe, Sebastian Götschel, and Robert Speck. Adaptive time step selection for spectral deferred correction. Numerical Algorithms, 100(1):353–367, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01964-z. See correction [SGS26].

Saupe:2026:CAT

[SGS26] Thomas Saupe, Sebastian Götschel, and Robert Speck. Correction to: Adaptive time step selection for spectral deferred correction. Numerical Algorithms, 101(3):2155, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02191-w. See [SGS25].

Shanazari:2012:ANT

[SH12] Kamal Shanazari and Mohammad Hosami. Adapting nodes in three dimensional irregular regions for meshless-type methods. Numerical Algorithms, 61(1):83–103, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=61&issue=1&
spage=83.

Skalna:2017:NAC

[SH17] Iwona Skalna and Milan Hladík. A new algorithm for Chebyshev minimum-error multiplication of reduced affine forms. Numerical Algorithms, 76(4):1131–1152, December 2017. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Skalna:2021:PSE

[SH21a] Iwona Skalna and Milan Hladík. On preconditioning and solving an extended class of interval parametric linear systems. Numerical Algorithms, 87(4):1535–1562, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01018-0.

Sugiura:2021:TCC

[SH21b] Hiroshi Sugiura and Takemitsu Hasegawa. A truncated Clenshaw–Curtis formula approximates integrals over a semiinfinite interval. Numerical Algorithms, 86(2):659–674, February 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00905-w.

Sugiura:2022:ECC

[SH22] Hiroshi Sugiura and Takemitsu Hasegawa. Extensions of Clenshaw–Curtis-type rules to integrals over a semi-infinite interval. Numerical Algorithms, 90(1):3–30, May 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01177-8.

Sugiura:2023:TFN

[SH23] Hiroshi Sugiura and Takemitsu Hasegawa. Two formulae with nodes related to zeros of Bessel functions for semiinfinite integrals: extending Gauss–Jacobi-type rules. Numerical Algorithms, 94(4):1949–1981, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01560-7.

Sugiura:2026:TGR

[SH26] Hiroshi Sugiura and Takemitsu Hasegawa. On two Gauss– Radau quadrature formulae and their weighted averagedrule: Is Gauss–Legendre always better than Radau? Numerical Algorithms, 101(4):2333–2356, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02089-7.

Shardlow:1997:IML

[Sha97] Tony Shardlow. Inertial manifolds and linear multi-step methods. Numerical Algorithms, 14(1–3):189–209, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/5/12/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

5/12/fulltext.pdf. Dynamical numerical analysis (Atlanta, GA, 1995).

Sharp:2002:RPB

[Sha02] P. W. Sharp. Requirements of a package for N-body simulations of the solar system. Numerical Algorithms, 31(1– 4):271–279, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
43/6/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/43/6/fulltext.pdf.

Sharp:2019:PBI

[Sha19] P. W. Sharp. The performance of the N-body integrator SSS. Numerical Algorithms, 81(4):1459–1472, August 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shi:2025:FDM

[SHD25] Siyuan Shi, Zhaopeng Hao, and Rui Du. A finite difference method for elliptic equations with the variable-order fractional derivative. Numerical Algorithms, 99(3):1463–1501, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01922-9.

Shekhtman:2000:DPR

[She00] Boris Shekhtman. On the density principle for rational functions. Numerical Algorithms, 25(1–4):341–346, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/3/12/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

3/12/fulltext.pdf. Mathematical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Shen:2015:PEE

[She15] Quan Shen. A posteriori error estimates of the Morley element for the fourth order elliptic eigenvalue problem. Numerical Algorithms, 68(3):455–466, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9854-8.

Schmutzhard:2015:NSL

[SHF15] Sebastian Schmutzhard, Tomasz Hrycak, and Hans G. Feichtinger. A numerical study of the Legendre–Galerkin method for the evaluation of the prolate spheroidal wave functions. Numerical Algorithms, 68(4):691–710, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9867-3.

Song:2022:SCS

[SHGL22] Guoting Song, Junhao Hu, Shuaibin Gao, and Xiaoyue Li. The strong convergence and stability of explicit approximations for nonlinear stochastic delay differential equations. Numerical Algorithms, 89(2):855–883, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01137-2.

Shi:1996:GPS

[Shi96] Yixun Shi. A globalization procedure for solving nonlinear systems of equations. Numerical Algorithms, 12(3–4):273–286, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shen:2018:CUL

[SHLY18] Weiping Shen, Yaohua Hu, Chong Li, and Jen-Chih Yao. Convergence of a Ulm-like method for square inverse singular value problems with multiple and zero singular values. Numerical Algorithms, 79(2):375–398, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shokri:2018:NEO

[Sho18] Ali Shokri. A new eight-order symmetric two-step multiderivative method for the numerical solution of second-order IVPs with oscillating solutions. Numerical Algorithms, 77(1):95–109, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Si:2012:SOM

[Si12] Zhiyong Si. Second order modified method of characteristics mixed defect-correction finite element method for time dependent Navier–Stokes problems. Numerical Algorithms, 59(2):271–300, February 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=2&spage=271.

Segal:2013:ISF

[SI13] Ben Segal and M. A. Iwen. Improved sparse Fourier approximation results: faster implementations and stronger guarantees. Numerical Algorithms, 63(2):239–263, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9621-7.

Shehu:2017:SCR

[SI17] Yekini Shehu and Olaniyi S. Iyiola. Strong convergence result for monotone variational inequalities. Numerical Algorithms, 76(1):259–282, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Shehu:2018:IAS

[SI18] Yekini Shehu and Olaniyi S. Iyiola. Iterative algorithms for solving fixed point problems and variational inequalities with uniformly continuous monotone operators. Numerical Algorithms, 79(2):529–553, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sidi:1994:CIR

[Sid94] Avram Sidi. Convergence of intermediate rows of minimal polynomial and reduced rank extrapolation tables. Numerical Algorithms, 6(3–4):229–244, March 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sidorov:2007:OIC

[Sid07] S. P. Sidorov. Optimal interpolation of convergent algebraic series. Numerical Algorithms, 44(3):273–279, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=44&issue=3&
spage=273.

Sidi:2017:VVG

[Sid17] Avram Sidi. On a vectorized version of a generalized Richardson extrapolation process. Numerical Algorithms, 74(3):937– 949, March 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-016-0178-8.

Sidi:2020:ACS

[Sid20a] Avram Sidi. Acceleration of convergence of some infinite sequences A_{n} whose asymptotic expansions involve fractional powers of n via the d̃^{(m)} transformation. Numerical Algorithms, 85(4):1409–1445, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00870-
z.

Sidi:2020:CSR

[Sid20b] Avram Sidi. A convergence study for reduced rank extrapolation on nonlinear systems. Numerical Algorithms, 84(3): 957–982, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00788-6.

Shehu:2016:IAS

[SIE16] Y. Shehu, O. S. Iyiola, and C. D. Enyi. An iterative algorithm for solving split feasibility problems and fixed point problems in Banach spaces. Numerical Algorithms, 72(4):835–864, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0069-4.

Simeon:1998:DPE

[Sim98] B. Simeon. DAEs and PDEs in elastic multibody systems. Numerical Algorithms, 19(1–4):235–246, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/16/15/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/16/

15/fulltext.pdf. Differential algebraic equations (Grenoble, 1997).

Sintamarian:2007:GEC

[Sîn07] Alina Sîntamarian. A generalization of Euler’s constant. Numerical Algorithms, 46(2):141–151, October 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=2&spage=141.

Shehu:2020:IMI

[SIO20] Yekini Shehu, Olaniyi S. Iyiola, and Ferdinard U. Ogbuisi. Iterative method with inertial terms for nonexpansive mappings: applications to compressed sensing. Numerical Algorithms, 83 (4):1321–1347, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Song:2014:ANP

[SJ14] Bo Song and Yao-Lin Jiang. Analysis of a new parareal algorithm based on waveform relaxation method for time-periodic problems. Numerical Algorithms, 67(3):599–622, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9810-z.

Song:2021:ATN

[SJW21] Bo Song, Yao-Lin Jiang, and Xiaolong Wang. Analysis of two new parareal algorithms based on the Dirichlet–Neumann/ Neumann–Neumann waveform relaxation method for the heat equation. Numerical Algorithms, 86(4):1685–1703, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00949-y.

Smyrlis:2004:LLS

[SK04] Yiorgos-Sokratis Smyrlis and Andreas Karageorghis. A linear least-squares MFS for certain elliptic problems. Numerical Algorithms, 35(1):29–44, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/49/A/1/abstract.htm.

Siriyan:2019:NGS

[SK19] Keerati Siriyan and Atid Kangtunyakarn. A new general system of variational inequalities for convergence theorem and application. Numerical Algorithms, 81(1):99–123, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Singh:2025:EAH

[SK25] Anshima Singh and Sunil Kumar. Error analysis of a highorder fully discrete method for two-dimensional time-fractional convection-diffusion equations exhibiting weak initial singularity. Numerical Algorithms, 99(1):251–284, May 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01877-x.

Skalna:2013:AMR

[Ska13] Iwona Skalna. Algorithm for min-range multiplication of affine forms. Numerical Algorithms, 63(4):601–614, August 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-012-9644-0.pdf.

S:2023:DQP

[SKA23] Aswin V. S., Riyasudheen T. K., and Ashish Awasthi. Differential quadrature parallel algorithms for solving systems of convection–diffusion and reaction models. Numerical Algorithms, 93(1):321–346, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01416-

Shao:2026:SAE

[SKD26] Zehui Shao, Saeed Kosari, and MohammadHossein Derakhshan. Stability analysis of an efficient hybrid numerical approach for solving the two-dimensional space/multitime fractional Bloch–Torrey model involving the Riesz fractional operator. Numerical Algorithms, 102(1):721–747, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02151-4.

Svelander:2018:RIS

[SKJ^{+}18] Frida Svelander, Gustav Kettil, Tomas Johnson, Andreas Mark, Anders Logg, and Fredrik Edelvik. Robust intersection of structured hexahedral meshes and degenerate triangle meshes with volume fraction applications. Numerical Algorithms, 77(4):1029–1068, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11075-017-0352-7.
pdf.

Sahu:2021:PPA

[SKK21] D. R. Sahu, Ajeet Kumar, and Shin Min Kang. Proximal point algorithms based on S-iterative technique for nearly asymptotically quasi-nonexpansive mappings and applications. Numerical Algorithms, 86(4):1561–1590, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00945-2.

Salihu:2025:SCT

[SKK25] Nasiru Salihu, Poom Kumam, and Wiyada Kumam. Some combined techniques of spectral conjugate gradient methods with applications to robotic and image restoration models. Numerical Algorithms, 100(1):529–551, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01970-1.

Singh:2025:FLW

[SKM25] Akanksha Singh, Ankur Kanaujiya, and Jugal Mohapatra. Fractional Legendre wavelet approach resolving multi-scale optimal control problems involving Caputo–Fabrizio derivative. Numerical Algorithms, 99(1):101–132, May 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01871-3.

Sharma:2020:PEE

[SKP20] Nisha Sharma, Morrakot Khebchareon, and Amiya K. Pani. A priori error estimates of expanded mixed FEM for Kirchhoff type parabolic equation. Numerical Algorithms, 83(1):125–147, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sapir:2021:NVS

[SKSS21] Luba Sapir, Tamara Kogan, Ariel Sapir, and Amir Sapir. Nonstationary vs. stationary iterative processes. Numerical Algorithms, 86(2):515–535, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00899-
5.

Shokri:2019:NFT

[SKTGR19] Ali Shokri, Mohammad Mehdizadeh Khalsaraei, Mortaza Tahmourasi, and Raquel Garcia-Rubio. A new family of threestage two-step p-stable multiderivative methods with vanished phase-lag and some of its derivatives for the numerical solution of radial Schrödinger equation and IVPs with oscillating solutions. Numerical Algorithms, 80(2):557–593, February 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Schwarz:2015:DSPb

[SL15a] Diana Estévez Schwarz and René Lamour. Diagnosis of singular points of properly stated DAEs using automatic differentiation. Numerical Algorithms, 70(4):777–805, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9973-x.

Schwarz:2015:DSPa

[SL15b] Diana Estévez Schwarz and René Lamour. Diagnosis of singular points of structured DAEs using automatic differentiation. Numerical Algorithms, 69(4):667–691, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9919-8.

Schwarz:2016:NPB

[SL16] Diana Estévez Schwarz and René Lamour. A new projector based decoupling of linear DAEs for monitoring singularities. Numerical Algorithms, 73(2):535–565, October 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0107-x.

Schwarz:2018:NAC

[SL18] Diana Estévez Schwarz and René Lamour. A new approach for computing consistent initial values and Taylor coefficients for DAEs using projector-based constrained optimization. Numerical Algorithms, 78(2):355–377, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Schwarz:2021:PEI

[SL21a] Diana Estévez Schwarz and René Lamour. Projected explicit and implicit Taylor series methods for DAEs. Numerical Algorithms, 88(2):615–646, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01051-
z;   http://link.springer.com/content/pdf/10.1007/
s11075-020-01051-z.pdf.

Shi:2021:SAT

[SL21b] Dongyang Shi and Chaoqun Li. Superconvergence analysis of two-grid methods for bacteria equations. Numerical Algorithms, 86(1):123–152, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00882-
0.

Slawig:2006:PCC

[Sla06] Thomas Slawig. PDE-constrained control using Femlab-Control of the Navier–Stokes equations. Numerical Algorithms, 42(2):107–126, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=42&issue=2&spage=107.

Shen:2011:NAS

[SLA11] Shujun Shen, Fawang Liu, and Vo Anh. Numerical approximations and solution techniques for the space–time Riesz–Caputo fractional advection-diffusion equation. Numerical Algorithms, 56(3):383–403, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=3&spage=383.

Shehu:2020:EPT

[SLD20] Yekini Shehu, Xiao-Huan Li, and Qiao-Li Dong. An efficient projection-type method for monotone variational inequalities in Hilbert spaces. Numerical Algorithms, 84(1):365–388, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Slevinsky:2025:FSR

[Sle25] Richard Mikaël Slevinsky. Fast and stable rational approximation of generalized hypergeometric functions. Numerical Algorithms, 98(2):587–624, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01808-w.

Sun:2022:SAS

[SLL22] Wumei Sun, Hongwei Liu, and Zexian Liu. Several accelerated subspace minimization conjugate gradient methods based on regularization model and convergence rate analysis for nonconvex problems. Numerical Algorithms, 91(4):1677–1719, December 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01319-6.

Sun:2023:RLM

[SLL23] Wumei Sun, Hongwei Liu, and Zexian Liu. A regularized limited memory subspace minimization conjugate gradient method for unconstrained optimization. Numerical Algorithms, 94(4):1919–1948, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01559-0.

Shen:2015:NSA

[SLLA15] S. Shen, F. Liu, Q. Liu, and V. Anh. Numerical simulation of anomalous infiltration in porous media. Numerical Algorithms, 68(3):443–454, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9853-9.

Si:2020:UOE

[SLT20] Zhiyong Si, Yanfang Lei, and Zhang Tong. Unconditional optimal error estimate of the projection/Lagrange–Galerkin finite element method for the Boussinesq equations. Numerical Algorithms, 83(2):669–700, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Samson:2013:NTS

[SLW13] Michael Daniel Samson, Huiyuan Li, and Li-Lian Wang. A new triangular spectral element method I: implementation and analysis on a triangle. Numerical Algorithms, 64(3):519–547, November 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9677-4.

Sun:2025:DQD

[SLZ25] Jianhua Sun, Ying Li, and Mingcui Zhang. Dual quaternion QR decompositon and its corresponding complex structurepreserving algorithms. Numerical Algorithms, 100(3):1315– 1331, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01989-4.

Stanimirovic:2010:AGD

[SM10] Predrag S. Stanimirović and Marko B. Miladinović. Accelerated gradient descent methods with line search. Numerical Algorithms, 54(4):503–520, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=4&spage=503.

Salkuyeh:2017:NRH

[SM17] Davod Khojasteh Salkuyeh and Mohsen Masoudi. A new relaxed HSS preconditioner for saddle point problems. Numerical Algorithms, 74(3):781–795, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0171-2.

Shary:2021:SIL

[SM21] Sergey P. Shary and Behnam Moradi. Solving interval linear least squares problems by PPS-methods. Numerical Algorithms, 87(1):41–75, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00958-x.

Sedaghat:2025:IRE

[SM25a] Salameh Sedaghat and Francisco Marcellán. Integral representations of Eta functions and fractional calculus. Numerical Algorithms, 99(1):491–504, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01885-
x.

Spiteri:2025:NSS

[SM25b] Pierre Spiteri and Frédéric Messine. Numerical schemes for the solution of the telegrapher equation. Numerical Algorithms, 100(4):1715–1738, December 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02099-5.

Spitaleri:1999:FDS

[SMA99] Rosa Maria Spitaleri, Riccardo March, and Daniele Arena. Finite difference solution of Euler equations arising in variational image segmentation. Numerical Algorithms, 21(1– 4):353–365, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
20/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/20/2/fulltext.pdf.

Spiteri:2002:EPO

[SMB02] P. Spiteri, J. C. Miellou, and J. M. Bahi. Evaluations of parameters for the optimization of S.S.O.R. and A.D.I. preconditioning. Numerical Algorithms, 29(1–3):249–265, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.lwwonline.com/
content/getfile/5058/38/14/abstract.htm;   http://
ipsapp008.lwwonline.com/content/getfile/5058/38/14/
fulltext.pdf.

Spiteri:2003:PAS

[SME03] Pierre Spiteri, Jean-Claude Miellou, and Didier El Baz. Parallel asynchronous Schwarz and multisplitting methods for a nonlinear diffusion problem. Numerical Algorithms, 33 (1–4):461–474, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
38/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/38/fulltext.pdf.

Smiley:1997:NBS

[Smi97] M. W. Smiley. Numerical bifurcation and stability analysis for steady-states of reaction diffusion equations. Numerical Algorithms, 14(1–3):211–225, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/13/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/13/fulltext.pdf. Dynami-

cal numerical analysis (Atlanta, GA, 1995).

Smietanski:2006:ANM

[Śmi06] Marek J. Śmietański. An approximate Newton method for non-smooth equations with finite max functions. Numerical Algorithms, 41(3):219–238, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=3&spage=219.

Smietanski:2009:CGN

[Śmi09] Marek J. Śmietański. Convergence of a generalized Newton and an inexact generalized Newton algorithms for solving nonlinear equations with nondifferentiable terms. Numerical Algorithms, 50(4):401–415, April 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=4&spage=401.

Smietanski:2013:PVI

[Śmi13] Marek J. Śmietański. A perturbed version of an inexact generalized Newton method for solving nonsmooth equations. Numerical Algorithms, 63(1):89–106, May 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9613-7;   http://link.springer.com/
content/pdf/10.1007/s11075-012-9613-7.pdf.

Shateyi:2014:NPS

[SMK14] S. Shateyi, S. S. Motsa, and Y. Khan. A new piecewise spectral homotopy analysis of the Michaelis–Menten enzymatic reactions model. Numerical Algorithms, 66(3):495–510, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9745-4.

Sharma:2025:HAT

[SMM25] Divya Sharma, Kajal Mittal, and R. K. Mohanty. High accuracy two-level compact implicit method in exponential form for 2D fourth order quasi-linear parabolic equations. Numerical Algorithms, 100(3):1239–1287, November 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01987-6.

Solhi:2024:EML

[SMN24] Erfan Solhi, Farshid Mirzaee, and Shiva Naserifar. Enhanced moving least squares method for solving the stochastic fractional Volterra integro-differential equations of Hammerstein type. Numerical Algorithms, 95(4):1921–1951, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01633-7.

Slimani:2020:ISF

[SMNZ20] S. Slimani, I. Medarhri, K. Najib, and A. Zine. Identification of the source function for a seawater intrusion problem in unconfined aquifer. Numerical Algorithms, 84(4):1565–1587, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00925-6.

Smoch:1999:SRA

[Smo99] L. Smoch. Some results about GMRES in the singular case. Numerical Algorithms, 22(2):193–212, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.

com/content/getfile/5058/22/6/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
22/6/fulltext.pdf.

Singh:2026:NVA

[SMR26] Ankit Singh, Vikas Maurya, and Manoj K. Rajpoot. Numerical valuation of American options with liquidity shocks using IMEX methods. Numerical Algorithms, 102(1):685– 720, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02148-z.

Shahraki:2018:WNP

[SMZMA18] M. Sayadi Shahraki, H. Mansouri, M. Zangiabadi, and N. Mahdavi-Amiri. A wide neighborhood primal-dual predictor-corrector interior-point method for symmetric cone optimization. Numerical Algorithms, 78(2):535–552, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Singh:2022:SEE

[SN22] Gautam Singh and Srinivasan Natesan. Superconvergence error estimates of discontinuous Galerkin time stepping for singularly perturbed parabolic problems. Numerical Algorithms, 90(3):1073–1090, July 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01222-6.

Seal:2024:NAN

[SN24] Aniruddha Seal and Srinivasan Natesan. A numerical approach for nonlinear time-fractional diffusion equation with generalized memory kernel. Numerical Algorithms, 97(2):539–565, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01714-7.

Soderlind:2002:ACA

[Söd02] Gustaf Söderlind. Automatic control and adaptive timestepping. Numerical Algorithms, 31(1–4):281–310, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/12/abstract.htm;   http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
12/fulltext.pdf.

Soleymani:2011:RJM

[Sol11] Fazlollah Soleymani. Revisit of Jarratt method for solving nonlinear equations. Numerical Algorithms, 57(3):377–388, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=3&
spage=377.

Soleymani:2015:ESN

[Sol15] Fazlollah Soleymani. An efficient and stable Newton-type iterative method for computing generalized inverse A_{T,S}^{(2)}. Numerical Algorithms, 69(3):569–578, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9913-1.

Solary:2023:MPD

[Sol23] Maryam Shams Solary. From matrix polynomial to determinant of block Toeplitz–Hessenberg matrix. Numerical Algorithms, 94(3):1421–1434, November 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01541-w.

Sommariva:2005:FNA

[Som05] Alvise Sommariva. A fast Nyström–Broyden solver by Chebyshev compression. Numerical Algorithms, 38(1–3):47–60, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sonar:1993:DUS

[Son93] Th. Sonar. On the design of an upwind scheme for compressible flow on general triangulations. Numerical Algorithms, 4(1–2): 135–149, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Soodhalter:2015:BMA

[Soo15] Kirk M. Soodhalter. A block MINRES algorithm based on the band Lanczos method. Numerical Algorithms, 69(3):473– 494, July 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9907-z.

Sajjadi:2021:CAP

[SP21] Sayed Arsalan Sajjadi and Saeed Pishbin. Convergence analysis of the product integration method for solving the fourth kind integral equations with weakly singular kernels. Numerical Algorithms, 86(1):25–54, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00877-x.

Spalevic:2007:NGA

[Spa07] Miodrag M. Spalević. A note on generalized averaged Gaussian formulas. Numerical Algorithms, 46(3):253–264, November 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=46&issue=3&
spage=253.

Spalevic:2020:NGA

[Spa20] Miodrag M. Spalević. A note on generalized averaged Gaussian formulas for a class of weight functions. Numerical Algorithms, 85(3):977–993, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00848-x.

Spalevic:2024:MAG

[Spa24] Miodrag M. Spalević. Modified anti-Gaussian quadrature formulae of Chebyshev type. Numerical Algorithms, 95(3): 1347–1357, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01611-z.

Sahu:2026:SEE

[SPB26] Pradeep Kumar Sahu, Kuldip Singh Patel, and Muddun Bhuruth. Stability and error estimates of operator-splitting methods for pricing American option under regime-switching model with jumps. Numerical Algorithms, 102(2):1317–1353, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02178-7.

Spillane:2025:GSC

[Spi25] Nicole Spillane. GenEO spectral coarse spaces in SPD domain decomposition. Numerical Algorithms, 100(4):1987–2020, December 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02166-x.

Sprengel:1998:PIW

[Spr98] Frauke Sprengel. Periodic interpolation and wavelets on sparse grids. Numerical Algorithms, 17(1–2):147–169, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/31/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
31/7/fulltext.pdf.

Sprengel:2001:MAF

[Spr01] Frauke Sprengel. Multilevel algorithms for finite difference discretizations on sparse grids. Numerical Algorithms, 26 (2):111–121, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
1/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/1/2/fulltext.pdf.

Sriwastav:2026:RLC

[SPS26] Nikhil Sriwastav, R N Prajapati, and Mehakpreet Singh. A robust Legendre collocation scheme and convergence analysis for Lane–Emden multi-pantograph delay differential equations arising in stellar physics including Chandrasekhar’s white dwarf problem. Numerical Algorithms, 102(2):779–808, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02153-2.

Sahu:2020:NIT

[SPV20] D. R. Sahu, A. Pitea, and M. Verma. A new iteration technique for nonlinear operators as concerns convex programming and feasibility problems. Numerical Algorithms, 83(2):421– 449, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sharp:2013:HOE

[SQG13] Philip W. Sharp, Mohammad A. Qureshi, and Kevin R. Grazier. High order explicit Runge–Kutta Nyström pairs. Numerical Algorithms, 62(1):133–148, January 2013. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-012-9571-0/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=1&spage=133-148.

Sosa:2004:NAM

[SR04] Wilfredo Sosa and Fernanda M. P. Raupp. A new approach to a multicriteria optimization problem. Numerical Algorithms, 35(2–4):233–247, April 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/50/A/14/abstract.htm.

Studniarski:2006:ACS

[SR06] Marcin Studniarski and El Desouky Rahmo. Approximating Clarke’s subgradients of semismooth functions by divided differences. Numerical Algorithms, 43(4):385–392, December 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=43&issue=4&
spage=385.

Sayevand:2016:GFV

[SR16] K. Sayevand and M. R. Rostami. General fractional variational problem depending on indefinite integrals. Numerical Algorithms, 72(4):959–987, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0076-5.

Siettos:2022:NMA

[SR22] Constantinos Siettos and Lucia Russo. A numerical method for the approximation of stable and unstable manifolds of microscopic simulators. Numerical Algorithms, 89(3):1335–1368, March 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01155-0;   http://link.
springer.com/content/pdf/10.1007/s11075-021-01155-
0.pdf.

Singh:2024:NAB

[SR24a] P. K. Singh and S. Saha Ray. A numerical approach based on Pell polynomial for solving stochastic fractional differential equations. Numerical Algorithms, 97(4):1513–1534, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01760-9.

Song:2024:ROF

[SR24b] Junpeng Song and Hongxing Rui. Reduced-order finite element approximation based on POD for the parabolic optimal control problem. Numerical Algorithms, 95(3):1189–1211, March 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01605-x.

Sayana:2025:CPP

[SR25] K. J. Sayana and G. D. Reddy. A class of a posteriori parameter choice rules for filter-based regularization schemes. Numerical Algorithms, 98(2):767–795, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01815-x.

Sharma:2025:EIP

[SRK^{+}25] Puneet Sharma, Higinio Ramos, Vinay Kanwar, Ramandeep Behl, and Mithil Rajput. Efficient iterative procedures for approximating fixed points of contractive-type mappings with applications. Numerical Algorithms, 99(1):27–74, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01869-x.

Sun:1994:FEM

[SS94] Guang Fu Sun and Martin Stynes. Finite element methods on piecewise equidistant meshes for interior turning point problems. Numerical Algorithms, 8(1):111–129, ???? 1994. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sidi:1998:UBC

[SS98] Avram Sidi and Yair Shapira. Upper bounds for convergence rates of acceleration methods with initial iterations. Numerical Algorithms, 18(2):113–132, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/14/1/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
14/1/fulltext.pdf.

Sadkane:1999:IVB

[SS99] Miloud Sadkane and Roger B. Sidje. Implementation of a variable block Davidson method with deflation for solving large sparse eigenproblems. Numerical Algorithms, 20(2–3):217–240, June 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/18/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
18/3/fulltext.pdf.

Saylor:2001:AWG

[SS01a] Paul E. Saylor and Dennis C. Smolarski. Addendum to: Why Gaussian quadrature in the complex plane? Numerical Algorithms, 27(2):215–217, June 2001. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp007.lwwonline.com/content/
getfile/5058/35/3/abstract.htm;   http://ipsapp007.
lwwonline.com/content/getfile/5058/35/3/fulltext.pdf.

Saylor:2001:WGQ

[SS01b] Paul E. Saylor and Dennis C. Smolarski. Why Gaussian quadrature in the complex plane? Numerical Algorithms, 26(3):251–280, March 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
2/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/2/4/fulltext.pdf.

Sagona:2003:PEE

[SS03] Manuela Sagona and Alessandra Seghini. An a posteriori error estimate for a semi-Lagrangian scheme for Hamilton–Jacobi equations. Numerical Algorithms, 33(1–4):453–460, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/46/37/abstract.htm;   http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
37/fulltext.pdf.

Shi:2006:MGM

[SS06] Zhen-Jun Shi and Jie Shen. On memory gradient method with trust region for unconstrained optimization. Numerical Algorithms, 41(2):173–196, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=2&spage=173.

Slevinsky:2008:NTT

[SS08] Mikael Slevinsky and Hassan Safouhi. Numerical treatment of a twisted tail using extrapolation methods. Numerical Algorithms, 48(4):301–316, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=301.

Sharma:2010:NFM

[SS10] Janak Raj Sharma and Rajni Sharma. A new family of modified Ostrowski’s methods with accelerated eighth order convergence. Numerical Algorithms, 54(4):445–458, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=54&issue=4&
spage=445.

Sargolzaei:2011:AFO

[SS11a]  Parviz Sargolzaei and Fazlollah Soleymani.  Accurate

fourteenth-order methods for solving nonlinear equations. Numerical Algorithms, 58(4):513–527, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=513.

Sidje:2011:RAF

[SS11b] Roger B. Sidje and Yousef Saad. Rational approximation to the Fermi–Dirac function with applications in density functional theory. Numerical Algorithms, 56(3):455–479, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.

asp?genre=article&issn=1017-1398&volume=56&issue=3&
spage=455.

Sharma:2012:MCH

[SS12a] Janak Raj Sharma and Rajni Sharma. Modified Chebyshev– Halley type method and its variants for computing multiple roots. Numerical Algorithms, 61(4):567–578, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9551-4/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=4&spage=567-578.

Slevinsky:2012:CSN

[SS12b] Richard M. Slevinsky and Hassan Safouhi. A comparative study of numerical steepest descent, extrapolation, and sequence transformation methods in computing semi-infinite integrals. Numerical Algorithms, 60(2):315–337, June 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=2&
spage=315.

Slevinsky:2014:UPC

[SS14] Richard M. Slevinsky and Hassan Safouhi. Useful properties of the coefficients of the Slevinsky–Safouhi formula for differentiation. Numerical Algorithms, 66(3):457–477, July 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9743-6.

Shokri:2015:HPL

[SS15] Ali Shokri and Hosein Saadat. High phase-lag order trigonometrically fitted two-step Obrechkoff methods for the numerical solution of periodic initial value problems. Numerical Algorithms, 68(2):337–354, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9847-7.

Soheili:2016:IMN

[SS16] Ali R. Soheili and Fazlollah Soleymani. Iterative methods for nonlinear systems associated with finite difference approach in stochastic differential equations. Numerical Algorithms, 71(1):

89–102, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9986-5.

Siahkolaei:2021:PST

[SS21] Tahereh Salimi Siahkolaei and Davod Khojasteh Salkuyeh. On the parameter selection in the transformed matrix iteration method. Numerical Algorithms, 86(1):179–189, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00884-y.

Sidki:2023:RQO

[SS23a] S. Sidki and R. Sadaka. The recursive quasi-orthogonal polynomial algorithm. Numerical Algorithms, 92(1):945–971, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01369-w.

Slevinsky:2023:RAE

[SS23b] Richard M. Slevinsky and Hassan Safouhi. A recursive algorithm for an efficient and accurate computation of incomplete Bessel functions. Numerical Algorithms, 92(1):973–983, January 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01438-0.

Shakya:2024:AST

[SS24a] Pratibha Shakya and Rajen Kumar Sinha. Analysis for the space-time a posteriori error estimates for mixed finite element solutions of parabolic optimal control problems. Numerical Algorithms, 96(2):879–924, June 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01669-9.

Singh:2024:ECT

[SS24b] Harmandeep Singh and Janak Raj Sharma. An efficient class of Traub–Steffensen-type optimal order multiple root solvers. Numerical Algorithms, 96(4):1727–1754, August 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01683-x.

Singh:2024:FTS

[SS24c] Harmandeep Singh and Janak Raj Sharma. A fractional Traub–Steffensen-type method for solving nonlinear equations. Numerical Algorithms, 95(3):1103–1126, March 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01601-1.

Sun:2024:SCS

[SS24d] Tao Sun and Hai-Wei Sun. A stiff-cut splitting technique for stiff semi-linear systems of differential equations. Numerical Algorithms, 95(3):1387–1412, March 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01613-x.

Sahoo:2026:CFD

[SS26] Nirupam Sahoo and Randhir Singh. Compact finite difference schemes and error estimation for third-order Emden–Fowler equations. Numerical Algorithms, 101(2):847–882, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02025-9.

Sreelakshmi:2025:AFP

[SSA25] A Sreelakshmi, V P Shyaman, and Ashish Awasthi. An adaptive finite point scheme for the two-dimensional coupled Burgers’ equation. Numerical Algorithms, 99(4):1855–1870, August 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01936-3.

Solary:2025:SCI

[SSC25] Maryam Shams Solary and Stefano Serra-Capizzano. Spectral characterizations and integer powers of tridiagonal 2-Toeplitz matrices. Numerical Algorithms, 100(2):707–727, October 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01863-3.

Shi:2019:SIA

[SSH^{+}19a] Yanyan Shi, Yajuan Sun, Yang He, Hong Qin, and Jian Liu. Symplectic integrators with adaptive time step applied to runaway electron dynamics. Numerical Algorithms, 81(4):1295– 1309, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Suzuki:2019:ECR

[SSH19b] Tomohiro Suzuki, Hiroshi Sugiura, and Takemitsu Hasegawa. Estimating convergence regions of Schröder’s iteration formula: how the Julia set shrinks to the Voronoi boundary. Numerical Algorithms, 82(1):183–199, September 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Suzuki:2020:RCD

[SSH20] Tomohiro Suzuki, Hiroshi Sugiura, and Takemitsu Hasegawa. Regions of convergence and dynamics of Schröder-like iteration formulae as applied to complex polynomial equations with multiple roots. Numerical Algorithms, 85(1):133–144, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00806-7.

Singh:2023:SYE

[SSK23] Harmandeep Singh, Janak Raj Sharma, and Sunil Kumar. A simple yet efficient two-step fifth-order weighted-Newton method for nonlinear models. Numerical Algorithms, 93 (1):203–225, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01412-w.

Shi:2024:SCE

[SSL24] Hongling Shi, Minghui Song, and Mingzhu Liu. Strong convergence of explicit numerical schemes for stochastic differential equations with piecewise continuous arguments. Numerical Algorithms, 97(2):779–800, October 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01723-6.

Singer:2012:NMP

[SSN^{+}12] Sanja Singer, Sasa Singer, Vedran Novaković, Aleksandar Usćumlić, and Vedran Dunjko. Novel modifications of parallel Jacobi algorithms. Numerical Algorithms, 59(1):1–27, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=1&
spage=1.

Smoktunowicz:2015:TTR

[SSP15] Alicja Smoktunowicz, Agata Smoktunowicz, and Ewa Pawelec. The three-term recursion for Chebyshev polynomials is mixed forward-backward stable. Numerical Algorithms, 69(4): 785–794, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-014-9925-x.pdf.

Sahoo:2025:IFO

[SSR25] Nirupam Sahoo, Randhir Singh, and Higinio Ramos. An innovative fourth-order numerical scheme with error analysis for Lane–Emden–Fowler type systems. Numerical Algorithms, 99(1):411–439, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01882-0.

Sahu:2014:NLM

[SSS14] D. R. Sahu, Krishna Kumar Singh, and Vipin Kumar Singh. A Newton-like method for generalized operator equations in Banach spaces. Numerical Algorithms, 67(2):289–303, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9791-y.

Santos:2021:NTM

[SSS21] P. S. M. Santos, G. N. Silva, and R. C. M. Silva. Newtontype method for solving generalized inclusion. Numerical Algorithms, 88(4):1811–1829, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01096-
8.

Sousa:2022:QNT

[SSSS22] Leonardo A. Sousa, Susana Scheimberg, Pedro Jorge S. Santos, and Paulo Sérgio M. Santos. A quasi-Newton type method for equilibrium problems. Numerical Algorithms, 89(3):1129–1143, March 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01148-z.

Sakurai:1992:NFP

[SST92] Tetsuya Sakurai, Hiroshi Sugiura, and Tatsuo Torii. Numerical factorization of a polynomial by rational Hermite interpolation. Numerical Algorithms, 3(1–4):411–418, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Setzer:2008:RIR

[SST08] S. Setzer, G. Steidl, and T. Teuber. Restoration of images with rotated shapes. Numerical Algorithms, 48(1–3):49–66, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=48&issue=1&
spage=49.

Shen:2020:DTG

[SSYL20] Ruigang Shen, Shi Shu, Ying Yang, and Benzhuo Lu. A decoupling two-grid method for the time-dependent Poisson– Nernst–Planck equations. Numerical Algorithms, 83(4):1613– 1651, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Skeel:1992:LPE

[ST92] Robert D. Skeel and Hon-Wah Tam. Limits of parallelism in explicit ODE methods. Numerical Algorithms, 2(3–4):337–349, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Stynes:1998:FDA

[ST98] Martin Stynes and Lutz Tobiska. A finite difference analysis of a streamline diffusion method on a Shishkin mesh. Numerical Algorithms, 18(3–4):337–360, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/15/7/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
15/7/fulltext.pdf.

Sirotkin:1999:PTL

[ST99] V. Sirotkin and P. Tarvainen. Parallel two-level Schwarz methods for singularly perturbed elliptic problems. Numerical Algorithms, 22(2):129–156, December 1999. CODEN NUALEG.

ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
22/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/22/1/fulltext.pdf.

Stathopoulos:2002:RMG

[ST02] Andreas Stathopoulos and Shang-Hua Teng. Recovering mesh geometry from a stiffness matrix. Numerical Algorithms, 30(3–4):303–322, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
42/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/42/4/fulltext.pdf.

Senosiain:2017:TSS

[ST17] M. J. Senosiain and A. Tocino. Two-step strong order 1.5 schemes for stochastic differential equations. Numerical Algorithms, 75(4):973–1003, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Scott:2018:SCA

[ST18] Jennifer Scott and Miroslav Tuma. A Schur complement approach to preconditioning sparse linear least-squares problems with some dense rows. Numerical Algorithms, 79(4):1147– 1168, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-018-0478-2.pdf.

Simos:2021:EDR

[ST21] T. E. Simos and Ch. Tsitouras. Evolutionary derivation of Runge–Kutta pairs for addressing inhomogeneous linear problems. Numerical Algorithms, 87(2):511–525, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00976-9.

Scott:2022:NSA

[ST22] Jennifer Scott and Miroslav Tuma. A null-space approach for large-scale symmetric saddle point systems with a small and non zero (2, 2) block. Numerical Algorithms, 90(4):1639–1667, August 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01245-z.

Senosiain:2023:SMS

[ST23] M. J. Senosiain and A. Tocino. A survey of mean-square destabilization of multidimensional linear stochastic differential systems with non-normal drift. Numerical Algorithms, 93 (4):1543–1559, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01478-6.

Scott:2026:DRI

[ST26] Jennifer Scott and Miroslav Tuma. Developing robust incomplete Cholesky factorizations in half precision arithmetic. Numerical Algorithms, 101(1):543–564, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02015-x.

Steidl:1995:MAF

[Ste95] Gabriele Steidl. On multivariate attenuation factors. Numerical Algorithms, 9(3–4):245–261, ???? 1995. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Steihaug:2020:CSE

[Ste20] Trond Steihaug. Computational science in the eighteenth century. Test cases for the methods of Newton, Raphson, and Halley: 1685 to 1745. Numerical Algorithms, 83(4):1259–1275, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Stepien:2024:ASS

[Ste24] Łukasz Stepień. Adaptive step-size control for global approximation of SDEs driven by countably dimensional Wiener process. Numerical Algorithms, 96(4):1699–1725, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01682-y.

Stillfjord:2018:AHO

[Sti18] Tony Stillfjord. Adaptive high-order splitting schemes for large-scale differential Riccati equations. Numerical Algorithms, 78(4):1129–1151, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-017-
0416-8.pdf.

Stolovitch:1993:CVF

[Sto93] Laurent Stolovitch. On the computation of a versal family of matrices. Numerical Algorithms, 4(1–2):25–46, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Strom:1993:PP

[Str93] Kyrre Strøm. Products of B-patches. Numerical Algorithms, 4(4):323–337, May 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Strauss:1997:IPP

[Str97] Hans Strauss. On interpolation with products of positive definite functions. Numerical Algorithms, 15(2):153–165, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/7/2/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/7/
2/fulltext.pdf.

Strauss:2002:ACC

[Str02] Hans Strauss. Approximation with constraints by conditionally positive definite functions. Numerical Algorithms, 30(2):185–198, June 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
41/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/41/5/fulltext.pdf.

Struckmeier:2005:DSS

[Str05] Jens Struckmeier. A deterministic scheme for Smoluchowski’s coagulation equation based on binary grid refinement. Numerical Algorithms, 40(3):233–249, November 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=40&issue=3&spage=233.

Strakos:2009:MRU

[Str09] Zdenek Strakos. Model reduction using the Vorobyev moment problem. Numerical Algorithms, 51(3):363–379, July 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=51&issue=3&

spage=363. Tributes to Gene H. Golub, Part III.

Stuart:1997:PDC

[Stu97] A. M. Stuart. Probabilistic and deterministic convergence proofs for software for initial value problems. Numerical Algorithms, 14(1–3):227–260, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
5/14/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/5/14/fulltext.pdf. Dynami-

cal numerical analysis (Atlanta, GA, 1995).

Springer:2014:CEA

[SU14] Theresa Springer and Karsten Urban. Comparison of the EM algorithm and alternatives. Numerical Algorithms, 67(2): 335–364, October 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9794-8.

Sun:1994:CHI

[Sun94a] Xingping Sun. Cardinal Hermite interpolation using positive definite functions. Numerical Algorithms, 7(2–4):253–268, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sun:1994:FTC

[Sun94b] Xingping Sun. The fundamentality of translates of a continuous function on spheres. Numerical Algorithms, 8(1):131–134, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sutton:2009:CCC

[Sut09] Brian D. Sutton. Computing the complete CS decomposition. Numerical Algorithms, 50(1):33–65, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=50&issue=1&spage=33.

Sutton:2017:VEM

[Sut17] Oliver J. Sutton. The virtual element method in 50 lines of MATLAB. Numerical Algorithms, 75(4):1141–1159, August 2017. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11075-016-0235-3.pdf.

Sleijpen:1995:MCP

[Sv95] Gèrard L. G. Sleijpen and Henk A. van der Vorst. Maintaining convergence properties of BiCGstab methods in finite precision arithmetic. Numerical Algorithms, 10(3–4):203–223, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Serra:2026:HGM

[SV26] Ray Serra and Sandoel Vieira. A hybrid gradient method for vector optimization problems. Numerical Algorithms, 101 (1):373–391, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02009-9.

Sleijpen:1994:OHB

[SvF94] G. L. G. Sleijpen, H. A. van der Vorst, and D. R. Fokkema. bicgstab(l) and other hybrid Bi-CG methods. Numerical Algorithms, 7(1):75–109, June 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sommariva:2001:FSP

[SVZ01] Alvise Sommariva, Marco Vianello, and Renato Zanovello. Fast summation of power series with coefficients analytic at infinity. Numerical Algorithms, 27(1):77–87, May 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.lwwonline.
com/content/getfile/5058/33/4/abstract.htm;   http:
//ipsapp007.lwwonline.com/content/getfile/5058/33/
4/fulltext.pdf.

Sommariva:2005:ABC

[SVZ05] Alvise Sommariva, Marco Vianello, and Renato Zanovello. Adaptive bivariate Chebyshev approximation. Numerical Algorithms, 38(1–3):79–94, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sommariva:2008:NCC

[SVZ08] Alvise Sommariva, Marco Vianello, and Renato Zanovello. Nontensorial Clenshaw–Curtis cubature. Numerical Algorithms, 49(1–4):409–427, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=49&issue=1&spage=409.

Schaback:2000:AGT

[SW00] Robert Schaback and Holger Wendland. Adaptive greedy techniques for approximate solution of large RBF systems. Numerical Algorithms, 24(3):239–254, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Smoktunowicz:2005:IAH

[SW05] Alicja Smoktunowicz and Iwona Wróbel. On improving the accuracy of Horner’s and Goertzel’s algorithms. Numerical Algorithms, 38(4):243–258, April 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Song:2007:IUT

[SW07] Lunji Song and Yujiang Wu. Incremental unknowns in three-dimensional stationary problem. Numerical Algorithms, 46(2):153–171, October 2007. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=2&spage=153.

Schmitt:2010:PSE

[SW10a] Bernhard A. Schmitt and Rüdiger Weiner. Parallel start for explicit parallel two-step peer methods. Numerical Algorithms, 53(2–3):363–381, March 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=2&spage=363.

Shang:2010:LPF

[SW10b] Yueqiang Shang and Kun Wang. Local and parallel finite element algorithms based on two-grid discretizations for the transient Stokes equations. Numerical Algorithms, 54(2):195–218, June 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=54&
issue=2&spage=195.

Shi:2011:MNA

[SW11] Zhenjun Shi and Shengquan Wang. Modified nonmonotone Armijo line search for descent method. Numerical Algorithms, 57(1):1–25, May 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=1&spage=1.

Shao:2014:NSK

[SW14] Wenting Shao and Xionghua Wu. A numerical study for the KdV and the good Boussinesq equations using Fourier Chebyshev tau meshless method. Numerical Algorithms, 67(3):581– 597, November 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9809-5.

Shi:2019:EGT

[SW19] Wei Shi and Xinyuan Wu. Explicit Gautschi-type integrators for nonlinear multi-frequency oscillatory second-order initial value problems. Numerical Algorithms, 81(4):1275–1294, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Singh:2022:FVA

[SW22] Mehakpreet Singh and Gavin Walker. Finite volume approach for fragmentation equation and its mathematical analysis. Numerical Algorithms, 89(2):465–486, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01122-9;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01122-9.pdf.

Shao:2024:NNM

[SW24a] Xin-Hui Shao and Zhe Wang. The nonsmooth Newton’s method for the horizontal nonlinear complementarity problem. Numerical Algorithms, 96(1):75–103, May 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01640-8.

Sun:2024:RPA

[SW24b] Min Sun and Yiju Wang. A relaxed proximal ADMM method for block separable convex programming. Numerical Algorithms, 95(2):575–603, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01582-
1.

Sun:2026:EBL

[SW26] Guixiang Sun and Feng Wang. Error bounds for the linear complementarity problems of N_{1}-matrices and BN_{1}-matrices. Numerical Algorithms, 102(2):1049–1064, June 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02168-9.

Sidje:2008:FGC

[SWB08] Roger B. Sidje, Alan B. Williams, and Kevin Burrage. Fast generalized cross validation using Krylov subspace methods. Numerical Algorithms, 47(2):109–131, February 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=2&spage=109.

Sun:2020:IQI

[SWG20] Zhengjie Sun, Zongmin Wu, and Wenwu Gao. An iterated quasi-interpolation approach for derivative approximation. Numerical Algorithms, 85(1):255–276, September 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00812-9.

Song:2024:ANK

[SWJ24] Bo Song, Jing-Yi Wang, and Yao-Lin Jiang. Analysis of a new Krylov subspace enhanced parareal algorithm for time-periodic problems. Numerical Algorithms, 97(1):289–310, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01704-9.

Shao:2022:SBL

[SWS22] Xin-Hui Shao, Zhe Wang, and Hai-Long Shen. A signbased linear method for horizontal linear complementarity problems. Numerical Algorithms, 91(3):1165–1181, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01296-w.

Sauer:1996:RPL

[SX96] Thomas Sauer and Yuan Xu. Regular points for Lagrange interpolation on the unit disk. Numerical Algorithms, 12(3– 4):287–296, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Szyld:2000:CSA

[SX00] Daniel B. Szyld and Jian-Jun Xu. Convergence of some asynchronous nonlinear multisplitting methods. Numerical Algorithms, 25(1–4):347–361, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/11/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/11/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Song:2020:SVP

[SXHZ20] Xiongfeng Song, Wei Xu, Ken Hayami, and Ning Zheng. Secant variable projection method for solving nonnegative separable least squares problems. Numerical Algorithms, 85(2): 737–761, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00835-2.

Shen:2020:NSS

[SY20] Yunqiu Shen and Tjalling J. Ypma. Numerical solution of separable nonlinear equations with a singular matrix at the solution. Numerical Algorithms, 85(4):1195–1211, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00861-0.

Song:2014:SSO

[SYLT14] J. Song, Q. Yu, F. Liu, and I. Turner. A spatially second-order accurate implicit numerical method for the space and time fractional Bloch–Torrey equation. Numerical Algorithms, 66(4): 911–932, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9768-x.

Suo:2022:WCE

[SYZ22] Yongqiang Suo, Chenggui Yuan, and Shao-Qin Zhang. Weak convergence of Euler scheme for SDEs with low regular drift. Numerical Algorithms, 90(2):731–747, June 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01206-6.

Spedicato:1999:GIL

[SZ99] E. Spedicato and M. Zhu. A generalization of the implicit LU algorithm to an arbitrary initial matrix. Numerical Algorithms, 20(4):343–351, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
19/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/19/4/fulltext.pdf.

Sun:2020:ESO

[SZ20] Yabing Sun and Weidong Zhao. An explicit second-order numerical scheme for mean-field forward backward stochastic differential equations. Numerical Algorithms, 84(1):253–283, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sun:2021:NMM

[SZ21] Yabing Sun and Weidong Zhao. Numerical methods for meanfield stochastic differential equations with jumps. Numerical Algorithms, 88(2):903–937, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-01062-
w.

Shi:2023:UEE

[SZ23] Dongyang Shi and Houchao Zhang. Unconditional error estimates of linearized BDF2–Galerkin FEMs for nonlinear coupled Schrödinger-=Helmholtz equations. Numerical Algorithms, 92(3):1679–1705, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01360-

Sun:2024:MBP

[SZ24] Yabing Sun and Quan Zhou. Maximum bound principle for matrix-valued Allen–Cahn equation and integrating factor Runge–Kutta method. Numerical Algorithms, 97(1):391–429, September 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01708-5.

Szafraniec:1992:EBP

[Sza92] Franciszek Hugon Szafraniec. On extending backwards positive definite sequences. Numerical Algorithms, 3(1–4):419–425, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Szafraniec:2003:DMP

[Sza03] Franciszek Hugon Szafraniec. Detecting mass points of representing measures. Numerical Algorithms, 33(1–4):475–483, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/46/39/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
46/39/fulltext.pdf.

Sun:2023:EOM

[SZQS23] Jingwei Sun, Hong Zhang, Xu Qian, and Songhe Song. Up to eighth-order maximum-principle-preserving methods for the Allen–Cahn equation. Numerical Algorithms, 92(2):1041–1062, February 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01329-4.

Sun:2021:FTS

[SzS21] Hong Sun and Zhi zhong Sun. A fast temporal second-order compact ADI difference scheme for the 2D multi-term fractional wave equation. Numerical Algorithms, 86(2):761–797, February 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00910-z.

Sun:2011:GNI

[SZX11] Zhe Sun, Jinping Zeng, and Hongru Xu. Generalized Newton-iterative method for semismooth equations. Numerical Algorithms, 58(3):333–349, November 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=3&spage=333.

Szyld:2006:MPI

[Szy06] Daniel B. Szyld. The many proofs of an identity on the norm of oblique projections. Numerical Algorithms, 42(3–4):309–323, July 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=42&
issue=3&spage=309.

Thomas-Agnan:1996:CFR

[TA96] Christine Thomas-Agnan. Computing a family of reproducing kernels for statistical applications. Numerical Algorithms, 13 (1–2):21–32, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Toutounian:2013:RSP

[TA13] Faezeh Toutounian and Nasser Akhoundi. Recursive self preconditioning method based on Schur complement for Toeplitz matrices. Numerical Algorithms, 62(3):505–525, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9603-9/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=505-525.

Taghipour:2024:NSD

[TA24] M. Taghipour and H. Aminikhah. Numerical solution of distributed-order fractional Korteweg–de Vries equation via fractional Zigzag rising diagonal functions. Numerical Algorithms, 96(2):739–776, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01664-
0.

Tan:2026:BPC

[TA26] Jianguo Tan and Yaqi An. Boundary preservation and convergence of the linearly implicit Euler method with truncated Wiener process for the Wright–Fisher model. Numerical Algorithms, 101(4):2485–2503, April 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02096-8.

Tachev:2012:NEV

[Tac12] Gancho Tachev. New estimates in Voronovskaja’s theorem. Numerical Algorithms, 59(1):119–129, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=1&spage=119.

Tai:1992:GEP

[Tai92] Xue Cheng Tai. Global extrapolation with a parallel splitting method. Numerical Algorithms, 3(1–4):427–440, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Takahashi:2017:SCN

[Tak17] Wataru Takahashi. The split common null point problem for generalized resolvents in two Banach spaces. Numerical Algorithms, 75(4):1065–1078, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tamberg:2010:TES

[Tam10] Gert Tamberg. On truncation errors of some generalized Shannon sampling operators. Numerical Algorithms, 55(2– 3):367–382, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=55&issue=2&spage=367.

Taiwo:2021:HTI

[TAM21] A. Taiwo, T. O. Alakoya, and O. T. Mewomo. Halperntype iterative process for solving split common fixed point and monotone variational inclusion problem between Banach spaces. Numerical Algorithms, 86(4):1359–1389, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00937-2.

Tan:2017:SSL

[Tan17] Xue-Yuan Tan. Shifted SSOR-like preconditioner for non-Hermitian positive definite matrices. Numerical Algorithms, 75 (1):245–260, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tanaka:2020:GPS

[Tan20] Ken’ichiro Tanaka. Generation of point sets by convex optimization for interpolation in reproducing kernel Hilbert spaces. Numerical Algorithms, 84(3):1049–1079, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00792-w.

Tasche:1993:FAD

[Tas93] Manfred Tasche. Fast algorithms for discrete Chebyshev– Vandermonde transforms and applications. Numerical Algorithms, 5(1–4):453–464, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Tang:2019:NSI

[TB19] Yan Tang and Zhiqing Bao. New semi-implicit midpoint rule for zero of monotone mappings in Banach spaces. Numerical Algorithms, 81(3):853–878, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/content/pdf/10.1007/s11075-018-
0574-3.pdf.

Themistoclakis:2025:NGF

[TB25] Woula Themistoclakis and Marc Van Barel. A note on generalized Floater–Hormann interpolation at arbitrary distributions of nodes. Numerical Algorithms, 100(2):729–744, October 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01933-6.

Tokarzewski:1994:TPP

[TBA94] Stanisław Tokarzewski, Jerzy Bławzdziewicz, and Igor Andrianov. Two-point Padé approximants for formal Stieltjes series. Numerical Algorithms, 8(2–4):313–328, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tavares:2023:PSD

[TBC^{+}23] S. Tavares, C. P. Brás, A. L. Custódio, V. Duarte, and P. Medeiros. Parallel strategies for direct multisearch. Numerical Algorithms, 92(3):1757–1788, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01364-1.

Tharwat:2013:NCE

[TBY13] M. M. Tharwat, A. H. Bhrawy, and Ahmet Yildirim. Numerical computation of eigenvalues of discontinuous Sturm– Liouville problems with parameter dependent boundary conditions using sinc method. Numerical Algorithms, 63(1):27– 48, May 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9609-3.

Talbot:2005:APS

[TC05] C. J. Talbot and A. Crampton. Application of the pseudospectral method to 2D eigenvalue problems in elasticity. Numerical Algorithms, 38(1–3):95–110, March 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tan:2026:EAS

[TCO26] Bing Tan, Jiawei Chen, and Xiaoqing Ou. Extragradient algorithms for solving variational inequalities on Hadamard manifolds. Numerical Algorithms, 101(3):1915–1948, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02068-y.

Tapley:2019:NAR

[TCOA19] Benjamin Tapley, Elena Celledoni, Brynjulf Owren, and Helge I. Andersson. A novel approach to rigid spheroid models in viscous flows using operator splitting methods. Numerical Algorithms, 81(4):1423–1441, August 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tsai:2014:TDR

[TCW14] Angela Y. J. Tsai, Robert P. K. Chan, and Shixiao Wang. Two-derivative Runge–Kutta methods for PDEs using a novel discretization approach. Numerical Algorithms, 65(3):687–703, March 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9823-2.

Tunga:2009:CMB

[TD09] Burcu Tunga and Metin Demiralp. Constancy maximization based weight optimization in high dimensional model representation. Numerical Algorithms, 52(3):435–459, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=3&
spage=435.

Thong:2021:NSC

[TDC21] Duong Viet Thong, Vu Tien Dung, and Yeol Je Cho. A new strong convergence for solving split variational inclusion problems. Numerical Algorithms, 86(2):565–591, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00901-0.

Taneja:2024:NNA

[TDKB24] Komal Taneja, Komal Deswal, Devendra Kumar, and Dumitru Baleanu. Novel numerical approach for time fractional equations with nonlocal condition. Numerical Algorithms, 95(3): 1413–1433, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01614-w.

Tian:2024:PAI

[TDWL24] Zhaolu Tian, Xuefeng Duan, Nian-Ci Wu, and Zhongyun Liu. The parameterized accelerated iteration method for solving the matrix equation AXB = C. Numerical Algorithms, 97 (2):843–867, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01726-3.

Talibi:2003:NST

[TE03] M. El Alaoui Talibi and A. El Kacimi. Numerical simulation of the two-hydrodynamic film thickness. Numerical Algorithms, 33(1–4):241–250, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/

20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/20/fulltext.pdf.

Temme:1997:NAU

[Tem97] N. M. Temme. Numerical algorithms for uniform Airytype asymptotic expansions. Numerical Algorithms, 15 (2):207–225, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
7/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/7/5/fulltext.pdf.

Temsah:2008:SMS

[Tem08] R. S. Temsah. Spectral methods for some singularly perturbed third order ordinary differential equations. Numerical Algorithms, 47(1):63–80, January 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=1&spage=63.

Terekhov:2022:GLE

[Ter22] Andrew V. Terekhov. Generating the Laguerre expansion coefficients by solving a one-dimensional transport equation. Numerical Algorithms, 89(1):303–322, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01115-8.

Terekhov:2023:ECM

[Ter23] Andrew V. Terekhov. An extra-component method for evaluating fast matrix-vector multiplication with special functions. Numerical Algorithms, 92(4):2189–2217, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01383-y.

Tan:2000:NTR

[TF00] Jieqing Tan and Yi Fang. Newton–Thiele’s rational interpolants. Numerical Algorithms, 24(1–2):141–157, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/27/9/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/

27/9/fulltext.pdf. Computational methods from rational approximation theory (Wilrijk, 1999).

Tselepidis:2019:DAT

[TFPG19] N. A. Tselepidis, C. K. Filelis-Papadopoulos, and G. A. Gravvanis. Distributed algebraic tearing and interconnecting techniques. Numerical Algorithms, 82(3):809–842, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Fan:2016:RPS

[tFZyZ16] Hong tao Fan, Bing Zheng, and Xin yun Zhu. A relaxed positive semi-definite and skew-Hermitian splitting preconditioner for non-Hermitian generalized saddle point problems. Numerical Algorithms, 72(3):813–834, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0068-5.

Tang:2020:NSA

[TG20] Yan Tang and Aviv Gibali. New self-adaptive step size algorithms for solving split variational inclusion problems and its applications. Numerical Algorithms, 83(1):305–331, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Thong:2018:MSE

[TH18a] Duong Viet Thong and Dang Van Hieu. Modified subgradient extragradient method for variational inequality problems. Numerical Algorithms, 79(2):597–610, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Thong:2018:WSC

[TH18b] Duong Viet Thong and Dang Van Hieu. Weak and strong convergence theorems for variational inequality problems. Numerical Algorithms, 78(4):1045–1060, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Thong:2019:ISE

[TH19a] Duong Viet Thong and Dang Van Hieu. Inertial subgradient extragradient algorithms with line-search process for solving variational inequality problems and fixed point problems. Numerical Algorithms, 80(4):1283–1307, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Thong:2019:SEV

[TH19b] Duong Viet Thong and Dang Van Hieu. Some extragradientviscosity algorithms for solving variational inequality problems and fixed point problems. Numerical Algorithms, 82(3):761– 789, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tang:2023:MSP

[TH23] Shi-Ping Tang and Yu-Mei Huang. A matrix splitting preconditioning method for solving the discretized tempered fractional diffusion equations. Numerical Algorithms, 92(2):1311–1333, February 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01341-8.

OC:1997:PSS

[The97] The Organizing Committee, editor. Preface: Second Symposium on New Techniques for the Treatment of Sparse Matrices for Industrial Problems, volume 16(1) of Numerical Algorithms. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, December 1997. CODEN NUA-LEG. ISBN ???? ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). LCCN ???? URL http://ipsapp007.kluweronline.
com/content/getfile/5058/9/1/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/9/
1/fulltext.pdf.

Themistoclakis:2012:UAD

[The12] Woula Themistoclakis. Uniform approximation on [−1, 1] via discrete de la Vallée Poussin means. Numerical Algorithms, 60(4):593–612, August 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=4&spage=593.

Tan:2021:NMS

[THF21] Zengzhen Tan, Rong Hu, and Yaping Fang. A new method for solving split equality problems via projection dynamical systems. Numerical Algorithms, 86(4):1705–1719, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00950-5.

Thiry:1993:ESB

[Thi93] S. Thiry. Extremal signatures for bivariate Chebyshev approximation problems. Numerical Algorithms, 5(1–4):215–225, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Thron:1992:LPS

[Thr92] W. J. Thron. Limit periodic Schur algorithms, the case |γ| = 1, ∑ d_{n} < ∞. Numerical Algorithms, 3(1–4):441–450, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Thuy:2020:CHV

[THS20] Nguyen Thi Thu Thuy, Pham Thanh Hieu, and J. J. Strodiot. Convergence of a hybrid viscosity approximation method for finding zeros of m-accretive operators. Numerical Algorithms, 83(4):1591–1612, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tuyen:2019:SPM

[THT19] Truong Minh Tuyen, Nguyen Song Ha, and Nguyen Thi Thu Thuy. A shrinking projection method for solving the split common null point problem in Banach spaces. Numerical Algorithms, 81(3):813–832, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tian:2021:BPI

[Tia21] Yan Tian. Barycentric prolate interpolation and pseudospectral differentiation. Numerical Algorithms, 88(2):793–811, October 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01057-7;   http://link.
springer.com/content/pdf/10.1007/s11075-020-01057-
7.pdf.

Tirani:2002:PAE

[Tir02] R. Tirani. A parallel algorithm for the estimation of the global error in Runge–Kutta methods. Numerical Algorithms, 31 (1–4):311–318, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
17/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/17/fulltext.pdf.

Turcajova:1994:SPF

[TK94] Radka Turcajová and Jaroslav Kautsky. Shift products and factorizations of wavelet matrices. Numerical Algorithms, 8 (1):27–45, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Toprakseven:2024:WGF

[TKS24] Suayip Toprakseven, Aditya Kaushik, and Manju Sharma. A weak Galerkin finite element method for singularly perturbed problems with two small parameters on Bakhvalovtype meshes. Numerical Algorithms, 97(2):727–751, October 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01721-8.

Taddele:2023:SAA

[TKSG23] Guash Haile Taddele, Poom Kumam, Pongsakorn Sunthrayuth, and Anteneh Getachew Gebrie. Self-adaptive algorithms for solving split feasibility problem with multiple output sets. Numerical Algorithms, 92(2):1335–1366, February 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01343-6.

Tondji:2023:FRB

[TL23] Lionel Tondji and Dirk A. Lorenz. Faster randomized block sparse Kaczmarz by averaging. Numerical Algorithms, 93(4): 1417–1451, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01473-x.

Tran:2024:MIC

[TL24] Qiqi Tran and Jinjie Liu. Modified iterated Crank–Nicolson method with improved accuracy for advection equations. Numerical Algorithms, 95(4):1539–1560, April 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01619-5.

Tian:2022:CIA

[TLD22] Zhaolu Tian, Zhongyun Liu, and Yinghui Dong. The coupled iteration algorithms for computing PageRank. Numerical Algorithms, 89(4):1603–1637, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-021-01166-
x.

Thong:2023:SLC

[TLD^{+}23] Duong Viet Thong, Xiaoxiao Li, Qiao-Li Dong, Vu Tien Dung, and Nguyen Phuong Lan. Strong and linear convergence of projection-type method with an inertial term for finding minimum-norm solutions of pseudomonotone variational inequalities in Hilbert spaces. Numerical Algorithms, 92(4):2243– 2274, April 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01386-9.

Thong:2025:NAK

[TLR^{+}25] Duong Viet Thong, Xiao-Huan Li, Simeon Reich, Qiao-Li Dong, and Dang Huy Ngan. A new approach to the Korpelevich method for solving pseudomonotone equilibrium problems. Numerical Algorithms, 98(2):719–741, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01813-z.

Tang:2005:CES

[TM05] Yuehong Tang and Günter W. Mühlbach. Cardinal ECTsplines. Numerical Algorithms, 38(4):259–283, April 2005. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Triantafyllou:2010:RNS

[TM10] Dimitrios Triantafyllou and Marilena Mitrouli. On rank and null space computation of the generalized Sylvester matrix. Numerical Algorithms, 54(3):297–324, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=297.

Tebbens:2014:PBE

[TM14] Jurjen Duintjer Tebbens and Gérard Meurant. Prescribing the behavior of early terminating GMRES and Arnoldi iterations. Numerical Algorithms, 65(1):69–90, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9695-x.

Tebbens:2020:RNR

[TM20] Jurjen Duintjer Tebbens and Gérard Meurant. On the residual norms, the Ritz values and the harmonic Ritz values that can be generated by restarted GMRES. Numerical Algorithms, 84(4):1329–1352, August 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00846-z.

Tichy:2025:BCA

[TMS25] Petr Tichý, Gérard Meurant, and Dorota Simonová. Block CG algorithms revisited. Numerical Algorithms, 100(4):1527–1553, December 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02038-4.

Tuan:2010:SSH

[TN10] Nguyen Minh Tuan and Lan T. T. Nguyen. On solutions of a system of hereditary and self-referred partial-differential equations. Numerical Algorithms, 55(1):101–113, September 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=55&issue=1&
spage=101.

Turaci:2017:DTD

[TÖ17] Mukaddes Ökten Turaci and Turgut Özis. Derivation of threederivative Runge–Kutta methods. Numerical Algorithms, 74 (1):247–265, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0147-2.

Tanaka:2021:ELN

[TO21] Mirai Tanaka and Takayuki Okuno. Extension of the LP– Newton method to conic programming problems via semiinfinite representation. Numerical Algorithms, 86(3):1285– 1302, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00933-6.

Tommasini:1992:NAS

[Tom92] T. Tommasini. A new algorithm for special Vandermonde systems. Numerical Algorithms, 2(3–4):299–306, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tomasiello:2011:NSD

[Tom11] Stefania Tomasiello. Numerical stability of DQ solutions of wave problems. Numerical Algorithms, 57(3):289–312, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=3&
spage=289.

Tongnoi:2025:FBF

[Ton25] Buris Tongnoi. The forward–backward–forward algorithm with extrapolation from the past and penalty scheme for solving monotone inclusion problems and applications. Numerical Algorithms, 98(4):2113–2143, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01866-
0.

Toutounian:1998:SOS

[Tou98] Faezeh Toutounian. The stable A^{T}A-orthogonal s-step Orthomin(k) algorithm with the CADNA library. Numerical Algorithms, 17(1–2):105–119, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/31/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/31/10/fulltext.pdf.

Tovbis:1997:HCN

[Tov97] Alexander Tovbis. Homoclinic connections and numerical integration. Numerical Algorithms, 14(1–3):261–267, March 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/5/15/abstract.htm;   http:

//ipsapp007.kluweronline.com/content/getfile/5058/

5/15/fulltext.pdf. Dynamical numerical analysis (Atlanta, GA, 1995). See erratum [Tov98].

Tovbis:1998:EHC

[Tov98] Alexander Tovbis. Erratum: “Homoclinic connections and numerical integration” [Numer. Algorithms 14 (1997), no. 1, 261–267; MR 98c:58153]. Numerical Algorithms, 17 (3–4):355, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
12/9/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/12/9/fulltext.pdf. See [Tov97].

Tang:2026:IHS

[TOY26] Jiahuan Tang, Chen Ouyang, and Gonglin Yuan. An inertial Hestenes–Stiefel algorithm for nonsmooth convex optimization and its applications in machine learning. Numerical Algorithms, 101(2):1241–1271, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02040-w.

Thery:2022:ASW

[TPLB22] Sophie Thery, Charles Pelletier, Florian Lemarié, and Eric Blayo. Analysis of Schwarz waveform relaxation for the coupled Ekman boundary layer problem with continuously variable coefficients. Numerical Algorithms, 89(3):1145–1181, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01149-y.

Tang:2014:CSI

[TPY14] Yuchao Tang, Jigen Peng, and Shigang Yue. Cyclic and simultaneous iterative methods to matrix equations of the form A_{i}XB_{i} = F_{i}. Numerical Algorithms, 66(2):379–397, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9740-9.

Tan:2022:RSE

[TQC22] Bing Tan, Xiaolong Qin, and Sun Young Cho. Revisiting subgradient extragradient methods for solving variational inequalities. Numerical Algorithms, 90(4):1593–1615, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01243-1.

Tan:2024:AIA

[TQW24] Bing Tan, Xiaolong Qin, and Xianfu Wang. Alternated inertial algorithms for split feasibility problems. Numerical Algorithms, 95(2):773–812, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01589-8.

Tan:2021:TMI

[TQY21] Bing Tan, Xiaolong Qin, and Jen-Chih Yao. Two modified inertial projection algorithms for bilevel pseudomonotone variational inequalities with applications to optimal control problems. Numerical Algorithms, 88(4):1757–1786, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01093-x.

Traversoni:1993:ANS

[Tra93] Leonardo Traversoni. An algorithm for natural spline interpolation. Numerical Algorithms, 5(1–4):63–70, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Thong:2024:IMS

[TRCL24] Duong Viet Thong, Simeon Reich, Prasit Cholamjiak, and Le Dinh Long. Iterative methods for solving monotone variational inclusions without prior knowledge of the Lipschitz constant of the single-valued operator. Numerical Algorithms, 97(3):1267–1300, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01749-4.

Tijskens:2002:ADS

[TRRD02] E. Tijskens, D. Roose, H. Ramon, and J. De Baerdemaeker. Automatic differentiation for solving nonlinear partial differential equations: An efficient operator overloading approach. Numerical Algorithms, 30(3–4):259–301, August 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/42/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
42/3/fulltext.pdf.

Thong:2023:NPM

[TRSI23] Duong Viet Thong, Simeon Reich, Yekini Shehu, and Olaniyi S. Iyiola. Novel projection methods for solving variational inequality problems and applications. Numerical Algorithms, 93(3):1105–1135, July 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01457-
x.

Trummer:2024:SKO

[Tru24] Manfred R. Trummer. The Szegő kernel and oblique projections: conformal mapping of non-smooth regions. Numerical Algorithms, 95(2):929–942, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01594-
x.

Tanabe:1992:PSR

[TS92] Kunio Tanabe and Masahiko Sagae. Pivoting strategy for rankone modification of LDM^{t}-like factorization. Numerical Algorithms, 2(2):137–153, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tocino:2015:TSM

[TS15] A. Tocino and M. J. Senosiain. Two-step Milstein schemes for stochastic differential equations. Numerical Algorithms, 69 (3):643–665, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9918-9.

Tabrizidooz:2018:BPB

[TS18a] Hamid Reza Tabrizidooz and Khadigeh Shabanpanah. Bernstein polynomial basis for numerical solution of boundary value problems. Numerical Algorithms, 77(1):211–228, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tomovic:2018:COS

[TS18b] Tatjana V. Tomović and Marija P. Stanić. Construction of the optimal set of two or three quadrature rules in the sense of Borges. Numerical Algorithms, 78(4):1087–1109, August 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tsitouras:2007:RKI

[Tsi07] Ch. Tsitouras. Runge–Kutta interpolants for high precision computations. Numerical Algorithms, 44(3):291–307, March 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=44&issue=3&
spage=291.

Thong:2020:WSC

[TSI20] Duong Viet Thong, Yekini Shehu, and Olaniyi S. Iyiola. Weak and strong convergence theorems for solving pseudomonotone variational inequalities with non-Lipschitz mappings. Numerical Algorithms, 84(2):795–823, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00780-0.

Tsukerman:2002:FED

[Tsu02] Igor Tsukerman. Finite element differential–algebraic systems for eddy current problems. Numerical Algorithms, 31 (1–4):319–335, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
13/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/13/fulltext.pdf.

Taussky:2006:CTT

[TT06] Olga Taussky and John Todd. Cholesky, Toeplitz and the triangular factorization of symmetric matrices. Numerical Algorithms, 41(2):197–202, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=2&spage=197.

Taherian:2021:BGC

[TT21] A. Taherian and F. Toutounian. Block GPBi–CG method for solving nonsymmetric linear systems with multiple right-hand sides and its convergence analysis. Numerical Algorithms, 88 (4):1831–1850, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01097-7.

Tu:2026:MSM

[TTA26] Ho Phi Tu, Do Duy Thanh, and Phan Thanh An. Multivalued solution mappings of partially pseudomonotone variational inequality problems and applications. Numerical Algorithms, 102(2):901–921, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02160-3.

Thong:2020:SCE

[TTLD20] Duong Viet Thong, Nguyen Anh Triet, Xiao-Huan Li, and Qiao-Li Dong. Strong convergence of extragradient methods for solving bilevel pseudo-monotone variational inequality problems. Numerical Algorithms, 83(3):1123–1143, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tomar:2026:SON

[TTP26] Aditi Tomar, Lok Pati Tripathi, and Amiya Kumar Pani. A second order non-uniform IMEX-Alikhanov-FEM for timefractional PDEs and PIDEs with time-dependent coefficients. Numerical Algorithms, 101(2):951–991, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02029-5.

Telen:2021:BEM

[TTV21] Simon Telen, Sascha Timme, and Marc Van Barel. Backward error measures for roots of polynomials. Numerical Algorithms, 87(1):19–39, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00956-z.

Tang:2023:CSC

[TTXZ23] Chunming Tang, Wancheng Tan, Shajie Xing, and Haiyan Zheng. A class of spectral conjugate gradient methods for Riemannian optimization. Numerical Algorithms, 94(1):131–147, September 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01495-5.

Tuomela:1998:RSO

[Tuo98] Jukka Tuomela. On the resolution of singularities of ordinary differential systems. Numerical Algorithms, 19(1– 4):247–259, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/16/
20/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/16/20/fulltext.pdf. Differential

algebraic equations (Grenoble, 1997).

Turcajova:1994:FCL

[Tur94] Radka Turcajová. Factorizations and construction of linear phase paraunitary filter banks and higher multiplicity wavelets. Numerical Algorithms, 8(1):1–25, ???? 1994. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Turhan:2026:IDD

[Tur26] Nezihe Turhan. Implementation of divided differences to shape preservation. Numerical Algorithms, 102(1):365–408, May 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02134-5.

Themistoclakis:2017:GVP

[TV17] Woula Themistoclakis and Marc Van Barel. Generalized de la Vallée Poussin approximations on [−1, 1]. Numerical Algorithms, 75(1):1–31, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Themistoclakis:2019:UAS

[TV19] Woula Themistoclakis and Marc Van Barel. Uniform approximation on the sphere by least squares polynomials. Numerical Algorithms, 81(3):1089–1111, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Themistoclakis:2025:FPD

[TV25] Woula Themistoclakis and Marc Van Barel. A free parameter dependent family of polynomial wavelets on a compact interval. Numerical Algorithms, 100(4):1679–1714, December 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02098-6.

Thong:2020:NSC

[TVC20] Duong Viet Thong, Nguyen The Vinh, and Yeol Je Cho. New strong convergence theorem of the inertial projection and contraction method for variational inequality problems. Numerical Algorithms, 84(1):285–305, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tang:2019:NES

[TX19] Xiao Tang and Aiguo Xiao. New explicit stabilized stochastic Runge–Kutta methods with weak second order for stiff Itô stochastic differential equations. Numerical Algorithms, 82(2): 593–604, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tong:1996:LSS

[TY96] Charles H. Tong and Qiang Ye. A linear system solver based on a modified Krylov subspace method for breakdown recovery. Numerical Algorithms, 12(1–2):233–251, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Tokutome:2021:AAD

[TY21] Kimiki Tokutome and Toshihiro Yamada. Acceleration of automatic differentiation of solutions to parabolic partial differential equations: a higher order discretization. Numerical Algorithms, 86(2):593–635, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-020-00902-
z.

Teng:2020:ALP

[TYSY20] Yue Teng, Li Yang, Xiaoliang Song, and Bo Yu. An augmented Lagrangian proximal alternating method for sparse discrete optimization problems. Numerical Algorithms, 83(3): 833–866, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Teng:2024:HOB

[TZ24] Xueqing Teng and Hong Zhang. High-order L^{2}-bound-preserving Fourier pseudo-spectral schemes for the Allen–Cahn equation. Numerical Algorithms, 97(4):1859–1894, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01772-5.

Tang:2025:TSB

[TZ25] Jingyong Tang and Jinchuan Zhou. A two-step Broyden-like method for nonlinear equations. Numerical Algorithms, 98 (3):1085–1105, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01827-7.

Talaei:2024:EIR

[TZH24] Younes Talaei, Mahmoud A. Zaky, and Ahmed S. Hendy. An easy-to-implement recursive fractional spectral-Galerkin method for multi-term weakly singular Volterra integral equations with non-smooth solutions. Numerical Algorithms, 97(3): 1089–1117, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01742-3.

Uko:2009:GEV

[UA09] Livinus U. Uko and Ioannis K. Argyros. Generalized equations, variational inequalities and a weak Kantorovich theorem. Numerical Algorithms, 52(3):321–333, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=321.

Uhlig:2009:GCN

[Uhl09] Frank Uhlig. Geometric computation of the numerical radius of a matrix. Numerical Algorithms, 52(3):335–353, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=3&
spage=335.

Uhlig:2022:CUB

[Uhl22a] Frank Uhlig. Correction to: On the unitary blockdecomposability of 1-parameter matrix flows and static matrices. Numerical Algorithms, 89(3):1413–1414, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/
10.1007/s11075-021-01216-4;   http://link.springer.
com/content/pdf/10.1007/s11075-021-01216-4.pdf. See

[Uhl22b].

Uhlig:2022:UBD

[Uhl22b] Frank Uhlig. On the unitary block-decomposability of 1parameter matrix flows and static matrices. Numerical Algorithms, 89(2):529–549, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01124-

7. See correction [Uhl22a].

Usta:2018:MQI

[UL18] Fuat Usta and Jeremy Levesley. Multilevel quasi-interpolation on a sparse grid with the Gaussian. Numerical Algorithms, 77 (3):793–808, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Upadhyay:2025:CDT

[UPS25] B. B. Upadhyay, Rupesh K. Pandey, and Shubham Kumar Singh. On conjugate direction-type method for intervalvalued multiobjective quadratic optimization problems. Numerical Algorithms, 100(1):425–453, September 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01967-w.

Upadhyay:2026:IPP

[UPY26] Balendu Bhooshan Upadhyay, Subham Poddar, and Jen-Chih Yao. An inexact proximal point method with quasi-distance for quasiconvex multiobjective optimization problems on Riemannian manifolds. Numerical Algorithms, 102(2):1265–1315, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02177-8.

Ullah:2014:NSN

[USAF14] Malik Zaka Ullah, Fazlollah Soleymani, and A. S. Al-Fhaid. Numerical solution of nonlinear systems by a general class of iterative methods with application to nonlinear PDEs. Numerical Algorithms, 67(1):223–242, September 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9784-x.

Ushiyama:2024:PPC

[USM24] Kansei Ushiyama, Shun Sato, and Takayasu Matsuo. Properties and practicability of convergence-guaranteed optimization methods derived from weak discrete gradients. Numerical Algorithms, 96(3):1331–1362, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01790-
3.

Ueno:2007:NSB

[UTO07] Toshihide Ueno, Simon Truscott, and Masami Okada. New spline basis functions for sampling approximations. Numerical Algorithms, 45(1–4):283–293, August 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-007-9119-x.

Uchino:2024:AIR

[UTO24] Yuki Uchino, Takeshi Terao, and Katsuhisa Ozaki. Acceleration of iterative refinement for singular value decomposition. Numerical Algorithms, 95(2):979–1009, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01596-9.

Vaish:2020:GVI

[VA20] Rajat Vaish and Md. Kalimuddin Ahmad. Generalized viscosity implicit scheme with Meir–Keeler contraction for asymptotically nonexpansive mapping in Banach spaces. Numerical Algorithms, 84(3):1217–1237, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00798-
4.

Valkov:2014:FFV

[Val14] Radoslav Valkov. Fitted finite volume method for a generalized Black–Scholes equation transformed on finite interval. Numerical Algorithms, 65(1):195–220, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9701-3.

Valkov:2015:CFV

[Val15] Radoslav Valkov. Convergence of a finite volume element method for a generalized Black–Scholes equation transformed on finite interval. Numerical Algorithms, 68(1):61–80, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9838-8.

VanIseghem:1992:BCP

[Van92] Jeannette Van Iseghem. Best choice of the pole for the Padétype approximant of a Stieltjes function. Numerical Algorithms, 3(1–4):463–475, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

vanDamme:1993:ADA

[van93] Ruud van Damme. An algorithm for determining the approximation orders of multivariate periodic spline spaces. Numerical Algorithms, 5(1–4):71–81, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

VanIseghem:2002:VOL

[Van02] Jeannette Van Iseghem. Vector-orthogonality and Lanczostype methods. Numerical Algorithms, 29(1–3):267–279, March 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp008.lwwonline.com/
content/getfile/5058/38/15/abstract.htm;   http://
ipsapp008.lwwonline.com/content/getfile/5058/38/15/
fulltext.pdf.

VanIseghem:2003:VSC

[Van03] Jeannette Van Iseghem. Vector Stieltjes continued fraction and vector QD algorithm. Numerical Algorithms, 33 (1–4):485–498, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
40/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/40/fulltext.pdf.

VanDeun:2007:EPC

[Van07] Joris Van Deun. Eigenvalue problems to compute almost optimal points for rational interpolation with prescribed poles. Numerical Algorithms, 45(1–4):89–99, August 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=45&issue=1&spage=89.

Gorder:2012:CEH

[Van12] Robert A. Van Gorder. Control of error in the homotopy analysis of semi-linear elliptic boundary value problems. Numerical Algorithms, 61(4):613–629, December 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9554-1/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
61&issue=4&spage=613-629.

Gorder:2017:UHA

[Van17] Robert A. Van Gorder. On the utility of the homotopy analysis method for non-analytic and global solutions to nonlinear differential equations. Numerical Algorithms, 76(1):151– 162, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11075-016-0248-y.pdf.

VanGorder:2019:OHA

[Van19] Robert A. Van Gorder. Optimal homotopy analysis and control of error for implicitly defined fully nonlinear differential equations. Numerical Algorithms, 81(1):181–196, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-018-0540-0.pdf.

VanAssche:2025:GWV

[Van25] Walter Van Assche. A Golub–Welsch version for simultaneous Gaussian quadrature. Numerical Algorithms, 100(2):599–620, October 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01767-2.

VanBarel:1991:CNP

[VB91] Marc Van Barel and Adhemar Bultheel. The computation of non-perfect Padé–Hermite approximants. Numerical Algorithms, 1(3):285–304, September 1991. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

VanBarel:1992:GMT

[VB92] M. Van Barel and A. Bultheel. A general module-theoretic framework for vector M-padé and matrix rational interpolation. Numerical Algorithms, 3(1–4):451–461, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Varadhan:1996:ADS

[VBG96] Sowmini Varadhan, Michael W. Berry, and Gene H. Golub. Approximating dominant singular triplets of large sparse matrices via modified moments. Numerical Algorithms, 13(1–2): 123–152, ???? 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Varga:1992:SNR

[VC92] Richard S. Varga and Amos J. Carpenter. Some numerical results on best uniform rational approximation of x^{α} on [0, 1]. Numerical Algorithms, 2(2):171–185, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Varga:2000:ZPS

[VC00] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z). I. Numerical Algorithms, 25(1–4):363–375, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/2/fulltext.pdf. Mathemat-

ical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

VanBarel:2010:MCR

[VC10] Marc Van Barel and Andrey Chesnokov. A method to compute recurrence relation coefficients for bivariate orthogonal polynomials by unitary matrix transformations. Numerical Algorithms, 55(2–3):383–402, November 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=55&issue=2&spage=383.

vanderHouwen:1998:SMS

[vdHM98] P. J. van der Houwen and E. Messina. Splitting methods for second-order initial value problems. Numerical Algorithms, 18(3–4):233–257, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
15/1/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/15/1/fulltext.pdf.

vanderHouwen:2002:AFS

[vdHS02] P. J. van der Houwen and B. P. Sommeijer. Approximate factorization in shallow water applications. Numerical Algorithms, 31(1–4):337–360, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
14/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/14/fulltext.pdf.

vanderMee:2006:FCT

[vdMRS06] C. van der Mee, G. Rodriguez, and S. Seatzu. Fast computation of two-level circulant preconditioners. Numerical Algorithms, 41(3):275–295, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=41&issue=3&spage=275.

Vasconcelos:2013:JDT

[VdR13] Paulo B. Vasconcelos, Filomena D. d’Almeida, and Jose E. Roman. A Jacobi–Davidson type method with a correction equation tailored for integral operators. Numerical Algorithms, 64(1):85–103, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9656-9.

Verdiere:2012:NME

[VDVJB12] Nathalie Verdière, Lilianne Denis-Vidal, and Ghislaine Joly-Blanchard. A new method for estimating derivatives based on a distribution approach. Numerical Algorithms, 61(1):163–186, September 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=1&spage=163.

Vepstas:2008:EAA

[Vep08] Linas Vepstas. An efficient algorithm for accelerating the convergence of oscillatory series, useful for computing the polylogarithm and Hurwitz zeta functions. Numerical Algorithms, 47(3):211–252, March 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=3&spage=211.

Verlinden:1999:SRM

[Ver99] Pierre Verlinden. Stable rational modification of a weight. Numerical Algorithms, 22(2):183–192, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp007.kluweronline.com/content/getfile/
5058/22/5/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/22/5/fulltext.pdf.

Verner:2010:NOR

[Ver10] J. H. Verner. Numerically optimal Runge–Kutta pairs with interpolants. Numerical Algorithms, 53(2–3):383–396, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=383.

Verner:2014:ERK

[Ver14] J. H. Verner. Explicit Runge–Kutta pairs with lower stageorder. Numerical Algorithms, 65(3):555–577, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9783-y.

Verner:2024:NYN

[Ver24] J. H. Verner. Nullspaces yield new explicit Runge–Kutta pairs. Numerical Algorithms, 96(3):1363–1390, July 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01731-6.

vonGudenberg:2004:FFB

[vGK04] Jürgen Wolff von Gudenberg and Vladik Kreinovich. A full function-based calculus of directed and undirected intervals: Markov’s interval arithmetic revisited. Numerical Algorithms, 37(1–4):417–428, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/33/abstract.htm.

VanIseghem:1996:AFG

[VGM96] Jeannette Van Iseghem and Peter R. Graves-Morris. Approximation of a function given by its Laurent series. Numerical Algorithms, 11(1–4):339–351, March 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Orthogonal polynomials and numerical analysis (Luminy, 1995).

Vandebril:2006:SNS

[VGV06] Raf Vandebril, Gene Golub, and Marc Van Barel. A small note on the scaling of symmetric positive definite semiseparable matrices. Numerical Algorithms, 41(3):319–326, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=41&issue=3&
spage=319.

Verlinden:1992:AEW

[VH92] Pierre Verlinden and Ann Haegemans. An asymptotic expansion in wavelet analysis and its application to accurate numerical wavelet decomposition. Numerical Algorithms, 2(3–4): 287–298, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Valinejad:2010:VSS

[VH10] A. Valinejad and S. Mohammad Hosseini. A variable step-size control algorithm for the weak approximation of stochastic differential equations. Numerical Algorithms, 55(4):429–446, December 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
55&issue=4&spage=429.

Valinejad:2012:SCA

[VH12] A. Valinejad and S. Mohammad Hosseini. A stepsize control algorithm for SDEs with small noise based on stochastic Runge–Kutta Maruyama methods. Numerical Algorithms, 61(3):479–498, November 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=479.

Vignes:2004:DSA

[Vig04]  Jean Vignes.  Discrete stochastic arithmetic for vali-

dating results of numerical software. Numerical Algorithms, 37(1–4):377–390, December 2004. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://ipsapp009.kluweronline.com/IPS/content/
ext/x/J/5058/I/58/A/31/abstract.htm.

Vlcek:2019:PBB

[VL19] Jan Vlcek and Ladislav Luksan. Properties of the block BFGS update and its application to the limited-memory block BNS method for unconstrained minimization. Numerical Algorithms, 80(3):957–987, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Vong:2016:HOF

[VLCL16] Seakweng Vong, Pin Lyu, Xu Chen, and Siu-Long Lei. High order finite difference method for time-space fractional differential equations with Caputo and Riemann–Liouville derivatives. Numerical Algorithms, 72(1):195–210, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0041-3.

vanLent:2002:MWR

[vLV02] Jan van Lent and Stefan Vandewalle. Multigrid waveform relaxation for anisotropic partial differential equations. Numerical Algorithms, 31(1–4):361–380, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://ipsapp007.kluweronline.com/
content/getfile/5058/43/15/abstract.htm;   http://
ipsapp007.kluweronline.com/content/getfile/5058/43/
15/fulltext.pdf.

Verma:2017:PWU

[VM17] Ajeet Singh Verma and Pankaj Mathur. Polynomial wavelets on the unit circle. Numerical Algorithms, 76(4):977–992, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Voinea-Marinescu:2021:BFR

[VMMD21] Andreea-Paula Voinea-Marinescu, Liviu Marin, and Franck Delvare. BEM-Fading regularization algorithm for Cauchy problems in 2D anisotropic heat conduction. Numerical Algorithms, 88(4):1667–1702, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01090-
0.

Vulanovic:2018:NMS

[VN18] Relja Vulanović and Thái Anh Nhan. A numerical method for stationary shock problems with monotonic solutions. Numerical Algorithms, 77(4):1117–1139, April 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Voss:2000:SEL

[Vos00] Heinrich Voss. A symmetry exploiting Lanczos method for symmetric Toeplitz matrices. Numerical Algorithms, 25(1– 4):377–385, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/10/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/10/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Verma:2023:DNF

[VP23] Manuj Verma and Amit Priyadarshi. Dimensions of new fractal functions and associated measures. Numerical Algorithms, 94(2):817–846, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01521-0.

Vivas:2024:NNM

[VPA24] Hevert Vivas, Rosana Pérez, and Carlos A. Arias. A nonsmooth Newton method for solving the generalized complementarity problem. Numerical Algorithms, 95(2):551–574, ????

2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01581-2.

Vivas:2025:LCS

[VPA25] H. Vivas, R. Pérez, and C. Arias. A least-change secant algorithm for solving generalized complementarity problem. Numerical Algorithms, 99(1):75–100, May 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01870-4.

Verlinden:1997:EEM

[VPL97] P. Verlinden, D. M. Potts, and J. N. Lyness. Error expansions for multidimensional trapezoidal rules with Sidi transformations. Numerical Algorithms, 16(3–4):321–347, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/11/5/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
11/5/fulltext.pdf.

Vinko:2004:MBP

[VR04] Tamás Vinkó and Dietmar Ratz. A multidimensional branchand-prune method for interval global optimization. Numerical Algorithms, 37(1–4):391–399, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/38/abstract.htm.

Vasconcelos:2023:SDE

[VRM23] P. B. Vasconcelos, J. E. Roman, and J. M. A. Matos. Solving differential eigenproblems via the spectral Tau method. Numerical Algorithms, 92(3):1789–1811, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01366-z.

Vuong:2019:CET

[VS19] Phan Tu Vuong and Yekini Shehu. Convergence of an extragradient-type method for variational inequality with applications to optimal control problems. Numerical Algorithms, 81(1):269–291, May 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Vosoughi:2012:UMP

[VSA12] Hossein Vosoughi, Elyas Shivanian, and Saeid Abbasbandy. Unique and multiple PHAM series solutions of a class of nonlinear reactive transport model. Numerical Algorithms, 61(3):515–524, November 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=515.

vanderHouwen:1994:PAS

[vSv94] P. J. van der Houwen, B. P. Sommeijer, and W. A. van der Veen. Parallelism across the steps in iterated Runge–Kutta methods for stiff initial value problems. Numerical Algorithms, 8(2–4):293–312, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Vulanovic:2010:UNM

[VT10] Relja Vulanović and Ljiljana Teofanov. A uniform numerical method for semilinear reaction-diffusion problems with a boundary turning point. Numerical Algorithms, 54(4):431–444, August 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
54&issue=4&spage=431.

Vinh:2022:SOD

[VTV22] Le Van Vinh, Van Nam Tran, and Phan Tu Vuong. A second-order dynamical system for equilibrium problems. Numerical Algorithms, 91(1):327–351, September 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01264-4.

Vulanovic:1997:FOA

[Vul97] Relja Vulanović. Fourth order algorithms for a semilinear singular perturbation problem. Numerical Algorithms, 16 (2):117–128, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
10/4/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/10/4/fulltext.pdf.

Daele:2007:PSE

[VV07a] M. Van Daele and G. Vanden Berghe. P-stable exponentiallyfitted Obrechkoff methods of arbitrary order for second-order differential equations. Numerical Algorithms, 46(4):333–350, December 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
46&issue=4&spage=333.

Daele:2007:PSO

[VV07b] Marnix Van Daele and Guido Vanden Berghe. P-stable Obrechkoff methods of arbitrary order for second-order differential equations. Numerical Algorithms, 44(2):115–131, February 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=44&
issue=2&spage=115.

VandenBerghe:2011:SEF

[VV11] Guido Vanden Berghe and Marnix Van Daele. Symplectic exponentially-fitted four-stage Runge–Kutta methods of the Gauss type. Numerical Algorithms, 56(4):591–608, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=591.

Vermeylen:2025:SIF

[VV25] Charlotte Vermeylen and Marc Van Barel. Stability improvements for fast matrix multiplication. Numerical Algorithms, 100(2):645–683, October 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01806-y.

Vermeylen:2025:RSB

[VVL25] Charlotte Vermeylen, Nico Vervliet, and Lieven De Lathauwer. Reducing swamp behavior for the canonical polyadic decomposition problem by rank-1 freezing. Numerical Algorithms, 100 (2):831–859, October 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02111-y.

Buggenhout:2022:GOR

[VVV22] Niel Van Buggenhout, Marc Van Barel, and Raf Vandebril. Generation of orthogonal rational functions by procedures for structured matrices. Numerical Algorithms, 89(2):551–582, February 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01125-6.

Vale:2008:TFG

[VW08] Richard Vale and Shayne Waldron. Tight frames generated by finite nonabelian groups. Numerical Algorithms, 48(1–3): 11–27, July 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
48&issue=1&spage=11.

VanHuffel:1993:ETL

[VZ93] Sabine Van Huffel and Hong Yuan Zha. An efficient total least squares algorithm based on a rank-revealing two-sided orthogonal decomposition. Numerical Algorithms, 4(1–2):101–133, January 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Walz:1994:AEC

[Wal94] Guido Walz. Asymptotic expansions for classical and generalized divided differences including applications. Numerical Algorithms, 7(2–4):161–171, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Waldron:2006:COP

[Wal06] Shayne Waldron. Computing orthogonal polynomials on a triangle by degree raising. Numerical Algorithms, 42(2):171–179, June 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=42&
issue=2&spage=171.

Waldron:2007:HPP

[Wal07] Shayne Waldron. Hermite polynomials on the plane. Numerical Algorithms, 45(1–4):231–238, August 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-007-9085-3.

Wang:2015:PSP

[Wan15a] Caifang Wang. A parallel strategy and performance analysis of the EM-like reconstruction method for diffuse optical tomography. Numerical Algorithms, 68(2):377–391, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9852-x.

Wang:2015:CFD

[Wan15b] Yuan-Ming Wang. A compact finite difference method for a class of time fractional convection-diffusion-wave equations with variable coefficients. Numerical Algorithms, 70(3):625– 651, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-015-9965-x.

Wang:2017:CCA

[Wan17] Fenghui Wang. On the convergence of CQ algorithm with variable steps for the split equality problem. Numerical Algorithms, 74(3):927–935, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0177-9.

Wang:2018:PGM

[Wan18] Fenghui Wang. Polyak’s gradient method for split feasibility problem constrained by level sets. Numerical Algorithms, 77 (3):925–938, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2019:CNT

[Wan19] Yuan-Ming Wang. A Crank–Nicolson-type compact difference method and its extrapolation for time fractional Cattaneo convection-diffusion equations with smooth solutions. Numerical Algorithms, 81(2):489–527, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2022:NPI

[Wan22] Yan Wang. On nested Picard iterative integrators for highly oscillatory second-order differential equations. Numerical Algorithms, 91(4):1627–1651, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01317-8.

Wang:2024:DSQ

[Wan24] Hong Wang. A decentralized smoothing quadratic regularization algorithm for composite consensus optimization with non-Lipschitz singularities. Numerical Algorithms, 96(1):369– 396, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01650-6.

Wang:2025:ANS

[Wan25] Xiaofeng Wang. Analytical and numerical studies of the modified Kawahara equation with dual-power law nonlinearities. Numerical Algorithms, 98(3):1107–1139, March 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01828-6.

Watson:1992:LBA

[Wat92] G. A. Watson. Linear best approximation using a class of polyhedral norms. Numerical Algorithms, 2(3–4):321–335, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Watson:1993:MAP

[Wat93] G. A. Watson. On matrix approximation problems with Ky Fan k norms. Numerical Algorithms, 5(1–4):263–272, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Watson:1994:LBA

[Wat94] G. A. Watson. Linear best approximation using a class of k-major l_{p} norms. Numerical Algorithms, 8(1):135–146, ???? 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Watson:2006:FEC

[Wat06] G. Alistair Watson. Fitting enclosing cylinders to data in R^{n}. Numerical Algorithms, 43(2):189–196, October 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=43&issue=2&spage=189.

Worz:2017:GCD

[WB17] Sascha Wörz and Heinz Bernhardt. A global convergent derivative-free method for solving a system of non-linear equations. Numerical Algorithms, 76(1):109–124, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2025:PEEb

[WBY25] Shixi Wang, Hai Bi, and Yidu Yang. The a posteriori error estimates and an adaptive algorithm of the discontinuous Galerkin method for the modified transmission eigenvalue problem with absorbing media. Numerical Algorithms, 100(3):1085–1122, November 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01981-y.

Wu:2010:TPH

[WC10] Yongyan Wu and Kwok Fai Cheung. Two-parameter homotopy method for nonlinear equations. Numerical Algorithms, 53(4):555–572, April 2010. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=53&issue=4&spage=555.

Wu:2013:CAM

[WC13] Qingbiao Wu and Minhong Chen. Convergence analysis of modified Newton–HSS method for solving systems of nonlinear equations. Numerical Algorithms, 64(4):659–683, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9684-5.

Wan:2023:WHT

[WC23] Xiying Wan and Guanghui Cheng. Weighted hybrid truncated norm regularization method for low-rank matrix completion. Numerical Algorithms, 94(2):619–641, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01513-0.

Wang:2024:SOS

[WC24] Yibo Wang and Wanrong Cao. Strong 1.5 order scheme for fractional Langevin equation based on spectral approximation of white noise. Numerical Algorithms, 95(1):423–450, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01576-z.

Wang:2025:DDS

[WC25] Zheng Wang and Yuhao Cong. Delay-dependent stability of a class of Runge–Kutta methods for neutral differential equations. Numerical Algorithms, 98(4):1643–1666, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01846-4.

Wang:2015:CFG

[WCB15] Liqi Wang, Moody T. Chu, and Yu Bo. A computational framework of gradient flows for general linear matrix equations. Numerical Algorithms, 68(1):121–141, January 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9885-1.

Wu:2021:SMS

[WCD21] Longbin Wu, Zhong Chen, and Xiaohua Ding. A stable minimal search method for solving multi-order fractional differential equations based on reproducing kernel space. Numerical Algorithms, 87(4):1707–1727, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01026-0.

Williams:2015:NCT

[WCH15] Bryan Michael Williams, Ke Chen, and Simon P. Harding. A new constrained total variational deblurring model and its fast algorithm. Numerical Algorithms, 69(2):415–441, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9904-2.

Wang:2021:MSC

[WCHK21] Peng Wang, Yanzhao Cao, Xiaoying Han, and Peter Kloeden. Mean-square convergence of numerical methods for random ordinary differential equations. Numerical Algorithms, 87 (1):299–333, May 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00967-w.

Wang:2016:INU

[WCLW16] H. J. Wang, F. Chen, H. Liu, and Q. Wang. Incomplete Newton–Ulm method for large scale nonlinear equations. Numerical Algorithms, 72(2):409–424, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0052-0.

Wright:1994:MSS

[WCM94] R. Wright, J. Cash, and G. Moore. Mesh selection for stiff two-point boundary value problems. Numerical Algorithms, 7 (2–4):205–224, July 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2020:GUS

[WCW20] Xuezhong Wang, Maolin Che, and Yimin Wei. Global uniqueness and solvability of tensor complementarity problems for H_{+}-tensors. Numerical Algorithms, 84(2):567–590, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00769-9.

Wang:2025:GMR

[WCY25] Longfei Wang, Yu Chen, and Meijia Yang. Globally maximizing the ratio of two generalized quadratic matrix form functions over the Stiefel manifold. Numerical Algorithms, 99(4): 1943–1965, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01939-0.

Wang:2025:TSN

[WCZ25] Lu-Xin Wang, Yang Cao, and Chen-Can Zhou. Two-step nonlinear modulus-based matrix splitting iteration method for implicit complementarity problems. Numerical Algorithms, 100 (3):1289–1314, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01988-

Willemans:1995:NSF

[WD95] Karin Willemans and Paul Dierckx. Nonnegative surface fitting with Powell–Sabin splines. Numerical Algorithms, 9(3–4):263– 276, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Willemans:1996:SSD

[WD96] Karin Willemans and Paul Dierckx. Smoothing scattered data with a monotone Powell–Sabin spline surface. Numerical Algorithms, 12(1–2):215–232, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2023:LPS

[WD23] Xinhui Wang and Guangzhi Du. Local and parallel stabilized finite element methods based on two-grid discretizations for the Stokes equations. Numerical Algorithms, 93(1):67– 83, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01403-x.

Wang:2016:IPI

[WDL16] Ke Wang, Jingjing Di, and Don Liu. Improved PHSS iterative methods for solving saddle point problems. Numerical Algorithms, 71(4):753–773, April 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-0022-6.

Wang:2023:MLP

[WDL23] Xinhui Wang, Guangzhi Du, and Yi Li. A modified local and parallel finite element method for the coupled Stokes–Darcy model with the Beavers–Joseph interface condition. Numerical Algorithms, 93(2):815–831, June 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01442-4.

Wei:2004:EET

[WDY04] Musheng Wei, Alvaro R. De Pierro, and Jiahong Yin. Error estimates for two filters based on polynomial interpolation for recovering a function from its Fourier coefficients. Numerical Algorithms, 35(2–4):205–231, April 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/13/abstract.htm.

Wei:2017:SCF

[Wei17] Leilei Wei. Stability and convergence of a fully discrete local discontinuous Galerkin method for multi-term time fractional diffusion equations. Numerical Algorithms, 76(3):695– 707, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wei:2018:ANF

[Wei18] Leilei Wei. Analysis of a new finite difference/local discontinuous Galerkin method for the fractional Cattaneo equation. Numerical Algorithms, 77(3):675–690, March 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Weniger:1992:IBS

[Wen92] Ernst Joachim Weniger. Interpolation between sequence transformations. Numerical Algorithms, 3(1–4):477–486, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Weniger:2003:RAD

[Wen03] Ernst Joachim Weniger. A rational approximant for the digamma function. Numerical Algorithms, 33(1–4):499– 507, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/46/41/abstract.
htm;   http://ipsapp007.kluweronline.com/content/getfile/
5058/46/41/fulltext.pdf.

Weideman:2023:FNL

[WF23] J. A. C. Weideman and Bengt Fornberg. Fully numerical Laplace transform methods. Numerical Algorithms, 92(1): 985–1006, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01368-x.

Wakrim:1999:STF

[WG99] Mohamed Wakrim and Fattehallah Ghadi. On some three-field mixed systems. Numerical Algorithms, 21(1– 4):367–375, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
22/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/22/fulltext.pdf.

Wang:2013:RKT

[WG13] Xiaojie Wang and Siqing Gan. A Runge–Kutta type scheme for nonlinear stochastic partial differential equations with multiplicative trace class noise. Numerical Algorithms, 62(2): 193–223, February 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9568-8/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=2&spage=193-223.

Wang:2025:GWG

[WG25] Yue Wang and Fuzheng Gao. Generalized weak Galerkin finite element method for linear elasticity interface problems. Numerical Algorithms, 99(2):1005–1042, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01904-x.

Wang:2011:SCM

[WGK11] Xiuhua Wang, Chuanqing Gu, and Jisheng Kou. Semilocal convergence of a multipoint fourth-order super-Halley method in Banach spaces. Numerical Algorithms, 56(4):497–516, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=497.

Wang:2018:MDI

[WGZ18a] Jing Wang, Xue-Ping Guo, and Hong-Xiu Zhong. MN-DPMHSS iteration method for systems of nonlinear equations with block two-by-two complex Jacobian matrices. Numerical Algorithms, 77(1):167–184, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2018:AIL

[WGZ18b] Tongke Wang, Yuesheng Gu, and Zhiyue Zhang. An algorithm for the inversion of Laplace transforms using Puiseux expansions. Numerical Algorithms, 78(1):107–132, May 2018.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2025:IPD

[WGZ25] Jueyu Wang, Chao Gu, and Detong Zhu. An interiorpoint derivative-free algorithm for nonlinear complementarity problems. Numerical Algorithms, 100(1):395–424, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01966-x.

Walster:2004:UPF

[WH04] G. William Walster and Eldon Hansen. Using pillow functions to efficiently compute crude range tests. Numerical Algorithms, 37(1–4):401–415, December 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/58/A/32/abstract.htm.

Wang:2015:CLD

[WH15] Pengde Wang and Chengming Huang. A conservative linearized difference scheme for the nonlinear fractional Schrödinger equation. Numerical Algorithms, 69(3):625–641, July 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9917-x.

Wang:2022:CMS

[WHD22] Yifei Wang, Jin Huang, and Ting Deng. A combination method for solving multi-dimensional systems of Volterra integral equations with weakly singular kernels. Numerical Algorithms, 91 (2):473–504, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01270-6.

Wang:2024:NAS

[WHL24] Yifei Wang, Jin Huang, and Hu Li. A numerical approach for the system of nonlinear variable-order fractional Volterra integral equations. Numerical Algorithms, 95(4):1855–1877, April 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01630-w.

Waziri:2020:DPC

[WHS20] Mohammed Yusuf Waziri, Kabiru Ahmed Hungu, and Jamilu Sabi’u. Descent Perry conjugate gradient methods for systems of monotone nonlinear equations. Numerical Algorithms, 85(3): 763–785, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00836-1.

Wen:2023:APM

[WHS23] Chun Wen, Qian-Ying Hu, and Zhao-Li Shen. An adaptively preconditioned multi-step matrix splitting iteration for computing PageRank. Numerical Algorithms, 92(2):1213–1231, February 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01337-4.

Wang:2018:SNR

[WHZ^{+}18] Si Wang, Ting-Zhu Huang, Xi-Le Zhao, Jin-Jin Mei, and Jie Huang. Speckle noise removal in ultrasound images by firstand second-order total variation. Numerical Algorithms, 78 (2):513–533, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Willis:2012:AGH

[Wil12] Joshua L. Willis. Acceleration of generalized hypergeometric functions through precise remainder asymptotics. Numerical Algorithms, 59(3):447–485, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=447.

Wimp:1999:CCO

[Wim99] Jet Wimp. Connection coefficients, orthogonal polynomials and the WZ-algorithms. Numerical Algorithms, 21(1– 4):377–386, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
18/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/18/fulltext.pdf. Numerical

methods for partial differential equations (Marrakech, 1998).

Wimp:2000:HDS

[Wim00] Jet Wimp. Hankel determinants of some polynomials arising in combinatorial analysis. Numerical Algorithms, 24 (1–2):179–193, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/27/
11/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/27/11/fulltext.pdf. Computa-

tional methods from rational approximation theory (Wilrijk, 1999).

Witomski:1996:ACP

[Wit96] P. Witomski. Approximation of the capillarity problem by an augmented Lagrangian method. Numerical Algorithms, 12(3– 4):321–346, July 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2025:PEEa

[WJC25] Xuesong Wang, Shan Jiang, and Yao Cheng. Pointwise error estimate of the LDG method for 2D singularly perturbed reaction-diffusion problem. Numerical Algorithms, 100(1): 133–187, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01958-x.

Wang:2014:MHA

[WJW14] Fu-Sheng Wang, Jin-Bao Jian, and Chuan-Long Wang. A model-hybrid approach for unconstrained optimization problems. Numerical Algorithms, 66(4):741–759, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9757-0.

Williams:1993:NCF

[WK93] Jack Williams and Z. Kalogiratou. Nonlinear Chebyshev fitting from the solution of ordinary differential equations. Numerical Algorithms, 5(1–4):325–337, November 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for approximation, III (Oxford, 1992).

Wang:2012:SCM

[WK12] Xiuhua Wang and Jisheng Kou. Semilocal convergence of a modified multi-point Jarratt method in Banach spaces under general continuity condition. Numerical Algorithms, 60 (3):369–390, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=60&issue=3&spage=369.

Wang:2013:SCO

[WK13] Xiuhua Wang and Jisheng Kou. Semilocal convergence and Rorder for modified Chebyshev–Halley methods. Numerical Algorithms, 64(1):105–126, September 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9657-8.

Wang:2014:CFM

[WK14] Xiuhua Wang and Jisheng Kou. Convergence for a family of modified Chebyshev methods under weak condition. Numerical Algorithms, 66(1):33–48, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9722-y.

Wang:2015:CCM

[WK15] Xiuhua Wang and Jisheng Kou. Convergence for a class of multi-point modified Chebyshev–Halley methods under the relaxed conditions. Numerical Algorithms, 68(3):569–583, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9861-9.

Wang:2016:SCM

[WK16] Xiuhua Wang and Jisheng Kou. Semilocal convergence of multi-point improved super-Halley-type methods without the second derivative under generalized weak condition. Numerical Algorithms, 71(3):567–584, March 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-0010-x.

Wang:2017:SCF

[WK17] Xiuhua Wang and Jisheng Kou. Semilocal convergence on a family of root-finding multi-point methods in Banach spaces under relaxed continuity condition. Numerical Algorithms, 74 (3):643–657, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0165-0.

Wu:2020:BEL

[WK20] Xian-Ping Wu and Ri-Huan Ke. Backward errors of the linear complementarity problem. Numerical Algorithms, 83(3): 1249–1257, March 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2011:SCS

[WKG11] Xiuhua Wang, Jisheng Kou, and Chuanqing Gu. Semilocal convergence of a sixth-order Jarratt method in Banach spaces. Numerical Algorithms, 57(4):441–456, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=4&spage=441.

Wu:2025:EQC

[WKY25] Pengling Wu, Kit Ian Kou, and Zhaoyuan Yu. Efficient quaternion CUR method for low-rank approximation to quaternion matrix. Numerical Algorithms, 99(3):1503–1530, July 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01923-8.

Weideman:2000:QRB

[WL00] J. A. C. Weideman and D. P. Laurie. Quadrature rules based on partial fraction expansions. Numerical Algorithms, 24 (1–2):159–178, December 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/27/
10/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/27/10/fulltext.pdf. Computa-

tional methods from rational approximation theory (Wilrijk, 1999).

Wu:2017:MCS

[WL17] Shi-Liang Wu and Cui-Xia Li. Modified complex-symmetric and skew-Hermitian splitting iteration method for a class of complex-symmetric indefinite linear systems. Numerical Algorithms, 76(1):93–107, September 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2022:MBC

[WL22a] Yan Wang and Chenliang Li. A modulus-based cascadic multigrid method for elliptic variational inequality problems.

Numerical Algorithms, 90(4):1777–1791, August 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01251-1.

Wu:2022:NMB

[WL22b] Shiliang Wu and Liang Li. New modulus-based matrix splitting methods for implicit complementarity problem. Numerical Algorithms, 90(4):1735–1754, August 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01249-9.

Wang:2024:CMA

[WL24] Ting Wang and Hongwei Liu. A class of modified accelerated proximal gradient methods for nonsmooth and nonconvex minimization problems. Numerical Algorithms, 95(1):207– 241, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01569-y.

Wang:2025:GDR

[WL25] Dan Wang and Jicheng Li. General double-relaxation twosweep modulus-based matrix splitting iteration methods for horizontal linear complementarity problem. Numerical Algorithms, 98(4):1965–1983, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01860-
6.

Wang:2026:FCA

[WLG26] Tao Wang, Ying Li, and Wenbin Guo. A fast construction algorithm for compact singular value decomposition of dual quaternion matrix and its application. Numerical Algorithms, 101(2): 1073–1096, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02033-9.

Wang:2024:SSC

[WLJ24] Tongke Wang, Huan Lian, and Lu Ji. Singularity separation Chebyshev collocation method for weakly singular Volterra integral equations of the second kind. Numerical Algorithms, 95 (4):1829–1854, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01629-3.

Wang:2012:CUA

[WLL12] Pengfei Wang, Jianping Li, and Qian Li. Computational uncertainty and the application of a high-performance multiple precision scheme to obtaining the correct reference solution of Lorenz equations. Numerical Algorithms, 59(1):147–159, January 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=59&
issue=1&spage=147.

Wang:2026:SPA

[WLL26] Huanxin Wang, Junhan LYU, and Min Li. Stochastic parareal algorithm for stochastic differential equations. Numerical Algorithms, 102(1):267–289, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-025-02130-
9.

Wang:2021:NAD

[WLMA21] Ying Wang, Fawang Liu, Liquan Mei, and Vo V. Anh. A novel alternating-direction implicit spectral Galerkin method for a multi-term time-space fractional diffusion equation in three dimensions. Numerical Algorithms, 86(4):1443–1474, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00940-7.

Wang:2025:DNC

[WLS25] Zheng Wang, Shi-Hao Li, and Jian-Qing Sun. Discrete noncommutative hungry Toda lattice and its application in matrix computation. Numerical Algorithms, 99(3):1269–1288, July 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01915-8.

Wang:2024:AML

[WLW^{+}24] Tao Wang, Ying Li, Musheng Wei, Yimeng Xi, and Mingcui Zhang. Algebraic method for LU decomposition of dual quaternion matrix and its corresponding structure-preserving algorithm. Numerical Algorithms, 97(3):1367–1382, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01753-8.

Wen:2021:FSO

[WLY^{+}21] Cao Wen, Yang Liu, Baoli Yin, Hong Li, and Jinfeng Wang. Fast second-order time two-mesh mixed finite element method for a nonlinear distributed-order sub-diffusion model. Numerical Algorithms, 88(2):523–553, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01048-8.

Wang:2022:SAB

[WLZ22] Junjun Wang, Meng Li, and Yu Zhang. Superconvergence analysis of BDF–Galerkin FEM for nonlinear Schrödinger equation. Numerical Algorithms, 89(1):195–222, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01111-y.

Wang:2021:ADA

[WMCW21] Xuezhong Wang, Changxin Mo, Maolin Che, and Yimin Wei. Accelerated dynamical approaches for finding the unique positive solution of KS-tensor equations. Numerical Algorithms, 88(4):1787–1810, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01095-9.

Wang:2000:AUI

[WO00] Weichung Wang and Dianne P. O’Leary. Adaptive use of iterative methods in predictor–corrector interior point methods for linear programming. Numerical Algorithms, 25(1– 4):387–406, September 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
3/26/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/3/26/fulltext.pdf. Mathe-

matical journey through analysis, matrix theory and scientific computation (Kent, OH, 1999).

Wu:2018:PGM

[WPL18] Xianping Wu, Xiaofei Peng, and Wen Li. A preconditioned general modulus-based matrix splitting iteration method for linear complementarity problems of H-matrices. Numerical Algorithms, 79(4):1131–1146, December 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2025:EBC

[WPW25] Jinxuan Wang, Kejia Pan, and Xiaoxin Wu. An edgebased cascadic multigrid method for H(curl) problems. Numerical Algorithms, 99(3):1317–1342, July 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01917-6.

Wang:2025:OEE

[WPY25] Jiangxing Wang, Kejia Pan, and Xiaofeng Yang. Optimal error estimate of fully discrete HDG-IEQ scheme for the general type Cahn–Hilliard equation. Numerical Algorithms, 100(1): 455–485, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01968-9.

Wang:2010:BCF

[WQ10] Renhong Wang and Jiang Qian. On branched continued fractions rational interpolation over pyramid-typed grids. Numerical Algorithms, 54(1):47–72, May 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=47.

Wang:2023:SOP

[WQ23] Yongshuai Wang and Yi Qin. Second-order partitioned method and adaptive time step algorithms for the nonstationary Stokes–Darcy equations. Numerical Algorithms, 94(1): 413–457, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01507-y.

Wang:2020:AAL

[WQL20] Tongke Wang, Meng Qin, and Huan Lian. The asymptotic approximations to linear weakly singular Volterra integral equations via Laplace transform. Numerical Algorithms, 85(2):

683–711, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00832-5.

Wang:2024:SOF

[WQZH24] R. Wang, L. Qiao, M. A. Zaky, and A. S. Hendy. A second-order finite difference scheme for nonlinear tempered fractional integrodifferential equations in three dimensions. Numerical Algorithms, 95(1):319–349, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01573-2.

Wright:1995:SFD

[Wri95] R. K. Wright. Spline fitting discontinuous functions given just a few Fourier coefficients. Numerical Algorithms, 9(1–2):157– 169, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Wright:2001:COD

[Wri01] W. M. Wright. The construction of order 4 DIMSIMs for ordinary differential equations. Numerical Algorithms, 26 (2):123–130, February 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
1/3/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/1/3/fulltext.pdf.

Wright:2002:EGL

[Wri02] W. M. Wright. Explicit general linear methods with inherent Runge–Kutta stability. Numerical Algorithms, 31 (1–4):381–399, December 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/43/
16/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/43/16/fulltext.pdf.

Wen:2017:MQC

[WRM17] Rui-Ping Wen, Fu-Jiao Ren, and Guo-Yan Meng. Modified quasi-Chebyshev acceleration to nonoverlapping parallel multisplitting method. Numerical Algorithms, 75(4):1123–1140, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2024:CSM

[WS24a] Yunxia Wang and Zhiyong Si. A convex splitting method for the time-dependent Ginzburg–Landau equation. Numerical Algorithms, 96(2):999–1017, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01672-
0.

Wei:2024:ESS

[WS24b] Siqi Wei and Raymond J. Spiteri. The effect of splitting strategy on qualitative property preservation. Numerical Algorithms, 96(3):1391–1421, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01730-
7.

Wang:2014:CAF

[WSK14] Xiuhua Wang, Dongyang Shi, and Jisheng Kou. Convergence analysis for a family of improved super-Halley methods under general convergence condition. Numerical Algorithms, 65(2): 339–354, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9708-9.

Wang:2024:FCD

[WSLL24] Qiu-Ya Wang, Zi-Hang She, Cheng-Xue Lao, and Fu-Rong Lin. Fractional centered difference schemes and banded preconditioners for nonlinear Riesz space variable-order fractional diffusion equations. Numerical Algorithms, 95(2):859–895, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01592-z.

Wu:2004:CGM

[WSY04] X. Wu, B. P. B. Silva, and J. Y. Yuan. Conjugate gradient method for rank deficient saddle point problems. Numerical Algorithms, 35(2–4):139–154, April 2004. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://ipsapp009.kluweronline.com/IPS/
content/ext/x/J/5058/I/50/A/2/abstract.htm.

Wang:2012:SFE

[WSY12] Kun Wang, Zhiyong Si, and Yanfang Yang. Stabilized finite element method for the viscoelastic Oldroyd fluid flows. Numerical Algorithms, 60(1):75–100, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=1&spage=75.

Wang:2021:NSP

[WSZ21] Xue-Xia Wang, Jian-Qing Sun, and Ying-Nan Zhang. A numerical study on the N-periodic wave solutions of two coupled bilinear equations. Numerical Algorithms, 88(2):711–728, October 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01054-w.

Wang:2008:GMI

[WT08] Xinghua Wang and Peipei Tang. A generalization of Müller’s iteration method based on standard information. Numerical Algorithms, 48(4):347–359, August 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=48&issue=4&spage=347.

Wu:2022:TSR

[Wu22] Wen-Ting Wu. On two-subspace randomized extended Kaczmarz method for solving large linear least-squares problems. Numerical Algorithms, 89(1):1–31, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01104-x.

Wang:2014:HAE

[WW14] Bin Wang and Xinyuan Wu. A highly accurate explicit symplectic ERKN method for multi-frequency and multidimensional oscillatory Hamiltonian systems. Numerical Algorithms, 65(3):705–721, March 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9811-y.

Wang:2019:GEB

[WW19] Bin Wang and Xinyuan Wu. Global error bounds of one-stage extended RKN integrators for semilinear wave equations. Numerical Algorithms, 81(4):1203–1218, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2021:TSO

[WWBM21] Ying Wang, Gang Wang, Linlin Bu, and Liquan Mei. Two second-order and linear numerical schemes for the multi-dimensional nonlinear time-fractional Schrödinger equation. Numerical Algorithms, 88(1):419–451, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01044-y.

Wu:2012:SNT

[WWD^{+}12] Rongxing Wu, Ji Wang, Jianke Du, Yuantai Hu, and Hongping Hu. Solutions of nonlinear thickness-shear vibrations of an infinite isotropic plate with the homotopy analysis method. Numerical Algorithms, 59(2):213–226, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=2&spage=213.

Wang:2024:SEB

[WWL24] Yuxuan Wang, Tongke Wang, and Huan Lian. The series expansions and blow-up time estimation for the solutions of convolution Volterra–Hammerstein integral equations. Numerical Algorithms, 95(2):637–663, ???? 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01584-z.

Wu:2021:OPA

[WWM21] Xinyuan Wu, Bin Wang, and Lijie Mei. Oscillation-preserving algorithms for efficiently solving highly oscillatory secondorder ODEs. Numerical Algorithms, 86(2):693–727, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00908-7.

Wang:2024:CVT

[WX24] Xiaofeng Wang and Jiayi Xu. Conformable vector Traub’s method for solving nonlinear systems. Numerical Algorithms, 97(4):1563–1582, December 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01762-
7.

Wang:2020:ANA

[WXQ20] Xiaoping Wang, Huanying Xu, and Haitao Qi. Analytical and numerical analysis of time fractional dual-phase-lag heat conduction during short-pulse laser heating. Numerical Algorithms, 85(4):1385–1408, December 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00869-
6.

Wang:2022:ITI

[WXT22] W. Y. Wang, F. Q. Xia, and K. Tu. Inertial-type incremental constraint projection method for solving variational inequalities without Lipschitz continuity. Numerical Algorithms, 89 (4):1769–1798, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01172-z.

Yang:2018:CCM

[wYN18] Wei wei Yang and Qin Ni. A cubically convergent method for solving the largest eigenvalue of a nonnegative irreducible tensor. Numerical Algorithms, 77(4):1183–1197, April 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2023:CNT

[WYP23] Xiaoliang Wang, Gonglin Yuan, and Liping Pang. A class of new three-term descent conjugate gradient algorithms for large-scale unconstrained optimization and applications to image restoration problems. Numerical Algorithms, 93(3):949– 970, July 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01448-y.

Wang:2021:WGF

[WYZ21] Xiaoshen Wang, Xiu Ye, and Shangyou Zhang. Weak Galerkin finite element methods with or without stabilizers. Numerical Algorithms, 88(3):1361–1381, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01079-9.

Wang:2022:NBA

[WYZ22] Qi Wang, Zhanwen Yang, and Chengchao Zhao. Numerical blow-up analysis of the explicit L_{1}-scheme for fractional ordinary differential equations. Numerical Algorithms, 89(1): 451–463, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-021-01121-w.

Wang:2011:UKF

[WZ11] Guoqiang Wang and Detong Zhu. A unified kernel function approach to primal-dual interior-point algorithms for convex quadratic SDO. Numerical Algorithms, 57(4):537–558, August 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=4&
spage=537.

Wang:2013:PAI

[WZ13a] Shan-Shan Wang and Guo-Feng Zhang. Preconditioned AHSS iteration method for singular saddle point problems. Numerical Algorithms, 63(3):521–535, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-012-9638-y.

Wang:2013:FST

[WZ13b] Xiaofeng Wang and Tie Zhang. A family of Steffensen type methods with seventh-order convergence. Numerical Algorithms, 62(3):429–444, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9597-3/;
http://www.springerlink.com/openurl.asp?genre=article&
issn=1017-1398&volume=62&issue=3&spage=429-444.

Wang:2015:EPI

[WZ15a] Xiaofeng Wang and Tie Zhang. Efficient n-point iterative methods with memory for solving nonlinear equations. Numerical Algorithms, 70(2):357–375, October 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9951-8.

Wathen:2015:SDK

[WZ15b] Andrew J. Wathen and Shengxin Zhu. On spectral distribution of kernel matrices related to radial basis functions. Numerical Algorithms, 70(4):709–726, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9970-0.

Wang:2016:CGP

[WZ16] Jueyu Wang and Detong Zhu. Conjugate gradient path method without line search technique for derivative-free unconstrained optimization. Numerical Algorithms, 73(4):957–983, December 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0124-9.

Wang:2019:SSC

[WZ19] Shanshan Wang and Luming Zhang. Split-step cubic Bspline collocation methods for nonlinear Schrödinger equations in one, two, and three dimensions with Neumann boundary conditions. Numerical Algorithms, 81(4):1531–1546, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-019-00762-2.

Wang:2022:PPE

[WZ22a] Qiming Wang and Zhaojie Zhou. A priori and a posteriori error analysis for virtual element discretization of elliptic optimal control problem. Numerical Algorithms, 90(3):989– 1015, July 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-021-01219-1.

Wei:2022:YRB

[WZ22b] Qiyuan Wei and Liwei Zhang. Yosida-regularization based differential equation approach to generalized equations with applications to nonlinear convex programming. Numerical Algorithms, 91(2):557–581, October 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01273-

Wang:2023:MLS

[WZ23a] Dongling Wang and Jun Zou. Mittag-Leffler stability of numerical solutions to time fractional ODEs. Numerical Algorithms, 92(4):2125–2159, April 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01379-8.

Wu:2023:LMB

[WZ23b] Xiang Wu and Kanjian Zhang. A limited-memory BFGSbased differential evolution algorithm for optimal control of nonlinear systems with mixed control variables and probability constraints. Numerical Algorithms, 93(2):493–542, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01425-5.

Wang:2026:SWC

[WZ26] Liqiong Wang and Qing Zhou. Strong and weak convergence rates of Euler–Maruyama scheme for stochastic differential equations with state-dependent Markovian switching driven by α -stable process. Numerical Algorithms, 101(4):2269–2299, April 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02087-9.

Wang:2023:DCP

[WZC23] Yue Wang, Youxing Zhao, and Hu Chen. Discrete comparison principle of a finite difference method for the multi-term time fractional diffusion equation. Numerical Algorithms, 93 (4):1581–1593, August 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01480-y.

Wang:2026:MIS

[WZC26] Shenghua Wang, Yueyao Zhang, and Yeo Je Cho. Modified inertial subgradient projection and contraction method for solving nonmonotone variational inequality problem in Hilbert space. Numerical Algorithms, 101(4):2819–2848, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02115-8.

Wang:2026:NAS

[WZL26] Yifei Wang, Li Zhang, and Hu Li. A numerical approach for solving fractional order pantograph mixed Volterra–Fredholm delay-integro-differential equations. Numerical Algorithms, 101 (1):273–299, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-02005-5.

Wang:2017:PHA

[WZQ17] Zhen Wang, Li Zou, and Yupeng Qin. Piecewise homotopy analysis method and convergence analysis for formally wellposed initial value problems. Numerical Algorithms, 76(2): 393–411, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Wang:2015:SOD

[WZQT15] Xiaofeng Wang, Tie Zhang, Weiyi Qian, and Mingyan Teng. Seventh-order derivative-free iterative method for solving nonlinear systems. Numerical Algorithms, 70(3):545–558, November 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-9960-2.

Wu:2014:NAC

[WZS14] Xiang Wu, Kanjian Zhang, and Changyin Sun. Numerical algorithm for a class of constrained optimal control problems of switched systems. Numerical Algorithms, 67(4):771–792, December 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9822-8.

Wang:2022:CSP

[WZVJ22] Gang Wang, Dong Zhang, Vasily. I. Vasiliev, and Tongsong Jiang. A complex structure-preserving algorithm for the full rank decomposition of quaternion matrices and its applications. Numerical Algorithms, 91(4):1461–1481, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01310-1.

Wang:2024:IPM

[WZW^{+}24] Chengjing Wang, Xile Zhao, Qingsong Wang, Zepei Ma, and Peipei Tang. An inexact proximal majorization-minimization algorithm for remote sensing image stripe noise removal. Numerical Algorithms, 97(3):1119–1139, November 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01743-2.

Wang:2007:EIR

[WZZ07a] Xiaoping Wang, Weizhong Zhang, and Liyan Zhang. Erratum: Intersection of a ruled surface with a free-form surface. Numerical Algorithms, 46(1):101–103, September 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=1&spage=101.

See [WZZ07b].

Wang:2007:IRS

[WZZ07b] Xiaoping Wang, Weizhong Zhang, and Liyan Zhang. Intersection of a ruled surface with a free-form surface. Numerical Algorithms, 46(1):85–100, September 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=46&issue=1&

spage=85. See erratum [WZZ07a].

Wu:2015:INM

[WZZ15] Jia Wu, Liwei Zhang, and Yi Zhang. An inexact Newton method for stationary points of mathematical programs constrained by parameterized quasi-variational inequalities. Numerical Algorithms, 69(4):713–735, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9922-0.

Wang:2016:NCT

[WZZ16] Peiyuan Wang, Haiyun Zhou, and Jianjun Zhou. A new convergence theorem of a projection algorithm with variable steps for variational inequalities. Numerical Algorithms, 72(3):769– 773, July 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0066-7.

Wang:2026:IGF

[WZZ26] Xiuli Wang, Ran Zhang, and Shangyou Zhang. Interpolated Galerkin finite elements on rectangular and cuboid meshes for the Poisson equation. Numerical Algorithms, 101(2):933–949, February 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02028-6.

Wang:2024:IPD

[WZZQ24] Jing Wang, Yuting Zhang, Danchen Zhu, and Lingzhi Qian. An interior penalty discontinuous Galerkin reduced order model for the variable coefficient advection-diffusion-reaction equation. Numerical Algorithms, 97(1):243–270, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01702-x.

Xu:2025:PEB

[XC25] Qichen Xu and Tyler Chen. A posteriori error bounds for the block-Lanczos method for matrix function approximation. Numerical Algorithms, 98(2):903–927, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01819-7.

Xie:2023:SCB

[XCD23] Zhongbing Xie, Gang Cai, and Qiao-Li Dong. Strong convergence of Bregman projection method for solving variational inequality problems in reflexive Banach spaces. Numerical Algorithms, 93(1):269–294, May 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01414-
8.

Xu:2015:HBA

[XCLA15] Dali Xu, Jifeng Cui, Shijun Liao, and A. Alsaedi. A HAMbased analytic approach for physical models with an infinite number of singularities. Numerical Algorithms, 69(1):59–74, May 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9881-5.

Xu:2021:ENI

[XCY21] Chen Xu, Chuanjun Chen, and Xiaofeng Yang. Efficient, noniterative, and decoupled numerical scheme for a new modified binary phase-field surfactant system. Numerical Algorithms, 86(2):863–885, February 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00915-8.

Xu:2025:UES

[XF25] Zhuangzhi Xu and Yayun Fu. Unconditional energy stability and maximum principle preserving scheme for the Allen– Cahn equation. Numerical Algorithms, 99(1):355–376, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01880-2.

Xu:2020:TCM

[XH20] Fei Xu and Qiumei Huang. A type of cascadic multigrid method for coupled semilinear elliptic equations. Numerical Algorithms, 83(2):485–510, February 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xue:2021:SOD

[XH21] Dandan Xue and Yanren Hou. A second-order decoupled algorithm with different subdomain time steps for the nonstationary Stokes/Darcy model. Numerical Algorithms, 88 (3):1137–1182, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01070-4.

Xiong:2025:VRM

[XH25] Xiangtuan Xiong and Jingjing Han. A variational regularization method for solving the non-characteristic Cauchy problem in multiple dimensions. Numerical Algorithms, 98(2):563–586, February 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01805-z.

Xie:2007:IEP

[XHZ07] Dongxiu Xie, Ningjun Huang, and Qin Zhang. An inverse eigenvalue problem and a matrix approximation problem for symmetric skew-Hamiltonian matrices. Numerical Algorithms, 46(1):23–34, September 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=46&issue=1&spage=23.

Xu:2014:QBS

[XL14] Xiao-Ping Xu and Feng-Gong Lang. Quintic B-spline method for function reconstruction from integral values of successive subintervals. Numerical Algorithms, 66(2):223–240, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9731-x.

Xing:2024:TLI

[XL24] Xin Xing and Demin Liu. A two-level iterative method with Newton-type linearization for the stationary micropolar fluid equations. Numerical Algorithms, 97(1):475–501, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01711-w.

Xu:1993:CMA

[XLC93] Yuan Xu, W. A. Light, and E. W. Cheney. Constructive methods of approximation by ridge functions and radial functions. Numerical Algorithms, 4(3):205–223, March 1993. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xu:2022:ENM

[XLG22] Zhenhua Xu, Zhanmei Lv, and Hongrui Geng. Efficient numerical methods for Cauchy principal value integrals with highly oscillatory integrands. Numerical Algorithms, 91(3): 1287–1314, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01302-1.

Xu:2026:MNR

[XLL26] Haobo Xu, Jiehui Lu, and Fang Li. Mixed noise removal via generalized Gaussian scale mixture modeling and plugand-play deep prior. Numerical Algorithms, 101(4):2659–2686, April 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02103-y.

Xie:2020:MND

[XLW20] Fang Xie, Rong-Fei Lin, and Qing-Biao Wu. Modified Newton–DSS method for solving a class of systems of nonlinear equations with complex symmetric Jacobian matrices. Numerical Algorithms, 85(3):951–975, November 2020.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00847-y.

Xie:2016:MPD

[XM16] Ya-Jun Xie and Chang-Feng Ma. A modified positivedefinite and skew-Hermitian splitting preconditioner for generalized saddle point problems from the Navier–Stokes equation. Numerical Algorithms, 72(1):243–258, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0043-1.

Xue:2023:SSO

[XP23] Menglong Xue and Liping Pang. A strong sequential optimality condition for cardinality-constrained optimization problems. Numerical Algorithms, 92(3):1875–1904, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01371-2.

Xia:2024:TFT

[XQZ24] Sijia Xia, Duo Qiu, and Xiongjun Zhang. Tensor factorization via transformed tensor–tensor product for image alignment. Numerical Algorithms, 95(3):1251–1289, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01607-9.

Xu:2022:OCT

[XSL22] Jiali Xu, Haiyan Su, and Zhilin Li. Optimal convergence of three iterative methods based on nonconforming finite element discretization for 2D/3D MHD equations. Numerical Algorithms, 90(3):1117–1151, July 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01224-4.

Xiao:2016:HSO

[XT16] Aiguo Xiao and Xiao Tang. High strong order stochastic Runge–Kutta methods for Stratonovich stochastic differential equations with scalar noise. Numerical Algorithms, 72(2):259– 296, June 2016. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-015-0044-0.

Xu:2007:IRO

[XTH07] Yuan Xu, Oleg Tischenko, and Christoph Hoeschen. Image reconstruction by OPED algorithm with averaging. Numerical Algorithms, 45(1–4):179–193, August 2007. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-007-9089-z.

Xu:2019:SOD

[Xu19] Da Xu. Second-order difference approximations for Volterra equations with the completely monotonic kernels. Numerical Algorithms, 81(3):1003–1041, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xue:1995:OPS

[Xue95] Guo Liang Xue. On an open problem in spherical facility location. Numerical Algorithms, 9(1–2):1–12, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Xu:2017:NCT

[XW17] Chang-Lin Xu and Guo-Yin Wang. A novel cognitive transformation algorithm based on Gaussian cloud model and its application in image segmentation. Numerical Algorithms, 76(4): 1039–1070, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xiao:2018:NSS

[XW18] X. Y. Xiao and X. Wang. A new single-step iteration method for solving complex symmetric linear systems. Numerical Algorithms, 78(2):643–660, June 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xie:2026:FTT

[XWC26] Mengyan Xie, Qing-Wen Wang, and Jie Chen. Fixedtime TGNN model with the nonlinear activation function for online solution of Sylvester tensor equation. Numerical Algorithms, 101(2):1023–1047, February 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02031-x.

Xie:2025:MPM

[XWL25] Zhongbing Xie, Huanqin Wu, and Liya Liu. Modified projection method and strong convergence theorem for solving variational inequality problems with non-Lipschitz operators. Numerical Algorithms, 98(4):1755–1780, April 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01851-7.

Xu:2024:LPM

[XWX24] Fei Xu, Bingyi Wang, and Manting Xie. Local and parallel multigrid method for semilinear Neumann problem with nonlinear boundary condition. Numerical Algorithms, 96(1):185– 210, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01643-5.

Xiao:2019:PRI

[XWY19] X. Y. Xiao, X. Wang, and H. W. Yin. Preconditioned Richardson iteration for augmented linear systems. Numerical Algorithms, 82(3):843–867, November 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xu:2016:EHO

[XX16] Zhenhua Xu and Shuhuang Xiang. On the evaluation of highly oscillatory finite Hankel transform using special functions. Numerical Algorithms, 72(1):37–56, May 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0033-3.

Xie:2017:CPA

[XXW17] Pengpeng Xie, Hua Xiang, and Yimin Wei. A contribution to perturbation analysis for total least squares problems. Numerical Algorithms, 75(2):381–395, June 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xia:2025:AHI

[XY25] Yande Xia and Yun-Bo Yang. An Arrow–Hurwicz iterative method based on charge-conservation for the stationary inductionless magnetohydrodynamic system. Numerical Algorithms, 98(2):1045–1084, February 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01825-9.

Xiao-yong:2016:LAN

[XyJl16] Zhang Xiao-yong and Li Jun-lin. Laguerre approximation with negative integer and its application for the delay differential equation. Numerical Algorithms, 72(4):1065–1087, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0080-9.

Xie:2014:AAS

[XYZ14] Wei-Si Xie, Yu-Fei Yang, and Bo Zhou. An ADMM algorithm for second-order TV-based MR image reconstruction. Numerical Algorithms, 67(4):827–843, December 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9826-z.

Xu:2025:TFB

[XZ25] Xuejun Xu and Shangyou Zhang. Three families of C^{1}– P_{2m+1} Bell finite elements on triangular meshes. Numerical Algorithms, 99(2):717–734, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01894-
w.

Xia:2012:NLF

[XZL12] Peng Xia, Shugong Zhang, and Na Lei. The Neville-like form of the Fitzpatrick algorithm for rational interpolation. Numerical Algorithms, 61(1):105–120, September 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=1&spage=105.

Xu:2020:CMM

[XZP^{+}20] Weiwei Xu, Lei Zhu, Xiaofei Peng, Hao Liu, and Junfeng Yin. A class of modified modulus-based synchronous multisplitting iteration methods for linear complementarity problems. Numerical Algorithms, 85(1):1–21, September 2020.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00799-3.

Xiong:2013:SGM

[XZW13] Xiangtuan Xiong, Xiaochen Zhao, and Junxia Wang. Spectral Galerkin method and its application to a Cauchy problem of Helmholtz equation. Numerical Algorithms, 63(4):691–711, August 2013. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-012-9648-9.

Xu:2019:SCI

[XZZ19a] Yangyang Xu, Ruijuan Zhao, and Bing Zheng. Some criteria for identifying strong H-tensors. Numerical Algorithms, 80(4): 1121–1141, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xu:2019:TNP

[XZZ19b] Yangyang Xu, Ruijuan Zhao, and Bing Zheng. Two nonparameter iterative algorithms for identifying strong Htensors. Numerical Algorithms, 81(3):1113–1128, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Xie:2022:CAM

[XZZ22] Yingying Xie, Liuqiang Zhong, and Yuping Zeng. Convergence of an adaptive modified WG method for second-order elliptic problem. Numerical Algorithms, 90(2):789–808, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01209-3.

Yoshida:2024:CIB

[YA24] Toshio Yoshida and Yoshinori Adachi. Computation of incomplete beta function ratios I_{x}(a, b) with Deuflhard’s algorithm. Numerical Algorithms, 97(1):373–390, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01707-6.

Yakoubsohn:1994:AZA

[Yak94] Jean-Claude Yakoubsohn. Approximating the zeros of analytic functions by the exclusion algorithm. Numerical Algorithms, 6(1–2):63–88, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yakoubsohn:1995:UCC

[Yak95] Jean-Claude Yakoubsohn. A universal constant for the convergence of Newton’s method and an application to the classical homotopy method. Numerical Algorithms, 9(3–4):223–244, ???? 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yalamov:2001:ISB

[Yal01] Plamen Y. Yalamov. Improvement of some bounds on the stability of fast Helmholtz solvers. Numerical Algorithms, 26(1):11–20, January 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
4/2/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/4/2/fulltext.pdf.

Yan:1995:MMD

[Yan95] Xiao Pu Yan. Multistep methods for differential algebraic equations. Numerical Algorithms, 10(3–4):245–260, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2017:CMI

[Yan17] Yaguang Yang. CurveLP — a MATLAB implementation of an infeasible interior-point algorithm for linear programming. Numerical Algorithms, 74(4):967–996, April 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0180-1.

Yang:2018:TCE

[Yan18] Y. Yang. Two computationally efficient polynomial-iteration infeasible interior-point algorithms for linear programming. Numerical Algorithms, 79(3):957–992, November 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2022:EAS

[Yan22] Yaguang Yang. An efficient arc-search interior-point algorithm for convex quadratic programming with box constraints.

Numerical Algorithms, 91(2):711–748, October 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01279-x.

Yambangwai:2020:NMT

[YAT20] Damrongsak Yambangwai, Sukanya Aunruean, and Tanakit Thianwan. A new modified three-step iteration method for G-nonexpansive mappings in Banach spaces with a graph. Numerical Algorithms, 84(2):537–565, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00768-w.

Yao:2026:FRB

[YAY26] Yonghong Yao, Abubakar Adamu, and Jen-Chih Yao. Forward-reflected-backward algorithms with double inertial extrapolations for variational inequalities with applications to image processing. Numerical Algorithms, 101(4):2449–2484, April 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02095-9.

Yang:2021:CNM

[YBK^{+}21] Guoguo Yang, Kevin Burrage, Yoshio Komori, Pamela Burrage, and Xiaohua Ding. A class of new Magnus-type methods for semi-linear non-commutative Itô stochastic differential equations. Numerical Algorithms, 88(4):1641–1665, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01089-7.

Yang:2022:ESA

[YC22] Qingmin Yang and Xiao Shan Chen. On estimating the separation associated with the periodic continuous Sylvester equation. Numerical Algorithms, 90(4):1419–1435, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01235-1.

Yang:2017:SCG

[YCL17] Yueting Yang, Yuting Chen, and Yunlong Lu. A subspace conjugate gradient algorithm for large-scale unconstrained optimization. Numerical Algorithms, 76(3):813–828, November 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yin:2019:BCF

[YCW^{+}19] Weidi Yin, Songsong Cheng, Yiheng Wei, Jianmei Shuai, and Yong Wang. A bias-compensated fractional order normalized least mean square algorithm with noisy inputs. Numerical Algorithms, 82(1):201–222, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2026:NFO

[YCZ26] Xu Yang, Qing Chen, and Weidong Zhao. A new first order numerical scheme for nonlinear jump-diffusion problems and its strong convergence analysis. Numerical Algorithms, 102(2):847–874, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02156-z.

Yaman:2009:NRA

[YD09] Irem Yaman and Metin Demiralp. A new rational approximation technique based on transformational high dimensional model representation. Numerical Algorithms, 52(3):385–407, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
52&issue=3&spage=385.

Yuan:2026:SMI

[YD26] Yuqian Yuan and Fangfang Dou. Statistical methods for the inversion of the source term and initial values of the diffusion equation. Numerical Algorithms, 102(2):1355–1376, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02179-6.

Yang:2026:NAS

[YDW26] Naya Yang, Xuefeng Duan, and Qingwen Wang. A new algorithm for solving a class of matrix optimization problem arising in unsupervised feature selection. Numerical Algorithms, 101 (1):97–118, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01997-4.

Yang:2015:GPI

[YDWL15] Ai-Li Yang, Yan Dou, Yu-Jiang Wu, and Xu Li. On generalized parameterized inexact Uzawa methods for singular saddlepoint problems. Numerical Algorithms, 69(3):579–593, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9914-0.

Ye:1996:ABL

[Ye96] Qiang Ye. An adaptive block Lanczos algorithm. Numerical Algorithms, 12(1–2):97–110, April 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ye:2022:IPT

[Ye22] Minglu Ye. An infeasible projection type algorithm for nonmonotone variational inequalities. Numerical Algorithms, 89 (4):1723–1742, April 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-021-01170-1.

Yang:2022:AMR

[YF22] Chaoyi Yang and Brian C. Fabien. An adaptive mesh refinement method for indirectly solving optimal control problems. Numerical Algorithms, 91(1):193–225, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01259-1.

Ye:1997:DDF

[YH97] Xiu Ye and Charles A. Hall. A discrete divergence-free basis for finite element methods. Numerical Algorithms, 16 (3–4):365–380, December 1997. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.com/content/getfile/5058/
11/7/abstract.htm;   http://ipsapp007.kluweronline.
com/content/getfile/5058/11/7/fulltext.pdf.

Yang:2021:JSA

[YH21] Changqing Yang and Jianhua Hou. Jacobi spectral approximation for boundary value problems of nonlinear fractional pantograph differential equations. Numerical Algorithms, 86(3): 1089–1108, March 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00924-7.

Yang:2024:SCE

[YH24] Hongfu Yang and Jianhua Huang. Strong convergence and extinction of positivity preserving explicit scheme for the stochastic SIS epidemic model. Numerical Algorithms, 95(4):1475– 1502, April 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01617-7.

Yang:2025:RAF

[YHL25] Yi Yang, Jin Huang, and Hu Li. An α-robust analysis of finite element method for space–time fractional diffusion equation. Numerical Algorithms, 98(1):165–190, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01789-w.

Yang:2019:MMF

[YHQ19] Xiaobo Yang, Weizhang Huang, and Jianxian Qiu. Moving mesh finite difference solution of non-equilibrium radiation diffusion equations. Numerical Algorithms, 82(4):1409–1440, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yazdani:2026:OSD

[YHR26] Alireza Yazdani, Ronald D. Haynes, and Steven J. Ruuth. Optimized Schwarz domain decomposition algorithms for the closest point method on closed manifolds. Numerical Algorithms, 101(1):33–64, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01995-6.

Yao:2018:IPC

[YHS18] Shengwei Yao, Donglei He, and Lihua Shi. An improved Perry conjugate gradient method with adaptive parameter choice. Numerical Algorithms, 78(4):1255–1269, August 2018. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2020:ACE

[YHZ20] Zhao Yang, Rong Huang, and Wei Zhu. Accurate computations for eigenvalues of products of Cauchy-polynomial-Vandermonde matrices. Numerical Algorithms, 85(1):329–351, September 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00816-5.

Yang:2021:ASS

[YHZL21] Zhao Yang, Rong Huang, Wei Zhu, and Jianzhou Liu. Accurate solutions of structured generalized Kronecker product linear systems. Numerical Algorithms, 87(2):797–818, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00988-5.

Yamashita:2022:IIP

[YIY22] Makoto Yamashita, Einosuke Iida, and Yaguang Yang. An infeasible interior-point arc-search algorithm for nonlinear constrained optimization. Numerical Algorithms, 89(1):249–275, January 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01113-w.

Yang:2018:EUV

[YJ18] Yun-Bo Yang and Yao-Lin Jiang. An explicitly uncoupled VMS stabilization finite element method for the time-dependent Darcy–Brinkman equations in double-diffusive convection. Numerical Algorithms, 78(2):569–597, June 2018. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2021:UOE

[YJ21] Yun-Bo Yang and Yao-Lin Jiang. Unconditional optimal error estimates of linearized backward Euler Galerkin FEMs for nonlinear Schrödinger–Helmholtz equations. Numerical Algorithms, 86(4):1495–1522, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-00942-5.

Yin:2021:HTT

[YJJ^{+}21] Jianghua Yin, Jinbao Jian, Xianzhen Jiang, Meixing Liu, and Lingzhi Wang. A hybrid three-term conjugate gradient projection method for constrained nonlinear monotone equations with applications. Numerical Algorithms, 88(1):389–418, September 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-01043-z.

Yao:2025:CFT

[YJS25] Yonghong Yao, Lateef O. Jolaoso, and Yekini Shehu. C-FISTA type projection algorithm for quasi-variational inequalities. Numerical Algorithms, 98(4):1781–1798, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01852-6.

Yu:2015:NCN

[YJX15] Bo Yu, Xiaoyun Jiang, and Huanying Xu. A novel compact numerical method for solving the two-dimensional nonlinear fractional reaction-subdiffusion equation. Numerical Algorithms, 68(4):923–950, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9877-1.

Yang:2022:EDP

[YK22] Junxiang Yang and Junseok Kim. Energy dissipationpreserving time-dependent auxiliary variable method for the phase-field crystal and the Swift–Hohenberg models. Numerical Algorithms, 89(4):1865–1894, April 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01176-9.

Yamashita:2015:NSF

[YKY15] Takumi Yamashita, Kinji Kimura, and Yusaku Yamamoto. A new subtraction-free formula for lower bounds of the minimal singular value of an upper bidiagonal matrix. Numerical Algorithms, 69(4):893–912, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-014-9931-z.

Yin:2016:SFD

[YL16] Ping Yin and Jacques Liandrat. A smooth fictitious domain/multiresolution method for elliptic equations on general domains. Numerical Algorithms, 72(3):705–720, July 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0063-x.

Yang:2019:SCR

[YL19] Jun Yang and Hongwei Liu. Strong convergence result for solving monotone variational inequalities in Hilbert space. Numerical Algorithms, 80(3):741–752, March 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yi:2022:MEC

[YL22] Nianyu Yi and Hailiang Liu. A mass- and energy-conserved DG method for the Schrödinger–Poisson equation. Numerical Algorithms, 89(2):905–930, February 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01139-
0.

Yao:2011:GYA

[YLD11] Guozhu Yao, Anping Liao, and Xuefeng Duan. Positive definite solution of the matrix equation X = Q+A^{H}(I⊗X−C)^{δ}A. Numerical Algorithms, 56(3):349–361, March 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=56&issue=3&spage=349.

Yang:2020:CSE

[YLL20] Jun Yang, Hongwei Liu, and Guaiwei Li. Convergence of a subgradient extragradient algorithm for solving monotone variational inequalities. Numerical Algorithms, 84(1):389–405, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yuan:2022:GCB

[YLL22] Gonglin Yuan, Pengyuan Li, and Junyu Lu. The global convergence of the BFGS method with a modified WWP line search for nonconvex functions. Numerical Algorithms, 91(1):353–365, September 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01265-3.

Yuan:2012:PIS

[YLY12] Shifang Yuan, Anping Liao, and Guozhu Yao. Parameterized inverse singular value problem for anti-bisymmetric matrices. Numerical Algorithms, 60(3):501–522, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=60&issue=3&spage=501.

Yang:2023:NAL

[YLYZ23] Xiaochen Yang, Mengna Li, Zhanwen Yang, and Chiping Zhang. Numerical analysis of a linearly backward Euler method with truncated Wiener process for a stochastic SIS model. Numerical Algorithms, 93(2):563–579, June 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01427-3.

Yadav:2024:EPR

[YM24] Narendra Singh Yadav and Kaushik Mukherjee. Efficient parameter-robust numerical methods for singularly perturbed semilinear parabolic PDEs of convection-diffusion type. Numerical Algorithms, 96(2):925–973, June 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01670-2.

Yousefpour:2016:CSD

[You16] Rohollah Yousefpour. Combination of steepest descent and BFGS methods for nonconvex nonsmooth optimization. Numerical Algorithms, 72(1):57–90, May 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-015-0034-2.

Yun:2009:IMB

[YP09] Beong In Yun and Miodrag S. Petković. Iterative methods based on the signum function approach for solving nonlinear equations. Numerical Algorithms, 52(4):649–662, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=52&issue=4&
spage=649.

Yu:2023:DAD

[YP23] Yongchao Yu and Jigen Peng. DCACO: an algorithm for designing incoherent redundant matrices. Numerical Algorithms, 93(2):785–813, June 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01441-5.

Yang:2021:FLM

[YPL21] Fan Yang, Qu Pu, and Xiao-Xiao Li. The fractional Landweber method for identifying the space source term problem for timespace fractional diffusion equation. Numerical Algorithms, 87 (3):1229–1255, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01006-4.

Yuste:2016:FAR

[YQM16] Santos B. Yuste and J. Quintana-Murillo. Fast, accurate and robust adaptive finite difference methods for fractional diffusion equations. Numerical Algorithms, 71(1):207–228, January 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-9998-1.

Yserentant:1999:CGS

[Yse99] Harry Yserentant. Coarse grid spaces for domains with a complicated boundary. Numerical Algorithms, 21(1– 4):387–392, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/20/
13/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/20/13/fulltext.pdf.

Yang:2019:QBV

[YSLH19] Fan Yang, Ya-Ru Sun, Xiao-Xiao Li, and Can-Yun Huang. The quasi-boundary value method for identifying the initial value of heat equation on a columnar symmetric domain. Numerical Algorithms, 82(2):623–639, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2023:SPB

[YSLL23] Yining Yang, Ziyu Sun, Yang Liu, and Hong Li. Structurepreserving BDF2 FE method for the coupled Schrödinger– Boussinesq equations. Numerical Algorithms, 93(3):1243–1267, July 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01466-w.

Yu:2019:SPG

[YSXY19] Gaohang Yu, Yisheng Song, Yi Xu, and Zefeng Yu. Spectral projected gradient methods for generalized tensor eigenvalue complementarity problems. Numerical Algorithms, 80(4): 1181–1201, April 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yuan:2021:PWL

[Yua21] Long Yuan. A plane wave least squares method for the Maxwell equations in anisotropic media. Numerical Algorithms, 87(2): 873–894, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00991-w.

Yue:2025:TNN

[Yue25] Yukun Yue. Two novel numerical methods for gradient flows: generalizations of the Invariant Energy Quadratization method. Numerical Algorithms, 98(4):1667–1701, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01847-3.

Yang:2017:TJA

[YW17] Xu Yang and Xiaojie Wang. A transformed jump-adapted backward Euler method for jump-extended CIR and CEV models. Numerical Algorithms, 74(1):39–57, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-016-0137-4.

Yuan:2020:FWC

[YWS20] Gonglin Yuan, Xiaoliang Wang, and Zhou Sheng. Family weak conjugate gradient algorithms and their convergence analysis for nonconvex functions. Numerical Algorithms, 84(3):935– 956, July 2020. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-019-00787-7.

Yang:2012:BPI

[YWWR12] Aili Yang, Yujiang Wu, Yongqing Wu, and Dawei Ren. Biparameter incremental unknowns ADI iterative methods for elliptic problems. Numerical Algorithms, 60(3):483–499, July 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=60&issue=3&
spage=483.

Yang:2014:SCP

[YWX14] Ai-Li Yang, Yu-Jiang Wu, and Zhen-Jian Xu. The semiconvergence properties of MHSS method for a class of complex nonsymmetric singular linear systems. Numerical Algorithms, 66(4):705–719, August 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-013-9755-2.

Yang:2022:UGJ

[YWYN22] Zongze Yang, Jungang Wang, Zhanbin Yuan, and Yufeng Nie. Using Gauss–Jacobi quadrature rule to improve the accuracy of FEM for spatial fractional problems. Numerical Algorithms, 89(3):1389–1411, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01158-x.

Yu:2019:LHS

[YWZ19] Hao Yu, Boying Wu, and Dazhi Zhang. The Laguerre–Hermite spectral methods for the time-fractional sub-diffusion equations on unbounded domains. Numerical Algorithms, 82(4): 1221–1250, December 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

You:2011:AAP

[YX11] Xiangcheng You and Hang Xu. Analytical approximations for the periodic motion of the Duffing system with delayed feedback. Numerical Algorithms, 56(4):561–576, April 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=56&issue=4&
spage=561.

Yu:2018:CSF

[YXL18] Qiang Yu, Hang Xu, and Shijun Liao. Coiflets solutions for Föppl–von Kármán equations governing large deflection of a thin flat plate by a novel wavelet-homotopy approach. Numerical Algorithms, 79(4):993–1020, December 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2022:SIW

[YXS22] Shuping Yang, Xiangtuan Xiong, and Yue Sun. Stationary iterated weighted Tikhonov regularization method for identifying an unknown source term of time-fractional radial heat equation. Numerical Algorithms, 90(2):881–903, June 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01213-7.

Yue:2013:NIF

[YY13] Lina Yue and Yongjian Yang. A new integral filter algorithm for unconstrained global optimization. Numerical Algorithms, 63(3):419–430, July 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9630-6.

Yao:2024:NSA

[YY24] Zichen Yao and Zhanwen Yang. Numerical stability analysis of spatial-temporal fully discrete scheme for time-fractional delay Schrödinger equations. Numerical Algorithms, 97(3): 1237–1265, November 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01747-y.

Yu:2014:RCM

[YYD14] Yan Yu, Bo Yu, and Bo Dong. Robust continuation methods for tracing solution curves of parameterized systems. Numerical Algorithms, 65(4):825–841, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9716-9.

Yadav:2024:RFB

[YYK24] Rajesh Yadav, Deepak Kumar Yadav, and Alpesh Kumar. RBF-FD based some implicit-explicit methods for pricing option under regime-switching jump-diffusion model with variable coefficients. Numerical Algorithms, 97(2):645–685, October 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01719-2.

Yang:2015:HMN

[YYL15] Li Yang, Bo Yu, and YanXi Li. A homotopy method for nonlinear second-order cone programming. Numerical Algorithms, 68(2):355–365, February 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-014-9848-6.

Yin:2026:HAD

[YYL26] Jianghua Yin, Wen Ye, and Jun Li. A hybrid accelerated derivative-free projection method for solving nonlinear equations. Numerical Algorithms, 101(2):1097–1129, February 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-025-02034-8.

Yao:2023:ESN

[YYLX23] Tingfu Yao, Changlun Ye, Xianbing Luo, and Shuwen Xiang. An ensemble scheme for the numerical solution of a random transient heat equation with uncertain inputs. Numerical Algorithms, 94(2):643–668, October 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01514-z.

Yu:2021:FSL

[YYW21] Xuhong Yu, Xueqin Ye, and Zhongqing Wang. A fast solver of Legendre–Laguerre spectral element method for the Camassa– Holm equation. Numerical Algorithms, 88(1):1–23, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01028-y.

Yuan:2022:ATT

[YYZ22] Gonglin Yuan, Heshu Yang, and Mengxiang Zhang. Adaptive three-term PRP algorithms without gradient Lipschitz continuity condition for nonconvex functions. Numerical Algorithms, 91(1):145–160, September 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-022-01257-
3.

Yu:2011:CMM

[YZ11] Haixiong Yu and Jinping Zeng. A cascadic multigrid method for a kind of semilinear elliptic problem. Numerical Algorithms, 58(2):143–162, October 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=2&spage=143.

Yang:2017:FET

[YZ17] Hongli Yang and Xianyang Zeng. A feasible and effective technique in constructing ERKN methods for multi-frequency multidimensional oscillators in scientific computation. Numerical Algorithms, 76(3):761–782, November 2017. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Yang:2021:TGM

[YZ21] Jiming Yang and Jing Zhou. A two-grid method for discontinuous Galerkin approximations to nonlinear Sobolev equations. Numerical Algorithms, 86(4):1523–1541, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00943-4.

Yang:2023:TGC

[YZ23] Jiming Yang and Jing Zhou. A two-grid combined mixed finite element and discontinuous Galerkin method for a compressible miscible displacement problem. Numerical Algorithms, 94 (2):733–763, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01518-9.

Yao:2021:RDF

[YZBJ21] Teng-Teng Yao, Zhi Zhao, Zheng-Jian Bai, and Xiao-Qing Jin. A Riemannian derivative-free Polak–Ribiére–Polyak method for tangent vector field. Numerical Algorithms, 86(1):325–355, January 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00891-z.

Yang:2021:RBT

[YZH21] Min Yang, Yunong Zhang, and Haifeng Hu. Relationship between time-instant number and precision of ZeaD formulas with proofs. Numerical Algorithms, 88(2):883–902, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01061-x.

Yang:2020:LIM

[YZL20] Fan Yang, Yan Zhang, and Xiao-Xiao Li. Landweber iterative method for identifying the initial value problem of the time-space fractional diffusion-wave equation. Numerical Algorithms, 83(4):1509–1530, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Ye:2025:NTO

[YZL25] Zhijian Ye, Zhoushun Zheng, and Zhilin Li. New third-order convex splitting methods and analysis for the phase field crystal equation. Numerical Algorithms, 98(1):47–84, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01782-3.

Yuan:2024:APB

[YZLC24] Gonglin Yuan, Xiong Zhao, Kejun Liu, and Xiaoxuan Chen. An adaptive projection BFGS method for nonconvex unconstrained optimization problems. Numerical Algorithms, 95 (4):1747–1767, April 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01626-6.

Yang:2016:MTY

[YZLP16] Ximei Yang, Yinkui Zhang, Hongwei Liu, and Yonggang Pei. A Mizuno–Todd–Ye predictor–corrector infeasible-interior-point method for linear programming over symmetric cones. Numerical Algorithms, 72(4):915–936, August 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-015-0074-7.

Yan:2022:LIS

[YZLZ22] Jinliang Yan, Ling Zhu, Fuqiang Lu, and Sihui Zheng. Linearly implicit and second-order energy-preserving schemes for the modified Korteweg–de Vries equation. Numerical Algorithms, 91(4):1511–1546, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01312-z.

Yu:2025:OAS

[YZX25] Peizhao Yu, Fuheng Zhao, and Haoming Xin. Optimization algorithms for stabilization of multi-input vibration system with time delay using eigenvalues assignment technique. Numerical Algorithms, 99(2):839–871, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01899-5.

You:2014:OCR

[YZY^{+}14] Xiong You, Jinxi Zhao, Hongli Yang, Yonglei Fang, and Xinyuan Wu. Order conditions for RKN methods solving general second-order oscillatory systems. Numerical Algorithms, 66(1):147–176, May 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9728-5.

Yan:2017:NHO

[YZZL17] Jin-Liang Yan, Qian Zhang, Zhi-Yue Zhang, and Dong Liang. A new high-order energy-preserving scheme for the modified Korteweg–de Vries equation. Numerical Algorithms, 74(3): 659–674, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-016-0166-z.

Zaky:2020:PEE

[ZA20] Mahmoud A. Zaky and Ibrahem G. Ameen. A priori error estimates of a Jacobi spectral method for nonlinear systems of fractional boundary value problems and related Volterra– Fredholm integral equations with smooth solutions. Numerical Algorithms, 84(1):63–89, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zaghloul:2024:CFI

[ZA24] Mofreh R. Zaghloul and Leen Alrawas. Calculation of Fresnel integrals of real and complex arguments up to 28 significant digits. Numerical Algorithms, 96(2):489–506, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01654-2.

Zhou:2026:ECA

[ZAC26] Mi Zhou, Rizwan Anjum, and Yeol Je Cho. Equivalence and convergence analysis of fixed point iterative schemes using higher order averaged mappings. Numerical Algorithms, 101 (3):2101–2146, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02080-2.

Zafati:2022:DMM

[Zaf22] Eliass Zafati. Discussions on a macro multi-time scales coupling method: existence and uniqueness of the numerical solution and strict non-negativity of the Schur complement. Numerical Algorithms, 90(4):1389–1417, August 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01234-2.

Zaglia:1992:PRA

[Zag92] M. Redivo Zaglia. Particular rules for the Θ-algorithm. Numerical Algorithms, 3(1–4):353–369, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Zaghloul:2024:EMP

[Zag24a] Mofreh R. Zaghloul. Efficient multiple-precision computation of the scaled complementary error function and the Dawson integral. Numerical Algorithms, 95(3):1291–1308, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01608-8. See correction [Zag26].

Zaghloul:2024:ENA

[Zag24b] Mofreh R. Zaghloul. Efficient numerical algorithms for multiprecision and multi-accuracy calculation of the error functions and Dawson integral with complex arguments. Numerical Algorithms, 97(2):869–887, October 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01727-2.

Zaghloul:2026:CEM

[Zag26] Mofreh R. Zaghloul. Correction to: Efficient multiple-precision computation of the scaled complementary error function and the Dawson integral. Numerical Algorithms, 101(3):2151, March 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02057-1.

Zaky:2022:CTA

[ZAGD22] Mahmoud A. Zaky, Howayda Abo-Gabal, and Eid H. Doha. Computational and theoretical aspects of romanovski-Bessel polynomials and their applications in spectral approximations. Numerical Algorithms, 89(4):1567–1601, April 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01165-y.

Zahra:2009:ESS

[Zah09] W. K. Zahra. Exponential spline solutions for a class of two point boundary value problems over a semi-infinite range. Numerical Algorithms, 52(4):561–573, December 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=4&spage=561.

Zaabouli:2026:IGP

[ZAH26] Ziad Zaabouli, Lekbir Afraites, and Aissam Hadri. Identification of Gaussian and Poisson noise through non-convex optimization with a non-local PDE constraint of r[u]-Laplace type. Numerical Algorithms, 102(1):123–167, May 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02124-7.

Zakharova:2017:TVR

[Zak17] Anastasia Zakharova. Total variation reconstruction from quadratic measurements. Numerical Algorithms, 75(1):81–92, May 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zaslavski:2022:MCP

[Zas22] Alexander J. Zaslavski. The method of cyclic projections for closed convex sets in a Hilbert space under the presence of computational errors. Numerical Algorithms, 91(3):1427–1439, November 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01308-9.

Zayeni:2023:FRM

[ZBDK23] Hatem Zayeni, Amel Ben Abda, Franck Delvare, and Faten Khayat. Fading regularization MFS algorithm for the Cauchy problem associated with the two-dimensional Stokes equations. Numerical Algorithms, 94(3):1461–1488, November 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01543-8.

Zhang:2021:EDM

[ZBX21] Biao Zhang, Weiping Bu, and Aiguo Xiao. Efficient difference method for time-space fractional diffusion equation with Robin fractional derivative boundary condition. Numerical Algorithms, 88(4):1965–1988, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01102-
z.

Zhao:2026:CDC

[ZCC26] Junjian Zhao, Jiaxi Cai, and Yasong Chen. Comprehensive deep contrastive manifold regularized non-negative matrix factorization for multi-view clustering. Numerical Algorithms, 102(1):637–683, May 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02147-0.

Zhong:2015:PMN

[ZCG15] Hong-Xiu Zhong, Guo-Liang Chen, and Xue-Ping Guo. On preconditioned modified Newton–MHSS method for systems of nonlinear equations with complex symmetric Jacobian matrices. Numerical Algorithms, 69(3):553–567, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9912-2.

Zhang:2024:FGM

[ZCGS24] Xin Zhang, Jingya Chang, Zhili Ge, and Zhou Sheng. A family of gradient methods using Householder transformation with application to hypergraph partitioning. Numerical Algorithms, 95(2):897–927, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01593-y.

Zhou:2014:CRM

[ZCS14] Xiaojian Zhou, Xin Chen, and Yongzhong Song. On the convergence radius of the modified Newton method for multiple roots under the center-Hölder condition. Numerical Algorithms, 65(2):221–232, February 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-013-9702-2.

Zafar:2019:SAF

[ZCT19] Fiza Zafar, Alicia Cordero, and Juan R. Torregrosa. Stability analysis of a family of optimal fourth-order methods for multiple roots. Numerical Algorithms, 81(3):947–981, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhou:2024:MGV

[ZCTD24] Xiaolin Zhou, Gang Cai, Bing Tan, and Qiao-Li Dong. A modified generalized version of projected reflected gradient method in Hilbert spaces. Numerical Algorithms, 95(1):117– 147, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01566-1.

Zhang:2015:NQM

[ZD15] Jianhua Zhang and Hua Dai. A new quasi-minimal residual method based on a biconjugate A-orthonormalization procedure and coupled two-term recurrences. Numerical Algorithms, 70(4):875–896, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-015-9978-5.

Zhang:2017:NBP

[ZD17] Jian-Hua Zhang and Hua Dai. A new block preconditioner for complex symmetric indefinite linear systems. Numerical Algorithms, 74(3):889–903, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0175-y.

Zuo:2018:PTG

[ZD18] Liyun Zuo and Guangzhi Du. A parallel two-grid linearized method for the coupled Navier–Stokes–Darcy problem. Numerical Algorithms, 77(1):151–165, January 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2021:SCA

[ZD21] Tong Zhang and Mengmeng Duan. Stability and convergence analysis of stabilized finite element method for the Kelvin– Voigt viscoelastic fluid flow model. Numerical Algorithms, 87 (3):1201–1228, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01005-5.

Zheng:2020:FCA

[ZDSY20] Xiuyun Zheng, Xiaoliang Dong, Jiarong Shi, and Wei Yang. Further comment on another hybrid conjugate gradient algorithm for unconstrained optimization by Andrei. Numerical Algorithms, 84(2):603–608, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-019-00771-1.

Dang:2018:RTP

[zDYG18] Ya zheng Dang, Jian Yao, and Yan Gao. Relaxed two points projection method for solving the multiple-sets split equality problem. Numerical Algorithms, 78(1):263–275, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2026:TGD

[ZDZ26] Chongxin Zhang, Guangzhi Du, and Yuhong Zhang. Twogrid domain decomposition method for the coupled dualpermeability-Stokes system. Numerical Algorithms, 102(2): 1377–1402, June 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02180-z.

ZivariPiran:2010:EUA

[ZE10] Hossein ZivariPiran and Wayne H. Enright. An efficient unified approach for the numerical solution of delay differential equations. Numerical Algorithms, 53(2–3):397–417, March 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=53&issue=2&
spage=397.

Zahra:2012:QSS

[ZE12] Waheed K. Zahra and Samah M. Elkholy. Quadratic spline solution for boundary value problem of fractional order. Numerical Algorithms, 59(3):373–391, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=373.

Zeghdane:2025:CFA

[Zeg25a] Rebiha Zeghdane. Correction to: Fast and accurate numerical algorithm for solving stochastic Itô–Volterra integral equations. Numerical Algorithms, 99(2):837, June 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01914-9. See [Zeg25b].

Zeghdane:2025:FAN

[Zeg25b] Rebiha Zeghdane. Fast and accurate numerical algorithm for solving stochastic Itô–Volterra integral equations. Numerical Algorithms, 99(2):809–835, June 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01898-

6. See correction [Zeg25a].

Zhang:2022:MVE

[ZF22] Yadong Zhang and Minfu Feng. A mixed virtual element method for the time-fractional fourth-order subdiffusion equation. Numerical Algorithms, 90(4):1617–1637, August 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01244-0.

Zhang:2018:IST

[ZFC18] Lixin Zhang, Changjie Fang, and Shenglan Chen. An inertial subgradient-type method for solving single-valued variational inequalities and fixed point problems. Numerical Algorithms, 79(3):941–956, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2018:IBS

[ZFG18] Ju-Li Zhang, Hong-Tao Fan, and Chuan-Qing Gu. An improved block splitting preconditioner for complex symmetric indefinite linear systems. Numerical Algorithms, 77(2): 451–478, February 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-017-0323-z.

Zhang:2023:NLC

[ZFH23] Yongle Zhang, Limei Feng, and Yiran He. A new low-cost feasible projection algorithm for pseudomonotone variational inequalities. Numerical Algorithms, 94(2):1031–1054, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01622-w.

Zhu:2026:FPD

[ZFH26] Ting-Ting Zhu, Ya-Ping Fang, and Rong Hu. Fast primaldual algorithm with Tikhonov regularization for a linear equality constrained convex optimization problem. Numerical Algorithms, 101(1):393–422, January 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02010-2.

Zhao:2014:DDS

[ZFX14] Jingjun Zhao, Yan Fan, and Yang Xu. Delay-dependent stability analysis of symmetric boundary value methods for linear delay integro-differential equations. Numerical Algorithms, 65(1): 125–151, January 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-013-9698-7.

Zhao:2019:ADT

[ZFZ19] Jianxi Zhao, Qian Feng, and Lina Zhao. Alternating direction and Taylor expansion minimization algorithms for unconstrained nuclear norm optimization. Numerical Algorithms, 82 (1):371–396, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhengda:2009:LCF

[ZG09] Huang Zhengda and Ma Guochun. On the local convergence of a family of Euler–Halley type iterations with a parameter. Numerical Algorithms, 52(3):419–433, November 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=52&issue=3&spage=419.

Zheng:2012:RRS

[ZG12a] Lin Zheng and Chuanqing Gu. Recurrence relations for semilocal convergence of a fifth-order method in Banach spaces.

Numerical Algorithms, 59(4):623–638, April 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=4&spage=623.

Zheng:2012:SCS

[ZG12b] Lin Zheng and Chuanqing Gu. Semilocal convergence of a sixth-order method in Banach spaces. Numerical Algorithms, 61(3):413–427, November 2012. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=61&issue=3&spage=413.

Zhang:2025:NSM

[ZGK^{+}25] Jiansong Zhang, Liping Gao, Yuanshuo Kong, Mei Wang, and Guanqi Yang. A new splitting mixed finite element analysis of the viscoelastic wave equation. Numerical Algorithms, 99(1):229–249, May 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01876-y.

Zhang:2024:IDL

[ZGLH24] Bo Zhang, Yuelin Gao, Xia Liu, and Xiaoli Huang. Interval division and linearization algorithm for minimax linear fractional program. Numerical Algorithms, 95(2):839–858, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01591-0.

Zhao:2017:AFT

[ZH17] Longbin Zhao and Chengming Huang. An adaptive Filon-type method for oscillatory integrals without stationary points. Numerical Algorithms, 75(3):753–775, July 2017. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhao:2019:SAI

[ZH19] Jing Zhao and Dingfang Hou. A self-adaptive iterative algorithm for the split common fixed point problems. Numerical Algorithms, 82(3):1047–1063, November 2019. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zare:2022:EGN

[ZH22] Hossein Zare and Masoud Hajarian. An efficient Gauss– Newton algorithm for solving regularized total least squares problems. Numerical Algorithms, 89(3):1049–1073, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01145-2.

Zhao:2023:GQM

[ZH23] Longbin Zhao and Chengming Huang. The generalized quadrature method for a class of highly oscillatory Volterra integral equations. Numerical Algorithms, 92(3):1503–1516, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01350-7.

Zhao:2025:TPT

[ZH25] Zexiong Zhao and Chengming Huang. Two postprocessing techniques for barycentric rational collocation methods applied to weakly singular VIEs. Numerical Algorithms, 100(3): 1389–1406, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01992-9.

Zirakkar:2026:EIR

[ZH26] Behnaz Banaye Zirakkar and Alireza Hosseini. Enhancing image restoration using a discrete variational model with triangular mesh-based TV. Numerical Algorithms, 102(2):1207– 1235, June 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02175-w.

Zhao:1995:BIS

[Zha95] Kang Zhao. Best interpolation in seminorm with convex constraints. Numerical Algorithms, 9(1–2):141–156, May 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Algorithms for constrained approximation and optimization (Stowe, VT, 1993).

Zhang:2009:TMD

[Zha09] Li Zhang. Two modified Dai–Yuan nonlinear conjugate gradient methods. Numerical Algorithms, 50(1):1–16, January 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=50&issue=1&
spage=1.

Zhang:2011:TSM

[Zha11] Li-Li Zhang. Two-step modulus-based matrix splitting iteration method for linear complementarity problems. Numerical Algorithms, 57(1):83–99, May 2011. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=57&issue=1&spage=83.

Zhang:2015:MBA

[Zha15] Jian-Jun Zhang. MSSOR-based alternating direction method for symmetric positive-definite linear complementarity problems. Numerical Algorithms, 68(3):631–644, March 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9864-6.

Zhang:2019:NCC

[Zha19] Jiansong Zhang. A new combined characteristic mixed finite element method for compressible miscible displacement problem. Numerical Algorithms, 81(3):1157–1179, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2020:DFC

[Zha20] Li Zhang. A derivative-free conjugate residual method using secant condition for general large-scale nonlinear equations. Numerical Algorithms, 83(4):1277–1293, April 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2023:CBA

[Zha23] Xiaolong Zhang. Comparisons of best approximations with Chebyshev expansions for functions with logarithmic endpoint singularities. Numerical Algorithms, 94(3):1355–1379, November 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01538-5.

Zhang:2024:ESR

[Zha24] David Kai Zhang. An explicit 16-stage Runge–Kutta method of order 10 discovered by numerical search. Numerical Algorithms, 96(3):1243–1267, July 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01783-
2.

Zhang:2026:NBU

[Zha26] Jian-Jun Zhang. A new block upper triangular preconditioner for double saddle point systems arising from liquid crystal directors modeling. Numerical Algorithms, 101(2):803–822, February 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02023-x.

Zhang:2021:LHO

[ZHFW21] Guoyu Zhang, Chengming Huang, Mingfa Fei, and Nan Wang. A linearized high-order Galerkin finite element approach for two-dimensional nonlinear time fractional Klein– Gordon equations. Numerical Algorithms, 87(2):551–574, June 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00978-7.

Zeng:2023:RKM

[ZHSX23] Yun Zeng, Deren Han, Yansheng Su, and Jiaxin Xie. Randomized Kaczmarz method with adaptive stepsizes for inconsistent linear systems. Numerical Algorithms, 94(3):1403–1420, November 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01540-x.

Zhang:2015:IDN

[ZHT15] Y. Zhang, X. B. Hu, and H. W. Tam. Integrable discretization of nonlinear Schrödinger equation and its application with Fourier pseudo-spectral method. Numerical Algorithms, 69(4): 839–862, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9928-7.

Zhu:2015:FFM

[Zhu15] Xiaojing Zhu. A feasible filter method for the nearest lowrank correlation matrix problem. Numerical Algorithms, 69(4): 763–784, August 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9924-y.

Zhu:2021:CBF

[Zhu21] Yuanpeng Zhu. A class of blending functions with C^{∞} smoothness. Numerical Algorithms, 88(2):555–582, October 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-01049-7.

Zhang:2025:NNA

[ZHW25] Xueli Zhang, Jin Huang, and Xiaoxia Wen. A new numerical algorithm based on least squares method for solving stochastic Itô–Volterra integral equations. Numerical Algorithms, 98 (1):117–132, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01787-y.

Zhang:2020:NZN

[ZHY^{+}20] Yunong Zhang, Huanchang Huang, Min Yang, Yihong Ling, Jian Li, and Binbin Qiu. New zeroing neural dynamics models for diagonalization of symmetric matrix stream. Numerical Algorithms, 85(3):849–866, November 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00840-
5.

Zhou:2025:EIM

[ZHZ25] Hao Zhou, Yaozhong Hu, and Jingjun Zhao. Explicit– implicit methods for stochastic susceptible-infected-recovered model. Numerical Algorithms, 100(3):1199–1238, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01986-7.

Zillober:2001:GCN

[Zil01] Christian Zillober. Global convergence of a nonlinear programming method using convex approximations. Numerical Algorithms, 27(3):265–289, July 2001. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.lwwonline.com/content/getfile/5058/34/2/
abstract.htm;   http://ipsapp007.lwwonline.com/content/
getfile/5058/34/2/fulltext.pdf.

Zhang:2008:NAB

[ZJ08] Xiangchao Zhang and Xiangqian Jiang. Numerical analyses of the boundary effect of radial basis functions in 3D surface reconstruction. Numerical Algorithms, 47(4):327–339, April 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=47&issue=4&
spage=327.

Zhang:2014:NCO

[ZJ14] Hui Zhang and Yao-Lin Jiang. A note on the H^{1}-convergence of the overlapping Schwarz waveform relaxation method for the heat equation. Numerical Algorithms, 66(2):299–307, June 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9734-7.

Zhou:2023:TEM

[ZJ23] Shaobo Zhou and Hai Jin. The truncated Euler–Maruyama method for highly nonlinear stochastic differential equations with multiple time delays. Numerical Algorithms, 94(2): 581–617, October 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01512-1.

Zhang:2024:CSP

[ZJJW24] Dong Zhang, Tongsong Jiang, Chuan Jiang, and Gang Wang. A complex structure-preserving algorithm for computing the singular value decomposition of a quaternion matrix and its applications. Numerical Algorithms, 95(1):267–283, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01571-4.

Zahir:2025:QTL

[ZJR25] Alaeddine Zahir, Khalide Jbilou, and Ahmed Ratnani. Quaternion tensor low-rank approximation using a family of nonconvex norms. Numerical Algorithms, 100(4):2135–2161, December 2025. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-025-02226-2.

Zhou:2025:VBS

[ZJSS25] Bin Zhou, Jie Jiang, Yongzhong Song, and Hailin Sun. Variance-based stochastic projection gradient method for twostage co-coercive stochastic variational inequalities. Numerical Algorithms, 98(1):1–33, January 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01779-
y.

Zhang:2018:GLS

[ZJWF18] Hui Zhang, Xiaoyun Jiang, Chu Wang, and Wenping Fan. Galerkin–Legendre spectral schemes for nonlinear space fractional Schrödinger equation. Numerical Algorithms, 79(1):337– 356, September 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2020:CLS

[ZJZ20] Hui Zhang, Xiaoyun Jiang, and Rumeng Zheng. Chebyshev– Legendre spectral method and inverse problem analysis for the space fractional Benjamin–Bona–Mahony equation. Numerical Algorithms, 84(2):513–536, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00767-
x.

Zubik-Kowal:2002:NAT

[ZKD02] Barbara Zubik-Kowal and Penny J. Davies. Numerical approximation of time domain electromagnetic scattering from a thin wire. Numerical Algorithms, 30(1):25–36, May 2002. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/40/3/abstract.htm;   http:
//ipsapp007.kluweronline.com/content/getfile/5058/
40/3/fulltext.pdf.

Zubik-Kowal:2004:FSA

[ZKD04] Barbara Zubik-Kowal and Penny J. Davies. Fourier stability analysis of a numerical method for time domain electromagnetic scattering from a thin wire. Numerical Algorithms, 35 (1):121–130, January 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/49/A/8/
abstract.htm.

Zheng:2007:PSK

[ZL07] Hong Zheng and Jianlin Li. A practical solution for KKT systems. Numerical Algorithms, 46(2):105–119, October 2007. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=46&issue=2&
spage=105.

Zhong:2017:HAM

[ZL17] Xiaoxu Zhong and Shijun Liao. On the homotopy analysis method for backward/forward–backward stochastic differential equations. Numerical Algorithms, 76(2):487–519, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2022:IRM

[ZL22a] Hongwu Zhang and Yong Lv. Iteration regularization method for a sideways problem of time-fractional diffusion equation. Numerical Algorithms, 91(3):1145–1163, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01295-x.

Zhang:2022:FEM

[ZL22b] Jin Zhang and Yanhui Lv. Finite element method for singularly perturbed problems with two parameters on a Bakhvalovtype mesh in 2D. Numerical Algorithms, 90(1):447–475, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01194-7.

Zhang:2023:SPL

[ZL23] Jin Zhang and Xiaowei Liu. Supercloseness and postprocessing for linear finite element method on Bakhvalov-type meshes. Numerical Algorithms, 92(3):1553–1570, March 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01353-4.

Zietlow:2024:ATI

[ZL24] Christian Zietlow and Jörg K. N. Lindner. ADMM-TGV image restoration for scientific applications with unbiased parameter choice. Numerical Algorithms, 97(4):1481–1512, December 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-01759-2.

Zhang:2025:MIM

[ZL25] Hongwu Zhang and Yanhui Li. A modified iteration method for an inverse problem of diffusion equation with Laplace and Riesz–Feller space fractional operators. Numerical Algorithms, 100(1):1–34, September 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01951-4.

Zheng:2026:UCA

[ZL26a] Ning Zheng and Tianwen Li. A unified convergence analysis of random sketch methods for rank deficient linear systems. Numerical Algorithms, 101(1):703–739, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02297-1.

Zietlow:2026:CAT

[ZL26b] Christian Zietlow and Jörg K. N. Lindner. Correction to: ADMM-TGV image restoration for scientific applications with unbiased parameter choice. Numerical Algorithms, 101(4): 2939–2940, April 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02105-w.

Zhou:2023:FEA

[ZLCW23] Zhaojie Zhou, Jie Liu, Yanping Chen, and Qiming Wang. Finite element approximation of optimal control problem with integral fractional Laplacian and state constraint. Numerical Algorithms, 94(4):1983–2004, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01561-6.

Zhang:2013:DTZ

[ZLG^{+}13] Yunong Zhang, Zhen Li, Dongsheng Guo, Zhende Ke, and Pei Chen. Discrete-time ZD, GD and NI for solving nonlinear time-varying equations. Numerical Algorithms, 64(4):721– 740, December 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9690-7.

Zhang:2021:RTP

[ZLH21] Qiuyu Zhang, Jian Li, and Pengzhan Huang. Recovery type a posteriori error estimates for the conduction convection problem. Numerical Algorithms, 86(1):425–441, January 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00894-w.

Zhang:2022:SEF

[ZLH22] Li-Ping Zhang, Zi-Cai Li, and Hung-Tsai Huang. Spurious eigenvalue-free algorithms of the method of fundamental solutions for solving the Helmholtz equation in bounded multiply connected domains. Numerical Algorithms, 91(2):895–932, October 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01286-y.

Zhang:2021:CDZ

[ZLL^{+}21a] Yunong Zhang, Xiao Liu, Yihong Ling, Min Yang, and Huanchang Huang. Continuous and discrete zeroing dynamics models using JMP function array and design formula for solving time-varying Sylvester-transpose matrix inequality. Numerical Algorithms, 86(4):1591–1614, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-020-00946-1.

Zhao:2021:NSM

[ZLL21b] Ting Zhao, Hongwei Liu, and Zexian Liu. New subspace minimization conjugate gradient methods based on regularization model for unconstrained optimization. Numerical Algorithms, 87(4):1501–1534, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01017-1.

Zhu:2011:NDF

[ZLLC11] Jianguang Zhu, Hongwei Liu, Changhe Liu, and Weijie Cong. A nonmonotone derivative-free algorithm for nonlinear complementarity problems based on the new generalized penalized Fischer–Burmeister merit function. Numerical Algorithms, 58(4):573–591, December 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=58&issue=4&spage=573.

Zhang:2022:SAM

[ZLLH22] Li-Ping Zhang, Zi-Cai Li, Ming-Gong Lee, and Hung-Tsai Huang. Stability analysis of the method of fundamental solutions with smooth closed pseudo-boundaries for Laplace’s equation: better pseudo-boundaries. Numerical Algorithms, 89(3):1183–1222, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01150-5.

Zhang:2025:FLW

[ZLM25] Qifeng Zhang, Dongfang Li, and Wanying Mao. A family of linearly weighted-θ compact ADI schemes for sine-Gordon equations in high dimensions. Numerical Algorithms, 98(2): 797–838, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01816-w. See

correction [ZM26].

Zhang:2019:EPA

[ZLQT19] Ruili Zhang, Jian Liu, Hong Qin, and Yifa Tang. Energypreserving algorithm for gyrocenter dynamics of charged particles. Numerical Algorithms, 81(4):1521–1530, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-019-00739-1.

Zhou:2024:FSO

[ZLS24] Yongtao Zhou, Cui Li, and Martin Stynes. A fast second-order predictor-corrector method for a nonlinear time-fractional Benjamin–Bona–Mahony–Burgers equation. Numerical Algorithms, 95(2):693–720, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-023-01586-
x.

Zhang:2025:QDD

[ZLS25] Mingcui Zhang, Ying Li, and Jianhua Sun. QR decomposition of dual quaternion matrix and blind watermarking scheme. Numerical Algorithms, 99(4):1743–1762, August 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01930-9.

Zhang:2017:NSF

[ZLT^{+}17] H. Zhang, F. Liu, I. Turner, S. Chen, and Q. Yang. Numerical simulation of a Finite Moment Log Stable model for a European call option. Numerical Algorithms, 75(3):569–585, July 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhong:2025:TLM

[ZLT25] Liuqiang Zhong, Huilan Li, and Ming Tang. Two-level methods for solving higher order finite element discretizations of nonsymmetric and indefinite elliptic problem. Numerical Algorithms, 98(4):1897–1916, April 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://

link.springer.com/article/10.1007/s11075-024-01857-
1.

Zhuang:2016:GFE

[ZLTA16] P. Zhuang, F. Liu, I. Turner, and V. Anh. Galerkin finite element method and error analysis for the fractional cable equation. Numerical Algorithms, 72(2):447–466, June 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-015-0055-x.

Zheng:2017:RMB

[ZLV17] Hua Zheng, Wen Li, and Seakweng Vong. A relaxation modulus-based matrix splitting iteration method for solving linear complementarity problems. Numerical Algorithms, 74 (1):137–152, January 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0142-7.

Zhang:2013:AOD

[ZLW^{+}13] Xindong Zhang, Juan Liu, Juan Wen, Bo Tang, and Yinnian He. Analysis for one-dimensional time-fractional tricomitype equations by LDG methods. Numerical Algorithms, 63 (1):143–164, May 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9617-3.

Zhou:2009:PAC

[ZLWQ09] Liangmin Zhou, Lijing Lin, Yimin Wei, and Sanzheng Qiao. Perturbation analysis and condition numbers of scaled total least squares problems. Numerical Algorithms, 51(3):381–399, July 2009. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=1017-1398&volume=51&

issue=3&spage=381. Tributes to Gene H. Golub, Part III.

Zhang:2021:ABQ

[ZLWZ21] Yanzhen Zhang, Ying Li, Musheng Wei, and Hong Zhao. An algorithm based on QSVD for the quaternion equality constrained least squares problem. Numerical Algorithms, 87(4): 1563–1576, August 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01019-z.

Zhu:2022:LRB

[ZLZ22] Xiaogang Zhu, Jimeng Li, and Yaping Zhang. A local RBFsbased DQ approximation for Riesz fractional derivatives and its applications. Numerical Algorithms, 90(1):159–196, May 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-021-01183-w.

Zhang:2023:RUE

[ZLZ23] Fengxia Zhang, Ying Li, and Jianli Zhao. A real unconstrained equivalent problem of the quaternion equality constrained weighted least squares problem. Numerical Algorithms, 94 (1):73–91, September 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-022-01493-7.

Zaslavsky:1994:NSP

[ZM94] Boris G. Zaslavsky and Alexander Moskvin. Numerical solution of positive control problem via linear programming. Numerical Algorithms, 8(2–4):185–199, January 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zheng:2016:API

[ZM16] Qing-Qing Zheng and Chang-Feng Ma. Accelerated PMHSS iteration methods for complex symmetric linear systems. Numerical Algorithms, 73(2):501–516, October 2016. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-016-0105-z.

Zhang:2026:CFL

[ZM26] Qifeng Zhang and Wanying Mao. Correction to: A family of linearly weighted-θ compact ADI schemes for sine-Gordon equations in high dimensions. Numerical Algorithms, 101(3): 2147–2149, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02056-2. See

[ZLM25].

Zhou:2026:RAS

[ZMZ26] Zhihan Zhou, Zixiang Meng, and Fengxia Zhang. Real algorithms for special least squares solutions of the complex split quaternion matrix equations AX = B and AXC = B. Numerical Algorithms, 102(1):505–530, May 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02138-1.

Zhang:2017:SOC

[ZP17] Pu Zhang and Hai Pu. A second-order compact difference scheme for the fourth-order fractional sub-diffusion equation. Numerical Algorithms, 76(2):573–598, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zheng:2023:FAA

[ZP23] Zijun Zheng and Gang Pang. A fast accurate artificial boundary condition for the Euler–Bernoulli beam. Numerical Algorithms, 93(4):1685–1718, August 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01485-7.

Zheng:2025:FSO

[ZPEL25] Zijun Zheng, Gang Pang, Matthias Ehrhardt, and Baiyili Liu. A fast second-order absorbing boundary condition for the linearized Benjamin–Bona–Mahony equation. Numerical Algorithms, 98(4):2037–2080, April 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01864-2.

Zhou:2021:NMR

[ZPX21] Shenglong Zhou, Lili Pan, and Naihua Xiu. Newton method for l_{0}-regularized optimization. Numerical Algorithms, 88 (4):1541–1570, December 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-021-01085-x.

Zhang:2019:SDC

[ZQL^{+}19] Yunong Zhang, Zhiyuan Qi, Jian Li, Binbin Qiu, and Min Yang. Stepsize domain confirmation and optimum of ZeaD formula for future optimization. Numerical Algorithms, 81 (2):561–574, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2024:TOA

[ZQS24] Hong Zhang, Xu Qian, and Songhe Song. Third-order accurate, large time-stepping and maximum-principle-preserving schemes for the Allen–Cahn equation. Numerical Algorithms, 95(3):1213–1250, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01606-w.

Zhang:2022:LCS

[ZQzS22] Qifeng Zhang, Yifan Qin, and Zhi zhong Sun. Linearly compact scheme for 2D Sobolev equation with Burgers’ type nonlinearity. Numerical Algorithms, 91(3):1081–1114, November 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01293-z.

Zadvan:2017:NPS

[ZR17] Homa Zadvan and Jalil Rashidinia. Non-polynomial spline method for the solution of two-dimensional linear wave equations with a nonlinear source term. Numerical Algorithms, 74(2):289–306, February 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-016-0149-0.

Zhang:2011:HBC

[ZRZ11] Guo-Feng Zhang, Zhi-Ru Ren, and Yuan-Yuan Zhou. On HSS-based constraint preconditioners for generalized saddlepoint problems. Numerical Algorithms, 57(2):273–287, June 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=2&
spage=273.

Zakynthinaki:2003:SOA

[ZS03] M. S. Zakynthinaki and Y. G. Saridakis. Stochastic optimization for adaptive real-time wavefront correction. Numerical Algorithms, 33(1–4):509–520, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp007.kluweronline.com/content/getfile/5058/46/
42/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/42/fulltext.pdf.

Zhou:2008:BKS

[ZS08] Yunkai Zhou and Yousef Saad. Block Krylov–Schur method for large symmetric eigenvalue problems. Numerical Algorithms, 47(4):341–359, April 2008. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=4&spage=341.

Zolfaghari:2013:SPP

[ZS13] Reza Zolfaghari and Abdollah Shidfar. Solving a parabolic PDE with nonlocal boundary conditions using the Sinc method. Numerical Algorithms, 62(3):411–427, March 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9595-5/;   http://www.springerlink.
com/openurl.asp?genre=article&issn=1017-1398&volume=
62&issue=3&spage=411-427.

Zieniuk:2019:SBB

[ZS19] Eugeniusz Zieniuk and Krzysztof Szerszeń. A separation between the boundary shape and the boundary functions in the parametric integral equation system for the 3D Stokes equation. Numerical Algorithms, 80(3):753–780, March 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-018-0505-3.pdf.

Zheng:2022:TSD

[ZS22] Bo Zheng and Yueqiang Shang. A three-step defect-correction algorithm for incompressible flows with friction boundary conditions. Numerical Algorithms, 91(4):1483–1510, December 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01311-0.

Zhang:2018:PFE

[ZSF18] Qing Zhang, Haiyan Su, and Xinlong Feng. A partitioned finite element scheme based on Gauge–Uzawa method for timedependent MHD equations. Numerical Algorithms, 78(1):277– 295, May 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zarvan:2025:RNA

[ZSGJ25] Z. Zarvan, K. Sayevand, R. M. Ganji, and H. Jafari. A reliable numerical algorithm mixed with hypergeometric function for analyzing fractional variational problems. Numerical Algorithms, 98(4):2081–2112, April 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01865-1.

Zhang:2024:CSM

[ZSLZ24] Yuhang Zhang, Minghui Song, Mingzhu Liu, and Bowen Zhao. Convergence and stability of the Milstein scheme for stochastic differential equations with piecewise continuous arguments. Numerical Algorithms, 96(1):417–448, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-023-01652-4.

Zhang:2025:CAG

[ZSWW25] Kunyu Zhang, Hu Shao, Ting Wu, and Xiaoquan Wang. A class of accelerated GADMM-based method for multi-block nonconvex optimization problems. Numerical Algorithms, 98 (2):943–985, February 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-024-01821-z.

Zadeh:2019:WDG

[ZTW19] Najmeh Azizi Zadeh, Azita Tajaddini, and Gang Wu. Weighted and deflated global GMRES algorithms for solving large Sylvester matrix equations. Numerical Algorithms, 82(1):155–181, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhu:2019:SSD

[ZTZZ19] Beibei Zhu, Yifa Tang, Ruili Zhang, and Yihao Zhang. Symplectic simulation of dark solitons motion for nonlinear Schrödinger equation. Numerical Algorithms, 81(4):1485–1503, August 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zaman:2019:NEI

[ZuI19] Sakhi Zaman and Siraj ul Islam. On numerical evaluation of integrals involving oscillatory Bessel and Hankel functions. Numerical Algorithms, 82(4):1325–1343, December 2019. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zheng:2019:MMB

[ZV19] Hua Zheng and Seakweng Vong. A modified modulus-based matrix splitting iteration method for solving implicit complementarity problems. Numerical Algorithms, 82(2):573–592, October 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zheng:2021:TSM

[ZV21] Hua Zheng and Seakweng Vong. A two-step modulus-based matrix splitting iteration method for horizontal linear complementarity problems. Numerical Algorithms, 86(4):1791–1810, April 2021. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-020-00954-1.

Zhang:2012:INL

[ZW12a] Hongwu Zhang and Ting Wei. An improved non-local boundary value problem method for a Cauchy problem of the Laplace equation. Numerical Algorithms, 59(2):249–269, February 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=2&
spage=249.

Zhang:2012:NGF

[ZW12b] Yang Zhang and Kairong Wang. A new general form of conjugate gradient methods with guaranteed descent and strong global convergence properties. Numerical Algorithms, 60 (1):135–152, May 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=1017-
1398&volume=60&issue=1&spage=135.

Zhang:2014:TIM

[ZW14] H. W. Zhang and T. Wei. Two iterative methods for a Cauchy problem of the elliptic equation with variable coefficients in a strip region. Numerical Algorithms, 65(4):875–892, April 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-013-9719-6.

Zhou:2015:SRP

[ZW15] Haiyun Zhou and Peiyuan Wang. Some remarks on the paper “Strong convergence of a self-adaptive method for the split feasibility problem”. Numerical Algorithms, 70(2):333–339, October 2015. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-014-9949-2.

Zhao:2020:IPA

[ZW20] Jian-Xun Zhao and Song Wang. An interior penalty approach to a large-scale discretized obstacle problem with nonlinear constraints. Numerical Algorithms, 85(2):571–589, October 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-019-00827-2.

Zhang:2022:PAC

[ZW22] Pingping Zhang and Qun Wang. Perturbation analysis and condition numbers of mixed least squares-scaled total least squares problem. Numerical Algorithms, 89(3):1223–1246, March 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-021-01151-4.

Zhai:2019:ISH

[ZWFY19] Shuying Zhai, Zhifeng Weng, Xinlong Feng, and Jinyun Yuan. Investigations on several high-order ADI methods for timespace fractional diffusion equation. Numerical Algorithms, 82 (1):69–106, September 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhao:2018:CDG

[ZWG18] Wenling Zhao, Changyu Wang, and Yajing Gu. On the convergence of s-dependent GFR conjugate gradient method for unconstrained optimization. Numerical Algorithms, 78(3):721– 738, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zeng:2026:CDS

[ZWL26] Wenlong Zeng, Qing-Wen Wang, and Jianzhou Liu. Construction and decomposition of scaling matrices for S^{k}-SDD matrices and their application to linear complementarity problems. Numerical Algorithms, 101(3):1425–1449, March 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02047-3.

Zhang:2024:FRE

[ZWLZ24] Fengxia Zhang, Musheng Wei, Ying Li, and Jianli Zhao. The forward rounding error analysis of the partial pivoting quaternion LU decomposition. Numerical Algorithms, 96(1):267– 288, May 2024. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-023-01646-2.

Zhao:2021:BEM

[ZWW21] Yuying Zhao, Xiaojie Wang, and Mengchao Wang. On the backward Euler method for a generalized Ait–Sahalia-type rate model with Poisson jumps. Numerical Algorithms, 87(3): 1321–1341, July 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-01009-1.

Zuo:2025:RMR

[ZWW25] Qian Zuo, Nian-Ci Wu, and Yatian Wang. On the randomized multiple row-action methods for solving linear least-squares problems. Numerical Algorithms, 100(3):893–920, November 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-024-01972-z.

Zhai:2020:NAF

[ZWWW20] Shuying Zhai, Longyuan Wu, Jingying Wang, and Zhifeng Weng. Numerical approximation of the fractional Cahn– Hilliard equation by operator splitting method. Numerical Algorithms, 84(3):1155–1178, July 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-019-00795-
7.

Zhao:2019:EAH

[ZWX19] M. Zhao, H. Wu, and C. Xiong. Error analysis of HDG approximations for elliptic variational inequality: obstacle problem. Numerical Algorithms, 81(2):445–463, June 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhou:2022:HDT

[ZWX22] Bingzhen Zhou, Bo Wang, and Ziqing Xie. A hybridizable discontinuous triangular spectral element method on unstructured meshes and its hp-error estimates. Numerical Algorithms, 91(3):1231–1260, November 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01300-3.

Zhang:2024:MNP

[ZWXX24] Yuanyuan Zhang, Qingbiao Wu, Yao Xiao, and Zhewei Xie. Modified Newton–PBS method for solving a class of complex symmetric nonlinear systems. Numerical Algorithms, 96 (1):333–368, May 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01649-z.

Zhu:2022:TPB

[ZWY22] Jun-Li Zhu, Yu-Jiang Wu, and Ai-Li Yang. A two-parameter block triangular preconditioner for double saddle point problem arising from liquid crystal directors modeling. Numerical Algorithms, 89(3):987–1006, March 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-021-01142-
5.

Zhang:2025:QBR

[ZWZ25] Lu Zhang, Hongxia Wang, and Hui Zhang. Quantile-based random sparse Kaczmarz for corrupted and noisy linear systems. Numerical Algorithms, 98(3):1575–1610, March 2025. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01844-6.

Zhao:2014:CMF

[ZXF14] Jingjun Zhao, Jingyu Xiao, and Neville J. Ford. Collocation methods for fractional integro-differential equations with weakly singular kernels. Numerical Algorithms, 65(4):723–743, April 2014. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL http://link.springer.com/
article/10.1007/s11075-013-9710-2.

Zhu:2023:ENM

[ZXL23] Bi-Yun Zhu, Ai-Guo Xiao, and Xue-Yang Li. An efficient numerical method on modified space-time sparse grid for time-fractional diffusion equation with nonsmooth data. Numerical Algorithms, 94(4):1561–1596, December 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01547-4.

Zhang:2025:CRD

[ZXL25] Tao Zhang, Yong Xia, and Shiru Li. O(1/k^{2}) convergence rates of (dual–primal) balanced augmented Lagrangian methods for linearly constrained convex programming. Numerical Algorithms, 98(1):325–345, January 2025. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-024-01796-x.

Zhao:2015:DDS

[ZXLF15] Jingjun Zhao, Yang Xu, Xindi Li, and Yan Fan. Delaydependent stability of symmetric boundary value methods for second order delay differential equations with three parameters. Numerical Algorithms, 69(2):321–336, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9898-9.

Zhang:2011:CDT

[ZXRL11] Yunong Zhang, Lin Xiao, Gongqin Ruan, and Zhan Li. Continuous and discrete time Zhang dynamics for time-varying 4th root finding. Numerical Algorithms, 57(1):35–51, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=57&issue=1&
spage=35.

Zhang:2013:NMM

[ZY13a] Jun Zhang and Yu-Fei Yang. Nonlinear multigrid method for solving the anisotropic image denoising models. Numerical Algorithms, 63(2):291–315, June 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-012-9623-5.

Zheng:2013:AMB

[ZY13b] Ning Zheng and Jun-Feng Yin. Accelerated modulus-based matrix splitting iteration methods for linear complementarity problem. Numerical Algorithms, 64(2):245–262, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-012-9664-9.

Zhang:2021:CSE

[ZY21a] Chengjian Zhang and Xiaoqiang Yan. Convergence and stability of extended BBVMs for nonlinear delay-differential– algebraic equations with piecewise continuous arguments. Numerical Algorithms, 87(3):921–937, July 2021. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11075-020-00993-8.

Zhang:2021:MBI

[ZY21b] Jian-Jun Zhang and Wan-Zhou Ye. A modulus-based iterative method for sparse signal recovery. Numerical Algorithms, 88(1):165–190, September 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.

springer.com/article/10.1007/s11075-020-01035-z.

Zhang:2023:NLM

[ZY23] Biao Zhang and Yin Yang. A new linearized maximum principle preserving and energy stability scheme for the space fractional Allen–Cahn equation. Numerical Algorithms, 93(1):179– 202, May 2023. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01411-x.

Zhao:2023:RIN

[ZYBJ23] Zhi Zhao, Teng-Teng Yao, Zheng-Jian Bai, and Xiao-Qing Jin. A Riemannian inexact Newton dogleg method for constructing a symmetric nonnegative matrix with prescribed spectrum. Numerical Algorithms, 92(3):1951–1981, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01373-0.

Zhang:2017:PAC

[ZYGQ17] Jiansong Zhang, Danping Yang, Hui Guo, and Yan Qu. Parallel algorithm combined with mixed element procedure for compressible miscible displacement problem. Numerical Algorithms, 76(4):993–1019, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2022:NAI

[ZYJY22] Jiansong Zhang, Yun Yu, Bingjie Ji, and Yue Yu. Numerical analysis of incompressible wormhole propagation with mass-preserving characteristic mixed finite element procedure. Numerical Algorithms, 89(1):323–340, January 2022. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01116-7.

Zhu:2018:DQM

[ZYLN18] X. G. Zhu, Z. B. Yuan, F. Liu, and Y. F. Nie. Differential quadrature method for space-fractional diffusion equations on 2D irregular domains. Numerical Algorithms, 79(3):853–877, November 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2021:PMP

[ZYQ^{+}21] Hong Zhang, Jingye Yan, Xu Qian, Xianming Gu, and Songhe Song. On the preserving of the maximum principle and energy stability of high-order implicit–explicit Runge– Kutta schemes for the space-fractional Allen–Cahn equation. Numerical Algorithms, 88(3):1309–1336, November 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-021-01077-x.

Zeng:2017:ITC

[ZYW17] Xianyang Zeng, Hongli Yang, and Xinyuan Wu. An improved tri-coloured rooted-tree theory and order conditions for ERKN methods for general multi-frequency oscillatory systems. Numerical Algorithms, 75(4):909–935, August 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2021:MRM

[ZYW21] Wei-Hong Zhang, Ai-Li Yang, and Yu-Jiang Wu. Minimum residual modified HSS iteration method for a class of complex symmetric linear systems. Numerical Algorithms, 86(4): 1543–1559, April 2021. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-020-00944-3.

Zhou:2022:TGF

[ZYW22] Jie Zhou, Xing Yao, and Wansheng Wang. Two-grid finite element methods for nonlinear time-fractional parabolic equations. Numerical Algorithms, 90(2):709–730, June 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-021-01205-7.

Zhang:2023:FSO

[ZYW23] Chun-Hua Zhang, Jian-Wei Yu, and Xiang Wang. A fast second-order scheme for nonlinear Riesz space-fractional diffusion equations. Numerical Algorithms, 92(3):1813–1836, March 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/
10.1007/s11075-022-01367-y.

Zhang:2019:HON

[ZYX19] Haixiang Zhang, Xuehua Yang, and Da Xu. A high-order numerical method for solving the 2D fourth-order reaction– diffusion equation. Numerical Algorithms, 80(3):849–877, March 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhu:2010:RLF

[ZZ10] Peng Zhu and Shuzi Zhou. Relaxation Lax–Friedrichs sweeping scheme for static Hamilton–Jacobi equations. Numerical Algorithms, 54(3):325–342, July 2010. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=3&spage=325.

Zeng:2017:CEM

[ZZ17] Min-Li Zeng and Guo-Feng Zhang. Complex-extrapolated MHSS iteration method for singular complex symmetric linear systems. Numerical Algorithms, 76(4):1021–1037, December 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhou:2018:TSD

[ZZ18] Zhaojie Zhou and Chenyang Zhang. Time-stepping discontinuous Galerkin approximation of optimal control problem governed by time fractional diffusion equation. Numerical Algorithms, 79(2):437–455, October 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2019:GTT

[ZZ19] Hongwu Zhang and Xiaoju Zhang. Generalized Tikhonov-type regularization method for the Cauchy problem of a semi-linear elliptic equation. Numerical Algorithms, 81(3):833–851, July 2019. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhang:2022:RSP

[ZZ22a] Fengxia Zhang and Jianli Zhao. A real structure-preserving algorithm based on the quaternion QR decomposition for the quaternion equality constrained least squares problem. Numerical Algorithms, 91(4):1815–1827, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01323-w.

Zhang:2022:FSM

[ZZ22b] Min Zhang and Guo-Feng Zhang. Fast solution method and simulation for the 2D time-space fractional Black–Scholes equation governing European two-asset option pricing. Numerical Algorithms, 91(4):1559–1575, December 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01314-x.

Zhang:2023:NME

[ZZ23] Qian Zhang and Zhiyue Zhang. Nitsche’s method for elliptic Dirichlet boundary control problems on curved domains. Numerical Algorithms, 94(2):511–545, October 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01510-3.

Zeng:2024:IPI

[ZZ24] Min-Li Zeng and Zhong Zheng. An improved preconditioned inexact Uzawa method for elliptic optimal control problems. Numerical Algorithms, 97(2):503–538, October 2024. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-023-01712-9.

Zhou:2020:PQC

[ZZB20] Yongtao Zhou, Chengjian Zhang, and Luigi Brugnano. Preconditioned quasi-compact boundary value methods for spacefractional diffusion equations. Numerical Algorithms, 84(2): 633–649, June 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-019-00773-z.

Zhang:2015:DTL

[ZZH15] Tong Zhang, Xin Zhao, and Pengzhan Huang. Decoupled two level finite element methods for the steady natural convection problem. Numerical Algorithms, 68(4):837–866, April 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://link.springer.com/article/10.
1007/s11075-014-9874-4.

Zhang:2017:ARB

[ZZL17] Qi Zhang, Yangzhang Zhao, and Jeremy Levesley. Adaptive radial basis function interpolation using an error indicator. Numerical Algorithms, 76(2):441–471, October 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s11075-017-0265-5.pdf.

Zheng:2023:MBS

[ZZLV23] Hua Zheng, Yongxiong Zhang, Xiaoping Lu, and Seakweng Vong. Modulus-based synchronous multisplitting iteration methods for large sparse vertical linear complementarity problems. Numerical Algorithms, 93(2):711–729, June 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01436-2.

Zhang:2026:BRE

[ZZW26] Fengxia Zhang, Yuqing Zhang, and Musheng Wei. The backward rounding error analysis of the quaternion Givens QR decomposition. Numerical Algorithms, 101(1):199–220, January 2026. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-024-02002-8.

Zhu:2012:AAT

[ZZWK12] Shengfeng Zhu, Hancan Zhu, Qingbiao Wu, and Yasir Khan. An adaptive algorithm for the Thomas–Fermi equation. Numerical Algorithms, 59(3):359–372, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=59&issue=3&spage=359.

Zhao:2023:PAP

[ZZX^{+}23] Zhi Zhao, Qin Zeng, Yu-Nong Xu, Ya-Guan Qian, and Teng-Teng Yao. A projection algorithm for pseudomonotone vector fields with convex constraints on Hadamard manifolds. Numerical Algorithms, 93(3):1209–1223, July 2023. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-022-01464-y.

Zhang:2026:ATA

[ZZX26] Lei-Hong Zhang, Ya-Nan Zhang, and Ruwu Xiao. A AAAtype algorithm for the microwave duplexer filtering. Numerical Algorithms, 101(3):1607–1631, March 2026. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s11075-025-02058-0.

Zhang:2018:IAD

[ZZY18] Jian-Jun Zhang, Jian-Li Zhang, and Wan-Zhou Ye. An inexact alternating direction method of multipliers for the solution of linear complementarity problems arising from free boundary problems. Numerical Algorithms, 78(3):895–910, July 2018. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zhou:2024:HNE

[ZZY24] Ziyi Zhou, Haixiang Zhang, and Xuehua Yang. H^{1}-norm error analysis of a robust ADI method on graded mesh for threedimensional subdiffusion problems. Numerical Algorithms, 96 (4):1533–1551, August 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-023-01676-w.

Zhou:2026:NAD

[ZZY26] Jinhui Zhou, Yongkui Zou, and Xue Yang. Numerical approximation and dynamics of periodic solution in distribution of stochastic differential equations. Numerical Algorithms, 101 (1):343–372, January 2026. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL https://link.

springer.com/article/10.1007/s11075-025-02008-w.

Zhang:2020:UFE

[ZZZ20] Qian Zhang, Tengjin Zhao, and Zhiyue Zhang. Unfitted finite element for optimal control problem of the temperature in composite media with contact resistance. Numerical Algorithms, 84(1):165–180, May 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zheng:2022:VPI

[ZZZ22] Zhong Zheng, Min-Li Zeng, and Guo-Feng Zhang. A variant of PMHSS iteration method for a class of complex symmetric indefinite linear systems. Numerical Algorithms, 91(1):283–300, September 2022. CODEN NUALEG. ISSN 1017-1398 (print),

1572-9265 (electronic). URL https://link.springer.com/
article/10.1007/s11075-022-01262-6.